A Complete Bibliography of Statistics in Medicine (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

26 June 2021 Version 1.01

Title word cross-reference

$ [2587]. $107.50 [1166]. $110 [945, 871]. $12 [511]. $120 [2345]. $125 [2090, 872, 166]. $125.00 [146]. $125.00/£ [2462]. $13.50 [363]. $135 [2068]. $138 [827]. $144 [1443]. $150 [1970, 1935]. $150.00 [146]. $152.50 [554]. $162.50 [830]. $165.00 [1911]. $175 [1848]. $18 [2495, 527]. $18.95 [452]. $19.00 [21]. $19.95 [1936]. $20.50 [2379]. $21 [243]. $225.75 [1850]. $23.00 [1994]. $23.95 [2571]. $25 [1359, 2408]. $26 [2311]. $26.50 [2254]. $27 [1153]. $29.95 [2277]. 2 × 2 [205, 2424, 2270, 437, 1615, 2248, 499, 1879, 28, 1080, 1390, 2504, 438, 508, 1079, 2409, 768, 519, 306]. $30 [147, 1015]. $32.50 [667]. $34.50 [829]. $34.95 [526, 164, 325]. $35 [2055]. $35.00 [24]. $37.95 [2067]. $39.50 [1934]. $39.75 [1192]. $39.95 [1087, 1086, 2571]. $40 [1338]. $44.95 [2544]. $49.00 [20]. $49.95 [2254, 1861, 1912, 1811]. $50. [2296]. $50.00 [2313]. $54.00 [2379]. $54.95 [1936, 2546]. $55 [2181]. $59.50 [1888]. $59.95 [1898, 2157, 1846]. $64.95 [1960]. $65.95 [2276]. $71.50 [2167]. $72.95 [612]. $75 [787]. $76 [1825]. $79 [132]. $79.95 [2277].

$84.50 [1004]. $90 [597]. $93.50 [145]. $95 [1937]. $96 [346]. _{0} [1783, 652]. α [1148, 1645]. · [1812]. D [539]. F [2119]. K [438, 2424, 515, 1993, 1526, 1667, 1274]. P [344, 806, 516, 847, 1080, 726, 61, 1144, 1475, 2079, 2646]. Q [232, 2614]. R^{2} [2357, 2636]. S [2301, 539]. t [666, 470, 515, 2463, 1347, 1566, 1094, 1750, 2631]. U [2448].

-estimator [2301]. -interferon [1148]. -quantile [1094]. -ratio [515]. -sample [1274]. -spending [1645]. -squared [539]. -statistic [1566]. -test [2463, 1347, 2631, 2448]. -tests [1750]. -value [1475]. -values [61, 2079, 2646, 344, 806, 516, 847, 726].

//www.stams.strath.ac.uk/external/quercus [2278]. /Jena [1826].

0 [1961, 1901, 614, 2584, 434, 656, 770, 758, 1087, 827, 1812, 2253, 510, 945, 597, 1938, 1934, 2252, 2462, 553, 1798, 2397, 1826, 828, 2394, 1165, 1984, 1027, 2090, 2067, 2368, 1048, 1359, 2276, 1922, 2277, 452, 2367, 840, 2312, 598, 1983, 482, 2408, 2092, 1876, 362, 613, 2131, 2314, 871, 1887, 1014, 2039, 1026, 731, 2204, 2275, 2254, 2018, 1153, 2148, 786, 402, 1859, 1799, 2041, 892, 1786, 2070, 554, 2545, 1911, 2396, 2042, 759, 2369, 1878, 2379, 2038, 2134, 872, 2263, 1898, 2286, 1192, 2081, 2157, 1899, 2006, 2179, 1994, 1888, 1970, 933, 511, 2495]. 0 [1194, 944, 2573, 2378, 2345, 829, 1923, 1004, 2169, 788, 947, 2586, 2278, 1824, 1370, 1455, 1454, 1443, 2601, 2091, 1848, 1896, 2068, 841, 2599, 2007, 2194, 1972, 2461, 2572, 570, 2546, 2313, 1924, 612, 1951, 2040, 2417, 2418, 1847, 1933, 1037, 1278, 655, 946, 1849, 481, 2587, 2133, 1897, 2439, 1002, 2311, 668, 870, 2571, 1787, 1038, 2585, 526, 2150, 2093, 1982, 1971, 921, 1166, 1860, 1900, 2264, 1850, 1338, 1877, 2600, 2547, 1861, 2053, 2366, 528, 1108, 2383, 2170, 2181, 2168, 1981, 787, 2019, 1003, 2395, 2192, 2149, 2239, 1937, 527, 1277, 382, 1825]. 0 [2381, 2069, 1912, 830, 1025, 1846, 1788, 1315, 1314, 1960, 893, 1001, 920, 2008, 1164, 2265, 1811, 2178, 2382, 1935, 2167]. 0-07-023511-2 [1153]. 0-07-024268-2 [2545]. 0-0702-4268-2 [2264]. 0-11-691294-4 [362]. 0-12-212165-1 [481]. 0-12-262270-7 [1846]. 0-12-599450-8 [829]. 0-12-724962-1 [526]. 0-12-724963-X [526]. 0-13-202565-5 [870]. 0-19-261558-0 [434]. 0-19-261707-9 [788]. 0-19-261819-9 [510]. 0-19-261873-3 [1048]. 0-19-261880-6 [434]. 0-19-261918-7 [841]. 0-19-26199-3-4 [1370]. 0-19-262017-7 [1037]. 0-19-262125-4 [892]. 0-19-262205-6 [920]. 0-19-262280-3 [1003]. 0-19-262428-8 [1971]. 0-19-262619-1 [1897]. 0-19-262637X [2147]. 0-19-262670-1 [2147]. 0-19-505288-9 [1165]. 0-19-505473-3 [759]. 0-19-506120-9 [840]. 0-19-506766-5 [1025]. 0-19-506831-9 [1315]. 0-19-507490-4 [2254]. 0-19-852219-3 [2584]. 0-19-852221-5 [1194]. 0-19-852285-1 [1922]. 0-19-852331-9 [2093]. 0-19-852345-9 [2396]. 0-19-852359-9 [2149]. 0-19-853662-3 [787]. 0-19-853864-2 [1984]. 0-19-854850-8 [1849]. 0-1950-7207-3 [2178]. 0-1950-8607-4 [2311]. 0-262-19368-X [2275].

0-300-06299-0 [1982]. 0-306-44805-X [1888]. 0-309-04730-7 [1359]. 0-316-51854-9 [452]. 0-340-52921-0 [1164]. 0-340-52922-9 [1933]. 0-340-54937-8 [2494]. 0-340-59529-9 [1878]. 0-340-66240-9 [2573]. 0-340-67776-7 [2366]. 0-340-67785-6 [2572]. 0-340-71892-7 [2600]. 0-3406-768-6 [2278]. 0-387-94350-1 [1901]. 0-387-94483-4 [2053]. 0-387-94499-0 [1887]. 0-387-94516-4 [2395]. 0-387-94543-1 [2042]. 0-387-94610-1 [2546]. 0-387-94619-5 [2417]. 0-387-94727-2 [2314]. 0-387-94729-9 [2091]. 0-387-94803-1 [2276]. 0-387-94821-X [2418]. 0-387-94829-5 [2092]. 0-387-97872-0 [1108]. 0-3879-4661-6 [2286]. 0-38794716-7 [2170]. 0-393-04002-X [2313]. 0-412-05551-1 [2169]. 0-412-05611-9 [2382]. 0-412-06291-7 [2134]. 0-412-06341-7 [2069]. 0-412-07331-5 [2253]. 0-412-31750-8 [1038]. 0-412-34390-8 [528]. 0-412-39850-8 [668]. 0-412-44890-4 [1314]. 0-412-46530-2 [2168]. 0-412-4721-7 [2007]. 0-412-59930-9 [1860]. 0-412-59940-6 [1951]. 0-412-79540-X [2599]. 0-4127-3060-X [2381]. 0-442-00456-7 [656]. 0-444-81495-7 [1850]. 0-471-11338-7 [1983]. 0-471-12358-7 [2157]. 0-471-16018-0 [1960]. 0-471-17085-2 [2367]. 0-471-17912-4 [2462]. 0-471-19479-4 [2494]. 0-471-30316-X [2383]. 0-471-33869-9 [2585]. 0-471-52218-X [613]. 0-471-52912-5 [2167]. 0-471-54561-9 [1924]. 0-471-55761-7 [1027]. 0-471-55779-X [1972]. 0-471-57150-4 [1847]. 0-471-58465-7 [1278]. 0-471-58885-7 [1838]. 0-471-58925-X [1838]. 0-471-59375-3 [2150]. 0-471-61592-7 [1002]. 0-471-85301-1 [612]. 0-471-92391-5 [482]. 0-471-93113-6 [730]. 0-471-93460-5 [1812]. 0-471-93493-3 [946]. 0-471-93726-6 [2018]. 0-471-93934-X [1787]. 0-471-939870 [1961]. 0-471-94092-5 [1277]. 0-471-96026-8 [1824]. 0-471-96195-7 [1981]. 0-471-96201-5 [2148]. 0-471-96325-9 [2041]. 0-471-96936-2 [2586]. 0-471-97136-7 [2439]. 0-4711-6381-3 [2204]. 0-4711-6609-X [2312]. 0-4719-6020-9 [2345]. 0-4719-7488-9 [2263]. 0-4719-8089-7 [2394]. 0-521-38091-X [570]. 0-521-38385-4 [731]. 0-521-40195-X [1001]. 0-521-41041-X [770]. 0-521-41453-9 [947]. 0-521-42580-8 [947]. 0-521-45339-9 [1898]. 0-521-47382-9 [1936]. 0-521-47536-8 [1912]. 0-521-47928-2 [1936]. 0-521-58142-7 [2252]. 0-534-23328-7 [2070]. 0-534-23802-5 [1799]. 0-534-25776-3 [2040]. 0-534-97217-9 [1026]. 0-631-13149-3 [362]. 0-632-02911-0 [944]. 0-683-30074-1 [2571]. 0-684-80417-4 [1994]. 0-691-03794-9 [1934]. 0-697-14979-X [1876]. 0-697-20209-7 [1877]. 0-7083-1373-6 [2378]. 0-7123-0831-8 [1798]. 0-7123-0841-5 [2397]. 0-7216-8254-5 [598]. 0-7279-0757-3 [1014]. 0-7279-0904-5 [2039]. 0-7279-0907-X [1859]. 0-7279-0916-9 [2008]. 0-7279-1151-1 [2408]. 0-7486-0353-0 [921]. 0-7506-1345-9 [614]. 0-7619-0103-5 [2090]. 0-7619-0109-4 [1896]. 0-7619-0269-4 [2131]. 0-7619-0599-5 [2019]. 0-7619-5272-1 [2277]. 0-7619-5273-X [2277]. 0-7817-0118-X [1166]. 0-7923-1634-7 [827]. 0-7923-1874-9 [933]. 0-7923-3235-0 [1848]. 0-7923-3520-1 [1825]. 0-8014-2453-4 [362]. 0-8016-2078-3 [655]. 0-8018-5336-2 [1938].

0-8018-5759-7 [2379]. 0-8039-3676-1 [1788]. 0-8039-3896-9 [1900]. 0-8039-5707-6 [1899]. 0-8039-5710-6 [1923]. 0-8039-7039-0 [2265]. 0-8039-7374-8 [2194]. 0-8039-7409-4 [2461]. 0-8039-7918-5 [2239]. 0-8039-9055-3 [2547]. 0-8108-2355-1 [1004]. 0-8176-3632-3 [1166]. 0-8176-3809-1 [2081]. 0-8247-8268-9 [871]. 0-8247-8400-6 [597]. 0-8247-8636-X [1970]. 0-8247-8674-2 [2181]. 0-8247-8682-3 [872]. 0-8247-8798-6 [1192]. 0-8247-8892-3 [945]. 0-8247-9073-1 [1935]. 0-8247-9334-X [1911]. 0-8247-9341-2 [2068]. 0-8385-4541-4 [382]. 0-8385-6214-0 [2067]. 0-8493-4487-5 [786]. 0-8493-8242-4 [1861]. 0-8493-9461-9 [2587]. 0-85404-422-1 [2369]. 0-86542-870-0 [2368]. 0-86720-869-4 [2133]. 0-86720-898-6 [1454]. 0-869542-138-2 [893]. 0-87154-226-9 [1811]. 0-87389-367-0 [2006]. 0-87969-391-6 [1443]. 0-88167-736-1 [554]. 0-88167-767-1 [830]. 0-900511-168 [1455]. 0-90057476-3 [553]. 0-900767-68-5 [402]. 0-907-631-444 [828]. 0-910133-26-0 [1087]. 0-917227-06-9 [511]. 0-917227-07-7 [527]. 0-940600-39-0 [2495]. 0-943126-44-4 [2571]. 0-961-3921-2-6 [2192]. 0-961-39211-8 [758]. 0-9634884-0-6 [1338]. 0-9634884-1-4 [1786]. 0-9642236-6-X [2038]. 0-9646424-0-9 [1937]. 0316 [667]. 0316-76599-6 [667]. 047 [2082]. 047-194-0585 [2082]. 0471 [2043]. 0471-03900-4 [2043]. 0585 [2082]. 07-111 [1923]. 07-112 [1899]. 09 [450]. 09-00574-74-7 [450].

1 [1901, 602, 2276, 1922, 2277, 2408, 2602, 879, 554, 2042, 2369, 2081, 1085, 1004, 2169, 1154, 2546, 612, 481, 1950, 1897, 2571, 1825, 830, 1788, 663, 1935]. 1-85728-067-9 [1372]. 1-85728-068-7 [1372]. 1-873240-68-6 [1085]. 1-897761-03-1 [1950]. 10 [2644, 880]. 11 [891, 890, 2646]. 12 [2127, 1568, 2639, 1809, 2126]. 13 [2065, 1822]. 14 [1476, 1625, 1947, 2141, 1580, 1626, 1872]. 1454 [2645]. 15 [2089, 1871, 2393, 2003, 1948, 2005, 1874, 2429, 1869, 2004, 2002]. 16 [2649, 2128, 2343, 2638, 2129, 2332, 2130, 2145, 2165, 2331, 2248, 2237, 2164, 2066, 2190, 2635, 2250, 2143, 2333, 2354]. 168 [1455]. 17 [2506, 2648, 2473, 2474, 2633]. 17.15 [22]. 1704 [2639]. 1752 [2643]. 1773 [2644]. 18 [2641, 2636, 2559, 2642, 2560, 2637, 2640, 2643]. 18.99 [1985]. 184 [92]. 19.95 [365]. 19.95/$37.95 [947]. 19/20 [1443]. 1905 [2638]. 1930 [2462]. 1968 [1370]. 1968-85 [1370]. 1980s [410, 408]. 1988 [432]. 1989 [73, 153, 288]. 1990 [388, 2252]. 1990’s [125]. 1991 [308]. 1992 [435, 896, 890, 892]. 1993 [831, 1169, 1170, 1085]. 1994 [1458]. 1998 [2309].

2 [1938, 1984, 2367, 403, 1902, 2314, 364, 1153, 2545, 2586, 2068, 2132, 2264, 2181, 2168, 2054, 2180, 893, 1973]. 2-257-10310-6 [364]. 2-257-12310-7 [1902]. 20 [44]. 20/$50 [432]. 2000 [1897]. 21.50/$29.95 [598]. 21281 [1813]. 24 [451, 644]. 24-hour [621, 644]. 2401 [2649]. 25 [206, 327, 242, 207]. 25th [2408]. 262637X [2147]. 27.50 [433]. 27.70 [347]. 29.50 [23]. 2B [1933]. 2D [2009]. 2D/3D [2009]. 2DZ [402]. 2nd [1087, 2310, 2368, 1902, 1878, 511, 667, 788, 1002, 668, 870, 1630, 1935].

3 [2584, 945, 2310, 553, 1826, 1048, 2055, 403, 480, 401, 1014, 1962, 2204, 759, 872, 2179, 1000, 1086, 404, 2419, 2132, 2544, 2040, 655, 946, 2150, 2547, 2384, 787, 1003, 555, 2056, 2054, 2020, 2180, 2178]. 3-11-014936-2 [2132]. 3-437-20510-2 [2180]. 3-437-20511-0 [1826]. 3-437-20536-6 [1962]. 3-437-25100-7 [2056]. 3-471-11030-2 [2054]. 3-540-50490-7 [362]. 3-540-52090-2 [403]. 3-540-52571-8 [404]. 3-540-53022-3 [480]. 3-540-53621-3 [555]. 3-540-63351-0 [2179]. 3-540-94367-6 [2020]. 3-540-97798-8 [1086]. 3-5406-2399-X [2544]. 3-7643-3864-4 [2384]. 3-7643-5495-X [2419]. 3-8055-5367-6 [1000]. 3-8055-6276-4 [2055]. 30 [184]. 30/$48 [758]. 30/$54.50 [570]. 3003 [2642]. 31.80 [244]. 32.00 [381]. 35 [284]. 35.00 [43]. 36.00 [326]. 3D [2009]. 3QD [402]. 3rd [1026, 1687].

4 [2462, 2055, 362, 2131, 731, 2254, 892, 1786, 1994, 1862, 1370, 1896, 2461, 1847, 2380, 2133, 2311, 2043, 1861, 2053, 2384, 2395, 382, 1314, 2193, 1788]. 40.00 [348]. 414 [90]. 444 [828]. 44DM [404]. 45/$85 [770]. 47th [288]. 4th [452].

5 [1812, 2253, 2397, 2090, 598, 482, 2092, 2039, 2148, 786, 402, 1799, 1194, 2417, 1823, 870, 2019, 2239, 1277, 1025, 2167]. 5-FU [1055]. 50.00/$89.95 [731]. 55/$100 [1001]. 55/$89.95 [947].

6 [434, 597, 2009, 1962, 2018, 364, 2286, 1192, 1085, 1899, 1193, 1000, 667, 2378, 1923, 2278, 1371, 1454, 1443, 2572, 1951, 1338, 2192, 2020, 920]. 64.80/$119.25 [1000]. 656 [2647].

7 [656, 827, 2252, 2394, 1027, 1359, 1983, 1902, 2070, 450, 2379, 2134, 2157, 1870, 841, 2007, 1037, 1278, 2439, 1002, 1850, 1877, 2600, 2366, 2170, 1981, 527, 2069, 2056, 1846]. 7-8 [288]. 7.95 [144]. 70.50 [1973]. 74 [383]. 75/75 [562]. 76DM [91]. 798 [2641, 2640].

8 [758, 1798, 89, 1086, 404, 829, 947, 1824, 2194, 1849, 668, 1038, 1971, 528, 1912]. 80 [384]. 80th [2622]. 81 [1823]. 81-239-0161-5 [1823]. 85 [1370]. 872 [2648]. 879 [2646].

9 [614, 510, 1934, 2310, 1165, 452, 840, 871, 1026, 283, 2041, 2396, 1878, 2263, 1898, 933, 511, 2573, 2345, 788, 2091, 2647, 1924, 1933, 2380, 2587, 2585, 2093, 1860, 1900, 2645, 2149, 1937, 1315, 2008, 1811, 2382]. 9-8102-3066-4 [2380]. 9-8102-3111-3 [2310]. 9.95 [148, 285]. 90 [2009, 1862, 1154]. 90-5199-158-4 [1862]. 90-5199-210-6 [2009]. 90-6743-212-1 [1154]. 91 [1371]. 91-22-01587-6 [1371]. 92 [1193, 1813, 2193]. 92-832-2120-6 [1193]. 92-832-21281 [1813]. 92-832-2135-4 [2193]. 92-EN [1366]. 939870 [1961]. 981 [2602, 2601]. 981-02-1831-1 [2602]. 981-02-2454-0 [2601]. 99.00/$ A. [770]. Aage [2019]. Aalen [2137]. ABC [890, 559]. aberrations [585, 1394, 2613]. Aberystwyth [1950]. abnormalities [1380]. abortion [218]. abortions [218]. Abramson [112]. absolute [750, 311, 2325, 2449]. absorption [2490]. Abstracts [1256, 633]. Academic [827, 829, 1848, 481, 526, 1825, 1846, 1205]. Academy [22]. accelerate [2211]. accelerated [1767, 623]. acceleration [1409]. accept [849]. acceptance [1783, 652]. accessibility [1171]. accident [1438, 368, 1958]. according [1115, 1726]. account [2163]. Accounting [596]. accounts [601]. Accrual [2430]. accuracy [2588, 2642, 926, 292, 1895, 1578]. accurate [1880]. acid [1055]. Acknowledgement [186, 909]. acquired [1132]. across [1842, 2270, 141]. ACSF [1779]. act [1137]. action [985, 1156]. active [2247, 535, 905, 10]. active-factor [10]. activity [962, 776, 2582]. actual [1621]. actually [377, 935]. acuity [1210]. acute [375, 677, 2361]. Adam [148]. Adams [1372]. adaptations [764]. Adaptive [1208, 2443, 1357, 2479, 2541]. addendum [58]. additional [1450, 791]. additive [2425, 2137, 1211, 528]. Additivity [1394]. address [1793]. addresses [1232]. addressing [2226]. Adelin [597]. adenomatous [1929]. adequacy [1504, 165]. adherence [604, 779]. adjust [2514, 212]. adjusted [2458, 756, 1378, 1375, 639, 156, 2160, 1474, 2467, 687, 1680, 2003, 1552, 2459, 2002]. Adjusting [161, 198, 1843, 170, 2423, 595, 1185, 2605, 577, 1625, 2141]. Adjustment [2644, 332, 398, 2011, 654, 1475, 1868, 1447, 2570, 2557, 2617, 2606, 2339, 1244, 1996, 1018, 1330]. Adjustments [487, 659, 1906, 835]. Adjuvant [2104, 1378]. administrative [1914, 201]. admissions [2426]. adolescence [475]. adolescent [2558]. adolescents [689]. adult [836, 2337]. Advanced [827, 2132, 1933, 1788, 2116, 419, 1004]. Advances [409, 1922, 966, 2044, 1848, 2518, 2207, 1960]. advent [1588]. Adventures [2547]. Adverse [2515, 1561, 1157, 290, 907]. Affairs [828, 779]. affect [423]. affected [609]. Afifi [2381]. Africa [210]. after [1903, 1915, 157, 230, 1394, 726, 2173]. against [803, 1064, 1037, 2568, 1331]. Age [1375, 2085, 2171, 2172, 1161, 1210, 2437, 750, 2502, 1380, 198, 2232, 2359, 914, 1620, 2387, 1491, 2557, 2617, 2173, 2325, 313, 1162, 2139, 77, 2411, 1726, 2352, 549, 1244, 212, 1128, 342, 2457, 2161, 1350, 2233, 2024, 1889, 2473]. Age-adjusted [1375]. age-adjustment [2557, 1244]. Age-period-cohort [2171, 1491]. Age-related [1161, 1210, 750, 198, 2233]. age-specific [2232, 1620, 1162, 2139, 2411, 1128, 2161, 2024]. age-standardization [1380]. age-time [2352]. Agency [451, 2193]. Agents [2125, 495, 677]. ages [342]. aggregate [1657]. aggregated [2046]. Aggregation [857, 2390]. aging [22]. agreement [987, 425, 1852, 2155, 2451, 664, 1214, 885, 2118, 1425, 86, 2088, 53]. Agresti [1983, 612]. Ahlbom [511]. Ahn [1626]. aid [893]. Aided [2381]. AIDS [384, 1067, 2164, 1966, 1186, 1124, 1125, 1990, 371, 1126, 1519, 1979, 2061, 486, 929, 2371, 70, 2421, 1129, 1349, 958, 1802, 1837, 1011, 1066, 1453, 2044, 744, 1709, 1131, 2059, 574, 1139, 1711, 169, 1136, 682, 1140, 2484, 722, 1116, 1127, 1512, 1132, 1133, 1726, 1236, 1138, 1130, 330, 601, 1135, 1715, 1991, 1494, 1604, 48, 1134, 1012, 1321, 1890, 663, 1166]. AIDS-related [1890]. AIDS/HIV [574]. air [1544, 2508]. al [93]. Alain [2368, 1000]. Alan [1983, 2408, 612, 1897, 2043]. alarm [849]. Alaskan [1234]. Albany [13]. Albert [597, 383]. alcohol [1607, 1242, 921]. Aldershot [90]. Alderson [111]. Alert [907, 2224, 369]. Alessandra [2380]. Alexander [432, 526, 527, 2193]. Alfredo [1166]. Algebra [2081]. Algorithm [2157, 1554, 2607, 868, 665, 1661, 660]. Alldredge [945]. alleged [344]. allelic [2183]. allelic-loss [2183]. allergy [699]. Alligators [2038]. allocate [2245]. allocated [1489]. Allocating [1367, 1476, 1461]. allocation [2477, 1731, 538, 177, 2643, 1285, 1497, 2412, 8]. allow [1008]. Almost [1876]. Almqvist [1371]. alone [1055]. Alonzo [2642]. alpha [1056, 1057, 672]. Alternative [477, 228, 1095, 2305, 1644, 311, 1020, 1274, 1272, 2449, 2048, 139, 621, 1191, 623]. alternatives [988, 2568]. Altman [381, 21, 2039]. aluminium [1105]. Alvan [1982]. Alzheimer [394, 1334]. amalgam [1382]. ambulatory [834, 1041, 11]. America [1994, 2058, 2060]. American [1359, 2571]. Ames [164]. AMI [261]. among [278, 594, 2061, 486, 879, 1247, 2047, 1512, 1389, 1793, 129, 2088, 2060]. amplification [2145, 1884, 2190]. Amsterdam [2009, 1850, 574]. anaesthesiology [2500]. analgesia [1150]. analgesic [554]. analyse [1980, 2416, 1902, 364]. analysed [1335]. Analyses [147, 313, 2506, 706, 1046, 303, 2442, 983, 2218, 1596, 423, 733, 1792, 154, 2524, 833, 1804, 1680, 1956, 563, 2528, 2316, 701, 2527, 1406, 1285, 702]. Analysing [504, 849, 1103, 1752, 154, 2464, 1869, 424, 2117, 806, 993, 1112, 1770, 182, 1464, 1574, 964, 2358, 2447, 619, 904, 1768, 1319, 2214, 1495, 688, 84, 1478, 1917, 2065, 1961, 1085]. Analysis [1382, 1635, 1985, 675, 1812, 2512, 305, 393, 2277, 2558, 533, 609, 2092, 1640, 942, 2009, 2289, 2018, 534, 2148, 1837, 603, 1352, 1622, 575, 2042, 653, 1987, 2271, 1862, 585, 485, 2132, 524, 377, 2401, 2544, 295, 1679, 1514, 1034, 1605, 1298, 902, 2534, 642, 1849, 604, 2198, 1581, 1787, 779, 2585, 1982, 1900, 1759, 1042, 2547, 2366, 932, 2069, 1846, 1796, 334, 822, 1163, 2063, 280, 1381, 2529, 2321, 2472, 412, 827, 1696, 945, 597, 298, 1842, 2288, 1594, 1265, 1107, 1222, 1515, 2287, 530, 1437, 2343, 1208, 2270, 994, 1685, 2156, 1817, 728, 1646]. analysis [2320, 638, 1760, 14, 525, 1773, 2246, 2515, 1845, 1099, 215, 2503, 437, 55, 445, 1588, 1542, 1579, 486, 1633, 223, 537, 756, 1378, 754, 2519, 505, 1407, 1834, 1795, 180, 1417, 1943, 414, 2362, 1910, 1627, 998, 1056, 2393, 1584, 420, 1435, 952, 1249, 311, 1621, 449, 684, 915, 300, 652, 1697, 1590, 1650, 624, 206, 318, 984, 240, 2029, 2469, 661, 1057, 2282, 156, 2160, 2566, 2248, 637, 1466, 2216, 448, 20, 1091, 1172, 967, 1681, 443, 2324, 872, 1008, 502, 1446, 1556, 1623, 1524, 2121, 1879, 28, 1925, 368, 1202, 418, 1747]. analysis [110, 2387, 532, 947, 1200, 2522, 2639, 682, 143, 2530, 378, 844, 1658, 2226, 2377, 629, 2013, 1907, 1724, 1927, 586, 2003, 2363, 355, 1835, 1040, 1147, 2356, 2538, 2185, 2417, 2418, 24, 685, 77, 650, 394, 251, 742, 1473, 1881, 814, 1656, 1893, 1523, 164, 1306, 1687, 1133, 1267, 1386, 108, 2066, 2183, 2353, 868, 128, 543, 238, 662, 1232, 1586, 78, 2516, 374, 668, 870, 2581, 551, 1055, 755, 508, 207, 2386, 259, 2591, 884, 1079, 2629, 1479, 2554, 2171, 1755, 1756, 1084, 1776, 1535, 2610, 1661, 2508, 1543, 312, 1675, 1761, 861, 622, 2229, 834]. analysis [348, 943, 1424, 941, 2414, 501, 1555, 1463, 1238, 241, 2637, 794, 1511, 1790, 2500, 1114, 1952, 1954, 2017, 2563, 1344, 2033, 2436, 2513, 2116, 2108, 213, 2045, 2592, 621, 1191, 53, 1025, 898, 900, 1439, 1641, 627, 2532, 2031, 2626, 1497, 2619, 558, 623, 1092, 2002, 390, 1179, 2032, 1322, 1537, 1944, 1019, 1890, 2159, 291, 2470, 2115, 620, 1470, 503, 651, 2388, 2333, 2164, 1922, 2367, 1983, 2573, 1848, 1951, 2384, 2381, 2382, 613, 612, 1002, 2054, 1164, 2134, 2495, 1371, 2132, 1038, 44, 644]. Analytic [362, 276, 916, 1590, 972, 791, 1239]. analyzing [365]. Anastasios [89]. anatomical [454]. anatomically [1677]. Anatomy [347]. ancillary [575]. Anders [2462, 511, 2019]. Andersen [2544, 1108, 284]. Anderson [2417]. Andrew [482, 1902, 362, 364, 2600]. angina [14, 661]. Anguish [2585]. Animal [402, 1579, 983, 984]. animals [402]. Anne [2314]. Annie [2314]. Anniversary [2408]. Annotated [2296, 2571, 424, 1004]. Announcement [149, 185, 385, 219, 267, 435, 571, 645, 831, 1028, 1326, 1688, 2182, 1262, 911]. Announcements [599, 1005]. anonymous [1236]. anova [1094, 1882]. antecedent [456]. antenatal [1117]. Anthony [1085, 1370, 1933]. anthropometry [559, 890]. anti [985, 263, 2078]. anti-hypertension [2078]. anti-hypertensive [263]. anti-tumour [985]. antibiotic [913]. antibody [1267]. antidepressant [2230]. antigens [1856]. antihypertensive [644]. antiretroviral [1858]. any [2349, 2085]. apnoea [1945]. appear [1069]. appearance [1466]. Appleton [777, 382]. applicable [1926, 983]. Application [1966, 962, 102, 13, 415, 421, 36, 950, 1030, 964, 931, 489, 2607, 860, 2203, 640, 40, 821, 945, 1836, 477, 1746, 228, 371, 1979, 1748, 2270, 71, 1773, 1090, 2503, 1695, 413, 1918, 2232, 2284, 2230, 1375, 2362, 942, 462, 744, 1622, 575, 2616, 443, 2271, 574, 2425, 2323, 1620, 606, 1868, 1140, 37, 2578, 1033, 550, 2556, 1573, 1613, 2617, 2030, 927, 773, 1743, 2565, 2534, 2066, 2224, 738, 2337, 1480, 12, 547, 49, 1777, 1969, 392, 446, 665, 1494, 834, 1408, 1604, 1954, 2045, 306, 1239, 1334, 1750, 1295, 1184, 1321, 1779, 369, 558]. Applications [1681, 1899, 1923, 506, 783, 1924, 2384, 84, 1366, 607, 1046, 1977, 414, 1011, 1397, 1093, 422, 1136, 2206, 1242, 1694, 1893, 1776, 257, 866, 359, 2136, 2197, 2592, 1929, 2602, 1887, 871]. Applied [1901, 1382, 1053, 1172, 1914, 320, 985, 2047, 304, 2492, 140, 1511, 2032, 2367, 94, 326, 829]. Applying [2566, 2334, 812, 1154]. appraisal [1377, 184]. Approach [1849, 1158, 2186, 1186, 1647, 1762, 2156, 437, 533, 537, 754, 792, 1449, 1320, 2446, 1056, 2421, 580, 1852, 2490, 125, 300, 1492, 1636, 1059, 20, 443, 2324, 1524, 1781, 2211, 972, 2290, 2570, 999, 1302, 2257, 1147, 1156, 1272, 1077, 1694, 2411, 124, 2618, 591, 2509, 2103, 838, 2361, 263, 2171, 312, 774, 941, 1865, 2457, 1041, 899, 2143, 898, 1968, 1855, 390, 1058, 2375, 2376, 1057, 1812, 2396, 2546, 1860, 285, 325]. Approaches [1871, 845, 979, 1742, 1248, 1571, 2096, 362, 384, 1953, 1053, 1770, 2473, 1505, 1613, 791, 1293, 2172, 1761, 2318, 2113, 1463, 1239, 2505]. appropriate [1099, 1772]. appropriateness [1525, 2086]. Approximate [2063, 320, 2435, 2511, 1388, 2505]. Approximating [2568]. approximation [157, 726]. Approximations [852, 1094]. April [435]. arabinoside [1306]. arbitrary [311, 540]. Archie [2408]. architecture [1325]. area [1599, 559, 1387, 2294, 1658, 2606, 1659, 1541, 890, 1003, 1578, 2532]. area-based [1541]. areas [432, 1032, 1665, 2456]. Argentina [1537]. ARIMA [1958]. arm [1221, 1220]. armed [2641, 1068, 2400, 2640]. Arminger [1888]. Armitage [1960, 73, 69, 68]. Armpits [2038]. Arnold [1878, 2278, 242, 2572, 1933, 2494, 2600, 2366, 1164]. Arnold/Wiley [2494]. array [1159]. arrestees [1247]. arrhythmia [127]. Art [2275, 827]. artefacts [395]. arterial [1893]. arteriographic [2027]. artery [779]. arthritis [776, 761, 2470]. artificial [2154]. as-randomized [779]. Ash [2636]. Ashwini [2165]. Asia [1137]. aspects [1853, 2255, 619, 1063, 2076, 419, 2168]. assay [372]. Assays [1035, 1396, 2305, 551, 1796]. assembling [836]. assertion [686]. assess [2588, 1712, 2642, 2423, 2087]. Assessing [314, 2320, 1379, 2284, 531, 105, 1735, 1022, 1140, 2258, 2403, 1306, 995, 848, 1720, 1791, 855, 2410, 1717, 1806, 1821, 781, 1369, 1607, 1263, 1304, 1156, 1801, 2361, 928, 1152, 1578, 225, 1741, 676]. Assessment [1483, 2543, 1620, 1376, 2137, 1512, 1077, 1777, 411, 916, 1990, 2106, 2102, 27, 562, 2097, 1737, 2222, 1273, 2139, 1941, 2101, 2224, 2100, 2103, 751, 752, 259, 1964, 1181, 1968, 1184, 790, 558, 1015]. assessments [1572, 664, 926]. assignment [1208]. assisted [861]. associated [2102, 468, 1082, 11]. Association [363, 1359, 402, 1223, 2502, 373, 952, 1940, 1815, 1578, 2291, 1580, 1899]. Associations [1480, 1736, 443, 1306, 2337]. assuming [1990]. assumption [2597, 1369, 1801, 1331]. assumptions [1504, 1638, 1079]. asthma [2508]. asymptomatic [1802]. asymptotes [1472]. Asymptotic [2598, 1526, 2409, 1772, 1880]. ATBC [2578]. atherosclerosis [2027]. Atkinson [2643]. Atlas [362, 432, 132, 1370, 2608]. atlases [1233]. atomic [2325]. attack [1082]. attained [2437, 514]. attendances [1958]. attenuation [1046]. Attitudes [2300, 1469]. Attributable [2644, 1160, 398, 639, 654, 227, 1618, 1853, 1332]. attrition [761, 1213]. Audit [1950, 884]. augment [2438]. augmentation [1765]. Augmented [2062]. August [896]. Austin [739]. Australian [2096]. Auswahlverfahren [2180]. Author [1949, 2129, 431, 2475, 1810, 2127, 769, 2146, 2191, 2249, 34, 1873, 88, 2251, 2003, 2166, 131, 2238, 2634, 2142, 2005, 1875, 2144, 2355, 2344, 1476, 1625, 1626]. authorizations [1366]. Authors [2571, 1050]. authorship [2266]. autocorrelation [1550, 2390, 1549]. autocorrelations [925]. autologous [1408]. Autoregressive [1491, 2564, 1918, 474, 678, 241]. autosomal [54]. auxiliary [1103, 2065, 1011, 1127]. availability [1150, 1242]. Available [596, 1510]. AVAS [2161]. average [2359, 39, 747, 2374, 2508, 2615, 1181]. averaged [1734]. avoid [547, 846]. avoiding [311]. awards [1259]. awareness [1469]. AX [373]. AZT [1473].

B [1901, 770, 1812, 1359, 840, 598, 132, 1936, 554, 2369, 2248, 891, 145, 1970, 2601, 2068, 242, 2544, 2418, 2342, 1823, 668, 2366, 1947, 2239, 2178, 2382, 795, 1698, 1038]. b19 [1622]. Babbie [2547]. Back [1129, 371, 1126, 2421, 1349, 507, 1743, 1116, 1127, 2411, 1350]. back-calculation [2421, 1349, 507, 1743, 2411, 1350]. Back-projection [1129, 371, 1126, 1116, 1127]. Backcalculation [1128, 722, 330]. bacteria [1699]. bad [1198]. Bailar [1087]. Baker [1970]. Bala [1454]. balance [2189, 1100]. balanced [1208, 479, 1692, 913, 2349, 903, 8]. Baltimore [24]. Band [403]. bands [2377]. bane [553]. banging [2607]. Bangkok [1438]. Bank [1037]. Bank/WHO [1037]. Bar [402]. Barcelona [2508]. Barker [1004]. Barndorff [2168]. Barndorff-Nielsen [2168]. Barrie [1048]. Bart [2379]. Bartlett [1454, 2133]. base [38]. Based [1849, 1874, 1872, 1483, 1117, 1158, 520, 757, 1219, 979, 277, 1858, 2010, 521, 1780, 1445, 399, 1730, 2362, 2127, 2117, 1820, 1570, 727, 1059, 2616, 2468, 1513, 1078, 2330, 1770, 1413, 972, 581, 2363, 1809, 1383, 1345, 2326, 1156, 2606, 964, 659, 1694, 1293, 2295, 543, 1541, 1618, 928, 913, 263, 2629, 1356, 1616, 446, 2079, 1108, 509, 2049, 1388, 2633, 2592, 2567, 1684, 2555, 1058, 1083, 1013, 2126, 1796, 1470, 1027]. Basel [1458, 1000]. Baseline [281, 323, 2343, 1885, 499, 1008, 1410, 595, 768, 1100, 1478, 2031]. basic [1639, 667, 1629, 656, 1876, 2133, 2395, 382]. Basics [2461]. Basil [362]. Basing [111]. Basing-stoke [111]. basis [280, 1705, 1245]. batch [449, 448]. batch-to-batch [449, 448]. Bausell [1004]. Bayes [2382, 2351, 530, 792, 335, 1032, 632, 443, 733, 856, 631, 1710, 289, 2540, 1659, 1998, 935, 200, 489, 1007, 969, 1660, 2382]. Bayes/empirical [632]. Bayesian [1871, 2573, 1933, 1061, 2066, 325, 2063, 1647, 1926, 530, 1437, 1436, 994, 1685, 1773, 1178, 437, 1571, 2402, 2446, 420, 1510, 1854, 1999, 1886, 421, 1060, 566, 1053, 2248, 422, 1336, 1008, 1446, 1879, 1925, 418, 833, 1686, 1818, 2290, 2334, 1927, 1302, 2227, 1694, 931, 2411, 128, 2509, 2200, 1052, 884, 1079, 1766, 1221, 810, 774, 201, 1463, 812, 1415, 2500, 1114, 780, 1220, 2197, 2541, 2592, 1295, 2460, 1282, 1911, 2572]. Bayesian/likelihood [1818]. Bayesianism [800]. be [61, 118, 1892, 2278, 660, 959]. Beaty [1165]. Beck [401]. Becker [44]. behalf [770, 165, 1037]. Behavioral [1888]. behaviour [1712, 674, 1779]. Behavioural [210, 1849, 482]. Behind [693]. being [2627]. Bell [1876]. Belle [2312, 1278]. Belmont [365, 1799, 2070, 2040]. Bemmel [2179]. beneficiaries [252]. Benefit [553, 972, 2569]. benefits [352, 107]. Benhamou [1813]. Benichou [2644]. Bennett [348]. Beolchi [2009]. bereavement [230]. bereavement-is [230]. Bergen [137]. Bergman [570, 1001]. Berlin [403, 480, 362, 1887, 384, 91, 404, 2132, 555]. Bernard [1824, 400]. Bernice [1165]. Berry [1911, 166]. Bert [346, 830]. Bertil [1001]. best [350, 458, 919, 2582]. Beth [382]. between [602, 475, 559, 1751, 504, 908, 781, 952, 2490, 535, 1151, 427, 1371, 1658, 1074, 176, 2295, 2123, 890, 1480, 1754, 2554, 2174, 1664, 2508, 139, 1152, 775, 2633, 343, 2233, 1880, 2336, 869, 989]. Bias [9, 1884, 1273, 523, 1635, 1736, 1646, 1794, 2262, 531, 971, 160, 2535, 2260, 1522, 761, 2453, 233, 256, 547, 2335, 2026, 1805, 981, 1102, 2360, 2145, 2190]. biased [2477, 1885, 2643]. biased-coin [2477, 2643]. biases [2243, 548, 1216]. Bibliography [73, 2309, 424, 408]. Bijleveld [1154]. biliary [606]. Billard [1838]. Binary [1209, 1354, 2321, 2288, 2332, 2558, 1445, 2398, 2519, 2052, 468, 664, 1020, 2077, 1737, 1735, 2013, 1975, 1044, 1577, 181, 1043, 2237, 1077, 1959, 2028, 784, 2358, 884, 2255, 626, 729, 1939, 501, 1190, 103, 2078, 2328, 7, 2635, 1223, 1580, 2413, 969, 274, 2274, 1626]. binding [19]. binocular [341, 886]. binomial [1097, 174, 590, 321, 1394, 1345, 742, 1047, 2295, 1356, 2645, 2633, 741]. binormal [1623, 1895, 1805]. bioassay [725, 1272]. bioassays [868]. bioavailability [52, 872, 2428, 354, 517]. bioavailability/bioequivalence [52, 354, 517]. biochemical [985, 290]. Bioequivalence [610, 1821, 52, 562, 872, 301, 1781, 354, 517, 1273, 1304, 302, 772, 1181, 1741, 1968]. Biologic [1702]. Biological [2018, 1877, 1518, 210, 2564, 108]. Biology [2081, 2345, 1703, 731, 433]. biomarker [474, 678]. biomarkers [2423]. Biomathematics [384, 404]. biomedical [46]. Biometrie [1628, 2056]. Biometrische [1962]. Biometry [1838, 1850, 1960]. biopharmaceutical [289, 146, 871]. Biopharmacy [288]. Bioscientists [2278]. Biostatistical [1366, 2602, 1799, 1364, 1365, 2224, 2181, 2022, 1629, 245]. biostatistician [1367, 2315]. Biostatistics [59, 153, 388, 896, 1169, 1170, 1458, 2067, 2001, 1876, 2545, 2637, 1846, 452, 1845, 2628, 91, 1886, 957, 1690, 738, 2023, 1684, 1911, 2133, 2311, 2264, 1153, 655, 400, 382, 328, 1278, 1823, 1015]. Biostatistics. [2312]. Biplots [46]. Birkhäuser [2081, 2419, 1166, 2384]. Birth [687, 2313, 2437, 445, 2468, 957, 2085, 1535, 1664, 1906, 1677]. Birth-weight-adjusted [687]. birthdates [930]. birthday [2622]. births [2502]. birthweight [1742, 1600]. bisexual [574]. Bishop [1984]. Bivariate [2502, 1832, 2595, 2594, 1576, 1363, 1446, 1333, 1905, 2596, 2542, 2374, 1606, 1012, 1311, 898, 676, 1617, 2575, 2410, 2470]. Bjørn [284]. black [313]. Blackwell [2368, 362, 944, 284, 893]. bladder [995, 2498]. Bland [1971]. bleeding [217]. blind [2079, 1939]. Blinded [1806]. blindness [1607]. block [319]. blocks [913]. blood [505, 1676, 2015, 644, 2438, 2469, 661, 546, 99, 1893, 1159, 137, 834, 359, 1041, 621, 1191, 1092, 2388]. BLUPs [2582]. Blythe [2278]. BMJ [2408, 2039, 1859, 2008]. Board [2627, 693]. boards [711]. Boca [786, 2587]. body [1599, 2084]. Böhning [890]. bomb [2325]. bone [1903, 2271, 901, 2342, 2361]. Book [23, 1961, 1901, 614, 1985, 2584, 434, 432, 656, 770, 758, 1087, 827, 1812, 2253, 510, 945, 597, 1938, 1934, 2252, 2310, 2462, 381, 363, 553, 1798, 2397, 345, 1826, 828, 2394, 1165, 1984, 1027, 21, 2090, 2147, 2067, 596, 2368, 1048, 1359, 2276, 1922, 2277, 90, 452, 2367, 840, 2312, 2055, 146, 730, 94, 366, 598, 403, 1983, 482, 480, 132, 2408, 326, 1902, 2092, 1876, 244, 362, 613, 2131, 2314, 401, 871, 2602, 1887, 1014, 2039, 1026, 644, 384, 2009, 433, 91, 365, 731, 92, 1962, 2204, 2275, 2254, 2018, 364, 144, 1153, 2148, 346, 786]. Book [402, 1859, 1799, 206, 451, 2041, 892, 111, 1786, 2070, 1936, 554, 2545, 1911, 2396, 450, 2042, 759, 2369, 1878, 147, 2379, 20, 2038, 2134, 872, 2263, 1898, 2296, 2286, 145, 1192, 2081, 93, 2157, 1085, 1899, 2006, 2179, 1193, 110, 1000, 1994, 1888, 1970, 1862, 933, 511, 165, 2495, 328, 667, 1086, 1194, 944, 2573, 2378, 404, 2345, 829, 1923, 1004, 2169, 788, 947, 1154, 2586, 327, 2278, 22, 1824, 1370, 1371, 1455, 1372, 1454, 1443, 2601, 2091, 1848, 1896, 2068, 841, 2419, 2599, 2007, 2194, 2132, 242, 1972, 2461, 2572, 570, 2544, 2546, 2313, 2082, 1924]. Book [612, 1951, 2040, 2417, 2418, 1847, 1933, 1037, 1278, 655, 112, 24, 946, 2380, 243, 1849, 481, 1813, 1823, 164, 1687, 1015, 1950, 148, 2587, 285, 2133, 2494, 1897, 2439, 1002, 400, 1973, 2311, 668, 870, 2571, 1787, 1055, 1038, 2585, 526, 2150, 207, 2093, 1982, 1971, 921, 2043, 1166, 1860, 1900, 2264, 1838, 1850, 383, 1338, 1877, 2600, 2547, 1861, 2053, 2366, 528, 1108, 1701, 2383, 2170, 2384, 2181, 2168, 284, 1981, 325, 348, 787, 2019, 1003, 555, 166, 2395, 2192, 347, 43, 2149, 2239, 246, 1937, 527, 44, 1277, 382, 1825, 2381]. Book [1630, 2069, 1912, 621, 830, 1025, 2056, 2054, 1846, 1788, 1315, 1314, 1960, 2020, 2180, 893, 1001, 2193, 920, 2008, 184, 1164, 2265, 1811, 2178, 2382, 1704, 245, 1935, 2167, 1876]. Books [1087, 1950]. Bookstein [731]. Bootstrap [2479, 2375, 2376, 2556, 1291, 640, 2049, 338, 1388, 2228]. Boreham [1897]. Borgan [1108]. born [2061]. Borrowing [1747]. Bosten [667]. Boston [1087, 452, 1026, 2081, 1454, 481, 2133, 400, 526, 1166]. both [1177, 188, 952, 1351, 2327]. bound [2177, 2474]. boundaries [1059, 1432, 1058]. boundary [1940]. bounded [765, 2228]. bounding [1399]. bounds [1081, 1857]. Bourke [944]. Bouvier [2314]. bovismorbificans [369]. Bowel [2104]. Bowers [2041, 2439]. Box [596]. Boyle [362, 2473, 2193]. Bradford [739]. brain [1505]. Breaking [1347]. Breast [2102, 2096, 2104, 77, 1762, 2083, 1378, 1351, 2616, 1211, 2100, 2085, 1537, 1889, 454, 450]. breast-fed [2083]. Breen [1923]. Brenner [2128, 1948]. Brian [667, 2082, 2069, 1164]. Brillinger [1838]. Bristol [148]. Britain [1441]. British [363, 1798, 2397, 21, 401, 1014, 1544, 2084, 704]. Britt [2019]. Britton [362]. Broemeling [2066]. Brown [345, 452, 1876, 401, 667, 1877]. Brownian [735]. Bruce [1825]. Bryan [433, 947]. budgeting [1690]. BUGS [2564]. Building [831, 793, 640, 2026]. Bulman [363]. Buncher [1935]. burden [1111]. burn [681]. burn-care [681]. Busby [1950]. Butler [2369]. Butterworth [614, 1973]. Büttner [1822]. bypass [779]. Byrne [2571]. Byron [110].

C [434, 770, 1087, 896, 2310, 2089, 730, 1876, 362, 1026, 384, 92, 2127, 20, 1193, 1888, 933, 404, 1154, 1443, 2473, 1809, 2133, 921, 2043, 1877, 2645, 2643, 1277, 1788, 2382, 1626, 2126, 2333, 1935]. C.A.MAN [861]. CA [1899, 2040]. CABG [2497]. Caesarean [2425]. calculate [1554]. calculated [2576]. Calculating [587, 392, 361, 1718, 2244, 342]. Calculation [765, 1341, 2301, 2421, 1349, 157, 590, 2201, 1565, 507, 1743, 996, 2411, 275, 1350]. calculations [1783, 2598, 1568, 2202, 1355, 2203, 2460, 897]. calculator [665]. CALGB [1053]. calibration [2638, 1517, 2165, 1035, 1932]. California [2131, 365, 1799, 1900, 1656]. Cambridge [770, 1169, 1170, 2252, 731, 1936, 2369, 1898, 947, 570, 1630, 1912, 1001].

Campaign [2097]. Campbell [2368, 362, 285, 2008]. Can [1123, 1892, 2278]. Canada [146, 1055, 1271, 2103]. Canadian [705]. Cancer [770, 2152, 2106, 2102, 2096, 2097, 451, 2104, 1813, 2101, 2103, 2193, 2195, 2437, 102, 1762, 2095, 730, 1529, 1780, 403, 1542, 2014, 198, 771, 1201, 132, 587, 1378, 2284, 362, 2362, 335, 420, 1351, 92, 1067, 2289, 1802, 1863, 2117, 2109, 1011, 1066, 1453, 1352, 98, 696, 1053, 450, 36, 637, 422, 443, 1446, 93, 1925, 418, 914, 1521, 858, 972, 1370, 1443, 2556, 1522, 2173, 1211, 77, 1233, 2227, 2569, 1656, 2411, 1539, 1287, 200, 1015, 995, 389, 78, 2100, 2509, 1055, 1065, 1532, 49, 2085, 1051, 2432, 780, 2116, 2108, 1239, 558, 419]. cancer [1537, 2240, 1465, 1083, 2498, 1324, 1889, 454, 770, 704, 480, 451, 1193]. cancer-mortality [443]. cancers [2616]. candidate [2392]. canonical [1916]. capture [2048]. capture-recapture [2048]. capturing [713]. carcinogenesis [69, 1562]. carcinogenicity [1579]. carcinoma [1148]. CARDIA [2387]. cardiac [1907, 1077, 127, 12, 775, 1019]. Cardiff [2378]. cardiopulmonary [2274]. Cardiovascular [1862, 462, 1376, 547, 839, 1358, 796]. Cardon [933]. Cards [1896]. Care [2041, 2006, 1154, 1823, 2101, 2265, 434, 553, 413, 402, 20, 681, 687, 252, 2618, 1618, 2439, 1459]. career [1362]. carinii [1890]. Carl [1359]. Carlin [2382]. Carlo [433, 2169, 2455, 2611, 2467, 389, 1079, 2541]. Carroll [2007]. carry [629, 498, 2318]. carry-over [629, 498, 2318]. carryover [479, 353]. Carstensen [132]. CART [1814]. Cartwright [432]. Case [2644, 2465, 1860, 940, 803, 1647, 2223, 1842, 1438, 1736, 477, 228, 398, 979, 1762, 800, 370, 793, 399, 2230, 79, 311, 38, 276, 760, 1352, 654, 2616, 1214, 2324, 1425, 1392, 1064, 1620, 980, 611, 1299, 550, 2046, 1883, 1554, 178, 1246, 1329, 902, 1266, 177, 608, 476, 1841, 1224, 591, 1602, 2452, 928, 263, 488, 312, 589, 839, 1358, 1947, 1619, 835, 1160, 1978, 1917, 1469, 976, 715, 802, 2354, 1838, 1315]. case-base [38]. case-cohort [589, 839, 1358]. Case-Control [2644, 2465, 940, 2223, 1842, 477, 398, 979, 1762, 370, 793, 399, 79, 311, 276, 760, 1352, 654, 2616, 1620, 980, 611, 550, 1883, 178, 1329, 1266, 177, 608, 476, 1841, 1224, 928, 488, 312, 1619, 835, 1978, 1917, 976, 2354]. case-crossover [902]. case-distribution [2324]. case-only [2324, 928]. case-ontrol [228, 263]. Case-Study [1860]. case/control [1736]. casebook [2054]. cases [1979, 1931, 609, 608, 233, 1334, 1465]. Castillo [384]. Castillo-Chavez [384]. Categorical [1983, 2194, 2544, 864, 2518, 424, 425, 2317, 1832, 352, 505, 963, 1712, 1411, 1568, 1371, 378, 1333, 2226, 1355, 1034, 964, 1687, 2521, 526, 2322, 1331, 897, 1617, 612]. categories [1649, 393, 2565, 1179]. categorization [1768]. Catherine [1902, 364, 347]. Causal [2275, 1030, 2466, 2077, 824, 243, 1900, 184]. Causality [2020]. causation [810]. cause [1696, 1621, 2323]. caused [2026]. Causes [451, 510, 1677]. cautionary [1919, 241, 2250]. CBS [1823]. CD4 [2242, 1710, 743, 1473, 1118, 1389, 2272, 1031]. CD4-lymphocyte [743, 1118]. cell [2438, 1053, 1473, 1065, 1796]. Censored [2092, 1923, 2595, 2594, 2156, 521, 1695, 795, 1492, 2435, 1636, 2350, 1559, 1187, 1466, 1402, 2495, 1136, 577, 586, 2064, 604, 2596, 622, 359, 1789, 1604, 1790, 1448, 441, 1013]. censoring [823, 2566, 1565, 489, 591, 564, 1105, 977]. census [141, 860]. cent [2582]. Center [1937]. centile [1742, 560, 392, 1774]. centiles [750, 2084, 51, 1161]. central [1539]. Centre [153, 1458, 2455, 763, 2216, 1607, 2215, 2356, 2227, 2101, 2100, 2214, 2498]. centres [2356]. Century [2204, 22]. certain [2505]. Certainty [1934]. cervical [952, 884]. Cesar [388]. cessation [1607, 37, 824]. Chain [2169, 2186, 1351, 725]. Chairman [815]. challenge [551]. Challenges [171, 2518, 2095]. Chalmers [434, 2408, 2039, 20]. Chamant [1192]. Chamberlain [770]. Chambless [1424]. Chamge [595]. Chan [2474]. Chance [2310]. change [324, 504, 781, 885, 951, 681, 39, 595, 489, 526, 988, 747, 674, 1152, 1420, 2434, 1213, 1010, 2291, 989, 104, 2132, 1001]. change-corrected [885]. Changes [1588, 821, 1679, 2564, 601, 2322, 1210, 1489, 2555, 977]. changing [573, 2421, 1723, 2045]. Channeling [256]. Chaos [2168]. Chap [2367, 2181]. Chapman [1985, 2253, 381, 94, 433, 2134, 110, 2169, 327, 2599, 2007, 1951, 668, 1038, 1860, 528, 2168, 44, 2381, 2069, 1314, 2382]. Characteristic [1869, 17, 2140, 1495, 1963]. characteristics [1234, 1995, 362, 2615]. Characterization [1484]. characterize [1215]. Characterizing [1815]. Charles [656, 2070, 164]. chart [2302]. charts [1177, 2328, 1774]. Chastang [1902, 364]. Chatfield [1985]. Chatterjee [2054]. Chaudhary [2089]. Chavez [384]. checking [2499, 563]. Checkoway [345]. chemisch [1628]. chemisch-pharmazeutischen [1628]. Chemistry [2369]. chemotherapy [1378]. Chen [362]. Cheshire [758, 2192]. chest [1814]. Chestnut [511, 527]. Chi [1972]. Chi-Squared [1972]. Chichester [1961, 1812, 730, 482, 2018, 2041, 2263, 165, 2345, 2586, 2082, 946, 285, 1787, 1838, 1981, 1277]. chief [451]. child [2402, 1914]. childbirth [2437, 434]. Childhood [546, 1534, 2282, 2173, 1656, 1532, 1441, 1184, 2456]. children [2061, 505, 368, 136]. China [362]. Chinese [362]. Chlamydia [2145, 1884, 2190]. chlamydial [484]. choice [1429, 1138, 1244, 12, 1083, 8]. cholesterol [2469, 755, 2582]. Choosing [2047, 350]. chorionic [2515]. Chow [872]. Chris [1985, 2018, 2278, 1950]. Christensen [2134]. Christine [1359]. Christopher [1984, 2091, 246]. chromosome [1394, 995]. chromosome-loss [995]. chromosomes [204]. chronic [2338, 1903, 2446, 439, 2473, 179, 1459, 2171, 2172, 1245, 2228, 2273, 23]. Chung [872]. cigarette [1521]. circumference [2084]. circumventing [2437]. cirrhosis [606]. Citations [912]. City [828, 1991, 1931, 1242, 1664, 137, 1117]. city-level [1242]. CL [401]. CL-7279-4290-3 [401]. claims [255]. Clarendon [787, 245]. Clark [1897, 2381]. class [2118, 141, 204, 1975, 2306, 641, 1430, 53]. Classification [1967, 2565, 1665, 430, 229, 1481, 2543, 2422, 204, 1808, 2391, 561, 1890]. classifications [1603, 1948, 1682, 1179]. Classifying [1601, 2326]. Claude [1902, 364]. Clayton [1194]. clear [2222]. Cleveland [1786, 1338]. Cliff [362]. Clifford [1788]. Clinical [1961, 59, 153, 388, 767, 896, 1169, 1170, 1458, 2368, 2106, 1871, 1325, 1859, 624, 147, 145, 2586, 1848, 2103, 1055, 1784, 2383, 810, 1937, 1684, 2076, 280, 2195, 2285, 1797, 1158, 2472, 2466, 777, 281, 323, 1366, 72, 597, 298, 2300, 2152, 61, 294, 2641, 1070, 1746, 1594, 102, 305, 1515, 2095, 231, 1748, 800, 1208, 1646, 2320, 723, 638, 1090, 1702, 1178, 1071, 706, 2628, 101, 825, 1476, 1571, 1023, 1588, 960, 2014, 2267, 198, 771, 1462, 2442, 352, 764, 2398, 2114, 1843, 188, 414, 195, 2393, 1584, 1864, 420, 91, 2580, 535, 1644, 459]. clinical [1731, 428, 699, 1308, 1101, 1639, 157, 89, 1863, 2117, 2109, 300, 1151, 346, 786, 1886, 1590, 1837, 2445, 2196, 1011, 134, 462, 1318, 421, 1060, 603, 2029, 1908, 566, 98, 554, 709, 1596, 696, 1059, 1053, 1752, 1498, 565, 637, 423, 20, 1091, 514, 697, 100, 422, 1336, 211, 1446, 154, 2098, 379, 99, 2125, 891, 1925, 2535, 418, 636, 155, 1085, 1000, 1421, 667, 833, 749, 1086, 1831, 320, 1607, 1628, 2215, 2639, 351, 2099, 577, 567, 841, 2226, 170, 735, 887, 1391, 1680, 2003, 2531, 2556, 2484, 377, 234, 924, 1365, 2496, 1074, 743, 1679]. clinical [2647, 2030, 2086, 2453, 1156, 1061, 2112, 2164, 2486, 1171, 1460, 2534, 2589, 2227, 946, 1198, 996, 60, 1643, 814, 1306, 1419, 604, 1974, 1118, 130, 1138, 1310, 389, 2561, 78, 718, 1389, 350, 870, 1909, 704, 536, 67, 808, 1592, 2400, 259, 2629, 1479, 1072, 2527, 2, 1221, 254, 1468, 4, 1996, 1701, 660, 2229, 904, 1490, 1715, 2348, 411, 834, 158, 391, 943, 932, 1100, 1461, 1420, 1939, 1865, 2414, 903, 407, 959, 1063, 492, 196, 1051, 812, 347, 2640, 1678, 2308, 81, 6, 780, 1220, 2197, 382, 253, 2143, 2108, 2228, 830, 804, 1321, 1828, 1489]. clinical [2002, 390, 802, 1179, 1058, 2032, 2523, 702, 1714, 693, 1010, 2498, 1722, 1889, 66, 8, 104, 91, 92, 93, 667, 1937, 1277]. clinician [259]. Clinicians [2379]. Clogg [1888, 1788]. cloning [41, 183, 429]. Close [402, 1672]. Closed [491, 313, 273, 1084, 2162, 693]. Closed-form [491]. Closing [720, 817, 1260, 1671, 2620]. cloth [730, 2254, 1838]. Clues [846]. Cluster [456, 1556, 1870, 844, 1543, 1537, 1354, 2424, 2614, 362, 500, 1534, 1, 1548, 1907, 1541, 2390, 1661, 853, 845, 1734, 1954, 846, 1626]. cluster-specific [1734]. Clustered [2028, 1946, 1832, 1650, 967, 1668, 1044, 2187, 2447, 1776, 2448, 1223, 1580, 1734]. Clustering [1931, 502, 1557, 682, 1536, 1538, 1733, 2005, 850, 1555, 1430, 2004, 2413, 851, 854, 1428, 1540, 2193, 790]. clusters [1946, 847, 1547, 862, 1269, 1550, 1532, 501, 1184]. Co [452, 2380, 565, 2135]. co-operative [565]. co-twin [2135]. Coakes [2585]. cocaine [1247]. Cochrane [2408]. CODA [1772]. code [2598]. coefficient [2595, 445, 2374, 1042, 1906, 1190, 1606, 1549, 2505]. coefficients [1918, 2155, 1526, 2234, 1176, 2087, 1822]. Coggon [1859]. cognitive [674]. Cohen [1165, 279]. coherence [573]. Cohort [917, 1991, 875, 1762, 2558, 399, 335, 952, 160, 107, 1172, 914, 416, 2387, 2473, 337, 1491, 1563, 455, 313, 522, 2171, 2172, 589, 1490, 839, 1358, 1031, 663]. cohorts [2590]. coin [2477, 2643]. Cold [1443]. Cole [2133]. Colin [362, 325]. Collaborative [767, 1055]. collected [723, 254]. collection [2104, 2377]. College [1085, 2571]. Collett [1314]. Collinearity [558, 2437, 1793]. colon [684, 1324]. colony [2534]. colorectal [420, 1055, 2116, 419]. Colpaert [2005, 2004]. Columbia [1544]. Columbus [827]. Com [1902, 364]. Com-Nougue [364, 1902]. combination [2588, 61, 2642, 495, 724, 1488, 2489, 182, 1967, 536, 913]. combined [604]. Combining [2488, 1248, 2445, 1390, 791, 2391, 1615, 1308, 204, 2353, 1238, 2637, 2373, 1359]. Comment [117, 120, 123, 126, 31, 32, 193, 5, 463, 30, 191, 192, 29, 464, 190, 465, 471, 33, 116, 466, 189, 516, 2646]. Commentary [936, 2347, 2562]. comments [408, 1766, 1996]. Committee [704, 2315]. committees [708, 1228, 1281, 1062, 699, 713, 697, 1064, 705, 1654, 715]. common [1354, 2424, 1080, 2576, 473, 589, 1626]. commonsense [285]. Communication [2038]. communities [2246]. Community [2046, 1234, 2246, 1542, 1347, 2372, 105, 1570, 1919, 683, 197, 2250]. community-level [2372]. comorbidity [1263]. companies [2414]. Company [2313, 667]. Comparative [940, 399, 1932, 2177, 174, 733, 170, 2189, 1574, 350, 913, 2474, 2645, 1463, 1484, 2116, 2638, 2165]. compare [264, 1541, 621, 1191]. compared [1381, 1993, 2158]. Comparing [2011, 1718, 2350, 538, 2428, 1311, 561, 2431, 1423, 2285, 82, 1445, 576, 1730, 195, 299, 2234, 321, 1418, 1055, 2337, 886, 222, 1094, 2025, 1041, 2567, 2076, 1637]. Comparison [1219, 1649, 305, 888, 609, 2117, 1908, 639, 1020, 1856, 1447, 1211, 1866, 1356, 904, 794, 1750, 274, 1186, 2455, 361, 2641, 205, 979, 277, 1742, 2482, 2242, 1481, 2398, 2262, 983, 1045, 2155, 365, 2283, 1863, 1650, 760, 1397, 1953, 1059, 2543, 625, 1839, 1432, 2215, 2648, 1355, 1110, 440, 2013, 1577, 1274, 964, 1808, 1215, 1881, 2123, 2122, 1801, 304, 2187, 2464, 2516, 2454, 838, 782, 1343, 2400, 1268, 1105, 509, 1761, 137, 2548, 2640, 1223, 1580, 2413, 2563, 1344, 2434, 2016, 2162, 1058, 441, 2456, 2360, 2632]. Comparisons [535, 2452, 281, 323, 515, 1445, 19, 188, 2549, 42, 2551]. compartmental [1466]. competence [686]. competing [198, 1621, 38, 2389, 1856, 1833, 734, 2138, 2420]. complete [1907]. completely [2538, 1930]. completeness [1780]. Completion [2103]. Complex [1787, 2285, 2288, 494, 1988, 2583, 1779]. complexity [2617, 1074]. Compliance [2072, 2097, 2104, 2105, 2100, 2078, 1843, 1910, 2577, 2268, 1752, 2077, 1140, 2578, 2075, 2074, 2079, 2073, 1527]. complications [1464]. components [485, 1368, 1891, 543, 258, 359]. Composite [975, 1218, 1449, 1865, 2143]. composition [1394]. compound [1289]. comprehensive [660, 1490]. computation [520, 757, 1080, 1422]. Computational [619, 829, 1703, 2345]. Computer [2081, 861, 2598, 2545, 1553, 244, 1862, 2381]. Computer-Aided [2381]. Computing [2131, 1349, 308]. concentration [1291, 1527]. concept [607]. conception [2154]. Concepts [1876, 2379, 1048, 1364, 1310, 2253]. Conceptual [2546, 1853]. concerning [1852]. Concluding [1597]. Conclusions [1074, 1075]. conclusive [2385]. concomitancy [869]. concordance [2595]. Concorde [2555]. concurrent [2032]. condition [1903, 2446, 232]. Conditional [1783, 2451, 837, 1717, 2302, 1987, 980, 1404, 1528, 659, 742, 734, 2048, 729, 1341, 492, 1678]. Conditioned [736]. conditions [1659, 1245]. conducted [2106]. conference [250, 2393]. Confidence [236, 580, 1097, 1820, 1188, 727, 1413, 1482, 1785, 999, 1345, 1959, 235, 1047, 1619, 741, 1024, 1819, 520, 757, 21, 2089, 2611, 1609, 1144, 1718, 157, 2244, 2648, 2377, 2499, 1404, 2295, 2122, 2454, 273, 1291, 1081, 447, 1388, 2124, 2633, 1181, 1161, 1880, 2375, 2376, 2332, 2127, 1809, 2237, 2635, 2126, 1872]. confidentiality [2370, 1662]. configuration [853]. confirmatory [2399, 2207]. confounder [178, 865]. confounders [760, 1342]. confounding [607, 79, 1296, 824]. congenital [1380]. Congress [153]. congruence [1658]. Conjecture [2275]. Connecticut [758, 382, 1521]. connection [1802]. Conquest [1838]. consensus [767, 457]. consent [428, 2647, 66]. consideration [704]. Considerations [1138, 1070, 952, 2392, 1570, 1831, 990, 1044, 791, 476, 2224, 701, 2209, 1657, 1071]. considering [1350]. consists [2278]. constant [1992, 608]. constrained [1032]. constraints [2303, 1537]. Constructing [2611, 1291, 269, 447, 2375, 2376]. Construction [750, 1992, 2349, 1774, 1400, 1218, 580, 858, 999, 290, 369]. contact [951]. contain [1390]. contamination [1843]. Contending [2073]. context [1532]. contigency [668]. contiguity [1661]. Continents [1193]. contingency [491, 373, 28, 508, 1682, 306]. Continual [2489, 771, 1317, 2443, 1287]. continuity [28, 508]. Continuous [1678, 1773, 1832, 2371, 760, 2425, 1765, 887, 490, 2066, 2255, 541, 2348, 2329, 2025, 1209, 11, 1963]. Continuously [1940, 2140]. continuously-distributed [2140]. contour [1243]. contraceptive [213]. Contradictory [343]. contrast [2549, 2551]. contributions [2624, 2625]. Control [1572, 2644, 608, 940, 1717, 2223, 1842, 2152, 1736, 2247, 477, 398, 979, 1762, 370, 793, 515, 1149, 1445, 399, 2262, 1730, 2465, 79, 299, 311, 1698, 2438, 276, 760, 1352, 654, 2616, 1620, 2135, 980, 611, 550, 1883, 178, 1329, 1266, 177, 476, 1841, 1224, 928, 488, 312, 2229, 905, 1619, 835, 1978, 6, 2016, 1917, 804, 976, 1704, 2354, 451]. control-conditional [1717]. controlled [280, 2641, 2077, 1112, 2356, 175, 2075, 2400, 2554, 4, 2640, 40]. Controlling [1956, 865]. Controls [1640, 888, 609, 188, 535, 336, 357, 291]. Controversies [1201, 2527, 1476, 1461, 2214]. Convention [1458]. Conventional [1381]. conversational [686]. Cook [2066]. Cooper [1811]. Cooperative [2098, 1607]. Copenhagen [896]. Coping [215, 998]. cord [545]. corneal [128]. coronary [35, 13, 1246, 779, 755, 775, 2027]. corrected [2471, 546, 885]. Correcting [2578, 1186, 1995, 2555]. Correction [18, 150, 239, 307, 322, 380, 460, 568, 785, 997, 1088, 1119, 2037, 2034, 2559, 1870, 2035, 2036, 1785, 2406, 2407, 2235, 2440, 2236, 2560, 2074, 2405, 28, 508, 1871, 1869]. Corrections [592, 922, 1276, 1384, 1033, 1784, 1496, 1046]. Correlated [590, 424, 2332, 2163, 2130, 1397, 443, 320, 1418, 585, 1577, 2237, 1959, 1851, 1267, 2187, 523, 1268, 501, 579, 2635, 1651, 2159, 1637]. correlation [590, 427, 1868, 442, 1368, 1176, 1268, 626, 2374, 139, 1822, 2176, 2410, 2505]. correlations [1965, 546, 1893, 2452, 49, 343, 2233]. Correspondence [1172, 348]. Corter [1899]. cortical [1881]. cost [2089, 2611, 1609, 1627, 2612, 1421, 177, 476, 2610, 660]. cost-effective [1421, 660]. cost-effectiveness [2089, 2611, 1609, 1627, 2612, 2610]. costly [2259]. costs [2567, 2006]. could [920]. Council [1148, 704]. Count [2093, 794]. counter [188]. Countering [828]. counterpoint [2280]. counties [362, 2605, 1656]. counting [1108, 613]. Countries [1897, 146, 90, 132]. Counts [2213, 962, 491, 62, 524, 1710, 743, 1928, 996, 1473, 1118, 1389, 2272, 356]. county [1532, 688]. county-level [688]. Course [2070, 2054]. Court [325]. Courtroom [1824]. covariables [1918, 1410]. Covariance [55, 2301, 1340, 2226, 751, 752, 1424, 1511, 2272]. Covariate [659, 1552, 2231, 2458, 1647, 2477, 1915, 1510, 468, 1953, 1988, 2120, 1214, 1425, 1022, 1482, 1785, 2137, 2189, 2339, 837, 80, 1768, 2643, 1757, 1083, 953]. covariates [949, 876, 2512, 1096, 1816, 2482, 1676, 2362, 1435, 1636, 1131, 1681, 1735, 2576, 681, 1447, 1767, 1765, 2536, 1386, 175, 541, 2592, 1331, 331, 2240, 2388]. cover [245]. coverage [1219]. Cox [2369, 1797, 2242, 13, 756, 180, 1264, 2566, 1735, 1369, 681, 1147, 1353, 1801, 640, 904, 1768, 137, 1778, 1749, 1173, 1916, 623]. CPMP [1366]. Crawford [345]. Crawford-Brown [345]. CRC [786, 2587, 1861]. Creation [565]. Criteria [2266, 1723, 2222, 1818, 2305, 996, 2618]. critical [1176, 2480, 1822, 184]. Criticism [2351]. critique [841]. Cross [1559, 469, 900, 319, 1633, 1844, 1177, 1697, 917, 1188, 110, 1474, 1299, 1608, 2416, 378, 1514, 1507, 1942, 1304, 1162, 2415, 1692, 2349, 2591, 1479, 2026, 1947, 1478, 1968, 945, 596, 946]. Cross-over [1559, 319, 1633, 1844, 110, 1299, 1608, 2416, 378, 1514, 1507, 1942, 1304, 2415, 1692, 2349, 2591, 1479, 1947, 1478, 1968, 945]. Cross-over. [946]. cross-sectional [1177, 917, 1188, 1474, 1162]. Cross-validation [469, 900, 1697, 2026]. crossing [823]. Crossover [1023, 353, 525, 296, 134, 318, 240, 499, 653, 1009, 1008, 629, 297, 295, 902, 742, 479, 880, 768, 1036, 1181]. Crowder [1951]. crude [506, 539]. CT [2192]. CTG [1055]. CTR [767]. CTS [81]. cubic [1735, 1410, 1022]. culture [968]. cumulative [2614, 1320, 576, 1333, 1833, 2060, 342, 2328, 2032]. cure [521, 2120, 36, 1409, 2119, 2045, 1013]. cured [109]. current [2058, 1586]. curse [1772]. curtailing [980]. Curtis [2347, 1937]. curve [445, 2614, 792, 1815, 2363, 791, 1895, 87, 1145, 2115]. curve-based [2363]. Curves [1869, 1989, 559, 1742, 560, 1375, 468, 17, 1513, 1009, 972, 1856, 506, 2140, 890, 734, 392, 1081, 1495, 1651, 1805, 1321, 1637, 1963, 2068]. cuscore [1885]. CUSUM [2511]. cusums [541]. cutpoint [1697]. cutpoints [1475]. Cuzick [1003]. cviii [1937]. cycle [214]. cystic [1509]. cytosine [1306].

D [1901, 777, 656, 1812, 2394, 2343, 2332, 2312, 2055, 2130, 326, 1876, 362, 2602, 2039, 365, 2065, 2331, 1193, 2169, 2007, 1924, 2417, 2237, 1278, 481, 1823, 2066, 1897, 890, 1860, 1850, 1108, 1874, 1003, 166, 43, 2635, 1314, 920, 2008]. D. [777]. D.C [22]. D14.95 [94]. D15 [112]. D17.50 [165]. D27.50 [110].

D40 [93]. D85 [111]. Dale [1164]. Dalene [1911]. Daly [944]. Daniel [2006, 2571, 2239]. Danish [1695]. Data [1961, 2090, 1983, 1062, 644, 731, 713, 2369, 697, 711, 2132, 1907, 1724, 1951, 1849, 791, 2043, 701, 2547, 2383, 2384, 1250, 1277, 2069, 1314, 715, 2627, 2382, 1382, 280, 1117, 2518, 2321, 2186, 2472, 949, 1696, 1323, 1234, 324, 435, 708, 717, 2370, 424, 2512, 1990, 975, 2340, 1746, 371, 1126, 1979, 1265, 2597, 979, 2095, 2102, 393, 1685, 1989, 2595, 2594, 2156, 819, 1742, 2320, 723, 1760, 525, 706, 2242, 1946, 521, 1780, 533, 1695, 849, 1579, 1481, 1462, 1844, 352, 537, 1445, 2001, 764, 1980, 2096, 1379, 1177, 754, 2519, 795, 505, 2015]. data [1320, 2371, 2111, 2114, 362, 1313, 1450, 1115, 2362, 1910, 1965, 746, 1129, 1852, 823, 2580, 939, 456, 365, 415, 648, 534, 2117, 2109, 1189, 1492, 2435, 1636, 2350, 1886, 2110, 1559, 673, 1650, 1837, 766, 963, 318, 1709, 1908, 1352, 2104, 1953, 1988, 709, 993, 2243, 1020, 17, 2566, 1093, 2616, 1466, 1340, 2543, 1091, 1172, 967, 2059, 2468, 502, 1446, 2293, 1987, 2098, 1180, 883, 1914, 2330, 2501, 1775, 1202, 1668, 1402, 1525, 1601, 1770, 1831, 611, 1235, 1136, 2151, 1472, 2522, 1568, 2202, 2105, 585, 682, 972, 1732, 378, 577, 1140, 1333, 1658, 1328, 2226, 37, 1355]. data [47, 1033, 1765, 544, 887, 2046, 1391, 586, 1434, 2484, 2630, 234, 1883, 2401, 2364, 1147, 1514, 2047, 2064, 181, 612, 1034, 2538, 354, 1563, 2185, 237, 490, 2188, 1368, 2112, 1246, 2164, 1574, 1743, 964, 1116, 1127, 1512, 685, 1771, 1905, 650, 1694, 2140, 1893, 1523, 2415, 2107, 2537, 2521, 2028, 1133, 2607, 1386, 2198, 2352, 784, 1841, 2295, 995, 2183, 1243, 2224, 2187, 341, 848, 78, 2464, 718, 1002, 258, 2200, 1581, 1146, 2358, 2596, 1930, 2581, 705, 1038, 1602, 2492, 734, 140, 1343, 1592, 2075, 129, 1566, 884, 1130, 2447, 547, 1838, 1479, 1755, 1756, 1866, 2527]. data [473, 626, 1776, 2448, 2497, 2374, 1042, 254, 1105, 488, 474, 678, 622, 1319, 886, 762, 1494, 137, 834, 1657, 941, 1939, 501, 2113, 1612, 359, 1238, 1789, 2078, 241, 1406, 899, 1978, 794, 1511, 1604, 1790, 2322, 2500, 84, 40, 1448, 2633, 1012, 1223, 1580, 1734, 2413, 6, 1651, 2231, 2507, 2116, 2108, 1830, 2054, 2073, 1757, 1031, 627, 1181, 1210, 2532, 790, 1021, 1337, 1496, 2626, 1321, 487, 1828, 1254, 897, 1617, 1322, 693, 441, 2024, 1944, 1013, 2274, 291, 2115, 620, 1637, 2027, 2354, 526, 2092, 1786, 1898, 1923, 2544, 2093, 1900, 1338, 1861, 2366]. Data [2019, 1912, 346, 2495, 1371, 2417, 2418, 112, 1687, 44, 1025, 893, 20, 2599, 570]. Data-analytic [791, 972]. Data-dependent [1907]. data-derived [1481]. Database [2045, 261, 2283, 1242, 1237, 255]. databases [264, 262, 251, 260, 259, 263, 257, 1237, 253]. datafiles [1459]. datasets [1248, 1252, 1251]. date [2468, 212]. dates [1129]. David [510, 945, 2147, 2368, 1359, 366, 613, 1026, 1859, 2042, 165, 667, 1086, 1194, 1454, 570, 1951, 2164, 1015, 285, 2439, 1838, 2053, 1960, 1001, 1872, 1206, 1197]. Davidson [2090]. Dawson [382]. Dawson-Saunders [382]. Day [770, 1596, 1618]. Dc [596]. Deadly [1994]. DeAngelis [347]. Death [230, 1950, 1696, 1621, 1247, 901, 1077, 1244, 2138, 1019, 2378]. deaths [1803]. Debates [1938]. decade [412, 2044, 407]. December [73]. decennial [362]. Deception [1798, 2397]. deciding [990, 1978, 2354]. Decision [1396, 1208, 1178, 960, 1627, 421, 450, 154, 493, 155, 2222, 2618, 1283, 1408, 1415, 1805, 1282, 91]. decision-makers [91]. decision-theoretic [493]. decisions [1085, 1245]. deck [1754]. decline [1563, 1031]. decreasing [1730, 1274]. decrement [1711]. Dedication [266, 57]. defect [1535, 1664]. defects [1677]. defence [1424]. define [2486]. Defining [1308]. definition [601]. definitions [1160]. degree [373]. degrees [353]. Dekker [945, 597, 146, 871, 1911, 872, 1192, 1970, 2181, 166, 1935]. Delany [365]. delay [1979, 2371, 2471, 1494]. delayed [2442, 2430]. Delhi [1823]. delivery [2425]. dementia [2232]. Demographic [2600]. demonstrations [1258]. DEMP [1656]. Denis [1912]. Denmark [132]. denominators [1665]. densities [950]. density [2514, 71, 1625, 1656, 1185, 238, 392, 2141]. Dental [363]. Dentistry [363]. Department [1533, 845]. departments [1958]. departure [339]. departures [2466, 474]. Dependence [2253, 659, 2336]. dependency [2323]. Dependent [1126, 2194, 1163, 1382, 876, 2512, 2242, 180, 915, 468, 744, 1735, 1410, 1620, 681, 1907, 1565, 604, 1266, 784, 1964, 2159]. depending [1918]. depression [1681]. depressive [1030]. Derivation [1916]. derivative [1322]. derived [1178, 1481, 2306, 330, 343, 1651]. describe [232]. Describing [2325, 2121, 1767, 2564]. Description [419, 217, 2598, 862, 1916]. Descriptive [1922, 1813, 280]. Deshpande [2495]. Design [945, 1633, 1640, 952, 684, 2134, 2296, 1848, 2185, 2279, 2589, 164, 2587, 784, 1787, 1784, 1657, 1939, 162, 1846, 1150, 2259, 596, 2590, 413, 2302, 399, 1407, 2062, 2393, 1584, 79, 1101, 1240, 494, 917, 402, 664, 603, 1570, 1352, 554, 1596, 637, 2050, 514, 2443, 947, 1608, 2639, 295, 2153, 1272, 197, 1306, 1287, 1419, 374, 870, 2200, 1052, 913, 1479, 2554, 1969, 1221, 1468, 660, 932, 1420, 2432, 2124, 1954, 1253, 1478, 976, 1540, 1164, 2574, 2575, 796, 2431, 1048, 206, 872, 110, 207]. Designed [2006, 2069, 1731]. Designing [777, 1421, 931, 1917, 1503, 2196, 536, 777, 365]. Designs [503, 2195, 2477, 825, 1023, 2014, 2267, 1844, 500, 428, 1452, 276, 134, 760, 1296, 499, 653, 1009, 2444, 625, 173, 1110, 629, 297, 2556, 2647, 1304, 902, 2305, 742, 1891, 479, 1692, 2561, 1645, 1632, 928, 2349, 2479, 839, 1358, 768, 222, 959, 202, 2643, 1220, 1968, 2051, 1745, 66]. Desu [481]. detect [2468, 39, 474, 678, 1428]. detectability [142]. detectable [1351]. detected [1465]. Detecting [2121, 1429, 2145, 2001, 1884, 2569, 140, 850, 2190, 1430, 2613]. Detection [1892, 1269, 354, 1733, 1184, 2628, 52, 1396, 395, 1534, 773, 2492, 1555, 643, 2176, 851, 2005, 2004]. Detels [788]. determinants [210]. Determination [297, 1784, 1423, 2285, 1393, 2330, 1668, 1818, 567, 178, 1921, 476, 1403, 1895, 2454, 1468, 953]. determinations [766]. Determining [1450, 1162, 226, 2303, 1914, 179, 1624, 660, 2429, 1282]. deterministic [1236]. Developed [1897]. developing [1504]. Development [2147, 1377, 2125, 1242, 2506, 2525, 2488, 366, 445, 2218, 1148, 2211, 2334, 1233, 1310, 718, 1955, 812, 171, 146, 2263]. Developments [410, 1887, 101, 98, 100, 99, 738]. Deviance [1113]. deviation [293]. devices [226]. devil [938]. Dfl [1154]. diabetes [2565]. diabetic [1398, 1882, 1451]. diagnoses [351, 1512]. diagnosing [1814]. diagnosis [1966, 1510, 77, 935, 1134, 558]. Diagnostic [2528, 2588, 409, 666, 1603, 966, 2642, 1623, 1560, 457, 2363, 1867, 1948, 791, 1895, 1964, 53, 1963]. diagnostics [1942, 1693, 1749]. dialling [531]. dialogue [4]. dialysis [11]. Diane [2067]. Dianne [2164]. diary [754]. Dichotomization [2425]. dichotomous [2357, 2636, 2226, 237, 1330]. Dictionary [1936, 1850, 1937]. Died [1994]. Diego [1846]. Diet [362, 2469, 1469, 558]. diet-cancer [558]. diet-disease [1469]. dietary [925, 984, 926, 751, 752, 1469]. Dietz [1166]. Difference [1389, 2553, 781, 174, 2215, 1212, 2484, 2188, 2295, 2123, 1356, 2645, 1152, 899, 2124, 2633, 1880, 2076]. Differences [1246, 1784, 2332, 1615, 1371, 39, 1997, 2237, 1959, 1468, 1865, 1174, 2143, 977]. Different [929, 1997, 2488, 1742, 486, 760, 2347, 2570, 1638, 2346, 357, 274, 1179]. Differential [1736, 2403, 454, 227, 332, 1329]. differentials [313]. Differentiating [10]. differnces [2635]. differs [177]. difficulties [505, 905]. Diggle [245]. Digit [549, 531]. dilution [2305, 41, 183, 429, 1796, 2184]. dimensional [2139, 853]. dimensionality [1772]. diminish [857]. Direct [648, 1186, 2510, 2260, 2557, 517, 1156]. directed [2507]. directional [1434, 848]. directly [1820]. Directory [459]. disabilities [1240]. Disaster [1994]. disc [1838]. discharge [1236]. discordancy [2257]. discordant [1808]. discover [1346]. Discovery [1846]. discrete [2558, 1513, 1765, 473, 1944]. discrete-time [2558]. Discriminant [532, 2013, 1107, 1795, 238, 662, 2436, 1019]. Discrimination [985, 278, 594, 204]. discrimination- [594]. discrimination-the [278]. discs [2278]. Discussants [799, 1229, 1230, 617]. Discussion [265, 695, 698, 700, 703, 707, 710, 712, 714, 716, 719, 801, 805, 807, 809, 811, 813, 816, 961, 965, 970, 974, 978, 982, 986, 1284, 1286, 1288, 1290, 1292, 1294, 1297, 1301, 1303, 1305, 1307, 1309, 1312, 2478, 2481, 2483, 2485, 2487, 2491, 2493, 172, 1071, 101, 1829, 1725, 2080, 628, 1589, 1067, 1593, 2225, 624, 1716, 1585, 2217, 1595, 1057, 1059, 632, 2221, 1721, 2551, 1061, 1587, 889, 1591, 2212, 1713, 2219, 815, 708, 717, 1597, 50, 1073, 1176, 1822]. Disease [1816, 847, 1547, 856, 2378, 1548, 15, 2045, 23, 1966, 1117, 962, 314, 2579, 2128, 530, 1437, 1436, 2590, 1429, 1840, 2129, 504, 2503, 35, 13, 2614, 1433, 2338, 1903, 362, 939, 462, 2359, 879, 1567, 993, 575, 2616, 874, 1620, 862, 141, 776, 1732, 143, 2473, 2241, 2570, 1269, 1815, 1997, 455, 1246, 1162, 394, 1998, 1473, 358, 857, 1550, 1541, 2361, 973, 755, 1853, 1459, 2171, 2172, 859, 1552, 139, 1666, 775, 1657, 845, 1223, 1580, 44, 1334, 1469, 854, 1184, 1428, 855, 2456, 317, 2470, 43, 2193]. disease-marker [939]. disease-specific [1567]. Diseases [1898, 1037, 2615, 908, 699, 439, 2243, 375, 368, 54, 584, 1082, 2457, 1430, 1334, 2273, 2394, 510, 1037]. diskette [1901, 2547]. diskettes [786, 2585]. Disks [1973]. dismap [861]. dismapwin [1552, 1553]. disorder [1030]. disorders [2282]. Dispersion [2522, 835]. display [1254]. dissolution [2490, 2401]. dissolution/absorption [2490]. distance [2302]. distinct [1677]. distorted [1511]. distributed [1078, 2330, 2140, 1733, 2005, 2004]. Distribution [864, 1132, 432, 1979, 2371, 1399, 1820, 1363, 2324, 1769, 1545, 1485, 1727, 544, 1804, 517, 178, 1345, 968, 2028, 755, 1566, 2335, 518, 342, 774, 48, 1012, 1465]. Distribution-free [864, 544, 1804]. distributional [278, 594, 1638, 1079]. Distributions [2265, 1423, 362, 963, 138, 15, 1484, 2567, 1600, 540]. Distributors [1823]. disturbed [1322]. disturbed-highest-derivative-polynomial [1322]. disulfiram [1607]. diurnal [661]. Divergent [548, 1216, 2029]. DM [1826, 403, 2092, 1887, 384, 1962, 2042, 2286, 2179, 2091, 2132, 383, 2384, 2056, 2180]. DM118 [2081]. DM128 [1108]. DM148 [1166]. DM148/sfr134 [1166]. DM158 [362]. DM48 [328]. DM49 [2020]. DM55 [555]. DM58 [2395]. DM68 [2053]. DM84 [2170]. DM98 [480]. DNA [2145, 2190, 1825, 1884, 2276]. DNA-amplification [2145, 2190, 1884]. DNA-Sequencing [2276]. do [1203, 423, 2385, 920]. doctors [1099]. Does [2272, 1069]. dog [1638]. doing [1123]. Doll [1443, 69]. domains [2412]. Don [1981]. Donald [1359, 1911]. Données [1902, 364]. Donner [2250]. donor [2154]. Don’t [2571]. doors [693]. Dordrecht [827, 933]. Dorothy [22]. Dose [2385, 1285, 1497, 2152, 1357, 2247, 370, 2001, 746, 1249, 494, 1289, 1296, 2444, 495, 1287, 130, 389, 551, 1632, 1283, 1084, 919, 222, 1238, 162, 1422, 1239, 1295, 1855, 1282, 1503]. dose-effect [1296]. dose-response [1357, 370, 2001, 1249, 494, 495, 1084, 919, 1238, 162, 1422, 1239]. dose-toxicity [130]. doses [1287, 1527]. double [2079, 1939]. double-blind [2079, 1939]. doubly [1136, 586, 163, 1448]. doubly-censored [1136]. doubly-labelled [163]. Dougherty [164]. Douglas [381, 345, 21]. Dowdle [2394]. down [1149, 299, 1287, 1892, 649, 446]. downloaded [2278]. Draper [1359]. draw [1099]. Driven [2599, 2484]. drop [891, 636, 742, 2434]. drop-outs [891, 636, 2434]. dropout [2304]. dropouts [1573, 1771, 1865, 2143]. Drug [146, 2263, 1994, 2506, 2525, 602, 188, 1910, 644, 2052, 2490, 535, 2392, 2218, 661, 2268, 1498, 1000, 1488, 2211, 2222, 1247, 1140, 985, 182, 290, 2401, 1310, 718, 1389, 256, 907, 2220, 812, 255, 2073, 1527, 702]. drug-induced [255]. drugs [2489, 1156, 536]. DS [362]. dual [1240, 353]. due [1646, 105, 1767, 1997, 199]. Duffy [921]. Dunn [206, 207]. Dunstan [787]. Duodenal [147]. Dupin [1000]. Dupin-Spriet [1000]. duplicates [1634]. duration [1802, 577, 2531, 1514]. during [1150, 2000, 421, 1906]. Duxbury [1799, 2070, 2040]. Dwyer [759]. Dynamic [903, 2612, 574, 2281]. dynamics [602, 1698, 2279, 142, 76, 2027]. dysphagia [2288]. dysplasia [964]. dyspnoea [1868].

E19.99 [1314]. E8.95 [401]. E9 [2496]. each [1149, 1554]. ear [1618]. Early [1070, 1073, 99, 1990, 1783, 1062, 652, 1053, 100, 493, 170, 1051, 136, 1978, 2354, 1071]. Ease [2296]. Eastern [2098]. Easy [2296]. Eberhard [1962]. ECG [524]. echocardiography [12]. ECOG [89, 2098, 2]. ecologic [1104, 548, 1216]. economic [2210]. ectopic [1987]. Ed [871, 526, 90, 730, 403, 362, 384, 20, 22, 1848, 243, 921, 1166, 1838, 348, 555, 166, 347, 1825, 1912, 1315, 2178]. Edinburgh [921, 1991]. Edited [770, 788, 787]. Edition [1026, 148, 2264, 2008, 2147, 452, 2545, 511, 667, 944, 788, 1002, 400, 668, 2381, 2572]. editions [95, 247]. Editor [497, 221, 451, 666, 826, 654, 552, 1442, 611, 569, 143, 1455, 593, 582, 757, 2506, 75, 647, 2649, 2641, 2128, 2343, 2089, 2638, 2332, 2636, 1476, 2130, 2644, 2145, 2393, 2165, 283, 89, 2065, 2331, 2248, 2642, 891, 1568, 2639, 2648, 2473, 1625, 2647, 1809, 2237, 2164, 1948, 2066, 890, 2474, 2190, 2645, 880, 1947, 1874, 2646, 2429, 2637, 2640, 2141, 2633, 2635, 2643, 1580, 2250, 2143, 1822, 2002, 1626, 2126, 1872, 2333, 2354]. Editorial [658, 1262, 911, 406, 513, 1183, 1361, 1374, 2441, 1502, 209, 557, 2182, 670, 2548]. Editors [2571]. edn [1087, 2310, 2368, 1902, 1026, 1153, 1878, 1687, 1630, 1935, 870]. Eds [827, 2276, 840, 480, 759, 1443, 1823, 434, 770, 1087, 2394, 21, 2408, 362, 1887, 1014, 2039, 2393, 92, 2204, 554, 2369, 1898, 2179, 1193, 1888, 1862, 2495, 2573, 404, 2345, 2169, 1154, 2586, 2132, 570, 1037, 1015, 2150, 1166, 1850, 2168, 1003, 43, 1277, 1960, 1001, 2193, 920, 1811, 1935]. educational [2278]. Edward [758, 1878, 242, 1933, 2192, 1788, 1164]. Edwards [2053]. EEG [1918, 395]. Eerola [2020]. Effect [1462, 1864, 340, 1573, 853, 1102, 1966, 2529, 1806, 79, 2577, 203, 590, 2331, 661, 1131, 1296, 2268, 2077, 1721, 1521, 416, 227, 1620, 2151, 376, 1894, 517, 442, 2486, 824, 683, 1473, 197, 1266, 1720, 2497, 2229, 2231, 1791, 1757, 1010, 2184]. effective [1289, 1421, 660, 434]. Effectiveness [2408, 596, 2089, 2611, 1609, 1627, 2612, 1800, 2293, 1139, 1967, 2610]. Effects [2010, 486, 586, 1368, 1244, 1268, 1768, 1964, 2063, 864, 2437, 2455, 278, 594, 1222, 2287, 1817, 82, 2515, 630, 2230, 1621, 296, 396, 2244, 2029, 1953, 1214, 271, 1425, 2510, 1524, 914, 2387, 2260, 1482, 1785, 1394, 629, 1927, 1030, 2137, 295, 2538, 2423, 2326, 685, 2227, 1353, 1992, 784, 479, 848, 973, 256, 2591, 273, 747, 619, 80, 880, 1657, 1463, 1511, 2413, 2033, 1311, 1830, 969, 854, 1805, 1322, 2498, 2291, 10, 2199]. effects/longitudinal [619]. efficacy [1158, 1717, 1366, 1107, 2177, 1111, 2062, 535, 1718, 160, 211, 375, 951, 2222, 677, 2153, 2327, 642, 662, 263, 2474, 2220, 2229, 196, 226, 1311]. Efficiency [1682, 2482, 1885, 399, 105, 2408]. Efficient [1080, 1081, 80, 2152, 2590, 2029]. effort [1367]. effusion [179, 8]. elderly [1806, 1240, 1868, 1376, 290, 360, 840]. Eleanor [2383]. Elements [50, 1786]. elevation [2465]. eleven [2123]. Eleventh [388]. Elisa [1002]. Elisabeth [1371]. Elizabeth [24]. Elliott [1003]. Ellis [870]. Ellison [2369]. Elsevier [1850, 1850]. Elston [1809, 2126]. Elton [2127]. Elwood [184]. embedded [1783]. embryo [2199]. Emergence [2282]. emergency [1247, 1544, 1958]. Emerging [2525, 1137]. Emmanuel [2314]. Empirical [530, 42, 1659, 200, 1007, 969, 2382, 356, 1481, 792, 335, 1032, 760, 1953, 17, 632, 733, 856, 1710, 289, 2540, 1077, 489, 1801, 2229, 1660]. empirical-Bayes [733]. EN6 [402]. encouragement [1500]. endemic [1160]. endometrial [1352]. endpoint [1572, 2099, 2486, 2534, 604, 1974, 1138, 1996]. endpoints [2285, 2280, 1011, 1318, 439, 320, 2484, 743, 1118, 2316, 2629, 2245, 1714]. energy [475, 163]. Engel [2132]. England [432, 1876, 362, 1370, 1966, 396, 1562]. English [1003]. Enkin [434]. enrollment [1724]. enrolment [2347, 2346]. Ensuring [2107]. entities [15]. entry [1351, 996, 540]. Environment [2378, 828]. Environmental [308, 2038, 1662, 1669, 1435, 368, 783, 1440, 1522, 2339, 1666, 1003, 1239]. Environments [2131, 1547]. Envisioning [758]. EORTC [2099]. epidemic [675, 602, 1519, 1349, 2059, 574, 404, 485, 1130, 1137, 1912]. epidemics [677]. Epidemiologic [252, 840, 2256, 2175, 733, 1792, 1556, 2630, 1233, 738, 1025]. Epidemiological [1455, 2133, 916, 161, 1218, 362, 1965, 1907, 1434, 1329, 1486, 1243, 2224, 1793, 921, 837, 1238, 225, 1750, 558, 1981]. epidemiologically [1600]. Epidemiologische [1826]. Epidemiology [511, 243, 527, 510, 71, 1201, 1433, 480, 362, 1610, 335, 1435, 91, 1325, 408, 2624, 328, 667, 169, 2530, 783, 1440, 1611, 1166, 547, 860, 1003, 920, 1763, 2574, 2575, 2067, 2254, 1194, 1813, 2150, 1315, 2178, 345, 1165, 1048, 403, 384, 511, 246, 527]. epidural [1150]. EQS [2395]. equal [737]. equality [666, 2130, 1526, 2234, 1851]. equalizing [1656]. Equation [2461, 2395, 926, 964, 838, 312]. Equations [2379, 2163, 2558, 533, 1335, 1668, 1034, 1480, 202]. Equivalence [1874, 2124, 2259, 2177, 795, 1993, 2208, 1151, 1113, 867, 2330, 234, 1383, 1921, 2480, 1403, 2074, 1356, 2209, 2474, 1616, 905, 1745]. Eradication [2394]. Eric [244]. Erikssen [2429]. Erkki [1787]. Error [2007, 2242, 55, 1046, 2433, 1965, 2577, 1650, 279, 423, 2576, 173, 1033, 1765, 332, 2423, 605, 2516, 2358, 751, 752, 2255, 837, 1766, 1268, 330, 1493, 342, 774, 2231, 1830, 2631, 2574, 2575, 2375, 2291]. errors [1816, 1572, 2010, 1773, 2422, 1504, 1764, 2499, 176, 2449, 2066, 284, 1964, 2240, 206, 207]. escalation [2152, 2444, 1632]. essay [1518]. Essential [1860, 2384]. Essex [2067]. Essex-Saulie [2067]. establish [795, 1993, 1151, 1403]. Establishing [651, 1357, 2222]. establishments [2278]. Estève [1813]. Estimability [199]. estimate [2649, 216, 979, 1450, 1986, 1711, 1162, 968, 584, 775, 1330, 1350, 1471]. estimated [1892, 1727, 549]. Estimates [1117, 1872, 491, 1436, 1817, 486, 929, 1349, 396, 1113, 2609, 1296, 1413, 227, 1620, 1411, 539, 2606, 1998, 200, 212, 2048, 330, 1160, 2582, 1332]. Estimating [1842, 1599, 1979, 925, 1946, 1780, 1609, 2644, 1676, 2371, 1965, 2463, 2175, 1802, 588, 1289, 107, 1187, 2077, 2616, 315, 914, 2151, 926, 585, 484, 1342, 376, 1634, 1894, 507, 824, 478, 341, 2337, 140, 601, 2060, 339, 139, 886, 501, 2025, 202, 48, 1031, 1779, 1158, 398, 1762, 1175, 2163, 521, 2558, 533, 2062, 1335, 654, 1729, 1668, 545, 972, 2058, 1034, 2153, 2540, 1743, 964, 2569, 838, 1480, 1969, 2174, 312, 196, 1007, 1757, 1013, 2360, 2089, 2331]. Estimation [1354, 397, 939, 1351, 293, 38, 2389, 375, 1800, 2293, 2510, 218, 1970, 416, 1474, 1765, 930, 1044, 176, 517, 2486, 197, 522, 142, 67, 270, 518, 2374, 589, 2459, 359, 2161, 2434, 1334, 2281, 2336, 2261, 1635, 2424, 1989, 475, 71, 2402, 372, 792, 195, 580, 50, 1289, 2359, 744, 1953, 36, 2576, 2467, 1136, 2260, 1212, 1431, 999, 292, 1030, 2556, 2617, 138, 2188, 2504, 1368, 722, 1659, 2140, 489, 2352, 605, 1841, 2123, 1618, 2119, 523, 2542, 199, 392, 135, 839, 1358, 1657, 1939, 1190, 1606, 201, 51, 2231, 969, 2233, 2024, 1465, 2159, 2184, 2505, 1626]. estimator [205, 2594, 2301, 1032, 590, 2136]. estimators [277, 639, 2389, 2593, 2215, 544, 1710, 1891, 238, 509, 1749, 1213, 274]. Ethical [730, 701, 1070, 1071, 1752, 808]. Ethics [2315, 1691, 164]. Etiological [2125]. Eugene [597, 554]. Europe [706]. European [90, 2046, 1365]. Eurotext [147]. Eva [2601, 2164]. evaluate [262]. Evaluating [2612, 1824, 2570, 743, 685, 1236, 1118, 1782, 1018, 226, 2199, 1926, 1150, 723, 352, 2230, 1709, 951, 1504, 1142, 1867, 1507, 2569, 2353, 1420, 2637, 1701]. Evaluation [764, 672, 681, 2101, 121, 2335, 1191, 1749, 317, 2195, 25, 435, 2611, 966, 2145, 1943, 2052, 958, 673, 206, 211, 1884, 493, 550, 130, 1310, 207, 1283, 1700, 2190, 80, 674, 780, 2228, 356, 2273, 621]. Evaluations [578, 2210, 687]. evaluative [1449]. Evans [945]. Event [2530, 1561, 2320, 1834, 2127, 1576, 727, 867, 2293, 1769, 1831, 1418, 1835, 1809, 1974, 605, 2335, 342, 1952, 1830, 2619, 745, 2126]. event-time [2619]. events [334, 1836, 1407, 1834, 1557, 2213, 2486, 1533, 2534, 2589, 996, 1853, 1969, 2511]. Everitt [1936, 242, 668, 2366, 1164]. Evert [1849]. Evidence [1824, 2192, 516, 1200, 1592, 2220, 1701, 2646, 1677, 1684]. Exact [2177, 2433, 2377, 1921, 1412, 2409, 729, 519, 1422, 1483, 1219, 1131, 321, 1080, 1528, 2474]. examination [2000, 2435]. examinations [2500]. examine [1882]. Examining [2553, 272, 1620]. example [324, 2288, 2089, 1806, 1609, 760, 2341, 234, 2403, 1538, 779, 129, 136, 1952, 745, 1019, 1890]. examples [2598, 1412, 2555, 2314]. excess [2339, 1618, 1619, 2456, 1904]. exchange [1142]. Exhibition [153]. exist [1782]. existing [857]. exists [919]. exit [1351]. exloring [1079]. expectation [1144]. Expected [2158, 888, 1474, 2617, 199, 275, 2016]. expecting [2023]. expenditure [163]. Experience [704, 1062, 2469, 2105, 2107, 2101]. Experiences [2315, 2097, 1490]. Experiment [2252, 2000, 931, 2296]. Experimental [1329, 2587, 971, 494, 1306, 1069]. Experiments [2006, 72, 2247, 1579, 1633, 983, 2052, 365, 402, 984, 1156, 175, 2203, 1969, 641, 1114, 76, 1282, 2069, 945]. expert [1481]. expert-derived [1481]. Explained [1674, 1675, 2331, 47, 1894]. Explaining [2372, 2563]. explanation [2343, 2031]. Explanations [2192]. explanatory [1868, 1643, 2255, 2291]. exploit [1917]. exploiting [2184]. explorations [2243]. Exploratory [1542, 1433, 1535, 2308, 1881]. explore [2497]. Exploring [2608, 338, 482]. exponential [2359, 1482, 1785, 138, 953]. Exposure [951, 940, 2579, 979, 1762, 1450, 2175, 952, 575, 2269, 1342, 2570, 1033, 2046, 1815, 2173, 178, 522, 1293, 199, 1493, 312, 1657, 865, 1239, 1917, 1791]. exposure-disease [1815]. exposure-specific [979, 1762]. exposures [2590, 760, 733, 523, 2199]. expulsions [215, 998]. extended [1605, 2307, 2136]. Extending [875, 2163, 874, 1008]. extension [2595, 1287, 2436]. Extensions [652, 300, 2258, 2520, 2197, 2157]. external [942, 699, 1747, 2278, 1592, 1237]. extra [2649, 215, 998, 1986]. extra-Poisson [2649, 1986]. Extracting [2316]. extraction [783]. extraneous [135]. extrapolation [2467]. Eye [829, 526, 1602]. eye-tracking [1602]. eyes [1480]. Eylenbosch [43].

F [432, 363, 840, 433, 1870, 947, 1443, 1848, 148, 890, 2474, 2547, 880, 787, 1277, 2193]. F.-P [890]. Facey [2506]. Facilitated [1910]. facilities [1184]. factor [1836, 504, 234, 1956, 1246, 2534, 605, 1758, 2045, 10]. Factorial [134, 825, 724, 1488, 374]. factors [2351, 2102, 215, 998, 1892, 575, 1296, 1376, 179, 1997, 1211, 824, 642, 108, 2174, 135, 906, 1757, 11, 1882, 454]. Facts [1455]. faculty [250]. failure [2512, 393, 525, 1837, 318, 1953, 2389, 2271, 1770, 2151, 1212, 1767, 1727, 1147, 2164, 824, 734, 1604, 1790, 40, 213, 623, 1828, 2410]. failure-time [2151]. falciparum [1466]. Fallacies [144, 1455]. false [1892, 1956]. Falsificationism [391]. families [2457, 933]. family [1699, 1096, 2319, 173, 2457]. familywise [1572]. Farewell [2055, 1166]. Farrington [2645]. fat [984, 2582]. fatal [2138]. Fay [2649]. Fayers [2368]. FDA [718]. Feasibility [1243]. feasible [715]. features [1600]. febrile [2534]. fecundability [2360]. fed [2083]. Federation [402]. Feed [2156]. Feinleib [759]. Feinstein [1982]. female [1137]. female-to-male [1137]. females [2605]. Ferlay [1193]. fertile [214]. fertility [210, 211]. fertilization [1381, 820, 2199, 651]. Feuer [2649]. fewer [217]. FF [1902]. FF190 [364]. fibre [1469]. fibrosis [1509]. field [917, 951, 1601, 1037, 543, 2178]. Fieller [1181]. Fifteenth [1458]. figures [1455]. filaria [2323]. file [1236]. files [1235, 860]. fillings [1382]. final [1545]. final-size [1545]. finding [1287, 389, 1283, 1285, 1497]. finite [36, 2028]. Finkel [1900]. Finkelstein [2065, 2164]. Finland [132]. Finn [2417]. First [1957, 1197, 912, 1362, 2127, 1596, 727, 1466, 22, 1809, 1127, 2426, 2335, 342, 992, 2054, 893, 2126]. Fischer [1826, 1887, 1962, 2056, 2180]. Fisher [2312, 1278, 938, 348]. Fisherian [2380]. fit [1219, 1379, 580, 2341, 1839, 999, 2403, 927, 1777, 518]. fitted [2084, 2564, 784, 137, 2450]. Fitting [1009, 864, 949, 1742, 1792, 87]. Five [1193, 875]. fixed [2455, 2033]. Flammarion [1902, 364]. Fleming [613, 2331]. Flexible [2139, 2267, 652, 1513]. Floating [311]. floor [1511]. Florida [786]. flow [941]. fluvoxamine [2537]. focal [2027]. focused [1430, 1428, 1540]. foetal [1341]. folinic [1055]. Follies [144]. follow [1842, 2010, 2467, 1941, 1958, 547, 1715, 846, 540]. follow-up [1842, 2010, 2467, 1941, 1958, 547, 1715, 540]. following [2288, 2271, 524, 774, 775]. forces [106]. forecasting [2211]. forecasts [1958, 1130]. Forensic [1824]. Foreword [2072, 2298, 1121]. Forge [2547]. form [491, 1363, 2278]. formal [2224]. forming [334]. forms [2111]. formula [1186, 609, 1808, 2329]. Formulae [1870, 1468, 1784, 174, 1, 538, 1356, 2645]. formulating [2353, 2637]. formulation [335, 517]. formulations [913]. Forrest [450]. Forthofer [1846]. fortune [1054]. forward [2156]. Foundation [1811]. Foundations [1887, 935]. four [2641, 418, 1574, 1656, 2400, 1532, 1761, 2640]. four-armed [2641, 2400, 2640]. Fourier [238, 1041]. Fourteenth [1169, 1170]. Fourth [2545, 2264, 944]. Fox [90, 2131]. fraction [1188, 227, 1618, 1160]. Frailty [1904, 1096, 1387, 1767, 1343, 2457, 676, 2410, 2159]. frame [1240]. framework [2222, 1867, 1741]. Framingham [1375, 180, 1377, 1328, 1376]. France [308, 1902, 2193, 49, 369]. Francia [665]. Francisco [1990]. Franco [510]. Frank [1876, 1014, 1970]. Franklin [347]. Fraud [1798, 2628, 1014]. Fraumei [1443]. Fred [2090, 731]. Frederick [1087]. Free [596, 864, 2278, 544, 1804, 1401, 1293, 163]. free-knot [1401]. free-living [163]. Freedman [2393, 2552]. freedom [373, 353]. French [2573, 1779]. Frequency [2093]. frequentism [803, 802]. Frequentist [1053]. frequently [1996]. Frison [2639]. FU [1055]. Fuerst [2638, 2165]. FUFA [1055]. full [625]. fully [530]. function [1417, 1056, 1410, 2593, 2524, 141, 1401, 685, 1645, 232, 2542]. functional [1505, 2533, 1563]. functions [397, 1377, 576, 109, 1735, 1022, 173, 887, 1833, 2415, 2307]. Fund [432]. Fundamentals [1165, 400, 2181]. Further [822, 62, 50, 63, 2439]. futility [231]. future [2518, 125, 1271, 1233, 124, 2272, 1145, 2023]. Fuzzy [1924].

G [23, 1961, 381, 21, 2312, 2039, 2042, 2369, 1898, 891, 110, 1888, 1862, 2378, 2648, 1824, 2091, 2003, 1924, 1037, 1823, 1015, 2585, 1860, 2645, 2633, 44, 382, 2002, 2333]. GA0 [596]. GA0/PEMD [596]. GA0/PEMD-92-18 [596]. Gäbler [1962]. Gabriel [404]. gain [105, 1906]. Gainesville [2038]. gains [344]. Gaithersburg [596]. Galen [2133]. Gambia [1698]. game [938]. gamma [1820, 676]. Gao [1193]. Gardner [831, 21]. Gary [89]. gastroenterology [20]. Gaussian [2433, 1650, 834, 1322]. Gaver [1359]. GEE [2341, 2203, 2508, 2450]. Gefeller [2128]. Geller [2393]. Genant [2638, 2165]. gender [1782]. General [816, 596, 370, 362, 1136, 2597, 1980, 1156, 2452, 1430, 1741, 390]. Generalization [238, 1578]. generalizations [1562]. Generalized [746, 316, 2386, 1869, 2529, 520, 757, 470, 2163, 533, 55, 1320, 588, 1492, 1335, 315, 1668, 2425, 1683, 2290, 2047, 2520, 1034, 2536, 2325, 332, 478, 2306, 838, 2119, 988, 1480, 1495, 528]. genes [2392, 995]. Genetic [2583, 1165, 480, 2616, 933, 2364, 1334, 1763, 2276]. genetics [24]. genital [952]. genomes [1703]. genotype [1917]. Geoffrey [2147, 366, 892, 2157, 944, 2461]. Geographic [1669, 1234, 1432, 857, 1254, 362]. geographical [530, 71, 1433, 573, 1868, 1732, 49, 2456, 1003]. Geographically [2370, 1668, 2046]. geography [362]. geometric [941]. Geometry [731]. Georg [1001]. George [788, 1544, 1809, 2126]. Gerald [1278]. Gerber [2418]. Gergen [2546]. Gerhard [1887]. German [1129, 2100]. gestational [2502, 549]. gestations [2502]. get [1123]. Gibbs [2321, 1636, 1905, 775]. Gilford [22]. Gilks [2169]. Gill [1108]. Gini [1815]. gis [1669, 1534]. given [608, 121]. GIVIO [2101]. Glantz [1153, 2545, 2264]. Glasbey [2018]. Glasgow [144]. Glattre [132]. Glenn [2275]. Glidden [1829]. Global [1579, 1488, 2088, 2363, 84, 2162, 1617]. Glynn [1455]. Goel [827, 1359]. gold [2638, 2165, 1510, 1932]. Goldberg [2248]. Goldstein [1878]. Gompertz [2120]. Good [1000, 2586]. Goodman [2646]. Goodness [2341, 1219, 1379, 580, 1839, 999, 2403, 927, 518]. Goodness-of-fit [2341, 1219, 1379, 580, 1839, 999, 2403, 927]. Gordon [667]. Göteberg [1371]. Goteti [1454]. Gower [90]. grades [637, 775]. graduates [115]. Graeme [2586]. graft [1903, 2361, 1400]. graft-versus-host [1903, 2361]. Graham [2018, 206, 207]. graph [1545, 2273]. Graphical [468, 1369, 1328, 1379, 942, 2473, 1941, 2172, 2053, 158, 2584]. Graphics [758, 2192, 860]. Graphing [1786, 2019]. Graphs [631, 1560, 2507]. Grayson [1798, 2397]. Great [1441]. greater [1287]. Greehouse [1362]. Green [1950, 1538]. Greenhouse [1359, 1838]. Greenland [485]. Greenwood [1972]. Gregg [2178]. Griffin [23]. Group [767, 2106, 2102, 2096, 2408, 2039, 1859, 2422, 2098, 2524, 173, 1804, 924, 2101, 2103, 1055, 2008, 278, 594, 2262, 1730, 311, 89, 588, 2372, 565, 574, 577, 567, 2556, 2198, 2452, 140, 2, 1221, 2049, 2348, 2329, 2322, 1678, 1062, 697, 2107, 2116]. group-randomized [2198]. Grouped [1927, 1172, 1807, 1012]. groups [2285, 781, 1149, 1730, 195, 2097, 2549, 1577, 1116, 2198, 705, 1152, 2025, 2548, 1041, 76, 2551]. Growth [1177, 2115, 1836, 445, 2084, 2302, 1009, 316, 2012, 2307, 1341, 87, 1145, 2068]. Gruet [2314]. Grundbegriffe [2056]. Grundmann [362]. Gruttola [2331]. Gruyter [2132]. Gudmund [2546]. guidance [1366]. Guide [1798, 2147, 2314, 1972, 1846, 366, 1004, 2418, 729, 1164, 830]. guidebook [2278, 2278]. guideline [1365, 2496]. Guidelines [1690, 2571, 1062, 786, 749, 2334, 1082, 21, 402]. Gullberg [2313]. Gustav [1826, 1962, 2056, 2180]. GUSTO [2283]. Guy [93]. Guyatt [667]. Gwendolyn [2150].

H [896, 1165, 1027, 2128, 2638, 2129, 132, 1887, 2009, 92, 759, 2179, 2495, 667, 1154, 841, 1924, 1847, 1037, 655, 112, 1948, 348, 1869, 1960, 1811, 1783, 652]. Haaland [2019]. habits [1522]. Haccou [1849]. Hadgu [2145, 2190]. haematologic [2200]. haematology [305]. haemophilia [70, 2060]. haemophiliacs [486, 929, 234]. Haenszel [487]. Haggett [362]. Hague [1439]. Hakama [770]. Hald [2462]. Hall [1985, 2253, 381, 2067, 94, 433, 2134, 110, 2169, 327, 2599, 2007, 1951, 668, 1038, 1860, 528, 2168, 44, 2381, 2069, 1314, 2382]. Halley [2547]. Halperin [97]. Hambleton [2286]. Hamilton [402]. Hampton [1877]. hamster [1381]. Hand [1951]. Handbook [2133, 786, 1701, 1811, 2286, 2179, 1888]. Handcock [2054]. handling [891, 636, 2090]. hands [435]. Hans [759, 2419]. Harbor [1443]. hard [345, 245]. hardback [2147, 366, 1786, 1936, 947, 246].

Hardcastle [2369]. hardcover [2277, 2379]. Harmonic [661]. harmonization [1310]. Harold [365, 246]. Harrington [613]. Harris [597]. Harry [2253, 2252, 2165]. Harvey [345, 1878, 2311]. Hashemi [2248]. Hassard [655]. Hastie [528]. having [774]. Hawaii [1533]. Hay [1164]. Hayduk [1938]. Haynes [667]. Hazard [2593, 1647, 277, 1946, 576, 1999, 2359, 744, 1410, 2137, 138, 509, 1916, 1904]. hazardous [199]. hazards [949, 412, 2242, 1648, 1264, 2566, 2271, 1735, 1410, 1022, 1369, 681, 42, 1482, 1785, 1928, 1353, 1564, 1801, 2075, 2427, 1749, 540]. HB [1372]. Hback [365]. Hbk [2147]. head [2317, 2084, 2607]. head-banging [2607]. Health [308, 2312, 2408, 1876, 2041, 2042, 2125, 2006, 1278, 1823, 1973, 1629, 2265, 1483, 1234, 875, 2370, 553, 1125, 1990, 2597, 1248, 680, 598, 413, 1662, 1669, 1377, 1977, 456, 111, 874, 100, 2341, 165, 328, 1557, 1235, 788, 1154, 585, 22, 1454, 1658, 376, 2046, 2423, 164, 2618, 2439, 2103, 1459, 2554, 2511, 257, 845, 347, 43, 1245, 1253, 893, 2532, 1021, 1337, 1496, 2613, 101, 1533, 1950, 759, 788, 2147, 90, 366]. health-are [257]. health-related [2103]. healthy [1864, 379]. Healy [2309]. Heaney [164]. hearing [360, 1992]. Heart [180, 1328, 1376, 1019, 35, 13, 427, 99, 141, 1246, 755, 1459, 2457, 40, 2345]. heart-failure [40]. Heath [1897]. heavy [591]. Hedges [1359, 1811]. Heidelberg [2179]. height [689, 2084]. Heinemann [614, 1973]. Helath [828]. held [59, 153, 388, 896, 1169, 1170, 1458]. help [1069]. helpful [1814]. Helping [1099]. Helsinki [132]. Hempfling [890]. Henry [480]. hepatitis [1698]. Herod [2081]. Herts [402]. Hesketh [2366]. heterogeneity [2343, 2270, 1817, 2121, 1767, 2453, 1945, 2017, 2563, 2031, 1944, 2410, 2333]. heterogeneous [1797, 2186, 626]. heteroscedasticity [1549]. Heterosexual [1335, 602, 310]. Hettmansperger [2494]. Hib [1699]. hidden [2630]. Hierarchical [1977, 103, 2351, 1852, 2469, 1792, 631, 2227, 2411, 928, 1415, 2404, 2500, 1114, 1575]. hierarchically [1890]. high [1990, 713, 1712, 140]. high-risk [1712]. Higher [2505, 1968]. Higher-moment [2505]. higher-order [1968]. highest [1322]. Hilger [148]. Hill [1902, 364, 1153, 2545, 511, 2264, 527, 739]. Hills [1194]. Hinde [2600]. Hirsch [893]. Historical [98, 1513, 1518, 100, 6, 291, 101]. History [2462, 2579, 2446, 694, 993, 1392, 2319, 143, 2530, 973, 2470]. HIV [2402, 1723, 1966, 1117, 1717, 2280, 875, 602, 1803, 1125, 1990, 1126, 1649, 2061, 1695, 486, 929, 792, 70, 2062, 203, 1335, 2044, 879, 1712, 993, 874, 2468, 1513, 951, 1914, 574, 2484, 1679, 507, 2327, 722, 1743, 1127, 1512, 2279, 2139, 1998, 1473, 1694, 358, 1133, 233, 1389, 1166, 1130, 1141, 1720, 330, 601, 1128, 2060, 1137, 1991, 171, 310, 1484, 48, 2592, 76, 1350, 1031, 2281, 2184, 663]. HIV-1 [2402, 602, 879, 663]. HIV-infected [1389]. HIV-positive [486, 929]. HIV-related [1803]. HIV/AIDS [1125, 1990, 2044, 2484]. HMSO [362]. Hobart [1786, 1338]. hoc [1841]. Hoch [2043]. Hodgkin [795, 2045]. Hoel [1015]. Hoffmeister [759]. Holden [2345]. Holland [553, 2379, 788]. Holme [93]. home [1377]. homo [574]. homo/bisexual [574]. Homogeneity [337, 1045, 336, 438, 2409, 1920]. homogeneous [1511]. homogenetic [1471]. homoscedastic [1396]. homosexual [1695, 1389]. honour [2622]. hope [2075]. Hopkins [1938, 2394, 2379, 1085, 24, 1937]. Horgan [2018]. hormone [224, 643]. Horn [2180]. Horwood [870]. Hosmer [326, 165, 1872]. Hospital [450, 1991, 2257, 2617, 1236, 1018, 1007, 1890, 2274]. hospital-specific [1007]. Hospitals [553]. host [1903, 2361]. hot [1754]. hot-deck [1754]. Hotel [308]. Hothorn [555]. hour [834, 621, 1191, 644]. house [92]. household [1545]. Howe [2378]. Howell [1026]. Howson [325]. http [2278]. Huet [2314]. Hui [2165, 2143]. Hul [2638]. Human [1898, 1950, 1599, 2402, 1669, 2058, 204, 1505, 1815, 743, 24, 1881, 1118, 214, 454, 2360, 1825]. humans [2342]. humongous [251]. Huw [1085]. Hwang [1874]. hybrid [1344]. hyperkinetic [15]. Hypertension [839, 1358, 2320, 1806, 15, 2078]. hypertensive [263]. hypertensives [1092]. hypertext [1913]. hypospadias [324]. hypotheses [447, 244]. Hypothesis [4, 1241, 174, 1188, 918, 1543, 2645].

I.C.O.H [308]. Iain [2408]. IARC [451, 1193, 1813]. IBCSG [2102]. IBM [786]. IBM-PC [786]. Iceland [132]. ICG [767]. ICG-CTR [767]. ICH [2496]. ideas [1136, 2334, 812]. Identification [643, 1825, 224, 290, 1246, 1460]. identify [2257, 2162, 2456]. Identifying [2114, 214, 495]. Ignaz [2022]. Ignorability [2535]. ignorable [2597, 2015, 2536, 1755, 1756, 1759]. ignoring [2175]. II [2473, 1178, 1678, 102, 305, 723, 2014, 2267, 1101, 514, 2444, 1925, 350, 752, 2172, 7, 1346, 6, 780, 2575, 2376, 1722, 2632]. II/III [7, 1722]. III [1087, 1947, 1283, 1366, 1299, 7, 1722]. III/3630/92 [1366]. III/3630/92-EN [1366]. IIIrd [308]. Illinois [325]. illness [1111, 921]. illnesses [1711]. Illusion [814]. illustrated [261, 1264, 1328, 275]. illustration [2304, 2497]. illustrations [993, 394]. IMA [2276]. Image [2009, 2018]. Images [2192, 394]. imagination [1486]. Imaging [2009, 1505]. imbalance [659, 175]. immediate [567]. immunes [1187]. immunoassay [1856, 1920]. immunodeficiency [2402, 2058, 743, 1118]. Immunointervention [841]. immunologic [1723]. immunological [1998]. Impact [605, 2175, 89, 2484, 1941, 2516, 2, 1527, 676]. imperative [1752]. imperfect [142]. implement [1460]. Implications [494, 1233, 968, 603, 2639, 258, 1389, 2554, 932]. Importance [638, 943, 456, 64, 906, 1051]. important [1436, 2553]. impossible [1752]. imprecise [1548, 523]. improve [212]. Improved [679, 1841, 2295, 1632, 157, 1131, 2633]. improvements [1317, 2443]. improving [1450, 2207]. imputation [2015, 823, 679, 1510, 1131, 1391, 586, 2596, 1754, 257, 1758, 48, 2388]. imputed [2492]. in-hospital [1890, 2274]. in-patient [561]. inch [1901]. Incidence [1803, 1193, 1439, 1117, 1186, 1124, 1990, 1126, 979, 2061, 908, 1529, 2503, 2558, 1380, 132, 792, 2371, 1032, 983, 1551, 1976, 684, 648, 2471, 1104, 760, 1513, 502, 1078, 1521, 682, 1370, 1443, 506, 1509, 950, 1833, 1562, 2173, 2325, 1162, 1116, 1127, 1512, 2139, 1533, 2411, 1133, 2352, 1441, 601, 1128, 2060, 2432, 1882, 2456, 663]. incidences [1126, 82]. including [1966, 2247, 1476, 1709, 1625, 1092, 1626]. Incomplete [825, 2111, 628, 2110, 975, 1979, 1844, 2371, 1730, 648, 534, 2117, 993, 1340, 625, 1770, 544, 913, 1479, 201, 1757, 2470, 1637]. incompletely [1331]. inconclusive [1978, 2354]. inconsistency [1264]. incontinence [2581]. Incorporated [347]. Incorporating [1817, 1130, 6, 2233]. Incorporation [2319, 2033]. increase [888, 2016, 663]. increased [1542, 230, 1533]. increases [1139]. incubation [1990, 2421, 1139, 485, 1726, 1484, 1012]. independence [373, 1363, 1682, 837, 1917, 715]. independent [1603, 415, 1064, 1948, 791, 1266, 2123, 1232, 274]. Index [1872, 2084, 1091, 1560, 1413, 1607, 776, 1815, 850]. Indexes [2549, 2551]. indicators [314, 1377]. indices [2118, 1329, 1895]. indirect [2510]. Individual [301, 1874, 1821, 781, 548, 1216, 2293, 2189, 1616, 772, 1152, 2322, 2162, 1489]. individual-level [548]. individuals [1907, 2326, 1512, 1480, 1031, 11]. induced [1635, 303, 218, 2325, 255]. induction [1131]. industrial [49]. industrie [1628]. industry [711, 701, 1744, 1935]. inequalities [90]. infancy [475]. Infant [2403, 1234, 687, 108, 392, 1664]. Infants [2125, 445, 2083, 1466]. infarction [760, 524]. infected [2061, 1512, 1031, 1389]. infection [875, 1649, 1695, 792, 70, 939, 952, 203, 106, 1712, 2323, 484, 507, 2327, 722, 1743, 142, 2352, 1618, 2060, 76, 1350]. infections [1836, 2271, 1545]. Infectious [2394, 1898, 699, 375, 677, 1082, 44]. infectiousness [2062, 2153, 2327, 1720]. infectivity [1990, 203]. Inference [1933, 2380, 1321, 2063, 580, 1854, 1999, 2451, 2435, 1570, 2530, 1727, 999, 1269, 1302, 1833, 1694, 489, 762, 348, 527, 2541, 969, 1192, 2181, 2149, 243]. inferences [1099, 1852]. Inferring [1293]. Influence [793, 1296, 314, 1892, 2616, 178, 591, 454]. influences [1124]. influencing [215, 998, 642]. Informatics [2179]. Information [875, 1359, 1696, 1234, 2477, 319, 2302, 1669, 1129, 1308, 1510, 713, 2120, 2269, 867, 204, 735, 258, 2049, 1237, 2643, 2245, 1010, 1083, 758]. information-based [2049]. informative [2435, 489, 2434]. informatively [622]. Inherent [905, 745]. inhibitor [234]. initial [781, 1152, 1855, 989]. initiated [2527]. injection [1389]. injuries [545]. injury [2317, 1394, 1554, 2619]. Innovations [2540]. Innovative [260]. inquiry [347]. insemination [2154]. INSERM [1902]. Institute [288, 896, 828, 2038, 1898, 2495, 1154, 2103, 1055, 699, 98, 99]. Institutes [101, 100]. institutional [2227]. institutionalization [1377]. institutions [2426]. instruction [452]. Instructor [2296]. Instructors [2296]. Instrumental [2569, 1046, 761]. insulin [2387]. intake [984]. intakes [475, 925]. integer [2503]. integer-valued [2503]. intended [2466]. intensity [1791, 1350]. intensive [244]. Intentionally [625]. inter [319, 2155, 2335]. inter-block [319]. inter-event [2335]. inter-observer [2155]. interaction [602, 2320, 763, 1667, 1434, 374, 1909, 1782]. interactions [2244, 1735, 1022, 1942, 782]. interactive [1454]. interclass [2452]. Interdisciplinary [2101]. interest [1073]. Interfacing [860]. interferon [1148]. Interim [298, 706, 2442, 1056, 423, 2506, 2218, 300, 652, 1925, 127, 701, 1939, 1063, 1285, 1497, 702, 1057]. interlaboratory [292]. intermediate [1576, 439]. Internal [1175, 2631, 2632, 2433, 2463, 1624, 2429]. International [59, 153, 308, 388, 288, 896, 1169, 1170, 1458, 2102, 1026, 451, 20, 400, 2045, 2193, 2496, 1310, 59, 153, 388, 767, 896, 1169, 1170, 1458]. Internet [1913]. interobserver [425, 1852]. Interpretation [1915, 374, 893, 61, 638, 1017, 256, 943, 1054, 2073, 944, 1704]. Interpreting [2387, 2363, 1824]. interrater [86, 129]. Interval [2424, 1492, 2435, 1466, 2123, 1105, 1872, 2595, 2594, 1695, 2614, 580, 823, 2350, 2244, 2566, 1413, 999, 1907, 2064, 2188, 2504, 604, 2454, 1291, 1789, 794, 1604, 1790, 2124, 1880, 2505]. Interval-censored [2435, 2064, 1789, 1604, 1790]. intervals [520, 757, 21, 2089, 2611, 2332, 1821, 1609, 1144, 236, 1097, 1718, 157, 1820, 1188, 2648, 1482, 1785, 2499, 2237, 1959, 235, 1047, 2295, 2122, 273, 2335, 1341, 447, 1619, 1388, 2633, 2635, 741, 1181, 1161, 1024, 2024, 2375, 2376, 1819]. Intervention [231, 1870, 25, 2246, 2446, 673, 105, 1570, 672, 1, 1919, 964, 683, 197, 605, 1319, 2250, 76]. interventions [2284, 1347, 1139, 676]. interventlons [1037]. interview [1248]. intra [781, 1633, 1152, 1920]. intra-individual [781, 1152]. intra-run [1920]. intra-subject [1633]. intraclass [590, 442, 1368, 1893, 1176, 1822, 2505]. intraepithelial [952]. intrapartum [2425]. intravenous [602]. Introducing [730]. Introduction [2604, 1983, 2041, 956, 2081, 843, 1708, 2572, 2297, 1971, 1753, 616, 2395, 1655, 2239, 1703, 510, 2163, 1610, 1551, 418, 347, 2404, 527, 1684, 1985, 1982, 245, 511, 2544, 2053, 1846]. introductory [2496, 2396, 148, 1877]. Intuitive [2311]. invariant [2322]. inverse [753, 1404, 1528, 1921]. investigate [1731, 1287, 108, 255]. Investigating [1533, 1346, 2017, 2498, 1331, 2193, 2333]. Investigation [2339, 2614, 2465, 63, 751, 752]. investigations [1241, 2209, 845, 242]. investigators [91]. involved [935]. involvement [1690]. involving [515, 952, 604, 1895, 1285, 1497, 1179]. iodine [931]. IOS [1862]. Iowa [164]. irregularly [1638]. irreversible [1162]. Isabel [1370]. ISBN [1961, 1901, 614, 2584, 434, 656, 770, 758, 1087, 827, 1812, 2253, 510, 945, 597, 1938, 1934, 2252, 2310, 2462, 553, 1798, 2397, 1826, 828, 2394, 1165, 1984, 1027, 2090, 2147, 2067, 2368, 1048, 1359, 2276, 1922, 2277, 452, 2367, 840, 2312, 2055, 598, 403, 1983, 482, 480, 2408, 1902, 2092, 1876, 362, 613, 2131, 2314, 401, 871, 2602, 1887, 1014, 2039, 1026, 2009, 731, 1962, 2204, 2275, 2254, 2018, 364, 1153, 2148, 786, 402, 1859, 1799, 2041, 892, 1786, 2070, 1936, 554, 2545, 1911, 2396, 450, 2042, 759, 2369, 1878, 2379, 2038, 2134, 872, 2263]. ISBN [1898, 2286, 1192, 2081, 2157, 1085, 1899, 2006, 2179, 1193, 1000, 1994, 1888, 1970, 1862, 933, 511, 2495, 1086, 1194, 944, 2573, 2378, 404, 2345, 829, 1923, 1004, 2169, 788, 947, 1154, 2586, 2278, 1824, 1370, 1371, 1455, 1372, 1454, 1443, 2601, 2091, 1848, 1896, 2068, 841, 2419, 2599, 2007, 2194, 2132, 1972, 2461, 2572, 570, 2544, 2546, 2313, 2082, 1924, 612, 1951, 2040, 2417, 2418, 1847, 1933, 1037, 1278, 655, 946, 2380, 1849, 481, 1813, 1823, 1950, 2587, 2133, 2494, 1897, 2439, 1002, 2311, 668, 870, 2571, 1787, 1038, 2585, 526, 2150, 2093, 1982, 1971, 921]. ISBN [2043, 1166, 1860, 1900, 2264, 1838, 1850, 1338, 1877, 2600, 2547, 1861, 2053, 2366, 528, 1108, 2383, 2170, 2384, 2181, 2168, 1981, 787, 2019, 1003, 555, 2395, 2192, 2149, 2239, 1937, 527, 1277, 382, 1825, 2381, 2069, 1912, 830, 1025, 2056, 2054, 1846, 1788, 1315, 1314, 1960, 2020, 2180, 893, 1001, 2193, 920, 2008, 1164, 2265, 1811, 2178, 2382, 1935, 2167, 667]. ischaemic [1459]. ISE [365]. Isham [1898]. Islands [1533]. ISO [2009]. isobolograms [1156]. isotonic [2465, 865]. Israel [1931]. ISS [1554]. issue [2356]. Issues [1938, 1359, 1248, 2210, 2263, 2226, 2532, 2280, 2525, 717, 2223, 2106, 2102, 2302, 2519, 2208, 2097, 709, 1504, 169, 1263, 1724, 1940, 1199, 685, 1255, 1310, 1586, 67, 808, 1141, 1744, 1757, 1166]. Italian [486, 929, 1743]. Italy [602, 2101]. Item [2580, 2286, 1970]. items [1449, 2110]. Iterative [2390, 1891, 1880]. IV [402, 1813, 1962, 555]. Iversen [2546]. Ivo [1887]. ix [656, 510, 945, 2397, 244, 2314, 401, 384, 206, 145, 2573, 242, 1950, 668, 1787, 207, 2093, 2366, 44, 382].

J [434, 432, 770, 945, 1934, 345, 1165, 21, 2368, 1922, 90, 730, 1871, 401, 2393, 433, 92, 786, 2070, 2038, 891, 2286, 145, 2081, 2157, 2179, 1193, 1994, 2495, 944, 2573, 404, 829, 2169, 947, 1154, 327, 1568, 2639, 1454, 1443, 2007, 2194, 2003, 2417, 112, 243, 1813, 164, 285, 1787, 1038, 2585, 2093, 2005, 528, 2170, 2168, 348, 880, 1947, 787, 1003, 2429, 1869, 43, 2149, 2004, 1580, 2143, 2056, 920, 2008, 184, 2002, 245, 2333, 1935]. J-P [404]. jack [2371]. jack-knife [2371]. Jackknife [1794]. Jacques [2277, 2644]. James [144, 759, 1899, 944, 1847]. Jan [2313, 2019]. Jan-Aage [2019]. Janelle [165]. Japanese [2325]. Jean [147, 93]. Jeffrey [2043, 2054]. Jen [872]. Jen-Pei [872]. Jena [1826, 2180]. Jensen [132, 2601, 2168]. Jersey [1934, 1786, 1004, 1338, 347]. Jessica [2040]. Jewell [1166, 205]. Jillian [1897]. Jim [2396]. Jimma [2403]. job [575]. job-exposure [575]. Joe [2253]. Joel [1359, 1838]. John [1087, 827, 2506, 363, 1359, 482, 1902, 2092, 2131, 364, 346, 147, 1372, 870, 921, 1937]. Johns [1938, 2379, 24]. Johnson [2332, 2130, 2204, 2237, 2635, 255]. Joint [2437, 963, 1769, 178, 1156]. jointly [402]. Jolivet [2314]. Jonathon [786]. Jones [110, 2133, 1947]. Joness [1454]. Jørn [510]. Jorner [2019]. Josef [2586]. Joseph [2494]. Jost [2132]. Journal [21, 401, 1014]. journals [410, 2299, 1218]. Jr [656, 165, 1443, 1897]. judgements [1646]. Judith [1455]. Jukka [2004]. July [1458]. Jun [1359]. Junshi [362]. Junyao [362].

Kaldor [2002]. Kaplan [734, 2136]. Kaposi [2151]. Kappa [279, 2118, 361, 1219, 580, 999, 2504, 341, 2087, 886, 1190, 1606]. Kappa-like [2118]. kappa-type [2087]. Karel [510]. Karen [2506]. Karger [2055, 1000]. Karl [146, 871]. Karvonen [2005, 2004]. Kasovia [1454]. Kasovia-Schmitt [1454]. kcal [2582]. Keiding [1108]. Keirse [434]. Keith [841]. Kelley [656]. Kendall [2168, 2595, 1933]. Kenneth [243]. Kent [1026, 400]. Kenward [110]. kernel [392]. key [1420, 2414]. keyed [819]. Keying [2165]. Keynote [1232]. Khan [246]. Khoury [1165]. kidney [392, 1400]. Kinetics [931, 19]. Kingdom [2019]. Klar [165, 2250]. Klaus [1166]. Klein [827, 2092]. Kleinbaum [2042]. Kluwer [827, 933, 1848, 1825]. Knapp [1896]. knee [1660]. knife [2371]. knot [1401]. know [920]. knowledge [1469]. Knox [788]. Knudson [1562]. Koch [2552]. Kolman [2586]. Kotz [2204]. Koul [2495]. Kr [1371]. Kragh [1108]. Kramar [1902, 364]. Kramer [91]. Kreienbrock [1826]. Krishnamurty [1454]. Krishnan [2157]. Krzanowski [1922]. Kshirsagar [2068]. Kuhn [2009]. kurtosis [458]. Kyle [1315].

L [2584, 2147, 2332, 2312, 366, 2130, 2092, 2393, 2009, 731, 2204, 2347, 2369, 2248, 1193, 2495, 667, 1154, 2639, 2007, 2082, 2237, 1813, 1823, 2066, 2585, 1850, 2168, 555, 2429, 2637, 1937, 2635, 1811]. labelled [2468, 163]. labor [1150]. laboratories [656]. Laboratory [1443, 614, 656, 1702, 966, 402, 2625, 563, 1539, 2492, 1705, 1704, 291]. lack [1777, 2497]. lactation [753]. lags [1979]. Laird [2403]. Lance [1455]. Lancet [1455]. Landmark [731]. Lang [2571]. Lange [1838, 382]. Large [1999, 1647, 61, 2590, 1806, 2256, 456, 958, 917, 1235, 544, 1577, 2306, 263, 834, 1237, 2500]. large-sample [544]. larger [1104]. Larry [1359]. Lars [570, 1001]. larynx [93]. Laska [554]. lasso [1778]. last [591]. late [1334]. late-onset [1334]. Latency [2602, 2566, 1132, 1082]. Latent [2358, 53, 2118, 378, 1975, 641, 2078]. Latin [479, 913]. Lauritzen [2584]. Law [2002, 41, 183, 429]. lay [1072]. layouts [299]. LD50 [1272]. lead [505, 972, 1465, 2336]. lead-time [2336]. Leading [2204]. learning [1499, 505, 2283, 2042, 1896]. least [1792, 2449]. Leaverton [452]. Lecture [384, 404]. Lee [2393, 2572, 1002, 1846]. Leeds [432]. Lefevre [404]. legal [730, 1364]. Lehtonen [1787]. Leiden [1154]. Lemeshow [326, 165, 1872]. length [2340, 427, 2257, 2617, 233, 2615]. lengths [392]. lens [1480]. Lesley [1798, 2397]. Leslie [1938, 944]. less [2259, 1123]. Letter [2506, 2649, 2641, 2128, 666, 2343, 2089, 2638, 2332, 2636, 1476, 2130, 2644, 2145, 2393, 2165, 283, 89, 2065, 2331, 2248, 2642, 891, 1442, 611, 569, 1568, 2639, 2648, 143, 2473, 2647, 1809, 2237, 2164, 1948, 2066, 890, 2474, 2190, 2645, 880, 1947, 1874, 2646, 2429, 2637, 2640, 2004, 2141, 2633, 2635, 2643, 1580, 2250, 2143, 1822, 2002, 1626, 2126, 1872, 2333, 2354]. Letters [757, 826, 654, 552, 1625, 593, 582]. Leukaemia [432, 370, 1931, 1534, 1562, 1441, 2045, 2228, 1439, 1184, 2456, 432]. leukocyte [2438]. leukocyte-reduced [2438]. level [1746, 2163, 2154, 2372, 548, 1216, 545, 2046, 1242, 2012, 2307, 688, 2500]. levels [72, 505, 312, 2413, 2245]. Leventhal [145]. Lewis [2506, 1947]. LFP [37]. Li [362]. Libbey [147]. Library [1798, 2397]. licensing [2206, 2527]. licensure [2224]. Lidocaine [231]. Life [2587, 1205, 1163, 1746, 2095, 2106, 2102, 2096, 2111, 2114, 362, 1864, 2580, 915, 1863, 2117, 2109, 2097, 2110, 2104, 1567, 2160, 2098, 2105, 1454, 2099, 2112, 2107, 2101, 78, 2103, 199, 411, 2113, 736, 84, 2116, 2108, 2115, 104]. life-style [362]. lifestyle [828]. lifetime [1379, 2359, 2501, 2617]. lifetimes [1382]. ligand [19]. ligand-binding [19]. like [2118]. likelihood [2186, 520, 757, 205, 2128, 2594, 475, 1840, 2129, 1760, 560, 2084, 1551, 733, 1792, 1524, 1818, 181, 1605, 964, 2139, 2140, 304, 591, 838, 2542, 1754, 1190, 1606, 899, 51, 1173, 274]. likelihood-based [964]. likelihoods [2138, 2460]. Limburg [2072]. Limitations [1805, 261]. Limited [2194, 2412]. limiting [2305, 41, 183, 429]. limits [1396, 2127, 727, 1809, 1345, 1404, 2126]. Lin [2331, 1835]. Linden [2286]. Lindsey [2396, 2093, 2149]. Linear [2596, 1869, 1448, 822, 1635, 2529, 520, 757, 975, 425, 2163, 55, 1980, 2433, 1320, 628, 1396, 1976, 588, 1492, 1035, 315, 1402, 2201, 2290, 297, 2047, 2520, 176, 2536, 2538, 332, 773, 2564, 2415, 1992, 478, 2306, 868, 2386, 2048, 1268, 1777, 2539, 1495, 2582, 2404, 621, 1191, 2176, 1916, 627, 1322, 1847]. linear-logistic [1916]. Lineare [2419]. linearity [271, 1146]. Linearly [2029]. linkage [2010, 1235, 1236, 54, 1250]. linkages [1237]. linked [233, 76]. Linking [1234, 2224]. lipids [2387]. Lippert [759]. LISREL [2395, 1938]. List [250, 799, 671, 617, 1281, 2278]. lists [201]. Literature [1798, 1275, 1516, 2365, 2316, 2626]. Little [452, 667]. Liu [872]. liver [1180, 275]. lives [1857]. living [163]. Llewelyn [1085]. Lloyd [1278]. LMS [560]. Local [1989, 1658, 2601]. Localization [1550, 2390]. localized [2193]. Location [1320, 2597, 988, 1805]. Location-scale [1320]. locations [1548]. Lock [1014]. Log [425, 109, 1270, 517, 2048, 518, 2567, 1819, 2261]. Log-linear [425, 2048]. log-normal [517, 518, 2567, 1819, 2261]. Logistic [1499, 1988, 1266, 1761, 1331, 1872, 2321, 2502, 1107, 2156, 1760, 1794, 35, 326, 180, 746, 1915, 766, 1095, 1413, 1839, 2319, 2201, 42, 1033, 1412, 333, 1674, 2028, 128, 662, 928, 541, 774, 1693, 1606, 2413, 1332, 1916, 274]. logistical [701]. logit [864, 108]. Logrank [1691, 1797, 277, 538, 440, 2185, 564, 509, 1807]. Lon [933]. London [1985, 831, 381, 363, 553, 1798, 2397, 828, 21, 2277, 2408, 362, 401, 1014, 2039, 433, 92, 402, 450, 1878, 147, 93, 1085, 829, 1004, 2169, 327, 1455, 1372, 2599, 2007, 242, 1951, 1933, 2133, 2494, 668, 1038, 2600, 2366, 528, 2168, 2019, 2239, 44, 2069, 1314, 2008, 1164, 2382]. London/New [2494]. Long [2148, 2194, 77, 1715, 1889, 198, 524, 1749]. Long-term [2148, 1715, 1889, 198, 1749]. Longitudinal [2472, 689, 1417, 2403, 1951, 2020, 1322, 2321, 1323, 2288, 1989, 723, 1773, 2558, 1980, 1177, 1313, 2580, 939, 2445, 759, 1340, 891, 883, 1775, 625, 636, 1770, 829, 1472, 1544, 2522, 1140, 1710, 570, 39, 1573, 1771, 2066, 1581, 1146, 2358, 2012, 1930, 2581, 526, 747, 626, 1341, 619, 1042, 1105, 139, 622, 1420, 2078, 241, 794, 2272, 1774, 2231, 2434, 1001, 627, 1213, 1321, 620, 989]. longitudinally [233]. longitudinally-linked [233]. look [2561]. Lopez [1897]. Lorenz [2056, 1815, 2363]. Lorenzo [451]. loss [995, 2183, 540]. losses [1941, 736]. lost [199]. Lothar [1826]. Louis [655, 2382]. Louise [1838]. Loveday [1838]. low [2590, 445, 746, 2432]. low-birth-weight [445]. low-dose [746]. low-incidence [2432]. lower [344, 2177, 2474]. Ltd [614, 2368, 1455]. Lu [2638, 2165]. Luigi [2380]. lung [1053, 99, 1521, 1522, 1065, 49, 1239, 2240]. Lwanga [165]. lymphocyte [2342, 1031, 743, 1118]. lymphocytic [1562]. lymphoma [432, 795]. Lynch [480]. Lynne [1838]. Lyon [451, 1193, 1813, 2193].

M [1961, 434, 770, 1812, 2506, 2649, 2252, 1984, 2641, 2368, 1048, 1359, 2089, 2636, 362, 2009, 2065, 554, 2369, 2642, 891, 93, 2006, 2179, 1193, 1888, 1862, 829, 22, 2068, 2461, 2572, 2003, 2647, 1924, 2040, 2164, 481, 2133, 12, 2005, 1850, 1874, 1869, 2640, 527, 2004, 2008, 2002]. M-mode [12]. Maastricht [59, 153]. Machin [1812, 2368, 285]. Macintosh [786]. Macmillan [92, 111, 93]. MACS [663]. Madansky [383]. Made [2296, 89, 2]. magnetic [1505, 2533]. magnitude [2294]. Magnusson [570, 1001]. mailback [679]. Maillot [308]. make [2385, 2076]. makers [91]. making [421, 2618, 1415, 112]. malaria [1160]. male [1137]. males [2605]. malignant [2046]. Malker [132]. Maller [2148]. mammary [984]. man [841]. management [2506, 1913, 2218, 681, 1055, 1277, 2383]. managers [1701]. Mancia [1862]. mandated [1588]. Manfred [2180]. Manhattan [828]. Manly [433, 947]. Mann [2416, 2448]. Mann-Whitney [2416]. Manning [759, 2645]. Mantel [2622, 2624, 2625, 487, 2623]. Manton [23, 1924, 1771]. many [457]. map [1432, 1656, 860]. mapped [790]. Mapping [2276, 1254, 1438, 1816, 1433, 362, 856, 1505, 783, 1243, 2583, 859, 1552, 2240]. maps [1436, 2609, 858, 2607, 1244, 1703]. Marc [945]. Marcel [945, 597, 146, 871, 1911, 872, 1192, 1970, 2181, 166, 1935]. Margetts [1048]. Marginal [1834, 2521, 2048, 1399, 963, 1987, 1867, 1147, 2086, 1605, 964, 734, 886]. margins [306]. Marie [2314]. Marie-Anne [2314]. marital [230]. Marjorie [2265]. Mark [2054, 184, 584]. mark-recapture [584]. marked [1748]. Marker [877, 744, 1134, 1646, 939, 2331, 1140, 1894]. Marker-dependent [744]. Markers [876, 1943, 1723, 1067, 2289, 1066, 1453, 879, 1418, 1825]. market [2224]. marketing [1366, 907]. Markov [1966, 2186, 878, 1351, 993, 2268, 725, 2630, 507, 2327, 2470, 2169]. Markovian [82]. Marks [2252]. marriage [155]. marrow [1903, 2271, 901, 2342, 2361]. Martin [831, 21, 1951, 1971]. Marubini [1961]. Maruyama [2461]. Mary [2546]. Maryland [1937]. Maryon [1626]. masking [2370]. Mass [243, 2084]. Massachusetts [2275, 1848, 413]. Masthead [3, 26, 45, 56, 74, 96, 113, 133, 151, 167, 187, 208, 220, 248, 268, 286, 309, 329, 349, 367, 386, 405, 417, 436, 453, 461, 483, 496, 512, 529, 542, 556, 572, 583, 600, 615, 634, 646, 657, 669, 690, 721, 732, 740, 748, 778, 789, 797, 818, 832, 842, 863, 873, 881, 894, 910, 923, 934, 948, 954, 991, 1006, 1016, 1029, 1039, 1049, 1076, 1089, 1098, 1109, 1120, 1143, 1155, 1167, 1182, 1195, 1207, 1217, 1226, 1261, 1279, 1316, 1327, 1339, 1348, 1360, 1373, 1385, 1395, 1405, 1414, 1426].

Masthead [1444, 1456, 1467, 1477, 1487, 1501, 1508, 1520, 1530, 1558, 1569, 1582, 1598, 1614, 1631, 1642, 1652, 1663, 1673, 1689, 1706, 1719, 1728, 1739]. matched [477, 228, 370, 415, 1352, 336, 980, 2202, 550, 2188, 608, 729, 488, 1190]. matches [1347]. matching [105, 760, 2468, 1919, 2189, 683, 177, 2250]. materially [423]. Maternal [1906, 1380, 1914]. maternal-child [1914]. Math [1876]. Mathematical [2462, 494, 2495, 384]. Mathematics [23, 2081, 2313]. Mathur [2638, 2165]. matrices [575]. matrix [2301, 1340, 2087]. matter [2272]. matters [301]. Matthews [1934, 2055, 92]. maturation [1918]. Mauricio [2150]. Max [554, 97]. maxima [847]. maximin [2482]. Maximum [475, 181, 2140, 304, 1190, 1606, 51, 205, 2594, 1760, 2084, 733, 838, 2542, 715, 274]. maximum-likelihood [733]. Maxwell [365]. May [2332, 2130, 2237, 2635, 1782, 2309]. Maynard [2408]. McCloskey [2278]. McConkey [92]. McCormick [144]. McFadden [2383]. Mcgilvray [944]. McGraw [1153, 2545, 2264]. McGraw-Hill [1153, 2545]. McKean [2494]. McKinney [432]. McLachlan [2157]. MCMC [1440]. McNeil [1981]. McNemar [415, 581, 2258, 550, 1383]. Md [596]. Mean [330, 1203, 1946, 1351, 293, 427, 603, 109, 499, 2524, 340, 2639, 768, 932, 1036, 2459, 1819, 2261, 2273]. means [2330, 865, 1181, 1779]. measles [675, 573, 106, 485]. Measure [890, 2357, 559, 2636, 1111, 1864, 2524, 687, 1578]. measured [2242, 1636, 65, 2576, 1139, 1765, 2484, 1493, 312]. Measurement [2147, 1046, 1517, 751, 752, 1493, 2574, 2575, 1746, 366, 55, 781, 27, 1218, 1449, 1965, 2175, 644, 952, 2577, 206, 1033, 2423, 605, 2516, 2358, 523, 207, 2255, 837, 1766, 1268, 774, 1152, 1612, 163, 2231, 1830, 2291, 2292, 104, 2007, 1704]. measurements [949, 1676, 414, 1730, 65, 1567, 499, 1008, 2425, 2304, 2086, 176, 237, 490, 2464, 838, 1632, 1341, 158, 768, 641, 1774, 1478, 2233, 1796]. Measures [987, 47, 2472, 1265, 217, 533, 352, 1980, 2294, 1863, 2445, 603, 2029, 1397, 2160, 1093, 885, 16, 1769, 2387, 2428, 2639, 1804, 2363, 1043, 2538, 1246, 1574, 1905, 650, 1550, 2390, 2307, 2452, 2591, 2203, 458, 139, 2348, 932, 2414, 1094, 202, 2322, 1114, 426]. Measuring [1567, 64, 677, 129, 1452, 1504, 2327, 747, 1600]. mechanism [985]. mechanisms [1462]. Médecine [1902]. media [179, 8]. Median [929, 1092, 1212, 1447, 1482, 1785, 342, 2292]. medians [2501]. Medical [1934, 21, 401, 1936, 2179, 1370, 148, 2571, 704, 1971, 1860, 1314, 920, 410, 1275, 1516, 2299, 2365, 115, 1518, 381, 2488, 596, 470, 819, 2055, 1099, 118, 1218, 1014, 125, 283, 944, 1455, 252, 242, 2334, 1867, 355, 1367, 2047, 1928, 935, 121, 2528, 258, 2266, 1205, 1268, 284, 1237, 2567, 1148, 1087, 2067, 401, 2009, 285]. Medicare [261, 252]. medicinal [1366]. Medicine [23, 1994, 385, 409, 1748, 1702, 1627, 892, 20, 667, 665, 1705, 347, 1684, 184, 1704, 2506, 2649, 2641, 2128, 2343, 2089, 2638, 2129, 2332, 2636, 1476, 1871, 2130, 2644, 2145, 2393, 2165, 2127, 283, 89, 2065, 2331, 2248, 2642, 891, 1870, 1568, 2639, 2648, 2473, 2003, 1625, 2647, 1809, 2237, 2164, 1948, 2066, 890, 2571, 2005, 2474, 2190, 2645, 880, 1947, 1874, 2646, 2429, 1869, 2637, 2640, 2004, 2141, 2633, 2635, 2643, 1580, 2250, 2143, 1822, 2002, 1626, 2126, 1872, 2333, 2354, 144, 2559, 2560, 614, 787, 912]. medicines [1199, 263]. medium [574]. Medley [1898]. MEDSNAP. [1973]. Meelis [1849]. Meeting [59, 153, 388, 288, 896, 1169, 1170, 1458]. Meier [734, 2136]. Meinert [1937, 2347]. melanoma [2605, 2046, 1271]. Melvin [2092]. Melvyn [2378]. men [1695, 574, 1389]. Meng [2586]. Meningitis [2125]. menopausal [1211]. menstrual [217, 754, 685, 1386, 214]. mental [2597, 680, 445, 2341]. Menten [1683]. Meridien [308]. merits [1919, 2250]. Merletti [510]. Mervi [2020]. Meta [2529, 2287, 437, 1943, 1627, 1584, 1249, 1590, 984, 1200, 2353, 1232, 1586, 2637, 1239, 2626, 2063, 1842, 1594, 1222, 1515, 2343, 2270, 1817, 2320, 638, 2515, 1046, 303, 1588, 1378, 2393, 2469, 1596, 1524, 2121, 1747, 1927, 2528, 2316, 2386, 2229, 943, 1463, 1238, 1406, 1954, 2017, 2563, 2033, 2513, 898, 2031, 390, 1179, 2032, 2333]. meta-analyses [1046, 303, 1596, 2528, 2316, 1406]. Meta-analysis [2529, 2287, 437, 1943, 1627, 1584, 1249, 1590, 984, 1200, 2353, 1232, 1586, 2637, 2626, 2063, 1842, 1594, 1222, 1515, 2343, 2270, 1817, 2320, 638, 2515, 1588, 1378, 2393, 2469, 1524, 2121, 1747, 1927, 2386, 2229, 943, 1463, 1238, 1954, 2017, 2563, 2033, 2513, 898, 2031, 390, 1179, 2032, 2333]. Meta-analytic [1239, 1590]. metabolic [2469]. metabolites [216]. metastases [454]. metastatic [1148]. methadone [1700]. Method [2501, 1842, 1357, 371, 1806, 771, 560, 2015, 531, 1820, 1317, 9, 1112, 2443, 2576, 2201, 2260, 1142, 726, 1835, 1460, 1808, 1564, 130, 304, 1244, 536, 2447, 49, 1268, 1624, 1319, 1666, 2429, 87, 1978, 1388, 1778, 225, 1350, 2376, 1470, 663, 1660, 1637, 8, 2273, 2354]. Methoden [1826]. Methodik [1962]. Methodologic [1255]. Methodological [1218, 2207, 1613, 1166, 213, 651, 101, 98, 100, 99, 284]. methodologies [1940, 260]. Methodology [2312, 1498, 625, 1278, 1366, 409, 966, 2044, 1004, 841, 1365, 2565, 584, 166, 407, 1406, 1651, 481, 933]. Methods [288, 398, 1995, 1380, 781, 576, 1399, 733, 145, 788, 2064, 1813, 2133, 389, 1787, 259, 1861, 2170, 622, 1152, 1003, 688, 1641, 2382, 2373, 614, 1186, 916, 25, 1518, 424, 345, 361, 278, 594, 305, 2000, 1760, 680, 521, 2558, 27, 1844, 1379, 244, 2262, 414, 1730, 2011, 1627, 942, 1026, 644, 433, 1045, 2283, 1718, 2117, 1650, 276, 578, 1908, 1397, 1364, 566, 654, 1059, 422, 1336, 856, 1369, 2648, 2416, 1764, 2058, 204, 2489, 242, 440, 1464, 2013, 950, 1440, 179, 2499, 2364, 176, 181, 289, 2540, 773, 1574, 964, 1215, 968, 358, 1287, 2123, 2122]. methods [2464, 1002, 2358, 526, 1793, 129, 1629, 212, 1079, 2048, 2629, 1866, 2479, 626, 1776, 392, 1105, 1996, 555, 201, 162, 794, 246, 527, 1223, 1580, 643, 1007, 2563, 2541, 2162, 621, 1191, 1001, 1828, 2555, 1058, 1665, 2456, 977, 1013, 317, 620, 2644, 1166, 1799, 2599, 2600, 1981, 2167, 94, 2494, 1037, 2193]. metropolis [1438]. Metuchen [1004]. Michael [2368, 1048, 2276, 91, 111, 933, 1194, 285, 1897, 2178, 2309]. Michaelis [1683]. Michelle [2571]. microcomputers [401]. microsimulation [2284]. Mid [1144, 362, 1080, 586]. Mid- [1144, 1080]. Mid-1980 [362]. mid-point [586]. middle [1618]. migration [2342]. Mikhail [1972]. Miller [770, 327].

Milwaukee [2006]. Mimms [402]. Ming [2150]. minimize [2113]. minimum [2553, 609, 1289, 2449]. Minitab [482]. Minneapolis [2296]. minorities [1240]. mins [777]. Misclassification [278, 594, 1522, 1330, 331, 1621, 160, 227, 2570, 332, 1329]. Misconduct [1798, 1014]. misleading [2609]. mismeasured [2269]. misreporting [212]. misselection [161]. Missing [1696, 2095, 2098, 2537, 1647, 491, 2597, 2102, 1462, 2096, 505, 2015, 2111, 1115, 2109, 2110, 1988, 2243, 2484, 2364, 2536, 2538, 964, 1694, 1243, 1930, 2581, 1755, 1756, 2497, 80, 2113, 2108, 1331, 2388]. missing-data [2597, 1694]. missingness [2114]. mission [1752]. misspecification [1368, 1766, 1805]. Mitchell [554]. mixe [2347]. Mixed [1905, 973, 2163, 1980, 2338, 2268, 1818, 2290, 2538, 1992, 2386, 2591, 2539, 15, 834, 2413, 2436, 2231, 2176, 1741]. Mixed-effects [973, 1992]. mixes [2346]. Mixture [2362, 1769, 1602, 859, 1265, 2083, 2402, 1622, 1681, 141, 2028, 2581, 2119, 861, 1552, 774, 2460, 1451]. Mixtures [2427, 36, 1856]. ML [2159]. MLE [2391]. mode [12]. Model [1264, 1340, 1770, 2606, 1618, 446, 135, 1869, 822, 1966, 1635, 1797, 821, 1158, 2321, 2351, 2466, 949, 412, 1203, 2357, 1699, 1096, 820, 2340, 2597, 1222, 1817, 793, 82, 2163, 2636, 2242, 13, 2083, 1980, 2433, 1320, 1377, 576, 373, 69, 753, 1351, 365, 1976, 2399, 1648, 1576, 296, 2117, 1189, 917, 2359, 694, 2120, 2268, 2566, 1681, 1017, 1623, 2271, 2125, 1737, 574, 1410, 1402, 545, 914, 1521, 1839, 681, 1545, 2151, 1482, 1785, 378, 1333, 37, 485, 224, 2013, 1975, 1565, 2137, 761, 1573, 2403, 1507, 1562, 2538, 332, 1563, 507, 2327, 1368, 1928, 824, 1771]. model [2227, 1353, 659, 1992, 1293, 1801, 2426, 2119, 2591, 273, 1957, 601, 640, 1135, 474, 904, 1768, 2026, 137, 498, 880, 641, 1495, 1789, 2078, 2450, 310, 2500, 1145, 1778, 2231, 2045, 621, 1191, 1749, 1916, 1741, 969, 1805, 676, 623, 1617, 1929, 2410, 1451, 2470, 1324, 2115, 953, 1400]. Model-based [1770, 2606, 1618, 446, 1158, 2117, 659, 1293]. model-building [793]. model-free [1293]. Modeling [1519, 1888, 2395, 1866, 2461]. Modelling [2518, 2458, 602, 2579, 2232, 1387, 1903, 753, 106, 963, 2120, 2396, 1214, 1425, 1775, 2323, 2135, 1608, 2241, 1638, 1562, 360, 233, 1416, 2610, 1137, 1945, 2513, 1469, 11, 1210, 2529, 962, 2502, 1125, 425, 2163, 1832, 754, 2446, 2284, 2577, 2490, 2154, 1492, 1636, 1289, 1513, 725, 1180, 2425, 631, 2046, 950, 1867, 1401, 2521, 2307, 1766, 2171, 541, 1042, 2457, 136, 2404, 2272, 1904, 2027, 23, 2053, 2093, 1314]. Models [2253, 1814, 2602, 1887, 1729, 1899, 1194, 1923, 2194, 1710, 1726, 1166, 1912, 1788, 1872, 864, 2529, 520, 757, 878, 975, 477, 1746, 228, 607, 1265, 430, 370, 1773, 2256, 630, 229, 2503, 55, 1695, 1918, 1379, 756, 1676, 1320, 2230, 1417, 1977, 2362, 335, 746, 2580, 1249, 1999, 494, 588, 2435, 2244, 2469, 408, 1712, 1988, 1622, 993, 1020, 759, 1093, 2269, 1466, 315, 1095, 1513, 1792, 1987, 469, 1504, 1914, 2501, 1735, 316, 1022, 1769, 1711, 1413, 2319, 42, 1411, 2522, 926, 1767, 2473, 2304, 901, 1328, 2290, 182, 1033, 1765, 1927, 297, 2531, 2630, 1409]. models [1491, 2617, 2030, 2086, 2047, 2520, 927, 2536, 2325, 1611, 1211, 2564, 1077, 1398, 1905, 333, 200, 108, 478, 2306, 2066, 784, 1224, 2200, 1146, 2012, 2581, 1602, 973, 1343, 928, 2386, 751, 752, 2255, 2048, 837, 2172, 1755, 1756, 1772, 2427, 1128, 1493, 1759, 2174, 619, 1700, 2539, 1108, 678, 859, 1552, 774, 674, 1906, 1693, 137, 834, 1190, 103, 1612, 87, 241, 1415, 899, 162, 1223, 1580, 1734, 2413, 2436, 306, 1830, 159, 2176, 627, 1161, 1295, 561, 1322, 356, 2240, 2291, 2360, 2584, 1878, 2081, 2007, 1847, 1849, 528, 1898]. Modern [2286, 511, 1901, 2167]. modification [660, 2376]. modified [2185, 676]. Modifying [2050]. Moeschberger [2092]. Molenaar [1887]. Møller [132]. Mollison [1912]. moment [2505]. moments [852]. monitor [2059]. monitored [1940]. Monitoring [231, 1871, 1157, 2630, 704, 1068, 492, 2328, 2627, 702, 1561, 708, 717, 1594, 706, 1571, 849, 1062, 2393, 1584, 699, 1308, 1886, 1590, 713, 1060, 709, 1596, 696, 423, 697, 1336, 1112, 1392, 1064, 711, 351, 1804, 735, 887, 2589, 563, 1419, 718, 1052, 705, 127, 755, 808, 1592, 12, 701, 1955, 834, 1063, 1054, 1220, 2197, 226, 715, 1714, 2613, 693, 1010, 317, 1061]. Monk [482]. Monographs [1640, 92, 93]. monotherapy [724]. monotone [2001, 1755, 1756]. monotonic [919]. monotonically [865]. Monte [433, 2169, 2455, 2611, 2467, 389, 1079, 2541]. Moolgavkar [1562]. Moore [1994]. moral [715]. Moran [1625, 2141, 2514, 1185, 1549]. morbidity [573, 111, 1711, 1246, 1416]. Moreau [1902, 364]. Morgan [1038]. Morphology [543]. Morphometric [731]. Morris [1359]. Morrow [1037]. Mortality [362, 2458, 1234, 261, 908, 1529, 198, 1387, 2445, 443, 469, 16, 856, 141, 2605, 687, 2046, 1605, 1271, 313, 200, 604, 2607, 108, 2608, 1416, 1664, 2508, 839, 1358, 1991, 1018, 688, 1007, 1537, 2240, 1890, 2274, 1443, 111, 1897, 2608]. Mosby [655]. most [457]. Mosteller [1087, 764]. mother [2402]. motion [735]. motor [1881]. Motulsky [2311]. Mould [148]. moving [2508]. Mowinckel [2429]. Moya [2278]. MSDOS [786]. MSG [2273]. much [1997, 63]. Mueller [2395]. Muin [1165]. Muir [362, 1193, 1443]. Müller [1822]. Multi [1746, 2154, 993, 2216, 2046, 1398, 2618, 2012, 2307, 1763, 2455, 2247, 2163, 2444, 1607, 2215, 2356, 2227, 2101, 2100, 1068, 15, 2214, 2500, 1945, 2498, 2470]. multi-armed [1068]. Multi-centre [2216, 2455, 1607, 2215, 2356, 2227, 2101, 2100, 2214, 2498]. Multi-criteria [2618]. multi-dose [2247]. Multi-level [1746, 2154, 2046, 2012, 2307, 2163, 2500]. multi-normal [15]. Multi-stage [1763, 2444, 1945]. Multi-state [993, 1398, 2470]. multicenter [663]. Multicentre [2125, 2106, 764, 565, 1446, 1171, 103]. multifactor [940]. Multilevel [1093, 753, 1878]. multilogit [108]. Multimodality [2009]. multinomial [975, 1852, 519]. multinormal [320]. MultiODA [2274]. Multiple [1836, 294, 1515, 1241, 1445, 188, 823, 1993, 257, 2388, 962, 278, 1027, 2641, 594, 2287, 994, 1572, 515, 1603, 1995, 2256, 960, 1579, 1149, 1377, 1730, 299, 2471, 1837, 1318, 1073, 443, 733, 2271, 83, 1711, 749, 320, 1975, 1391, 2153, 442, 773, 1246, 2164, 1948, 742, 1893, 2596, 2454, 129, 2400, 1754, 2629, 473, 1221, 2539, 1996, 2138, 2640, 48, 1220, 2197, 2162, 1527, 1828, 2245, 1714, 2180].

Multiple-outcome [1515]. multiple-period [742]. multiplicative [1017, 1211]. multiplicative-model [1017]. multiply [2492]. multistage [69, 914, 487]. Multistate [1180, 1466, 901]. Multivariable [1504, 1019, 1982]. Multivariate [1922, 1318, 2416, 355, 394, 2492, 755, 626, 1776, 2539, 2381, 2420, 1635, 2340, 2597, 537, 1676, 628, 534, 2359, 1953, 2282, 653, 1009, 1831, 204, 544, 924, 1147, 2521, 2361, 1343, 1661, 762, 1774, 1641, 2253, 2277]. multiwave [413]. multiway [137]. mumps [106]. Murphy [41, 183, 429]. Murray [2587]. Musen [2179]. Must [61]. Musta [132]. mutation [2364, 1562]. myelogenous [2228]. myocardial [760, 524].

N [1901, 434, 770, 1359, 2393, 2204, 1823, 2646, 43, 2141, 2250, 1846]. Nîmes [388]. naive [2631]. Nandram [2248]. Narrative [2192]. nasopharyngeal [2556]. natal [505]. Nathan [266, 2622, 2624]. National [101, 2104, 22, 2103, 1055, 2615, 1248, 699, 98, 100, 99, 1779, 651]. NATO [827]. natural [2446, 143, 973, 1441]. naturally [1907]. Naukkarinen [2005, 2004]. nchs [673]. NCI [985]. NCIC [1055]. NCIC-CTG [1055]. Neal [1625]. Neale [933]. near [1429, 1536]. nearest [1667]. needle [1142]. negative [1638, 1345, 1102]. neglected [355]. neighbour [1667]. Neil [345]. NEJM [1087]. Nelson [1048]. Nenadic [656]. Neo [2380]. Neo-Fisherian [2380]. neonatal [1117, 108]. neoplasia [952]. nested [2413, 561]. net [50, 545]. Netherlands [933, 2395, 59, 856]. nets [1040]. Network [916, 1189]. Networks [1984, 2156, 2243, 2168]. neural [2156, 1189, 1040, 1984]. neuroimaging [2533]. neurons [968]. neutralization [551]. neutropenia [2534]. newborns [1416]. Newcastle [777]. Newcombe [2648, 2633]. Newell [1206, 1197]. Newton [1898]. NHLBI [694, 2457]. Nicholas [1166, 1838]. Nicolosi [1166]. Niels [1108, 44]. Nielsen [2168]. nih [709, 957]. Nikulin [1972]. Ninth [2008]. NMR [2043]. No [23, 1961, 1901, 1985, 2584, 656, 770, 1087, 827, 1812, 2253, 510, 1938, 1934, 2252, 2310, 2462, 381, 363, 553, 1798, 2397, 345, 1826, 828, 2394, 1984, 1027, 2090, 2147, 2067, 2368, 1048, 1359, 2276, 1922, 2277, 90, 452, 2367, 840, 2312, 2055, 146, 730, 94, 366, 403, 1983, 482, 480, 132, 2408, 326, 1902, 2092, 1876, 244, 362, 2131, 2314, 401, 871, 2602, 1887, 2039, 1026, 384, 2009, 91, 365, 731, 92, 1962, 2204, 2275, 2254, 2018, 364, 144, 2148, 346, 402, 1859, 1799, 206, 451, 2041, 892, 111, 1786, 2070, 1936, 2545, 1911, 2396, 450, 2042]. No [2369, 1878, 147, 2379, 2038, 2134, 872, 2263, 1898, 2296, 2286, 145, 1192, 2081, 93, 2157, 1085, 1899, 2006, 2179, 1193, 110, 1994, 1888, 1970, 1862, 511, 165, 2495, 328, 1086, 1194, 2573, 2378, 404, 2345, 829, 2169, 2586, 327, 2278, 22, 1824, 1371, 1455, 1372, 1454, 1443, 2601, 2091, 1848, 1896, 2068, 841, 2419, 2599, 2007, 2194, 2132, 242, 1972, 2461, 2572, 570, 2544, 2546, 2313, 2082, 1924, 1951, 2040, 2417, 2418, 1847, 1933, 1037, 1278, 655, 112, 24, 2380, 243, 1849, 481, 1813, 1823, 164, 1015, 1950, 148, 2587, 285, 2133, 2494, 1897, 2439]. No [400, 2311, 870, 2571, 1787, 1038, 2585, 526, 2150, 207, 2093, 1982, 1971, 2043, 1860, 1900, 2264, 1838, 1850, 383, 1338, 1877, 2600, 1861, 2053, 2366, 528, 1108, 2383, 2170, 2384, 2181, 2168, 284, 1981, 325, 348, 2019, 1003, 166, 2395, 2192, 347, 43, 2149, 2239, 246, 1937, 527, 44, 1277, 382, 1825, 2381, 2069, 1912, 830, 1025, 2056, 2054, 1846, 1788, 1315, 1314, 1960, 2020, 2180, 893, 1001, 2193, 2008, 184, 2265, 1811, 2178, 2382, 245, 1935, 2167, 614, 434, 432, 758, 945, 597, 1165, 21, 596, 598, 613, 1014, 433, 1153, 786, 554, 759]. no [20, 1000, 933, 667, 944, 1923, 1004, 947, 612, 946, 1002, 668, 921, 1166, 787, 555, 920, 1164]. Noah [43]. nominal [373, 1214, 1425, 129, 2087]. Non [1647, 1730, 1506, 1136, 1212, 1431, 1727, 1464, 1833, 2536, 1891, 1133, 928, 1755, 1756, 1861, 627, 1465, 822, 940, 371, 2594, 2611, 630, 521, 908, 2177, 1462, 795, 505, 2015, 1320, 1843, 2262, 628, 414, 679, 1976, 415, 174, 468, 1908, 1053, 1035, 17, 1752, 271, 379, 227, 2416, 1856, 2304, 2226, 2578, 1638, 332, 1928, 1274, 2564, 1329, 1659, 1992, 918, 868, 54, 1146, 1645, 779, 988, 2542, 1733, 2005, 2074, 2474, 1759, 665, 2079, 1490, 2645, 2138, 1204, 1578, 1604, 2004, 2016, 1880, 561, 977, 1013, 10, 1963, 2408]. non-adherence [779]. non-compliance [1843, 1752, 2578, 2074, 2079]. non-constant [1992]. non-decreasing [1730, 1274]. non-differential [227, 332, 1329]. non-fatal [2138]. Non-hierarchical [928]. non-Hodgkin [795]. Non-ignorable [2536, 2015, 1755, 1759]. non-independent [415]. Non-iterative [1891, 1880]. Non-linear [627, 1320, 628, 1976, 1035, 2564, 868]. non-linearity [271, 1146]. non-monotone [1755, 1756]. non-negative [1638]. non-nested [561]. non-normal [17]. non-normality [665]. non-optimal [1645]. Non-parametric [1647, 1730, 1506, 1136, 1212, 1431, 1727, 1464, 1833, 1133, 1465, 822, 940, 371, 2594, 2611, 630, 521, 414, 1908, 2416, 2226, 988, 2542, 1578, 1604, 2016, 977, 1013, 1963]. Non-Precise [1861]. non-proportional [1928]. non-random [1462, 2304, 54, 2408]. non-randomized [2262, 379, 1490]. non-randomly [505]. non-rare [1659]. Non-response [1755, 1756, 1759]. non-responses [679]. non-reversible [908, 468]. non-small-cell [1053]. non-specialists [1204]. non-specific [10]. non-uniformly [1733, 2005, 2004]. non-unity [174, 918, 2645]. non-zero [2177, 174, 2474, 2645]. Nonlinear [2314, 2007, 1612]. nonparametric [94, 2494, 2265]. Nonsense [1192]. Nordic [132]. Noreen [244]. Norell [2254]. normal [17, 204, 517, 543, 238, 270, 518, 15, 2567, 1819, 2261]. normality [339, 665, 1511]. normally [2330]. Norman [2147, 366]. Normand [2637]. norms [2012]. Norrell [511]. North [231]. North-Norwegian [231]. Norton [2313, 316]. Norton-Simon [316]. Norway [132, 137, 310, 1324, 1889]. Norwegian [231, 2105]. Nostrand [656]. Note [647, 75, 497, 1366, 223, 221, 557, 86, 611, 2496, 138, 442, 773, 2188, 2504, 591, 488, 241, 275]. Notes [384, 2296, 1404, 404]. Notifiable [2615]. Nougue [364, 1902]. November [308]. NSABP [2104]. Nuclear [1950, 1439, 1184]. Nuffield [553, 450]. null [174, 918, 2645]. Number [1370, 1647, 2246, 1347, 1729, 2468, 788, 1154, 1485, 930, 2349, 501]. Numbers [2313, 507, 1629, 1500]. nursing [1377]. nutrition [1469]. nutritional [1048, 559, 1033, 890]. NY [1443].

O [2128, 2129, 2133, 2168, 2395]. Oaks [2131, 1899, 1923, 1896, 1900, 1788, 2265]. obesity [2337]. observation [1145, 527]. Observational [1961, 1738, 1845, 264, 1913, 971, 107, 1709, 2151, 2530, 1030, 2531]. observations [1163, 2442, 2446, 915, 1078, 1212, 924, 2198, 1930, 2581, 579]. observed [1331]. observer [2155, 664, 1214, 2118, 1425]. observers [2088]. observing [1646]. obstructive [1459, 1945]. obtain [1012]. occasional [887]. Occupational [308, 1315, 345, 443, 1538, 403]. occurrence [2127, 451, 727, 1809, 342, 1929, 1451, 2126, 454]. October [831]. odd [2349]. odds [1354, 205, 2424, 638, 1320, 1450, 1915, 1080, 2499, 1841, 438, 523, 135, 338, 943, 2450, 1617, 1626]. Oden [2141]. Odom [1626]. Odom-Maryon [1626]. oesophageal [964]. oesophagus [92]. oestrogen [216]. Office [596]. Ogden [2384]. O’Hagan [1933]. Ohio [2240, 827]. old [2389]. Olsen [510, 920]. omega [672]. omission [1768]. omitted [1767, 1146, 547]. on-therapy [1646]. Oncology [1062, 697, 2098, 2107, 2116, 305, 1151, 2489, 145]. one [2302, 373, 299, 1151, 9, 2444, 1273, 1383, 853, 2162, 1750, 2008]. one-dimensional [853]. one-sample [1750]. one-sided [1151, 1273, 1383, 2162]. one-step [9]. one-way [299]. ongoing [2050]. only [1926, 1990, 2324, 928]. onset [70, 2359, 879, 2457, 1334]. ontrol [228, 263]. opacities [1480]. Open [325]. Opening [692, 1231]. Operating [2586, 1869, 1995, 2140, 1495, 1963]. operation [708]. operative [565, 1077]. operator [17]. ophthalmic [1397]. ophthalmologic [1617]. Opportunities [1359]. Optimal [2014, 2267, 1795, 1731, 1452, 2444, 2557, 2153, 563, 12, 1101, 1697, 2303, 297, 1272, 2305, 931, 2200, 1645, 2051, 1019, 1890]. optimally [1475]. Optimization [2480]. Optimum [2477, 177, 476, 2643]. Option [2292, 377]. Option-3 [2292]. O’Quigley [1902, 364]. oral [2500]. order [849, 293, 2185, 1287, 1480, 338, 1968]. ordered [2518, 2317, 1832, 532, 1568, 1333, 2226, 1355, 1907, 964, 1682, 2322, 897, 1179, 1617]. Ordinal [2322, 1651, 1788, 430, 393, 229, 2001, 1795, 1320, 296, 1506, 1397, 1352, 1095, 2125, 1525, 2135, 2202, 1371, 1573, 2030, 927, 2565, 2358, 1493, 1776, 1319, 880, 1174, 2500, 1734, 1210]. organ [1298]. Organization [537, 2125]. organizing [1281]. Ørnulf [1108]. Ørsted [896]. orthogonal [723]. Osborn [363]. Ostroff [1454]. other [1186, 1124, 146, 276, 1127, 868, 541]. otitis [179, 8]. outbreak [1789]. outbreaks [375]. outcome [2357, 1515, 2590, 2636, 952, 2077, 2125, 2576, 227, 765, 1306, 2554, 729, 2348, 2329, 2430, 2162, 2228]. outcomes [2287, 2256, 2558, 1832, 2230, 2052, 1157, 1073, 1077, 742, 996, 2539, 1906, 1939, 2025, 2328, 7, 1106, 1209, 1220, 2197, 2162]. Outlier [52, 2492, 643, 2292]. outliers [2257]. outlying [354]. outs [891, 636, 2434]. ova [1381]. ovary [2616]. over-the-counter [188]. over. [946]. overall [638, 1733, 2005, 943, 2004, 2016]. overdispersion [141, 2447]. overdose [2152]. overlap [1432]. overlapping [1724, 2412]. oversight [252]. Overview [1798, 618, 257, 1828, 620]. overviews [67, 1406]. overweight [2606]. ovulation [216]. owing [523]. OXFORD [944, 23, 614, 2584, 434, 510, 345, 1165, 1984, 2147, 2368, 1048, 1922, 840, 366, 1876, 362, 2254, 206, 892, 2396, 759, 1194, 788, 1370, 841, 1037, 112, 1849, 1897, 1973, 2311, 207, 2093, 1971, 284, 348, 787, 1003, 347, 43, 2149, 246, 1025, 1315, 893, 920, 184, 2178, 245, 788].

P [432, 770, 827, 2368, 1359, 94, 1871, 403, 2092, 362, 613, 2165, 2369, 404, 2473, 1037, 2066, 2494, 890, 2150, 1166, 2645, 1003, 2429, 1960, 1822, 893, 2193, 2167, 777]. P.O [596]. PA [2571]. Pace [2380]. package [1553]. packages [861]. paediatric [2432]. Page [2133, 2265]. Pages [1961, 510, 1359, 871, 2296, 1193, 511, 1194, 2586, 2439, 528, 1025, 1315, 1811, 23, 1901, 614, 1985, 2584, 434, 432, 656, 770, 758, 1087, 827, 1812, 2253, 945, 597, 1938, 1934, 2252, 2310, 2462, 381, 363, 553, 1798, 2397, 345, 1826, 828, 2394, 1165, 1984, 1027, 21, 2090, 2147, 2067, 596, 2368, 1048, 2276, 1922, 2277, 90, 452, 2367, 840, 2312, 2055, 146, 730, 94, 366, 598, 403, 1983, 482, 480, 132, 2408, 326, 1902, 2092, 1876, 244, 362, 613, 2131, 2314, 401, 2602, 1887, 1014, 2039, 1026, 384, 2009, 433, 91, 365, 731, 92, 1962, 2204]. pages [2275, 2254, 2018, 364, 144, 1153, 2148, 346, 786, 402, 1859, 1799, 206, 451, 2041, 892, 111, 1786, 2070, 1936, 554, 2545, 1911, 2396, 450, 2042, 759, 2369, 1878, 147, 2379, 20, 2038, 2134, 872, 2263, 1898, 2286, 145, 1192, 2081, 93, 2157, 1085, 1899, 2006, 2179, 110, 1000, 1994, 1888, 1970, 1862, 933, 165, 2495, 328, 667, 1086, 944, 2573, 2378, 404, 2345, 829, 1923, 1004, 2169, 788, 947, 1154, 327, 2278, 22, 1824, 1370, 1371, 1455, 1372, 1454, 1443, 2601, 2091, 1848, 1896, 2068, 841, 2419, 2599, 2007, 2194, 2132, 242, 1972, 2461, 2572, 570]. pages [2544, 2546, 2313, 2082, 1924, 612, 1951, 2040, 2417, 2418, 1847, 1933, 1037, 1278, 655, 112, 24, 946, 2380, 243, 1849, 481, 1813, 1823, 164, 1015, 1950, 148, 2587, 285, 2133, 2494, 1897, 1002, 400, 2311, 668, 870, 2571, 1787, 1038, 2585, 526, 2150, 207, 2093, 1982, 1971, 921, 2043, 1166, 1860, 1900, 2264, 1838, 1850, 383, 1338, 1877, 2600, 2547, 1861, 2053, 2366, 1108, 2383, 2170, 2384, 2181, 2168, 284, 1981, 325, 348, 787, 2019, 1003, 555, 166, 2395, 2192, 347, 43, 2149, 2239, 246, 1937, 527, 44, 1277, 382, 1825, 2381, 2069, 1912]. pages [830, 2056, 2054, 1846, 1788, 1314, 1960, 2020, 2180, 893, 1001, 2193, 920, 2008, 184, 1164, 2265, 2178, 2382, 245, 1935, 2167]. Pahkinen [1787]. pain [1814]. pair [415, 729]. pair-matched [729]. paired [1150, 2246, 1347, 1020, 2202, 1212, 1371, 234, 181, 1368, 2295, 1566, 2328, 2124, 2633, 1954, 1344, 441, 2410]. paired-sample [2124]. pairs [1347, 1352, 836, 2468, 2188, 136]. pairwise [425]. Pakistan [212]. Palace [388]. Paling [2038]. Panel [708, 717, 2132, 1900, 413, 1328, 761]. panencephalitis [396]. Panfilov [2345]. panic [828]. Paper [1899, 1923, 2571, 730, 1829, 2254, 1838, 245]. Paperback [2312, 2147, 2277, 1859, 1936, 2379, 404, 947, 1896, 1037, 1823, 246, 366]. papers [101, 912, 628, 632, 418, 2266]. papillomavirus [952]. paradigms [2073]. paradox [163]. parallel [1856, 2025]. Parameter [1970, 2609, 518, 1330]. parameterization [477, 228]. parameters [1946, 236, 1999, 315, 1017, 292, 1401, 2540, 2564, 935, 478, 2231, 2281, 2375, 2376, 1920, 2360, 2027]. Parametric [521, 1712, 2138, 2016, 977, 1013, 822, 940, 2649, 1647, 371, 2594, 2611, 630, 486, 929, 1379, 414, 1730, 1576, 2435, 1506, 1986, 1908, 1953, 2269, 1623, 1781, 1136, 1212, 2416, 1764, 1431, 1727, 2290, 2226, 1464, 1833, 1133, 2352, 988, 2542, 1772, 1759, 1578, 1604, 1830, 390, 1465, 1963, 2149]. parasites [1466]. parasitic [142]. Parati [1862]. Paris [308, 288, 1902, 364, 147]. Parkin [1193]. Parmar [1812]. part [1914, 751, 752]. partial [2259, 2156, 591, 2088, 976]. partially [544, 1807]. Participants [1050]. Participatns [799]. particular [158]. particularly [1732]. partition [1907, 2565]. partitioning [1661, 1018]. partner [2243]. partners [1485]. party [1366]. parvovirus [1622]. past [2044]. patch [2497]. pathogenesis [1135, 15]. paths [877]. patient [20, 301, 1391, 1942, 1408, 2430, 2073, 1527, 561, 540]. Patients [1994, 1926, 1476, 70, 36, 901, 1680, 2003, 761, 2617, 2158, 1287, 1490, 1461, 275, 2002, 419, 1890, 2274]. Patricia [1454]. Patrick [450]. Patsy [1849]. Pattern [1984, 2581, 1021, 1337, 1496]. pattern-mixture [2581]. patterns [2288, 217, 2083, 1115, 2608, 1955, 1041, 1031, 855]. Paul [614, 452, 2586, 1973]. PB [1372]. pbk [2147, 365]. PC [786, 1554]. pcnonlin [435]. Peace [146, 871]. Pearce [345]. pectoris [14]. pediatric [1151]. Pedigree [24]. Pedotti [1862]. Pei [872]. PEMD-92-18 [596]. Penalized [1173, 560, 2084, 1551, 1513, 1792, 2139]. penetration [1381]. people [930, 1439, 2378]. Pepe [2642]. peptic [2158]. per-act [1137]. percentage [595]. percentages [1779]. Percentile [1267, 1289]. percentiles [1992]. perform [2316]. Performance [2597, 1872, 2357, 2636, 240, 469, 1413, 438, 1964]. performances [2363]. perinatal [687]. perinatally [1132]. period [1990, 1844, 427, 240, 1009, 1008, 914, 2387, 2473, 485, 629, 1491, 1132, 742, 1726, 1479, 2171, 2172, 747, 1012, 1478, 1350]. period-cohort [2473, 2172]. periodic [2589]. periodontal [1223, 1580, 317]. peritoneal [11]. peritonitis [11]. permits [2444]. permutation [1572]. Person [2246, 546, 1072]. Person-time [2246]. Personal [58, 1476, 68, 1461]. Personalities [2204]. persons [1240, 2060]. perspective [2612, 2118, 844, 1533, 259, 1072, 2380, 365, 1055]. Perspectives [1823, 1070, 1071, 1669, 2537, 718, 810]. Persson [2019]. Peter [2368, 362, 759, 93, 667, 1848, 2599, 2572, 325, 245, 73, 68]. Peto [362, 1897]. Petr [144]. Petre [480]. Pett [2265]. pH [1452]. pH-profiles [1452]. Pharmaceutical [1640, 1935, 385, 1639, 711, 701, 1744, 866, 166, 2414]. pharmaco [2210]. pharmaco-economic [2210]. pharmacodynamic [435, 1295]. pharmacoepidemiological [2223]. pharmacoepidemiology [1613, 2224]. Pharmacoepidemiology. [2082]. pharmacokinetic [435, 1293, 1632, 1866, 1295, 2375, 2376]. pharmacokinetic/pharmacodynamic [435, 1295]. pharmacokinetics [1298]. pharmacology [723]. pharmacovigilance [2224, 306].

Pharmakologischer [1962]. pharmazeutischen [1628]. phase [2195, 2152, 102, 305, 723, 1178, 2014, 2267, 771, 1351, 1101, 1854, 566, 1317, 637, 514, 2444, 1925, 2443, 1110, 2489, 2561, 350, 1283, 214, 1947, 7, 1346, 1678, 6, 2430, 2376, 2431, 1299, 780, 1722]. phases [2488]. phenomenon [379]. Philadelphia [598, 2571]. Philip [2022]. Philippines [475]. physical [2582]. physician [376, 2412]. Physicians [1085, 2571]. physiological [941, 869]. Physiology [1500]. Picard [404]. Pickles [787]. piecewise [1482, 1785]. Pierre [147]. Pignon [147]. pilot [1175, 1393, 2433, 2463, 1157, 766, 1624, 2429, 2631, 2632]. Pine [2547]. Pinol [2368]. Pitfalls [745, 1455, 213, 846]. Pitkäniemi [2005, 2004]. Place [402]. placebo [188, 535, 40]. placebo-controlled [40]. Plainview [1443]. plan [1419]. Planned [514, 2032]. Planning [280, 1639, 1300, 2348, 2414, 1519, 402, 1310, 2214, 6, 1253]. plans [2207]. plant [1439, 1184]. plaque [1452]. plasma [2387]. Plasmodium [1466]. platelet [642]. Play [1722, 2398, 1691, 686]. play-the-winner [2398]. Playing [1896]. Plenum [1888]. plots [1402, 426]. Plotting [901]. Plus [2314, 777, 786, 1055, 1901]. Pneumocystis [1890]. Pneumonia [2125, 1890]. pock [62]. pocket [665]. pocket-calculator [665]. Pocock [2639]. point [2280, 324, 1748, 1429, 2614, 2465, 586, 1536, 138, 1541, 1666, 702]. point- [1541]. point/counterpoint [2280]. points [2414]. poisoning [1538]. Poisson [2649, 215, 1144, 998, 236, 1986, 1078, 742, 2447, 137, 1929]. policies [2076]. policy [1519, 22]. pollutants [1793]. Pollution [1950, 1544, 1440, 2508, 1666]. polychotomous [477, 228, 1794]. polynomial [723, 753, 1009, 1322]. Polynomials [1472, 2377]. polyps [684, 1929]. polytomous [1977]. Pooled [2184, 180, 642]. Pooling [357, 9]. Population [1874, 1734, 2458, 602, 2514, 979, 1662, 2362, 1104, 2616, 1800, 2293, 1781, 1139, 22, 1634, 1625, 1345, 968, 1643, 543, 1185, 2200, 928, 263, 1733, 2005, 273, 1866, 1616, 584, 1535, 772, 310, 2004, 2141, 1945, 1350, 1295, 1855, 2375, 2376, 2412]. Population-averaged [1734]. population-based [979, 2362, 2616, 543, 928, 263]. populations [1032, 1526, 544, 1815, 1997, 2153, 2540]. Porte [308]. Portnoy [554]. positive [188, 1892, 495, 1956, 1127, 1102, 486, 929]. possible [1646, 290, 2499, 342]. post [849, 1211, 235, 108, 1841, 2224, 907, 1465, 2336]. post-alarm [849]. post-lead-time [1465, 2336]. post-market [2224]. post-marketing [907]. post-menopausal [1211]. post-neonatal [108]. post-test [235]. Poster [633, 1257, 1670]. postpartum [1620]. Potential [2360, 1943, 1011, 1941, 1539, 310]. Potentials [261]. potroom [1105]. Potters [402]. Powell [92, 93, 1193]. Power [1797, 737, 2052, 1668, 1831, 550, 1565, 1784, 851, 344, 1219, 1783, 728, 2598, 2433, 1864, 1691, 1080, 980, 990, 1355, 376, 39, 765, 2189, 683, 1851, 1541, 782, 1468, 853, 1511, 1422, 2568, 2228, 1184, 1428, 1540, 1805]. powerful [83, 2079]. Poynard [147]. practicable [1452]. Practical [949, 1812, 435, 2147, 2106, 730, 2314, 2208, 709, 1475, 2133, 128, 1065, 2113, 1406, 1054, 2102, 366, 2302, 2519, 1648, 1317, 2464, 729, 1063, 1968, 184, 381, 786, 1799, 1951, 1787, 2069]. Practice [2573, 1860, 253, 1000, 1658, 1074, 1052, 2500, 1859, 2169, 2586, 1823]. practices [696]. Pre [505, 1748, 1628, 1156, 2224, 1991]. Pre-AIDS [1991]. pre-clinical [1748, 1628, 1156]. pre-market [2224]. Pre-natal [505]. Precise [1861]. precisely [1892]. Precision [917, 1078, 1805]. preclinical [1717, 1351]. predict [2289, 1451]. Predicting [1381, 985, 455, 128, 1890, 2274, 10, 2357, 2636, 35, 2425]. Prediction [1125, 606, 1857, 1415, 1966, 1519, 1976, 469, 2125, 1202, 1271, 1881, 2337, 1957, 1578, 202, 2272, 1145]. predictions [1078, 901]. Predictive [1693, 159, 2128, 666, 1840, 2129, 1676, 1510, 2451, 1302, 1077, 1102]. predictor [13, 2229]. predictors [1914, 1764, 2326, 2582]. Preface [1827, 1196, 798, 2094, 152, 168, 387, 635, 895, 1168, 1457, 1740, 2021, 2621, 1122, 691, 955, 2603, 1707, 114, 1583, 2071, 249, 1531, 2205, 2517, 2057, 1427, 1280, 2476, 1653, 287]. preference [549]. preferences [1507]. prefix [531]. pregnancy [434, 1987, 1906, 745, 981]. preliminary [271, 733, 211, 1287]. preliminary-testing [733]. Prem [827, 1359]. premature [524]. premenopausal [1378]. Prentice [2067]. prescription [1561]. Prescriptions [383]. presence [1377, 1965, 648, 2244, 1103, 2065, 2389, 2422, 218, 340, 1342, 295, 824, 489, 135, 80, 977]. present [1634, 1865, 2143, 2204]. Presentation [1259, 611, 2112, 488, 346]. presentations [1257, 1670]. Presenters [1229, 1230]. Presenting [2302]. preserve [2370]. preserving [2151, 824]. prespecification [2399]. prespectives [1362]. Press [23, 2584, 434, 770, 758, 510, 1938, 1934, 2252, 345, 1165, 1984, 2147, 1048, 1922, 840, 366, 362, 2009, 731, 92, 2275, 2254, 144, 346, 786, 1799, 206, 892, 111, 1786, 2070, 1936, 554, 2396, 759, 2379, 20, 1898, 145, 93, 2006, 1888, 1862, 1194, 2378, 829, 1004, 788, 947, 22, 1372, 1443, 841, 570, 2040, 1037, 112, 24, 1849, 481, 164, 2587, 1897, 2311, 526, 207, 2093, 1982, 1971, 921, 1166, 1338, 2547, 1861, 348, 787, 1003, 2192, 347, 43, 2149, 246, 1912, 830, 1025, 1846, 1315, 1001, 184, 2178, 245]. pressure [1676, 2015, 644, 661, 546, 1893, 137, 834, 359, 1041, 621, 1191, 1092, 2388]. prevalence [1990, 2128, 1840, 2129, 1780, 2558, 587, 1903, 1854, 588, 2058, 1345, 455, 2606, 1659, 1694, 2352, 232, 330, 601, 201, 2592, 2184, 2431]. prevalent [1563, 522, 1031]. prevented [1188, 227]. Prevention [2104, 25, 2628, 1111, 673, 1073, 1681, 672, 686, 1154]. Preventive [1154, 680, 892]. Price [1961, 1901, 614, 1985, 777, 770, 1087, 827, 1812, 2253, 510, 597, 1934, 2252, 2310, 2462, 381, 363, 1798, 2397, 345, 1826, 2394, 1984, 1027, 2090, 2147, 2067, 596, 2368, 1048, 1359, 2276, 1922, 2277, 452, 2367, 840, 2312, 2055, 146, 730, 94, 366, 598, 403, 1983, 482, 480, 132, 2408, 326, 1902, 2092, 1876, 244, 362, 613, 2131, 2314, 401, 871, 2602, 1887, 2039, 384, 2009, 91, 365, 731, 1962, 2204, 2254, 2018, 364, 144, 2148, 346, 1859, 1799, 206, 451, 2041, 892, 111, 1786, 2070, 1936, 554, 2545, 1911, 2396, 450, 2042, 2369, 1878, 147, 2379]. Price [2134, 872, 2263, 1898, 2296, 2286, 145, 1192, 2081, 93, 2157, 1085, 1899, 2006, 2179, 110, 1994, 1888, 1970, 1862, 511, 2495, 328, 1086, 1194, 2573, 2378, 404, 2345, 829, 1923, 2169, 2586, 327, 1824, 1370, 1371, 1455, 1454, 1443, 2601, 2091, 1848, 1896, 2068, 841, 2419, 2599, 2007, 2194, 2132, 242, 1972, 2461, 2572, 2544, 2546, 2313, 2082, 1924, 612, 1951, 2040, 2417, 2418, 1847, 1933, 1278, 112, 2380, 243, 1849, 481, 1813, 1823, 164, 1950, 148, 2587, 285, 2133, 2494, 1897, 2439, 400, 1973, 2311, 870, 2571, 1787, 2585, 526, 2150, 207, 2093, 1982]. Price [1971, 2043, 1860, 1900, 2264, 1838, 1850, 383, 1338, 1877, 2600, 2547, 1861, 2053, 2366, 528, 1108, 2383, 2170, 2384, 2181, 2168, 284, 1981, 325, 348, 2019, 166, 2395, 2192, 347, 2149, 2239, 1937, 527, 1277, 382, 1825, 2381, 2069, 1912, 830, 1025, 2056, 2054, 1846, 1788, 1315, 1314, 1960, 2020, 2180, 893, 2193, 2008, 184, 2265, 1811, 2178, 2382, 1935, 2167, 23, 434, 432, 656, 758, 945, 553, 1165, 21, 1014, 1026, 433, 1153, 786, 759, 20, 1193, 1000, 933, 667, 944, 1247, 1004, 788, 947, 1154, 2278, 22, 570, 1037, 655, 24, 946, 1015]. price [1002, 668, 1038, 921, 1166, 787, 1003, 555, 43, 246, 44, 1001, 920, 1164, 245, 165]. primary [606]. primer [656, 91, 328, 1704, 1153, 2545, 2264]. Prince [1544]. Princeton [1934]. Principal [543]. Principles [526, 2395, 163, 1639, 1628, 2496, 2090, 2380, 2587]. prior [867, 1513, 780, 2245]. priorities [2618]. prioritization [1744]. priors [1436, 630, 1766, 2460]. Priscilla [1972]. Probabilistic [1235, 2363, 2168, 2020]. Probabilities [901, 852, 2389, 320, 484, 605, 1356, 886]. Probability [2070, 1758, 1219, 173, 722, 1808, 235, 734, 1137, 492, 1678, 2379]. probit [2282, 1333, 355]. problem [2437, 2577, 2109, 1035, 1432, 234, 779, 988, 353, 561]. Problems [264, 303, 1202, 251, 949, 2095, 1218, 2098, 1440, 1242, 1065, 263, 1766, 2113, 163, 1054, 651, 730, 1001]. procedure [987, 2641, 960, 1149, 764, 763, 652, 292, 355, 649, 442, 350, 2400, 564, 2511, 640, 904, 2640, 1741]. Procedures [2586, 666, 1445, 580, 1615, 2438, 725, 83, 999, 1273, 563, 1236, 2528, 2203, 473, 1084, 1478, 1282, 2632]. Proceedings [827]. Process [1640, 2438, 1748, 2524, 2526, 2527, 2078]. processes [334, 2338, 613, 404, 1108]. processing [1907, 2043]. produced [777]. producer [777]. products [1366, 2438]. Professionals [1973, 164, 2554, 2041, 598, 328, 2439]. professions [347]. Professors [2552]. profile [1527]. profiles [1638, 1298, 1452]. profiling [2011]. progesterone [216]. Prognostic [140, 2543, 1504, 1418, 1956, 12, 906, 1758, 1330, 1757, 454, 1400]. program [452, 1806, 786, 2545, 1460, 1552]. Programme [1037, 1700, 2500]. programmes [2612, 1142]. Progress [2252, 480, 407]. progression [1966, 444, 2579, 486, 2289, 722, 1998, 139, 1223, 1580, 1882, 2273]. Project [2104, 1744]. projecting [601]. Projection [2061, 1104, 371, 1126, 1129, 1116, 1127]. projections [574, 1656, 358, 330]. projects [1690]. proliferative [1091]. Prologue [1227]. promises [251]. promising [1346]. Propensity [2262]. Properties [590, 850, 278, 594]. prophylactic [951, 2327]. prophylaxis [964]. Proportion [109, 1874, 1097, 2331, 1187, 2648, 1894, 1047, 2122, 1616, 1334, 741]. proportional [949, 412, 2242, 576, 1648, 1999, 1264, 2566, 2271, 1735, 1410, 1022, 1369, 681, 42, 1482, 1785, 176, 1928, 1353, 1564, 1801, 2427, 2450, 1749, 1916, 540, 1324]. proportionality [2450]. proportionate [16]. Proportions [1784, 2332, 2130, 609, 737, 1045, 321, 1302, 2237, 1959, 1851, 2295, 2123, 2187, 1468, 2124, 2633, 2635, 1344, 1880, 1024]. proposal [2514, 1150, 272]. proposals [704]. Prorok [770]. Proschan [2641, 2640]. Prospective [2189, 728, 918, 1470]. prospectively [2032]. prostate [973]. Protecting [1662]. protective [2062]. protein [475]. proteins [1035]. protocol [681]. protocols [1724]. Proulx [164]. provided [252]. provider [2011]. Providing [2220]. Provincial [553, 450]. proving [2330, 1356]. provisional [351]. Proximity [2366]. pseudo [1760, 1298, 899]. pseudo-likelihood [1760]. pseudo-profiles [1298]. Psychiatric [2150, 2282, 2426]. psychiatry [821]. Psychologist [1164]. psychology [1026]. psychopharmacologic [494]. psychophysical [1969]. psychotherapy [444]. Publ [2008]. Public [1245, 1234, 875, 1125, 1990, 456, 874, 1235, 788, 1154, 585, 1658, 1253, 2613]. Publication [451, 1193, 2461, 2193]. Publications [2309, 2090, 2277, 2131, 1898, 1899, 944, 1370, 1896, 2194, 1171, 1813, 1900, 284, 2019, 2239, 1788, 893, 920, 2265, 1206, 73, 362]. Published [402, 165, 2316]. Publishers [1876, 1454, 1848, 1823, 2133, 1877, 1825]. Publishing [2408, 2039, 1859, 2380, 2571]. Pukkala [132]. pulmonary [1459]. pulse [942, 869]. pulses [224, 1215, 643]. Puna [132]. purity [1247]. purposes [402]. pushback [763]. putative [1440, 2339]. Putting [2310]. puzzle [846]. PWS [1026, 400]. PWS-Kent [1026, 400].

Q [2573]. QMMP [261]. QOL [2100]. Quadrat [1555]. Qualification [2369]. qualifier [232]. qualitative [763, 234, 1909, 782]. Quality [2102, 2096, 156, 2160, 2006, 2099, 1872, 1746, 2095, 2106, 756, 1378, 2111, 2114, 2580, 1863, 2117, 2109, 2097, 2110, 2438, 713, 2104, 1567, 2098, 1413, 1474, 2467, 2105, 570, 2112, 2107, 2101, 1236, 78, 2103, 411, 2459, 2113, 84, 2513, 2116, 2108, 2115, 104]. quality-adjusted [1378, 2467]. quality-of-life [2103]. Quan [2143]. quantal [1038]. Quantification [1934]. Quantifying [160, 2548]. quantile [1094, 865]. quantiles [1329]. Quantitative [1518, 1899, 1923, 1788, 1627, 2549, 2550, 20, 1623, 2551, 2224, 1500, 2273]. Quantities [2192]. quasi [2186, 1792, 321, 1605, 1094, 136]. quasi-exact [321]. quasi-likelihood [2186, 1792, 1605]. quasi-symmetry [136]. QUECUS [2278]. Queensberry [402]. QUERCUS [2278, 2278]. Quest [1934]. questionnaire [709, 926, 976]. questionnaires [679].

R [777, 432, 758, 2394, 2147, 1359, 366, 613, 401, 2393, 92, 2127, 2331, 111, 2369, 2248, 2286, 145, 2081, 933, 667, 1004, 2169, 1455, 1443, 1896, 2007, 570, 2546, 1809, 1823, 148, 2587, 1982, 528, 2384, 348, 1003, 2395, 2192, 1580, 1277, 1846, 1822, 1001, 2126, 1935]. Rabe [2366]. Rabe-Hesketh [2366]. Rachid [1902, 364]. Radhakrishna [2310]. Radiation [370, 1394, 2173, 2325, 1441]. radiation-induced [2325]. radioimmunoassay [1291]. radioimmunoassays [868]. radiotherapy [2288]. Raghavarao [481]. raised [1429]. Random [630, 1042, 306, 2063, 864, 2529, 2455, 2340, 1222, 2287, 1817, 1946, 445, 1462, 399, 531, 296, 1524, 1545, 1371, 2304, 1927, 1401, 2538, 2326, 1928, 2342, 543, 2390, 54, 2454, 1930, 619, 678, 880, 1857, 1463, 1174, 736, 2231, 1830, 1331, 969, 1322, 2408]. random-digit [531]. random-effects [1222, 1463]. Randomization [1870, 2555, 1208, 500, 456, 1570, 1, 2079, 903, 804, 433]. Randomization-based [2555, 1570, 2079]. Randomized [428, 147, 1490, 66, 281, 323, 231, 2089, 825, 1609, 2398, 795, 1407, 2111, 1843, 2262, 2577, 2399, 2110, 2372, 1596, 2077, 379, 972, 2226, 2578, 1680, 2003, 2484, 683, 2198, 779, 1065, 2075, 2554, 1221, 660, 904, 1768, 959, 1051, 2432, 1346, 40, 1954, 2116, 2002, 390, 693, 2647]. randomly [505]. Randy [2133]. range [1287]. ranges [290, 269, 270, 518, 2161, 1161, 2233]. ranging [222]. rank [1107, 470, 2151, 1409, 824, 662, 1448, 2568]. rank-sum [470, 2568]. ranking [1858]. rankings [1806, 462]. ranks [2540]. Ranta [2004]. Rao [2310]. rare [2590, 1659, 1430, 2456]. Rasch [1850, 1887]. rate [820, 1542, 2359, 423, 867, 1914, 1345, 489, 2426, 2119, 2229, 839, 1358, 1420, 899, 2631, 1019, 1471]. raters [457, 129]. rates [2649, 979, 1762, 2590, 521, 1885, 2294, 1032, 1551, 1892, 648, 1820, 1986, 1104, 2120, 36, 951, 1732, 2605, 2473, 2557, 39, 1997, 1956, 1605, 1271, 1743, 1512, 1998, 2411, 200, 857, 1974, 2103, 1244, 212, 2171, 2172, 747, 1128, 1535, 1664, 1082, 775, 1018, 688, 1106, 1388, 1007, 2434, 11, 855, 1213, 1537, 356, 1013, 1660, 2431, 2373, 651]. Ratiff [656]. rating [1174]. ratio [1354, 520, 757, 216, 205, 2424, 277, 638, 515, 1946, 1450, 1915, 16, 2330, 1080, 1212, 138, 1921, 608, 1841, 509, 589, 839, 1358, 943, 1678, 1181, 1617, 1626]. rationale [2048]. ratios [940, 2128, 2089, 1840, 2129, 2611, 1609, 38, 950, 2499, 438, 135, 338]. Ratkowsky [945]. Raton [786, 2587]. Raven [346, 554, 145, 1166, 830]. Raymond [1813, 1877]. Re [1883, 1283, 1420, 2527, 1939]. re-analyses [2527]. re-estimation [1939]. Re-evaluating [1420]. Re-evaluation [1283]. Re-using [1883]. reach [1085]. reaction [1265]. reactions [1561, 290, 907]. reader [1154, 1171, 1486]. reading [442]. Realistic [1541, 2318]. reality [814]. really [377]. Reasoning [2507, 325]. reassessment [771, 1317, 2443, 2489, 1287]. recalculation [2000]. recall [1736]. recapture [2048, 584]. received [377]. Receiver [17, 1869, 2140, 1495, 1963]. receiving [2274]. receptivity [2199]. recessive [54]. Recognition [1984]. recollection [68]. recombinant [1035]. recommendation [1295]. recommendations [767, 1937]. recommended [1283]. Reconciling [1232]. record [2010, 1236, 233]. recording [1459]. Recovery [319]. Recurrent [2619, 2525, 1407, 1834, 1464, 2486, 2534, 2589, 1853, 1106, 1952, 1830]. recurring [1835]. red [2438]. Redesign [2346, 2347]. reduce [684, 2026]. reduced [2438]. Reducing [761, 1504, 2260]. Reduction [1111, 1794, 2262, 2062, 2327, 2292]. Rees [1860]. referees [1281]. Reference [2369, 2588, 750, 560, 2084, 311, 2642, 290, 269, 270, 518, 1341, 1705, 158, 2161, 51, 1161, 2233, 2024]. refinement [650]. Reflections [2408]. Reform [2252]. Refusal [686]. Regenstrief [266]. regimen [1910, 1069]. regimens [2073]. region [1958]. regional [1513, 1550, 2390, 1021, 1337, 1496]. regions [857]. register [930]. Registrar [362]. Registries [767, 459, 736]. registry [324, 2010, 1780, 1236, 651]. registry-based [2010]. Regression [2256, 1695, 2314, 2244, 967, 499, 2576, 2419, 768, 1036, 1790, 2228, 1830, 1944, 1872, 2167, 822, 1797, 821, 2502, 412, 876, 1647, 820, 2340, 607, 1222, 2287, 2156, 1499, 1760, 1773, 1794, 2163, 1517, 1379, 756, 1676, 180, 1977, 2465, 1915, 1289, 766, 661, 1988, 2268, 1020, 271, 733, 1095, 1735, 1022, 1369, 1402, 1413, 1839, 2319, 2201, 42, 2387, 1411, 340, 1764, 1333, 2570, 2377, 1033, 1765, 761, 1867, 1147, 181, 927, 773, 1928, 1353, 1412, 333, 1564, 1674, 2028, 1267, 1266, 2449, 2066, 784, 1224, 128, 1146, 2596, 973, 2255, 1268, 1777, 2427, 1759, 541]. regression [1042, 640, 774, 904, 1319, 137, 1606, 865, 1209, 1511, 2413, 1651, 1332, 1331, 1173, 558, 623, 274, 561, 1617, 1575, 1929, 1324, 2134, 326, 327, 1923, 2194]. regressions [272]. Regret [1560]. regulating [211]. regulation [1199]. regulatory [2207, 2526]. rehabilitation [347]. Reinecke [2132]. Reinhard [1861]. Reinhold [656]. reject [849]. Rejoinder [472, 194, 467, 937, 34]. relapse [901]. relapsing [962]. relapsing-remitting [962]. Related [827, 750, 510, 1803, 198, 1133, 2103, 347, 865, 1161, 1210, 2233, 1890, 2027]. Relation [180, 1898, 1912, 539, 919, 1882]. relationship [1357, 1751, 504, 2001, 2490, 1296, 16, 1247, 2570, 1074, 176, 1754, 1664, 2508, 558, 989]. Relationships [908, 821, 1669, 494, 130, 1239, 184]. Relative [50, 899, 2458, 876, 1842, 205, 397, 728, 888, 370, 13, 311, 1731, 174, 1431, 337, 2499, 2325, 2188, 918, 523, 2174, 853, 906, 2645, 1332, 558]. relatives [609]. relaxation [631]. Relevance [686]. relevant [713]. reliability [1635, 206, 1342, 207, 129, 2087, 2051]. reliable [885, 457]. Remarks [692, 720, 817, 1260, 1671, 1231, 2620]. Remembrances [97]. remission [901]. remitting [962]. REML [1090, 2159]. renal [1148]. renewal [334, 2338]. Reparametrization [1918]. reparametrizations [1648]. repeat [2198]. Repeated [65, 603, 932, 334, 864, 2472, 1836, 1746, 1265, 533, 1980, 414, 1730, 952, 2289, 1506, 963, 2029, 2160, 1093, 1737, 1775, 1769, 2524, 2387, 2428, 2304, 1804, 726, 1573, 2030, 2086, 1034, 1043, 2538, 1304, 237, 490, 1928, 1574, 964, 1905, 2464, 2307, 838, 2591, 2203, 2348, 2414, 1094, 641, 1612, 202, 2500, 1114, 426, 1796, 2639]. repeated-measures [1265]. repeated-series [1775]. repeatedly [1636]. repeats [65]. replacement [1660]. replication [516, 2468, 2646]. Reply [2146, 2191, 2249, 2003, 2142, 2005, 2144, 1949, 2129, 431, 1476, 2475, 1810, 2127, 769, 1873, 88, 2251, 1625, 2166, 131, 2238, 2634, 1875, 1424, 2355, 2344, 1626]. Report [2571, 565, 2045]. reported [2499, 342]. Reporting [2471, 833, 1494, 1979, 2294, 272, 1613, 2353, 2637, 1489]. reports [2224, 2613]. representation [468, 158]. representations [2389, 1881]. reprinted [2381]. reprocessing [1439]. reproductive [1211]. requirement [490]. requirements [644, 500, 2155, 79, 918]. Resampling [2629, 2026, 640, 1027]. Resampling-based [2629, 1027]. Research [432, 827, 2106, 2408, 362, 2097, 1148, 451, 1037, 1813, 1823, 2571, 704, 2547, 1981, 2315, 2193, 470, 261, 819, 2230, 1977, 494, 1325, 1011, 46, 2625, 2098, 1004, 947, 1455, 1367, 1690, 289, 1233, 946, 1643, 260, 738, 1586, 254, 347, 253, 893, 1001, 1811, 1798, 1359, 2277, 2131, 1314, 2265, 381, 596, 1014, 829, 570, 526, 284, 347, 345, 145, 164]. Researchers [2586]. resected [1055]. residential [1542]. Residual [1402, 426, 607, 295, 2176]. residuals [750]. resonance [1505, 2533]. resource [2230]. Resources [511, 243, 527]. respiratory [1417, 368, 1105]. responder [379, 2516]. Response [122, 119, 2552, 2286, 1970, 1357, 2288, 393, 370, 1751, 2001, 746, 531, 2580, 1249, 494, 653, 1683, 495, 2385, 567, 1905, 1038, 1755, 1756, 1084, 1777, 1957, 1759, 919, 1420, 501, 103, 1238, 162, 1422, 1734, 2413, 1191, 1239, 969, 1179, 1503, 1010, 2291, 10, 621]. Responses [1874, 864, 2317, 2163, 2398, 1795, 679, 296, 1506, 709, 2135, 1638, 1044, 1577, 2030, 1043, 964, 1267, 1616, 880, 1190]. restricted [1410, 1086]. Restriction [2260]. restrictions [2185]. resulting [1856]. Results [362, 761, 663, 822, 61, 793, 806, 9, 1941, 389, 1232, 1866]. resuscitation [2274]. retinopathy [1398, 1882, 1451]. retrospective [562, 745]. reversal [1054]. reversible [314, 908, 468]. Review [23, 1961, 1901, 614, 1985, 2584, 777, 434, 432, 656, 770, 758, 1087, 827, 1812, 2253, 510, 945, 597, 1938, 1934, 2252, 2310, 2462, 381, 363, 553, 1798, 2397, 345, 1826, 828, 2394, 1165, 1984, 1027, 21, 2090, 2067, 596, 2368, 1048, 1359, 2276, 1922, 90, 452, 2367, 840, 2312, 2055, 146, 730, 94, 366, 598, 403, 2644, 1983, 482, 480, 132, 2408, 326, 1902, 2092, 1876, 244, 613, 2131, 2314, 401, 871, 2602, 1887, 1014, 2039, 1026, 644, 384, 2009, 433, 91, 365, 731, 92, 1962, 2204, 2275, 2254, 2018, 364, 144, 1153, 2148, 346, 786, 402]. Review [1859, 1799, 206, 451, 2041, 892, 111, 1786, 2070, 554, 2545, 1911, 2396, 450, 2042, 759, 2369, 1878, 147, 2379, 20, 2038, 2134, 872, 2263, 1898, 2296, 2286, 145, 1192, 2081, 93, 2157, 1085, 1899, 2006, 2179, 1193, 110, 1000, 1994, 1888, 1970, 1862, 933, 511, 165, 2495, 328, 667, 1086, 1194, 944, 2573, 2378, 404, 2345, 829, 1923, 1004, 2169, 788, 1154, 2586, 327, 2278, 22, 1824, 1370, 1371, 1455, 1372, 1454, 1443, 2601, 2091, 1848, 1896, 2068, 841, 2419, 2599, 2007, 2194, 2132, 242, 1972, 2461, 2572, 570, 2544, 2546, 2313, 2082, 1924, 612, 1951, 2040]. Review [2417, 2418, 1847, 1933, 1037, 1278, 655, 112, 24, 946, 2380, 243, 1849, 481, 1813, 1823, 164, 1687, 1015, 1950, 148, 2587, 285, 2133, 2494, 1897, 2439, 1002, 400, 1973, 2311, 668, 870, 1787, 1055, 1038, 2585, 2150, 207, 2093, 1982, 1971, 921, 2043, 1860, 1900, 2264, 1838, 1850, 383, 1338, 1877, 2600, 2547, 1861, 2053, 2366, 528, 1108, 1701, 2383, 2170, 2384, 2181, 2168, 284, 1981, 325, 348, 787, 2019, 1003, 555, 166, 2395, 2192, 347, 43, 2149, 2239, 246, 1937, 527, 44, 1277, 382, 1825, 2381, 1630, 2069, 1912, 621, 830, 1025, 2056]. Review [2054, 1846, 1788, 1315, 1314, 1960, 2020, 2180, 893, 1001, 2193, 920, 2008, 184, 1164, 2265, 1811, 2178, 2382, 1704, 245, 1935, 2167, 398, 2317, 2001, 362, 654, 1567, 733, 924, 2520, 2537, 389, 1468, 1494, 2039]. Reviewers [2571]. reviewing [2299]. Reviews [2147, 2277, 362, 1936, 947, 2571, 526, 1166, 95, 247]. revised [1990, 2008]. revising [1300]. revisited [1648, 2216, 1758, 1036, 610, 183, 429]. rhabdomyosarcoma [502]. rhDNase [1952]. rheumatoid [776, 761, 2470]. rhinitis [2030]. rhythm [661]. rhythms [1702]. Richard [945, 362, 1923, 1037, 1897, 1108, 893]. Richardson [2169]. Ricketts [432]. Riedwyl [2419]. Riegelman [893]. Rienzo [1862]. right [489]. Rijckevorsel [1154]. ring [215, 998]. Ripley [1901]. Risk [2644, 2038, 839, 1358, 1332, 2437, 916, 876, 1842, 205, 1649, 398, 530, 1437, 2343, 397, 1429, 728, 370, 504, 35, 13, 2232, 1377, 2011, 2465, 952, 311, 203, 1615, 174, 396, 38, 2359, 1712, 639, 654, 575, 574, 218, 1711, 1620, 858, 230, 1431, 2241, 1376, 179, 337, 1997, 2325, 1921, 1211, 455, 1246, 1643, 1539, 1015, 2339, 918, 605, 2224, 755, 140, 1853, 2085, 2174, 446, 584, 853, 589, 2645, 1666, 1657, 1619, 1018, 899, 2017, 2045, 76, 11, 2031, 1882, 558, 2333]. risk-adjustment [2011]. Risk. [1372]. Risks [2038, 1483, 352, 1731, 1621, 38, 575, 2389, 2499, 1074, 1833, 2173, 734, 2138, 2420]. Risto [1787]. Robert [840, 2131, 2648, 164, 2633, 382, 893]. Robertson [92, 2278, 1824, 2473]. Robin [93]. Robust [1571, 1336, 762, 2318, 1478, 2482, 2301, 1615, 927, 1891, 1564, 2449, 868, 1682, 1749, 1871, 2494]. Robustness [1511, 1805]. ROC [17, 2140, 1963, 1623, 791, 1895, 1081, 1578, 1651, 1805, 1637]. Rødland [2429]. Roeback [1424]. Roger [788]. Role [1596, 1594, 2628, 673, 806, 2303, 1767, 735, 743, 1118, 1592, 15, 686]. role-play [686]. Rolf [2019, 2056]. Rom [1874]. Rome [20]. Ronald [2134, 841]. Rondel [1277]. room [1247, 1544]. Rose [892]. Rosner [89, 400]. Ross [2148]. Rosser [1934]. Rothman [243]. routinely [254, 2582]. Rovine [829]. Royal [2369, 1085]. rubella [106]. Rudinger [2180, 1001]. Rule [1858, 2398, 562, 442, 492]. Rule-based [1858]. rules [1070, 1071, 706, 493, 60]. run [376, 660, 2615, 1920]. run-in [376, 660]. Running [777]. Rupees [1823]. Ruppert [2007]. Rusanen [2005, 2004]. Russell [554, 1811]. Ryan [1838, 2167].

S [363, 1027, 2343, 2276, 2089, 2638, 326, 362, 2314, 91, 2204, 1786, 1936, 2369, 1193, 1086, 2573, 2169, 2639, 1443, 242, 1972, 2418, 668, 2585, 2043, 2474, 1338, 2366, 2170, 2168, 2646, 2637, 1580, 1277, 1825, 2069, 2054, 1846, 1788, 2333, 1901]. S-L [2637]. S-Plus [2314, 1901]. S. [369]. Sackett [667]. sacrifice [983]. Safety [2627, 537, 1308, 211, 1310, 2224, 263]. SAGE [2090, 2277, 2131, 1899, 1923, 1896, 2194, 2461, 1900, 2019, 2239, 1788, 1811]. Saharan [210]. Salle [325]. Salsburg [1086]. Salvan [2380]. same [2198]. Sample [1423, 2285, 728, 2317, 1178, 195, 500, 2155, 766, 867, 2330, 1, 1923, 2202, 567, 1043, 178, 237, 490, 1528, 996, 302, 918, 1403, 1895, 2454, 1784, 2329, 447, 1330, 2051, 897, 540, 2412, 953, 1483, 1717, 2000, 1175, 1806, 2482, 1393, 609, 2230, 737, 2463, 2175, 580, 79, 1731, 2392, 1999, 174, 1506, 786, 917, 590, 1188, 1113, 1300, 2468, 416, 165, 538, 2201, 1818, 999, 581, 550, 544, 440, 950, 1044, 1409, 1577, 539, 765, 1383, 1921, 1274, 2606, 1162, 177, 608, 476, 543, 438, 988, 2203, 1356, 729, 1468, 2645, 1939, 2025, 40, 2124, 1750, 274, 1092, 1870, 1568]. Sample [2368, 481]. Sample-selected [1923]. sampled [1638]. sampler [2321, 775]. Samples [2265, 1797, 1187, 836, 585, 1956, 1893, 177, 270]. Sampling [413, 2243, 1082, 1354, 2424, 2515, 1240, 1636, 1404, 1528, 1921, 1905, 968, 1523, 1761, 866, 1779, 1763, 1626]. Samprit [2054]. Samuel [266, 1362]. San [1846, 1990]. Sandvik [2429]. Santos [1370]. sarcoma [2151]. Saskatchewan [1459]. Satellite [288]. Saulie [2067]. Saunders [598, 382]. scale [425, 1320, 1214, 1425, 1401, 988, 854]. Scales [2147, 994, 366, 1218, 1449, 1567, 2087, 1174, 343]. scan [852, 1534, 1546, 853, 851]. Scarecrow [1004]. Schach [1826]. Schelp [890]. scheme [1159, 2292]. schemes [1858, 2543]. schizophrenic [1265]. Schlattmann [890]. Schmitt [1454]. Schoenfeld [2065, 2164]. Schoenfelder [346]. School [2571]. Schwartz [2636]. Schweizerhalle [368]. Science [2252, 2368, 2277, 1640, 402, 1824, 2544, 1850, 2395, 155, 667, 2526]. Sciences [1902, 1876, 2204, 2042, 1899, 1923, 1788, 1518, 2312, 2018, 1936, 1888, 1278, 2587, 2239, 482, 1454, 166]. Scientific [717, 1281, 2310, 1798, 2602, 451, 1193, 2601, 2380, 1813, 284, 893, 2193, 831, 402, 944, 347, 1245, 1798, 2397, 325]. Scientist [1015]. sclerosing [396]. sclerosis [962, 2471]. score [520, 757, 35, 2262, 2350, 1565, 1554, 918, 1224, 2228, 2273]. scores [723, 2303, 765, 140, 1332]. Scotland [1130]. Scott [365, 2586, 2194]. Scratch [2041, 2439]. screen [985, 1293, 1945, 1465, 450]. screen-detected [1465]. Screening [1123, 770, 2284, 1351, 1892, 107, 972, 2258, 1975, 2565, 2569, 1694, 446, 137, 1408, 780, 2592, 2626, 770, 553]. search [1460, 2565]. Searching [510]. Seasonal [1124, 821, 2030]. seasonality [1803]. Secic [2571]. Second [1792, 148, 412, 2147, 2572, 1480]. second-order [1480]. Second-stage [1792]. secondary [1082, 2045]. sectional [1177, 917, 1188, 1474, 1162]. secular [1139]. Seeing [2040]. Segmented [1764]. segregation [54]. select [1224]. Selected [685, 348, 1475, 1923, 1968]. Selecting [919, 1697]. Selection [971, 2304, 2453, 399, 1449, 2230, 1340, 2625, 327, 2013, 2606, 273, 135, 959, 1778, 1855]. Self [2196, 2042, 1401, 452, 1499]. Self-designing [2196]. self-instruction [452]. Self-learning [2042, 1499]. Self-modelling [1401]. Selvin [1799, 1025]. Selwyn [2587]. semen [1381]. Semi [414, 2269, 2352, 2649, 576, 1648, 1986, 1953, 443, 1764, 2290, 1772, 1759]. semi-Bayes [443]. Semi-parametric [414, 2269, 2352, 2649, 1986, 1953, 1764, 2290, 1772, 1759]. semi-proportional [576, 1648]. semiannual [565]. Semmelweis [2022]. Sempos [246]. Senn [2263, 946]. Senoussi [1902, 364]. sense [112, 1192]. sensitivities [1383]. Sensitivity [579, 2259, 2128, 1840, 2129, 2304, 1940, 1306, 140, 1079, 1945, 1331, 1439]. sensory [1969]. Separation [1174]. Sepsis [2125]. September [59, 153, 388, 288, 1169, 1170]. sequel [1542]. sequence [1844, 1346]. sequences [2416, 1703]. Sequencing [2276]. Sequential [735, 60, 1058, 1745, 1714, 2477, 1783, 871, 157, 89, 1392, 2524, 173, 577, 567, 1804, 2556, 924, 649, 1419, 2480, 870, 564, 1969, 2, 1221, 2049, 7, 1678, 2643, 1220, 2033, 1750, 2523, 1059]. Sequentially [1692]. serial [534, 650, 2352, 2176]. Series [1985, 1640, 821, 962, 2503, 362, 395, 2042, 1093, 1775, 368, 1899, 1923, 2504, 1233, 1215, 238, 526, 1793, 1079, 1041, 643, 369, 2032, 869, 245]. serious [2095]. seroconversion [486, 929, 1131, 48, 663]. seroconverters [879]. serologic [1777, 1957]. serological [1622, 2468]. seroprevalence [1117]. seroreversion [1957]. serum [2387]. service [2341]. Services [2408, 2597, 1977]. Session [288, 628, 624, 632]. set [2576, 2537]. sets [1885, 1615, 2369, 1924]. setting [1364, 2618, 2162]. settings [1978, 2354]. seven [2648, 2122]. Seventeenth [2204]. several [1219, 515, 1149, 1445, 2117, 2350, 1080, 913, 263, 2409, 2088, 312, 2016, 441, 2373]. severity [545, 1464, 1554, 2361]. sexual [1712, 2243, 1485, 674, 1779]. sexually [2243, 584]. SFR [2419]. sfr134 [1166]. Shafer [2275]. shall [2356]. shape [1485, 1401, 2091]. Shapiro [665]. shared [1677]. Sharp [2333]. Shein [872]. Shein-Chung [872]. shelf [1857]. shelf-lives [1857]. SHEP [1806]. shift [1401, 2568]. shifts [678]. Shih [2143]. Shihadeh [1788]. Shonkwiler [2081]. Short [95, 247, 574, 1124, 1990, 606, 358, 330, 1082]. short-latency [1082]. short-term [1124, 606, 330]. shortcomings [2206]. Shott [598]. should [115, 118, 125, 124, 920]. Shrink [2006]. Shukla [1823]. Shuster [786, 1994]. sickness [1538]. side [82]. sided [1644, 1151, 1273, 1383, 2122, 2162, 1750, 2648]. Siegfried [1826]. sign [1224]. Signals [1862]. Significance [1161, 72, 580, 2549, 2550, 83, 1086, 2551, 999, 726, 995, 1909, 49, 4, 519, 2245]. significance-testing [580, 999]. significant [2162]. Signs [2125]. Silva [1370]. SIMEX [2576]. Similar [1874, 1616]. Similarity [1899]. Simon [1994, 316]. Simonoff [2170, 2054]. Simple [1615, 336, 2499, 2030, 2581, 1388, 1096, 277, 2163, 1976, 958, 2201, 726, 773, 128, 2447, 1777, 509, 498, 1021, 1337, 1496, 2456, 1637]. Simplified [2024]. simulated [1908, 1866]. Simulation [1575, 2611, 771, 1697, 2234, 1187, 2593, 1694, 1419, 2049, 1494, 222, 1094, 2228, 2592, 1213, 796]. simulation-based [1694, 2592]. simulations [1541]. simulator [81]. Simultaneous [2605, 1116, 2377]. Simultaneously [1636, 352]. since [2437, 939, 522]. Singapore [2310, 2602, 2601, 2380, 2585]. single [987, 2488, 205, 983, 2248, 1879, 2648, 2122, 1220]. single-arm [1220]. single-sacrifice [983]. Sir [450, 739]. site [510]. sites [773, 454]. situation [2574, 2575]. situations [888]. Siu [2165]. six [1218]. Sixteen [539]. Size [2368, 1870, 1423, 2285, 2000, 1175, 728, 2317, 1393, 2433, 609, 195, 2463, 2175, 1864, 580, 500, 2155, 79, 2392, 174, 786, 917, 590, 766, 109, 1113, 1300, 2330, 416, 1, 165, 538, 2201, 1545, 1818, 2202, 567, 1355, 999, 581, 550, 440, 1634, 1044, 539, 765, 1043, 1383, 178, 490, 1528, 1921, 1162, 659, 996, 968, 481, 608, 476, 918, 1403, 1895, 2454, 2452, 273, 2203, 1356, 729, 1341, 584, 1624, 2645, 2329, 1939, 2025, 2429, 1330, 1805, 2051, 897, 953]. Sizes [1784, 1483, 1178, 737, 1188, 867, 1568, 237, 302, 1468, 2348, 2329, 447, 2025, 540]. Sizing [1974]. skewed [2024]. skewness [458]. skill [686]. Skrabanck [144]. SLACK [347]. sleep [1945]. Sloan [2006]. slopes [1511]. Small [2091, 1198, 438, 2265, 1483, 1717, 2502, 2246, 1542, 2294, 1662, 737, 1032, 1347, 2052, 1053, 1658, 1956, 2606, 1659, 2390, 1065, 1003, 2532, 274, 1665]. small-area [2294, 1003]. smears [884]. Smith [2573, 2068, 1037, 2333]. smoke [1522, 1239]. Smoking [1897, 2558, 1521, 37, 1522, 824, 1906]. Smoller [328]. Smooth [1963, 1989, 1321]. smoother [2161]. Smoothing [1323, 560, 1732, 1012, 2242, 1551, 2605, 2509, 1882, 2170]. Snashall [510]. Sobel [1888]. Social [828, 2277, 2131, 1899, 1888, 1923, 2547, 2239, 1788, 141, 108]. Societies [132]. Society [59, 153, 388, 896, 1169, 1170, 1458, 2369, 385]. Software [1258, 2369, 1556, 1557, 2278, 860, 1555]. sojourn [1351]. solid [2173, 2325]. solutions [2098]. Solving [846]. Some [2519, 1362, 300, 2624, 1317, 100, 169, 990, 204, 544, 1523, 718, 1645, 1793, 2255, 1766, 1141, 1996, 1744, 2214, 1063, 2197, 2541, 1757, 607, 1476, 422, 625, 2520, 791, 473, 257, 1461, 162, 1484, 2456]. Sons [482]. Sort [2038]. Sorting [1554]. sounds [136]. source [1429, 2465, 1440, 1522, 2048, 342, 2017, 2333]. Sources [1743, 981, 1536, 1246, 2339, 1666, 1004]. South [402]. Southwest [1062, 697, 2107, 2116]. space [1437, 1548, 1667, 2241, 1434, 1733, 2005, 2004]. space-time [1437, 1548, 1434, 1733, 2005, 2004]. spaced [474, 678]. Spain [2508]. span [625]. Spanish [1494]. sparse [1842, 502, 899]. Spatial [1269, 968, 1538, 2509, 1441, 1661, 1435, 1534, 1431, 1881, 200, 1550, 848, 2390, 2608, 1955, 1535, 1664, 1555, 1549, 855, 790, 1021, 1337, 1496]. Spatio [2240]. Spatio-temporal [2240]. Speakers [799, 617]. Special [1252, 1037, 1251, 1270]. specialists [1204]. specific [979, 1762, 2232, 1567, 1711, 1620, 507, 1162, 2139, 2411, 2352, 1128, 269, 2161, 1734, 1007, 2024, 10]. specifications [1420]. specificity [2259, 2128, 1840, 2129, 579, 1945]. spect [394]. spectrum [1918]. speech [136]. Speed [2276]. spending [1056, 1057, 173, 887, 1645]. Spiegelhalter [2169]. Spilker [346, 830]. spinal [545]. spline [1735, 1410, 1022, 1401, 2307, 1882]. splines [1323, 1513]. sponsor [2527, 702]. Sponsored [308, 2097, 2414]. spontaneous [218, 1613]. spread [2243, 310]. Sprent [94, 2599]. Spriet [1000]. Spring [1443]. Springer [1901, 2276, 403, 480, 2092, 362, 2314, 1887, 384, 91, 2042, 2286, 2179, 328, 1086, 404, 2091, 2544, 2546, 2417, 2418, 383, 2053, 1108, 2170, 555, 2395, 2020]. Springer-Verlag [1901, 403, 480, 2314, 1887, 384, 91, 2179, 328, 1086, 404, 2091, 2417, 2418, 383, 1108, 555, 2020]. SPSS [2418, 2547, 1554, 2585]. SPSS-PC [1554]. spurious [1736, 857]. Square [2008, 479, 1159, 330]. Squared [1972, 539]. squares [1792, 2449, 1692, 913]. Srivastava [1823]. St [655]. Stability [1640, 449, 448, 1001]. Staff [1228, 1654, 671]. Staffan [2254, 511]. stage [1762, 1760, 2014, 1101, 276, 240, 1792, 2444, 724, 507, 1645, 1761, 1945, 1763, 2376, 1889, 503, 866]. stage-specific [507]. Stallard [23]. Stamenovic [2164]. Standard [2586, 2458, 2638, 2165, 1993, 293, 1510, 279, 1932, 2375, 2376]. standardization [1380, 648]. Standardized [1411, 1106, 1820, 315, 1017, 478, 1388]. standards [689, 1210]. Stangl [1911]. Stanley [1027, 165, 1872]. Stanton [1153, 2545, 2264]. Stapleton [1847]. starting [2614, 1072]. Stat [2296]. Stat-Ease [2296]. State [827, 2531, 164, 2186, 2338, 993, 1398, 1236, 15, 845, 2470]. States [2252, 596, 2058, 2608, 2060, 444, 474, 1124, 1803, 573, 836, 601]. statics [656]. statistic [1219, 1115, 580, 852, 1534, 1418, 1355, 999, 440, 927, 1546, 1566, 853, 851]. Statistical [334, 2280, 2299, 288, 2223, 2090, 1359, 1646, 680, 1844, 19, 1640, 2204, 439, 1878, 1091, 1192, 636, 1888, 1557, 2091, 1263, 2599, 1505, 2533, 179, 2544, 1199, 2206, 2496, 1611, 1847, 1171, 2380, 1473, 1893, 1310, 78, 1973, 127, 808, 212, 1705, 2384, 2168, 2149, 1630, 2108, 1540, 1677, 1489, 25, 435, 1203, 1699, 21, 409, 94, 966, 1218, 871, 384, 753, 2283, 1639, 1650, 206, 2044, 578, 408, 862, 169, 292, 2356, 2417, 2418, 2526, 1306, 1687, 2183, 1801, 2494, 2528, 258, 207, 1141, 860, 2209, 1543, 674, 1744, 407, 621, 1191, 225, 159, 2532, 790]. statistical [1828, 419, 356, 317, 2360, 2273, 2027, 288, 891, 614, 2131, 2314, 1026, 759, 2263, 1194, 242, 1924, 1849, 1813, 1002, 526, 1861, 1108, 348, 555, 166, 246, 1025, 1788, 893]. statistician [1205, 2627]. statisticians [2210, 2266, 383]. Statistics [1901, 410, 1275, 1516, 2365, 2462, 146, 1640, 125, 1936, 2396, 2042, 2495, 2278, 1896, 2572, 2526, 1860, 2019, 2239, 2265, 115, 381, 1836, 470, 482, 118, 1662, 401, 293, 283, 1863, 174, 421, 603, 2029, 856, 1548, 829, 1154, 2639, 2416, 22, 2484, 355, 121, 285, 2316, 2645, 932, 1204, 84, 2054, 1164, 2506, 2649, 2641, 2128, 2343, 2089, 2638, 2129, 2332, 2636, 1476, 1871, 2130, 2644, 2145, 2393, 2165, 2127, 283, 89, 2065, 2331, 2248, 2642, 891, 1870, 1568, 2639, 2648, 2473, 2003, 1625, 2647, 1809, 2237, 2164, 1948, 2066, 890, 2571, 2005, 2474, 2190, 2645]. Statistics [880, 1947, 1874, 2646, 2429, 1869, 2637, 2640, 2004, 2141, 2633, 2635, 2643, 1580, 2250, 2143, 1822, 2002, 1626, 2126, 1872, 2333, 2354, 2070, 2040, 1933, 148, 1971, 1877, 2170, 1087, 2055, 111, 944, 2310, 363, 21, 598, 912, 2559, 1859, 2041, 2369, 1454, 2546, 2439, 2560, 787, 2008, 1935]. Statistique [1902, 364]. Stats [1876]. status [2590, 1392, 2058, 2364]. stay [2257, 2617]. steady [474]. Stearns [2089]. Steed [2585]. Steenland [1315]. Stefanski [2007]. Steffen [2584]. Stein [347]. Step [299, 1149, 9, 1410, 87, 1741]. Step-down [299, 1149]. Stephen [1014, 2263, 165, 946]. stepwise [1107, 1572, 272, 662, 2162]. Stereology [2601]. sterilization [685]. Stern [2043, 1003]. Steve [1799, 2278, 1025]. Steven [1900]. Stevens [255]. Stevens-Johnson [255]. Stewart [553]. stillbirths [108]. stimation [931]. stimulating [2534]. Stine [2131]. Stochastic [444, 980, 358, 1737, 631, 2602, 404]. Stockholm [435]. stoke [111]. Stone [2070, 2339, 1666]. Stopping [1053, 749, 1069, 1070, 1071, 706, 1062, 1073, 493, 1940, 60, 1072, 492, 1051, 1978, 2354]. strata [177]. strata-matching [177]. strategic [2432]. strategies [25, 2247, 793, 1723, 958, 534, 673, 672, 2226, 2430, 1285, 1497, 1938]. strategy [1926, 596, 892, 836, 1391, 660, 780, 2197]. Stratification [1115, 577, 1841]. stratified [361, 500, 276, 476, 918, 1403, 85, 282, 2374, 589, 899, 1978, 487, 2354]. stratum [177, 608]. stratum-matched [608]. Streiner [2147, 366]. strength [1747]. Strengthening [473]. Strike [614, 1973]. Strom [2082]. Strömberg [2354]. Structural [2461, 2395, 2151, 1390, 926, 2533, 824, 2048]. Structure [1898, 1912, 1913, 751, 752, 2272]. Structured [1849, 1547]. structures [1635]. structuring [526]. STS [2376]. Student [1026, 2278, 400]. students [125]. Studies [2644, 1870, 1838, 1315, 940, 1717, 2455, 2223, 1842, 1096, 1836, 1736, 477, 228, 398, 1175, 161, 723, 2010, 62, 525, 1773, 1845, 2256, 1099, 1913, 2558, 2177, 413, 52, 223, 2096, 2015, 737, 2175, 1045, 79, 2399, 971, 684, 1854, 160, 107, 38, 276, 206, 318, 664, 760, 2469, 1709, 1188, 336, 654, 1729, 759, 1317, 1214, 548, 1216, 1738, 2324, 1425, 2443, 416, 933, 1, 1620, 2135, 1488, 980, 2467, 169, 2416, 2385, 143, 1328, 2207, 292, 581, 550, 1030, 2531, 337, 1883, 39, 2364, 1833, 2356, 1043, 354, 517, 178, 313, 1329, 2279, 683, 1808]. studies [302, 197, 791, 1266, 177, 608, 476, 918, 2066, 479, 304, 1224, 2224, 389, 1895, 1232, 2454, 1632, 207, 928, 263, 2591, 357, 837, 1268, 747, 2474, 729, 1624, 312, 589, 2049, 906, 1758, 1761, 1490, 905, 1666, 1657, 2429, 1619, 103, 835, 1238, 1978, 1604, 1484, 1422, 2434, 1239, 225, 1917, 1750, 976, 1540, 2051, 1213, 558, 745, 981, 1745, 1617, 2631, 1503, 2574, 2575, 540, 796, 989, 1470, 503, 2373, 2632, 2354, 1961, 2368, 2020, 597, 872, 165, 947, 1003]. Study [2102, 2125, 577, 1860, 1253, 1647, 1438, 2597, 1762, 370, 793, 261, 840, 413, 771, 2433, 362, 1435, 952, 2392, 1157, 1697, 105, 2234, 1352, 1187, 17, 9, 443, 733, 2593, 2387, 1607, 1544, 2578, 376, 2046, 1522, 1573, 2189, 1574, 1941, 1287, 1419, 389, 1602, 913, 1532, 139, 2229, 1494, 222, 2457, 1094, 1463, 310, 2513, 2615, 1213, 419, 1575, 663, 2431, 13, 1375, 180, 1377, 1328, 1376, 2403]. studying [724, 973, 1660]. Stuttgart [1826, 1962, 2056, 2180]. style [362]. sub [210]. sub-Saharan [210]. Subacute [396]. Subcloning [429]. subgroups [1513, 1460]. subject [781, 1633, 340, 1891, 1152, 1322]. subjective [1806, 462]. subjects [742, 1041, 163]. subsampling [161]. subset [1926, 420, 327]. success [1381, 820, 2207, 651]. succession [652]. such [2456]. sucking [2083]. sudden [1019]. Suggestions [2112]. sum [470, 2013, 2449, 2328, 2568]. summarizing [72, 723, 1313, 883, 734]. Summary [1863, 709, 1115, 603, 2029, 1596, 2639, 687, 901, 2484, 2363, 2538, 650, 791, 2316, 137, 932]. Summit [1786, 1338]. Sumpf [1850]. sums [236]. superficial [2498]. superiority [1993]. superiority/equivalence [1993]. supplement [362]. suppression [127]. suppressor [995]. surface [1599, 1683, 495]. surgery [1077, 1055, 779]. Surgical [2104, 2328]. Surrogate [1011, 1066, 1453, 2280, 1450, 1943, 2331, 743, 1894, 1118, 1010, 1067]. Surveillance [1536, 1955, 2615, 875, 578, 1709, 874, 2059, 585, 2630, 649, 907, 601, 2511, 1245, 2613, 43]. survey [1248, 1218, 1988, 1622, 696, 1474, 950, 2626, 1779, 2412]. surveys [531, 1622, 1443, 487, 1787]. Survie [1902, 364]. Survival [1961, 412, 827, 2367, 14, 182, 1680, 2403, 1040, 1314, 822, 1382, 1423, 2285, 314, 1696, 2455, 994, 2156, 2482, 1845, 1751, 1946, 521, 2558, 1695, 756, 1378, 399, 1375, 613, 2362, 1399, 311, 1576, 50, 1189, 1697, 1492, 2435, 1636, 468, 1103, 624, 2065, 109, 1908, 1059, 156, 2160, 1466, 2543, 1681, 1446, 1914, 1202, 606, 1474, 2467, 972, 1447, 1482, 1785, 143, 577, 901, 37, 47, 2158, 2185, 77, 2227, 1002, 2509, 2316, 232, 1343, 1776, 2497, 906, 1675, 762, 137, 2459, 835, 275, 1749, 900, 676, 623, 1058, 1714, 1465, 2336, 1944, 1013, 540, 796]. survival [2159, 1324, 1889, 1471, 953, 2388, 1400, 2003, 2002, 827, 1812, 2092, 2148, 2042]. survival-based [399]. survival-data [1202]. survive [1205]. survivor [2542]. survivors [2325, 2148]. Survivorship [597]. Susan [598]. susceptibility [1721, 2153, 2327, 928, 1019]. susceptible [1350]. Susie [553]. suspected [1542]. suspension [1940, 1306]. Svensson [1371]. SW7 [402]. Swanson [401]. Sweden [435, 132, 1534, 1184]. Swedish [1371]. sweeps [1392]. Swerdlow [1370]. Swinscow [2008]. Swiss [2106]. switch [1680, 2003, 2002]. Switzerland [2055, 2419]. Sylvia [328]. Sylvie [2314]. Symmetry [136]. Symposium [25, 671, 616, 1672, 2072]. symptomatic [358]. symptoms [1105]. syndrome [1892, 649, 446, 255]. synergy [2174]. synthesis [596, 1627, 1643, 1586, 1811]. Synthetic [858]. syphilis [1777, 1957]. System [2615, 1234, 849, 1556, 233, 1916, 184, 2613]. systematic [1371, 1174, 2039]. systemic [1293, 2027]. systems [1669, 673, 2557, 907, 1955, 561]. systolic [1806].

T [432, 2638, 2367, 2055, 480, 362, 2127, 2331, 891, 1193, 1809, 2417, 1038, 2150, 1166, 528, 2181, 246, 1580, 2008, 2382, 1626, 2126, 2333, 992]. t-test [992]. table [2248, 1879, 306]. Tables [2368, 1784, 1163, 491, 205, 2424, 2270, 437, 373, 1615, 915, 28, 1080, 1390, 2504, 438, 508, 1682, 1079, 2409, 1468, 137, 519, 736, 1630, 668]. Tacq [2277]. tail [852]. Takashi [1015]. tale [1919, 2250]. Talking [1164]. Tango [850]. Tarragon [144]. taught [118]. Tautu [480]. Teach [2571, 115, 118]. Teaching [1204, 283, 121, 920]. technique [940, 1313, 883, 1600]. Techniques [2277, 2092, 1970, 2132, 1788, 1234, 370, 2611, 14, 2242, 1885, 1340, 985, 1171, 1881, 1694, 333, 1523, 304, 2497, 541, 2026, 1954, 2592]. technologies [226]. telephone [531]. telephone-prefix [531]. temporal [1538, 605, 1555, 2240]. Ten [1958, 1580, 912]. Ten-year [1958]. Tens [1994]. Tenth [59, 153]. teratologic [1045]. teratology [2028]. term [1124, 198, 2148, 574, 606, 524, 77, 358, 330, 1715, 1749, 1889]. terminating [1834, 1835]. termination [170]. Terminology [1937]. terms [2579, 926]. Terri [1165]. Terry [2276]. Test [449, 174, 448, 1909, 1381, 1797, 1717, 520, 757, 2588, 2259, 61, 102, 305, 277, 1429, 2482, 2398, 2433, 1225, 2463, 1347, 415, 1526, 2468, 2642, 1525, 321, 538, 1667, 1394, 581, 2258, 550, 1565, 2257, 927, 2185, 1383, 1528, 1921, 1127, 235, 2339, 791, 918, 54, 1930, 85, 282, 1065, 1130, 473, 1084, 2448, 665, 509, 992, 1807, 1604, 2124, 1964, 2568, 2567, 1021, 1337, 1496, 487, 2631, 663, 2645]. tested [959]. Testing [2455, 1561, 324, 2247, 1363, 1146, 175, 1616, 1784, 1100, 1865, 2450, 977, 869, 1920, 2259, 294, 2270, 1995, 244, 373, 737, 580, 1045, 1993, 588, 733, 2422, 83, 858, 1628, 999, 726, 924, 1512, 2480, 1159, 49, 1543, 1468, 4, 1082, 447, 519, 2162, 2184, 1874, 2143, 1972, 1027]. Tests [763, 1868, 1432, 1434, 2180, 2588, 2455, 1219, 666, 1783, 888, 1572, 515, 966, 1241, 2177, 1579, 2145, 1149, 2001, 1115, 1129, 299, 1644, 89, 1506, 2350, 1318, 2234, 1188, 1113, 336, 271, 1884, 2642, 1623, 1080, 1839, 1086, 724, 1488, 2258, 544, 1975, 1409, 2363, 1867, 1273, 1274, 1851, 1801, 1541, 438, 782, 988, 1682, 2409, 1356, 2474, 2, 2190, 835, 1678, 40, 1448, 1430, 2016, 2162, 1123, 1750, 1161, 854, 1428, 441, 2291, 1963, 2130]. text [1499, 1454, 2042]. textbook [788, 2150]. Textbooks [1640]. Thall [1848]. Their [1898, 1912, 2147, 366, 2602, 1892, 402, 258, 253]. them [2385, 2226]. theorem [935]. theoretic [493]. theoretical [738]. Theory [477, 228, 2091, 1933, 1412, 1178, 36, 1686, 1658, 1772, 1408, 2523, 2286, 1970, 404]. Therapeutic [2252, 2209, 234, 1507, 263, 1356, 2073]. therapeutically [1289]. therapies [1709]. Therapy [704, 1646, 2230, 1863, 2109, 262, 724, 601]. Therese [1000]. Thierry [1902, 364, 147]. thinking [2533, 1500]. Third [1153, 400, 2381]. Thirteenth [896]. Thomas [656, 613, 2165, 20, 1994, 1896, 655, 2133, 2494, 2571, 2167]. Thompson [841, 24, 2333]. Thorofare [347]. thoughts [2308]. Thousand [2131, 1899, 1923, 1896, 1900, 1788, 2265]. Thousands [1994]. Three [352, 2641, 979, 2014, 1844, 1101, 1692, 350, 1055, 2400, 1479, 518, 2640, 1741]. three- [2641, 2400, 2640]. three-parameter [518]. three-period [1844, 1479]. three-squares [1692]. three-stage [2014, 1101]. three-step [1741]. Threshold [1700, 2466, 823, 225, 369]. threshold-crossing [823]. thresholds [360, 1805]. Thriyambakam [2157]. Thun [1897]. thyroid [931]. Tibshirani [528]. Tierexperimente [1962]. Tiku [1850]. Time [962, 1985, 70, 1849, 747, 821, 2437, 777, 876, 2512, 1437, 393, 994, 2320, 525, 2246, 1751, 2242, 2503, 2558, 486, 929, 180, 2421, 939, 1351, 468, 1837, 318, 109, 1953, 1131, 1093, 2293, 2271, 1735, 368, 1410, 1022, 1770, 1620, 1831, 681, 1548, 1667, 2151, 1418, 972, 1767, 1727, 2241, 182, 1030, 1434, 1367, 1147, 2173, 2325, 2164, 824, 1127, 1215, 522, 1386, 2352, 734, 1793, 1733, 2005, 850, 2335, 626, 269, 158, 2459, 2414, 1604, 1790, 2004, 1134, 643, 2045, 1791, 2619, 369, 623, 1828, 745, 981, 1010, 1465, 2336, 1944, 2159, 526, 245]. time-by-covariate [1022]. time-changing [2045]. time-dependent [876, 2512, 2242, 1735, 1410, 1620, 681]. Time-period [747]. time-points [2414]. time-specific [269]. Time-Structured [1849]. time-to-event [2293, 1831, 745]. time-varying [1030, 1386]. times [1265, 2482, 1576, 652, 2435, 1131, 2271, 1769, 1212, 1447, 1482, 1785, 1727, 2213, 2480, 48, 1058, 2410, 1059]. Timing [2402, 1969, 1055]. Timmereck [2133]. tissue [968]. titration [223]. TNO [1154]. tobacco [1522, 1538, 686, 1239]. Todd [2384]. Tohen [2150]. Tokyo [1015]. Tolerance [1821]. Tolley [1924]. Tomatis [451]. tool [2026]. Toolbox [1037]. Toolkit [1973]. Tools [2314, 731, 1433]. Topics [827]. Torestad [1001]. total [372]. toxicity [1157, 637, 725, 130, 1311]. toxicokinetics [1298]. toxicology [555]. trachomatis [2145, 1884, 2190]. Tracking [879, 1712, 546, 1602, 882]. tracts [141]. Traffic [1438]. trails [281, 1628]. Training [1037, 2500]. traits [2583]. trajectory [313]. transformation [1107, 1915, 1270, 662]. transformations [271, 990]. transfusion [1133]. transfusion-related [1133]. transfusions [642]. Transient [396, 1353, 2288, 232, 1677]. transition [210, 2426, 775, 2470]. transitional [2186, 963]. transmission [1699, 2402, 1698, 1335, 1914, 484, 1137]. transmitted [2243, 584]. transplant [275]. transplantation [1903, 2271, 1180, 901, 2361, 775]. transplants [128]. Trapp [382]. treated [36, 1092]. treating [2454]. Treatment [1942, 1966, 2466, 2477, 888, 2320, 1806, 730, 1751, 2628, 1476, 1149, 763, 2262, 1730, 195, 1993, 1723, 958, 2331, 240, 2029, 661, 2428, 2215, 2416, 170, 1030, 2484, 377, 1679, 1894, 2453, 2486, 742, 1479, 1777, 1957, 1782, 80, 2229, 1461, 1865, 2643, 2016, 2143, 1285, 1497, 1489, 2076, 2555, 1010, 1083, 10]. treatment-by-centre [763]. Treatment-patient [1942]. treatment-time [2320]. treatments [1926, 2488, 515, 264, 1445, 795, 299, 1993, 1723, 1151, 749, 182, 1967, 1680, 2003, 1403, 2516, 350, 2349, 1782, 959, 1051, 1346, 780, 2016, 1311, 621, 1191, 2002]. tree [904, 1890, 1470, 1899]. tree-based [1470]. Trees [882, 1481, 2391]. trend [1803, 1225, 773, 85, 282]. Trends [1529, 1271, 1649, 638, 2059, 682, 2139, 678, 943, 1443]. Trent [1958]. Trial [767, 1871, 2104, 1848, 2195, 777, 2488, 231, 2089, 2317, 1571, 1609, 764, 795, 2062, 420, 2577, 535, 296, 157, 1863, 1151, 1148, 1837, 421, 806, 1053, 1300, 2216, 1681, 1446, 2444, 1421, 1607, 2215, 351, 841, 1110, 1940, 2556, 234, 2030, 2153, 2164, 2486, 964, 2534, 2227, 1419, 604, 2561, 705, 127, 779, 1065, 1592, 2527, 1221, 2497, 660, 904, 880, 1408, 519, 2078, 2432, 7, 1054, 40, 2116, 2228, 1321, 2555, 2376, 2498, 8, 231, 777]. Trials [1961, 2096, 147, 946, 2103, 1784, 1937, 280, 2285, 1797, 1158, 2472, 2466, 323, 1366, 298, 2300, 2152, 61, 294, 2641, 1746, 1594, 102, 305, 1515, 2095, 2106, 2102, 800, 1208, 1858, 2320, 638, 14, 2246, 680, 1090, 1178, 706, 2628, 101, 825, 1476, 1023, 1588, 1111, 960, 2014, 2267, 198, 771, 319, 1462, 2442, 352, 2398, 1407, 2111, 2114, 1843, 188, 414, 195, 2393, 1584, 644, 1864, 2580, 456, 1644, 459, 1731, 2208, 428, 699, 958, 2392, 1308, 1101, 1639, 89, 2117, 2109, 300, 174, 1886, 2110, 1559, 1590, 917, 2372, 2445, 624, 2196, 1011, 1066]. trials [1453, 134, 462, 1318, 1060, 240, 603, 2029, 1570, 1113, 566, 98, 709, 1596, 2347, 696, 1752, 2077, 1498, 2269, 565, 637, 423, 867, 2050, 1091, 514, 697, 1073, 100, 422, 1336, 1112, 2098, 379, 99, 891, 1925, 2535, 418, 636, 1747, 155, 538, 833, 749, 1831, 724, 1488, 320, 2428, 1299, 1608, 2639, 972, 378, 2099, 577, 567, 1919, 2226, 2578, 2489, 170, 735, 887, 297, 1391, 1680, 2003, 2484, 377, 924, 1365, 2496, 1074, 743, 1679, 2647, 1514, 1507, 1942, 2086, 2453, 1037, 1061, 2112, 1171, 2589, 1198, 2569, 996, 60, 1643, 814, 2346, 1974, 2198, 1118, 130]. trials [1403, 605, 1138, 1310, 389, 78, 2100, 718, 350, 870, 704, 1052, 1055, 536, 1065, 67, 808, 2075, 1068, 2400, 2629, 1479, 1141, 2554, 2074, 1072, 1283, 2, 2079, 1468, 4, 1996, 810, 2229, 1490, 1768, 1319, 2645, 1715, 2348, 411, 834, 158, 353, 391, 943, 932, 1036, 1100, 1461, 1947, 2214, 1420, 1939, 1865, 2414, 903, 407, 959, 1063, 492, 196, 1051, 812, 1346, 2640, 1678, 2308, 81, 6, 1220, 2197, 2430, 1954, 2250, 2033, 2143, 2108, 804, 1828, 1489, 2002, 2076, 390, 802, 1179, 2032, 2523, 702, 1714, 693, 1010, 1722, 66, 104, 1067, 2383, 786, 554]. trials [110, 1000, 1086, 830]. Triangular [1065, 102, 305, 2398]. tribute [2022]. Trichopoulos [920]. triglyceride [1638]. tropical [1037]. Truncated [2092, 1923, 1929]. truncation [1391]. Trust [553, 450]. Truth [2310]. Tsay [1935]. Tsiatis [89]. Tsodikov [2602]. Tsuang [2150]. tubal [685]. tuberculosis [1789]. Tucker [1359]. Tudying [989]. Tufte [758, 2192]. Tugwell [667]. Tukey [764]. Tulinius [132]. Tumor [2602]. tumour [1579, 983, 109, 316, 985, 2325, 474, 678, 1408]. tumours [984]. Tuomilehto [2005, 2004]. Tutorial [1845, 2001, 1886, 2637]. twenty [22]. twenty-first [22]. twin [2502, 836, 2135]. twins [933]. Two [2338, 1914, 724, 1409, 1577, 2122, 1784, 1423, 2285, 2186, 434, 491, 1649, 1762, 277, 888, 1760, 2482, 2130, 1844, 373, 737, 195, 2155, 1644, 1854, 1157, 2097, 1506, 2549, 276, 240, 1953, 1009, 1008, 2330, 321, 1418, 2416, 2551, 1110, 2489, 544, 629, 1997, 176, 1562, 1273, 313, 1116, 2139, 1851, 742, 1403, 2561, 2516, 1645, 782, 988, 1479, 2174, 1468, 509, 1761, 1094, 2025, 2548, 87, 40, 2413, 1478, 621, 1191, 1750, 1181, 1880, 274, 2076, 2376, 869, 796, 503, 2431, 866, 2648]. two-by-two [724]. two-dimensional [2139]. two-mutation [1562]. Two-part [1914]. two-period [240, 1009, 1008, 629, 1478]. two-phase [1854, 2431]. Two-Sample [1784, 1409, 2482, 737, 1506, 988, 1468, 40]. two-sequence [1844]. Two-sided [2122, 1644, 1750, 2648]. Two-stage [724, 1762, 1760, 276, 240, 1645, 1761, 2376, 503, 866]. Two-state [2338, 2186]. two-step [87]. two-treatment [240, 742, 1479]. two-way [491, 373]. type [2402, 1225, 423, 173, 85, 282, 2087, 2631]. types [2114, 2570, 1480].

U [2354]. U.K [2394, 402, 2278, 2381, 1912]. U.K. [2315]. U.S [1994]. U.S. [689, 2605, 1233, 860, 1664]. U.S.A [2090, 596, 511, 2194, 243]. UCL [1372]. UICC [770]. UK [70]. UK£ [2069]. ulcer [2158, 147]. Ulf [2019]. ultrasonographic [534]. unanimity [442]. unavailable [1510]. unbalanced [1980, 299, 538, 501, 1285, 1497]. Unbiased [1332, 2582]. unblinding [2050, 1420]. uncertainty [2284, 2175, 1560, 1743]. unconditional [729, 1341, 519]. uncontrolled [1679]. uncorrelated [1268]. underestimating [1074]. undergraduates [115, 283, 121]. underlying [2017, 2333]. Understanding [2270, 2055, 1669, 87, 655, 2239]. UNDP [1037]. UNDP/World [1037]. unequal [2452, 2329, 2025]. unequally [474, 678]. unexpected [2023]. unhyphenated [2623]. unidentified [736]. unified [2324, 1112, 2523]. uniform [2349]. uniform-balanced [2349]. uniformity [1543]. uniformly [1733, 2005, 2004]. unifying [917]. Union [132]. Unit [828, 1870, 1, 2274]. unit-weighted [2274]. United [2252, 596, 2608, 2019, 1124, 1803, 573, 836, 2058, 601, 2060]. units [2390, 1114]. unity [174, 918, 2645]. univariate [1697, 2377, 2337, 2574]. Univeristy [841]. Universities [402]. University [23, 2584, 777, 434, 770, 510, 1169, 1170, 1938, 1934, 2252, 345, 1165, 1984, 2147, 1048, 1922, 840, 366, 362, 731, 2254, 206, 892, 1936, 2396, 759, 2379, 20, 1898, 1899, 1194, 2378, 1923, 788, 947, 570, 1037, 112, 24, 1849, 164, 1897, 2311, 207, 2093, 1982, 1971, 921, 348, 1003, 347, 43, 2149, 246, 1912, 1025, 1315, 1001, 184, 2178]. unknown [1291]. unmatched [611]. Unobserved [2482]. unpaired [1566, 1344]. unpublished [1729]. unrestricted [953]. unseen [416]. unspecified [1363]. Unusual [2265]. unweighted [2013]. up-and-down [1287]. update [428, 2647, 66, 2397]. updated [949, 1406]. updating [1400]. upon [581, 1383]. uptake [931]. Urbach [325]. urban [1535]. urinary [216, 2581]. US$32.95 [2418]. US$40.00 [2006]. US$59.95 [2069]. Usage [1937]. Use [2469, 1709, 1737, 887, 1835, 1304, 722, 1656, 2448, 1666, 775, 1825, 1010, 470, 1393, 35, 1644, 699, 402, 462, 408, 1513, 1623, 1086, 1247, 990, 108, 705, 1159, 860, 541, 2049, 241, 2582, 686, 1749, 1537, 2375, 796, 291, 2147, 366]. used [278, 594, 2289, 1419, 1996, 1094]. useful [2357, 430, 888, 2636, 229, 2550, 687, 498, 1600, 623]. users [602]. uses [1087, 944, 253]. Using [2588, 216, 723, 1913, 1980, 637, 2059, 262, 1711, 351, 2556, 2364, 1924, 289, 1127, 1224, 551, 547, 2497, 2174, 584, 2547, 254, 474, 678, 255, 1882, 2245, 1451, 2351, 750, 1234, 1561, 435, 1990, 1265, 161, 2611, 264, 2558, 2402, 352, 537, 764, 763, 2463, 1129, 753, 939, 1351, 293, 588, 766, 2359, 1104, 760, 603, 2029, 1908, 1131, 1622, 2120, 575, 867, 1987, 1735, 856, 1668, 1022, 2425, 1534, 2576, 914, 538, 1831, 173, 681, 2387, 2151, 2639, 378, 1140, 485, 2013, 297, 2630, 234, 1815, 1883, 1554, 1034, 1562, 2325, 507, 1246, 2606, 2564, 824, 1512, 2139]. using [1077, 2565, 1905, 650, 1694, 2415, 1287, 200, 2028, 260, 1550, 1236, 1266, 918, 238, 591, 2516, 2390, 1389, 2307, 1645, 1632, 751, 752, 2075, 913, 263, 12, 1866, 2335, 392, 1661, 674, 338, 137, 932, 2391, 1657, 1237, 163, 1415, 2328, 48, 643, 2592, 1031, 1637, 2642, 2055, 2006]. usual [305]. uterine [2199]. utilities [2308]. utility [958]. Utilization [1252, 1251, 2230, 2341]. utilizing [1723]. Utts [2040]. UV [2046]. Uwe [2132].

V [1359, 2055, 2127, 2331, 1898, 2081, 2495, 2345, 1809, 108, 2381, 2008, 1811, 2126, 432, 511, 2181, 527]. VA [1607]. vaccination [2612, 1800, 2293, 2510, 1720]. vaccine [1717, 2062, 2052, 1718, 160, 1113, 2269, 211, 375, 1721, 677, 2153, 2327, 1141, 1408, 171]. vaccines [951, 2327]. vaginal [215, 998]. Validation [1640, 926, 333, 1125, 372, 1697, 439, 469, 2026, 900, 2574, 2575, 1400]. Validity [1459, 1941, 1851]. Valsecchi [1961]. value [1450, 1892, 1475, 804, 159, 989]. valued [2503, 2567]. values [1647, 61, 344, 2128, 666, 1840, 2129, 1676, 1510, 806, 1988, 516, 847, 726, 2480, 1243, 2079, 1705, 80, 2646, 1134, 1102]. variability [1817, 888, 62, 215, 1633, 372, 998, 427, 546, 340, 2423, 857, 605, 359, 2016, 1019]. variable [1646, 203, 1103, 2065, 378, 1342, 761, 1643, 1974, 1546, 547, 2457, 1330, 1778, 1010]. Variables [1788, 1046, 1104, 2576, 1868, 858, 532, 2013, 1554, 2569, 1266, 1224, 1146, 2358, 12, 2255, 1493, 501, 274, 2291, 2194]. Variance [1186, 2134, 258, 533, 2371, 2372, 485, 1368, 1891, 1992, 374, 359, 1920, 1471]. variances [2428, 585, 2329, 1114]. variates [590]. Variation [1840, 141, 2031, 2649, 530, 1437, 781, 2294, 449, 1986, 1526, 2234, 448, 1431, 47, 1674, 1441, 2554, 1675, 1152, 1018, 1660, 2128, 2343, 2129]. Variations [2173, 475, 925, 2325, 1726]. varies [1791]. various [2215, 454, 2632]. Varley [1277]. varying [1464, 1030, 1386, 1906]. varying-coefficient [1906]. VAT [777]. vector [1915]. Vedel [2601]. velocity [689, 2302]. Venables [1901]. ventricular [524]. Venzon [1562]. Verification [1237, 1102]. Verifying [819, 2590]. Verlag [1901, 1826, 403, 480, 2092, 362, 2314, 1887, 384, 91, 1962, 2179, 328, 1086, 404, 2091, 2419, 2417, 2418, 383, 1108, 555, 2020, 2180]. Verma [1823]. Vernon [1166]. Version [1973]. versus [530, 1208, 521, 1149, 1903, 917, 514, 1792, 2526, 2361, 2048, 4, 686, 1013]. Verwandtes [2419]. very [1990]. veterans [779]. VI [1193, 614, 1165, 1899, 2495, 1455, 1443, 2439, 921]. via [1572, 1289, 1605, 2596, 1906, 2161]. viability [2199]. Video [777]. Viertl [1861]. view [1476, 1461, 2432, 702]. viewed [2118]. viewpoint [921]. Vignaux [1824]. vii [553, 366, 1983, 1859, 1194, 1154, 2418, 2168]. viii [23, 403, 364, 1000, 404, 1923, 1004, 841, 284, 166, 1277, 1315, 1164, 234]. villus [2515]. violation [1331]. violations [1369]. viral [2279]. virologic [1723]. virological [1796]. Virtala [2005, 2004]. virus [2402, 1698, 2058, 743, 1118, 551]. Visit [2484]. Visit-driven [2484]. visits [1247, 1544, 773]. Visual [790, 1601, 543, 1210, 2192]. visualization [2415]. Visualizing [1338]. vital [2649, 1986, 1392, 2557, 212]. vitamin [964]. vitro [1381, 820, 2490, 985, 2199, 651]. vivo [2490]. Voelkl [2418]. VOL [526, 2276, 92, 93, 1933, 1813, 362, 1898, 404, 1443]. Vollandt [2180]. Volume [1193, 788]. volumes [434]. Vuylsteek [510].

W [1901, 614, 553, 2394, 2332, 598, 2130, 326, 1876, 244, 1887, 1362, 2018, 2369, 2286, 2081, 165, 2169, 788, 2068, 2313, 2417, 2571, 1877, 2168, 43, 2635, 1872, 2313]. Wadsworth [365]. Wald [2567]. Wales [432, 362, 2378, 1370, 1950, 1966, 396, 1562]. walk [678]. Walker [527]. Wallace [840, 841]. Wallgren [2019]. Walter [553, 2343, 788, 2132]. Wang [1192]. ward [2469]. Ware [2403]. Warren [2237]. Washington [596, 22]. Wassertheil [328]. Wassertheil-Smoller [328]. waste [1439]. Water [92, 163]. Water-house [92]. Waterman [2276]. Waternaux [1359]. Wavelets [2384]. way [491, 2259, 373, 299]. Webb [1277]. weekdays [925]. weekends [925]. Wei [1829, 1835]. Weibull [2501, 2335]. Weichung [2143]. Weighing [1200]. weight [445, 2084, 687, 2356, 1906]. Weighted [1760, 2558, 236, 2607, 1754, 1024, 2274]. weighting [1449, 2557]. Weir [1825]. Weissfeld [1835]. welcome [2604, 1655]. Welfare [402]. well [2278].

Wellcome [831]. Wells [1014, 968]. Westfall [1027]. Whelan [1193]. Where [995, 888, 118, 1978, 2354]. whether [660]. Which [458, 1418, 996, 2076]. white [2605, 313]. Whitehead [1568, 870]. Whitney [2416, 2448]. WHO [165, 1037, 118, 1680, 2003, 742, 2002]. WI [2006]. wideband [942]. width [2454]. Wiksell [1371]. Wilcoxon [1225, 1525, 1355, 85, 282, 2568]. Wilcoxon-type [1225, 85, 282]. Wiley [1961, 1812, 2462, 2394, 1027, 2367, 2312, 730, 1983, 482, 326, 244, 613, 2204, 2018, 2148, 2041, 2263, 2157, 165, 2573, 2345, 2586, 1824, 1972, 2082, 1924, 612, 1847, 1278, 946, 285, 2494, 2439, 1002, 1787, 2585, 2150, 2043, 1838, 2383, 1277, 2054, 1960, 2167]. Wilkins [2571]. will [660]. William [656, 1786, 2237, 1338]. Williams [730, 2571]. window [1546]. Windows [2547]. Wings [1950]. winner [2398, 1691, 1722]. Withdrawal [564]. within [432, 2488, 1762, 475, 546, 1139, 340, 1891, 1480, 1114]. within-person [546]. within-subject [340, 1891]. within-units [1114]. Without [2379, 2585, 2638, 2165, 2050, 1932, 1420]. Wittes [145]. Wojtek [1922]. women [1117, 1992, 1389]. Woodbury [1924, 1771]. Woolson [840]. Work [482, 510, 831, 1523, 2310]. Work-book [482]. work-related [510]. Workbook [2254]. workers [1105, 656]. workgroup [1866]. working [1366, 383]. workplace [1542]. Workshop [827, 308, 435, 875, 1596, 283]. World [2310, 2602, 2125, 2601, 1037, 2380, 288]. Worrying [2038]. Worst [1994]. Worth [2038]. would [89, 2]. Wright [2239]. www.stams.strath.ac.uk [2278].

X [2277, 1838, 827, 1934, 2394, 2368, 94, 362, 872, 1086, 327, 2091, 2132, 1951, 2585, 1981, 770, 2312, 1876, 613, 2275, 1859, 1911, 2038, 1888, 1970, 2419, 2599, 1972, 570, 2544, 2313, 2418, 1787, 526, 1166, 2383, 2381, 1001]. xi [2252, 1048, 1359, 840, 1014, 2041, 2396, 759, 2586, 655, 2494, 1108, 2383, 184, 245]. Xian [2148]. xii [1798, 2090, 146, 598, 480, 91, 892, 2042, 1085, 165, 2345, 1372, 1896, 1972, 2572, 570, 481, 1823, 148, 285, 1002, 2150, 2170, 2384, 325, 787, 246, 2069, 1960, 893, 2178]. xiii [1901, 1985, 2253, 597, 2367, 326, 613, 433, 346, 829, 164, 2600, 347, 1788, 920]. xiv [730, 1876, 20, 944, 1924, 526, 1971, 1025]. xix [90, 24, 2167]. XL [2179]. xv [452, 2408, 362, 1799, 2286, 1192, 2169, 22, 2601, 612, 946, 400, 528]. xvi [1961, 770, 2602, 365, 1153, 2148, 451, 111, 2545, 2134, 2006, 328, 947, 2461, 2040, 1982, 1166, 1001, 1811, 2382]. xvii [2310, 2462, 1984, 1027, 362, 1026, 731, 2369, 2263, 2157, 667, 1370, 2380, 1849, 347, 43, 1314, 2193]. xviii [345, 482, 2546, 383, 348, 1003, 2149, 1912, 1935]. xx [871, 2378, 1038]. xxi [554, 1824, 2417, 2311, 2381]. xxii [1938, 2312, 362, 788, 1037, 1278]. xxiii [2204, 1898]. xxiv [450, 1454, 2007]. xxv [933, 830]. xxvii [1087]. xxviii [2395]. xxx [2194]. xxxi [1193]. xxxiv [434].

Yakovlev [2602]. Yale [1982]. Yanagawa [1015]. Yandell [2069]. Yates [28, 508]. Ye [2638, 2165]. year [694, 655, 1958, 1957, 1324]. Yeargers [2081]. years [1197, 912, 1864, 1362, 199, 1210]. Ying [2165]. York [1901, 1117, 656, 945, 597, 2462, 345, 828, 1165, 1984, 1027, 2276, 2367, 840, 2312, 146, 1983, 326, 244, 613, 2314, 871, 2204, 2254, 1153, 346, 206, 554, 2545, 1911, 759, 872, 2286, 145, 1192, 2157, 1994, 1888, 1970, 328, 1086, 2573, 2091, 2068, 1972, 2546, 2313, 1924, 612, 2417, 2418, 1847, 1278, 112, 2494, 1002, 2311, 870, 2150, 207, 2093, 2043, 1166, 383, 2053, 1108, 2181, 166, 347, 246, 830, 2054, 1315, 1960, 2020, 2180, 1811, 2178, 1935, 2167, 1236]. Youden [1560]. Young [1027, 1439, 2125].

Zahner [2150]. Zanchetti [1862]. Zealand [2096]. Zelen [2647]. zero [2177, 174, 2474, 2645, 2567]. zeros [1390]. Zhou [2148]. zidovudine [2151, 1473].

References

Hsieh:1988:SSF

[1] F. Y. Hsieh. Sample size formulae for intervention studies with the cluster as unit of randomization. Statistics in Medicine, 7(11):1195–1201, November 1988. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [1870].

Rosner:1989:IGS

[2] Gary L. Rosner and Anastasios A. Tsiatis. The impact that group sequential tests would have made on ECOG clinical trials. Statistics in Medicine, 8(4):505–516, April 1989. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [89].

Anonymous:1990:Ma

[3] Anonymous. Masthead. Statistics in Medicine, 9(3):fmi, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salsburg:1990:HVS

[4] David Salsburg. Hypothesis versus significance testing for controlled clinical trials: a dialogue. Statistics in Medicine, 9(3):201–211, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1990:C

[5] L. Freedman and D. Byar. Comment. Statistics in Medicine, 9(3):213– 214, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:1990:IHC

[6] Peter F. Thall and Richard Simon. Incorporating historical control data in planning phase II clinical trials. Statistics in Medicine, 9(3):215–228, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Storer:1990:SPI

[7] Barry E. Storer. A sequential phase II/III trial for binary outcomes. Statistics in Medicine, 9(3):229–235, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zielhuis:1990:CBA

[8] G. A. Zielhuis, H. Straatman, A. E. Van ’t Hof-Grootenboer, H. J. J. Van Lier, G. H. Rach, and P. Van Den Broek. The choice of a balanced allocation method for a clinical trial in otitis media with effusion. Statistics in Medicine, 9(3):237–246, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1990:BOS

[9] Sander Greenland and Alberto Salvan. Bias in the one-step method for pooling study results. Statistics in Medicine, 9(3):247–252, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Young:1990:PRT

[10] Michael A. Young and Louis F. Fogg. Predicting response to treatment: Differentiating active-factor and non-specific effects. Statistics in Medicine, 9(3):253–261, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vonesh:1990:MPR

[11] Edward F. Vonesh. Modelling peritonitis rates and associated risk factors for individuals on continuous ambulatory peritoneal dialysis. Statistics in Medicine, 9(3):263–271, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Puterman:1990:OCP

[12] Martin L. Puterman, Peter Schumacher, and George G. S. Sandor. Optimal choice of prognostic variables with an application to cardiac monitoring using m-mode echocardiography. Statistics in Medicine, 9(3):273– 286, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:1990:ACM

[13] Hwa-Gan H. Chang, Lloyd L. Lininger, Joseph T. Doyle, Patricia A. Maccubbin, and Richard B. Rothenberg. Application of the Cox model as a predictor of relative risk of coronary heart disease in the Albany Study. Statistics in Medicine, 9(3):287–292, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bristol:1990:SAT

[14] D. R. Bristol and J. V. Castellana. Survival analysis techniques in angina pectoris trials. Statistics in Medicine, 9(3):293–299, March 1990. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schork:1990:DEM

[15] Nicholas J. Schork, Alan B. Weder, M. Anthony Schork, David R. Bassett, and Stevo Julius. Disease entities, mixed multi-normal distributions, and the role of the hyperkinetic state in the pathogenesis of hypertension. Statistics in Medicine, 9(3):301–314, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hansen:1990:PMR

[16] Eva Stottrup Hansen. The proportionate mortality ratio and its relationship with measures of mortality. Statistics in Medicine, 9(3):315–323, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goddard:1990:ROC

[17] M. J. Goddard and I. Hinberg. Receiver operator characteristic (ROC) curves and non-normal data: an empirical study. Statistics in Medicine, 9(3):325–337, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Ca

[18] Anonymous. Correction. Statistics in Medicine, 9(3):339, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cologne:1990:SCL

[19] J. B. Cologne, P. M. Mendelman, D. O. Chaffin, G. Dunn, John B. Cologne, Paul M. Mendelman, and Donald O. Chaffin. Statistical comparisons of ligand-binding kinetics. Statistics in Medicine, 9(3):341–342, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grambsch:1990:BRD

[20] Patricia Grambsch. Book review: Data analysis for clinical medicine: The quantitative approach to patient care in gastroenterology. Thomas C. Chalmers (ed.), International University Press, Rome, 1988. no. of pages: xiv + 217. price: $49.00. Statistics in Medicine, 9(3):343–344, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baum:1990:BRS

[21] Michael Baum and Richard Sylvester. Book review: Statistics with confidence: Confidence intervals and statistical guidelines. Martin J. Gardner and Douglas G. Altman (eds), British Medical Journal, London, 1989. no. of pages: 140. price: £7.95 ($19.00). Statistics in Medicine, 9(3):344– 345, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kalache:1990:BRA

[22] Alex Kalache. Book review: The aging population in the twenty-first century; statistics for health policy. Dorothy M. Gilford (ed.), National Academy Press, Washington D.C., 1988. No. of pages: xv + 323. price: £17.15. Statistics in Medicine, 9(3):346–347, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:1990:BRC

[23] Odd O. Aalen. Book review: Chronic disease modelling. K. G. Manton and E. Stallard, Mathematics in Medicine 2, Griffin and Oxford University Press, 1988. No. of pages: viii + 279. price: £29.50. Statistics in Medicine, 9(3):347–348, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Majumder:1990:BRP

[24] Partha P. Majumder. Book review: Pedigree analysis in human genetics. Elizabeth A. Thompson, Johns Hopkins University Press, Baltimore, 1986. No. of pages: xix + 213. price: £22.50 ($35.00). Statistics in Medicine, 9(3):348–349, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:SSM

[25] Anonymous. Symposium on statistical methods for evaluation of intervention and prevention strategies. Statistics in Medicine, 9(3):350, March 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Mb

[26] Anonymous. Masthead. Statistics in Medicine, 9(4):fmi, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chinn:1990:AMM

[27] Susan Chinn. The assessment of methods of measurement. Statistics in Medicine, 9(4):351–362, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haviland:1990:YCC

[28] Mark G. Haviland. Yates’s correction for continuity and the analysis of 2×2 contingency tables. Statistics in Medicine, 9(4):363–367, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mantel:1990:C

[29] Nathan Mantel. Comment. Statistics in Medicine, 9(4):369–370, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:1990:C

[30] Samuel W. Greenhouse. Comment. Statistics in Medicine, 9(4):371– 372, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnard:1990:C

[31] George Barnard. Comment. Statistics in Medicine, 9(4):373–375, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagostino:1990:C

[32] Ralph B. D’agostino. Comment. Statistics in Medicine, 9(4):377–378, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Overall:1990:C

[33] John E. Overall. Comment. Statistics in Medicine, 9(4):379–382, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haviland:1990:AR

[34] Mark G. Haviland. Author’s rejoinder. Statistics in Medicine, 9(4): 383, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chambless:1990:ULR

[35] Lloyd E. Chambless, Annette J. Dobson, C. C. Patterson, and Byron Raines. On the use of a logistic risk score in predicting risk of coronary heart disease. Statistics in Medicine, 9(4):385–396, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gordon:1990:ATF

[36] N. H. Gordon. Application of the theory of finite mixtures for the estimation of ‘cure’ rates of treated cancer patients. Statistics in Medicine, 9(4):397–407, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koehler:1990:ALS

[37] Kenneth J. Koehler and Paul G. McGovern. An application of the LFP survival model to smoking cessation data. Statistics in Medicine, 9(4): 409–421, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flanders:1990:ERR

[38] W. Dana Flanders, Rebecca Dersimonian, and Philip Rhodes. Estimation of risk ratios in case-base studies with competing risks. Statistics in Medicine, 9(4):423–435, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lefante:1990:PDD

[39] John J. Lefante. The power to detect differences in average rates of change in longitudinal studies. Statistics in Medicine, 9(4):437–446, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tandon:1990:ANT

[40] P. K. Tandon, Herbert Stander, and Richard P. Schwarz, Jr. Application of new two-sample tests to data from a randomized placebo-controlled heart-failure trial. Statistics in Medicine, 9(4):447–456, April 1990. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Staszewski:1990:MLLa

[41] Robert Staszewski. Murphy’s law of limiting dilution cloning. Statistics in Medicine, 9(4):457–461, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ingram:1990:ECP

[42] D. D. Ingram, J. C. Kleinman, and P. Mock. Empirical comparisons of proportional hazards and logistic regression models. Statistics in Medicine, 9(4):463–464, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stroup:1990:BRS

[43] Donna F. Stroup. Book review: Surveillance in health and disease. W. J. Eylenbosch and N. D. Noah (eds), Oxford University Press, Oxford, 1988. No. of pages: xvii + 286. price: £35.00. Statistics in Medicine, 9 (4):465–466, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thacker:1990:BRA

[44] Stephen B. Thacker. Book review: Analysis of infectious disease data. Niels G. Becker, Chapman and Hall, London, 1989. No. of pages: ix + 224. price: £20. Statistics in Medicine, 9(4):468, April 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Mc

[45] Anonymous. Masthead. Statistics in Medicine, 9(5):fmi, May 1990. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gabriel:1990:BBR

[46] K. Ruben Gabriel and Charles L. Odoroff. Biplots in biomedical research. Statistics in Medicine, 9(5):469–485, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:1990:MEV

[47] Edward L. Korn and Richard Simon. Measures of explained variation for survival data. Statistics in Medicine, 9(5):487–503, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:1990:EDT

[48] Jeremy M. G. Taylor, Alvaro Muñoz, Sue M. Bass, Alfred J. Saah, Joan S. Chmiel, and Lawrence A. Kingsley. Estimating the distribution of times from HIV seroconversion to AIDS using multiple imputation. Statistics in Medicine, 9(5):505–514, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:1990:MTS

[49] Sylvia Richardson. A method for testing the significance of geographical correlations with application to industrial lung cancer in France. Statistics in Medicine, 9(5):515–528, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Esteve:1990:RSE

[50] J. Estève, E. Benhamou, M. Croasdale, and L. Raymond. Relative survival and the estimation of net survival: Elements for further discussion. Statistics in Medicine, 9(5):529–538, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1990:MLE

[51] M. L. Thompson and G. B. Theron. Maximum likelihood estimation of reference centiles. Statistics in Medicine, 9(5):539–548, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:1990:ODB

[52] Shein-Chung Chow and Siu-Keung Tse. Outlier detection in bioavailability/bioequivalence studies. Statistics in Medicine, 9(5):549–558, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Uebersax:1990:LCA

[53] John S. Uebersax and William M. Grove. Latent class analysis of diagnostic agreement. Statistics in Medicine, 9(5):559–572, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pal:1990:TNR

[54] Nabendu Pal and Partha P. Majumder. On a test of non-random segregation under linkage for autosomal recessive diseases. Statistics in Medicine, 9(5):573–582, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carroll:1990:CAG

[55] R. J. Carroll, Stephen Senn, and Raymond J. Carroll. Covariance analysis in generalized linear measurement error models. Statistics in Medicine, 9(5):583–586, May 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Md

[56] Anonymous. Masthead. Statistics in Medicine, 9(6):fmi, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McPherson:1990:D

[57] Klim McPherson, Ted Colton, Laurence Freedman, Tony Johnson, and David Machin. Dedication. Statistics in Medicine, 9(6):587–589, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coloton:1990:PA

[58] Ted Coloton. Personal addendum. Statistics in Medicine, 9(6):591–593, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:ISCa

[59] Anonymous. International Society for Clinical Biostatistics Tenth International Meeting held 11–14 September 1989 at Maastricht, The Netherlands. Statistics in Medicine, 9(6):594, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McPherson:1990:SSR

[60] Klim McPherson. Sequential stopping rules in clinical trials. Statistics in Medicine, 9(6):595–600, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnard:1990:MCT

[61] G. A. Barnard. Must clinical trials be large? The interpretation of p-values and the combination of test results. Statistics in Medicine, 9 (6):601–614, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Breslow:1990:FSV

[62] Norman Breslow. Further studies in the variability of pock counts. Statistics in Medicine, 9(6):615–626, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mantel:1990:HMF

[63] Nathan Mantel and M. Adrian Gross. How much further investigation? Statistics in Medicine, 9(6):627–631, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Healy:1990:MI

[64] M. J. R. Healy. Measuring importance. Statistics in Medicine, 9(6):633– 637, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finney:1990:RMW

[65] David J. Finney. Repeated measurements: What is measured and what repeats? Statistics in Medicine, 9(6):639–644, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zelen:1990:RCD

[66] M. Zelen. Randomized consent designs for clinical trials: an update. Statistics in Medicine, 9(6):645–656, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2647].

Pocock:1990:EIC

[67] Stuart J. Pocock and Michael D. Hughes. Estimation issues in clinical trials and overviews. Statistics in Medicine, 9(6):657–671, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schneiderman:1990:PAP

[68] Marvin A. Schneiderman. Peter Armitage: a personal recollection. Statistics in Medicine, 9(6):673–676, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Day:1990:ADM

[69] N. E. Day. The Armitage–Doll multistage model of carcinogenesis. Statistics in Medicine, 9(6):677–679, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Darby:1990:TIH

[70] S. C. Darby, R. Doll, B. Thakrar, C. R. Rizza, and D. R. Cox. Time from infection with HIV to onset of AIDS in patients with haemophilia in the UK. Statistics in Medicine, 9(6):681–689, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bithell:1990:ADE

[71] J. F. Bithell. An application of density estimation to geographical epidemiology. Statistics in Medicine, 9(6):691–701, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anscombe:1990:SCE

[72] F. J. Anscombe. The summarizing of clinical experiments by significance levels. Statistics in Medicine, 9(6):703–708, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:BPP

[73] Anonymous. Bibliography. Publications of Peter Armitage to December 1989. Statistics in Medicine, 9(6):709–716, June 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Me

[74] Anonymous. Masthead. Statistics in Medicine, 9(7):fmi, July 1990. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:EN

[75] Anonymous. Editor’s note. Statistics in Medicine, 9(7):717, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanDruten:1990:HID

[76] J. A. M. Van Druten, A. G. M. Reintjes, J. C. Jager, S. H. Heisterkamp, M. J. J. C. Poos, R. A. Coutinho, M. G. W. Dijkgraaf, and E. J. Ruitenberg. HIV infection dynamics and intervention experiments in linked risk groups. Statistics in Medicine, 9(7):721–736, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marubini:1990:LTS

[77] E. Marubini, G. Mezzanotte, P. Boracchi, and U. Veronesi. Long term survival analysis: Breast cancer and age at diagnosis. Statistics in Medicine, 9(7):737–748, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olschewski:1990:SAQ

[78] M. Olschewski and M. Schumacher. Statistical analysis of quality of life data in cancer clinical trials. Statistics in Medicine, 9(7):749–763, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Drescher:1990:DCC

[79] Karsten Drescher, Jürgen Timm, and Karl-Heinz Jöckel. The design of case-control studies: The effect of confounding on sample size requirements. Statistics in Medicine, 9(7):765–776, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:1990:EET

[80] Michael Schemper and Terry L. Smith. Efficient evaluation of treatment effects in the presence of missing covariate values. Statistics in Medicine, 9(7):777–784, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:1990:CCT

[81] D. W. Taylor and E. G. Bosch. CTS: a clinical trials simulator. Statistics in Medicine, 9(7):787–801, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bristol:1990:MMC

[82] David R. Bristol and H. I. Patel. A Markovian model for comparing incidences of side effects. Statistics in Medicine, 9(7):803–809, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hochberg:1990:MPP

[83] Yosef Hochberg and Yoav Benjamini. More powerful procedures for multiple significance testing. Statistics in Medicine, 9(7):811–818, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tandon:1990:AGS

[84] P. K. Tandon. Applications of global statistics in analysing quality of life data. Statistics in Medicine, 9(7):819–827, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pee:1990:SWT

[85] David Yu-Wu Pee and Laurence S. Freedman. A stratified Wilcoxon-type test for trend. Statistics in Medicine, 9(7):829–834, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirji:1990:NIA

[86] Karim F. Hirji and Michael H. Rosove. A note on interrater agreement. Statistics in Medicine, 9(7):835–839, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stanek:1990:TSM

[87] Edward J. Stanek III. A two-step method for understanding and fitting growth curve models. Statistics in Medicine, 9(7):841–851, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ingram:1990:AR

[88] D. D. Ingram and J. C. Kleinman. Author’s reply. Statistics in Medicine, 9(7):853, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Facey:1990:LEI

[89] Karen M. Facey and John Whitehead. Letter to the Editor: The impact that group sequential tests would have made on ECOG clinical trials. Gary Rosner and Anastasios A. Tsiatis, Statistics in Medicine, 8, 505– 516 (1989). Statistics in Medicine, 9(7):853–854, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2].

Black:1990:BRH

[90] Nick Black. Book review: Health inequalities in European countries. J. Fox (ed), Gower, Aldershot, 1989. No. of pages: xix + 414. Statistics in Medicine, 9(7):855–856, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1990:BRC

[91] Allan Donner. Book review: Clinical epidemiology and biostatistics. A primer for clinical investigators and decision-makers. Michael S. Kramer, Springer-Verlag, Berlin, 1988 No, of pages: xii + 286. Price: 76DM. Statistics in Medicine, 9(7):856, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Earlam:1990:BRC

[92] Richard Earlam. Book review: Clinical cancer monographs. Vol. 1: Cancer of the oesophagus. Matthews, H. R., Water-house, J. A. H., Powell, J., Robertson, J. E. and McConkey, C. (eds,), Macmillan Press, London, 1987. No. of pages: 184. Statistics in Medicine, 9(7):856–857, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heitjan:1990:BRC

[93] Daniel F. Heitjan. Book review: Clinical cancer monographs. Vol. 2: Cancer of the larynx. Peter E. Robin, Jean Powell, Guy M. Holme el al., Macmillan Press, London, 1989. No. of pages: 238. Price: £D40. Statistics in Medicine, 9(7):857–858, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1990:BRA

[94] M. J. Campbell. Book review: Applied nonparametric statistical methods. P. Sprent, Chapman and Hall, 1989. No. of pages: x + 259. Price:

£D14.95. Statistics in Medicine, 9(7):858–859, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1990:SRN

[95] Douglas G. Altman. Short reviews of new editions. Statistics in Medicine, 9(7):861, July 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Mf

[96] Anonymous. Masthead. Statistics in Medicine, 9(8):fmi, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1990:RMH

[97] Ted Colton, Samuel W. Greenhouse, Marvin Zelen, Edmund A. Gehan, David Demets, James H. Ware, Tavia Gordon, John M. Lachin, and Janet Wittes. Remembrances of Max Halperin. Statistics in Medicine, 9 (8):863–870, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gehan:1990:HMD

[98] Edmund A. Gehan and Marvin A. Schneiderman. Historical and methodological developments in clinical trials at the national cancer institute. Statistics in Medicine, 9(8):871–880, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halperin:1990:EMD

[99] Max Halperin, David L. Demets, and James H. Ware. Early methodological developments for clinical trials at the national heart, lung and blood institute. Statistics in Medicine, 9(8):881–892, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:1990:SHM

[100] Samuel W. Greenhouse. Some historical and methodological developments in early clinical trials at the national institutes of health. Statistics in Medicine, 9(8):893–901, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byar:1990:DPH

[101] David P. Byar. Discussion of papers on ‘Historical and methodological developments in clinical trials at the National Institutes of Health’. Statistics in Medicine, 9(8):903–906, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bellissant:1990:ATT

[102] Eric Bellissant, Jacques Benichou, and Claude Chastang. Application of the triangular test to phase II cancer clinical trials. Statistics in Medicine, 9(8):907–917, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skene:1990:HMM

[103] A. M. Skene and J. C. Wakefield. Hierarchical models for multicentre binary response studies. Statistics in Medicine, 9(8):919–929, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zwinderman:1990:MCQ

[104] Aeilko H. Zwinderman. The measurement of change of quality of life in clinical trials. Statistics in Medicine, 9(8):931–942, August 1990. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1990:AGE

[105] L. S. Freedman, S. B. Green, and D. P. Byar. Assessing the gain in efficiency due to matching in a community intervention study. Statistics in Medicine, 9(8):943–952, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:1990:MFI

[106] C. P. Farrington. Modelling forces of infection for measles, mumps and rubella. Statistics in Medicine, 9(8):953–967, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flanders:1990:EBS

[107] W. Dana Flanders and Ira M. Longini, Jr. Estimating benefits of screening from observational cohort studies. Statistics in Medicine, 9(8):969– 980, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murrells:1990:ULM

[108] T. J. Murrells, T. M. F. Smith, and D. Machin. The use of logit models to investigate social and biological factors in infant mortality. v: a multilogit analysis of stillbirths, neonatal and post-neonatal mortality. Statistics in Medicine, 9(8):981–998, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gamel:1990:PCM

[109] John W. Gamel, Ian W. McLean, and Saul H. Rosenberg. Proportion cured and mean log survival time as functions of tumour size. Statistics in Medicine, 9(8):999–1006, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hills:1990:BRD

[110] Michael Hills. Book review: Design and analysis of cross-over trials. Byron Jones Michael G. Kenward, Chapman and Hall, 1989. No. of pages:

340. Price: £D27.50. Statistics in Medicine, 9(8):1007, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fox:1990:BRM

[111] John Fox. Book review: Mortality, morbidity and health statistics. Michael R. Alderson Macmillan Press, Basing-stoke. No. of pages: xvi + 501. Price: £D85. Statistics in Medicine, 9(8):1008, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Macrae:1990:BRM

[112] K. D. Macrae. Book review: Making sense of data. J. H. Abramson Oxford University Press, New York, 1988. No. of pages: 326. Price: £D15. Statistics in Medicine, 9(8):1008–1009, August 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Mg

[113] Anonymous. Masthead. Statistics in Medicine, 9(9):fmi, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gardner:1990:P

[114] M. J. Gardner and A. O. Hughes. Preface. Statistics in Medicine, 9 (9):1011–1012, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Appleton:1990:WSS

[115] Davi D. R. Appleton. What statistics should we teach medical undergraduates and graduates? Statistics in Medicine, 9(9):1013–1021, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peters:1990:C

[116] Tim J. Peters. Comment. Statistics in Medicine, 9(9):1023–1027, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Cb

[117] Anonymous. Comment. Statistics in Medicine, 9(9):1029, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clayden:1990:WST

[118] A. David Clayden. Who should teach medical statistics, when, how and where should it be taught? Statistics in Medicine, 9(9):1031–1037, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1990:R

[119] Michael J. Campbell. Response. Statistics in Medicine, 9(9):1039–1041, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Cc

[120] Anonymous. Comment. Statistics in Medicine, 9(9):1043–1044, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:1990:EST

[121] Robert G. Newcombe. Evaluation of statistics teaching given to medical undergraduates. Statistics in Medicine, 9(9):1045–1055, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:1990:R

[122] Deborah Ashby. Response. Statistics in Medicine, 9(9):1057–1060, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Cd

[123] Anonymous. Comment. Statistics in Medicine, 9(9):1061–1062, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:1990:HWS

[124] Gordon D. Murray. How we should approach the future. Statistics in Medicine, 9(9):1063–1068, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evans:1990:SMS

[125] Stephen J. W. Evans. Statistics for medical students in the 1990’s: How should we approach the future? Statistics in Medicine, 9(9):1069–1075, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Ce

[126] Anonymous. Comment. Statistics in Medicine, 9(9):1077–1078, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pawitan:1990:SIM

[127] Yudi Pawitan and Alfred Hallstrom. Statistical interim monitoring of the cardiac arrhythmia suppression trial. Statistics in Medicine, 9(9):1081– 1090, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OHagan:1990:PBA

[128] A. O’Hagan, E. G. Woodward, and L. C. Moodaley. Practical Bayesian analysis of a simple logistic regression: Predicting corneal transplants. Statistics in Medicine, 9(9):1091–1101, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Posner:1990:MIR

[129] Karen L. Posner, Paul D. Sampson, Robert A. Caplan, Richard J. Ward, and Frederick W. Cheney. Measuring interrater reliability among multiple raters: an example of methods for nominal data. Statistics in Medicine, 9(9):1103–1115, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nakamura:1990:MED

[130] Tsuyoshi Nakamura, Kouhei Akazawa, and Yoshiaki Nose. A method for the evaluation of dose-toxicity relationships in clinical trials. Statistics in Medicine, 9(9):1117, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mather:1990:AR

[131] Frances J. Mather, Richard M. Simon, Gary M. Clark, and Daniel D. Authors’ reply. Statistics in Medicine, 9(9):1118, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cliff:1990:BRA

[132] Andrew D. Cliff. Book review: Atlas of cancer incidence in the Nordic countries. Ø. Møller Jensen, B. Carstensen, E. Glattre, B. Malker, E. Pukkala and H. Tulinius, Nordic Cancer Union (The Cancer Societies of Denmark, Finland, Iceland, Norway and Sweden), Puna Musta, Helsinki, 1988. No. of pages: 206. Price: £47 ($79). Statistics in Medicine, 9 (9):1119–1120, September 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Mh

[133] Anonymous. Masthead. Statistics in Medicine, 9(10):fmi, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fletcher:1990:FDC

[134] D. J. Fletcher, S. M. Lewis, and J. N. S. Matthews. Factorial designs for crossover clinical trials. Statistics in Medicine, 9(10):1121–1129, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schall:1990:MSE

[135] Robert Schall and Walter Zucchini. Model selection and the estimation of odds ratios in the presence of extraneous factors. Statistics in Medicine, 9(10):1131–1141, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stasny:1990:SQS

[136] Elizabeth A. Stasny and Harold R. Bauer. Symmetry and quasisymmetry: an example in modelling pairs of sounds from children’s early speech. Statistics in Medicine, 9(10):1143–1155, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Selmer:1990:CPR

[137] Randi Selmer. A comparison of Poisson regression models fitted to multiway summary tables and Cox’s survival model using data from a blood pressure screening in the city of Bergen, Norway. Statistics in Medicine, 9(10):1157–1165, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1990:NPE

[138] Kung-Jong Lui and Phil Rhodes. A note on point estimation of the hazard ratio in exponential distributions. Statistics in Medicine, 9(10): 1167–1173, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schluchter:1990:ECB

[139] Mark D. Schluchter. Estimating correlation between alternative measures of disease progression in a longitudinal study. Statistics in Medicine, 9 (10):1175–1188, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Phillips:1990:PSD

[140] Andrew N. Phillips, Simon G. Thompson, and Stuart J. Pocock. Prognostic scores for detecting a high risk group: Estimating the sensitivity when applied to new data. Statistics in Medicine, 9(10):1189–1198, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jarjoura:1990:VHD

[141] David Jarjoura and Everett Logue. Variation in heart disease mortality across census tracts as a function of overdispersion and social class mixture. Statistics in Medicine, 9(10):1199–1209, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:1990:EPI

[142] Nico J. D. Nagelkerke, Ruth N. Chunge, and Stephen N. Kinoti. Estimation of parasitic infection dynamics when detectability is imperfect. Statistics in Medicine, 9(10):1211–1219, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1990:LES

[143] Niels Keiding, Matthew W. Knuiman, Avital Cnaan, and Louise Ryan. Letter to the editor survival analysis in natural history studies of disease. Statistics in Medicine, 9(10):1221–1222, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evans:1990:BRF

[144] Pamela Evans. Book review: Follies and Fallacies in Medicine. Petr Skrabanck and James McCormick, Tarragon Press, Glasgow, 1989. No. of pages: 170. Price: £7.95. Statistics in Medicine, 9(10):1223–1224, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haybittle:1990:BRR

[145] John Haybittle. Book review: Research Methods in Clinical Oncology. B. J. Leventhal and R. E. Wittes, Raven Press, New York, 1988. No. of pages: ix + 246. Price: $93.50. Statistics in Medicine, 9(10):1224, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buncher:1990:BRB

[146] C. Ralph Buncher. Book review: Biopharmaceutical Statistics for Drug Development. Karl E. Peace, Marcel Dekker, New York, 1987. No of pages: xii + 640. Price: $125.00 US and Canada; $150.00 other countries. Statistics in Medicine, 9(10):1224–1225, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gotzsche:1990:BRD

[147] Peter C. Gøtzsche. Book review: Duodenal Ulcer. Analyses of 293 Randomized Clinical Trials. Thierry Poynard and Jean Pierre Pignon, John Libbey Eurotext, London and Paris, 1989. No. of pages: 149. Price: $30. Statistics in Medicine, 9(10):1225–1226, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:1990:BRI

[148] Gordon D. Murray. Book review: Introductory Medical Statistics. Second Edition. R. F. Mould. Adam Hilger, Bristol, 1989. No. of pages: xii + 192. Price: £9.95. Statistics in Medicine, 9(10):1226–1227, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Aa

[149] Anonymous. Announcement. Statistics in Medicine, 9(10):1228, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Cf

[150] Anonymous. Correction. Statistics in Medicine, 9(10):1228, October 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Mi

[151] Anonymous. Masthead. Statistics in Medicine, 9(11):fmi, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1990:Pa

[152] Ted Colton, Laurence Freedman, Tony Johnson, and David Machin. Preface. Statistics in Medicine, 9(11):iii, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:ISCb

[153] Anonymous. International Society for Clinical Biostatistics Tenth International Meeting held at Maastricht Exhibition and Congress Centre, Maastricht 11–14 September 1989. Statistics in Medicine, 9(11):iv, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Habbema:1990:ACD

[154] J. D. F. Habbema, P. M. M. Bossuyt, D. W. J. Dippel, S. Marshall, and J. Hilden. Analysing clinical decision analyses. Statistics in Medicine, 9(11):1229–1242, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:1990:MCT

[155] Jørgen Hilden and J. Dik F. Habbema. The marriage of clinical trials and clinical decision science. Statistics in Medicine, 9(11):1243–1257, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glasziou:1990:QAS

[156] P. P. Glasziou, R. J. Simes, and R. D. Gelber. Quality adjusted survival analysis. Statistics in Medicine, 9(11):1259–1276, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Facey:1990:IAC

[157] Karen M. Facey and John Whitehead. An improved approximation for calculation of confidence intervals after a sequential clinical trial. Statistics in Medicine, 9(11):1277–1285, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1990:GRC

[158] S. J. Senn, P. Auclair, and Samuel Johnson. The graphical representation of clinical trials with particular reference to measurements over time. Statistics in Medicine, 9(11):1287–1302, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanHouwelingen:1990:PVS

[159] J. C. Van Houwelingen and S. Le Cessie. Predictive value of statistical models. Statistics in Medicine, 9(11):1303–1325, November 1990. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:1990:QMB

[160] Conor Patrick Farrington. Quantifying misclassification bias in cohort studies of vaccine efficacy. Statistics in Medicine, 9(11):1327–1337, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blettner:1990:AMU

[161] Maria Blettner and Heiko Becher. Adjusting for misselection using subsampling in epidemiological studies. Statistics in Medicine, 9(11):1339– 1352, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strijbosch:1990:DMS

[162] Leo W. G. Strijbosch, Ronald J. M. M. Does, and Willem Albers. Design methods for some dose-response models. Statistics in Medicine, 9(11): 1353–1363, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Speakman:1990:PPP

[163] J. R. Speakman. Principles, problems and a paradox with the measurement of energy expenditure of free-living subjects using doubly-labelled water. Statistics in Medicine, 9(11):1365–1380, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Milner:1990:BRR

[164] Philip C. Milner. Book review: Research for health professionals: Design, analysis and ethics. Robert Proulx Heaney and Charles J. Dougherty, Iowa State University Press, Ames, Iowa, 1988. No. of pages: xiii + 268. Price: $34.95. Statistics in Medicine, 9(11):1381, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1990:BRA

[165] F. Y. Hsieh and Andrew A. Liu. Book review: Adequacy of sample size in health studies. Stanley Lemeshow, David W. Hosmer, Jr., Janelle Klar and Stephen K. Lwanga published on behalf of WHO by Wiley, Chichester, 1990. No. of pages: xii + 233. Price: £ D17.50. Statistics in Medicine, 9(11):1382, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:1990:BRS

[166] Weichung J. Shih. Book review: Statistical methodology in the pharmaceutical sciences. D. A. Berry (ed.), Marcel Dekker, New York, 1990. No. of pages: viii + 578. Price: $125. Statistics in Medicine, 9(11):1382–1383, November 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Mj

[167] Anonymous. Masthead. Statistics in Medicine, 9(12):fmi, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1990:Pb

[168] T. Colton, L. Freedman, A. Johnson, and D. Machin. Preface. Statistics in Medicine, 9(12):1385, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jewell:1990:SSI

[169] Nicholas P. Jewell. Some statistical issues in studies of the epidemiology of AIDS. Statistics in Medicine, 9(12):1387–1416, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lagakos:1990:AET

[170] S. W. Lagakos, L. L-Y. Lim, and J. M. Robins. Adjusting for early treatment termination in comparative clinical trials. Statistics in Medicine, 9(12):1417–1424, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stablein:1990:CHV

[171] Donald M. Stablein. Challenges of HIV vaccine development. Statistics in Medicine, 9(12):1425–1431, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Billard:1990:D

[172] L. Billard. Discussion. Statistics in Medicine, 9(12):1433–1437, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hwang:1990:GSD

[173] Irving K. Hwang, Weichung J. Shih, and John S. De Cani. Group sequential designs using a family of type I error probability spending functions. Statistics in Medicine, 9(12):1439–1445, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:1990:TSS

[174] Conor P. Farrington and Godfrey Manning. Test statistics and sample size formulae for comparative binomial trials with null hypothesis of nonzero risk difference or non-unity relative risk. Statistics in Medicine, 9 (12):1447–1454, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2645].

Permutt:1990:TIC

[175] Thomas Permutt. Testing for imbalance of covariates in controlled experiments. Statistics in Medicine, 9(12):1455–1462, December 1990. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Linnet:1990:ELR

[176] Kristian Linnet. Estimation of the linear relationship between the measurements of two methods with proportional errors. Statistics in Medicine, 9(12):1463–1473, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:1990:OAS

[177] Jun-Mo Nam and Thomas R. Fears. Optimum allocation of samples in strata-matching case-control studies when cost per sample differs from stratum to stratum. Statistics in Medicine, 9(12):1475–1483, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1990:SSD

[178] Kung-Jong Lui. Sample size determination for case-control studies: The influence of the joint distribution of exposure and confounder. Statistics in Medicine, 9(12):1485–1493, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Le:1990:SMD

[179] Chap T. Le and Bruce R. Lindgren. Statistical methods for determining risk factors of chronic otitis media with effusion. Statistics in Medicine, 9(12):1495–1500, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:1990:RPL

[180] Ralph B. D’Agostino, Mei-Ling Lee, Albert J. Belanger, L. Adrienne Cupples, Keaven Anderson, and William B. Kannel. Relation of pooled logistic regression to time dependent Cox regression analysis: The Framingham Heart Study. Statistics in Medicine, 9(12):1501–1515, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipsitz:1990:MLR

[181] Stuart R. Lipsitz, Nan M. Laird, and David P. Harrington. Maximum likelihood regression methods for paired binary data. Statistics in Medicine, 9(12):1517–1525, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kubler:1990:STM

[182] Jürgen Kübler and Martin Schumacher. Survival time models for analysing drug combination treatments. Statistics in Medicine, 9(12): 1527–1539, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Staszewski:1990:MLLb

[183] Robert Staszewski. Murphy’s law of limiting dilution cloning revisited. Statistics in Medicine, 9(12):1541, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

West:1990:BRC

[184] Robert West. Book review: Causal relationships in medicine. A practical system for critical appraisal. J. Mark Elwood, Oxford University Press, 1988. No. of pages: xi + 332. Price: £30. Statistics in Medicine, 9 (12):1543, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Ab

[185] Anonymous. Announcement. Statistics in Medicine, 9(12):1544, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1990:Ac

[186] Anonymous. Acknowledgement. Statistics in Medicine, 9(12):1545–1546, December 1990. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Ma

[187] Anonymous. Masthead. Statistics in Medicine, 10(1):fmi, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:1991:MCC

[188] Ralph B. D’Agostino and Timothy C. Heeren. Multiple comparisons in over-the-counter drug clinical trials with both positive and placebo controls. Statistics in Medicine, 10(1):1–6, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zagar:1991:C

[189] Anthony J. Zagar. Comment. Statistics in Medicine, 10(1):7–8, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OBrien:1991:C

[190] Peter C. O’Brien. Comment. Statistics in Medicine, 10(1):9–11, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koch:1991:C

[191] Gary G. Koch. Comment. Statistics in Medicine, 10(1):13–16, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laska:1991:C

[192] Eugene M. Laska and Morris J. Meisner. Comment. Statistics in Medicine, 10(1):17–20, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duncan:1991:C

[193] Xsdavid B. Duncan and Dennis O. Dixon. Comment. Statistics in Medicine, 10(1):21–26, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:1991:R

[194] Ralph B. D’Agostino and Timothy C. Heeren. Rejoinder. Statistics in Medicine, 10(1):27–31, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Day:1991:SSE

[195] Simon J. Day and David F. Graham. Sample size estimation for comparing two or more treatment groups in clinical trials. Statistics in Medicine, 10(1):33–43, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sommer:1991:EEC

[196] Alfred Sommer and Scott L. Zeger. On estimating efficacy from clinical trials. Statistics in Medicine, 10(1):45–52, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mickey:1991:EDE

[197] Ruth M. Mickey, Gregory D. Goodwin, and Michael C. Costanza. Estimation of the design effect in community intervention studies. Statistics in Medicine, 10(1):53–64, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheuvart:1991:AAR

[198] B. Cheuvart and L. Ryan. Adjusting for age-related competing mortality in long-term cancer clinical trials. Statistics in Medicine, 10(1):65–77, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:1991:EEE

[199] James Robins and Sander Greenland. Estimability and estimation of expected years of life lost due to a hazardous exposure. Statistics in Medicine, 10(1):79–93, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mollie:1991:EBE

[200] Annie Mollie and Sylvia Richardson. Empirical Bayes estimates of cancer mortality rates using spatial models. Statistics in Medicine, 10(1):95–112, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1991:BMP

[201] Philip J. Smith. Bayesian methods for prevalence estimation from incomplete administrative lists. Statistics in Medicine, 10(1):113–118, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stanek:1991:EPE

[202] Edward J. Stanek III and Gregory Kline. Estimating prediction equations in repeated measures designs. Statistics in Medicine, 10(1):119–130, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eisenberg:1991:EVI

[203] Bennett Eisenberg. The effect of variable infectivity on the risk of HIV infection. Statistics in Medicine, 10(1):131–139, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kirby:1991:SMC

[204] S. P. J. Kirby, C. M. Theobald, J. Piper, and A. D. Carothers. Some methods of combining class information in multivariate normal discrimination for the classification of human chromosomes. Statistics in Medicine, 10(1):141–149, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:1991:CML

[205] Stan Becker. A comparison of maximum likelihood and Jewell’s estimator of the odds ratio and relative risk in single 2 × 2 tables. Statistics in Medicine, 10(1):151–153, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleiss:1991:BRD

[206] Joseph L. Fleiss. Book review: Design and analysis of reliability studies: The statistical evaluation of measurement errors. Graham Dunn, Oxford University Press, New York, 1989. No. of pages: ix + 198. Price: £25. Statistics in Medicine, 10(1):155–156, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piegorsch:1991:BRD

[207] Walter W. Piegorsch. Book review: Design and analysis of reliability studies: The statistical evaluation of measurement errors. Graham Dunn, Oxford University Press, New York, 1989. No. of pages: ix + 198. Price: £25. Statistics in Medicine, 10(1):156–158, January 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Mb

[208] Anonymous. Masthead. Statistics in Medicine, 10(2):fmi, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farley:1991:E

[209] Timothy M. M. Farley and David Machin. Editorial. Statistics in Medicine, 10(2):159, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frank:1991:BBD

[210] Odile Frank and John Bongaarts. Behavioural and biological determinants of fertility transition in sub-Saharan Africa. Statistics in Medicine, 10(2):161–175, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Griffin:1991:PCE

[211] P. D. Griffin and W. R. Jones. The preliminary clinical evaluation of the safety and efficacy of a fertility regulating vaccine. Statistics in Medicine, 10(2):177–190, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pullum:1991:SMA

[212] Thomas W. Pullum. Statistical methods to adjust for date and age misreporting to improve estimates of vital rates in Pakistan. Statistics in Medicine, 10(2):191–200, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Trussell:1991:MPA

[213] James Trussell. Methodological pitfalls in the analysis of contraceptive failure. Statistics in Medicine, 10(2):201–220, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1991:IFP

[214] Patrick Royston. Identifying the fertile phase of the human menstrual cycle. Statistics in Medicine, 10(2):221–240, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1991:CEP

[215] M. J. Campbell, D. Machin, and C. D’Arcangues. Coping with extra Poisson variability in the analysis of factors influencing vaginal ring expulsions. Statistics in Medicine, 10(2):241–254, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baird:1991:URU

[216] Donna Day Baird, Clarice R. Weinberg, Allen J. Wilcox, D. Robert Mc-Connaughey, and Paul I. Musey. Using the ratio of urinary oestrogen and progesterone metabolites to estimate day of ovulation. Statistics in Medicine, 10(2):255–266, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Belsey:1991:DMB

[217] Elizabeth M. Belsey and Nancy Carlson. The description of menstrual bleeding patterns: Towards fewer measures. Statistics in Medicine, 10(2): 267–284, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:1991:ESA

[218] J. Hilden, J. Modvig, M. T. Damsgaard, and L. Schmidt. Estimation of the spontaneous abortion risk in the presence of induced abortions. Statistics in Medicine, 10(2):285–297, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Aa

[219] Anonymous. Announcement. Statistics in Medicine, 10(2):299, February 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Mc

[220] Anonymous. Masthead. Statistics in Medicine, 10(3):fmi, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1991:EN

[221] T. Colton, L. S. Freedman, A. L. Johnson, and D. Machin. Editors’ note. Statistics in Medicine, 10(3):301, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheiner:1991:SSC

[222] Lewis B. Sheiner, Yukiya Hashimoto, and Stuart L. Beal. A simulation study comparing designs for dose ranging. Statistics in Medicine, 10 (3):303–321, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang-Stein:1991:NAT

[223] Christy Chuang-Stein and Weichung Joseph Shih. A note on the analysis of titration studies. Statistics in Medicine, 10(3):323–328, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kushler:1991:MIH

[224] Robert H. Kushler and Morton B. Brown. A model for the identification of hormone pulses. Statistics in Medicine, 10(3):329–340, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ulm:1991:SMA

[225] K. Ulm. A statistical method for assessing a threshold in epidemiological studies. Statistics in Medicine, 10(3):341–349, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tolley:1991:DEM

[226] Elizabeth A. Tolley, Grant W. Somes, Elaine S. Willey, and Stephanie J. Phelps. Determining efficacy of monitoring devices: Evaluating new technologies. Statistics in Medicine, 10(3):351–360, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1991:END

[227] Chung heng Hsieh. The effect of non-differential outcome misclassification on estimates of the attributable and prevented fraction. Statistics in Medicine, 10(3):361–373, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becher:1991:APP

[228] Heiko Becher. Alternative parameterization of polychotomous models: Theory and application to matched case-ontrol studies. Statistics in Medicine, 10(3):375–382, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1991:OMU

[229] M. Karen Campbell, Allan Donner, and Karen M. Webster. Are ordinal models useful for classification? Statistics in Medicine, 10(3):383–394, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jagger:1991:DAM

[230] Carol Jagger and Christopher J. Sutton. Death after marital bereavement-is the risk increased? Statistics in Medicine, 10(3):395– 404, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berntsen:1991:MRC

[231] Rolf Franck Berntsen, Knut Rasmussen, and Jan F. Bjørnstad. Monitoring a randomized clinical trial for futility: The North-Norwegian Lidocaine Intervention Trial. Statistics in Medicine, 10(3):405–412, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pepe:1991:QSF

[232] Margaret Sullivan Pepe, Gary Longton, and Mark Thornquist. A qualifier Q for the survival function to describe the prevalence of a transient condition. Statistics in Medicine, 10(3):413–421, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nedelman:1991:MLB

[233] Jerry Nedelman, Alice Burns, Joan Cleary, Dan Gordon, Philip Vernon, and Charles E. Lawrence. Modelling length bias in a longitudinallylinked record system of HIV cases. Statistics in Medicine, 10(3):423–431, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1991:PTE

[234] Martin L. Lee and Jeanne M. Lusher. The problem of therapeutic equivalence with paired qualitative data: an example from a clinical trial using haemophiliacs with an inhibitor to factor VIII. Statistics in Medicine, 10 (3):433–441, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Monsour:1991:CIP

[235] Michael J. Monsour, Arthur T. Evans, and Lawrence L. Kupper. Confidence intervals for post-test probability. Statistics in Medicine, 10(3): 443–456, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobson:1991:CIW

[236] Annette J. Dobson, Kari Kuulasmaa, Ernst Eberle, and Josef Scherer. Confidence intervals for weighted sums of Poisson parameters. Statistics in Medicine, 10(3):457–462, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1991:SSR

[237] Kung-Jong Lui. Sample sizes for repeated measurements in dichotomous data. Statistics in Medicine, 10(3):463–472, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Odom-Maryon:1991:GND

[238] Tamara Odom-Maryon, Bryan Langholz, Joyce Niland, and Stanley Azen. Generalization of normal discriminant analysis using Fourier series density estimators. Statistics in Medicine, 10(3):473–485, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Ca

[239] Anonymous. Correction. Statistics in Medicine, 10(3):487, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freeman:1991:PTS

[240] P. R. Freeman. The performance of the two-stage analysis of twotreatment, two-period crossover trials. Statistics in Medicine, 10(3):489– 491, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stanek:1991:CNU

[241] Edward J. Stanek and Susan M. Shetterly. A cautionary note on the use of autoregressive models in analysis of longitudinal data. Statistics in Medicine, 10(3):492–495, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachenbruch:1991:BRS

[242] Peter A. Lachenbruch. Book review: Statistical methods for medical investigations. B. S. Everitt, Edward Arnold, London, 1989. No. of pages: ix + 195. Price: £25. Statistics in Medicine, 10(3):497–498, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McPherson:1991:BRC

[243] Klim McPherson. Book review: Causal inference. Kenneth J. Rothman (ed), Epidemiology Resources Inc., Mass., U.S.A., 1988. No. of pages:

207. Price: $21. Statistics in Medicine, 10(3):498–499, March 1991. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crowder:1991:BRC

[244] Martin Crowder. Book review: Computer intensive methods for testing hypotheses. Eric W. Noreen, Wiley, New York, 1989. No. of pages: ix + 229. Price: £31.80. Statistics in Medicine, 10(3):499–500, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeger:1991:BRT

[245] Scott L. Zeger. Book review: Time series, a biostatistical introduction. Peter J. Diggle, Clarendon Press, Oxford, 1990. No. of pages: xi + 257. price: £35 (hard cover) £15 (paper). Statistics in Medicine, 10(3):501, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Swan:1991:BRS

[246] A. V. Swan. Book review: Statistical methods in epidemiology. Harold A. Khan and Christopher T. Sempos, Oxford University Press, New York 1989. No. of pages: xii + 292. price: £19.50 (paperback) £38.00 (hardback). Statistics in Medicine, 10(3):501–502, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1991:SRN

[247] D. G. Altman. Short reviews of new editions. Statistics in Medicine, 10 (3):502–503, March 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Md

[248] Anonymous. Masthead. Statistics in Medicine, 10(4):fmi, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hui:1991:P

[249] Siu L. Hui, Clem McDonald, Barry Katz, and Charles M. Clark. Preface. Statistics in Medicine, 10(4):505–506, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:LCF

[250] Anonymous. List of conference faculty. Statistics in Medicine, 10(4):507– 508, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McDonald:1991:AHD

[251] Clement J. McDonald and Siu L. Hui. The analysis of humongous databases: Problems and promises. Statistics in Medicine, 10(4):511– 518, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Krakauer:1991:EOM

[252] Henry Krakauer and R. Clifton Bailey. Epidemiologic oversight of the medical care provided to Medicare beneficiaries. Statistics in Medicine, 10(4):521–540, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tierney:1991:PDT

[253] William M. Tierney and Clement J. McDonald. Practice databases and their uses in clinical research. Statistics in Medicine, 10(4):541–557, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Safran:1991:URC

[254] Charles Safran. Using routinely collected data for clinical research. Statistics in Medicine, 10(4):559–564, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strom:1991:UCD

[255] Brian L. Strom, Jeffrey L. Carson, Allan C. Halpern, Rita Schinnar, Ellen Sim Snyder, Paul D. Stolley, Michele Shaw, Hugh H. Tilson, Michael Joseph, Wanju S. Dai, Diane Chen, Robert S. Stern, Ulf Bergman, and Frank Lundin. Using a claims database to investigate drug-induced Stevens-Johnson syndrome. Statistics in Medicine, 10(4): 565–576, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petri:1991:CBI

[256] H. Petri and J. Urquhart. Channeling bias in the interpretation of drug effects. Statistics in Medicine, 10(4):577–581, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:1991:MIH

[257] Donald B. Rubin and Nathaniel Schenker. Multiple imputation in healthare databases: an overview and some applications. Statistics in Medicine, 10(4):585–598, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Orav:1991:VCT

[258] E. J. Orav, T. A. Louis, R. H. Palmer, and E. A. Wright. Variance components and their implications for statistical information in medical data. Statistics in Medicine, 10(4):599–616, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pryor:1991:MAA

[259] David B. Pryor and Kerry L. Lee. Methods for the analysis and assessment of clinical databases: The clinician’s perspective. Statistics in Medicine, 10(4):617–628, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moses:1991:IMR

[260] Lincoln E. Moses. Innovative methodologies for research using databases. Statistics in Medicine, 10(4):629–633, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blumberg:1991:PLD

[261] Mark S. Blumberg. Potentials and limitations of database research illustrated by the QMMP AMI Medicare mortality study. Statistics in Medicine, 10(4):637–646, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hlatky:1991:UDE

[262] Mark A. Hlatky. Using databases to evaluate therapy. Statistics in Medicine, 10(4):647–652, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Psaty:1991:ASP

[263] Bruce M. Psaty, Thomas D. Koepsell, David Siscovick, Pat Wahl, James P. Logerfo, Thomas S. Inui, and Edward H. Wagner. An approach to several problems in using large databases for population-based caseontrol studies of the therapeutic efficacy and safety of anti-hypertensive medicines. Statistics in Medicine, 10(4):653–662, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byar:1991:PUO

[264] David P. Byar. Problems with using observational databases to compare treatments. Statistics in Medicine, 10(4):663–666, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:D

[265] Anonymous. Discussion. Statistics in Medicine, 10(4):667–670, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:DSN

[266] Anonymous. Dedication. Samuel Nathan Regenstrief: 1910–1988. Statistics in Medicine, 10(4):671, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Ab

[267] Anonymous. Announcement. Statistics in Medicine, 10(4):673, April 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Me

[268] Anonymous. Masthead. Statistics in Medicine, 10(5):fmi, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1991:CTS

[269] Patrick Royston. Constructing time-specific reference ranges. Statistics in Medicine, 10(5):675–690, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1991:ERR

[270] P. Royston and J. N. S. Matthews. Estimation of reference ranges from normal samples. Statistics in Medicine, 10(5):691–695, May 1991. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grambsch:1991:ETP

[271] Patricia M. Grambsch and Peter C. O’Brien. The effects of transformations and preliminary tests for non-linearity in regression. Statistics in Medicine, 10(5):697–709, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauck:1991:PER

[272] Walter W. Hauck and Rei Miike. A proposal for examining and reporting stepwise regressions. Statistics in Medicine, 10(5):711–715, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Regal:1991:EMS

[273] Ronald R. Regal and Ernest B. Hook. The effects of model selection on confidence intervals for the size of a closed population. Statistics in Medicine, 10(5):717–721, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whaley:1991:CDM

[274] Fredrick S. Whaley. Comparison of different maximum likelihood estimators in a small sample logistic regression with two independent binary variables. Statistics in Medicine, 10(5):723–731, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomsen:1991:NCE

[275] Birthe Lykke Thomsen, Niels Keiding, and Douglas G. Altman. A note on the calculation of expected survival, illustrated by the survival of liver transplant patients. Statistics in Medicine, 10(5):733–738, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flanders:1991:AMT

[276] W. Dana Flanders and Sander Greenland. Analytic methods for twostage case-control studies and other stratified designs. Statistics in Medicine, 10(5):739–747, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:1991:CTS

[277] G. Berry, R. M. Kitchin, and P. A. Mock. A comparison of two simple hazard ratio estimators based on the logrank test. Statistics in Medicine, 10(5):749–755, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baron:1991:MAM

[278] Anna E. Barön. Misclassification among methods used for multiple group discrimination-the effects of distributional properties. Statistics in Medicine, 10(5):757–766, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Garner:1991:SEC

[279] J. Barry Garner. The standard error of Cohen’s kappa. Statistics in Medicine, 10(5):767–775, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abt:1991:PCC

[280] Klaus Abt. Planning controlled clinical trials on the basis of descriptive data analysis. Statistics in Medicine, 10(5):777–795, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:BCRa

[281] Anonymous. Baseline comparisons in randomized clinical trails. Statistics in Medicine, 10(5):797–799, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pee:1991:SWT

[282] D. Y. W. Pee and Laurence S. Freedmann. A stratified Wilcoxon-type test for trend. Statistics in Medicine, 10(5):799–800, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evans:1991:LEW

[283] Stephen J. W. Evans and Simon J. Day. Letter to the Editor: Workshop on teaching statistics to medical undergraduates. Statistics in Medicine, 9, 1011–1080 (1990). Statistics in Medicine, 10(5):801–802, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schor:1991:BRM

[284] Stanley Schor. Book review: Methodological errors in medical research. Bjørn Andersen, Blackwell Scientific Publications, Oxford, 1990. No. of pages: viii + 270. Price: £35. Statistics in Medicine, 10(5):803–804, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:1991:BRM

[285] Robert G. Newcombe. Book review: Medical statistics. A commonsense approach. Michael J. Campbell and David Machin, Wiley, Chichester, 1990. No. of pages: xii + 172. Price: £9.95. Statistics in Medicine, 10 (5):804–805, May 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Mf

[286] Anonymous. Masthead. Statistics in Medicine, 10(6):fmi, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zipfel:1991:P

[287] Andreas Zipfel. Preface. Statistics in Medicine, 10(6):807, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:SMB

[288] Anonymous. Statistical Methods in Biopharmacy Satellite Meeting of the 47th world Session of the International Statistical Institute, Paris 7-8 September 1989. Statistics in Medicine, 10(6):809, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louis:1991:UEB

[289] Thomas A. Louis. Using empirical Bayes methods in biopharmaceutical research. Statistics in Medicine, 10(6):811–829, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawrence:1991:CBR

[290] C. J. Lawrence and V. F. Trewin. The construction of biochemical reference ranges and the identification of possible adverse drug reactions in the elderly. Statistics in Medicine, 10(6):831–837, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yoshimura:1991:UHC

[291] Isao Yoshimura and Makoto Hayashi. On the use of historical controls in the analysis of laboratory data. Statistics in Medicine, 10(6):839– 841, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kristiansen:1991:SPE

[292] Sören Kristiansen. A statistical procedure for the estimation of accuracy parameters in interlaboratory studies. Statistics in Medicine, 10(6):843– 854, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunstan:1991:EMS

[293] F. D. J. Dunstan and A. B. J. Nix. Estimation of the mean and standard deviation using order statistics. Statistics in Medicine, 10(6):855– 869, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:1991:MTC

[294] P. Bauer. Multiple testing in clinical trials. Statistics in Medicine, 10 (6):871–890, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lehmacher:1991:ACD

[295] W. Lehmacher. Analysis of the crossover design in the presence of residual effects. Statistics in Medicine, 10(6):891–899, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ezzet:1991:REM

[296] Farkad Ezzet and John Whitehead. A random effects model for ordinal responses from a crossover trial. Statistics in Medicine, 10(6):901–907, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [880].

Lasserre:1991:DOD

[297] Virginie Lasserre. Determination of optimal designs using linear models in crossover trials. Statistics in Medicine, 10(6):909–924, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1991:IAC

[298] Peter Armitage. Interim analysis in clinical trials. Statistics in Medicine, 10(6):925–937, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunnett:1991:SMT

[299] Charles W. Dunnett and Ajit C. Tamhane. Step-down multiple tests for comparing treatments with a control in unbalanced one-way layouts. Statistics in Medicine, 10(6):939–947, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Falissard:1991:SEN

[300] Bruno Falissard and Joseph Lellouch. Some extensions to a new approach for interim analysis in clinical trials. Statistics in Medicine, 10(6):949– 957, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauck:1991:IBW

[301] Walter W. Hauck and Sharonerson Anderson. Individual bioequivalence: What matters to the patient. Statistics in Medicine, 10(6):959–960, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Metzler:1991:SSB

[302] Carl M. Metzler. Sample sizes for bioequivalence studies. Statistics in Medicine, 10(6):961–970, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chalmers:1991:PIM

[303] Thomas C. Chalmers. Problems induced by meta-analyses. Statistics in Medicine, 10(6):971–980, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nix:1991:MLT

[304] A. B. J. Nix and F. D. J. Dunstan. Maximum likelihood techniques applied to method comparison studies. Statistics in Medicine, 10(6): 981–988, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benichou:1991:API

[305] Jacques Benichou, Eric Bellissant, and Claude Chastang. Analysis of phase II clinical trials in haematology and oncology: Comparison of the triangular test to the usual methods. Statistics in Medicine, 10(6):989– 990, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tubert:1991:RMM

[306] Pascale Tubert and Bernard Begaud. Random models for margins of a 2 × 2 contingency table and application to pharmacovigilance. Statistics in Medicine, 10(6):991–999, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Cb

[307] Anonymous. Correction. Statistics in Medicine, 10(6):1001, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:COE

[308] Anonymous. Computing in Occupational and Environmental Health II-Ird International Workshop Sponsored by the I.C.O.H. Hotel Meridien, Porte Maillot, Paris, France 7–8 November, 1991. Statistics in Medicine, 10(6):1002, June 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Mg

[309] Anonymous. Masthead. Statistics in Medicine, 10(7):fmi, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stigum:1991:PSH

[310] Hein Stigum, J. K. Grønnesby, P. Magnus, J. M. Sundet, and L. S. Bakketeig. The potential for spread of HIV in the heterosexual population in Norway: a model study. Statistics in Medicine, 10(7):1003–1023, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Easton:1991:FAR

[311] D. F. Easton, J. Peto, and A. G. A. G. Babiker. Floating absolute risk: an alternative to relative risk in survival and case-control analysis avoiding an arbitrary reference group. Statistics in Medicine, 10(7):1025– 1035, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sato:1991:EEA

[312] Tosiya Sato. An estimating equation approach for the analysis of casecontrol studies with exposure measured at several levels. Statistics in Medicine, 10(7):1037–1042, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Manton:1991:ABW

[313] Kenneth G. Manton, Eric Stallard, and Steve Wing. Analyses of black and white differentials in the age trajectory of mortality in two closed cohort studies. Statistics in Medicine, 10(7):1043–1059, July 1991. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1991:AIR

[314] Per Kragh Andersen, Lars Sommer Hansen, and Niels Keiding. Assessing the influence of reversible disease indicators on survival. Statistics in Medicine, 10(7):1061–1067, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1991:ESP

[315] Sander Greenland. Estimating standardized parameters from generalized linear models. Statistics in Medicine, 10(7):1069–1074, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heitjan:1991:GNS

[316] Daniel F. Heitjan. Generalized Norton-Simon models of tumour growth. Statistics in Medicine, 10(7):1075–1088, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:1991:ESM

[317] Mark C. K. Yang, Ronald G. Marks, William B. Clark, and Ingvar Magnusson. Evaluation of statistical methods for monitoring periodontal disease. Statistics in Medicine, 10(7):1089–1097, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

France:1991:AFT

[318] Lesley A. France, John A. Lewis, and Richard Kay. The analysis of failure time data in crossover studies. Statistics in Medicine, 10(7):1099– 1113, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chi:1991:RIB

[319] Eric M. Chi. Recovery of inter-block information in cross-over trials. Statistics in Medicine, 10(7):1115–1122, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

James:1991:AMP

[320] Susan James. Approximate multinormal probabilities applied to correlated multiple endpoints in clinical trials. Statistics in Medicine, 10 (7):1123–1135, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirji:1991:QET

[321] Karim F. Hirji, Shu-Jane Tan, and Robert M. Elashoff. A quasi-exact test for comparing two binomial proportions. Statistics in Medicine, 10 (7):1137–1153, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Cc

[322] Anonymous. Correction. Statistics in Medicine, 10(7):1155, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:BCRb

[323] Anonymous. Baseline comparisons in randomized clinical trials. Statistics in Medicine, 10(7):1157–1159, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:TCP

[324] Anonymous. Testing for a change point in registry data with an example of hypospadias. Statistics in Medicine, 10(7):1159–1160, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1991:BRSa

[325] Stephen Senn. Book review: Scientific reasoning, the Bayesian approach. Colin Howson and Peter Urbach, Open Court, La Salle Illinois, 1989. No. of pages: xii + 312. Price: $34.95. Statistics in Medicine, 10(7):1161– 1162, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:1991:BRA

[326] T. J. Cole. Book review: Applied logistic regression. D. W. Hosmer and S. Lemeshow, Wiley, New York, 1989. No. of pages: xiii + 307. Price: £36.00. Statistics in Medicine, 10(7):1162–1163, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jovanovic:1991:BRS

[327] Borko D. Jovanovic. Book review: Subset selection in regression. A. J. Miller, Chapman and Hall, London, 1990. No. of pages: x + 229. Price: £25. Statistics in Medicine, 10(7):1164–1165, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1991:BRB

[328] A. O. Hughes. Book review: Biostatistics and epidemiology a primer for health professionals. Sylvia Wassertheil-Smoller, Springer-Verlag, New York, 1990. No. of pages: xvi + 119. Price: DM48. Statistics in Medicine, 10(7):1165–1166, July 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Mh

[329] Anonymous. Masthead. Statistics in Medicine, 10(8):fmi, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberg:1991:MSE

[330] P. S. Rosenberg, M. H. Gail, and D. Pee. Mean square error of estimates of HIV prevalence and short-term AIDS projections derived by backcalculation. Statistics in Medicine, 10(8):1167–1180, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walker:1991:MC

[331] Alexander M. Walker and Stephan F. Lanes. Misclassification of covariates. Statistics in Medicine, 10(8):1181–1196, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:1991:AND

[332] Xinhua Liu and Kung-Yee Liang. Adjustment for non-differential misclassification error in the generalized linear model. Statistics in Medicine, 10(8):1197–1211, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:1991:VTL

[333] Michael E. Miller, Siu L. Hui, and William M. Tierney. Validation techniques for logistic regression models. Statistics in Medicine, 10(8):1213– 1226, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:1991:SAR

[334] Odd O. Aalen and Einar Husebye. Statistical analysis of repeated events forming renewal processes. Statistics in Medicine, 10(8):1227–1240, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Desouza:1991:EBF

[335] Cynthia M. Desouza. An empirical Bayes formulation of cohort models in cancer epidemiology. Statistics in Medicine, 10(8):1241–1256, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gart:1991:STH

[336] John J. Gart. Simple tests of homogeneity of controls in matched studies. Statistics in Medicine, 10(8):1257–1266, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Le:1991:HRR

[337] Chap T. Le, Bruce R. Lindgren, Daniel Zelterman, and Andrew J. Umen. Homogeneity of the relative risk in cohort studies. Statistics in Medicine, 10(8):1267–1272, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwartzbaum:1991:EOO

[338] Judith A. Schwartzbaum and Joseph G. Hirschberg. Exploring the order of odds ratios using the bootstrap. Statistics in Medicine, 10(8):1273– 1281, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1991:EDN

[339] Patrick Royston. Estimating departure from normality. Statistics in Medicine, 10(8):1283–1293, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:1991:ERM

[340] William D. Johnson and Varghese T. George. Effect of regression to the mean in the presence of within-subject variability. Statistics in Medicine, 10(8):1295–1302, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oden:1991:EKB

[341] Neal L. Oden. Estimating kappa from binocular data. Statistics in Medicine, 10(8):1303–1311, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scalia-Tomba:1991:PSE

[342] Gianpaolo Scalia-Tomba and Johan Giesecke. A possible source of error in calculating median age at first occurrence of an event from a cumulative distribution of reported ages. Statistics in Medicine, 10(8): 1313–1314, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1991:CCB

[343] M. J. Thompson, D. J. Hand, and B. S. Everitt. Contradictory correlations between derived scales. Statistics in Medicine, 10(8):1315–1319, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnard:1991:AGP

[344] G. A. Barnard. ON alleged gains in power from lower P -values. Statistics in Medicine, 10(8):1321–1322, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Axelson:1991:BRR

[345] Olav Axelson. Book review: Research methods in occupational epidemiology. Harvey Checkoway, Neil E. Pearce and Douglas J. Crawford-Brown, Oxford University Press, New York, 1989. No. of pages: xviii + 344. Price: £30 (hard). Statistics in Medicine, 10(8):1323–1324, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fayers:1991:BRP

[346] Peter Fayers. Book review: Presentation of clinical data. Bert Spilker and John Schoenfelder, Raven Press, New York, 1990. No. of pages: xiii + 553. Price: $96. Statistics in Medicine, 10(8):1324–1325, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strachan:1991:BRI

[347] David P. Strachan. Book review: An introduction to clinical research. Catherine DeAngelis (ed.), Oxford University Press, New York, 1990. No. of pages: xvii + 166. Price: £15. Anatomy of clinical research. an introduction to scientific inquiry in medicine, rehabilitation and related health professions. Franklin Stein, SLACK Incorporated, Thorofare, New Jersey, 1989. No. of pages: xiii + 269. Price: £27.70. Statistics in Medicine, 10(8):1325–1326, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1991:BRSb

[348] Stephen Senn. Book review: Statistical inference and analysis: Selected Correspondence of R. A. Fisher. J. H. Bennett (ed), Oxford University Press, Oxford, 1990. No. of pages: xviii + 380. Price: £40.00. Statistics in Medicine, 10(8):1326, August 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Mi

[349] Anonymous. Masthead. Statistics in Medicine, 10(9):fmi, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmer:1991:CPI

[350] Christopher R. Palmer. A comparative phase II clinical trials procedure for choosing the best of three treatments. Statistics in Medicine, 10 (9):1327–1340, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karon:1991:UPD

[351] John M. Karon. Using provisional diagnoses in monitoring a clinical trial. Statistics in Medicine, 10(9):1341–1348, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang-Stein:1991:TMS

[352] Christy Chuang-Stein, Noel R. Mohberg, and Michaline S. Sinkula. Three measures for simultaneously evaluating benefits and risks using categorical data from clinical trials. Statistics in Medicine, 10(9):1349–1359, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1991:CTD

[353] S. J. Senn and H. Hildebrand. Crossover trials, degrees of freedom, the carryover problem and its dual. Statistics in Medicine, 10(9):1361–1374, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:1991:DOD

[354] Jen-Pei Liu and Chung-Sing Weng. Detection of outlying data in bioavailability/bioequivalence studies. Statistics in Medicine, 10(9):1375–1389, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:1991:MPA

[355] E. Lesaffre and G. Molenberghs. Multivariate probit analysis: a neglected procedure in medical statistics. Statistics in Medicine, 10(9):1391–1403, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wolfe:1991:EES

[356] Robert A. Wolfe, Gina R. Petroni, Catherine G. McLaughlin, and Laurence F. McMahon, Jr. Empirical evaluation of statistical models for counts or rates. Statistics in Medicine, 10(9):1405–1416, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raghunathan:1991:PCD

[357] T. E. Raghunathan. Pooling controls from different studies. Statistics in Medicine, 10(9):1417–1426, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mode:1991:SMS

[358] Charles J. Mode, Daniel Fife, and Steven M. Troy. Stochastic methods for short term projections of symptomatic HIV disease. Statistics in Medicine, 10(9):1427–1440, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1991:EVC

[359] L. Smith and B. Rosner. Estimation of variance components for censored data with applications to blood pressure variability. Statistics in Medicine, 10(9):1441–1452, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morrell:1991:MHT

[360] Christopher H. Morrell and Larry J. Brant. Modelling hearing thresholds in the elderly. Statistics in Medicine, 10(9):1453–1464, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barlow:1991:CMC

[361] William Barlow, Mei-Ying Lai, and Stanley P. Azen. A comparison of methods for calculating a stratified kappa. Statistics in Medicine, 10 (9):1465–1472, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuzick:1991:BRG

[362] Jack Cuzick. Book reviews: Geographic epidemiology: a cluster of new publications. cancer mapping. P. Boyle, C. S. Muir and E. Grundmann (eds) (Recent Results in Cancer Research, vol. 114), Springer Verlag, Berlin, 1989. No. of pages: x + 277. Price: DM158. ISBN 3-540-50490-7. Atlas of disease distributions: Analytic approaches to epidemiological data. Andrew D. Cliff and Peter Haggett, Basil Blackwell, Oxford, 1988. No. of pages: xv + 300. Price: £100. ISBN 0-631-13149-3. Mortality and geography: a review in the Mid-1980’s England and Wales. M. Britton (ed). The Registrar General’s decennial supplement for England and Wales (series DS no. 9) HMSO, London, 1990, No. of pages: xxii + 223. Price: £17.50. ISBN 0-11-691294-4. Diet, life-style and mortality in China: a study of the characteristics of 65 Chinese counties. Chen Junshi, T. Colin Campbell, Li Junyao and Richard Peto, Oxford University Press, Oxford, 1990. No. of pages: xvii + 894. Price: £95. ISBN 0-8014-2453-4. Statistics in Medicine, 10(9):1473–1474, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:1991:BRS

[363] Deborah Ashby. Book review: Statistics in Dentistry. John S. Bulman and John F. Osborn, British Dental Association, London, 1989, No. of pages: 118, Price: £6.95, $13.50. Statistics in Medicine, 10(9):1475–1476, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Esteve:1991:BRA

[364] J. Esteve. Book review: Analyse Statistique Des Données de Survie. Catherine Hill, Catherine Com-Nougue, Andrew Kramar, Thierry Moreau, John O’Quigley, Rachid Senoussi and Claude Chastang, Flammarion, Paris, 1990. No. of pages: viii + 204. Price: FF190. ISBN 2-257-10310-6. Statistics in Medicine, 10(9):1476–1477, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunn:1991:BRD

[365] Graham Dunn. Book review: Designing experiments and analyzing data. A model comparison perspective. Scott E. Maxwell and Harold D. Delany, Wadsworth, Belmont, California, 1990. No. of pages: xvi + 902. Price: Hback; ISE (pbk) £19.95. Statistics in Medicine, 10(9):1477–1478, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1991:BRH

[366] M. J. Campbell. Book review: Health measurement scales. A practical guide to their development and use. David L. Streiner, Geoffrey R. Norman, Oxford University Press, Oxford, 1989. No. of pages: vii + 175. Price: £25 hardback, £15 paperback. Statistics in Medicine, 10(9):1478– 1479, September 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Mj

[367] Anonymous. Masthead. Statistics in Medicine, 10(10):fmi, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Helfenstein:1991:EAS

[368] Ulrich Helfenstein, Ursula Ackermann-Liebrich, Charlotte Braun-Fahrländer, and Hans Urs Wanner. The environmental accident at ‘Schweizerhalle’ and respiratory diseases in children: a time series analysis. Statistics in Medicine, 10(10):1481–1492, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Watier:1991:TSC

[369] Laurence Watier, Sylvia Richardson, and Bruno Hubert. A time series construction of an alert threshold with application to S. bovismorbificans in France. Statistics in Medicine, 10(10):1493–1509, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blettner:1991:RDL

[370] Maria Blettner and John D. Boice, Jr. Radiation dose and leukaemia risk: General relative risk techniques for dose-response models in a matched case-control study. Statistics in Medicine, 10(10):1511–1526, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:1991:MNP

[371] Niels G. Becker, Lyndsey F. Watson, and John B. Carlin. A method of non-parametric back-projection and its application to AIDS data. Statistics in Medicine, 10(10):1527–1542, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:1991:ETV

[372] Shein-Chung Chow and Siu-Keung Tse. On the estimation of total variability in assay validation. Statistics in Medicine, 10(10):1543–1553, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis:1991:AOD

[373] Charles S. Davis. AX one degree of freedom nominal association model for testing independence in two-way contingency tables. Statistics in Medicine, 10(10):1555–1563, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ottenbacher:1991:IIF

[374] Kenneth J. Ottenbacher. Interpretation of interaction in factorial analysis of variance design. Statistics in Medicine, 10(10):1565–1571, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haber:1991:EVE

[375] Michael Haber, Ira M. Longini, Jr., and M. Elizabeth Halloran. Estimation of vaccine efficacy in outbreaks of acute infectious diseases. Statistics in Medicine, 10(10):1573–1584, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lang:1991:EER

[376] Janet M. Lang, Julie E. Buring, Bernard Rosner, Nancy Cook, and Charles H. Hennekens. Estimating the effect of the run-in on the power of the physicians’ health study. Statistics in Medicine, 10(10):1585–1593, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1991:ACT

[377] Young Jack Lee, Jonas H. Ellenberg, Deborah G. Hirtz, and Karin B. Nelson. Analysis of clinical trials by treatment actually received: Is it really an option? Statistics in Medicine, 10(10):1595–1605, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kenward:1991:ACD

[378] Michael G. Kenward and Byron Jones. The analysis of categorical data from cross-over trials using a latent variable model. Statistics in Medicine, 10(10):1607–1619, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hallstrom:1991:HRP

[379] Alfred P. Hallstrom, H. Leon Greene, Melissa L. Huther, and The Cast Investigators. The healthy responder phenomenon in non-randomized clinical trials. Statistics in Medicine, 10(10):1621–1631, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Cd

[380] Anonymous. Correction. Statistics in Medicine, 10(10):1633, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:1991:BRP

[381] Deborah Ashby. Book review: Practical statistics for medical research. Douglas G. Altman, Chapman and Hall, London, 1991. No. of pages: 611. Price: £32.00. Statistics in Medicine, 10(10):1635–1636, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1991:BRB

[382] Simon Thompson. Book review: Basic and clinical biostatistics. Beth Dawson-Saunders and Robert G. Trapp, Appleton and Lange, Connecticut, 1990. No. of pages: ix + 329. Price: £21.35. ISBN: 0-8385-4541-4. Statistics in Medicine, 10(10):1636–1637, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1991:BRP

[383] Patrick Royston. Book review: Prescriptions for working statisticians. Albert Madansky, Springer-Verlag, New York, 1988. No. of pages: xviii + 295. Price: DM 74. Statistics in Medicine, 10(10):1637–1638, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dietz:1991:BRM

[384] K. Dietz. Book review: Mathematical and statistical approaches to AIDS epidemiology. C. Castillo-Chavez (ed.), Lecture Notes in Biomathematics 83, Springer-Verlag, Berlin, 1989. No. of pages: ix + 405. Price: DM 80. Statistics in Medicine, 10(10):1638–1639, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:ASP

[385] Anonymous. Announcement society of pharmaceutical medicine. Statistics in Medicine, 10(10):A1, October 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Mk

[386] Anonymous. Masthead. Statistics in Medicine, 10(11):fmi, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1991:P

[387] Ted Colton, Laurence Freedmann, Tony Johnson, and David MacHin. Preface. Statistics in Medicine, 10(11):1643, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:ISC

[388] Anonymous. International Society for Clinical Biostatistics, Eleventh International Meeting held at Cesar Palace, Nîmes 18–21 September 1990. Statistics in Medicine, 10(11):1645, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OQuigley:1991:MDF

[389] John O’Quigley and Sylvie Chevret. Methods for dose finding studies in cancer clinical trials: a review and results of a Monte Carlo study. Statistics in Medicine, 10(11):1647–1664, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1991:GPA

[390] Anne Whitehead and John Whitehead. A general parametric approach to the meta-analysis of randomized clinical trials. Statistics in Medicine, 10(11):1665–1677, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1991:FCT

[391] S. J. Senn. Falsificationism and clinical trials. Statistics in Medicine, 10 (11):1679–1692, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rossiter:1991:CCC

[392] J. E. Rossiter. Calculating centile curves using kernel density estimation methods with application to infant kidney lengths. Statistics in Medicine, 10(11):1693–1701, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berridge:1991:AFT

[393] Damon M. Berridge and John Whitehead. Analysis of failure time data with ordinal categories of response. Statistics in Medicine, 10(11):1703– 1710, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCrory:1991:MAS

[394] S. J. McCrory and I. Ford. Multivariate analysis of spect images with illustrations in Alzheimer’s disease. Statistics in Medicine, 10(11):1711– 1718, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunstan:1991:DAE

[395] F. D. J. Dunstan and R. W. Marshall. The detection of artefacts in EEG series. Statistics in Medicine, 10(11):1719–1731, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:1991:SSP

[396] C. P. Farrington. Subacute sclerosing panencephalitis in England and Wales: Transient effects and risk estimates. Statistics in Medicine, 10 (11):1733–1744, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bithell:1991:ERR

[397] J. F. Bithell. Estimation of relative risk functions. Statistics in Medicine, 10(11):1745–1751, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benichou:1991:MAE

[398] Jacques Benichou. Methods of adjustment for estimating the attributable risk in case-control studies: a review. Statistics in Medicine, 10(11): 1753–1773, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2644].

Commenges:1991:CES

[399] Daniel Commenges and Thierry Moreau. Comparative efficiency of a survival-based case-control design and a random selection cohort design. Statistics in Medicine, 10(11):1775–1782, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osborn:1991:BRF

[400] John Osborn. Book review: Fundamentals of biostatistics. Bernard Rosner, Third edition, PWS-Kent, Boston, 1990. No. of pages: xv + 655. Price: £14.95 (International Student Edition). Statistics in Medicine, 10 (11):1783, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Daly:1991:BRM

[401] Leslie E. Daly. Book review: Medical statistics on microcomputers. R. A. Brown and J. Swanson Beck. British Medical Journal, London, 1990. No. of pages: ix + 103. Price: E8.95. ISBN CL-7279-4290-3. Statistics in Medicine, 10(11):1783–1785, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Festing:1991:BRG

[402] Michael F. W. Festing. Book review: Guidelines on the care of laboratory animals and their use for scientific purposes. IV planning and design of experiments. Published jointly by The Laboratory Animals Science Association, 20 Queensberry Place, London SW7 2DZ, U.K. and the Universities Federation for Animal Welfare, 8 Hamilton Close, South Mimms, Potters Bar, Herts, EN6 3QD, U.K., 1990. No. of pages: 21. ISBN 0-900767-68-5. Statistics in Medicine, 10(11):1785, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carpenter:1991:BRO

[403] Lucy Carpenter. Book review: Occupational cancer epidemiology. P. Band (ed.), Springer-Verlag, Berlin, 1990 No. of pages: viii + 193. Price: DM 118. ISBN: 3-540-52090-2. Statistics in Medicine, 10(11):1786–1787, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Isham:1991:BRS

[404] Valerie Isham. Book review: Stochastic processes in epidemic theory. (Lecture notes in biomathematics vol. 86). J-P. Gabriel, C. Lefevre and P. Picard (eds), Springer-Verlag, Berlin, 1990. No. of pages: viii + 197. Price: 44DM (paperback). ISBN: 3-540-52571-8. Statistics in Medicine, 10(11):1787–1788, November 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1991:Ml

[405] Anonymous. Masthead. Statistics in Medicine, 10(12):fmi, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1991:E

[406] Ted Colton, Laurence Freedman, Tony Johnson, and David MacHin. Editorial. Statistics in Medicine, 10(12):i–ii, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simon:1991:DPS

[407] Richard Simon. A decade of progress in statistical methodology for clinical trials. Statistics in Medicine, 10(12):1789–1817, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:1991:BCU

[408] Mitchell H. Gail. A bibliography and comments on the use of statistical models in epidemiology in the 1980s. Statistics in Medicine, 10(12): 1819–1885, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:1991:ASM

[409] Colini B. Begg. Advances in statistical methodology for diagnostic medicine in the 1980’s. Statistics in Medicine, 10(12):1887–1895, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1991:SMJ

[410] Douglas G. Altman. Statistics in medical journals: Developments in the 1980s. Statistics in Medicine, 10(12):1897–1913, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schumacher:1991:AQL

[411] Martin Schumacher, Manfred Olschewski, and Gabi Schulgen. Assessment of quality of life in clinical trials. Statistics in Medicine, 10(12): 1915–1930, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1991:SAS

[412] Per Kraghersen Andersen. Survival analysis 1982–1991: The second decade of the proportional hazards regression model. Statistics in Medicine, 10(12):1931–1941, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chappell:1991:SDM

[413] Rick Chappell. Sampling design of multiwave studies with an application to the Massachusetts health care panel study. Statistics in Medicine, 10 (12):1945–1958, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis:1991:SPN

[414] Charles S. Davis. Semi-parametric and non-parametric methods for the analysis of repeated measurements with applications to clinical trials. Statistics in Medicine, 10(12):1959–1980, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eliasziw:1991:AMT

[415] Michael Eliasziw and Allan Donner. Application of the McNemar test to non-independent matched pair data. Statistics in Medicine, 10(12): 1981–1991, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoover:1991:USC

[416] D. R. Hoover, A. Muñoz, V. Carey, N. Odaka, J. M. G. Taylor, J. S. Chmiel, J. Armstrong, and S. H. Vermund. The unseen sample in cohort studies: Estimation of its size and effect. Statistics in Medicine, 10(12): 1993–2003, December 1991. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Ma

[417] Anonymous. Masthead. Statistics in Medicine, 11(1):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Herson:1992:BAC

[418] Jay Herson. Bayesian analysis of cancer clinical trials: an introduction to four papers. Statistics in Medicine, 11(1):1–3, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wieand:1992:DSA

[419] Samuel Wieand and Stephen Cha. Description of the statistical aspects of a study for advanced colorectal cancer patients. Statistics in Medicine, 11(1):5–11, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dixon:1992:BSA

[420] Dennis O. Dixon and Richard Simon. Bayesian subset analysis in a colorectal cancer clinical trial. Statistics in Medicine, 11(1):13–22, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1992:ABS

[421] Laurence S. Freedman and David J. Spiegelhalter. Application of Bayesian statistics to decision making during a clinical trial. Statistics in Medicine, 11(1):23–35, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:1992:SAB

[422] Joel B. Greenhouse. On some applications of Bayesian methods in cancer clinical trials. Statistics in Medicine, 11(1):37–53, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:1992:IAM

[423] A. Lawrence Gould. Interim analyses for monitoring clinical trials that do not materially affect the type I error rate. Statistics in Medicine, 11(1):55–66, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:1992:ABM

[424] M. Ashby, J. M. Neuhaus, W. W. Hauck, P. Bacchetti, D. C. Heilbron, N. P. Jewell, M. R. Segal, and R. E. Fusaro. An annotated bibliography of methods for analysing correlated categorical data. Statistics in Medicine, 11(1):67–99, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:1992:LLM

[425] Mark P. Becker and Alan Agresti. Log-linear modelling of pairwise interobserver agreement on a categorical scale. Statistics in Medicine, 11 (1):101–114, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weiss:1992:RPR

[426] Robert E. Weiss and Carlos G. Lazaro. Residual plots for repeated measures. Statistics in Medicine, 11(1):115–124, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleiss:1992:CBH

[427] J. L. Fleiss, J. T. Bigger, Jr., and L. M. Rolnitzky. The correlation between heart period variability and mean period length. Statistics in Medicine, 11(1):125–129, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1992:RCD

[428] Susan S. Ellenberg. Randomized consent designs for clinical trials: an update. Statistics in Medicine, 11(1):131–132, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Staszewski:1992:MLL

[429] Robert Staszewski. Murphy’s law of limiting dilution cloning revisited: Subcloning. Statistics in Medicine, 11(1):132–133, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bell:1992:OMU

[430] Robert Bell. Are ordinal models useful for classification? Statistics in Medicine, 11(1):133–134, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1992:AR

[431] M. Karen Campbell, Allan Donner, and Karen M. Webster. Authors’ reply. Statistics in Medicine, 11(1):134, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1992:BRL

[432] Douglas Altman and John Bithell. Book review: 1. Leukaemia and lymphoma. An atlas of distribution within areas of England and Wales 1984– 1988. R. A. Cartwright, F. E. Alexander, P. A. McKinney and T. J. Ricketts, Leukaemia Research Fund, Leeds, 1990, no. of pages: v + 161. price: £20/$50. Statistics in Medicine, 11(1):135–136, ???? 1992. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobson:1992:BRR

[433] Annette J. Dobson. Book review: 2. Randomization and Monte Carlo methods in biology. Bryan F. J. Manly, Chapman and Hall, London, 1991. no. of pages: xiii + 281. price: £27.50. Statistics in Medicine, 11(1):136– 137, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1992:BRE

[434] Douglas G. Altman. Book review: 3. Effective care in pregnancy and childbirth. I. Chalmers, M. Enkin and M. J. N. C. Keirse (eds), Oxford University Press, Oxford, 1989. no. of pages: xxxiv + 1516 (in two volumes). price: £225. ISBN: 0-19-261558-0 and 0-19-261880-6. Statistics in Medicine, 11(1):137, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:APS

[435] Anonymous. Announcement; practical statistical evaluation of pharmacokinetic/pharmacodynamic data: a hands on workshop using pcnonlin Stockholm, Sweden 21–24 April 1992. Statistics in Medicine, 11(1):139, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mb

[436] Anonymous. Masthead. Statistics in Medicine, 11(2):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carlin:1992:MAT

[437] John B. Carlin. Meta-analysis for 2 × 2 tables: a Bayesian approach. Statistics in Medicine, 11(2):141–158, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:1992:SSP

[438] S. R. Paul and A. Donner. Small sample performance of tests of homogeneity of odds ratios in K 2 × 2 tables. Statistics in Medicine, 11 (2):159–165, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1992:SVI

[439] Laurence S. Freedman, Barry I. Graubard, and Arthur Schatzkin. Statistical validation of intermediate endpoints for chronic diseases. Statistics in Medicine, 11(2):167–178, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lakatos:1992:CSS

[440] Edward Lakatos and K. K. Gordon Lan. A comparison of sample size methods for the logrank statistic. Statistics in Medicine, 11(2):179–191, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Woolson:1992:CST

[441] Robert F. Woolson and Thomas W. O’Gorman. A comparison of several tests for censored paired data. Statistics in Medicine, 11(2):193–208, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1992:NEI

[442] Kung-Jong Lui. A note on the effect of the intraclass correlation in the multiple reading procedure with a unanimity rule. Statistics in Medicine, 11(2):209–218, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1992:SBA

[443] Sander Greenland. A semi-Bayes approach to the analysis of correlated multiple associations, with an application to an occupational cancermortality study. Statistics in Medicine, 11(2):219–230, ???? 1992. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Badalamenti:1992:SPN

[444] Anthony F. Badalamenti, Robert J. Langs, and Marc Kessler. Stochastic progression of new states in psychotherapy. Statistics in Medicine, 11(2): 231–242, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carter:1992:RCG

[445] Randy L. Carter, Michael B. Resnick, Mario Ariet, Gwowen Shieh, and Edward F. Vonesh. A random coefficient growth curve analysis of mental development in low-birth-weight infants. Statistics in Medicine, 11(2): 243–256, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1992:MBS

[446] Patrick Royston and Simon G. Thompson. Model-based screening by risk with application to Down’s syndrome. Statistics in Medicine, 11 (2):257–268, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seigel:1992:SSC

[447] Daniel Seigel. Sample sizes for constructing confidence intervals and testing hypotheses. Statistics in Medicine, 11(2):269, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goutis:1992:TBB

[448] Costas Goutis, Tom Fearn, Shein-Chung Chow, and Jun Shao. Test for batch-to-batch variation in stability analysis. Statistics in Medicine, 11 (2):269–270, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ekholm:1992:TBB

[449] Anders Ekholm and T. Timothy Chen. Test for batch-to-batch variation in stability analysis. Statistics in Medicine, 11(2):271–275, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glasziou:1992:BRB

[450] Paul Glasziou. Book review: Breast cancer: The decision to screen. Sir Patrick Forrest, Nuffield Provincial Hospital Trust, London, 1990. No. of pages: xxiv + 233. Price: £15. ISBN: 09-00574-74-7. Statistics in Medicine, 11(2):277–278, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Forman:1992:BRC

[451] David Forman. Book review: Cancer: Causes, occurrence and control. Lorenzo Tomatis (Editor in chief), International Agency for Research on Cancer, Lyon, 1990. IARC Scientific Publication No. 100. No. of pages: xvi + 352. Price: £24. Statistics in Medicine, 11(2):278–279, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bland:1992:BRR

[452] J. M. Bland. Book review: A review of biostatistics: a program for self-instruction (4th edition). Paul Leaverton, Little, Brown and Co., Boston, 1991. No. of pages: xv + 111. Price: $18.95. ISBN: 0-316-51854-9. Statistics in Medicine, 11(2):279–280, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mc

Zedeler:1992:DIP

[454] Karin Zedeler, Niels Keiding, and Claus Kamby. Differential influence of prognostic factors on the occurrence of metastases at various anatomical sites in human breast cancer. Statistics in Medicine, 11(3):281–294, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mackisack:1992:PPD

[455] M. Mackisack, A. J. Dobson, and C. R. Heathcote. Predicting the prevalence of a disease in a cohort at risk. Statistics in Medicine, 11(3):295– 305, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duffy:1992:CRL

[456] S. W. Duffy, M. C. South, and N. E. Day. Cluster randomization in large public health trials: The importance of antecedent data. Statistics in Medicine, 11(3):307–316, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:1992:HMR

[457] Helena Chmura Kraemer. How many raters? toward the most reliable diagnostic consensus. Statistics in Medicine, 11(3):317–331, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1992:WMS

[458] Patrick Royston. Which measures of skewness and kurtosis are best? Statistics in Medicine, 11(3):333–343, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Easterbrook:1992:DRC

[459] Philippa J. Easterbrook. Directory of registries of clinical trials. Statistics in Medicine, 11(3):345–423, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Ca

[460] Anonymous. Correction. Statistics in Medicine, 11(3):425, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Md

Follmann:1992:USR

[462] Dean Follmann, Janet Wittes, and Jeffrey A. Cutler. The use of subjective rankings in clinical trials with an application to cardiovascular disease. Statistics in Medicine, 11(4):427–437, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelber:1992:C

[463] Richard D. Gelber and Aron Goldhirsch. Comment. Statistics in Medicine, 11(4):439–441, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morgan:1992:C

[464] Timothy M. Morgan and Curt D. Furberg. Comment. Statistics in Medicine, 11(4):443–445, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OBrien:1992:Ca

[465] Peter C. O’Brien. Comment. Statistics in Medicine, 11(4):447–449, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Remington:1992:C

[466] Richard D. Remington. Comment. Statistics in Medicine, 11(4):451– 452, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:1992:R

[467] Dean Follmann, Janet Wittes, and Jeffrey A. Cutler. Rejoinder. Statistics in Medicine, 11(4):453–454, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feuer:1992:GRS

[468] Eric J. Feuer, Benjamin F. Hankey, Jeffrey J. Gaynor, Margaret N. Wesley, Stuart G. Baker, and Jason S. Meyer. Graphical representation of survival curves associated with a binary non-reversible time dependent covariate. Statistics in Medicine, 11(4):455–474, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hadorn:1992:CVP

[469] David C. Hadorn, David Draper, William H. Rogers, Emmett B. Keeler, and Robert H. Brook. Cross-validation performance of mortality prediction models. Statistics in Medicine, 11(4):475–489, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blair:1992:UGG

[470] R. Clifford Blair and Jorge G. Morel. On the use of the generalized t and generalized rank-sum statistics in medical research. Statistics in Medicine, 11(4):491–501, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OBrien:1992:Cb

[471] Peter C. O’Brien. Comment. Statistics in Medicine, 11(4):503–505, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blair:1992:R

[472] R. Clifford Blair and Jorge G. Morel. Rejoinder. Statistics in Medicine, 11(4):507–509, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rom:1992:SSC

[473] Dror M. Rom. Strengthening some common multiple test procedures for discrete data. Statistics in Medicine, 11(4):511–514, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schlain:1992:UAM

[474] Brian R. Schlain, Philip T. Lavin, and Cheryl L. Hayden. Using an autoregressive model to detect departures from steady states in unequally spaced tumour biomarker data. Statistics in Medicine, 11(4):515–532, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bhargava:1992:MLE

[475] Alok Bhargava and Howarth Bouis. Maximum likelihood estimation of between and within variations in energy and protein intakes from infancy to adolescence for The Philippines. Statistics in Medicine, 11(4):533– 545, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:1992:OSS

[476] Jun-Mo Nam and Thomas R. Fears. Optimum sample size determination in stratified case-control studies with cost considerations. Statistics in Medicine, 11(4):547–556, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baughman:1992:APP

[477] Andrew L. Baughman and H. Becher. Alternative parameterization of polychotomous models: Theory and application to matched case-control studies. Statistics in Medicine, 11(4):557–559, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelder:1992:ESP

[478] J. A. Nelder, Sander Greenland, and R. G. Newcombe. Estimating standardized parameters from generalized linear models. Statistics in Medicine, 11(4):559–560, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:1992:LSD

[479] R. G. Newcombe. Latin square designs for crossover studies balanced for carryover effects. Statistics in Medicine, 11(4):560, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Claus:1992:BRR

[480] Elizabeth B. Claus. Book review: Recent progress in the genetic epidemiology of cancer. Henry T. Lynch and Petre Tautu (Eds), Springer-Verlag, Berlin, 1991. No. of pages: xii + 162. Price: DM98. ISBN: 3-540-53022-3. Statistics in Medicine, 11(4):561–562, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mellor:1992:BRS

[481] B. Mellor. Book review: Sample size methodology. M. M. Desu and D. Raghavarao, Academic Press, Boston, 1990. No. of pages: xii + 135. Price: £21.50. ISBN: 0-12-212165-1. Statistics in Medicine, 11(4):562– 563, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clark:1992:BRE

[482] Peter Clark. Book review: Exploring statistics with Minitab. A Workbook for the behavioural sciences. Andrew Monk, John Wiley and Sons, Chichester, 1991. No. of pages: xviii + 247. Price: £11.95. ISBN: 0-471-92391-5. Statistics in Medicine, 11(4):563–564, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Me

Katz:1992:ETP

[484] Barry P. Katz. Estimating transmission probabilities for chlamydial infection. Statistics in Medicine, 11(5):565–577, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kronborg:1992:AIP

[485] Dorte Kronborg, Bettina Hansen, and Peter Aaby. Analysis of the incubation period for measles in the epidemic in Greenland in 1951 using a variance components model. Statistics in Medicine, 11(5):579–590, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiarotti:1992:EDP

[486] Flavia Chiarotti, Massimo Palombi, Nicola Schinaia, Alessandro Ghirardini, and Luca Prospero. Effects of different parametric estimates of seroconversion time on analysis of progression to AIDS among Italian HIV-positive haemophiliacs. Statistics in Medicine, 11(5):591–601, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weerasekera:1992:AMH

[487] D. R. Weerasekera and S. Bennett. Adjustments to the Mantel–Haenszel test for data from stratified multistage surveys. Statistics in Medicine, 11(5):603–616, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sasieni:1992:NPM

[488] Peter Sasieni. A note on the presentation of matched case-control data. Statistics in Medicine, 11(5):617–620, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mori:1992:AEB

[489] Motomi Mori, George G. Woodworth, and Robert F. Woolson. Application of empirical Bayes inference to estimation of rate of change in the presence of informative right censoring. Statistics in Medicine, 11 (5):621–631, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1992:SSR

[490] Kung-Jong Lui and William G. Cumberland. Sample size requirement for repeated measurements in continuous data. Statistics in Medicine, 11(5):633–641, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:1992:CFE

[491] Stuart G. Baker, William F. Rosenberger, and Rebecca Dersimonian. Closed-form estimates for missing counts in two-way contingency tables. Statistics in Medicine, 11(5):643–657, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snapinn:1992:MCT

[492] Steven M. Snapinn. Monitoring clinical trials with a conditional probability stopping rule. Statistics in Medicine, 11(5):659–672, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heitjan:1992:DTE

[493] Daniel F. Heitjan, Peter S. Houts, and Harold A. Harvey. A decisiontheoretic evaluation of early stopping rules. Statistics in Medicine, 11 (5):673–683, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faraone:1992:MMC

[494] Stephen V. Faraone, John C. Simpson, and Walter A. Brown. Mathematical models of complex dose-response relationships: Implications for experimental design in psychopharmacologic research. Statistics in Medicine, 11(5):685–702, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hung:1992:IPD

[495] H. M. James Hung. On identifying a positive dose-response surface for combination agents. Statistics in Medicine, 11(5):703–711, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mf

[496] Anonymous. Masthead. Statistics in Medicine, 11(6):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:EN

[497] Anonymous. Editors’ note. Statistics in Medicine, 11(6):713, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1992:SCM

[498] S. J. Senn. Is the ‘simple carry-over’ model useful? Statistics in Medicine, 11(6):715–726, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grender:1992:RTM

[499] Julie Myers Grender, William D. Johnson, and Robert C. Elston. Regression toward the mean in 2 × 2 crossover designs with baseline measurements. Statistics in Medicine, 11(6):727–741, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1992:SSR

[500] Allan Donner. Sample size requirements for stratified cluster randomization designs. Statistics in Medicine, 11(6):743–750, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shoukri:1992:ENC

[501] M. M. Shoukri and S. W. Martin. Estimating the number of clusters for the analysis of correlated binary response variables from unbalanced data. Statistics in Medicine, 11(6):751–760, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grimson:1992:CSD

[502] Roger C. Grimson, Tim E. Aldrich, and J. Wanzer Drane. Clustering in sparse data and an analysis of rhabdomyosarcoma incidence. Statistics in Medicine, 11(6):761–768, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:1992:DAT

[503] L. P. Zhao and S. Lipsitz. Designs and analysis of two-stage studies. Statistics in Medicine, 11(6):769–782, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cain:1992:ARB

[504] Kevin C. Cain, Richard A. Kronmal, and Andrzej S. Kosinski. Analysing the relationship between change in a risk factor and risk of disease. Statistics in Medicine, 11(6):783–797, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Conaway:1992:PNB

[505] Mark R. Conaway, Christine Waternaux, Elizabeth Allred, David Bellinger, and Alan Leviton. Pre-natal blood lead levels and learning difficulties in children: an analysis of non-randomly missing categorical data. Statistics in Medicine, 11(6):799–811, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:1992:ACI

[506] Edward L. Korn and Frederick J. Dorey. Applications of crude incidence curves. Statistics in Medicine, 11(6):813–829, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longini:1992:ESS

[507] Ira M. Longini, Jr., Robert H. Byers, Nancy A. Hessol, and Wai Y. Tan. Estimating the stage-specific numbers of HIV infection using a Markov model and back-calculation. Statistics in Medicine, 11(6):831– 843, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peritz:1992:YCC

[508] Eric Peritz and Mark G. Haviland. Yates’s correction for continuity and the analysis of 2×2 contingency tables. Statistics in Medicine, 11(6):845– 847, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sato:1992:CTS

[509] Tosiya Sato and G. Berry. A comparison of two simple hazard ratio estimators based on the logrank test. Statistics in Medicine, 11(6):847– 848, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:BRS

[510] Anonymous. Book review: Searching for causes of work-related diseases. An introduction to epidemiology at the work site. Jørn Olsen, Franco Merletti, David Snashall and Karel Vuylsteek, Oxford University Press, Oxford, 1991. No. of Pages: ix + 87. Price: £10.95. ISBN: 0-19-261819-9. Statistics in Medicine, 11(6):849, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Howel:1992:BRI

[511] Denise Howel. Book review: Introduction to modern epidemiology. Anders Ahlbom and Staffan Norrell, 2nd edition, Epidemiology Resources Inc., Chestnut Hill, U.S.A., 1990. No. of Pages: v + 102. Price: $12. ISBN: 0-917227-06-9. Statistics in Medicine, 11(6):849, ???? 1992. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mg

[512] Anonymous. Masthead. Statistics in Medicine, 11(7):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1992:E

[513] Ted Colton, Laurence Freedman, Tony Johnson, and David Machin. Editorial. Statistics in Medicine, 11(7):851, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1992:PVA

[514] Stephanie J. Green and Steve Dahlberg. Planned versus attained design in phase II clinical trials. Statistics in Medicine, 11(7):853–862, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brant:1992:RTM

[515] Larry J. Brant, David B. Duncan, and Dennis O. Dixon. k-ratio t tests for multiple comparisons involving several treatments and a control. Statistics in Medicine, 11(7):863–873, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goodman:1992:CRV

[516] Steven N. Goodman. A comment on replication, P -values and evidence. Statistics in Medicine, 11(7):875–879, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2646].

Liu:1992:EDF

[517] Jen-Pei Liu and Chung-Sing Weng. Estimation of direct formulation effect under log-normal distribution in bioavailability/bioequivalence studies. Statistics in Medicine, 11(7):881–896, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1992:ERR

[518] Patrick Royston. Estimation, reference ranges and goodness of fit for the three-parameter log-normal distribution. Statistics in Medicine, 11 (7):897–912, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shuster:1992:EUT

[519] Jonathan J. Shuster. Exact unconditional tables for significance testing in the 2 × 2 multinomial trial. Statistics in Medicine, 11(7):913–922, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alho:1992:CLR

[520] Juha M. Alho. On the computation of likelihood ratio and score test based confidence intervals in generalized linear models. Statistics in Medicine, 11(7):923–930, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cantor:1992:PVN

[521] Alan B. Cantor and Jonathan J. Shuster. Parametric versus nonparametric methods for estimating cure rates based on censored survival data. Statistics in Medicine, 11(7):931–937, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Munoz:1992:ETS

[522] Alvaro Muñoz, Vincent Carey, Jeremy M. G. Taylor, Joan S. Chmiel, Lawrence Kingsley, Mark van Raden, and Donald R. Hoover. Estimation of time since exposure for a prevalent cohort. Statistics in Medicine, 11 (7):939–952, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Phillips:1992:BRO

[523] Andrew N. Phillips and George Davey Smith. Bias in relative odds estimation owing to imprecise measurement of correlated exposures. Statistics in Medicine, 11(7):953–961, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laor:1992:APV

[524] Arie Laor and Ayala Cohen. Analysis of premature ventricular counts in long term ECG following myocardial infarction. Statistics in Medicine, 11(7):963–973, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bristol:1992:AFT

[525] David R. Bristol, John A. Lewis, and Richard Kay. The analysis of failure time data in crossover studies. Statistics in Medicine, 11(7):975– 977, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pickles:1992:BRS

[526] Andrew Pickles. Book reviews: Statistical methods in longitudinal research. Alexander von Eye (Ed), Academic Press, Boston, 1990. VOL.

1: Principles and structuring change. No. of pages: xiv + 256. Price: $34.95, ISBN: 0-12-724962-1, Vol. 2: Time series and categorical longitudinal data. No. of pages: xiv + 312. Price: $34.95, ISBN: 0-12-724963-X. Statistics in Medicine, 11(7):979–980, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:1992:BRO

[527] Toshiro Tango. Book review: Observation and inference: an introduction to the methods of epidemiology. Alexander M. Walker, Epidemiology Resources Inc., Chestnut Hill, 1991. No. of pages: v + 173. Price: $18. ISBN: 0-917227-07-7. Statistics in Medicine, 11(7):980–981, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sasieni:1992:BRG

[528] Peter Sasieni. Book review: Generalized additive models. T. J. Hastie and R. J. Tibshirani, Chapman and Hall, London, 1990. No. of Pages: xv + 335. Price: £25. ISBN: 0-412-34390-8. Statistics in Medicine, 11 (7):981–982, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mh

[529] Anonymous. Masthead. Statistics in Medicine, 11(8):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernardinelli:1992:EBV

[530] Luisa Bernardinelli and Cristina Montomoli. Empirical Bayes versus fully Bayesian analysis of geographical variation in disease risk. Statistics in Medicine, 11(8):983–1007, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diehr:1992:ARB

[531] Paula Diehr, Thomas D. Koepsell, Allen Cheadle, and Bruce M. Psaty. Assessing response bias in random-digit dialling surveys: The telephoneprefix method. Statistics in Medicine, 11(8):1009–1021, ???? 1992. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnston:1992:DAW

[532] B. Johnston and S. S. Seshia. Discriminant analysis when all variables are ordered. Statistics in Medicine, 11(8):1023–1032, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carr:1992:AVR

[533] Gregory J. Carr and Eric M. Chi. Analysis of variance for repeated measures data: a generalized estimating equations approach. Statistics in Medicine, 11(8):1033–1040, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Espeland:1992:ASS

[534] Mark A. Espeland, Robert P. Byington, Donald Hire, Vicki G. Davis, Tyler Hartwell, and Jeffrey Probstfield. Analysis strategies for serial multivariate ultrasonographic data that are incomplete. Statistics in Medicine, 11(8):1041–1056, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunnett:1992:CBN

[535] Charles W. Dunnett and Ajit C. Tamhane. Comparisons between a new drug and active and placebo controls in an efficacy clinical trial. Statistics in Medicine, 11(8):1057–1063, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pigeon:1992:MDC

[536] Joseph G. Pigeon, Margaret Diponzio Copenhaver, and James P. Whipple. A method of designing clinical trials for combination drugs. Statistics in Medicine, 11(8):1065–1074, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang-Stein:1992:OAS

[537] Christy Chuang-Stein, Noel R. Mohberg, and David M. Musselman. Organization and analysis of safety data using a multivariate approach. Statistics in Medicine, 11(8):1075–1089, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1992:CSS

[538] F. Y. Hsieh. Comparing sample size formulae for trials with unbalanced allocation using the logrank test. Statistics in Medicine, 11(8):1091– 1098, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lehr:1992:SSS

[539] Robert Lehr. Sixteen S-squared over D-squared: a relation for crude sample size estimates. Statistics in Medicine, 11(8):1099–1102, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yateman:1992:SSP

[540] Nigel A. Yateman and Allan M. Skene. Sample sizes for proportional hazards survival studies with arbitrary patient entry and loss to followup distributions. Statistics in Medicine, 11(8):1103–1113, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1992:UCO

[541] Patrick Royston. The use of cusums and other techniques in modelling continuous covariates in logistic regression. Statistics in Medicine, 11 (8):1115–1129, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mi

[542] Anonymous. Masthead. Statistics in Medicine, 11(9):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oden:1992:MNV

[543] Neal Oden. Morphology of the normal visual field in a population-based random sample: Principal components analysis. Statistics in Medicine, 11(9):1131–1150, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:1992:SLS

[544] John M. Lachin. Some large-sample distribution-free estimators and tests for multivariate partially incomplete data from two populations. Statistics in Medicine, 11(9):1151–1170, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holford:1992:MEL

[545] Theodore R. Holford and Michael B. Bracken. A model for estimating level and net severity of spinal cord injuries. Statistics in Medicine, 11 (9):1171–1186, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gillman:1992:CBP

[546] Matthew W. Gillman, Nancy R. Cook, Bernard Rosner, Laurel A. Beckett, Denis A. Evans, Mary Ellen Keough, James O. Taylor, and Charles H. Hennekens. Childhood blood pressure tracking correlations corrected for within-person variability. Statistics in Medicine, 11(9): 1187–1194, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rehm:1992:UFD

[547] J. Rehm, G. Arminger, and L. Kohlmeier. Using follow-up data to avoid omitted variable bias: an application to cardiovascular epidemiology. Statistics in Medicine, 11(9):1195–1208, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1992:DBE

[548] Sander Greenland. Divergent biases in ecologic and individual-level studies. Statistics in Medicine, 11(9):1209–1223, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pickering:1992:DPE

[549] R. M. Pickering. Digit preference in estimated gestational age. Statistics in Medicine, 11(9):1225–1238, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:1992:PSS

[550] John M. Lachin. Power and sample size evaluation for the McNemar test with application to matched case-control studies. Statistics in Medicine, 11(9):1239–1251, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parker:1992:UVC

[551] R. A. Parker and M. A. Pallansch. Using the virus challenge dose in the analysis of virus neutralization assays. Statistics in Medicine, 11(9): 1253–1262, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1992:LE

[552] Sander Greenland and Stephen Senn. Letters to the editor. Statistics in Medicine, 11(9):1263–1265, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashton:1992:BRS

[553] John Ashton. Book review: Screening in health care. Benefit or bane? Walter W. Holland and Susie Stewart, the Nuffield Provincial Hospitals Trust, London, 1990. No. of pages: vii + 233. price: £15. ISBN 0-90057476-3. Statistics in Medicine, 11(9):1267, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelber:1992:BRD

[554] Richard D. Gelber and Aron Goldhirsch. Book review: The design of analgesic clinical trials. Mitchell B. Max, Russell K. Portnoy and Eugene M. Laska (eds.), Raven Press, New York, 1991. no. of pages: xxi + 729. Price: $152.50. ISBN: 0-88167-736-1. Statistics in Medicine, 11(9):1267– 1268, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sherman:1992:BRS

[555] Claire D. Sherman and Christopher J. Portier. Book review: Statistical methods in toxicology. L. Hothorn (ed.), Springer-Verlag, Berlin, 1991. no. of pages: iv + 159. price: DM55: ISBN: 3-540-53621-3. Statistics in Medicine, 11(9):1269–1270, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mj

[556] Anonymous. Masthead. Statistics in Medicine, 11(10):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1992:EN

[557] L. S. Freedman. Editorial note. Statistics in Medicine, 11(10):1271, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wax:1992:CDR

[558] Yohanan Wax. Collinearity diagnosis for a relative risk regression analysis: an application to assessment of diet-cancer relationship in epidemiological studies. Statistics in Medicine, 11(10):1273–1287, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bohning:1992:ABC

[559] Dankmar Böhning, Axel Hempfling, Frank-Peter Schelp, and Peter Schlattmann. The area between curves (ABC) — measure in nutritional anthropometry. Statistics in Medicine, 11(10):1289–1304, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [890].

Cole:1992:SRC

[560] T. J. Cole and P. J. Green. Smoothing reference centile curves: The LMS method and penalized likelihood. Statistics in Medicine, 11(10): 1305–1319, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Willan:1992:CPC

[561] Andrew R. Willan, William Ross, and Thomas A. Mackenzie. Comparing in-patient classification systems: a problem of non-nested regression models. Statistics in Medicine, 11(10):1321–1331, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobbins:1992:RAR

[562] Thomas W. Dobbins and Balasamy Thiyagarajan. A retrospective assessment of the 75/75 rule in bioequivalence. Statistics in Medicine, 11 (10):1333–1342, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McShane:1992:OCP

[563] Lisa M. McShane and Bruce W. Turnbull. Optimal checking procedures for monitoring laboratory analyses. Statistics in Medicine, 11(10):1343– 1357, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:1992:WCS

[564] De Juran Richardson. Withdrawal censoring and the sequential logrank procedure. Statistics in Medicine, 11(10):1359–1366, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goodman:1992:CSR

[565] Phyllis J. Goodman, John J. Crowley, and Carl Benson. Creation of a semiannual report for a multicentre co-operative clinical trials group. Statistics in Medicine, 11(10):1367–1376, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gatsonis:1992:BMP

[566] Constantine Gatsonis and Joel B. Greenhouse. Bayesian methods for phase I clinical trials. Statistics in Medicine, 11(10):1377–1389, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:1992:SSD

[567] Kyungmann Kim and David L. Demets. Sample size determination for group sequential clinical trials with immediate response. Statistics in Medicine, 11(10):1391–1399, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Cb

[568] Anonymous. Correction. Statistics in Medicine, 11(10):1401, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Janosky:1992:LE

[569] J. E. Janosky. Letter to the editor. Statistics in Medicine, 11(10):1403, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lavori:1992:BRD

[570] Philip W. Lavori. Book review: 1. Data quality in longitudinal research. David Magnusson and Lars R. Bergman (eds), Cambridge University Press, 1990. No. of pages: xii + 285. price: £30/$54.50 ISBN: 0-521-38091-X. Statistics in Medicine, 11(10):1405, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Aa

[571] Anonymous. Announcement. Statistics in Medicine, 11(10):1407, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mk

[572] Anonymous. Masthead. Statistics in Medicine, 11(11):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cliff:1992:CGC

[573] Andrew D. Cliff, Peter Haggett, Donna F. Stroup, and Elizabeth Cheney. The changing geographical coherence of measles morbidity in the United States, 1962–88. Statistics in Medicine, 11(11):1409–1424, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heisterkamp:1992:SMT

[574] S. H. Heisterkamp, B. J. de Haan, J. C. Jager, J. A. M. van Druten, and J. C. M. Hendriks. Short and medium term projections of the AIDS/HIV epidemic by a dynamic model with an application to the risk group of homo/bisexual men in Amsterdam. Statistics in Medicine, 11(11):1425– 1441, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilks:1992:ADR

[575] Walter R. Gilks and Sylvia Richardson. Analysis of disease risks using ancillary risk factors, with application to job-exposure matrices. Statistics in Medicine, 11(11):1443–1463, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dabrowska:1992:MCC

[576] Dorota M. Dabrowska, Kjell A. Doksum, Nicholas J. Feduska, Robert Husing, and Padraic Neville. Methods for comparing cumulative hazard functions in a semi-proportional hazard model. Statistics in Medicine, 11 (11):1465–1476, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:1992:SDG

[577] Kyungmann Kim. Study duration for group sequential clinical trials with censored survival data adjusting for stratification. Statistics in Medicine, 11(11):1477–1488, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frisen:1992:EMS

[578] M. Frisén. Evaluations of methods for statistical surveillance. Statistics in Medicine, 11(11):1489–1502, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1992:SSC

[579] Philip J. Smith and Alula Hadgu. Sensitivity and specificity for correlated observations. Statistics in Medicine, 11(11):1503–1509, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1992:GFA

[580] Allan Donner and Michael Eliasziw. A goodness-of-fit approach to inference procedures for the kappa statistic: Confidence interval construction, significance-testing and sample size estimation. Statistics in Medicine, 11 (11):1511–1519, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachenbruch:1992:SSS

[581] Peter A. Lachenbruch. On the sample size for studies based upon McNemar’s test. Statistics in Medicine, 11(11):1521–1525, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vaeth:1992:LE

[582] Michael Væth and J. Barry Garner. Letters to the editor. Statistics in Medicine, 11(11):1527–1532, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Ml

[583] Anonymous. Masthead. Statistics in Medicine, 11(12):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:1992:UMR

[584] Gail Rubin, David Umbach, Shwu-Fang Shyu, and Carlos Castillo-Chavez. Using mark-recapture methodology to estimate the size of a population at risk for sexually transmitted diseases. Statistics in Medicine, 11 (12):1533–1549, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kafadar:1992:AAP

[585] Karen Kafadar and Donna F. Stroup. Analysis of aberrations in public health surveillance data: Estimating variances on correlated samples. Statistics in Medicine, 11(12):1551–1568, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Law:1992:EMP

[586] C. Gordon Law and Ron Brookmeyer. Effects of mid-point imputation on the analysis of doubly censored data. Statistics in Medicine, 11(12): 1569–1578, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coldman:1992:CPC

[587] Andrew J. Coldman, Mary L. McBride, and Terry Braun. Calculating the prevalence of cancer. Statistics in Medicine, 11(12):1579–1589, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:1992:EPG

[588] C. P. Farrington. Estimating prevalence by group testing using generalized linear models. Statistics in Medicine, 11(12):1591–1597, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sato:1992:ECR

[589] Tosiya Sato. Estimation of a common risk ratio in stratified case-cohort studies. Statistics in Medicine, 11(12):1599–1605, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:1992:CBV

[590] Ziding Feng and James E. Grizzle. Correlated binomial variates: Properties of estimator of intraclass correlation and its effect on sample size calculation. Statistics in Medicine, 11(12):1607–1614, ???? 1992. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oman:1992:NLC

[591] Paul Oman. A note on last case influence using the partial likelihood approach under heavy censoring. Statistics in Medicine, 11(12):1615– 1618, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Cc

[592] Anonymous. Corrections. Statistics in Medicine, 11(12):1619, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1992:LE

[593] Philip J. Smith. Letters to the editor. Statistics in Medicine, 11(12): 1621–1622, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bello:1992:MAM

[594] A. L. Bello and Anna E. Barón. Misclassification among methods used for for multiple group discrimination- the effects of distributional properties. Statistics in Medicine, 11(12):1623–1624, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:1992:ABC

[595] J. N. S. Matthews, M. J. Campbell, and Lee Kaiser. Adjusting for baseline: Chamge or percentage change. Statistics in Medicine, 11(12):1624– 1626, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:1992:BRC

[596] Colin B. Begg. Book review: 1. Cross design synthesis: a new strategy for medical effectiveness research. United States General Accounting Office, (GA0/PEMD-92-18), Washington, Dc, 1992. no. of pages: 121. Price: Free (Available from US General Accounting Office, P.O. Box 6015, Gaithersburg, Md 20877, U.S.A.). Statistics in Medicine, 11(12): 1627–1628, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Antes:1992:BRS

[597] Gerd Antes. Book review: 2. Survivorship analysis for clinical studies. Eugene K. Harris and Adelin Albert, Marcel Dekker, New York, 1991. no. of pages: xiii + 200. Price: $90. ISBN: 0-8247-8400-6. Statistics in Medicine, 11(12):1628–1630, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1992:BRS

[598] Mike Campbell. Book review: 3. Statistics for health professionals. Susan Shott, W. B. Saunders, Philadelphia, 1990. no. of pages: xii + 418. Price: £21.50/$29.95. ISBN: 0-7216-8254-5. Statistics in Medicine, 11(12):1630, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Ab

[599] Anonymous. Announcements. Statistics in Medicine, 11(12):1631–1632, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mm

[600] Anonymous. Masthead. Statistics in Medicine, 11(13):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberg:1992:EHP

[601] P. S. Rosenberg, M. H. Gail, and R. J. Carroll. Estimating HIV prevalence and projecting AIDS incidence in the United States: a model that accounts for therapy and changes in the surveillance definition of AIDS. Statistics in Medicine, 11(13):1633–1655, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arca:1992:EDH

[602] Massimo Arcà, Carlo A. Perucci, and Teresa Spadea. The epidemic dynamics of HIV-1 in Italy: Modelling the interaction between intravenous drug users and heterosexual population. Statistics in Medicine, 11(13): 1657–1684, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frison:1992:RMC

[603] Lars Frison and Stuart J. Pocock. Repeated measures in clinical trials: Analysis using mean summary statistics and its implications for design. Statistics in Medicine, 11(13):1685–1704, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2639].

Moye:1992:ACT

[604] Lemuel A. Moyé, Barry R. Davis, and C. Morton Hawkins. Analysis of a clinical trial involving a combined mortality and adherence dependent interval censored endpoint. Statistics in Medicine, 11(13):1705–1717, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neaton:1992:IME

[605] James D. Neaton and Glenn E. Bartsch. Impact of measurement error and temporal variability on the estimation of event probabilities for risk factor intervention trials. Statistics in Medicine, 11(13):1719–1729, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1992:PST

[606] Michael D. Hughes, Claire L. Raskino, Stuart J. Pocock, Maria Rosa Biagini, and Andrew K. Burroughs. Prediction of short-term survival with an application in primary biliary cirrhosis. Statistics in Medicine, 11 (13):1731–1745, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becher:1992:CRC

[607] Heiko Becher. The concept of residual confounding in regression models and some applications. Statistics in Medicine, 11(13):1747–1758, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:1992:CSS

[608] Jun-Mo Nam and Thomas R. Fears. Control sample size when cases are given in constant ratio stratum-matched case-control studies. Statistics in Medicine, 11(13):1759–1766, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Commenges:1992:CPA

[609] Daniel Commenges and Luc Letenneur. Comparison of the proportions of affected relatives of cases and controls: Analysis and minimum sample size formula. Statistics in Medicine, 11(13):1767–1776, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheiner:1992:BR

[610] Lewis B. Sheiner. Bioequivalence revisited. Statistics in Medicine, 11 (13):1777–1788, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Inskip:1992:LEN

[611] Hazel M. Inskip. Letter to the editor. a note on the presentation on unmatched case-control data. Statistics in Medicine, 11(13):1789–1790, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipsitz:1992:BRC

[612] Stuart R. Lipsitz. Book review: Categorical data analysis. Alan Agresti, Wiley, New York, 1990. no. of pages: xv + 558. Price: £48.50, $72.95. ISBN: 0-471-85301-1. Statistics in Medicine, 11(13):1791–1792, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuzick:1992:BRC

[613] Jack Cuzick. Book review: Counting processes and survival analysis. Thomas R. Fleming and David P. Harrington, Wiley, New York, 1991. no. of pages: xiii + 429. Price: £39.95. ISBN: 0-471-52218-X. Statistics in Medicine, 11(13):1792–1793, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1992:BRS

[614] Adelin Albert. Book review: Statistical methods in laboratory medicine. Paul W. Strike, Butterworth–Heinemann Ltd., Oxford, 1991. no. of pages: vi + 542., Price: £70. ISBN: 0-7506-1345-9. Statistics in Medicine, 11(13):1793–1795, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mn

[615] Anonymous. Masthead. Statistics in Medicine, 11(14-15):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Selwyn:1992:IS

[616] Murray R. Selwyn. Introduction to symposium. Statistics in Medicine, 11(14-15):1797, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagostino:1992:LSD

[617] R. B. D’agostino, N. Lange, L. M. Ryan, and M. R. Selwyn. List of speakers and discussants. Statistics in Medicine, 11(14-15):1799, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagostino:1992:O

[618] Ralph B. D’agostino, Nicholas Lange, and Louise Ryan. Overview. Statistics in Medicine, 11(14-15):1801–1805, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:1992:CAA

[619] Donald B. Rubin. Computational aspects of analysing random effects/ longitudinal models. Statistics in Medicine, 11(14-15):1809–1821, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zeger:1992:OMA

[620] Scott L. Zeger and Kung-Yee Liang. An overview of methods for the analysis of longitudinal data. Statistics in Medicine, 11(14-15):1825– 1839, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turney:1992:BRE

[621] Elizabeth A. Turney, Ingrid A. Amara, Gary G. Koch, and William H. Stewart. Book review: Evaluation of alternative statistical methods for linear model analysis to compare two treatments for 24-hour blood pressure response. Statistics in Medicine, 11(14-15):1843–1860, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schluchter:1992:MAI

[622] Mark D. Schluchter. Methods for the analysis of informatively censored longitudinal data. Statistics in Medicine, 11(14-15):1861–1870, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:1992:AFT

[623] L. J. Wei. The accelerated failure time model: a useful alternative to the Cox regression model in survival analysis. Statistics in Medicine, 11(14-15):1871–1879, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fisher:1992:DSC

[624] Lloyd D. Fisher. Discussion of session 1: Clinical trials, survival analysis. Statistics in Medicine, 11(14-15):1881–1885, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Helms:1992:IIL

[625] Ronald W. Helms. Intentionally incomplete longitudinal designs: I. methodology and comparison of some full span designs. Statistics in Medicine, 11(14-15):1889–1913, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:1992:MMB

[626] B. Rosner. Multivariate methods for binary longitudinal data with heterogeneous correlation over time. Statistics in Medicine, 11(14-15):1915– 1928, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vonesh:1992:NLM

[627] Edward F. Vonesh. Non-linear models for the analysis of longitudinal data. Statistics in Medicine, 11(14-15):1929–1954, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagostino:1992:DPS

[628] Ralph B. D’agostino. Discussion of papers of session 2: Incomplete, multivariate, non-linear. Statistics in Medicine, 11(14-15):1955–1963, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laird:1992:ATP

[629] Nan M. Laird, Jonathan Skinner, and Michael Kenward. An analysis of two-period crossover designs with carry-over effects. Statistics in Medicine, 11(14-15):1967–1979, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Butler:1992:REM

[630] Steven M. Butler and Thomas A. Louis. Random effects models with non-parametric priors. Statistics in Medicine, 11(14-15):1981–2000, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lange:1992:GSR

[631] Nicholas Lange. Graphs and stochastic relaxation for hierarchical Bayes modelling. Statistics in Medicine, 11(14-15):2001–2016, ???? 1992. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:1992:DPS

[632] A. Lawrence Gould. Discussion of papers of session 3: Bayes/empirical Bayes. Statistics in Medicine, 11(14-15):2017–2023, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:PA

[633] Anonymous. Poster abstracts. Statistics in Medicine, 11(14-15):2027– 2040, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Mo

[634] Anonymous. Masthead. Statistics in Medicine, 11(16):fmi, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1992:P

[635] Ted Colton, Laurence Freedman, Tony Johnson, and David Machin. Preface. Statistics in Medicine, 11(16):2041, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heyting:1992:SHD

[636] A. Heyting, J. T. B. M. Tolboom, and J. G. A. Essers. Statistical handling of drop-outs in longitudinal clinical trials. Statistics in Medicine, 11 (16):2043–2061, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [891].

Gordon:1992:UTG

[637] N. H. Gordon and J. K. V. Willson. Using toxicity grades in the design and analysis of cancer phase I clinical trials. Statistics in Medicine, 11 (16):2063–2075, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brand:1992:ITI

[638] R. Brand and H. Kragt. Importance of trends in the interpretation of an overall odds ratio in the meta-analysis of clinical trials. Statistics in Medicine, 11(16):2077–2082, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gefeller:1992:CAA

[639] Olaf Gefeller. Comparison of adjusted attributable risk estimators. Statistics in Medicine, 11(16):2083–2091, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauerbrei:1992:BRP

[640] Willi Sauerbrei and Martin Schumacher. A bootstrap resampling procedure for model building: Application to the Cox regression model. Statistics in Medicine, 11(16):2093–2109, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skene:1992:LCM

[641] A. M. Skene and S. A. White. A latent class model for repeated measurements experiments. Statistics in Medicine, 11(16):2111–2122, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:1992:AFI

[642] Jane P. Matthews and Kally Yuen. Analysis of factors influencing the efficacy of pooled platelet transfusions. Statistics in Medicine, 11(16): 2123–2132, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:1992:IPH

[643] Guy Thomas, Genevieve Plu, and Jean-Christophe Thalabard. Identification of pulses in hormone time series using outlier detection methods. Statistics in Medicine, 11(16):2133–2145, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dickson:1992:BRH

[644] Dorothy Dickson and Joerg Hasford. Book review: 24-Hour blood pressure measurement in antihypertensive drug trials: Data requirements and methods of analysis. Statistics in Medicine, 11(16):2147–2158, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1992:Ac

[645] Anonymous. Announcement. Statistics in Medicine, 11(16):i, ???? 1992. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Ma

[646] Anonymous. Masthead. Statistics in Medicine, 12(1):fmi, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:EN

[647] Anonymous. Editor’s note. Statistics in Medicine, 12(1):1, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emerson:1993:DSI

[648] Scott S. Emerson and Julia C. Emerson. Direct standardization of incidence rates in the presence of incomplete data. Statistics in Medicine, 12 (1):3–12, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lie:1993:NSP

[649] Rolv Terje Lie, Ivar Heuch, and Lorentz M. Irgens. A new sequential procedure for surveillance of Down’s syndrome. Statistics in Medicine, 12(1):13–25, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:1993:RAS

[650] J. N. S. Matthews. A refinement to the analysis of serial data using summary measures. Statistics in Medicine, 12(1):27–37, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

deMouzon:1993:ENV

[651] Jacques de Mouzon, Annie Bachelot, and Alfred Spira. Establishing a national in vitro fertilization registry: Methodological problems and analysis of success rates. Statistics in Medicine, 12(1):39–50, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Falissard:1993:SPI

[652] Bruno Falissard and Joseph Lellouch. The succession procedure for interim analysis: Extensions for early acceptance of H_{0} and for flexible times of analysis. Statistics in Medicine, 12(1):51–67, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grender:1993:ACD

[653] Julie Myers Grender and William D. Johnson. Analysis of crossover designs with multivariate response. Statistics in Medicine, 12(1):69–89, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gefeller:1993:LEM

[654] Olaf Gefeller, Geir Egil Eide, and Jacques Benichou. Letters to the editor methods of adjustment for estimating the attributable risk in case-control studies: a review. Statistics in Medicine, 12(1):91–96, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Macintyre:1993:BRU

[655] Celia Macintyre. Book review: Understanding biostatistics. Thomas H. Hassard, Mosby year book, St. Louis, 1991. No. of pages: xi + 324. price: £21.00. ISBN: 0-8016-2078-3. Statistics in Medicine, 12(1):97, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1993:BRBa

[656] Douglas G. Altman. Book review: Basic statics for laboratories. A primer for laboratory workers. William D. Kelley, Thomas A. Ratiff, Jr. and Charles Nenadic, New York, van Nostrand Reinhold, 1992. No. of pages: ix + 175. price: £32.50. ISBN: 0-442-00456-7. Statistics in Medicine, 12(1):98, January 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mb

[657] Anonymous. Masthead. Statistics in Medicine, 12(2):fmi, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:E

[658] Anonymous. Editorial. Statistics in Medicine, 12(2):99, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maxwell:1993:CIC

[659] Scott E. Maxwell. Covariate imbalance and conditional size: Dependence on model-based adjustments. Statistics in Medicine, 12(2):101–109, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schechtman:1993:CAD

[660] Kenneth B. Schechtman and Mae E. Gordon. A comprehensive algorithm for determining whether a run-in strategy will be a cost-effective design modification in a randomized clinical trial. Statistics in Medicine, 12(2): 111–128, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gaffney:1993:HRA

[661] Michael Gaffney, Colin Taylor, and Elizabeth Cusenza. Harmonic regression analysis of the effect of drug treatment on the diurnal rhythm of blood pressure and angina. Statistics in Medicine, 12(2):129–142, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ogorman:1993:ERT

[662] Thomas W. O’Gorman and Robert F. Woolson. On the efficacy of the rank transformation in stepwise logistic and discriminant analysis. Statistics in Medicine, 12(2):143–151, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1993:MTR

[663] Susan Y. J. Zhou, Lawrence A. Kingsley, Jeremy M. G. Taylor, Joan S. Chmiel, David Y. He, and Donald R. Hoover. A method to test for a recent increase in HIV-1 seroconversion incidence: Results from the multicenter AIDS cohort study (MACS). Statistics in Medicine, 12(2): 153–164, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1993:DOA

[664] Laurence S. Freedman, Mahesh K. B. Parmar, and Stuart G. Baker. The design of observer agreement studies with binary assessments. Statistics in Medicine, 12(2):165–179, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1993:PCA

[665] Patrick Royston. A pocket-calculator algorithm for the Shapiro–Francia test for non-normality: an application to medicine. Statistics in Medicine, 12(2):181–184, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bennett:1993:LET

[666] B. M. Bennett and J. Jamart. Letter to the editor. “on tests for equality of predictive values for t diagnostic procedures”. Statistics in Medicine, 12(2):185–186, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1993:BRC

[667] A. O. Hughes. Book review: Clinical epidemiology a basic science for clinical medicine (2nd edition). David L. Sackett, R. Brian Haynes, Gordon H. Guyatt, Peter Tugwell, Little Brown and company, Bosten, 1991. no. of pages: xvii + 441. price: £23.95, $32.50. ISBn: 0316-76599-6. Statistics in Medicine, 12(2):187–188, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmer:1993:BRA

[668] Christopher R. Palmer. Book review: The analysis of contigency tables. 2nd edition, B. S. Everitt, Chapman and Hall, London, 1992. no. of pages: ix + 164. price: £22.50. ISBN: 0-412-39850-8. Statistics in Medicine, 12 (2):188–189, January 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mc

[669] Anonymous. Masthead. Statistics in Medicine, 12(3-4):fmi, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reynolds:1993:E

[670] Gladys Reynolds and Theodore Colton. Editorial. Statistics in Medicine, 12(3-4):191–192, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:LSS

[671] Anonymous. List of symposium staff. Statistics in Medicine, 12(3-4):193, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hopkins:1993:EAO

[672] Donald R. Hopkins. Evaluation: The alpha and omega of intervention and prevention strategies. Statistics in Medicine, 12(3-4):197–205, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feinleib:1993:RND

[673] Manning Feinleib. The role of nchs data systems in the evaluation of intervention and prevention strategies. Statistics in Medicine, 12(3-4): 207–215, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schnell:1993:ESB

[674] Daniel J. Schnell, Christine Galavotti, and Kevin R. O’reilly. An evaluation of sexual behaviour change using statistical and cognitive models. Statistics in Medicine, 12(3-4):219–227, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Allen:1993:AME

[675] L. J. Allen, T. Lewis, C. F. Martin, M. A. Jones, C. K. Lo, M. Stamp, G. Mundel, and A. B. Way. Analysis of a measles epidemic. Statistics in Medicine, 12(3-4):229–239, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wassell:1993:BSM

[676] James T. Wassell and M. L. Moeschberger. A bivariate survival model with modified gamma frailty for assessing the impact of interventions. Statistics in Medicine, 12(3-4):241–248, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longini:1993:MVE

[677] Ira M. Longini, Jr., M. Elizabeth Halloran, Michael Haber, and Robert T. Chen. Measuring vaccine efficacy from epidemics of acute infectious agents. Statistics in Medicine, 12(3-4):249–263, February 1993. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schlain:1993:UAR

[678] Brian R. Schlain, Philip T. Lavin, and Cheryl L. Hayden. Using autoregressive and random walk models to detect trends and shifts in unequally spaced tumour biomarker data. Statistics in Medicine, 12(3-4):265–279, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Drane:1993:IIN

[679] J. Wanzer Drane, Donna Richter, and Carleen Stoskopf. Improved imputation of non-responses to mailback questionnaires. Statistics in Medicine, 12(3-4):283–288, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:1993:SMP

[680] C. Hendricks Brown. Statistical methods for preventive trials in mental health. Statistics in Medicine, 12(3-4):289–300, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ichida:1993:EPC

[681] J. M. Ichida, J. T. Wassell, M. D. Keller, and L. W. Ayers. Evaluation of protocol change in burn-care management using the Cox proportional hazards model with time-dependent covariates. Statistics in Medicine, 12(3-4):301–310, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kafadar:1993:AAI

[682] Karen Kafadar and John M. Karon. An analysis of AIDS incidence data by clustering trends. Statistics in Medicine, 12(3-4):311–326, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martin:1993:EMP

[683] Donald C. Martin, Paula Diehr, Edward B. Perrin, and Thomas D. Koepsell. The effect of matching on the power of randomized community intervention studies. Statistics in Medicine, 12(3-4):329–338, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emerson:1993:DAS

[684] Scott S. Emerson, Daniel L. McGee, Brian Fennerty, Lee Hixson, Harinder Garewal, and David Alberts. Design and analysis of studies to reduce the incidence of colon polyps. Statistics in Medicine, 12(3-4): 339–351, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martinez-Schnell:1993:EET

[685] Beverly Martinez-Schnell, Lynne S. Wilcox, Herbert B. Peterson, Patricia M. Jamison, and Joyce M. Hughes. Evaluating the effects of tubal sterilization on menstrual function: Selected issues in data analysis. Statistics in Medicine, 12(3-4):355–363, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sussman:1993:RAV

[686] Steve Sussman, Alan W. Stacy, Clyde W. Dent, Catherine Burciaga, Dee Burton, and Brian R. Flay. Refusal assertion versus conversational skill role-play competence: Relevance to prevention of tobacco use. Statistics in Medicine, 12(3-4):365–376, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kiely:1993:BWA

[687] John L. Kiely and Joel C. Kleinman. Birth-weight-adjusted infant mortality in evaluations of perinatal care: Towards a useful summary measure. Statistics in Medicine, 12(3-4):377–392, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stevenson:1993:MAC

[688] John M. Stevenson and David R. Olson. Methods for analysing countylevel mortality rates. Statistics in Medicine, 12(3-4):393–401, February 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berkey:1993:LHV

[689] Catherine S. Berkey, Douglas W. Dockery, Xiaobin Wang, David Wypij, and Benjamin Ferris, Jr. Longitudinal height velocity standards for U.S. adolescents. Statistics in Medicine, 12(3-4):403–414, February 1993. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Md

[690] Anonymous. Masthead. Statistics in Medicine, 12(5-6):fmi, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1993:P

[691] Susan S. Ellenberg, Nancy Geller, Richard Simon, and Salim Yusuf. Preface. Statistics in Medicine, 12(5-6):415, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:OR

[692] Anonymous. Opening remarks. Statistics in Medicine, 12(5-6):417, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:1993:BCD

[693] Janet Wittes. Behind closed doors: The data monitoring board in randomized clinical trials. Statistics in Medicine, 12(5-6):419–424, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friedman:1993:NMY

[694] Lawrence Friedman. The NHLBI model: a 25 year history. Statistics in Medicine, 12(5-6):425–431, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Da

[695] Anonymous. Discussion. Statistics in Medicine, 12(5-6):433–434, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

George:1993:SMP

[696] Stephen L. George. A survey of monitoring practices in cancer clinical trials. Statistics in Medicine, 12(5-6):435–450, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1993:DMC

[697] Stephanie Green and John Crowley. Data monitoring committees for Southwest Oncology Group clinical trials. Statistics in Medicine, 12(5-6):451–455, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Db

[698] Anonymous. Discussion. Statistics in Medicine, 12(5-6):457–459, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1993:UEM

[699] Susan S. Ellenberg, Maureen W. Myers, William C. Blackwelder, and Daniel F. Hoth. The use of external monitoring committees in clinical trials of the national institute of allergy and infectious diseases. Statistics in Medicine, 12(5-6):461–467, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dc

[700] Anonymous. Discussion. Statistics in Medicine, 12(5-6):469, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rockhold:1993:DMI

[701] Frank W. Rockhold and Gregory G. Enas. Data monitoring and interim analyses in the pharmaceutical industry: Ethical and logistical considerations. Statistics in Medicine, 12(5-6):471–479, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williams:1993:MCT

[702] George W. Williams, Robert L. Davis, Albert J. Getson, A. Lawrence Gould, Irving K. Hwang, Holly Matthews, Weichung Joseph Shih, Steven M. Snapinn, and Karen L. Walton-Bowen. Monitoring of clinical trials and interim analyses from a drug Sponsor’s point of view. Statistics in Medicine, 12(5-6):481–492, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dd

[703] Anonymous. Discussion. Statistics in Medicine, 12(5-6):493–495, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parmar:1993:MCT

[704] Mahesh K. B. Parmar and David Machin. Monitoring clinical trials: Experience of, and proposals under consideration by, the Cancer Therapy Committee of the British Medical Research Council. Statistics in Medicine, 12(5-6):497–504, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pater:1993:UDM

[705] Joseph L. Pater. The use of data monitoring committees in Canadian trial groups. Statistics in Medicine, 12(5-6):505–508, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyse:1993:IAS

[706] Marc Buyse. Interim analyses, stopping rules and data monitoring in clinical trials in Europe. Statistics in Medicine, 12(5-6):509–520, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:De

[707] Anonymous. Discussion. Statistics in Medicine, 12(5-6):521–525, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:PDO

[708] Anonymous. Panel discussion: The operation of data monitoring committees. Statistics in Medicine, 12(5-6):527–542, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geller:1993:PID

[709] Nancy L. Geller and Mario Stylianou. Practical issues in data monitoring of clinical trials: Summary of responses to a questionnaire at nih. Statistics in Medicine, 12(5-6):543–551, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Df

[710] Anonymous. Discussion. Statistics in Medicine, 12(5-6):553, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Herson:1993:DMB

[711] Jay Herson. Data monitoring boards in the pharmaceutical industry. Statistics in Medicine, 12(5-6):555–561, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dg

[712] Anonymous. Discussion. Statistics in Medicine, 12(5-6):563, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleming:1993:DMC

[713] Thomas R. Fleming. Data monitoring committees and capturing relevant information of high quality. Statistics in Medicine, 12(5-6):565–570, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dh

[714] Anonymous. Discussion. Statistics in Medicine, 12(5-6):571–573, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walters:1993:DMC

[715] Leroy Walters. Data monitoring committees: The moral case for maximum feasible independence. Statistics in Medicine, 12(5-6):575–580, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Di

[716] Anonymous. Discussion. Statistics in Medicine, 12(5-6):581–582, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:PDS

[717] Anonymous. Panel discussion: Scientific issues in data monitoring. Statistics in Medicine, 12(5-6):583–600, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oneill:1993:SFP

[718] Robert T. O’neill. Some FDA perspectives on data monitoring in clinical trials in drug development. Statistics in Medicine, 12(5-6):601–608, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dj

[719] Anonymous. Discussion. Statistics in Medicine, 12(5-6):609–614, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:CR

[720] Anonymous. Closing remarks. Statistics in Medicine, 12(5-6):615–616, March 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Me

[721] Anonymous. Masthead. Statistics in Medicine, 12(7):fmi, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marion:1993:UBE

[722] Stephen A. Marion and Martin T. Schechter. Use of backcalculation for estimation of the probability of progression from HIV infection to AIDS. Statistics in Medicine, 12(7):617–631, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bradstreet:1993:UOP

[723] Thomas E. Bradstreet. Using orthogonal polynomial scores in summarizing and evaluating longitudinal data collected in phase I and II clinical pharmacology studies. Statistics in Medicine, 12(7):633–643, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hung:1993:TST

[724] H. M. James Hung. Two-stage tests for studying monotherapy and combination therapy in two-by-two factorial trials. Statistics in Medicine, 12 (7):645–660, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hadjinicola:1993:MCM

[725] George Hadjinicola and Larry Goldstein. Markov chain modelling of bioassay toxicity procedures. Statistics in Medicine, 12(7):661–674, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1993:VAR

[726] Young J. Lee and Hui Quan. P -values after repeated significance testing: a simple approximation method. Statistics in Medicine, 12(7):675–684, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

George:1993:CLB

[727] Varghese T. George and Robert C. Elston. Confidence limits based on the first occurrence of an event. Statistics in Medicine, 12(7):685–690, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1809, 2126, 2127].

Blackwelder:1993:SSP

[728] William C. Blackwelder. Sample size and power for prospective analysis of relative risk. Statistics in Medicine, 12(7):691–698, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1993:ECU

[729] Patrick Royston. Exact conditional and unconditional sample size for pair-matched studies with binary outcome: a practical guide. Statistics in Medicine, 12(7):699–712, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyse:1993:BRI

[730] Marc Buyse. Book review: Introducing new treatment for cancer: Practical, Ethical and legal Problems, C. J. Williams (ed), Wiley, Chichester, 1992. No. of pages: xiv + 492. Price: £24.95 (paper), 0-471-93113-6 (cloth). Statistics in Medicine, 12(7):713–714, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunn:1993:BRM

[731] Graham Dunn. Book review: Morphometric Tools for Landmark Data: Geometry and Biology. Fred L. Bookstein, Cambridge University Press, Cambridge, 1991. No. of pages: xvii + 435, Price: £50.00/$89.95. ISBN 0-521-38385-4. Statistics in Medicine, 12(7):714–715, April 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mf

[732] Anonymous. Masthead. Statistics in Medicine, 12(8):fmi, April 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1993:MEA

[733] Sander Greenland. Methods for epidemiologic analyses of multiple exposures: a review and comparative study of maximum-likelihood, preliminary-testing, and empirical-Bayes regression. Statistics in Medicine, 12(8):717–736, April 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pepe:1993:KMM

[734] Margaret Sullivan Pepe and Motomi Mori. Kaplan–Meier, marginal or conditional probability curves in summarizing competing risks failure time data? Statistics in Medicine, 12(8):737–751, April 30, 1993. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lan:1993:SMC

[735] K. K. Gordon Lan and David M. Zucker. Sequential monitoring of clinical trials: The role of information and Brownian motion. Statistics in Medicine, 12(8):753–765, April 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tallis:1993:CLT

[736] G. M. Tallis, P. Leppard, and T. J. O’Neill. Conditioned life tables from registries with unidentified random losses. Statistics in Medicine, 12(8): 767–775, April 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis:1993:PTP

[737] Alan B. Davis. Power of testing proportions in small two-sample studies when sample sizes are equal. Statistics in Medicine, 12(8):777–787, April 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Obrams:1993:ADT

[738] G. Iris Obrams and Marthana Hjortland. The application of developments in theoretical biostatistics to epidemiologic research. Statistics in Medicine, 12(8):789–793, April 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Doll:1993:SAB

[739] Richard Doll. Sir Austin Bradford Hill, 1897–1991. Statistics in Medicine, 12(8):795–808, April 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mg

[740] Anonymous. Masthead. Statistics in Medicine, 12(9):fmi, May 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vollset:1993:CIB

[741] Stein Emil Vollset. Confidence intervals for a binomial proportion. Statistics in Medicine, 12(9):809–824, May 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McKnight:1993:CAT

[742] Barbara McKnight and Stephen K. Van Den Eeden. A conditional analysis for two-treatment multiple-period crossover designs with binomial or Poisson outcomes and subjects who drop out. Statistics in Medicine, 12 (9):825–834, May 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:1993:ERC

[743] D. Y. Lin, M. A. Fischl, and D. A. Schoenfeld. Evaluating the role of CD4-lymphocyte counts as surrogate endpoints in human immunodeficiency virus clinical trials. Statistics in Medicine, 12(9):835–842, May 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fusaro:1993:MDH

[744] Robert E. Fusaro, Jens P. Nielsen, and Thomas H. Scheike. Markerdependent hazard estimation: an application to AIDS. Statistics in Medicine, 12(9):843–865, May 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weinberg:1993:PIR

[745] Clarice R. Weinberg, Donna Day Baird, and Andrew S. Rowland. Pitfalls inherent in retrospective time-to-event studies: The example of time to pregnancy. Statistics in Medicine, 12(9):867–879, May 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Devidas:1993:GLM

[746] Meenakshi Devidas, E. Olusegun George, and Daniel Zelterman. Generalized logistic models for low-dose response data. Statistics in Medicine, 12(9):881–892, May 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roe:1993:TPE

[747] Denise J. Roe and Edward L. Korn. Time-period effects in longitudinal studies measuring average rates of change. Statistics in Medicine, 12(9): 893–900, May 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mh

[748] Anonymous. Masthead. Statistics in Medicine, 12(10):fmi, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1993:SGC

[749] Michael D. Hughes. Stopping guidelines for clinical trials with multiple treatments. Statistics in Medicine, 12(10):901–915, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1993:CAR

[750] Douglas G. Altman. Construction of age-related reference centiles using absolute residuals. Statistics in Medicine, 12(10):917–924, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Plummer:1993:MEDa

[751] Martyn Plummer and David Clayton. Measurement error in dietary assessment: an investigation using covariance structure models. Part I. Statistics in Medicine, 12(10):925–935, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Plummer:1993:MEDb

[752] Martyn Plummer and David Clayton. Measurement error in dietary assessment: an investigation using covariance structure models. Part II. Statistics in Medicine, 12(10):937–948, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Drewett:1993:MLU

[753] R. F. Drewett and H. Goldstein. Modelling lactation using an inverse polynomial in a multilevel statistical model. Statistics in Medicine, 12 (10):949–954, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collett:1993:MAA

[754] D. Collett and N. Weerasooriya. A modelling approach to the analysis of menstrual diary data. Statistics in Medicine, 12(10):955–965, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Percy:1993:MAC

[755] David F. Percy and Trevor J. Hine. Multivariate analysis of cholesterol distribution for monitoring the risk of coronary heart disease. Statistics in Medicine, 12(10):967–974, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:1993:CRM

[756] Bernard F. Cole, Richard D. Gelber, and Aron Goldhirsch. Cox regression models for quality adjusted survival analysis. Statistics in Medicine, 12(10):975–987, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alho:1993:LEC

[757] Juha M. Alho and Salomon Minkin. Letters to the Editor on the computation of likelihood ratio and score test based confidence intervals in generalized linear models. Statistics in Medicine, 12(10):989, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1993:BRE

[758] Douglas G. Altman. Book review: Envisioning information. Edward R. Tufte, Graphics Press, Cheshire, Connecticut, 1990. no. of pages: 126. price: £30/$48. ISBN: 0-961-39211-8. Statistics in Medicine, 12(10): 991–992, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldstein:1993:BRS

[759] Harvey Goldstein. Book review: Statistical models for longitudinal studies of health. James H. Dwyer, Manning Feinleib, Peter Lippert, and Hans Hoffmeister (Eds.), Oxford University Press, New York, 1992. no. of pages: xi + 385. price: £50. ISBN: 0-19-505473-3. Statistics in Medicine, 12(10):992, May 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friedlander:1993:CDM

[760] Yechiel Friedlander, Dafna Lev Merom, and Jeremy D. Kark. A comparison of different matching designs in case-control studies: an empirical example using continuous exposures, continuous confounders and incidence of myocardial infarction. Statistics in Medicine, 12(11):993–1004, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leigh:1993:RAB

[761] J. Paul Leigh, Michael M. Ward, and James F. Fries. Reducing attrition bias with an instrumental variable in a regression model: Results from a panel of rheumatoid arthritis patients. Statistics in Medicine, 12(11):

1005–1018, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Segal:1993:RIM

[762] Mark R. Segal and John M. Neuhaus. Robust inference for multivariate survival data. Statistics in Medicine, 12(11):1019–1031, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ciminera:1993:TQT

[763] Joseph L. Ciminera, Joseph F. Heyse, Ha H. Nguyen, and John W. Tukey. Tests for qualitative treatment-by-centre interaction using a ‘pushback’ procedure. Statistics in Medicine, 12(11):1033–1045, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ciminera:1993:EMC

[764] Joseph L. Ciminera, Joseph F. Heyse, Ha H. Nguyen, and John W. Tukey. Evaluation of multicentre clinical trial data using adaptations of the Mosteller–Tukey procedure. Statistics in Medicine, 12(11):1047– 1061, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:1993:CPS

[765] Emmanuel Lesaffre, Ilse Scheys, Jürgen Fröhlich, and Erich Bluhmki. Calculation of power and sample size with bounded outcome scores. Statistics in Medicine, 12(11):1063–1078, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flack:1993:SSD

[766] Virginia F. Flack and T. Lynn Eudey. Sample size determinations using logistic regression with pilot data. Statistics in Medicine, 12(11):1079– 1084, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:ICG

[767] Anonymous. International Collaborative Group on Clinical Trial Registries (ICG-CTR) consensus recommendations. Statistics in Medicine, 12(11):1085, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1993:RTM

[768] Stephen Senn. Regression toward the mean in 2 × 2 crossover designs with baseline measurements. Statistics in Medicine, 12(11):1086–1087, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grender:1993:AR

[769] Julie Myers Grender, William D. Johnson, and Robert C. Elston. Authors’ reply. Statistics in Medicine, 12(11):1088–1089, June 1993. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1993:BRBb

[770] Douglas Altman and Eugenio Paci. Book review: A. B. Miller, J. Chamberlain, N. E. Day, M. Hakama and P. C. Prorok (eds.), Cancer screening:

1. Cancer screening. Edited by A. B. Miller, J. Chamberlain, N. E. Day, M. Hakama and P. C. Prorok. Cambridge, Cambridge University Press (on behalf of UICC), 1991. No. of pages: xvi + 438. Price: £45/$85. ISBN: 0-521-41041-X. Statistics in Medicine, 12(11):1091–1092, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chevret:1993:CRM

[771] Sylvie Chevret. The continual reassessment method in cancer phase I clinical trials: a simulation study. Statistics in Medicine, 12(12):1093– 1108, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schall:1993:PIB

[772] Robert Schall and Herman G. Luus. On population and individual bioequivalence. Statistics in Medicine, 12(12):1109–1124, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1993:NAS

[773] Kung-Jong Lui and Duane Steffey. A note on the application of simple linear regression methods for trend detection at multiple sites and visits. Statistics in Medicine, 12(12):1125–1139, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmid:1993:BAL

[774] Christopher H. Schmid and Bernard Rosner. A Bayesian approach to logistic regression models having measurement error following a mixture distribution. Statistics in Medicine, 12(12):1141–1153, June 1993. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sharples:1993:UGS

[775] Linda D. Sharples. Use of the Gibbs sampler to estimate transition rates between grades of coronary disease following cardiac transplantation. Statistics in Medicine, 12(12):1155–1169, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1993:IDA

[776] P. W. Jones, M. F. M. Ziade, M. J. Davis, and P. T. Dawes. An index of disease activity in rheumatoid arthritis. Statistics in Medicine, 12 (12):1171–1181, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altaian:1993:VRD

[777] Douglas Altaian and Deborah Ashby. Video review: 1. D. R. Appleton (producer), Designing a clinical trial: I. Designing a clinical trial. Video produced by D. R. Appleton, University of Newcastle, 1991. Running time: 17 mins. Price: £37, plus p and p and VAT. Statistics in Medicine, 12(12):1183–1184, June 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mi

[778] Anonymous. Masthead. Statistics in Medicine, 12(13):fmi, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peduzzi:1993:ARP

[779] Peter Peduzzi, Janet Wittes, Katherine Detre, and Theodore Holford. Analysis as-randomized and the problem of non-adherence: an example from the veterans affairs randomized trial of coronary artery bypass surgery. Statistics in Medicine, 12(13):1185–1195, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:1993:BSS

[780] Peter F. Thall and Elihu H. Estey. A Bayesian strategy for screening cancer treatments prior to Phase II clinical evaluation. Statistics in Medicine, 12(13):1197–1211, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chambless:1993:MAD

[781] Lloyd E. Chambless and John R. Roeback. Methods for assessing difference between groups in change when initial measurement is subject to intra-individual variation. Statistics in Medicine, 12(13):1213–1237, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See comments [1152, 1424].

Piantadosi:1993:CPT

[782] S. Piantadosi and M. H. Gail. A comparison of the power of two tests for qualitative interactions. Statistics in Medicine, 12(13):1239–1248, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:1993:AEM

[783] A. B. Lawson and F. L. R. Williams. Applications of extraction mapping in environmental epidemiology. Statistics in Medicine, 12(13):1249–1258, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neuhaus:1993:DEB

[784] John M. Neuhaus and Mark R. Segal. Design effects for binary regression models fitted to dependent data. Statistics in Medicine, 12(13):1259– 1268, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:C

[785] Anonymous. Correction. Statistics in Medicine, 12(13):1269, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fayers:1993:BRP

[786] Peter Fayers. Book review: Practical handbook of sample size guidelines for clinical trials. Jonathon J. Shuster, CRC Press Inc., Boca Raton, Florida, 1993. no. of pages: 212, plus program on diskettes for IBM-PC (MSDOS) or Macintosh. price: £51. ISBN: 0-8493-4487-5. Statistics in Medicine, 12(13):1271–1272, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shanmugam:1993:BRS

[787] Ramalingam Shanmugam. Book review: Statistics in medicine. Edited by F. Dunstan and J. Pickles. Oxford, Clarendon Press, 1991. no. of pages: xii + 266. price: $75. ISBN: 0-19-853662-3. Statistics in Medicine, 12(13):1272–1273, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1993:BRO

[788] David R. Jones. Book review: Oxford textbook of public health. Volume 2: Methods of public health. 2nd edition. Edited by Walter W. Holland, Roger Detels and George Knox, Oxford University Press, Oxford, 1991. number of pages: xxii + 563. price: £70. ISBN: 0-19-261707-9. Statistics in Medicine, 12(13):1273–1274, ???? 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mj

[789] Anonymous. Masthead. Statistics in Medicine, 12(14):fmi, July 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1993:VSA

[790] S. D. Walter. Visual and statistical assessment of spatial clustering in mapped data. Statistics in Medicine, 12(14):1275–1291, July 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moses:1993:CIS

[791] Lincoln E. Moses, David Shapiro, and Benjamin Littenberg. Combining independent studies of a diagnostic test into a summary ROC curve: Data-analytic approaches and some additional considerations. Statistics in Medicine, 12(14):1293–1316, July 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Commenges:1993:EBA

[792] D. Commenges and B. Etcheverry. An empirical Bayes approach to the estimation of the incidence curve of HIV infection. Statistics in Medicine, 12(14):1317–1324, July 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blettner:1993:IMB

[793] Maria Blettner and Willi Sauerbrei. Influence of model-building strategies on the results of a case-control study. Statistics in Medicine, 12(14): 1325–1338, July 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stukel:1993:CMA

[794] Thérèse A. Stukel. Comparison of methods for the analysis of longitudinal interval count data. Statistics in Medicine, 12(14):1339–1351, July 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Com-Nougue:1993:HEE

[795] C. Com-Nougue, C. Rodary, and C. Patte. How to establish equivalence when data are censored: a randomized trial of treatments for B non-Hodgkin lymphoma. Statistics in Medicine, 12(14):1353–1364, July 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yateman:1993:USD

[796] Nigel A. Yateman and Allan M. Skene. The use of simulation in the design of two cardiovascular survival studies. Statistics in Medicine, 12 (14):1365–1372, July 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mk

[797] Anonymous. Masthead. Statistics in Medicine, 12(15-16):fmi, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:1993:P

[798] Deborah Ashby. Preface. Statistics in Medicine, 12(15-16):1373–1374, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:LPS

[799] Anonymous. List of participatns, speakers and discussants. Statistics in Medicine, 12(15-16):1375–1376, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:1993:CBC

[800] Donald A. Berry. A case for Bayesianism in clinical trials. Statistics in Medicine, 12(15-16):1377–1393, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dk

[801] Anonymous. Discussion. Statistics in Medicine, 12(15-16):1395–1404, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1993:CFC

[802] John Whitehead. The case for frequentism in clinical trials. Statistics in Medicine, 12(15-16):1405–1413, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:CAF

[803] Anonymous. The case against frequentism. Statistics in Medicine, 12 (15-16):1415–1419, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Urbach:1993:VRC

[804] Peter Urbach. The value of randomization and control in clinical trials. Statistics in Medicine, 12(15-16):1421–1431, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dl

[805] Anonymous. Discussion. Statistics in Medicine, 12(15-16):1433–1441, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freeman:1993:RVA

[806] Peter R. Freeman. The role of P -values in analysing trial results. Statistics in Medicine, 12(15-16):1443–1452, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dm

[807] Anonymous. Discussion. Statistics in Medicine, 12(15-16):1453–1458, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:1993:SEI

[808] Stuart J. Pocock. Statistical and ethical issues in monitoring clinical trials. Statistics in Medicine, 12(15-16):1459–1469, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dn

[809] Anonymous. Discussion. Statistics in Medicine, 12(15-16):1471–1475, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schaffner:1993:CTC

[810] Kenneth F. Schaffner. Clinical trials and causation: Bayesian perspectives. Statistics in Medicine, 12(15-16):1477–1494, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Do

[811] Anonymous. Discussion. Statistics in Medicine, 12(15-16):1495–1499, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Spiegelhalter:1993:ABI

[812] David J. Spiegelhalter, Laurence S. Freedman, and Mahesh K. B. Parmar. Applying Bayesian ideas in drug development and clinical trials. Statistics in Medicine, 12(15-16):1501–1511, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Dp

[813] Anonymous. Discussion. Statistics in Medicine, 12(15-16):1513–1517, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meier:1993:IRA

[814] Paul Meier. Illusion and reality in the analysis of clinical trials. Statistics in Medicine, 12(15-16):1519, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:CD

[815] Anonymous. Chairman’s discussion. Statistics in Medicine, 12(15-16): 1521, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:GD

[816] Anonymous. General discussion. Statistics in Medicine, 12(15-16):1523– 1531, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1993:CR

[817] Peter Armitage. Closing remarks. Statistics in Medicine, 12(15-16): 1533–1534, August 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Ml

[818] Anonymous. Masthead. Statistics in Medicine, 12(17):fmi, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blumenstein:1993:VKM

[819] Brent A. Blumenstein. Verifying keyed medical research data. Statistics in Medicine, 12(17):1535–1542, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baeten:1993:RMR

[820] S. Baeten, A. Bouckaert, E. Loumaye, and K. Thomas. A regression model for the rate of success of in vitro fertilization. Statistics in Medicine, 12(17):1543–1553, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1993:MSC

[821] Paul S. Albert. A model for seasonal changes in time series regression relationships: With an application in psychiatry. Statistics in Medicine, 12(17):1555–1568, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:1993:FRN

[822] Odd O. Aalen. Further results on the non-parametric linear regression model in survival analysis. Statistics in Medicine, 12(17):1569–1588, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dorey:1993:MIT

[823] Frederick J. Dorey, Roderick J. A. Little, and Nathaniel Schenker. Multiple imputation for threshold-crossing data with interval censoring. Statistics in Medicine, 12(17):1589–1603, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mark:1993:ECE

[824] Steven D. Mark and James M. Robins. Estimating the causal effect of smoking cessation in the presence of confounding factors using a rank preserving structural failure time model. Statistics in Medicine, 12(17): 1605–1628, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byar:1993:IFD

[825] David P. Byar, Agnes M. Herzberg, and Wai-Yuan Tan. Incomplete factorial designs for randomized clinical trials. Statistics in Medicine, 12 (17):1629–1641, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fayers:1993:LE

[826] Peter Fayers. Letters to the editor. Statistics in Medicine, 12(17):1643, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1993:BRS

[827] Per Kragh Andersen. Book review: Survival analysis: State of the art: Proceedings of the NATO Advanced Research Workshop on Survival Analysis and Related Topics, Columbus, Ohio. John P. Klein and Prem K. Goel Eds. Kluwer Academic, Dordrecht, 1992. No. of pages: x + 451. Price: £79, $138. ISBN: 0-7923-1634-7. Statistics in Medicine, 12(17):1645, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailar:1993:BRH

[828] John C. Bailar III. Book review: Helath, lifestyle and environment: Countering the panic. Social Affairs Unit, London and Manhattan Institute, New York City, 1991. No. of pages: 152. ISBN: 0-907-631-444. Statistics in Medicine, 12(17):1645–1647, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jagger:1993:BRA

[829] Carol Jagger. Book review: Applied computational statistics in longitudinal research. M. J. Rovine and A. von Eye, Academic Press, London, 1991. No. of pages: xiii + 237. Price: $34.50. ISBN: 0-12-599450-8. Statistics in Medicine, 12(17):1647–1648, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tyson:1993:BRG

[830] Jon Tyson and Kathleen Kennedy. Book review: Guide to clinical trials. Bert Spilker, Raven Press. New York, 1991. No. of pages: xxv + 1156. Price: $162.50. ISBN: 0-88167-767-1. Statistics in Medicine, 12(17):1648, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:ASW

[831] Anonymous. Announcement: The scientific work of Martin Gardner. Wellcome Building, London, 21 October 1993. Statistics in Medicine, 12(17):1649, September 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mm

[832] Anonymous. Masthead. Statistics in Medicine, 12(18):fmi, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1993:RBA

[833] Michael D. Hughes. Reporting Bayesian analyses of clinical trials. Statistics in Medicine, 12(18):1651–1663, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Selwyn:1993:ALG

[834] Murray R. Selwyn and Donna M. Difranco. The application of large Gaussian mixed models to the analysis of 24 hour ambulatory blood pressure monitoring data in clinical trials. Statistics in Medicine, 12(18): 1665–1682, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1993:DTA

[835] Philip J. Smith. Dispersion tests and adjustments for survival and casecontrol studies. Statistics in Medicine, 12(18):1683–1691, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldberg:1993:SAS

[836] Jack Goldberg, William G. Henderson, Seth A. Eisen, William True, Viswanathan Ramakrishnan, Michael J. Lyons, and Ming T. Tsuang. A strategy for assembling samples of adult twin pairs in the United States. Statistics in Medicine, 12(18):1693–1702, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:1993:CIM

[837] Sylvia Richardson and Walter R. Gilks. Conditional independence models for epidemiological studies with covariate measurement error. Statistics in Medicine, 12(18):1703–1722, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:1993:CGE

[838] Taesung Park. A comparison of the generalized estimating equation approach with the maximum likelihood approach for repeated measurements. Statistics in Medicine, 12(18):1723–1732, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schouten:1993:RRR

[839] Evert G. Schouten, Jacqueline M. Dekker, Frans J. Kok, Saskia Le Cessie, Hans C. Van Houwelingen, Jan Pool, and Jan P. Vandnbroucke. Risk ratio and rate ratio estimation in case-cohort designs: Hypertension and cardiovascular mortality. Statistics in Medicine, 12(18):1733–1745, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brant:1993:BRE

[840] Larry J. Brant. Book review: The epidemiologic study of the elderly. Robert B. Wallace and Robert F. Woolson Eds., Oxford University Press, New York, 1992. No. of pages: xi + 387. Price: £60. ISBN: 0-19-506120-9. Statistics in Medicine, 12(18):1747–1748, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knipschild:1993:BRI

[841] Paul Knipschild. Book review: Immunointervention in man. A critique of clinical trial methodology. Keith H. Wallace and Ronald A. Thompson, Oxford Univeristy Press, Oxford, 1991. No. of pages: viii + 221. Price: £60. ISBN: 0-19-261918-7. Statistics in Medicine, 12(18):1748– 1749, September 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mn

[842] Anonymous. Masthead. Statistics in Medicine, 12(19-20):fmi, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacquez:1993:I

[843] Geoffrey M. Jacquez. Introduction. Statistics in Medicine, 12(19-20): 1751, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kheifets:1993:CAP

[844] Leeka I. Kheifets. Cluster analysis: a perspective. Statistics in Medicine, 12(19-20):1755–1756, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1993:ADC

[845] Daniel Smith and Raymond Neutra. Approaches to disease cluster investigations in a state health department. Statistics in Medicine, 12(19-20):

1757–1762, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wartenberg:1993:SCP

[846] Daniel Wartenberg and Michael Greenberg. Solving the cluster puzzle: Clues to follow and pitfalls to avoid. Statistics in Medicine, 12(19-20): 1763–1770, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grimson:1993:DCM

[847] Roger C. Grimson. Disease clusters, maxima, and P -values. Statistics in Medicine, 12(19-20):1773–1794, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oden:1993:ADE

[848] Neal L. Oden. Assessing directional effects in spatial data. Statistics in Medicine, 12(19-20):1795–1805, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1993:APA

[849] R. Chen, R. R. Connelly, and N. Mantel. Analysing post-alarm data in a monitoring system in order to accept or reject the alarm. Statistics in Medicine, 12(19-20):1807–1812, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rayens:1993:PTI

[850] Mary Kay Rayens and Richard J. Kryscio. Properties of Tango’s index for detecting clustering in time. Statistics in Medicine, 12(19-20):1813–1827, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wallenstein:1993:PSS

[851] Sylvan Wallenstein, Joseph Naus, and Joseph Glaz. Power of the scan statistic for detection of clustering. Statistics in Medicine, 12(19-20): 1829–1843, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glaz:1993:ATP

[852] Joseph Glaz. Approximations for the tail probabilities and moments of the scan statistic. Statistics in Medicine, 12(19-20):1845–1852, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sahu:1993:ERR

[853] Sujit K. Sahu, Robert B. Bendel, and Cristina P. Sison. Effect of relative risk and cluster configuration on the power of the one-dimensional scan statistic. Statistics in Medicine, 12(19-20):1853–1865, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waller:1993:EST

[854] Lance A. Waller and Bruce W. Turnbull. The effects of scale on tests for disease clustering. Statistics in Medicine, 12(19-20):1869–1884, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1993:ASP

[855] S. D. Walter. Assessing spatial patterns in disease rates. Statistics in Medicine, 12(19-20):1885–1894, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heisterkamp:1993:DMU

[856] S. H. Heisterkamp, G. Doornbos, and M. Gankema. Disease mapping using empirical Bayes and Bayes methods on mortality statistics in The Netherlands. Statistics in Medicine, 12(19-20):1895–1913, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morris:1993:AEG

[857] Robert D. Morris and Rajika L. Munasinghe. Aggregation of existing geographic regions to diminish spurious variability of disease rates. Statistics in Medicine, 12(19-20):1915–1929, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacquez:1993:SCV

[858] Geoffrey M. Jacquez and Leeka I. Kheifets. Synthetic cancer variables and the construction and testing of synthetic risk maps. Statistics in Medicine, 12(19-20):1931–1942, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schlattmann:1993:MMD

[859] Peter Schlattmann and Dankmar Böhning. Mixture models and disease mapping. Statistics in Medicine, 12(19-20):1943–1950, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rizzardi:1993:IUC

[860] M. Rizzardi, M. S. Mohr, D. W. Merrill, and S. Selvin. Interfacing U.S. census map files with statistical graphics software: Application and use in epidemiology. Statistics in Medicine, 12(19-20):1953–1964, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schlattmann:1993:CPP

[861] Peter Schlattmann and Dankmar Böhning. Computer packages C.A.MAN computer assisted mixture analysis and dismap. Statistics in Medicine, 12(19-20):1965, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacquez:1993:SDD

[862] Geoffrey M. Jacquez. The statistical description of disease clusters. Statistics in Medicine, 12(19-20):1967–1968, October 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mo

[863] Anonymous. Masthead. Statistics in Medicine, 12(21):fmi, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agresti:1993:DFF

[864] Alan Agresti. Distribution-free fitting of logit models with random effects for repeated categorical responses. Statistics in Medicine, 12(21):1969– 1987, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stromberg:1993:CCM

[865] Ulf Strömberg. Controlling for a confounder monotonically related to exposure by means of isotonic quantile regression. Statistics in Medicine, 12(21):1989–1998, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:1993:TSS

[866] Jun Shao and Shein-Chung Chow. TWO-stage sampling in pharmaceutical applications. Statistics in Medicine, 12(21):1999–2008, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:1993:SSE

[867] A. Lawrence Gould. Sample sizes for event rate equivalence trials using prior information. Statistics in Medicine, 12(21):2009–2023, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Normolle:1993:ARN

[868] Daniel P. Normolle. AN algorithm for robust non-linear analysis of radioimmunoassays and other bioassays. Statistics in Medicine, 12(21): 2025–2042, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:1993:TCB

[869] Mark C. K. Yang and P. V. Rao. Testing concomitancy between two physiological pulse series. Statistics in Medicine, 12(21):2043–2055, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmer:1993:BRD

[870] Judy L. Palmer. Book review: The design and analysis of sequential clinical trials, 2nd edn. John Whitehead, Ellis Horwood, New York, 1992. No. of pages: 298. Price: £48. ISBN 0-13-202565-5. Statistics in Medicine, 12(21):2057–2059, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Day:1993:BRB

[871] Simon Day. Book review: Biopharmaceutical sequential statistical applications. Karl E. Peace (Ed.). Marcel Dekker, New York, 1992. No. of Pages: xx + 348. Price: $110. ISBN 0-8247-8268-9. Statistics in Medicine, 12(21):2059, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grieve:1993:BRD

[872] A. P. Grieve. Book review: Design and analysis of bioavailability and bioequivalence studies. Shein-Chung Chow and Jen-Pei Liu, Marcel Dekker, New York, 1992. No. of pages: x + 416. Price: $125. ISBN 0-8247-8682-3. Statistics in Medicine, 12(21):2059–2060, November 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mp

[873] Anonymous. Masthead. Statistics in Medicine, 12(22):fmi, November 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gore:1993:EPH

[874] S. M. Gore and V. T. Farewell. Extending public health surveillance of HIV disease. Statistics in Medicine, 12(22):2061–2064, November 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:EPH

[875] Anonymous. Extending public health surveillance of HIV infection: Information from a five cohort workshop. Statistics in Medicine, 12(22): 2065–2085, November 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:MTD

[876] Anonymous. Markers as time-dependent covariates in relative risk regression. Statistics in Medicine, 12(22):2087–2098, November 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:MPa

[877] Anonymous. Marker paths. Statistics in Medicine, 12(22):2099–2126, November 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:MMa

[878] Anonymous. Markov models. Statistics in Medicine, 12(22):2127–2130, November 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Galai:1993:TMO

[879] Noya Galai, Alvaro Muñoz, Kan Chen, Vincent J. Carey, Joan Chmiel, and Susan Y. J. Zhou. Tracking of markers and onset of disease among HIV-1 seroconverters. Statistics in Medicine, 12(22):2133–2145, November 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1993:LER

[880] Stephen Senn, Farkad Ezzet, and John Whitehead. Letter to the Editor: A random effects model for ordinal responses from a crossover trial, by F. Ezzet and J. Whitehead, Statistics in Medicine, 10, 901–907 (1991). Statistics in Medicine, 12(22):2147–2151, November 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [296].

Anonymous:1993:Mq

[881] Anonymous. Masthead. Statistics in Medicine, 12(23):fmi, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Segal:1993:TT

[882] Mark Robert Segal and Ira B. Tager. Trees and tracking. Statistics in Medicine, 12(23):2153–2168, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hathaway:1993:TSL

[883] Dale K. Hathaway and Ralph B. D’agostino. A technique for summarizing longitudinal data. Statistics in Medicine, 12(23):2169–2178, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raab:1993:BAB

[884] Gillian M. Raab and Robert A. Elton. Bayesian analysis of binary data from an audit of cervical smears. Statistics in Medicine, 12(23):2179– 2189, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guggenmoos-Holzmann:1993:HRC

[885] Irene Guggenmoos-Holzmann. HOW reliable are change-corrected measures of agreement? Statistics in Medicine, 12(23):2191–2205, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schouten:1993:EKB

[886] Hubert J. A. Schouten. Estimating kappa from binocular data and comparing marginal probabilities. Statistics in Medicine, 12(23):2207–2217, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lan:1993:USF

[887] K. K. Gordon Lan, William F. Rosenberger, and John M. Lachin. Use of spending functions for occasional or continuous monitoring of data in clinical trials. Statistics in Medicine, 12(23):2219–2231, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blair:1993:CTT

[888] R. Clifford Blair and Shlomo Sawilowsky. Comparison of two tests useful in situations where treatment is expected to increase variability relative to controls. Statistics in Medicine, 12(23):2233–2243, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Obrien:1993:D

[889] Peter C. O’brien. Discussion. Statistics in Medicine, 12(23):2245–2246, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osman:1993:LEA

[890] Magued I. Osman. Letter to the Editor: The area between curves ABC — Measure in nutritional anthropometry. D. Böhning, A. Hempfling, F.-P. Schelp and P. Schlattmann, Statistics in Medicine, 11, 1289–1304 1992. Statistics in Medicine, 12(23):2247–2248, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [559].

Harris:1993:LES

[891] R. I. Harris, A. Heyting, J. T. M. Tolboom, and J. G. A. Essers. Letter to the Editor: Statistical handling of drop-outs in longitudinal clinical trials, by Heyting, A.,Tolboom, J. T. B. M. and Essers, J. G. A. Statistics in Medicine, 11, 2043–2061 (1992). Statistics in Medicine, 12(23):2248– 2250, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [636].

Fowler:1993:BRS

[892] Godfrey Fowler. Book review: The strategy of preventive medicine, Geoffrey Rose, Oxford University Press, Oxford, 1992. No. of pages: xii + 135. Price: £17.50. ISBN 0-19-262125-4. Statistics in Medicine, 12(23): 2251, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verhulst:1993:BRS

[893] Steven J. Verhulst. Book review: Statistical first aid: Interpretation of health research data. Robert P. Hirsch and Richard K. Riegelman, Blackwell Scientific Publications, Oxford, 1992. No. of pages: xii + 409. Price: £25. ISBN 0-869542-138-2. Statistics in Medicine, 12(23):2252, December 15, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:Mr

[894] Anonymous. Masthead. Statistics in Medicine, 12(24):fmi, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1993:P

[895] Ted Colton, Laurence Freedman, Tony Johnson, and David Machin. Preface. Statistics in Medicine, 12(24):2253, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:ISC

[896] Anonymous. International Society for Clinical Biostatistics Thirteenth International Meeting held at H. C. Ørsted Institute Copenhagen, 17–21 August 1992. Statistics in Medicine, 12(24):2255, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1993:SSC

[897] John Whitehead. Sample size calculations for ordered categorical data. Statistics in Medicine, 12(24):2257–2271, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanHouwelingen:1993:BAM

[898] Hans C. Van Houwelingen, Koos H. Zwinderman, and Theo Stijnen. A bivariate approach to meta-analysis. Statistics in Medicine, 12(24):2273– 2284, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stijnen:1993:RRR

[899] Theo Stijnen and Hans C. Van Houwelingen. Relative risk, risk difference and rate difference models for sparse stratified data: a pseudo likelihood approach. Statistics in Medicine, 12(24):2285–2303, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verweij:1993:CVS

[900] Pierre J. M. Verweij and Hans C. Van Houwelingen. Cross-validation in survival analysis. Statistics in Medicine, 12(24):2305–2314, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klein:1993:PSP

[901] John P. Klein, Niels Keiding, and Edward A. Copelan. Plotting summary predictions in multistate survival models: Probabilities of relapse and death in remission for bone marrow transplantation patients. Statistics in Medicine, 12(24):2315–2332, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:1993:ACC

[902] Roger J. Marshall and Rodney T. Jackson. Analysis of case-crossover designs. Statistics in Medicine, 12(24):2333–2341, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Signorini:1993:DBR

[903] D. F. Signorini, O. Leung, R. J. Simes, E. Beller, V. J. Gebski, and T. Callaghan. Dynamic balanced randomization for clinical trials. Statistics in Medicine, 12(24):2343–2350, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmoor:1993:CCM

[904] Claudia Schmoor, Kurt Ulm, and Martin Schumacher. Comparison of the Cox model and the regression tree procedure in analysing a randomized clinical trial. Statistics in Medicine, 12(24):2351–2366, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1993:IDA

[905] Stephen Senn. Inherent difficulties with active control equivalence studies. Statistics in Medicine, 12(24):2367–2375, December 30, 1993. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:1993:RIP

[906] M. Schemper. The relative importance of prognostic factors in studies of survival. Statistics in Medicine, 12(24):2377–2382, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Praus:1993:ASP

[907] Michaela Praus, Fritz Schindel, Reinhard Fescharek, and Sabine Schwarz. Alert systems for post-marketing surveillance of adverse drug reactions. Statistics in Medicine, 12(24):2383–2393, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Capocaccia:1993:RBI

[908] Riccardo Capocaccia. Relationships between incidence and mortality in non-reversible diseases. Statistics in Medicine, 12(24):2395–2415, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1993:A

[909] Anonymous. Acknowledgement. Statistics in Medicine, 12(24):2417– 2418, December 30, 1993. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Ma

[910] Anonymous. Masthead. Statistics in Medicine, 13(1):fmi, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colto:1994:EA

[911] Theodore Colto, Tony Johnson, and David Machin. Editorial announcement. Statistics in Medicine, 13(1):1–2, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1994:SMC

[912] M. J. Campbell and S. A. Julious. Statistics in Medicine: Citations of papers in the first ten years. Statistics in Medicine, 13(1):3–10, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:1994:CSS

[913] Philip Prescott, Joyce Ryan, and Paul Shuttleworth. A comparative study of several antibiotic formulations using a design based on a combination of balanced incomplete blocks and Latin squares. Statistics in Medicine, 13(1):11–21, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holford:1994:EAP

[914] Theodore R. Holford, Zhongxin Zhang, and Lisa A. McKay. Estimating age, period and cohort effects using the multistage model for cancer. Statistics in Medicine, 13(1):23–41, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eriksson:1994:ALT

[915] Bo Eriksson and Ragnar Adell. On the analysis of life tables for dependent observations. Statistics in Medicine, 13(1):43–51, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altmann:1994:NAM

[916] Michael Altmann, Beeling C. Wee, Keith Willard, Denton Peterson, and Laël C. Gatewood. Network analytic methods for epidemiological risk assessment. Statistics in Medicine, 13(1):53–60, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feldman:1994:CVC

[917] Henry A. Feldman and Sonja M. McKinlay. Cohort versus crosssectional design in large field trials: Precision, sample size, and a unifying model. Statistics in Medicine, 13(1):61–78, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:1994:SSR

[918] Jun-Mo Nam. Sample size requirements for stratified prospective studies with null hypothesis of non-unity relative risk using the score test. Statistics in Medicine, 13(1):79–86, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Russek-Cohen:1994:SBD

[919] Estelle Russek-Cohen and Richard M. Simon. Selecting the best dose when a monotonic dose-response relation exists. Statistics in Medicine, 13(1):87–95, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1994:BRT

[920] S. D. Walter. Book review: Teaching epidemiology: what you should know and what you could do. J. Olsen and D. Trichopoulos (eds.), Oxford Medical Publications, Oxford, 1992. no. of pages: xiii + 340. price: £40. ISBN 0-19-262205-6. Statistics in Medicine, 13(1):97–98, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Preston-Martin:1994:BRA

[921] Susan Preston-Martin. Book review: Alcohol and illness: The epidemiological viewpoint. John C. Duffy (ed.), Edinburgh University Press, 1992. no. of pages: vi + 154. price: £25. ISBN 0-7486-0353-0. Statistics in Medicine, 13(1):98, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Ca

[922] Anonymous. Corrections. Statistics in Medicine, 13(1):99, January 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mb

[923] Anonymous. Masthead. Statistics in Medicine, 13(2):fmi, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1994:GST

[924] Jae Won Lee. Group sequential testing in clinical trials with multivariate observations: a review. Statistics in Medicine, 13(2):101–111, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bhargava:1994:EVA

[925] Alok Bhargava, Ronald Forthofer, Susie McPherson, and Milton Nichaman. Estimating the variations and autocorrelations in dietary intakes on weekdays and weekends. Statistics in Medicine, 13(2):113– 126, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaaks:1994:EAD

[926] Rudolf Kaaks, Elio Riboli, Jacques Estève, Anne Linda Van Kappel, and Wija A. Van Staveren. Estimating the accuracy of dietary questionnaire assessments: Validation in terms of structural equation models. Statistics in Medicine, 13(2):127–142, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipsitz:1994:RGF

[927] Stuart R. Lipsitz and John F. Buoncristiani. A robust goodness-of-fit test statistic with application to ordinal regression models. Statistics in Medicine, 13(2):143–152, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piegorsch:1994:NHL

[928] Walter W. Piegorsch, Clarice R. Weinberg, and Jack A. Taylor. Nonhierarchical logistic models and case-only designs for assessing susceptibility in population-based case-control studies. Statistics in Medicine, 13(2):153–162, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiarotti:1994:MTS

[929] Flavia Chiarotti, Massimo Palombi, Nicola Schinaia, Alessandro Ghirardini, and Rino Bellocco. Median time from seroconversion to AIDS in Italian HIV-positive haemophiliacs: Different parametric estimates.

Statistics in Medicine, 13(2):163–175, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Larsen:1994:ENP

[930] Severin Olesen Larsen. Estimation of the number of people in a register from the number of birthdates. Statistics in Medicine, 13(2):177–183, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Merle:1994:DOE

[931] Y. Merlé, F. Mentré, A. Mallet, and A. H. Aurengo. Designing an optimal experiment for Bayesian stimation: Application to the kinetics of iodine thyroid uptake. Statistics in Medicine, 13(2):185–196, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1994:RMCa

[932] Stephen Senn, Lars Frison, and Stuart Pocock. Repeated measures in clinical trials: Analysis using mean summary statistics and its implications for design. Statistics in Medicine, 13(2):197–198, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hopper:1994:BRM

[933] John L. Hopper. Book review: Methodology for genetic studies of twins and families. Michael C. Neale and Lon R. Cardon, Kluwer, Dordrecht, The Netherlands, 1992. no. of pages: xxv + 496. price: £99.00/$ 169.00. ISBN 0-7923-1874-9. Statistics in Medicine, 13(2):199–200, January 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mc

[934] Anonymous. Masthead. Statistics in Medicine, 13(3):fmi, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miettinen:1994:FMD

[935] Olli S. Miettinen and J. Jaime Caro. Foundations of medical diagnosis: What actually are the parameters involved in Bayes’ theorem? Statistics in Medicine, 13(3):201–209, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:1994:C

[936] Colin B. Begg. Commentary. Statistics in Medicine, 13(3):211–212, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miettinen:1994:R

[937] Olli S. Miettinen and J. Jaime Caro. Rejoinder. Statistics in Medicine, 13(3):213–215, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1994:FGD

[938] Stephen Senn. Fisher’s game with the devil. Statistics in Medicine, 13 (3):217–230, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dubin:1994:ETS

[939] Neil Dubin, Simeon Berman, Michael Marmor, Brett Tindall, Don Des Jarlais, and Mimi Kim. Estimation of time since infection using longitudinal disease-marker data. Statistics in Medicine, 13(3):231–244, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aickin:1994:CER

[940] Mikel Aickin, Cheryl Ritenbaugh, Earl Surwit, and Frank Meyskens, Jr. Comparative exposure ratios: a non-parametric, multifactor technique for case-control studies. Statistics in Medicine, 13(3):245–260, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:1994:GAA

[941] Weichung Joseph Shih. A geometric approach to the analysis of physiological flow data. Statistics in Medicine, 13(3):261–273, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Denby:1994:AWE

[942] Lorraine Denby, Colin L. Mallows, and Seymour G. Blank. Analysis of the wideband external pulse: an application of graphical methods. Statistics in Medicine, 13(3):275–291, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1994:ITI

[943] Stephen Senn and Ronald Brand. Importance of trends in the interpretation of an overall odds ratio in the meta-analysis of clinical trials. Statistics in Medicine, 13(3):293–296, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hutton:1994:BRI

[944] Jane L. Hutton. Book review: Interpretation and uses of medical statistics, fourth edition. Leslie E. Daly, Geoffrey J. Bourke and James Mcgilvray, Blackwell scientific Publications, OXFORD, 1991. no. of pages: xiv + 440. price: £24.95. ISBN 0-632-02911-0. Statistics in Medicine, 13(3):297–298, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:BRC

[945] Anonymous. Book review: Cross-over experiments. Design, analysis and application. David A. Ratkowsky, Marc A. Evans and J. Richard Alldredge, Marcel Dekker, New York, 1993. no. of pages: ix + 446. price: $110. ISBN 0-8247-8892-3. Statistics in Medicine, 13(3):298, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:1994:BRC

[946] J. N. S. Matthews. Book review: Cross-over. Trials in clinical research. Stephen Senn, Wiley, Chichester, 1993. no. of pages pages: xv + 266. price: £24.95. ISBN 0-471-93493-3. Statistics in Medicine, 13(3):298– 300, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1994:BRD

[947] David R. Jones. Book reviews: The design and analysis of research studies. Bryan F. J. Manly, Cambridge University Press, Cambridge, 1992. no. of pages: xvi + 353. price: £55/$89.95 hardback, ISBN 0-521-41453-9; £19.95/$37.95 paperback, ISBN 0-521-42580-8. Statistics in Medicine, 13(3):300, February 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Md

[948] Anonymous. Masthead. Statistics in Medicine, 13(4):fmi, February 28, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1994:PPF

[949] Douglas G. Altman and Bianca L. De Stavola. Practical problems in fitting a proportional hazards model to data with updated measurements of the covariates. Statistics in Medicine, 13(4):301–341, February 28, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lavange:1994:ASS

[950] Lisa M. Lavange, Lynette L. Keyes, Gary G. Koch, and Peter A. Margolis. Application of sample survey methods for modelling ratios to incidence densities. Statistics in Medicine, 13(4):343–355, February 28, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halloran:1994:EEC

[951] M. Elizabeth Halloran, Ira M. Longini, Jr., Michael J. Haber, Claudio J. Struchiner, and Robert C. Brunet. Exposure efficacy and change in contact rates in evaluating prophylactic HIV vaccines in the field. Statistics in Medicine, 13(4):357–377, February 28, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duffy:1994:DAC

[952] S. W. Duffy, T. E. Rohan, and J. R. McLaughlin. Design and analysis considerations in a cohort study involving repeated measurement of both exposure and outcome: The association between genital papillomavirus infection and risk of cervical intraepithelial neoplasia. Statistics in Medicine, 13(4):379–390, February 28, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhen:1994:SSD

[953] Boguang Zhen and James R. Murphy. Sample size determination for an exponential survival model with an unrestricted covariate. Statistics in Medicine, 13(4):391–397, February 28, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Me

[954] Anonymous. Masthead. Statistics in Medicine, 13(5-7):fmi, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1994:P

[955] Jonas H. Ellenberg, Mitchell H. Gail, and Richard M. Simon. Preface. Statistics in Medicine, 13(5-7):399, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:1994:I

[956] Samuel W. Greenhouse. Introduction. Statistics in Medicine, 13(5-7): 401, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Healy:1994:NBB

[957] Bernadine Healy. NIH and the birth of biostatistics. Statistics in Medicine, 13(5-7):403, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1994:ULS

[958] Susan S. Ellenberg and Mary A. Foulkes. The utility of large, simple trials in the evaluation of AIDS treatment strategies. Statistics in Medicine, 13(5-7):405–415, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simon:1994:NDS

[959] R. Simon, P. F. Thall, and S. S. Ellenberg. New designs for the selection of treatments to be tested in randomized clinical trials. Statistics in Medicine, 13(5-7):417–429, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1994:MDP

[960] T. Timothy Chen and Richard Simon. A multiple decision procedure in clinical trials. Statistics in Medicine, 13(5-7):431–446, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Da

[961] Anonymous. Discussion. Statistics in Medicine, 13(5-7):447–451, March– April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1994:TSM

[962] Paul S. Albert, Henry F. McFarland, Mary E. Smith, and Joseph A. Frank. Time series for modelling counts from a relapsing-remitting disease: Application to modelling disease activity in multiple sclerosis. Statistics in Medicine, 13(5-7):453–466, March–April 15–15, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:1994:MTJ

[963] Dean Follmann. Modelling transitional and joint marginal distributions in repeated categorical data. Statistics in Medicine, 13(5-7):467–477, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mark:1994:CLB

[964] Steven D. Mark and Mitchell H. Gail. A comparison of likelihood-based and marginal estimating equation methods for analysing repeated ordered categorical responses with missing data: Application to an intervention trial of vitamin prophylaxis for oesophageal dysplasia. Statistics in Medicine, 13(5-7):479–493, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Db

[965] Anonymous. Discussion. Statistics in Medicine, 13(5-7):495–498, March– April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1994:ASM

[966] Gregory Campbell. Advances in statistical methodology for the evaluation of diagnostic and laboratory tests. Statistics in Medicine, 13(5-7): 499–508, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graubard:1994:RAC

[967] Barry I. Graubard and Edward L. Korn. Regression analysis with clustered data. Statistics in Medicine, 13(5-7):509–522, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McShane:1994:SDN

[968] Lisa M. McShane and Meg Palmatier. Spatial distribution of neurons in tissue culture wells: Implications for sampling methods to estimate population size. Statistics in Medicine, 13(5-7):523–540, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waclawiw:1994:EBE

[969] Myron A. Waclawiw and Kung-Yee Liang. Empirical Bayes estimation and inference for the random effects model with binary response. Statistics in Medicine, 13(5-7):541–551, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Dc

[970] Anonymous. Discussion. Statistics in Medicine, 13(5-7):553–556, March– April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1994:SBO

[971] Jonas H. Ellenberg. Selection bias in observational and experimental studies. Statistics in Medicine, 13(5-7):557–567, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kafadar:1994:DAA

[972] Karen Kafadar and Philip C. Prorok. A data-analytic approach for estimating lead time and screening benefit based on survival curves in randomized cancer screening trials. Statistics in Medicine, 13(5-7):569–586, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pearson:1994:MER

[973] Jay D. Pearson, Christopher H. Morrell, Patricia K. Landis, H. Ballentine Carter, and Larry J. Brant. Mixed-effects regression models for studying the natural history of prostate disease. Statistics in Medicine, 13(5-7): 587–601, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Dd

[974] Anonymous. Discussion. Statistics in Medicine, 13(5-7):603–607, March– April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:1994:CLM

[975] Stuart G. Baker. Composite linear models for incomplete multinomial data. Statistics in Medicine, 13(5-7):609–622, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wacholder:1994:PQD

[976] Sholom Wacholder, Raymond J. Carroll, David Pee, and Mitchell H. Gail. The partial questionnaire design for case-control studies. Statistics in Medicine, 13(5-7):623–634, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:1994:TDC

[977] Margaret C. Wu, Sally Hunsberger, and David Zucker. Testing for differences in changes in the presence of censoring: Parametric and nonparametric methods. Statistics in Medicine, 13(5-7):635–646, March– April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:De

[978] Anonymous. Discussion. Statistics in Medicine, 13(5-7):647–649, March– April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benichou:1994:CTA

[979] Jacques Benichou and Sholom Wacholder. A comparison of three approaches to estimate exposure-specific incidence rates from populationbased case-control data. Statistics in Medicine, 13(5-7):651–661, March– April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hunsberger:1994:SCC

[980] Sally Hunsberger, Paul Sorlie, and Nancy L. Geller. Stochastic curtailing and conditional power in matched case-control studies. Statistics in Medicine, 13(5-7):663–670, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weinberg:1994:SBS

[981] Clarice R. Weinberg, Donna D. Baird, and Allen J. Wilcox. Sources of bias in studies of time to pregnancy. Statistics in Medicine, 13(5-7): 671–681, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Df

[982] Anonymous. Discussion. Statistics in Medicine, 13(5-7):683–687, March– April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dinse:1994:CTI

[983] Gregg E. Dinse. A comparison of tumour incidence analyses applicable in single-sacrifice animal experiments. Statistics in Medicine, 13(5-7): 689–708, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1994:MAA

[984] Laurence S. Freedman. Meta-analysis of animal experiments on dietary fat intake and mammary tumours. Statistics in Medicine, 13(5-7):709– 718, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koutsoukos:1994:DTA

[985] Antonis D. Koutsoukos, Lawrence V. Rubinstein, David Faraggi, Richard M. Simon, Sivaram Kalyandrug, John N. Weinstein, Kurt W. Kohn, and Kenneth D. Paull. Discrimination techniques applied to the NCI in vitro anti-tumour drug screen: Predicting biochemical mechanism of action. Statistics in Medicine, 13(5-7):719–730, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Dg

[986] Anonymous. Discussion. Statistics in Medicine, 13(5-7):731–736, March– April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartko:1994:MAS

[987] John J. Bartko. Measures of agreement: a single procedure. Statistics in Medicine, 13(5-7):737–745, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Podgor:1994:NPG

[988] Marvin J. Podgor and Joseph L. Gastwirth. On non-parametric and generalized tests for the two-sample problem with location and scale change alternatives. Statistics in Medicine, 13(5-7):747–758, March–April 15– 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zerbe:1994:TRB

[989] Gary O. Zerbe, Margaret C. Wu, and David M. Zucker. Tudying the relationship between change and initial value in longitudinal studies. Statistics in Medicine, 13(5-7):759–768, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kingman:1994:SPC

[990] Albert Kingman and Gary Zion. Some power considerations when deciding to use transformations. Statistics in Medicine, 13(5-7):769–783, March–April 15–15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mf

[991] Anonymous. Masthead. Statistics in Medicine, 13(8):fmi, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1994:FT

[992] Stephen Senn and William Richardson. The first t-test. Statistics in Medicine, 13(8):785–803, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gentleman:1994:MSM

[993] R. C. Gentleman, J. F. Lawless, J. C. Lindsey, and P. Yan. Multi-state Markov models for analysing incomplete disease history data with illustrations for HIV disease. Statistics in Medicine, 13(8):805–821, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berzuini:1994:BAS

[994] Carlo Berzuini and David Clayton. Bayesian analysis of survival on multiple time scales. Statistics in Medicine, 13(8):823–838, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newton:1994:ASC

[995] Michael A. Newton, Shi-Qi Wu, and Catherine A. Reznikoff. Assessing the significance of chromosome-loss data: Where are suppressor genes for bladder cancer? Statistics in Medicine, 13(8):839–858, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McMahon:1994:SSC

[996] Robert P. McMahon, Michael Proschan, Nancy L. Geller, Peter H. Stone, and George Sopko. Sample size calculation for clinical trials in which entry criteria and outcomes are counts of events. Statistics in Medicine, 13(8):859–870, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Cb

[997] Anonymous. Correction. Statistics in Medicine, 13(8):871, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demetrio:1994:CEP

[998] C. C. B. Demetrio, M. S. Ridout, M. J. Campbell, and D. Machin. Coping with extra Poisson variability in the analysis of factors influencing vaginal ring expulsions. Statistics in Medicine, 13(8):873–876, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:1994:GFA

[999] Helena Chmura Kraemer, Daniel A. Bloch, A. Donner, and M. Eliasziw. A goodness-of-fit approach to inference procedures for the kappa statistic: Confidence interval construction, significance-testing and sample size estimation. Statistics in Medicine, 13(8):876–880, April 30, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hinke:1994:BRG

[1000] Axel Hinke. Book review: Good practice of clinical drug trials. Alain Spriet and Therese Dupin-Spriet, Karger, Basel, 1992. no. of pages: viii + 280. price: £64.80/$119.25. ISBN 3-8055-5367-6. Statistics in Medicine, 13(8):881–882, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VonEye:1994:BRP

[1001] Alexander Von Eye. Book review: Problems and methods in longitudinal research — stability and change, David Magnusson, Lars R. Bergman, Georg Rudinger and Bertil Torestad (eds.), Cambridge University Press, Cambridge, 1991. No of pages: xvi + 351. price: £55/$100. ISBN 0-521-40195-X. Statistics in Medicine, 13(8):882–883, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oquigley:1994:BRS

[1002] John O’Quigley. Book review: Statistical methods for survival data analysis. (2nd edition). Elisa Lee, John Wiley, New York, 1992. no. of pages: xii + 482. price: £47.50. ISBN: 0-471-61592-7. Statistics in Medicine, 13 (8):883–884, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheppard:1994:BRG

[1003] Lianne Sheppard. Book review: Geographical and environmental epidemiology: Methods for small-area studies. P. Elliott, J. Cuzick, D. English and R. Stern (eds.), Oxford University Press, Oxford, 1992. No. of pages: xviii + 382. price: £45. ISBN: 0-19-262280-3. Statistics in Medicine, 13(8):884, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:1994:BRA

[1004] Tony Johnson. Book review: Advanced research methodology: an annotated guide to sources. R. Barker Bausell, The Scarecrow Press, Inc. Metuchen, New Jersey and London, 1991. no. of pages: viii + 903. price: £63.40 ($84.50). ISBN: 0-8108-2355-1. Statistics in Medicine, 13(8):885– 886, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Aa

[1005] Anonymous. Announcements. Statistics in Medicine, 13(8):887–888, April 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mg

[1006] Anonymous. Masthead. Statistics in Medicine, 13(9):fmi, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomas:1994:EBM

[1007] Neal Thomas, Nicholas T. Longford, and John E. Rolph. Empirical Bayes methods for estimating hospital-specific mortality rates. Statistics in Medicine, 13(9):889–903, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grieve:1994:EBA

[1008] Andrew P. Grieve. Extending a Bayesian analysis of the two-period crossover to allow for baseline measurements. Statistics in Medicine, 13 (9):905–929, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grender:1994:FMP

[1009] Julie Myers Grender and William D. Johnson. Fitting multivariate polynomial growth curves in two-period crossover designs. Statistics in Medicine, 13(9):931–943, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:1994:USI

[1010] Margaret C. Wu, K. K. Gordon Lan, and John E. Connett. Use of surrogate information time for monitoring the effect of treatment on the change in a response variable in clinical trials. Statistics in Medicine, 13 (9):945–953, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleming:1994:SAE

[1011] Thomas R. Fleming, Ross L. Prentice, Margaret S. Pepe, and David Glidden. Surrogate and auxiliary endpoints in clinical trials, with potential applications in cancer and AIDS research. Statistics in Medicine, 13 (9):955–968, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:1994:SGB

[1012] Jeremy M. G. Taylor and Yun Chon. Smoothing grouped bivariate data to obtain the incubation period distribution of AIDS. Statistics in Medicine, 13(9):969–981, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yakovlev:1994:PVN

[1013] Andrej Yu. Yakovlev, Alan B. Cantor, and Jonathan J. Shuster. Parametric versus non-parametric methods for estimating cure rates based on censored survival data. Statistics in Medicine, 13(9):983–986, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demets:1994:BRF

[1014] David L. Demets. Book review: Fraud and misconduct in medical research. Stephen Lock and Frank Wells (eds), London, British Medical Journal, 1993. no. of pages: xi + 202. price: £12.95. ISBN: 0-7279-0757-3. Statistics in Medicine, 13(9):987–988, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moolgavkar:1994:BRB

[1015] Suresh H. Moolgavkar. Book review: Biostatistics in cancer risk assessment. David G. Hoel and Takashi Yanagawa (eds), Scientist, Inc., Tokyo, 1991. No. of pages: 176. price: $30. Statistics in Medicine, 13(9):988, May 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mh

[1016] Anonymous. Masthead. Statistics in Medicine, 13(10):fmi, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1994:IMM

[1017] Sander Greenland and George Maldonado. The interpretation of multiplicative-model parameters as standardized parameters. Statistics in Medicine, 13(10):989–999, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1994:ERA

[1018] David W. Smith. Evaluating risk adjustment by partitioning variation in hospital mortality rates. Statistics in Medicine, 13(10):1001–1013, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yarnold:1994:HRV

[1019] Paul R. Yarnold, Robert C. Soltysik, and Gary J. Martin. Heart rate variability and susceptibility for sudden cardiac death: an example of multivariable optimal discriminant analysis. Statistics in Medicine, 13 (10):1015–1021, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glynn:1994:CAR

[1020] Robert J. Glynn and Bernard Rosner. Comparison of alternative regression models for paired binary data. Statistics in Medicine, 13(10): 1023–1036, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1994:STS

[1021] S. D. Walter. A simple test for spatial pattern in regional health data. Statistics in Medicine, 13(10):1037–1044, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hess:1994:ATC

[1022] Kenneth R. Hess. Assessing time-by-covariate interactions in proportional hazards regression models using cubic spline functions. Statistics in Medicine, 13(10):1045–1062, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carriere:1994:CDC

[1023] K. C. Carriere. Crossover designs for clinical trials. Statistics in Medicine, 13(10):1063–1069, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waller:1994:CIW

[1024] Jennifer L. Waller, Cheryl L. Addy, Kirby L. Jackson, and Carol Z. Garrison. Confidence intervals for weighted proportions. Statistics in Medicine, 13(10):1071–1082, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ulm:1994:BRS

[1025] Kurt Ulm. Book review: Statistical analysis of epidemiologic data. Steve Selvin, Oxford University Press, Oxford, 1991. No Of Pages: xiv + 375. Price: £27.50. ISBN: 0-19-506766-5. Statistics in Medicine, 13(10):1083, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dewey:1994:BRS

[1026] Michael Dewey. Book review: Statistical methods for psychology (3rd edn). David C. Howell, Boston, PWS-Kent, 1992. No. of pages: xvii + 693. price: £17.95 (International Student Edition). ISBN: 0-534-97217-9. Statistics in Medicine, 13(10):1084, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:1994:BRR

[1027] Peter Bauer. Book review: Resampling-based multiple testing. Peter H. Westfall, S. Stanley Young, Wiley, New York, 1993. No. of pages: xvii + 340. Price: £41.95. ISBN: 0-471-55761-7. Statistics in Medicine, 13(10): 1084–1086, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Ab

[1028] Anonymous. Announcement. Statistics in Medicine, 13(10):1087, May 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mi

[1029] Anonymous. Masthead. Statistics in Medicine, 13(11):fmi, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lavori:1994:CET

[1030] Philip W. Lavori, Ree Dawson, and Timothy B. Mueller. Causal estimation of time-varying treatment effects in observational studies: Application to depressive disorder. Statistics in Medicine, 13(11):1089–1100, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vittinghoff:1994:EPC

[1031] Eric Vittinghoff, Hina Mehta Malani, and Nicholas P. Jewell. Estimating patterns of CD4 lymphocyte decline using data from a prevalent cohort of HIV infected individuals. Statistics in Medicine, 13(11):1101–1118, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Devine:1994:CEB

[1032] Owen J. Devine and Thomas A. Louis. A constrained empirical Bayes estimator for incidence rates in areas with small populations. Statistics in Medicine, 13(11):1119–1133, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuha:1994:CEM

[1033] Jouni Kuha. Corrections for exposure measurement error in logistic regression models with an application to nutritional data. Statistics in Medicine, 13(11):1135–1148, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipsitz:1994:ARC

[1034] Stuart R. Lipsitz, Kyungmann Kim, and Lueping Zhao. Analysis of repeated categorical data using generalized estimating equations. Statistics in Medicine, 13(11):1149–1163, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Giltinan:1994:ARP

[1035] David M. Giltinan and Marie Davidian. Assays for recombinant proteins: a problem in non-linear calibration. Statistics in Medicine, 13(11):1165– 1179, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1994:RMCb

[1036] Stephen Senn, Julie Myers Grender, William D. Johnson, and Robert C. Elston. Regression to the mean and crossover trials revisited. Statistics in Medicine, 13(11):1181–1186, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Macfarlane:1994:BRM

[1037] Sarah B. J. Macfarlane. Book review: Methods for field trials of interventlons against tropical diseases. A Toolbox, P. G. Smith and Richard H. Morrow (eds.), Oxford University Press on behalf of UNDP/World Bank/WHO Special Programme for Research and Training in Tropical Diseases, 1991. No. of pages: xxii + 326. price: £15 (paperback). ISBN: 0-19-262017-7. Statistics in Medicine, 13(11):1187–1188, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paul:1994:BRA

[1038] Sudhir R. Paul. Book review: Analysis of quantal response data B. J. T. Morgan, Chapman and Hall, London, 1992. No. of pages: xx + 511. price: £35. ISBN 0-412-31750-8. Statistics in Medicine, 13(11):1188, June 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mj

[1039] Anonymous. Masthead. Statistics in Medicine, 13(12):fmi, June 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liestbl:1994:SAN

[1040] Knut Liestbl, Per Kragh Andersen, and Ulrich Andersen. Survival analysis and neural nets. Statistics in Medicine, 13(12):1189–1200, June 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Somes:1994:FSA

[1041] Grant W. Somes, G. A. Harshfield, K. L. Arheart, and S. T. Miller. A Fourier series approach for comparing groups of subjects on ambulatory blood pressure patterns. Statistics in Medicine, 13(12):1201–1210, June 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rutter:1994:ALD

[1042] Carolyn M. Rutter and Robert M. Elashoff. Analysis of longitudinal data: Random coefficient regression modelling. Statistics in Medicine, 13(12):1211–1231, June 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipsitz:1994:SSR

[1043] Stuart R. Lipsitz and Garrett M. Fitzmaurice. Sample size for repeated measures studies with binary responses. Statistics in Medicine, 13(12): 1233–1239, June 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1994:ESS

[1044] Eric W. Lee and Neil Dubin. Estimation and sample size considerations for clustered binary responses. Statistics in Medicine, 13(12):1241–1252, June 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1994:CMT

[1045] Allan Donner, Michael Eliasziw, and Neil Klar. A comparison of methods for testing homogeneity of proportions in teratologic studies. Statistics in Medicine, 13(12):1253–1264, June 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carroll:1994:MEI

[1046] R. J. Carroll and L. A. Stefanski. Measurement error, instrumental variables and corrections for attenuation with applications to metaanalyses. Statistics in Medicine, 13(12):1265–1282, June 30, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:1994:CIB

[1047] Robert G. Newcombe and Stein Emil Vollset. Confidence intervals for a binomial proportion. Statistics in Medicine, 13(12):1283–1285, June 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beresford:1994:BRD

[1048] Shirley A. A. Beresford. Book review: Design concepts in nutritional epidemiology. Barrie M. Margetts and Michael Nelson, Oxford University Press, Oxford, 1991. No. of pages: xi + 420. Price: £30.00. ISBN: 0-19-261873-3. Statistics in Medicine, 13(12):1287–1288, June 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mk

[1049] Anonymous. Masthead. Statistics in Medicine, 13(13-14):fmi, July 15– 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:PA

[1050] Anonymous. Participants and authors. Statistics in Medicine, 13(13-14):1289–1291, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Souhami:1994:CIE

[1051] Robert L. Souhami. The clinical importance of early stopping of randomized trials in cancer treatments. Statistics in Medicine, 13(13-14):1293– 1295, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parmar:1994:CTB

[1052] Mahesh K. B. Parmar, David J. Spiegelhalter, Laurence S. Freedman, and Chart Steering Committee. The chart trials: Bayesian design and monitoring in practice. Statistics in Medicine, 13(13-14):1297–1312, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

George:1994:SCT

[1053] Stephen L. George, Chengchang Li, Donald A. Berry, and Mark R. Green. Stopping a clinical trial early: Frequentist and Bayesian approaches applied to a CALGB trial in non-small-cell lung cancer. Statistics in Medicine, 13(13-14):1313–1327, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sylvester:1994:RFP

[1054] Richard Sylvester, Harry Bartelink, Robert Rubens, and EORTC Breast Cancer Cooperative Group. A reversal of fortune: Practical problems in the monitoring and interpretation of an EORTC breast cancer trial. Statistics in Medicine, 13(13-14):1329–1335, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pater:1994:BRT

[1055] Joseph L. Pater. Book review: Timing the collaborative analysis of three trials comparing 5-FU plus folinic acid (FUFA) to surgery alone in the management of resected colorectal cancer: a National Cancer Institute of Canada Clinical Trials Group (NCIC-CTG) perspective. Statistics in Medicine, 13(13-14):1337–1340, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demets:1994:IAA

[1056] David L. Demets and K. K. Gordon Lan. Interim analysis: The alpha spending function approach. Statistics in Medicine, 13(13-14):1341–1352, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geller:1994:DIA

[1057] Nancy L. Geller. Discussion of ‘Interim analysis: The alpha spending approach’. Statistics in Medicine, 13(13-14):1353–1356, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1994:SMB

[1058] John Whitehead. Sequential methods based on the boundaries approach for the clinical comparison of survival times. Statistics in Medicine, 13 (13-14):1357–1368, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

George:1994:DSM

[1059] Stephen L. George. Discussion of ‘Sequential methods based on the boundaries approach for the clinical comparison of survival times’. Statistics in Medicine, 13(13-14):1369–1370, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1994:WWH

[1060] Laurence S. Freedman, David J. Spiegelhalter, and Mahesh K. B. Parmar. The what, why and how of Bayesian clinical trials monitoring. Statistics in Medicine, 13(13-14):1371–1383, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machin:1994:DWW

[1061] David Machin. Discussion of ‘The what, why and how of Bayesian clinical trials monitoring’. Statistics in Medicine, 13(13-14):1385–1389, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crowley:1994:DMC

[1062] J. Crowley, S. Green, P. Y. Liu, and M. Wolf. Data monitoring committees and early stopping guidelines: The Southwest Oncology Group experience. Statistics in Medicine, 13(13-14):1391–1399, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simon:1994:SPA

[1063] Richard Simon. Some practical aspects of the interim monitoring of clinical trials. Statistics in Medicine, 13(13-14):1401–1409, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harrington:1994:CAI

[1064] David Harrington, John Crowley, Stephen L. George, Thomas Pajak, Carol Redmond, and Sam Wieand. The case against independent monitoring committees. Statistics in Medicine, 13(13-14):1411–1414, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pignon:1994:TTR

[1065] Jean-Pierre Pignon, Michèle Tarayre, Ariane Auquier, Rodrigo Arriagada, Thierry le Chevalier, Pierre Ruffié, Alain Rivière, Isabelle Monnet, Pierre Chomy, and Claude Tuchais. Triangular test and randomized trials: Practical problems in a small cell lung cancer trial. Statistics in Medicine, 13(13-14):1415–1421, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleming:1994:SMA

[1066] Thomas R. Fleming. Surrogate markers in AIDS and cancer trials. Statistics in Medicine, 13(13-14):1423–1435, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1994:DSM

[1067] Susan Ellenberg. Discussion of ‘Surrogate markers in AIDS and cancer trials’. Statistics in Medicine, 13(13-14):1437–1440, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:1994:MMA

[1068] Michael A. Proschan, Dean A. Follmann, and Nancy L. Geller. Monitoring multi-armed trials. Statistics in Medicine, 13(13-14):1441–1452, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wieand:1994:SWE

[1069] Sam Wieand, Georgene Schroeder, and Judith Rich O’Fallon. Stopping when the experimental regimen does not appear to help. Statistics in Medicine, 13(13-14):1453–1458, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baum:1994:ESR

[1070] M. Baum, J. Houghton, and K. Abrams. Early stopping rules — clinical perspectives and ethical considerations. Statistics in Medicine, 13(13-14):1459–1469, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burke:1994:DES

[1071] G. Burke. Discussion of ‘Early stopping rules — clinical perspectives and ethical considerations’. Statistics in Medicine, 13(13-14):1471–1472, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robinson:1994:LPP

[1072] Jean Robinson. A lay person’s perspective on starting and stopping clinical trials. Statistics in Medicine, 13(13-14):1473–1477, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1994:ESP

[1073] Sylvan B. Green and Laurence S. Freedman. Early stopping of prevention trials when multiple outcomes are of interest: a discussion. Statistics in Medicine, 13(13-14):1479–1484, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liberati:1994:CRB

[1074] Alessandro Liberati. Conclusions. 1: The relationship between clinical trials and clinical practice: The risks of underestimating its complexity. Statistics in Medicine, 13(13-14):1485–1491, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

ONeill:1994:C

[1075] Robert T. O’Neill. Conclusions. 2. Statistics in Medicine, 13(13-14): 1493–1499, July 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Ml

[1076] Anonymous. Masthead. Statistics in Medicine, 13(15):fmi, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:1994:APM

[1077] Guillermo Marshall, Frederick L. Grover, William G. Henderson, and Karl E. Hammermeister. Assessment of predictive models for binary outcomes: an empirical approach using operative death from cardiac surgery. Statistics in Medicine, 13(15):1501–1511, August 15, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hakulinen:1994:PIP

[1078] Timo Hakulinen and Tadeusz Dyba. Precision of incidence predictions based on Poisson distributed observations. Statistics in Medicine, 13 (15):1513–1523, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raghunathan:1994:MCM

[1079] T. E. Raghunathan. Monte Carlo methods for exloring sensitivity to distributional assumptions in a Bayesian analysis of a series of 2 × 2 tables. Statistics in Medicine, 13(15):1525–1538, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirji:1994:EPC

[1080] Mkarim F. Hirji, Man-Lai Tang, Stein E. Vollset, and Robert M. Elashoff. Efficient power computation for exact and mid-P tests for the common odds ratio in several 2 × 2 tables. Statistics in Medicine, 13(15):1539– 1549, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schafer:1994:ECB

[1081] Helmut Schafer. Efficient confidence bounds for ROC curves. Statistics in Medicine, 13(15):1551–1561, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schork:1994:SGT

[1082] Nicholas J. Schork. Sampling guidelines for testing secondary attack rates associated with short-latency infectious diseases. Statistics in Medicine, 13(15):1563–1573, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yakovlev:1994:CCT

[1083] Andrej Yu. Yakovlev, Roman E. Goot, and Tamara T. Osipova. The choice of cancer treatment based on covariate information. Statistics in Medicine, 13(15):1575–1581, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rom:1994:CTP

[1084] Dror M. Rom, Robert J. Costello, and Laura T. Connell. On closed test procedures for dose-response analysis. Statistics in Medicine, 13 (15):1583–1596, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:1994:BRA

[1085] Jørgen Hilden. Book review: Analysing how we reach clinical decisions Huw Llewelyn and Anthony Hopkins, Royal College of Physicians of London, London, 1993. No. of pages: xii + 166. Price £18.95. ISBN 1-873240-68-6. Statistics in Medicine, 13(15):1597–1598, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1994:BRU

[1086] Michael D. Hughes. Book review: The use of restricted significance tests in clinical trials. David S. Salsburg, Springer-Verlag, New York, 1992. No. of pages: x + 173. Price: $39.95. ISBN 3-540-97798-8. Statistics in Medicine, 13(15):1598, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1994:BRM

[1087] Douglas G. Altman. Book review: Medical uses of statistics. 2nd edn., John C. Bailar III and Frederick Mosteller (eds.), NEJM Books, Boston, 1992. No. of pages: xxvii + 449. Price: $39.95. ISBN: 0-910133-26-0. Statistics in Medicine, 13(15):1599, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Cc

[1088] Anonymous. Correction. Statistics in Medicine, 13(15):1600, August 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mm

[1089] Anonymous. Masthead. Statistics in Medicine, 13(16):fmi, August 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:1994:ARC

[1090] H. K. Brown and R. A. Kempton. The application of REML in clinical trials. Statistics in Medicine, 13(16):1601–1617, August 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grambsch:1994:SAP

[1091] Patricia Grambsch, Thomas A. Louis, Roberd M. Bostick, Greg A. Grandits, Lisa Fosdick, Mohamed Darif, and John D. Potter. Statistical analysis of proliferative index data in clinical trials. Statistics in Medicine, 13(16):1619–1634, August 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:1994:MAB

[1092] Ian R. White, Nishi Chaturvedi, and Paul M. McKeigue. Median analysis of blood pressure for a sample including treated hypertensives. Statistics in Medicine, 13(16):1635–1641, August 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldstein:1994:MTS

[1093] Harvey Goldstein, Michael J. R. Healy, and Jon Rasbash. Multilevel time series models with applications to repeated measures data. Statistics in Medicine, 13(16):1643–1655, August 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Silverberg:1994:SSC

[1094] Arthur R. Silverberg. A simulation study comparing two approximations for a quasi t-quantile, used in repeated measures anova. Statistics in Medicine, 13(16):1657–1663, August 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1994:AMO

[1095] Sander Greenland. Alternative models for ordinal logistic regression. Statistics in Medicine, 13(16):1665–1677, August 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Babiker:1994:SFM

[1096] Abdel Babiker and Jack Cuzick. A simple frailty model for family studies with covariates. Statistics in Medicine, 13(16):1679–1692, August 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwardes:1994:CIB

[1097] Michael Edwardes and Stein Emil Vollset. Confidence intervals for a binomial proportion. Statistics in Medicine, 13(16):1693–1698, August 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mn

[1098] Anonymous. Masthead. Statistics in Medicine, 13(17):fmi, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burton:1994:HDD

[1099] Paul R. Burton. Helping doctors to draw appropriate inferences from the analysis of medical studies. Statistics in Medicine, 13(17):1699–1713, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1994:TBB

[1100] Stephen Senn. Testing for baseline balance in clinical trials. Statistics in Medicine, 13(17):1715–1726, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ensign:1994:OTS

[1101] Lisa Garnsey Ensign, Edmund A. Gehan, Douglas S. Kamen, and Peter F. Thall. An optimal three-stage design for phase II clinical trials. Statistics in Medicine, 13(17):1727–1736, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1994:EVB

[1102] Xiao-Hua Zhou. Effect of verification bias on positive and negative predictive values. Statistics in Medicine, 13(17):1737–1745, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finkelstein:1994:ASP

[1103] Dianne M. Finkelstein and David A. Schoenfeld. Analysing survival in the presence of an auxiliary variable. Statistics in Medicine, 13(17):1747– 1754, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2065].

Frey:1994:PIR

[1104] Carolin M. Frey, Eric J. Feuer, and Marilyn J. Timmel. Projection of incidence rates to a larger population using ecologic variables. Statistics in Medicine, 13(17):1755–1770, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Samuelsen:1994:ICL

[1105] Sven Ove Samuelsen and Johny Kongerud. Interval censoring in longitudinal data of respiratory symptoms in aluminium potroom workers: a comparison of methods. Statistics in Medicine, 13(17):1771–1780, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stukel:1994:SRR

[1106] Thérèse A. Stukel, Robert J. Glynn, Elliott S. Fisher, Sandra M. Sharp, Grace Lu-Yao, and John E. Wennberg. Standardized rates of recurrent outcomes. Statistics in Medicine, 13(17):1781–1791, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bello:1994:ERT

[1107] A. L. Bello, Thomas W. O’Gorman, and Robert F. Woolson. On the efficacy of the rank transformation in stepwise logistic and discriminant analysis. Statistics in Medicine, 13(17):1793–1795, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sasieni:1994:BRS

[1108] Peter Sasieni. Book review: Statistical models based on counting processes. Per Kragh Andersen, Ørnulf Borgan, Richard D. Gill and Niels Keiding, Springer-Verlag, New York, 1993. No. of pages: xi + 767. Price: DM128. ISBN: 0-387-97872-0. Statistics in Medicine, 13(17):1797–1798, September 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mo

[1109] Anonymous. Masthead. Statistics in Medicine, 13(18):fmi, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:1994:CTP

[1110] Edward L. Korn, Douglas Midthune, T. Timothy Chen, Lawrence V. Rubinstein, Michaele C. Christian, and Richard M. Simon. A comparison of two phase I trial designs. Statistics in Medicine, 13(18):1799–1806, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chang:1994:RBI

[1111] M. N. Chang, H. A. Guess, and J. F. Heyse. Reduction in burden of illness: a new efficacy measure for prevention trials. Statistics in Medicine, 13(18):1807–1814, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grossman:1994:UMM

[1112] Jason Grossman, Mahesh K. B. Parmar, David J. Spiegelhalter, and Laurence S. Freedman. A unified method for monitoring and analysing controlled trials. Statistics in Medicine, 13(18):1815–1826, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Garner:1994:DES

[1113] J. Barry Garner. Deviance estimates of sample size for equivalence tests in vaccine trials. Statistics in Medicine, 13(18):1827–1839, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

TenHave:1994:BHA

[1114] Thomas R. Ten Have and Vernon M. Chinchilli. Bayesian hierarchical analysis of within-units variances in repeated measures experiments. Statistics in Medicine, 13(18):1841–1852, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dawson:1994:SSS

[1115] Jeffrey D. Dawson. Stratification of summary statistic tests according to missing data patterns. Statistics in Medicine, 13(18):1853–1863, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marschner:1994:SBP

[1116] Ian C. Marschner, Stuart V. Newstead, and Niels G. Becker. Simultaneous back-projection of AIDS incidence data for two or more groups. Statistics in Medicine, 13(18):1865–1879, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ades:1994:EDI

[1117] A. E. Ades and G. F. Medley. Estimates of disease incidence in women based on antenatal or neonatal seroprevalence data: HIV in New York City. Statistics in Medicine, 13(18):1881–1894, September 30, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:1994:ERC

[1118] Nico J. D. Nagelkerke, Anthony Mwatha, and D. Y. Lin. Evaluating the role of CD4-lymphocyte counts as surrogate endpoints in human immunodeficiency virus clinical trials. Statistics in Medicine, 13(18): 1895–1896, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Cd

[1119] Anonymous. Correction. Statistics in Medicine, 13(18):1897, September 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mp

[1120] Anonymous. Masthead. Statistics in Medicine, 13(19-20):fmi, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomsen:1994:F

[1121] D. L. Thomsen, Jr. Foreword. Statistics in Medicine, 13(19-20):1899– 1900, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dietz:1994:P

[1122] Klaus Dietz, Peter Bacchetti, and Marcello Pagano. Preface. Statistics in Medicine, 13(19-20):1901–1903, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:1994:STC

[1123] Xin M. Tu, Eugene Litvak, and Marcello Pagano. Screening tests: Can we get more by doing less? Statistics in Medicine, 13(19-20):1905–1919, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bacchetti:1994:SOS

[1124] Peter Bacchetti. Seasonal and other short-term influences on United States AIDS incidence. Statistics in Medicine, 13(19-20):1921–1931, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailey:1994:PVP

[1125] Norman T. J. Bailey. Prediction and validation in the public health modelling of HIV/AIDS. Statistics in Medicine, 13(19-20):1933–1943, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:1994:DHI

[1126] Niels G. Becker and Xu Chao. Dependent HIV incidences in backprojection of AIDS incidence data. Statistics in Medicine, 13(19-20): 1945–1958, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marschner:1994:UTF

[1127] Ian C. Marschner. Using time of first positive HIV test and other auxiliary data in back-projection of AIDS incidence. Statistics in Medicine, 13(19-20):1959–1974, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberg:1994:BMA

[1128] Philip S. Rosenberg. Backcalculation models of age-specific HIV incidence rates. Statistics in Medicine, 13(19-20):1975–1990, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dietz:1994:BPG

[1129] Klaus Dietz, Johannes Seydel, and Bernhard Schwartländer. Backprojection of German AIDS data using information on dates of tests. Statistics in Medicine, 13(19-20):1991–2008, October 15–30, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raab:1994:IHT

[1130] Gillian M. Raab, Katherine L. Fielding, and Gwen Allardice. Incorporating HIV test data into forecasts of the AIDS epidemic in Scotland. Statistics in Medicine, 13(19-20):2009–2020, October 15–30, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gauvreau:1994:ECI

[1131] Kimberlee Gauvreau, Victor Degruttola, Marcello Pagano, and Rino Bellocco. The effect of covariates on the induction time of AIDS using improved imputation of exact seroconversion times. Statistics in Medicine, 13(19-20):2021–2030, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mawhinney:1994:DLP

[1132] Samantha Mawhinney and Marcello Pagano. Distribution of the latency period for perinatally acquired AIDS. Statistics in Medicine, 13(19-20): 2031–2042, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Motika:1994:NPA

[1133] Milena Motika, Niels G. Becker, Brenton Wylie, Jennifer Learmont, and John Kaldor. Non-parametric analysis of transfusion-related HIV and AIDS incidence data. Statistics in Medicine, 13(19-20):2043–2058, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:1994:MVT

[1134] Jeremy M. G. Taylor and Dong Kee Kim. Marker values at the time of an AIDS diagnosis. Statistics in Medicine, 13(19-20):2059–2066, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schenzle:1994:MAP

[1135] Dieter Schenzle. A model for AIDS pathogenesis. Statistics in Medicine, 13(19-20):2067–2079, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jewell:1994:NPE

[1136] Nicholas P. Jewell. Non-parametric estimation and doubly-censored data: General ideas and applications to AIDS. Statistics in Medicine, 13(19-20):2081–2095, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Satten:1994:MFM

[1137] Glen A. Satten, Timothy D. Mastro, and Ira M. Longini, Jr. Modelling the female-to-male per-act HIV transmission probability in an emerging epidemic in Asia. Statistics in Medicine, 13(19-20):2097–2106, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Neaton:1994:CCC

[1138] James D. Neaton, Deborah N. Wentworth, Frank Rhame, Carlton Hogan, Donald I. Abrams, and Lawrence Deyton. Considerations in choice of a clinical endpoint for AIDS clinical trials. Statistics in Medicine, 13(19-20):2107–2125, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoover:1994:EII

[1139] Donald R. Hoover, Alvaro Muñtoz, Yanhua He, Jeremy M. G. Taylor, Lawrence Kingsley, Joan S. Chmiel, and Alfred Saah. The effectiveness of interventions on incubation of AIDS as measured by secular increases within a population. Statistics in Medicine, 13(19-20):2127–2139, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:1994:ADC

[1140] Hyungjin Myra Kim and Stephen W. Lagakos. Assessing drug compliance using longitudinal marker data, with application to AIDS. Statistics in Medicine, 13(19-20):2141–2153, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rida:1994:SSI

[1141] Wasima N. Rida and Dale N. Lawrence. Some statistical issues in HIV vaccine trials. Statistics in Medicine, 13(19-20):2155–2177, October 15– 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaplan:1994:MEN

[1142] Edward H. Kaplan. A method for evaluating needle exchange programmes. Statistics in Medicine, 13(19-20):2179–2187, October 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mq

[1143] Anonymous. Masthead. Statistics in Medicine, 13(21):fmi, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cohen:1994:MCI

[1144] Geoffrey R. Cohen and Shu-Ying Yang. Mid-p confidence intervals for the Poisson expectation. Statistics in Medicine, 13(21):2189–2203, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tian:1994:PFO

[1145] Jason J. Tian, Rakesh Shukla, and C. Ralph Buncher. On prediction of future observation in growth curve model. Statistics in Medicine, 13 (21):2205–2217, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palta:1994:TOV

[1146] Mari Palta, Tzy-Jyun Yao, and Raja Velu. Testing for omitted variables and non-linearity in regression models for longitudinal data. Statistics in Medicine, 13(21):2219–2231, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:1994:CRA

[1147] D. Y. Lin. Cox regression analysis of multivariate failure time data: The marginal approach. Statistics in Medicine, 13(21):2233–2247, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fayers:1994:DMR

[1148] P. M. Fayers, P. A. Cook, D. Machin, N. Donaldson, J. Whitehead, A. Ritchie, R. T. D. Oliver, and P. Yuen. On the development of the Medical Research Council trial of α-interferon in metastatic renal carcinoma. Statistics in Medicine, 13(21):2249–2260, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cheung:1994:SPM

[1149] Siu Hung Cheung and Burt Holland. A step-down procedure for multiple tests of treatment versus control in each of several groups. Statistics in Medicine, 13(21):2261–2267, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:1994:PAD

[1150] Stuart G. Baker and Karen S. Lindeman. The paired availability design: a proposal for evaluating epidural analgesia during labor. Statistics in Medicine, 13(21):2269–2278, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Falk:1994:HEE

[1151] Raymond W. Falk. How to establish equivalence between treatments: a one-sided clinical trial in pediatric oncology. Statistics in Medicine, 13 (21):2279–2280, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1994:MAD

[1152] Stephen Senn, Lloyd E. Chambless, and John R. Roeback. Methods for assessing difference between groups in change when initial measurement is subject to intra-individual variation. Statistics in Medicine, 13(21): 2280–2285, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [781, 1424].

Evans:1994:BRP

[1153] Stephen J. W. Evans. Book review: I. Primer of biostatistics. Third edn., Stanton A. Glantz, McGraw-Hill Inc., New York, 1992. no. of pages: xvi + 440. price: $27. ISBN: 0-07-023511-2. Statistics in Medicine, 13(21): 2287–2288, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1994:BRR

[1154] David R. Jones. Book review: A reader on applying statistics in public health and prevention. J. L. A. van Rijckevorsel and C. C. J. H. Bijleveld (eds), TNO Institute of Preventive Health Care, Leiden, 1992. number of pages: vii + 217. price: Dfl 49.50. ISBN: 90-6743-212-1. Statistics in Medicine, 13(21):2288, November 15, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Mr

[1155] Anonymous. Masthead. Statistics in Medicine, 13(22):fmi, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machado:1994:DGA

[1156] Stella G. Machado and Gregory A. Robinson. A direct, general approach based on isobolograms for assessing the joint action of drugs in preclinical experiments. Statistics in Medicine, 13(22):2289–2309, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Etzioni:1994:MPT

[1157] Ruth Etzioni and Margaret Sullivan Pepe. Monitoring of a pilot toxicity study with two adverse outcomes. Statistics in Medicine, 13(22):2311– 2321, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1994:MBA

[1158] Jeffrey M. Albert and David L. Demets. On a model-based approach to estimating efficacy in clinical trials. Statistics in Medicine, 13(22):2323– 2335, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Phatarfod:1994:USA

[1159] R. M. Phatarfod and Aidan Sudbury. The use of a square array scheme in blood testing. Statistics in Medicine, 13(22):2337–2343, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1994:AFE

[1160] Tom Smith, Joanna Armstrong Schellenberg, and Richard Hayes. Attributable fraction estimates and case definitions for malaria in endemic. Statistics in Medicine, 13(22):2345–2358, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wade:1994:ARR

[1161] A. M. Wade and A. E. Ades. Age-related reference ranges: Significance tests for models and confidence intervals for centiles. Statistics in Medicine, 13(22):2359–2367, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marschner:1994:DSC

[1162] Ian C. Marschner. Determining the size of a cross-sectional sample to estimate the age-specific incidence of an irreversible disease. Statistics in Medicine, 13(22):2369–2381, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:1994:ALT

[1163] Odd O. Aalen, Bo Eriksson, and Ragnar Adell. On the analysis of life tables for dependent observations. Statistics in Medicine, 13(22):2383– 2384, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

West:1994:BRT

[1164] Robert West and Martin Bland. Book review: Talking about statistics: a Psychologist’s guide to design and analysis. Brian Everitt and Dale Hay, Edward Arnold, London, 1992. no. of pages: viii + 130. price: £12.99. ISBN: 0-340-52921-0. Statistics in Medicine, 13(22):2385, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barrett:1994:BRF

[1165] Jenny Barrett. Book review: Fundamentals of genetic epidemiology. Muin J. Khoury, Terri H. Beaty and Bernice H. Cohen, Oxford University Press, New York, 1993. no. of pages: vi + 383. price: £35. ISBN: 0-19-505288-9. Statistics in Medicine, 13(22):2386–2387, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Raab:1994:BRA

[1166] Gillian M. Raab and Daniela De Angelis. Book reviews: AIDS epidemiology: Methodological issues. Nicholas P. Jewell, Klaus Dietz and Vernon T. Farewell (eds), Birkhäuser, Boston, 1992. no. of pages: xvi + 402. price: DM148/sfr134. ISBN: 0-8176-3632-3. HIV epidemiology: Models and methods. Alfredo Nicolosi (ed), Raven Press, New York, no. of pages: xvi + 367. price: $107.50. ISBN: 0-7817-0118-X. Statistics in Medicine, 13(22):2387–2388, November 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:Ms

[1167] Anonymous. Masthead. Statistics in Medicine, 13(23-24):fmi, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1994:P

[1168] Ted Colton, Tony Johnson, and David Machin. Preface. Statistics in Medicine, 13(23-24):2389, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:ISCa

[1169] Anonymous. International Society for Clinical Biostatistics Fourteenth International Meeting held at the University of Cambridge Cambridge 21–24 September 1993. Statistics in Medicine, 13(23-24):2391, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1994:ISCb

[1170] Anonymous. International Society for Clinical Biostatistics Fourteenth International Meeting held at the University of Cambridge Cambridge 21–24 September 1993. Statistics in Medicine, 13(23-24):2391, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marsh:1994:PMC

[1171] Marta J. Marsh and Barbara S. Hawkins. Publications from multicentre clinical trials: Statistical techniques and accessibility to the reader. Statistics in Medicine, 13(23-24):2393–2406, December 15–30, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grassi:1994:CAA

[1172] M. Grassi and S. Visentin. Correspondence analysis applied to grouped cohort data. Statistics in Medicine, 13(23-24):2407–2425, December 15– 30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verweij:1994:PLC

[1173] Pierre J. M. Verweij and Hans C. Van Houwelingen. Penalized likelihood in Cox regression. Statistics in Medicine, 13(23-24):2427–2436, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Svensson:1994:SSR

[1174] Elisabeth Svensson and Sture Holm. Separation of systematic and random differences in ordinal rating scales. Statistics in Medicine, 13(23-24): 2437–2453, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Birkett:1994:IPS

[1175] Martin A. Birkett and Simon J. Day. Internal pilot studies for estimating sample size. Statistics in Medicine, 13(23-24):2455–2463, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muller:1994:CDI

[1176] Reinhold Müller and Petra Büttner. A critical discussion of intraclass correlation coefficients. Statistics in Medicine, 13(23-24):2465–2476, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1822].

Cole:1994:GCB

[1177] T. J. Cole. Growth charts for both cross-sectional and longitudinal data. Statistics in Medicine, 13(23-24):2477–2492, December 15–30, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brunier:1994:SSP

[1178] Hazel C. Brunier and John Whitehead. Sample sizes for phase ii clinical trials derived from Bayesian decision theory. Statistics in Medicine, 13 (23-24):2493–2502, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1994:MAC

[1179] Anne Whitehead and Nicola M. B. Jones. A meta-analysis of clinical trials involving different classifications of response into ordered categories. Statistics in Medicine, 13(23-24):2503–2515, December 15–30, 1994. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hansen:1994:MML

[1180] B. E. Hansen, J. Thorogood, J. Hermans, R. J. Ploeg, J. H. Van Bockel, and J. C. Van Houwelingen. Multistate modelling of liver transplantation data. Statistics in Medicine, 13(23-24):2517–2529, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vuorinen:1994:FCI

[1181] Jouni Vuorinen and Juhani Tuominen. Fieller’s confidence intervals for the ratio of two means in the assessment of average bioequivalence from crossover data. Statistics in Medicine, 13(23-24):2531–2545, December 15–30, 1994. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Ma

[1182] Anonymous. Masthead. Statistics in Medicine, 14(1):fmi, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1995:Ea

[1183] Ted Colton, Tony Johnson, and David Machin. Editorial. Statistics in Medicine, 14(1):1–2, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waller:1995:DAC

[1184] Lance A. Waller, Bruce W. Turnbull, Göran Gustafsson, Ulf Hjalmars, and Bertil Andersson. Detection and assessment of clusters of disease: an application to nuclear power plant facilities and childhood leukaemia in Sweden. Statistics in Medicine, 14(1):3–16, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oden:1995:AMP

[1185] Neal Oden. Adjusting Moran’s I for population density. Statistics in Medicine, 14(1):17–26, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1625].

Alioum:1995:DAC

[1186] Ahmadou Alioum and Daniel Commenges. A direct approach for correcting AIDS incidence: Variance formula and comparison with other methods. Statistics in Medicine, 14(1):27–38, January 15, 1995. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghitany:1995:EPI

[1187] M. E. Ghitany, R. Maller, and S. Zhou. Estimating the proportion of immunes in censored samples: a simulation study. Statistics in Medicine, 14(1):39–49, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gargiullo:1995:CIH

[1188] Paul M. Gargiullo, Richard B. Rothenberg, and Hoyt G. Wilson. Confidence intervals, hypothesis tests, and sample sizes for the prevented fraction in cross-sectional studies. Statistics in Medicine, 14(1):51–72, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faraggi:1995:NNM

[1189] David Faraggi and Richard Simon. A neural network model for survival data. Statistics in Medicine, 14(1):73–82, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shoukri:1995:MLE

[1190] M. M. Shoukri, S. W. Martin, and I. U. H. Mian. Maximum likelihood estimation of the kappa coefficient from models of matched binary responses. Statistics in Medicine, 14(1):83–99, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turney:1995:EAS

[1191] Elizabeth A. Turney. Evaluation of alternative statistical methods for linear model analysis to compare two treatments for 24 hour blood pressure response. Statistics in Medicine, 14(1):101–102, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Healy:1995:BRS

[1192] M. J. R. Healy. Book review: Sense and Nonsense of Statistical Inference. Chamant Wang and Marcel Dekker, New York, 1993. No. of pages: xv + 244. Price: $39.75. ISBN: 0-8247-8798-6. Statistics in Medicine, 14(1): 103–104, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hill:1995:BRC

[1193] Catherine Hill. Book review: Cancer Incidence in Five Continents. Volume VI. D. M. Parkin, C. S. Muir, S. L. Whelan, Y. T. Gao, J. Ferlay and J. Powell (eds), IARC Scientific Publication No 120, Lyon IARC, 1992. No. of Pages: xxxi + 1033. price: £120. ISBN 92-832-2120-6. Statistics in Medicine, 14(1):104, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1995:BRS

[1194] A. O. Hughes. Book review: Statistical Models in Epidemiology. David Clayton and Michael Hills, Oxford University Press, Oxford, 1993. No. of Pages: vii + 367. Price: £30. ISBN: 0-19-852221-5. Statistics in Medicine, 14(1):104–105, January 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mb

[1195] Anonymous. Masthead. Statistics in Medicine, 14(2):fmi, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:P

[1196] Anonymous. Preface. Statistics in Medicine, 14(2):109–110, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1995:DNF

[1197] Peter Armitage. David Newell: The first 65 years. Statistics in Medicine, 14(2):111–114, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:1995:SCT

[1198] John N. S. Matthews. Small clinical trials: Are they all bad? Statistics in Medicine, 14(2):115–126, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:1995:SIR

[1199] John A. Lewis. Statistical issues in the regulation of medicines. Statistics in Medicine, 14(2):127–136, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1995:MAW

[1200] David R. Jones. Meta-analysis: Weighing the evidence. Statistics in Medicine, 14(2):137–149, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chilvers:1995:CCE

[1201] Clair E. D. Chilvers. Controversies in cancer epidemiology. Statistics in Medicine, 14(2):151–159, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Henderson:1995:PPS

[1202] Robin Henderson. Problems and prediction in survival-data analysis. Statistics in Medicine, 14(2):161–184, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Appleton:1995:WDW

[1203] David R. Appleton. What do we mean by a statistical model? Statistics in Medicine, 14(2):185–197, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simpson:1995:TSN

[1204] Judy M. Simpson. Teaching statistics to non-specialists. Statistics in Medicine, 14(2):199–208, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:1995:LAM

[1205] Stuart J. Pocock. Life as an academic medical statistician and how to survive it. Statistics in Medicine, 14(2):209–222, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:DNP

[1206] Anonymous. David Newell’s publications 1954–1993. Statistics in Medicine, 14(2):223–228, January 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mc

[1207] Anonymous. Masthead. Statistics in Medicine, 14(3):fmi, February 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:1995:AAV

[1208] Donald A. Berry and Stephen G. Eick. Adaptive assignment versus balanced randomization in clinical trials: a decision analysis. Statistics in Medicine, 14(3):231–246, February 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Suissa:1995:BRC

[1209] Samy Suissa and Lucie Blais. Binary regression with continuous outcomes. Statistics in Medicine, 14(3):247–255, February 15, 1995. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wade:1995:ARS

[1210] A. M. Wade, A. E. Ades, A. T. Salt, R. Jayatunga, and P. M. Sonksen. Age-related standards for ordinal data: Modelling the changes in visual acuity from 2 to 9 years of age. Statistics in Medicine, 14(3):257–266, February 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lund:1995:CAM

[1211] Eiliv Lund. Comparison of additive and multiplicative models for reproductive risk factors and post-menopausal breast cancer. Statistics in Medicine, 14(3):267–274, February 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:1995:NPE

[1212] Sin-Ho Jung and John Q. Su. Non-parametric estimation for the difference or ratio of median failure times for paired observations. Statistics in Medicine, 14(3):275–281, February 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang-Clow:1995:SSE

[1213] Fong Wang-Clow, Mary Lange, Nan M. Laird, and James H. Ware. A simulation study of estimators for rates of change in longitudinal studies with attrition. Statistics in Medicine, 14(3):283–297, February 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graham:1995:MCE

[1214] Patrick Graham. Modelling covariate effects in observer agreement studies: The case of nominal scale agreement. Statistics in Medicine, 14(3): 299–310, February 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mauger:1995:CMC

[1215] David T. Mauger, Morton B. Brown, and Robert H. Kushler. A comparison of methods that characterize pulses in a time series. Statistics in Medicine, 14(3):311–325, February 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1995:DBE

[1216] Sander Greenland. Divergent biases in ecologic and individual level studies. Statistics in Medicine, 14(3):327–330, February 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Md

[1217] Anonymous. Masthead. Statistics in Medicine, 14(4):fmi, February 28, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coste:1995:MSP

[1218] Joël Coste, Jacques Fermanian, and Alain Venot. Methodological and statistical problems in the construction of composite measurement scales: a survey of six medical and epidemiological journals. Statistics in Medicine, 14(4):331–345, February 28, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Basu:1995:CSG

[1219] Srabashi Basu and Ayanendranath Basu. Comparison of several goodness-of-fit tests for the kappa statistic based on exact power and coverage probability. Statistics in Medicine, 14(4):347–356, February 28, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:1995:BSM

[1220] Peter F. Thall, Richard M. Simon, and Elihu H. Estey. Bayesian sequential monitoring designs for single-arm clinical trials with multiple outcomes. Statistics in Medicine, 14(4):357–379, February 28, 1995. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:1995:BGS

[1221] Gary L. Rosner and Donald A. Berry. A Bayesian group sequential design for a multiple arm randomized clinical trial. Statistics in Medicine, 14 (4):381–394, February 28, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berkey:1995:RER

[1222] C. S. Berkey, D. C. Hoaglin, F. Mosteller, and G. A. Colditz. A randomeffects regression model for meta-analysis. Statistics in Medicine, 14(4): 395–411, February 28, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

TenHave:1995:AMP

[1223] Thomas R. Ten Have, J. Richard Landis, and Susan L. Weaver. Association models for periodontal disease progression: a comparison of methods for clustered binary data. Statistics in Medicine, 14(4):413–429, February 28, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OGorman:1995:USS

[1224] Thomas W. O’Gorman. Using sign score regression models to select variables in case-control studies. Statistics in Medicine, 14(4):431–443, February 28, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuzick:1995:WTT

[1225] Jack Cuzick. A Wilcoxon-type test for trend. Statistics in Medicine, 14 (4):445–446, February 28, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Me

[1226] Anonymous. Masthead. Statistics in Medicine, 14(5-7):fmi, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:1995:P

[1227] G. David Williamson. Prologue. Statistics in Medicine, 14(5-7):447–448, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:SC

[1228] Anonymous. Staff and committees. Statistics in Medicine, 14(5-7):449– 450, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:PDa

[1229] Anonymous. Presenters and discussants. Statistics in Medicine, 14(5-7): 450–453, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:PDb

[1230] Anonymous. Presenters and discussants. Statistics in Medicine, 14(5-7): 450–453, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dowdle:1995:OR

[1231] Walter R. Dowdle. Opening remarks. Statistics in Medicine, 14(5-7):455, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olkin:1995:KAM

[1232] Ingram Olkin. Keynote addresses. meta-analysis: Reconciling the results of independent studies. Statistics in Medicine, 14(5-7):457–472, March– April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mason:1995:DSU

[1233] Thomas J. Mason. The development of the series of U.S. cancer atlases: Implications for future epidemiologic research. Statistics in Medicine, 14(5-7):473–479, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andes:1995:LPH

[1234] Nancy Andes and Jack E. Davis. Linking public health data using geographic information system techniques: Alaskan community characteristics and infant mortality. Statistics in Medicine, 14(5-7):481–490, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jaro:1995:PLL

[1235] Matthew A. Jaro. Probabilistic linkage of large public health data files. Statistics in Medicine, 14(5-7):491–498, March–April 15–15, 1995. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muse:1995:EQA

[1236] Alison G. Muse, Jaromir Mikl, and Perry F. Smith. Evaluating the quality of anonymous record linkage using deterministic procedures with the new York state AIDS registry and a hospital discharge file. Statistics in Medicine, 14(5-7):499–509, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shevchenko:1995:VIL

[1237] Ivan P. Shevchenko, John T. Lynch, Angela S. Mattie, and Lori L. Reed-Fourquet. Verification of information in a large medical database using linkages with external databases. Statistics in Medicine, 14(5-7):511–530, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1995:CDR

[1238] S. Jay Smith, Samue L. P. Caudill, Karen K. Steinberg, and Stephen B. Thacker. On combining dose-response data from epidemiological studies by meta-analysis. Statistics in Medicine, 14(5-7):531–544, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tweedie:1995:MAA

[1239] R. L. Tweedie and K. L. Mengersen. Meta-analytic approaches to doseresponse relationships, with application in studies of lung cancer and exposure to environmental tobacco smoke. Statistics in Medicine, 14(5-7):545–569, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ezzati:1995:DFD

[1240] Trena M. Ezzati, Keith Hoffman, David R. Judkins, James T. Massey, and Thomas F. Moore. A dual frame design for sampling elderly minorities and persons with disabilities. Statistics in Medicine, 14(5-7): 571–583, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Caudill:1995:MHT

[1241] Samuel P. Caudill and Robert H. Hill. Multiple hypothesis tests in multiple investigations. Statistics in Medicine, 14(5-7):585–589, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mackinnon:1995:DAC

[1242] David P. Mackinnon, Richard Scribner, and Katharine A. Taft. Development and applications of a city-level alcohol availability and alcohol problems database. Statistics in Medicine, 14(5-7):591–604, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nobre:1995:FCM

[1243] Flavio F. Nobre and Marcia M. A. De Macedo. Feasibility of contour mapping epidemiological data with missing values. Statistics in Medicine, 14(5-7):605–613, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pickle:1995:ECA

[1244] Linda W. Pickle and Andrew A. White. Effects of the choice of ageadjustment method on maps of death rates. Statistics in Medicine, 14 (5-7):615–627, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thacker:1995:PHS

[1245] Stephen B. Thacker, Donna F. Stroup, Richard B. Rothenberg, and Ross C. Brownson. Public health surveillance for chronic conditions: a scientific basis for decisions. Statistics in Medicine, 14(5-7):629–641, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Madans:1995:DMM

[1246] Jennifer H. Madans, Cynthia A. Reuben, Sandra T. Rothwell, and Mark S. Eberhardt. Differences in morbidity measures and risk factor identification using multiple data sources: The case of coronary heart disease. Statistics in Medicine, 14(5-7):643–653, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hyatt:1995:PPC

[1247] Raymond R. Hyatt, Jr. and William Rhodes. The price and purity of cocaine: The relationship to emergency room visits and death, and to drug use among arrestees. Statistics in Medicine, 14(5-7):655–668, March– April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Botman:1995:CNH

[1248] Steven L. Botman and Susan S. Jack. Combining national health interview survey datasets: Issues and approaches. Statistics in Medicine, 14(5-7):669–677, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dumouchel:1995:MAD

[1249] William Dumouchel. Meta-analysis for dose-response models. Statistics in Medicine, 14(5-7):679–685, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schwartz:1995:DL

[1250] Harvey A. Schwartz. Data linkage. Statistics in Medicine, 14(5-7):687– 689, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stroup:1995:USD

[1251] Nancy E. Stroup. Utilization 1: Special datasets. Statistics in Medicine, 14(5-7):691–692, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hertzberg:1995:USD

[1252] Vicki Stover Hertzberg. Utilization 2: Special datasets. Statistics in Medicine, 14(5-7):693, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Truman:1995:SDP

[1253] Benedict I. Truman. Study design and public health planning. Statistics in Medicine, 14(5-7):695–696, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:1995:MGD

[1254] Andrew A. White. Mapping and geographic display of data. Statistics in Medicine, 14(5-7):697–699, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Milliken:1995:MI

[1255] George A. Milliken. Methodologic issues. Statistics in Medicine, 14(5-7): 701–702, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Aa

[1256] Anonymous. Abstracts. Statistics in Medicine, 14(5-7):703–708, March– April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:PP

[1257] Anonymous. Poster presentations. Statistics in Medicine, 14(5-7):709– 710, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:SD

[1258] Anonymous. Software demonstrations. Statistics in Medicine, 14(5-7): 711, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:1995:PA

[1259] G. D. Williamson. Presentation of awards. Statistics in Medicine, 14(5-7):713–714, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1995:CR

[1260] Theodore Colton. Closing remarks. Statistics in Medicine, 14(5-7):715– 717, March–April 15–15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mf

[1261] Anonymous. Masthead. Statistics in Medicine, 14(8):fmi, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:EA

[1262] Anonymous. Editorial announcement. Statistics in Medicine, 14(8):719, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kraemer:1995:SIA

[1263] Helena Chmura Kraemer. Statistical issues in assessing comorbidity. Statistics in Medicine, 14(8):721–733, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ford:1995:MII

[1264] Ian Ford, John Norrie, and Susan Ahmadi. Model inconsistency, illustrated by the Cox proportional hazards model. Statistics in Medicine, 14 (8):735–746, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Belin:1995:ARM

[1265] Thomas R. Belin and Donald B. Rubin. The analysis of repeatedmeasures data on schizophrenic reaction times using mixture models. Statistics in Medicine, 14(8):747–768, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:1995:LRC

[1266] Nico J. D. Nagelkerke, Stephen Moses, Francis A. Plummer, Robert C. Brunham, and David Fish. Logistic regression in case-control studies: The effect of using independent as dependent variables. Statistics in Medicine, 14(8):769–775, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moulton:1995:PRA

[1267] Lawrence H. Moulton, Elizabeth A. Holt, Jayakaran S. Job, and Neal A. Halsey. Percentile regression analysis of correlated antibody responses. Statistics in Medicine, 14(8):777–787, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rifkin:1995:ECU

[1268] Robert D. Rifkin. Effects of correlated and uncorrelated measurement error on linear regression and correlation in medical method comparison studies. Statistics in Medicine, 14(8):789–798, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kulldorff:1995:SDC

[1269] Martin Kulldorff and Neville Nagarwalla. Spatial disease clusters: Detection and inference. Statistics in Medicine, 14(8):799–810, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keene:1995:LTS

[1270] Oliver N. Keene. The log transformation is special. Statistics in Medicine, 14(8):811–819, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Macneill:1995:TMM

[1271] I. B. Macneill, J. M. Elwood, D. Miller, and Y. Mao. Trends in mortality from melanoma in Canada and prediction of future rates. Statistics in Medicine, 14(8):821–839, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Markus:1995:AAO

[1272] Richard A. Markus, Jess Frank, Susan Groshen, and Stanley P. Azen. An alternative approach to the optimal design of an LD50 bioassay. Statistics in Medicine, 14(8):841–852, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:1995:BTO

[1273] Jen-Pei Liu and Chung-Sing Weng. Bias of two one-sided tests procedures in assessment of bioequivalence. Statistics in Medicine, 14(8):853–861, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahrer:1995:CTS

[1274] Joanne M. Mahrer and Rhonda C. Magel. A comparison of tests for the k-sample, non-decreasing alternative. Statistics in Medicine, 14(8): 863–871, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1995:SML

[1275] Douglas Altman. Statistics in the medical literature: 1. Statistics in Medicine, 14(8):873–874, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Ca

[1276] Anonymous. Corrections. Statistics in Medicine, 14(8):875, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Therasse:1995:BRC

[1277] P. Therasse. Book review: Clinical Data Management. R. K. Rondel, S. A. Varley and C. F. Webb (eds), Wiley, Chichester, 1993. No. of pages: viii + 295. Price: £35. ISBN 0-471-94092-5. Statistics in Medicine, 14 (8):877, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machin:1995:BRB

[1278] David Machin. Book review: Biostatistics a Methodology for the Health Sciences. Lloyd D. Fisher and Gerald van Belle, Wiley, New York, 1993. No of pages: xxii + 991. Price: £62. ISBN: 0-471-58465-7. Statistics in Medicine, 14(8):878–879, April 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mg

[1279] Anonymous. Masthead. Statistics in Medicine, 14(9):fmi, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moccatti:1995:P

[1280] Dominique Moccatti. Preface. Statistics in Medicine, 14(9):881–882, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:SOC

[1281] Anonymous. Scientific and organizing committees and list of referees. Statistics in Medicine, 14(9):883, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1995:BDP

[1282] John Whitehead and Hazel Brunier. Bayesian decision procedures for dose determining experiments. Statistics in Medicine, 14(9):885–893, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rohmel:1995:RED

[1283] Jachim Röhmel. Re-evaluation of decision on the recommended dose from phase iii dose finding trials. Statistics in Medicine, 14(9):895, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Da

[1284] Anonymous. Discussion. Statistics in Medicine, 14(9):897–899, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wargenau:1995:DFS

[1285] Manfred Wargenau. Dose finding strategies involving interim analyses and unbalanced treatment allocation. Statistics in Medicine, 14(9):901– 907, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Db

[1286] Anonymous. Discussion. Statistics in Medicine, 14(9):909, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moller:1995:ECR

[1287] Susanne Møller. An extension of the continual reassessment methods using a preliminary up-and-down design in a dose finding study in cancer patients, in order to investigate a greater range of doses. Statistics in Medicine, 14(9):911–922, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Dc

[1288] Anonymous. Discussion. Statistics in Medicine, 14(9):923, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Filloon:1995:EMT

[1289] Thomas G. Filloon. Estimating the minimum therapeutically effective dose of a compound via regression modelling and percentile estimation. Statistics in Medicine, 14(9):925–932, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Dd

[1290] Anonymous. Discussion. Statistics in Medicine, 14(9):933, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosen:1995:CBC

[1291] Ori Rosen and Ayala Cohen. Constructing a bootstrap confidence interval for the unknown concentration in radioimmunoassay. Statistics in Medicine, 14(9):935–952, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:De

[1292] Anonymous. Discussion. Statistics in Medicine, 14(9):953, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nedelman:1995:ISE

[1293] Jerry R. Nedelman, Adel H. Karara, Cheng-Tao Chang, Ekaterina Gibiansky, Scott McDonald, and William T. Robinson. Inferring systemic exposure from a pharmacokinetic screen: model-free and model-based approaches. Statistics in Medicine, 14(9):955–968, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Df

[1294] Anonymous. Discussion. Statistics in Medicine, 14(9):969–970, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wakefield:1995:ABP

[1295] Jon Wakefield and Amy Racine-Poon. An application of Bayesian population pharmacokinetic/pharmacodynamic models to dose recommendation. Statistics in Medicine, 14(9):971–986, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Girard:1995:ICF

[1296] Pascal Girard, Silvy Laporte-Simitsidis, Patrick Mismetti, Hervé Decousus, and Jean-Pierre Boissel. Influence of confounding factors on designs for dose-effect relationship estimates. Statistics in Medicine, 14 (9):987–1005, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Dg

[1297] Anonymous. Discussion. Statistics in Medicine, 14(9):1007–1008, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mager:1995:APP

[1298] Harry Mager and Gernot Göller. Analysis of pseudo-profiles in organ pharmacokinetics and toxicokinetics. Statistics in Medicine, 14(9):1009– 1024, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1995:CCT

[1299] B. Jones and J. A. Lewis. The case for cross-over trials in Phase III. Statistics in Medicine, 14(9):1025–1038, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1947].

Gould:1995:PRS

[1300] A. Lawrence Gould. Planning and revising the sample size for a trial. Statistics in Medicine, 14(9):1039–1051, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Dh

[1301] Anonymous. Discussion. Statistics in Medicine, 14(9):1053–1055, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lecoutre:1995:BPA

[1302] Bruno Lecoutre, Gérard Derzko, and Jean-Marie Grouin. Bayesian predictive approach for inference about proportions. Statistics in Medicine, 14(9):1057–1063, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Di

[1303] Anonymous. Discussion. Statistics in Medicine, 14(9):1065, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:1995:URC

[1304] Jen-Pei Liu. Use of the repeated cross-over designs in assessing bioequivalence. Statistics in Medicine, 14(9):1067–1078, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Dj

[1305] Anonymous. Discussion. Statistics in Medicine, 14(9):1079–1080, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Minkin:1995:ASS

[1306] Salomon Minkin and Andrew Chow. Assessing sensitivity in suspension to cytosine arabinoside: statistical analysis, associations with clinical outcome and experimental design. Statistics in Medicine, 14(9):1081– 1095, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Dk

[1307] Anonymous. Discussion. Statistics in Medicine, 14(9):1097, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Enas:1995:DMC

[1308] Gregory G. Enas and David J. Goldstein. Defining, monitoring and combining safety information in clinical trials. Statistics in Medicine, 14 (9):1099–1111, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Dl

[1309] Anonymous. Discussion. Statistics in Medicine, 14(9):1113–1116, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

ONeill:1995:SCP

[1310] Robert T. O’Neill. Statistical concepts in the planning and evaluation of drug safety from clinical trials in drug development: issues of international harmonization. Statistics in Medicine, 14(9):1117–1127, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tuber-Bitter:1995:CBE

[1311] Pascale Tuber-Bitter, Daniel A. Bloch, and Jean-Pierre Raynauld. Comparing the bivariate effects of toxicity and efficacy of treatments. Statistics in Medicine, 14(9):1129–1141, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Dm

[1312] Anonymous. Discussion. Statistics in Medicine, 14(9):1143, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dagostiwo:1995:TSL

[1313] Ralph B. Dagostiwo. A technique for summarizing longitudinal data. Statistics in Medicine, 14(9):1145–1146, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanHouwelingen:1995:BRM

[1314] Hans Van Houwelingen. Book review: Modelling Survival Data in Medical Research. D. Collett, Chapman and Hall, London, 1994. No of pages: xvii + 347. Price: E19.99. ISBN 0-412-44890-4. Statistics in Medicine, 14(9):1147–1148, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanHouwelingen:1995:BRC

[1315] Hans Van Houwelingen. Book review: Case Studies in Occupational Epidemiology. Kyle Steenland (ed.), Oxford University, Press, New York, 1993. No. of Pages: viii + 208. Price: £27.50. ISBN: 0-19-506831-9. Statistics in Medicine, 14(9):1148, May 15–30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mh

[1316] Anonymous. Masthead. Statistics in Medicine, 14(11):fmi, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goodman:1995:SPI

[1317] Steven N. Goodman, Marianna L. Zahurak, and Steven Piantadosi. Some practical improvements in the continual reassessment method for phase I studies. Statistics in Medicine, 14(11):1149–1161, June 15, 1995. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:1995:MTM

[1318] Dean Follmann. Multivariate tests for multiple endpoints in clinical trials. Statistics in Medicine, 14(11):1163–1175, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schnell:1995:RMA

[1319] Daniel J. Schnell, Elvin Magee, and John R. Sheridan. A regression method for analysing ordinal data from intervention trials. Statistics in Medicine, 14(11):1177–1189, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:1995:LSC

[1320] Christopher Cox. Location-scale cumulative odds models for ordinal data: a generalized non-linear model approach. Statistics in Medicine, 14(11):1191–1203, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:1995:ISC

[1321] Yongxiao Wang and Jeremy M. G. Taylor. Inference for smooth curves in longitudinal data with application to an AIDS clinical trial. Statistics in Medicine, 14(11):1205–1218, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:1995:LDA

[1322] P. David Wilson. Longitudinal data analysis for linear Gaussian models with random disturbed-highest-derivative-polynomial subject effects. Statistics in Medicine, 14(11):1219–1233, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anderson:1995:SSL

[1323] Stewart J. Anderson and Richard H. Jones. Smoothing splines for longitudinal data. Statistics in Medicine, 14(11):1235–1248, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zahl:1995:PRM

[1324] Per-Henrik Zahl. A proportional regression model for 20 year survival of colon cancer in Norway. Statistics in Medicine, 14(11):1249–1261, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feinstein:1995:CEA

[1325] A. R. Feinstein. Clinical epidemiology: The architecture of clinical research. Statistics in Medicine, 14(11):1263, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Ab

[1326] Anonymous. Announcement. Statistics in Medicine, 14(11):1264, June 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mi

[1327] Anonymous. Masthead. Statistics in Medicine, 14(12):fmi, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klein:1995:GMP

[1328] John P. Klein, Niels Keiding, and Svend Kreiner. Graphical models for panel studies, illustrated on data from the Framingham Heart Study. Statistics in Medicine, 14(12):1265–1290, June 30, 1995. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marinos:1995:EQE

[1329] Aristides Th. Marinos, Anastasia J. Tzonou, and Menelaos E. Karantzas. Experimental quantiles of epidemiological indices in case-control studies with non-differential misclassification. Statistics in Medicine, 14(12): 1291–1306, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tavare:1995:MPD

[1330] C. J. Tavaré, E. L. Sobel, and F. H. Gilles. Misclassification of a prognostic dichotomous variable: Sample size and parameter estimate adjustment. Statistics in Medicine, 14(12):1307–1314, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vach:1995:LRI

[1331] Werner Vach and Maria Blettner. Logistic regression with incompletely observed categorical covariates — investigating the sensitivity against violation of the missing at random assumption. Statistics in Medicine, 14(12):1315–1329, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tomasson:1995:RSL

[1332] Helgi Tomasson. Risk scores from logistic regression: Unbiased estimates of relative and attributable risk. Statistics in Medicine, 14(12):1331– 1339, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:1995:BCP

[1333] Kyungmann Kim. A bivariate cumulative probit regression model for ordered categorical data. Statistics in Medicine, 14(12):1341–1352, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tzourio:1995:EPG

[1334] C. Tzourio, J. Maccario, and A. Alpérovitch. Estimation of the proportion of genetic cases in late-onset diseases: an application to Alzheimer’s disease. Statistics in Medicine, 14(12):1353–1363, June 30, 1995. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fielding:1995:HTH

[1335] Katherine L. Fielding, Raymond P. Brettle, Sheila M. Gore, Fiona O’Brien, Rona Wyld, J. Roy Robertson, and Rosina Weightman. Heterosexual transmission of HIV analysed by generalized estimating equations. Statistics in Medicine, 14(12):1365–1378, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:1995:RBM

[1336] Joel B. Greenhouse and Larry Waserman. Robust Bayesian methods for monitoring clinical trials. Statistics in Medicine, 14(12):1379–1391, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1995:STS

[1337] S. D. Walter. A simple test for spatial pattern in regional health data. Statistics in Medicine, 14(12):1393–1394, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1995:BRV

[1338] Patrick Royston. Book review: Visualizing Data. William S. Cleveland, Hobart Press, Summit, New Jersey, 1993. No. of pages: 360. Price: $40. ISBN: 0-9634884-0-6. Statistics in Medicine, 14(12):1395–1396, June 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mj

[1339] Anonymous. Masthead. Statistics in Medicine, 14(13):fmi, July 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grady:1995:MST

[1340] James J. Grady and Ronald W. Helms. Model selection techniques for the covariance matrix for incomplete longitudinal data. Statistics in Medicine, 14(13):1397–1416, July 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Royston:1995:CUC

[1341] Patrick Royston. Calculation of unconditional and conditional reference intervals for foetal size and growth from longitudinal measurements. Statistics in Medicine, 14(13):1417–1436, July 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:1995:ERE

[1342] Mimi Y. Kim, Bernard S. Pasternack, Raymond J. Carroll, Karen L. Koenig, and Paolo G. Toniolo. Estimating the reliability of an exposure variable in the presence of confounders. Statistics in Medicine, 14(13): 1437–1446, July 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pickles:1995:CFM

[1343] Andrew Pickles and Robert Crouchley. A comparison of frailty models for multivariate survival data. Statistics in Medicine, 14(13):1447–1461, July 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thomson:1995:HPU

[1344] P. C. Thomson. A hybrid paired and unpaired analysis for the comparison of proportions. Statistics in Medicine, 14(13):1463–1470, July 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1995:CLP

[1345] Kung-Jong Lui. Confidence limits for the population prevalence rate based on the negative binomial distribution. Statistics in Medicine, 14 (13):1471–1477, July 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strauss:1995:ISR

[1346] Nurith Strauss and Richard Simon. Investigating a sequence of randomized phase II trials to discover promising treatments. Statistics in Medicine, 14(13):1479–1489, July 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diehr:1995:BMP

[1347] Paula Diehr, Donald C. Martin, Thomas Koepsell, and Allen Cheadle. Breaking the matches in a paired t-test for community interventions when the number of pairs is small. Statistics in Medicine, 14(13):1491–1504, July 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mk

[1348] Anonymous. Masthead. Statistics in Medicine, 14(14):fmi, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ding:1995:CBC

[1349] Ye Ding. Computing back-calculation estimates of AIDS epidemic. Statistics in Medicine, 14(14):1505–1512, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verdecchia:1995:BCM

[1350] Arduino Verdecchia and Angela B. Mariotto. A back-calculation method to estimate the age and period HIV infection intensity, considering the susceptible population. Statistics in Medicine, 14(14):1513–1530, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duffy:1995:EMS

[1351] Stephen W. Duffy, Hsiu-Hsi Chen, Laszlo Tabar, and Nicholas E. Day. Estimation of mean sojourn time in breast cancer screening using a Markov chain model of both entry to and exit from the preclinical detectable phase. Statistics in Medicine, 14(14):1531–1543, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganguly:1995:AOD

[1352] S. S. Ganguly and U. V. Naik-Nimbalkar. Analysis of ordinal data in a study of endometrial cancer under a matched pairs case-control design. Statistics in Medicine, 14(14):1545–1552, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mauger:1995:TEC

[1353] Elizabeth A. Mauger, Robert A. Wolfe, and Friedrich K. Port. Transient effects in the Cox proportional hazards regression model. Statistics in Medicine, 14(14):1553–1565, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahn:1995:ECO

[1354] Chul Ahn and Tamara Odom-Maryon. Estimation of a common odds ratio under binary cluster sampling. Statistics in Medicine, 14(14):1567– 1576, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1626].

Kolassa:1995:CSP

[1355] John E. Kolassa. A comparison of size and power calculations for the Wilcoxon statistic for ordered categorical data. Statistics in Medicine, 14(14):1577–1581, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roebruck:1995:CTS

[1356] Peter Roebruck and Andreas Kühn. Comparison of tests and sample size formulae for proving therapeutic equivalence based on the difference of binomial probabilities. Statistics in Medicine, 14(14):1583–1594, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:1995:AME

[1357] P. Bauer and J. Röhmel. An adaptive method for establishing a doseresponse relationship. Statistics in Medicine, 14(14):1595–1607, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schouten:1995:RRR

[1358] E. G. Schouten, J. M. Dekker, F. J. Kok, S. Lecessie, H. C. Van Houwelingen, J. Pool, and J. P. Vandenbroucke. Risk ratio and rate ratio estimation in case-cohort designs: Hypertension and cardiovascular mortality. Statistics in Medicine, 14(14):1609–1610, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berlin:1995:BRC

[1359] Jesse A. Berlin. Book review: Combining Information: Statistical Issues and Opportunities for Research. David Draper, Donald P. Gaver Jun, Prem K. Goel, Joel B. Greenhouse, Larry V. Hedges, Carl N. Morris, John R. Tucker and Christine M. Waternaux, American Statistical Association, 1993. No. of Pages: xi + 217. Price: $25. ISBN: 0-309-04730-7. Statistics in Medicine, 14(14):1611–1612, July 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Ml

[1360] Anonymous. Masthead. Statistics in Medicine, 14(15):fmi, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1995:Eb

[1361] Ted Colton, Tony Johnson, and David Machin. Editorial. Statistics in Medicine, 14(15):1613, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1995:SPC

[1362] Jonas H. Ellenberg. Some prespectives on the career of Samuel W. Greehouse: The first 75 years. Statistics in Medicine, 14(15):1615–1619, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geisser:1995:TIW

[1363] Seymour Geisser and Wesley Johnson. Testing independence when the form of the bivariate distribution is unspecified. Statistics in Medicine, 14(15):1621–1639, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gastwirth:1995:BCM

[1364] Joseph L. Gastwirth and Samuel W. Greenhouse. Biostatistical concepts and methods in the legal setting. Statistics in Medicine, 14(15):1641– 1653, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:1995:BMC

[1365] John A. Lewis, David R. Jones, and Joachim Röhmel. Biostatistical methodology in clinical trials — a European guideline. Statistics in Medicine, 14(15):1655–1657, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:BMC

[1366] Anonymous. Biostatistical methodology in clinical trials in applications for marketing authorizations for medicinal products. CPMP working party on efficacy of medicinal products note for guidance III/3630/92-EN. Statistics in Medicine, 14(15):1659–1682, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesser:1995:BMR

[1367] Martin L. Lesser and Robert A. Parker. The biostatistician in medical research: Allocating time and effort. Statistics in Medicine, 14(15):1683– 1692, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:1995:EMM

[1368] Robert H. Lyles and Lloyd E. Chambless. Effects of model misspecification in the estimation of variance components and intraclass correlation for paired data. Statistics in Medicine, 14(15):1693–1706, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hess:1995:GMA

[1369] Kenneth R. Hess. Graphical methods for assessing violations of the proportional hazards assumption in Cox regression. Statistics in Medicine, 14(15):1707–1723, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1995:BRAa

[1370] Niels Keiding, Peter Yuen, and David Machin. Book review: Atlas of cancer incidence in England and Wales: 1968-85. Anthony Swerdlow and Isabel dos Santos Silva, Oxford Medical Publications, Oxford 1993. Number of pages: xvii + 275. Price: £170. ISBN: 0-19-26199-3-4. Statistics in Medicine, 14(15):1725–1726, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1995:BRAb

[1371] Niels Keiding and J. K. Lindsey. Book review: Analysis of systematic and random differences between paired ordinal categorical data. Elisabeth Svensson, Almqvist and Wiksell, Göteberg, 1993. No. of pages: 562. Price: 382 Swedish Kr. ISBN: 91-22-01587-6. Statistics in Medicine, 14 (15):1726–1727, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1995:BRR

[1372] Niels Keiding and Stephen Senn. Book review: Risk. John Adams, UCL Press, London, 1995. No. of pages xii + 228. ISBN 1-85728-067-9 HB, 1-85728-068-7 PB. Statistics in Medicine, 14(15):1727–1728, August 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mm

[1373] Anonymous. Masthead. Statistics in Medicine, 14(16):fmi, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1995:Ec

[1374] Ted Colton, Tony Johnson, and David Machin. Editorial. Statistics in Medicine, 14(16):1729, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cupples:1995:AAS

[1375] L. Adrienne Cupples, David R. Gagnon, Ratna Ramaswamy, and Ralph B. D’Agostino. Age-adjusted survival curves with application in the Framingham Study. Statistics in Medicine, 14(16):1731–1744, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Larson:1995:ACR

[1376] Martin G. Larson. Assessment of cardiovascular risk factors in the elderly: The Framingham Heart Study. Statistics in Medicine, 14(16): 1745–1756, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:1995:DHR

[1377] Ralph B. D’Agostino, Albert J. Belanger, Elizabeth W. Markson, Maggie Kelly-Hayes, and Philip A. Wolf. Development of health risk appraisal functions in the presence of multiple indicators: The Framingham Study nursing home institutionalization model. Statistics in Medicine, 14(16):1757–1770, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:1995:QAS

[1378] Bernard F. Cole and Richard D. Gelber. A quality-adjusted survival meta-analysis of adjuvant chemotherapy for premenopausal breast cancer. Statistics in Medicine, 14(16):1771–1784, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cohen:1995:AGF

[1379] Ayala Cohen and Ofra Barnett. Assessing goodness-of-fit of parametric regression models for lifetime data — graphical methods. Statistics in Medicine, 14(16):1785–1795, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carothers:1995:MMA

[1380] Andrew D. Carothers. Methods for maternal age-standardization of the incidence of congenital abnormalities. Statistics in Medicine, 14 (16):1797–1806, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abt:1995:PSV

[1381] Markus Abt, Erwin Strehler, Bernd Rosenbusch, and Karl Sterzik. Predicting the success of in vitro fertilization: Conventional semen analysis compared to the hamster ova penetration test. Statistics in Medicine, 14 (16):1807–1818, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:1995:ADS

[1382] Odd O. Aalen, Espen Bjertness, and Torleif Sønju. Analysis of dependent survival data applied to lifetimes of amalgam fillings. Statistics in Medicine, 14(16):1819–1829, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:1995:SSO

[1383] Ying Lu and Judy A. Bean. On the sample size for one-sided equivalence of sensitivities based upon McNemar’s test. Statistics in Medicine, 14 (16):1831–1839, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Cb

[1384] Anonymous. Corrections. Statistics in Medicine, 14(16):1841, August 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mn

[1385] Anonymous. Masthead. Statistics in Medicine, 14(17):fmi, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murphy:1995:AMD

[1386] Susan A. Murphy, Gillian R. Bentley, and Mary Ann O’Hanesian. An analysis for menstrual data with time-varying covariates. Statistics in Medicine, 14(17):1843–1857, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Congdon:1995:MFA

[1387] Peter Congdon. Modelling frailty in area mortality. Statistics in Medicine, 14(17):1859–1874, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Swift:1995:SCI

[1388] Michael B. Swift. Simple confidence intervals for standardized rates based on the approximate bootstrap method. Statistics in Medicine, 14(17): 1875–1888, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Paik:1995:DCI

[1389] Myunghee Cho Paik, Melissa Dowd Begg, Wafaa El-Sadr, Jack Gorman, and Zena Stein. Difference in clinical implications of CD4 counts among HIV-infected homosexual men and injection drug using men and women. Statistics in Medicine, 14(17):1889–1900, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:1995:CTC

[1390] William D. Johnson and Warren L. May. Combining 2×2 tables that contain structural zeros. Statistics in Medicine, 14(17):1901–1911, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lavori:1995:MIS

[1391] Philip W. Lavori, Ree Dawson, and David Shera. A multiple imputation strategy for clinical trials with truncation of patient data. Statistics in Medicine, 14(17):1913–1925, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hallstrom:1995:TVS

[1392] Al Hallstrom, Ruth McBride, and Rich Moore. Toward vital status sweeps: a case history in sequential monitoring. Statistics in Medicine, 14(17):1927–1931, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Browne:1995:UPS

[1393] Richard H. Browne. On the use of a pilot sample for sample size determination. Statistics in Medicine, 14(17):1933–1940, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Komornik:1995:ATC

[1394] J. Komornik, M. Mikulecky, and M. Beno. Additivity test for composition of binomial effects in chromosome aberrations after radiation injury. Statistics in Medicine, 14(17):1941–1947, September 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mo

[1395] Anonymous. Masthead. Statistics in Medicine, 14(18):fmi, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunne:1995:DDL

[1396] Adrian Dunne. Decision and detection limits for linear homoscedastic assays. Statistics in Medicine, 14(18):1949–1959, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gange:1995:CMC

[1397] Stephen J. Gange, Kathryn L. P. Linton, Alastair J. Scott, David L. Demets, and Ronald Klein. A comparison of methods for correlated ordinal measures with ophthalmic applications. Statistics in Medicine, 14(18):1961–1974, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:1995:MSM

[1398] Guillermo Marshall and Richard H. Jones. Multi-state models and diabetic retinopathy. Statistics in Medicine, 14(18):1975–1983, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dignam:1995:MBM

[1399] James J. Dignam, Lisa A. Weissfeld, and Stewart J. Anderson. Methods for bounding the marginal survival distribution. Statistics in Medicine, 14(18):1985–1998, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanHouwelingen:1995:CVU

[1400] Hans C. van Houwelingen and Jane Thorogood. Construction, validation and updating of a prognostic model for kidney graft survival. Statistics in Medicine, 14(18):1999–2008, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindstrom:1995:SMR

[1401] Mary J. Lindstrom. Self-modelling with random shift and scale parameters and a free-knot spline shape function. Statistics in Medicine, 14 (18):2009–2021, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hillis:1995:RPC

[1402] Stephen L. Hillis. Residual plots for the censored data linear regression model. Statistics in Medicine, 14(18):2023–2036, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nam:1995:SSD

[1403] Jun-Mo Nam. Sample size determination in stratified trials to establish the equivalence of two treatments. Statistics in Medicine, 14(18):2037– 2049, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1995:NCC

[1404] Kung-Jong Lui. Notes on conditional confidence limits under inverse sampling. Statistics in Medicine, 14(18):2051–2056, September 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mp

[1405] Anonymous. Masthead. Statistics in Medicine, 14(19):fmi, October 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stewart:1995:PMM

[1406] Lesley A. Stewart and Cochrane Working Group On Meta-Analysis Using Individual Patient Data. Practical methodology of meta-analyses (overviews) using updated individual patient data. Statistics in Medicine, 14(19):2057–2079, October 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:1995:DAR

[1407] Richard J. Cook. The design and analysis of randomized trials with recurrent events. Statistics in Medicine, 14(19):2081–2098, October 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shannon:1995:ADT

[1408] William D. Shannon, John Bryant, Theodore F. Logan, and Roger Day. An application of decision theory to patient screening for an autologous tumour vaccine trial. Statistics in Medicine, 14(19):2099–2110, October 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1995:TSR

[1409] Jae Won Lee. Two-sample rank tests for acceleration in cure models. Statistics in Medicine, 14(19):2111–2118, October 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Herndon:1995:RCS

[1410] James E. Herndon II and Frank E. Harrell, Jr. The restricted cubic spline as baseline hazard in the proportional hazards model with step function time-dependent covariables. Statistics in Medicine, 14(19):2119–2129, October 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joffe:1995:SEC

[1411] Marshall M. Joffe and Sander Greenland. Standardized estimates from categorical regression models. Statistics in Medicine, 14(19):2131–2141, October 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehta:1995:ELR

[1412] Cyrus R. Mehta and Nitin R. Patel. Exact logistic regression: Theory and examples. Statistics in Medicine, 14(19):2143–2160, October 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hosmer:1995:CIE

[1413] David W. Hosmer and Stanley Lemeshow. Confidence interval estimates of an index of quality performance based on logistic regression models. Statistics in Medicine, 14(19):2161–2172, October 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1872].

Anonymous:1995:Mq

[1414] Anonymous. Masthead. Statistics in Medicine, 14(20):fmi, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stangl:1995:PDM

[1415] Dalene K. Stangl. Prediction and decision making using Bayesian hierarchical models. Statistics in Medicine, 14(20):2173–2190, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ritter:1995:MMM

[1416] C. Ritter, A. Bouckaert, M. van Lierde, and I. Theunissen. Modelling mortality and morbidity of newborns. Statistics in Medicine, 14(20): 2191–2204, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:1995:LMA

[1417] Ralph B. D’Agostino, Jr., David Sparrow, Scott Weiss, and Bernard Rosner. Longitudinal models for analysis of respiratory function. Statistics in Medicine, 14(20):2205–2216, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:1995:SCT

[1418] Sin-Ho Jung, Sam Wieand, and Stephen S. Cha. A statistic for comparing two correlated markers which are prognostic for time to an event. Statistics in Medicine, 14(20):2217–2225, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

MoriBrooks:1995:SSU

[1419] Maria Mori Brooks, Al Hallstrom, and Monika Peckova. A simulation study used to design the sequential monitoring plan for a clinical trial. Statistics in Medicine, 14(20):2227–2237, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:1995:RED

[1420] Weichung Joseph Shih and A. Lawrence Gould. Re-evaluating design specifications of longitudinal clinical trials without unblinding when the key response is rate of change. Statistics in Medicine, 14(20):2239–2248, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hornberger:1995:DCE

[1421] John C. Hornberger, Byron Wm. Brown, Jr., and Jerry Halpern. Designing a cost-effective clinical trial. Statistics in Medicine, 14(20):2249– 2259, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:1995:EPC

[1422] Man-Lai Tang, Karim F. Hirji, and Stein E. Vollset. Exact power computation for dose-response studies. Statistics in Medicine, 14(20):2261– 2272, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahnn:1995:SSD

[1423] Sang Ahnn and Stewart J. Anderson. Sample size determination for comparing more than two survival distributions. Statistics in Medicine, 14(20):2273–2282, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1995:DAC

[1424] Stephen Senn. In defence of analysis of covariance: a reply to Chambless and Roeback. Statistics in Medicine, 14(20):2283–2285, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [781, 1152].

Guggenmoss-Holzmann:1995:MCE

[1425] Irene Guggenmoss-Holzmann. Modelling covariate effects in observer agreement studies: The case of nominal scale agreement. Statistics in Medicine, 14(20):2285–2288, October 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mr

[1426] Anonymous. Masthead. Statistics in Medicine, 14(21-22):fmi, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:1995:P

[1427] Andrew B. Lawson, Lance A. Waller, and Annibale Biggeri. Preface. Statistics in Medicine, 14(21-22):2289, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waller:1995:PFT

[1428] Lance A. Waller and Andrew B. Lawson. The power of focused tests to detect disease clustering. Statistics in Medicine, 14(21-22):2291–2308, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bithell:1995:CTD

[1429] J. F. Bithell. The choice of test for detecting raised disease risk near a point source. Statistics in Medicine, 14(21-22):2309–2322, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:1995:CTD

[1430] Toshiro Tango. A class of tests for detecting ‘general’ and ‘focused’ clustering of rare diseases. Statistics in Medicine, 14(21-22):2323–2334, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kelsall:1995:NPE

[1431] Julia E. Kelsall and Peter J. Diggle. Non-parametric estimation of spatial variation in relative risk. Statistics in Medicine, 14(21-22):2335–2342, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacquez:1995:MCP

[1432] Geoffrey M. Jacquez. The map comparison problem: Tests for the overlap of geographic boundaries. Statistics in Medicine, 14(21-22):2343–2361, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cislaghi:1995:ETD

[1433] Cesare Cislaghi, Annibale Biggeri, Mario Braga, Corrado Lagazio, and Marco Marchi. Exploratory tools for disease mapping in geographical epidemiology. Statistics in Medicine, 14(21-22):2363–2381, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:1995:TDS

[1434] Andrew B. Lawson and Jean Francois Viel. Tests for directional spacetime interaction in epidemiological data. Statistics in Medicine, 14(21-22):2383–2391, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donnelly:1995:SAC

[1435] Christl A. Donnelly. The spatial analysis of covariates in a study of environmental epidemiology. Statistics in Medicine, 14(21-22):2393–2409, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernardinelli:1995:BED

[1436] L Bernardinelli, D. Clayton, and C. Montomoli. Bayesian estimates of disease maps: How important are priors? Statistics in Medicine, 14(21-22):2411–2431, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernardinelli:1995:BAS

[1437] L. Bernardinelli, D. Clayton, C. Pascutto, C. Montomoli, M. Ghislandi, and M. Songini. Bayesian analysis of space-time variation in disease risk. Statistics in Medicine, 14(21-22):2433–2443, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ayuthya:1995:TAM

[1438] Rampai Suksawasdi Na Ayuthya and Dankmar Böhning. Traffic accident mapping in Bangkok metropolis: a case study. Statistics in Medicine, 14(21-22):2445–2458, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Viel:1995:ILY

[1439] Jean-François Viel, Dominique Pobel, and André Carré. Incidence of leukaemia in young people around the La Hague nuclear waste reprocessing plant: a sensitivity analysis. Statistics in Medicine, 14(21-22): 2459–2472, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawson:1995:MMP

[1440] Andrew B. Lawson. MCMC methods for putative pollution source problems in environmental epidemiology. Statistics in Medicine, 14(21-22): 2473–2485, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:1995:SVN

[1441] Sylvia Richardson, Christine Monfort, Martyn Green, Gerald Draper, and Colin Muirhead. Spatial variation of natural radiation and childhood leukaemia incidence in great Britain. Statistics in Medicine, 14(21-22): 2487–2501, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haybittle:1995:LE

[1442] J. Haybittle, P. Yuen, and D. Machin. Letter to the editor. Statistics in Medicine, 14(21-22):2503–2505, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1995:BRT

[1443] Niels Keiding. Book review: Trends in cancer incidence and mortality (Cancer surveys vol. 19/20). R. Doll, J. F. Fraumei, Jr. and C. S. Muir (Eds), Cold Spring Harbor Laboratory Press, Plainview NY, 1994. No. of pages: vi + 583. Price $144. ISBN: 0-87969-391-6. Statistics in Medicine, 14(21-22):2507–2508, November 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Ms

[1444] Anonymous. Masthead. Statistics in Medicine, 14(23):fmi, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang-Stein:1995:MCP

[1445] Christy Chuang-Stein and Donald M. Tong. Multiple comparisons procedures for comparing several treatments with a control based on binary data. Statistics in Medicine, 14(23):2509–2522, December 15, 1995. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gustafson:1995:BAB

[1446] Paul Gustafson. A Bayesian analysis of bivariate survival data from a multicentre cancer clinical trial. Statistics in Medicine, 14(23):2523– 2535, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karrison:1995:CMS

[1447] Theodore Karrison. Comparison of median survival times with adjustment for covariates. Statistics in Medicine, 14(23):2537–2553, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tang:1995:LRT

[1448] Ming-Xin Tang, Wei-Yann Tsai, Karen Marder, and Richard Mayeux. Linear rank tests for doubly censored data. Statistics in Medicine, 14 (23):2555–2563, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coste:1995:NAS

[1449] Joël Coste, Eric Walter, and Alain Venot. A new approach to selection and weighting of items in evaluative composite measurement scales. Statistics in Medicine, 14(23):2565–2580, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dahm:1995:DVA

[1450] P. Fred Dahm, Mitchel H. Gail, Philip S. Rosenberg, and David Pee. Determining the value of additional surrogate exposure data for improving the estimate of an odds ratio. Statistics in Medicine, 14(23):2581–2598, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Young:1995:UMM

[1451] Philip Young, Byron Morgan, Peter Sonksen, Sebastian Till, and Charles Williams. Using a mixture model to predict the occurrence of diabetic retinopathy. Statistics in Medicine, 14(23):2599–2608, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fedorov:1995:OPD

[1452] V. V. Fedorov, P. Hackl, and W. G. Müller. Optimal and practicable designs for measuring plaque pH-profiles. Statistics in Medicine, 14(23): 2609–2617, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fleming:1995:SMA

[1453] T. Fleming, N. E. Day, S. W. Duffy, and Laszlo Tabar. Surrogate markers in AIDS and cancer trials. Statistics in Medicine, 14(23):2619–2620, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1995:BRS

[1454] Niels Keiding and Jørgen Hilden. Book review: Statistics an interactive text for the health and life sciences. Goteti Bala Krishnamurty, Patricia Kasovia-Schmitt and David J. Ostroff, Joness and Bartlett Publishers, Boston, 1995. No. of pages: xxiv + 709. Price £14.95. ISBN: 0-86720-898-6. Statistics in Medicine, 14(23):2621–2622, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1995:BRF

[1455] Niels Keiding and Jørgen Hilden. Book review: Facts, figures, and fallacies. Epidemiological pitfalls in medical research. The Lance (editor Judith R. Glynn), The Lancet Ltd., London, 1994. No. of pages: vi + 28. Price £5.95. ISBN: 0-900511-168. Statistics in Medicine, 14(23):2622– 2623, December 15, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:Mt

[1456] Anonymous. Masthead. Statistics in Medicine, 14(24):fmi, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1995:P

[1457] Ted Colton, Tony Johnson, and David Machin. Preface. Statistics in Medicine, 14(24):2625, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1995:ISC

[1458] Anonymous. International Society for Clinical Biostatistics Fifteenth International Meeting held at the Convention Centre Basel 25–29 July 1994. Statistics in Medicine, 14(24):2626, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rawson:1995:VRI

[1459] Nigel S. B. Rawson and Edith Malcolm. Validity of the recording of ischaemic heart disease and chronic obstructive pulmonary disease in the Saskatchewan health care datafiles. Statistics in Medicine, 14(24): 2627–2643, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:1995:PIS

[1460] Roger J. Marshall. A program to implement a search method for identification of clinical subgroups. Statistics in Medicine, 14(24):2645–2659, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1995:PVS

[1461] Stephen Senn. A personal view of some controversies in allocating treatment to patients in clinical trials. Statistics in Medicine, 14(24):2661– 2674, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1476].

Choi:1995:ENR

[1462] S. C. Choi and I. L. Lu. Effect of non-random missing data mechanisms in clinical trials. Statistics in Medicine, 14(24):2675–2684, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1995:BAR

[1463] Teresa C. Smith, David J. Spiegelhalter, and Andrew Thomas. Bayesian approaches to random-effects meta-analysis: a comparative study. Statistics in Medicine, 14(24):2685–2699, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lancar:1995:NPM

[1464] Rémi Lancar, Andrew Kramar, and Christine Haie-Meder. Nonparametric methods for analysing recurrent complications of varying severity. Statistics in Medicine, 14(24):2701–2712, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:1995:NPE

[1465] Jian-Lun Xu and Philip C. Prorok. Non-parametric estimation of the post-lead-time survival distribution of screen-detected cancer cases. Statistics in Medicine, 14(24):2715–2725, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gottschau:1995:ICS

[1466] Adam Gottschau and Birthe Hogh. Interval censored survival data and multistate compartmental models in the analysis of first appearance of Plasmodium falciparum parasites in infants. Statistics in Medicine, 14 (24):2727–2736, December 30, 1995. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Ma

[1467] Anonymous. Masthead. Statistics in Medicine, 15(1):fmi, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sahai:1996:FTD

[1468] Hardeo Sahai and Anwer Khurshid. Formulae and tables for the determination of sample sizes and power in clinical trials for testing differences in proportions for the two-sample design: a review. Statistics in Medicine, 15(1):1–21, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Variyam:1996:MNK

[1469] Jayachandran N. Variyam, James Blaylock, and David M. Smallwood. Modelling nutrition knowledge, attitudes, and diet-disease awareness: the case of dietary fibre. Statistics in Medicine, 15(1):23–35, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhang:1996:TBM

[1470] Heping Zhang, Theodore Holford, and Michael B. Bracken. A tree-based method of analysis for prospective studies. Statistics in Medicine, 15(1):

37–49, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhao:1996:HEV

[1471] Guolong Zhao. The homogenetic estimate for the variance of survival rate. Statistics in Medicine, 15(1):51–60, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1996:PAL

[1472] Richard H. Jones. Polynomials with asymptotes for longitudinal data. Statistics in Medicine, 15(1):61–74, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mcneil:1996:SAZ

[1473] Alexander J. Mcneil and Sheila M. Gore. Statistical analysis of zidovudine (AZT) effect on CD4 cell counts in HIV disease. Statistics in Medicine, 15(1):75–92, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hwang:1996:EEQ

[1474] Jing-Shiang Hwang, Jau-Yih Tsauo, and Jung-Der Wang. Estimation of expected quality adjusted survival by cross-sectional survey. Statistics in Medicine, 15(1):93–102, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilsenbeck:1996:PVA

[1475] Susan Galloway Hilsenbeck and Gary M. Clark. Practical p-value adjustment for optimally selected cutpoints. Statistics in Medicine, 15(1): 103–112, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cappelleri:1996:LEI

[1476] Joseph C. Cappelleri and Stephen Senn. Letter to the Editor including author’s reply: A personal view of some controversies in allocating treatment to patients in clinical trials, by Stephen Senn, Statistics in Medicine, 14, 2661–2674 (1995). Statistics in Medicine, 15(1):113–116, January 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1461].

Anonymous:1996:Mb

[1477] Anonymous. Masthead. Statistics in Medicine, 15(2):fmi, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsai:1996:RPA

[1478] K. T. Tsai and H. L. Patel. Robust procedures for analysing a two-period cross-over design with baseline measurements. Statistics in Medicine, 15 (2):117–126, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:1996:AID

[1479] Barbra A. Richardson and Virginia F. Flack. The analysis of incomplete data in the three-period two-treatment cross-over design for clinical trials. Statistics in Medicine, 15(2):127–143, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Podgor:1996:ATL

[1480] Marvin J. Podgor and Rita Hiller. Associations of types of lens opacities between and within eyes of individuals: an application of second-order generalized estimating equations. Statistics in Medicine, 15(2):145–156, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chiogna:1996:ECE

[1481] M. Chiogna, D. J. Spiegelhalter, R. C. G. Franklin, and K. Bull. An empirical comparison of expert-derived and data-derived classification trees. Statistics in Medicine, 15(2):157–169, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karrison:1996:CIM

[1482] Theodore Karrison. Confidence intervals for median survival times under a piecewise exponential model with proportional hazards covariate effects. Statistics in Medicine, 15(2):171–182, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abt:1996:ASH

[1483] Klaus Abt and Albert Gülich. Assessment of small health risks based on exact sample sizes. Statistics in Medicine, 15(2):183–195, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:1996:CHI

[1484] Wai-Yuan Tan, Sho Rong Lee, and Si Chin Tang. Characterization of HIV incubation distributions and some comparative studies. Statistics in Medicine, 15(2):197–220, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kault:1996:SDN

[1485] David Kault. The shape of the distribution of the number of sexual partners. Statistics in Medicine, 15(2):221–230, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelson:1996:EIR

[1486] Lorene M. Nelson. The epidemiological imagination. a reader. Statistics in Medicine, 15(2):231–232, January 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mc

[1487] Anonymous. Masthead. Statistics in Medicine, 15(3):fmi, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hung:1996:GTC

[1488] H. M. James Hung. Global tests for combination drug studies in factorial trials. Statistics in Medicine, 15(3):233–247, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:1996:SRC

[1489] Ian R. White and Stuart J. Pocock. Statistical reporting of clinical trials with individual changes from allocated treatment. Statistics in Medicine, 15(3):249–262, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmoor:1996:RNR

[1490] Claudia Schmoor, Manfred Olschewski, and Martin Schumacher. Randomized and non-randomized patients in clinical trials: experiences with comprehensive cohort studies. Statistics in Medicine, 15(3):263–271, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1996:AAP

[1491] Wen C. Lee and Ruey S. Lin. Autoregressive age-period-cohort models. Statistics in Medicine, 15(3):273–281, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:1996:ICS

[1492] C. P. Farrington. Interval censored survival data: a generalized linear modelling approach. Statistics in Medicine, 15(3):283–292, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:1996:MEM

[1493] Bernard A. Rosner. Measurement error models for ordinal exposure variables measured with error. Statistics in Medicine, 15(3):293–303, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sellero:1996:RDR

[1494] C. Sánchez Sellero, E. Vázquez Fernández, W. González Manteiga, X. L. Otero, X. Hervada, E. Fernández, and X. A. Taboada. Reporting delay: a review with a simulation study and application to Spanish AIDS data. Statistics in Medicine, 15(3):305–321, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1996:GLM

[1495] Philip J. Smith, Theodore J. Thompson, Michael M. Engelgau, and William H. Herman. A generalized linear model for analysing receiver operating characteristic curves. Statistics in Medicine, 15(3):323–333, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1869].

Walter:1996:CST

[1496] S. D. Walter. Corrections: a simple test for spatial pattern in regional health data. Statistics in Medicine, 15(3):335–336, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wargenau:1996:DFS

[1497] Manfred Wargenau and Gerhard Hommel. Dose finding strategies involving interim analysis and unbalanced treatment allocation. Statistics in Medicine, 15(3):335–336, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldman:1996:MCD

[1498] Anne I. Goldman. Methodology of clinical drug trials. Statistics in Medicine, 15(3):337–338, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blettner:1996:LRS

[1499] Maria Blettner. Logistic regression. a self-learning text. Statistics in Medicine, 15(3):339, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Volund:1996:PNE

[1500] Aage Vølund. Physiology by numbers. an encouragement to quantitative thinking. Statistics in Medicine, 15(3):340, February 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Md

[1501] Anonymous. Masthead. Statistics in Medicine, 15(4):fmi, February 29, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:1996:E

[1502] Ralph B. D’Agostino. Editorial. Statistics in Medicine, 15(4):341, February 29, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:1996:DSD

[1503] Weng Kee Wong and Peter A. Lachenbruch. Designing studies for dose response. Statistics in Medicine, 15(4):343–359, February 29, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harrell:1996:MPM

[1504] Frank E. Harrell, Jr., Kerry L. Lee, and Daniel B. Mark. Multivariable prognostic models: issues in developing models, evaluating assumptions and adequacy, and measuring and reducing errors. Statistics in Medicine, 15(4):361–387, February 29, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lange:1996:SAH

[1505] Nicholas Lange. Statistical approaches to human brain mapping by functional magnetic resonance imaging. Statistics in Medicine, 15(4):389–428, February 29, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:1996:NPT

[1506] Michael P. Fay and Chris Gennings. Non-parametric two-sample tests for repeated ordinal responses. Statistics in Medicine, 15(4):429–442, February 29, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1996:MCT

[1507] J. K. Lindsey and B. Jones. A model for cross-over trials evaluating therapeutic preferences. Statistics in Medicine, 15(4):443–447, February 29, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Me

[1508] Anonymous. Masthead. Statistics in Medicine, 15(5):fmi, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kosorok:1996:ICF

[1509] Michael R. Kosorok, Wen-Hsiang Wei, and Philip M. Farrell. The incidence of cystic fibrosis. Statistics in Medicine, 15(5):449–462, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Epstein:1996:BIP

[1510] Leonardo D. Epstein, Alvaro Muñoz, and David He. Bayesian imputation of predictive values when covariate information is available and gold standard diagnosis is unavailable. Statistics in Medicine, 15(5):463–476, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sullivan:1996:RPA

[1511] Lisa M. Sullivan and Ralph B. D’Agostino. Robustness and power of analysis of covariance applied to data distorted from normality by floor effects: homogeneous regression slopes. Statistics in Medicine, 15(5): 477–496, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marschner:1996:AHT

[1512] Ian C. Marschner. Assessment of HIV testing rates among HIV infected individuals using incidence data on HIV and AIDS diagnoses. Statistics in Medicine, 15(5):497–511, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1996:HHI

[1513] Sander Greenland. Historical HIV incidence modelling in regional subgroups: use of flexible discrete models with penalized splines based on prior curves. Statistics in Medicine, 15(5):513–525, March 15, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1996:ACT

[1514] J. K. Lindsey, B. Jones, and J. A. Lewis. Analysis of cross-over trials for duration data. Statistics in Medicine, 15(5):527–535, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berkey:1996:MOM

[1515] C. S. Berkey, J. J. Anderson, and D. C. Hoaglin. Multiple-outcome meta-analysis of clinical trials. Statistics in Medicine, 15(5):537–557, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1996:SML

[1516] Editor: Douglas Altman. Statistics in the medical literature: 2. Statistics in Medicine, 15(5):559, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carroll:1996:MRC

[1517] R. J. Carroll. Measurement, regression and calibration. Statistics in Medicine, 15(5):561–562, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1996:QMB

[1518] Peter Armitage. Quantitative methods in biological and medical sciences: a historical essay. Statistics in Medicine, 15(5):562–563, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:1996:MAE

[1519] Niels G. Becker. Modeling the AIDS epidemic: planning, policy and prediction. Statistics in Medicine, 15(5):563–564, March 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mf

[1520] Anonymous. Masthead. Statistics in Medicine, 15(6):fmi, March 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holford:1996:MEC

[1521] Theodore R. Holford, Zhongxin Zhang, Tongzhang Zheng, and Lisa A. Mckay. A model for the effect of cigarette smoking on lung cancer incidence in Connecticut. Statistics in Medicine, 15(6):565–580, March 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1996:MSH

[1522] Peter N. Lee and Barbara A. Forey. Misclassification of smoking habits as a source of bias in the study of environmental tobacco smoke and lung cancer. Statistics in Medicine, 15(6):581–605, March 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:1996:STA

[1523] Michael E. Miller, Margaret K. James, Carl D. Langefeld, Mark A. Espeland, Jay A. Freedman, Douglas K. Martin, and David M. Smith. Some techniques for the analysis of work sampling data. Statistics in Medicine, 15(6):607–618, March 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hardy:1996:LAM

[1524] Rebecca J. Hardy and Simon G. Thompson. A likelihood approach to meta-analysis with random effects. Statistics in Medicine, 15(6):619–629, March 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilton:1996:AWT

[1525] Joan F. Hilton. The appropriateness of the Wilcoxon test in ordinal data. Statistics in Medicine, 15(6):631–645, March 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feltz:1996:ATE

[1526] Carol J. Feltz and G. Edward Miller. An asymptotic test for the equality of coefficients of variation from k populations. Statistics in Medicine, 15(6):647–658, March 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:1996:IPC

[1527] Wenping Wang, Francis Husan, and Shein-Chung Chow. The impact of patient compliance on drug concentration profile in multiple doses. Statistics in Medicine, 15(6):659–669, March 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1996:SSE

[1528] Kung-Jong Lui. Sample size for the exact conditional test under inverse sampling. Statistics in Medicine, 15(6):671–678, March 30, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Capocaccia:1996:TCI

[1529] Riccardo Capocaccia. Trends in cancer incidence and mortality. Statistics in Medicine, 15(6):679–680, March 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mg

[1530] Anonymous. Masthead. Statistics in Medicine, 15(7-9):fmi, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacques:1996:P

[1531] Geoff Jacques, Roger Grimson, Leeka Kheifets, Lance Waller, and Dan Wartenberg. Preface. Statistics in Medicine, 15(7-9):681–682, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reynolds:1996:FCS

[1532] Peggy Reynolds, Daniel F. Smith, Enid Satariano, David O. Nelson, Lynn R. Goldman, and Raymond R. Neutra. The four county study of childhood cancer: clusters in context. Statistics in Medicine, 15(7-9): 683–697, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Maskarinec:1996:III

[1533] Gertraud Maskarinec. Investigating increased incidence of events in the Islands: a Hawaii Department of Health perspective. Statistics in Medicine, 15(7-9):699–705, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hjalmars:1996:CLS

[1534] Ulf Hjalmars, Martin Kulldorff, Göran Gustafsson, and Neville Nagarwalla. Childhood leukaemia in Sweden: using GIS and a spatial scan statistic for cluster detection. Statistics in Medicine, 15(7-9):707–715, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rushton:1996:ESA

[1535] Gerard Rushton and Panos Lolonis. Exploratory spatial analysis of birth defect rates in an urban population. Statistics in Medicine, 15(7-9): 717–726, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Le:1996:SCN

[1536] Nhu D. Le, A. John Petkau, and Rhonda Rosychuk. Surveillance of clustering near point sources. Statistics in Medicine, 15(7-9):727–740, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wojdyla:1996:CAC

[1537] Daniel Wojdyla, Leonor Poletto, Cristina Cuesta, Clara Badler, and Maria Elena Passamonti. Cluster analysis with constraints: its use with breast cancer mortality rates in Argentina. Statistics in Medicine, 15(7-9):741–746, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mcknight:1996:STC

[1538] Robert H. Mcknight, Richard J. Kryscio, Jonathan R. Mays, and George C. Rodgers, Jr. Spatial and temporal clustering of an occupational poisoning: the example of green tobacco sickness. Statistics in Medicine, 15(7-9):747–757, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Modan:1996:PRC

[1539] Baruch Modan, Zvia Blumstein, Osnat Luxenburg, Ilya Novikov, and Joshua Shemer. A potential risk of cancer in a central laboratory. Statistics in Medicine, 15(7-9):759–763, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waller:1996:SPD

[1540] Lance A. Waller. Statistical power and design of focused clustering studies. Statistics in Medicine, 15(7-9):765–782, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oden:1996:RPS

[1541] Neal Oden, Geoffrey Jacquez, and Roger Grimson. Realistic power simulations compare point- and area-based disease cluster tests. Statistics in Medicine, 15(7-9):783–806, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1996:EAS

[1542] Rina Chen. Exploratory analysis as a sequel to suspected increased rate of cancer in a small residential or workplace community. Statistics in Medicine, 15(7-9):807–816, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saferis:1996:CAT

[1543] Viktoras Saferis and Leonardas A. Vilkauskas. Cluster analysis by testing the statistical hypothesis of uniformity. Statistics in Medicine, 15(7-9): 817–821, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jorgensen:1996:LSE

[1544] Bent Jørgensen, Søren Lundbye-Christensen, Xue-Kun Song, and Li Sun. A longitudinal study of emergency room visits and air pollution for Prince George, British Columbia. Statistics in Medicine, 15(7-9):823–836, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Islam:1996:RGM

[1545] M. N. Islam, C. D. O’Shaughnessy, and B. Smith. A random graph model for the final-size distribution of household infections. Statistics in Medicine, 15(7-9):837–843, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagarwalla:1996:SSV

[1546] Neville Nagarwalla. A scan statistic with a variable window. Statistics in Medicine, 15(7-9):845–850, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grimson:1996:DCS

[1547] Roger C. Grimson and Neal Oden. Disease clusters in structured environments. Statistics in Medicine, 15(7-9):851–871, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacquez:1996:DCS

[1548] Geoffrey M. Jacquez. Disease cluster statistics for imprecise space-time locations. Statistics in Medicine, 15(7-9):873–885, April 15, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

WaldhOR:1996:SAC

[1549] Thomas Waldhör. The spatial autocorrelation coefficient Moran’s I under heteroscedasticity. Statistics in Medicine, 15(7-9):887–892, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Munasinghe:1996:LDC

[1550] Rajika L. Munasinghe and Robert D. Morris. Localization of disease clusters using regional measures of spatial autocorrelation. Statistics in Medicine, 15(7-9):893–905, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Downer:1996:ISI

[1551] Robert G. Downer. An introduction to smoothing incidence rates by penalized likelihood. Statistics in Medicine, 15(7-9):907–917, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schlattmann:1996:CAM

[1552] Peter Schlattmann, Ekkehart Dietz, and Dankmar Böhning. Covariate adjusted mixture models and disease mapping with the program dismapwin. Statistics in Medicine, 15(7-9):919–929, April 15, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schlattmann:1996:CPP

[1553] Peter Schlattmann. The computer package dismapwin. Statistics in Medicine, 15(7-9):931, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Linn:1996:SVE

[1554] S. Linn. Sorting variables for each case: a new algorithm to calculate injury severity score (ISS) using SPSS-PC. Statistics in Medicine, 15(7-9):933–937, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skelton:1996:QAS

[1555] Adam G. Skelton. Quadrat analysis software for the detection of spatial or temporal clustering. Statistics in Medicine, 15(7-9):939–941, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hall:1996:CSS

[1556] H. I. Hall, C. V. Lee, and W. E. Kaye. Cluster: a software system for epidemiologic cluster analysis. Statistics in Medicine, 15(7-9):943–950, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacquez:1996:SSC

[1557] Geoffrey M. Jacquez. Statistical software for the clustering of health events. Statistics in Medicine, 15(7-9):951–952, April 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mh

[1558] Anonymous. Masthead. Statistics in Medicine, 15(10):fmi, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feingold:1996:CTC

[1559] Marcia Feingold and Brenda Wilson Gillespie. Cross-over trials with censored data. Statistics in Medicine, 15(10):953–967, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:1996:RGD

[1560] Jørgen Hilden and Paul Glasziou. Regret graphs, diagnostic uncertainty and Youden’s index. Statistics in Medicine, 15(10):969–986, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andrew:1996:TAR

[1561] J. E. Andrew, P. Prescott, T. M. F. Smith, W. H. W. Inman, and K. Kubota. Testing for adverse reactions using prescription event monitoring. Statistics in Medicine, 15(10):987–1002, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:1996:MLL

[1562] M. P. Little, C. R. Muirhead, and C. A. Stiller. Modelling lymphocytic leukaemia incidence in England and Wales using generalizations of the two-mutation model of carcinogenesis of Moolgavkar, Venzon and Knudson. Statistics in Medicine, 15(10):1003–1022, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:1996:FDM

[1563] Xinhua Liu, Wei-Yann Tsai, and Yaakov Stern. A functional decline model for prevalent cohort data. Statistics in Medicine, 15(10):1023– 1032, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Minder:1996:RMP

[1564] C. E. Minder and T. Bednarski. A robust method for proportional hazards regression. Statistics in Medicine, 15(10):1033–1047, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1996:PCS

[1565] Seung-Yeoun Lee. Power calculation for a score test in the dependent censoring model. Statistics in Medicine, 15(10):1049–1058, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:1996:DUS

[1566] Michael A. Proschan. On the distribution of the unpaired t-statistic with paired data. Statistics in Medicine, 15(10):1059–1063, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelber:1996:MDR

[1567] Richard D. Gelber. Measuring disease: a review of disease-specific quality of life measurements scales. Statistics in Medicine, 15(10):1067–1068, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Julious:1996:LES

[1568] Steven A. Julious and Michael J. Campbell. Letter to the Editor: Sample sizes calculations for ordered categorical data, by J. Whitehead, Statistics in Medicine, 12, 2257–2272 (1993). Statistics in Medicine, 15(10):1065– 1066, May 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mi

[1569] Anonymous. Masthead. Statistics in Medicine, 15(11):fmi, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:1996:DCR

[1570] Mitchell H. Gail, Steven D. Mark, Raymond J. Carroll, Sylvan B. Green, and David Pee. On design considerations and randomization-based inference for community intervention trials. Statistics in Medicine, 15(11): 1069–1092, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carlin:1996:RBA

[1571] Bradley P. Carlin and Daniel J. Sargent. Robust Bayesian approaches for clinical trial monitoring. Statistics in Medicine, 15(11):1093–1106, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [1871].

Blair:1996:CFE

[1572] R. Clifford Blair, James F. Troendle, and Roy W. Beck. Control of familywise errors in multiple endpoint assessments via stepwise permutation tests. Statistics in Medicine, 15(11):1107–1121, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:1996:EDL

[1573] Emmanuel Lesaffre, Geert Molenberghs, and Lode Dewulf. Effect of dropouts in a longitudinal study: an application of a repeated ordinal model. Statistics in Medicine, 15(11):1123–1141, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Manor:1996:CSF

[1574] Orly Manor and Jeremy D. Kark. A comparative study of four methods for analysing repeated measures data. Statistics in Medicine, 15(11): 1143–1159, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Witte:1996:SSH

[1575] John S. Witte and Sander Greenland. Simulation study of hierarchical regression. Statistics in Medicine, 15(11):1161–1170, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Epstein:1996:BPM

[1576] Leonardo D. Epstein and Alvaro Muñoz. A bivariate parametric model for survival and intermediate event times. Statistics in Medicine, 15(11): 1171–1185, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1996:TSC

[1577] Eric W. Lee. Two sample comparison for large groups of correlated binary responses. Statistics in Medicine, 15(11):1187–1197, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1996:MAA

[1578] Warren D. Smith, Robert C. Dutton, and N. Ty Smith. A measure of association for assessing prediction accuracy that is a generalization of non-parametric ROC area. Statistics in Medicine, 15(11):1199–1215, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1996:GTA

[1579] James J. Chen. Global tests for analysis of multiple tumour data from animal carcinogenicity experiments. Statistics in Medicine, 15(11):1217– 1225, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

TenHave:1996:LEA

[1580] Thomas R. Ten Have, J. Richard Landis, and Susan Weaver. Letter to the Editor: Association models for periodontal disease progression: a comparison of methods for clustered binary data, by T. R. Ten Have, J. R. Landis and S. Weaver, Statistics in Medicine 14, 413–430 (1995). Statistics in Medicine, 15(11):1227–1229, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmgren:1996:ALD

[1581] Juni Palmgren. Analysis of longitudinal data. Statistics in Medicine, 15(11):1231–1232, June 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mj

[1582] Anonymous. Masthead. Statistics in Medicine, 15(12):fmi, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geller:1996:P

[1583] Nancy Geller, Laurence Freedman, Young Jack Lee, and Rebecca Dersimonian. Preface. Statistics in Medicine, 15(12):1233, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dersimonian:1996:MAD

[1584] Rebecca Dersimonian. Meta-analysis in the design and monitoring of clinical trials. Statistics in Medicine, 15(12):1237–1248, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1996:D

[1585] Freedman, Simon, Dersimonian, Pocock, Demets, Brown, Chalmers, Louis, Clemens, Collins, Feinstein, George, Olkin, Ellenberg, and Lee. Discussion. Statistics in Medicine, 15(12):1249–1252, June 30, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olkin:1996:MAC

[1586] Ingram Olkin. Meta-analysis: current issues in research synthesis. Statistics in Medicine, 15(12):1253–1257, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meier:1996:D

[1587] Meier, Olkin, Begg, George, Ellenberg, Collins, Freedman, Pocock, Chalmers, and Lee. Discussion. Statistics in Medicine, 15(12):1259–1262, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chalmers:1996:CCT

[1588] Thomas C. Chalmers and Joseph Lau. Changes in clinical trials mandated by the advent of meta-analysis. Statistics in Medicine, 15(12): 1263–1268, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Demets:1996:D

[1589] Demets, Temple, Chalmers, Begg, Collins, Olkin, Freedman, Ellenberg, Geller, Berendes, Feinstein, George, Meier, and Lee. Discussion. Statistics in Medicine, 15(12):1269–1272, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feinstein:1996:MAM

[1590] Alvan R. Feinstein. Meta-analysis and meta-analytic monitoring of clinical trials. Statistics in Medicine, 15(12):1273–1280, June 30, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:1996:D

[1591] Pocock, Feinstein, Brown, Louis, Begg, Meier, Demets, Olkin, and Lee. Discussion. Statistics in Medicine, 15(12):1281–1283, June 30, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:1996:REE

[1592] Stuart J. Pocock. The role of external evidence in data monitoring of a clinical trial. Statistics in Medicine, 15(12):1285–1293, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ellenberg:1996:D

[1593] Ellenberg, Pocock, Lee, Olkin, Collins, Meier, Louis, George, Simon, and Lee. Discussion. Statistics in Medicine, 15(12):1295–1297, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:1996:RMA

[1594] Colin B. Begg. The role of meta-analysis in monitoring clinical trials. Statistics in Medicine, 15(12):1299–1306, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Friedman:1996:D

[1595] Friedman, Begg, Feinstein, Chalmers, Meier, Freedman, Pocock, Ellenberg, Brown, Louis, Collins, Olkin, Dersimonian, and Lee. Discussion. Statistics in Medicine, 15(12):1307–1311, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geller:1996:RMA

[1596] Nancy L. Geller. Role of meta-analyses in design and monitoring of randomized clinical trials: summary of first day of the workshop. Statistics in Medicine, 15(12):1313–1315, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:1996:CD

[1597] Brown, Friedman, Temple, Collins, Simon, Feinstein, Chalmers, Begg, Pocock, Freedman, Meier, George, Demets, Ellenberg, Clemens, Berendes, and Lee. Concluding discussion. Statistics in Medicine, 15(12): 1317–1323, June 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mk

[1598] Anonymous. Masthead. Statistics in Medicine, 15(13):fmi, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailey:1996:ESA

[1599] B. J. R. Bailey and G. L. Briars. Estimating the surface area of the human body. Statistics in Medicine, 15(13):1325–1332, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Umbach:1996:TME

[1600] David M. Umbach and Allen J. Wilcox. A technique for measuring epidemiologically useful features of birthweight distributions. Statistics in Medicine, 15(13):1333–1348, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilton:1996:CVF

[1601] Sterling Hilton, Joanne Katz, and Scott Zeger. Classifying visual field data. Statistics in Medicine, 15(13):1349–1364, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pauler:1996:MME

[1602] Donna K. Pauler, Michael D. Escobar, John A. Sweeney, and Joel Greenhouse. Mixture models for eye-tracking data: a case study. Statistics in Medicine, 15(13):1365–1376, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brenner:1996:HIM

[1603] Hermann Brenner. How independent are multiple ‘independent’ diagnostic classifications? Statistics in Medicine, 15(13):1377–1386, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1948].

Sun:1996:NPT

[1604] Jianguo Sun. A non-parametric test for interval-censored failure time data with application to AIDS studies. Statistics in Medicine, 15(13): 1387–1395, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:1996:AMR

[1605] Wang-Shu Lu and Robert K. Tsutakawa. Analysis of mortality rates via marginal extended quasi-likelihood. Statistics in Medicine, 15(13): 1397–1407, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shoukri:1996:MLE

[1606] M. M. Shoukri and I. U. H. Mian. Maximum likelihood estimation of the kappa coefficient from bivariate logistic regression. Statistics in Medicine, 15(13):1409–1419, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

James:1996:IAB

[1607] Kenneth E. James, Daniel A. Bloch, Kelvin K. Lee, Helena C. Kraemer, and Richard K. Fuller. An index for assessing blindness in a multicentre clinical trial: disulfiram for alcohol cessation — a VA cooperative study. Statistics in Medicine, 15(13):1421–1434, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1996:MDC

[1608] B. Jones and A. N. Donev. Modelling and design of cross-over trials. Statistics in Medicine, 15(13):1435–1446, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chaudhary:1996:ECI

[1609] Mohammad A. Chaudhary and Sally C. Stearns. Estimating confidence intervals for cost-effectiveness ratios: an example from a randomized trial. Statistics in Medicine, 15(13):1447–1458, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2089].

Darby:1996:IE

[1610] Sarah C. Darby. An introduction to epidemiology. Statistics in Medicine, 15(13):1459, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louis:1996:SME

[1611] Thomas A. Louis. Statistical models in epidemiology. Statistics in Medicine, 15(13):1459–1461, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skovgaard:1996:NMR

[1612] L. T. Skovgaard. Nonlinear models for repeated measurement data. Statistics in Medicine, 15(13):1462–1463, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:1996:MAP

[1613] S. C. Lewis. Methodological approaches in pharmacoepidemiology — application to spontaneous reporting. Statistics in Medicine, 15(13): 1464, July 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Ml

[1614] Anonymous. Masthead. Statistics in Medicine, 15(14):fmi, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emerson:1996:SRP

[1615] John D. Emerson, David C. Hoaglin, and David C. Hoaglin. Simple robust procedures for combining risk differences in sets of 2 × 2 tables. Statistics in Medicine, 15(14):1465–1488, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rom:1996:TIP

[1616] Dror M. Rom and Eunhee Hwang. Testing for individual and population equivalence based on the proportion of similar responses. Statistics in Medicine, 15(14):1489–1505, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:1996:GOR

[1617] John Williamson and Kyungmann Kim. A global odds ratio regression model for bivariate ordered categorical data from ophthalmologic studies. Statistics in Medicine, 15(14):1507–1518, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oja:1996:MBE

[1618] Hannu Oja, Olli Pekka Alho, and Esa Läärä. Model-based estimation of the excess fraction (attributable fraction): day care and middle ear infection. Statistics in Medicine, 15(14):1519–1534, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siskind:1996:CIE

[1619] Victor Siskind. Confidence intervals for the excess risk in case-control studies. Statistics in Medicine, 15(14):1535–1544, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1996:APT

[1620] Chung-Cheng Hsieh and Shou-Jen Lan. Assessment of postpartum timedependent disease risk in case-control studies: an application for examining age-specific effect estimates. Statistics in Medicine, 15(14):1545–1556, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ebrahimi:1996:EMA

[1621] Nader Ebrahimi. The effects of misclassification of the actual cause of death in competing risks analysis. Statistics in Medicine, 15(14):1557– 1566, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gay:1996:ASS

[1622] Nigel J. Gay. Analysis of serological surveys using mixture models: application to a survey of parvovirus b19. Statistics in Medicine, 15(14): 1567–1573, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hanley:1996:UBM

[1623] James A. Hanley. The use of the ‘binormal’ model for parametric ROC analysis of quantitative diagnostic tests. Statistics in Medicine, 15(14): 1575–1585, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sandvik:1996:MDS

[1624] Leiv Sandvik, Jan Erikssen, Petter Mowinckel, and Einar Andreas Rødland. A method for determining the size of internal pilot studies. Statistics in Medicine, 15(14):1587–1590, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2429].

Lawson:1996:LEI

[1625] A. Lawson and N. Oden. Letters to the Editor including author’s reply: Adjusting Moran’s I for population density, by Neal Oden, Statistics in Medicine, 14, 17–26 (1995). Statistics in Medicine, 15(14):1591–1594, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1185, 2141].

Yu:1996:LEI

[1626] Kai Fun Yu. Letter to the Editor including authors’ reply: Estimation of a common odds ratio under binary cluster sampling, by C. Ahn and T. Odom-Maryon, Statistics in Medicine, 14, 1567–1576 (1995). Statistics in Medicine, 15(14):1595–1600, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [1354].

Deeks:1996:MAD

[1627] Jonathan J. Deeks. Meta-analysis, decision analysis and costeffectiveness analysis: methods for quantitative synthesis in medicine. Statistics in Medicine, 15(14):1601–1602, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1996:BCP

[1628] David R. Jones. Biometrie in der chemisch-pharmazeutischen industrie: testing principles in clinical and pre-clinical trails. Statistics in Medicine, 15(14):1602–1603, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pugh:1996:HNB

[1629] W. J. Pugh. Health and numbers — basic biostatistical methods. Statistics in Medicine, 15(14):1603–1604, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tjur:1996:BRN

[1630] Tue Tjur. Book review: New Cambridge Statistical Tables (2nd edn). Statistics in Medicine, 15(14):1604, July 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mm

[1631] Anonymous. Masthead. Statistics in Medicine, 15(15):fmi, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Piantadosi:1996:IDD

[1632] Steven Piantadosi and Guanghan Liu. Improved designs for dose escalation studies using pharmacokinetic measurements. Statistics in Medicine, 15(15):1605–1618, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chinchilli:1996:DAI

[1633] Vernon M. Chinchilli and James D. Esinhart. Design and analysis of intra-subject variability in cross-over experiments. Statistics in Medicine, 15(15):1619–1634, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laska:1996:EPS

[1634] Eugene M. Laska, Morris Meisner, Joseph A. Wanderling, and H. B. Kushner. Estimating population size when duplicates are present. Statistics in Medicine, 15(15):1635–1646, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aickin:1996:AMR

[1635] Mikel Aickin and Cheryl Ritenbaugh. Analysis of multivariate reliability structures and the induced bias in linear model estimation. Statistics in Medicine, 15(15):1647–1661, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faucett:1996:SMC

[1636] Cheryl L. Faucett and Duncan C. Thomas. Simultaneously modelling censored survival data and repeatedly measured covariates: a Gibbs sampling approach. Statistics in Medicine, 15(15):1663–1685, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1996:SMC

[1637] Xiao H. Zhou and Constantine A. Gatsonis. A simple method for comparing correlated ROC curves using incomplete data. Statistics in Medicine, 15(15):1687–1693, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lambert:1996:MIS

[1638] Philippe Lambert. Modelling irregularly sampled profiles of non-negative dog triglyceride responses under different distributional assumptions. Statistics in Medicine, 15(15):1695–1708, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eriksson:1996:PPC

[1639] Sven Eriksson. Planning pharmaceutical clinical trials. basic statistical principles. Statistics in Medicine, 15(15):1709–1710, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeBritoPinheiro:1996:SDA

[1640] Flávio De Brito Pinheiro. Statistical design and analysis in pharmaceutical science — validation process controls and stability statistics: Textbooks and monographs series 143. Statistics in Medicine, 15(15):1710– 1711, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VonRosen:1996:MMA

[1641] Dietrich Von Rosen. Methods of multivariate analysis. Statistics in Medicine, 15(15):1711–1712, August 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mn

[1642] Anonymous. Masthead. Statistics in Medicine, 15(16):fmi, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mcintosh:1996:PRE

[1643] Martin W. Mcintosh. The population risk as an explanatory variable in research synthesis of clinical trials. Statistics in Medicine, 15(16):1713– 1728, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunnett:1996:AUT

[1644] Charles W. Dunnett and Michael Gent. An alternative to the use of two-sided tests in clinical trials. Statistics in Medicine, 15(16):1729– 1738, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Patel:1996:SON

[1645] H. I. Patel. Some optimal and non-optimal two-stage designs using an α-spending function. Statistics in Medicine, 15(16):1739–1745, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boateng:1996:SAP

[1646] Francis Boateng, Allan Sampson, and Barry Schwab. Statistical analysis of possible bias of clinical judgements due to observing an on-therapy marker variable. Statistics in Medicine, 15(16):1747–1755, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arjas:1996:NPB

[1647] Elja Arjas and Liping Liu. Non-parametric Bayesian approach to hazard regression: a case study with a large number of missing covariate values. Statistics in Medicine, 15(16):1757–1770, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eide:1996:SPH

[1648] Geir Egil Eide, Ernst Omenaas, and Amund Gulsvik. The semiproportional hazards model revisited: practical reparametrizations. Statistics in Medicine, 15(16):1771–1777, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Becker:1996:CTH

[1649] Niels G. Becker and Ji-Sheng Cui. Comparison of trends in HIV infection for two risk categories. Statistics in Medicine, 15(16):1779–1791, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:1996:CSM

[1650] Ziding Feng, Dale Mclerran, and James Grizzle. A comparison of statistical methods for clustered data analysis with Gaussian error. Statistics in Medicine, 15(16):1793–1806, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Toledano:1996:ORM

[1651] Alicia Y. Toledano and Constantine Gatsonis. Ordinal regression methodology for ROC curves derived from correlated data. Statistics in Medicine, 15(16):1807–1826, August 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mo

[1652] Anonymous. Masthead. Statistics in Medicine, 15(17):fmi, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1996:P

[1653] S. J. Smith. Preface. Statistics in Medicine, 15(17):1827–1828, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1996:SC

[1654] S. J. Smith. Staff and committees. Statistics in Medicine, 15(17):1829– 1834, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snider:1996:IW

[1655] Dixie. E. Snider. Introduction and welcome. Statistics in Medicine, 15 (17):1835–1836, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Merrill:1996:UDE

[1656] D. W. Merrill, S. Selvin, E. R. Close, and H. H. Holmes. Use of density equalizing map projections (DEMP) in the analysis of childhood cancer in four California counties. Statistics in Medicine, 15(17):1837–1848, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sheppard:1996:DCE

[1657] Lianne Sheppard, Ross L. Prentice, and Mary Anne Rossing. Design considerations for estimation of exposure effects on disease risk, using aggregate data studies. Statistics in Medicine, 15(17):1849–1858, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kirby:1996:TCB

[1658] Russell S. Kirby. Toward congruence between theory and practice in small area analysis and local public health data. Statistics in Medicine, 15(17):1859–1866, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martuzzi:1996:EBE

[1659] Marco Martuzzi and Paul Elliott. Empirical Bayes estimation of small area prevalence of non-rare conditions. Statistics in Medicine, 15(17): 1867–1873, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1996:EBM

[1660] Xiao-Hua Zhou, B. P. Katz, E. Holleman, C. A. Melfi, and R. Dittus. An empirical Bayes method for studying variation in knee replacement rates. Statistics in Medicine, 15(17):1875–1884, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

SaCarvalho:1996:SPU

[1661] Marilia Sá Carvalho, Oswaldo Gonçalves Cruz, and Flávio Fonseca Nobre. Spatial partitioning using multivariate cluster analysis and a contiguity algorithm. Statistics in Medicine, 15(17):1885–1894, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:1996:PCS

[1662] Lawrence H. Cox. Protecting confidentiality in small population health and environmental statistics. Statistics in Medicine, 15(17):1895–1905, September 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mp

[1663] Anonymous. Masthead. Statistics in Medicine, 15(18):fmi, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rushton:1996:SRB

[1664] Gerard Rushton, Rajesh Krishnamurthy, Diane Krishnamurti, Panos Lolonis, and Hu Song. The spatial relationship between infant mortality and birth defect rates in a U.S. city. Statistics in Medicine, 15(18): 1907–1919, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittie:1996:CMD

[1665] Peggy S. Wittie, Wanzer Drane, and Tim E. Aldrich. Classification methods for denominators in small areas. Statistics in Medicine, 15(18): 1921–1926, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shaddick:1996:USM

[1666] Gavin Shaddick and Paul Elliott. Use of Stone’s method in studies of disease risk around point sources of environmental pollution. Statistics in Medicine, 15(18):1927–1934, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacquez:1996:NNT

[1667] Geoffrey M. Jacquez. A k nearest neighbour test for space–time interaction. Statistics in Medicine, 15(18):1935–1949, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hendricks:1996:PDG

[1668] Scott A. Hendricks, James T. Wassell, James W. Collins, and Suzanne L. Sedlak. Power determination for geographically clustered data using generalized estimating equations. Statistics in Medicine, 15(18):1951–1960, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Croner:1996:GIS

[1669] Charles M. Croner, Jonathan Sperling, and Frederick R. Broome. Geographic information systems (gis): new perspectives in understanding human health and environmental relationships. Statistics in Medicine, 15(18):1961–1977, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:PP

[1670] Anonymous. Poster presentations. Statistics in Medicine, 15(18):1979– 1981, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1996:CR

[1671] Ted Colton. Closing remarks. Statistics in Medicine, 15(18):1983–1984, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1996:CS

[1672] S. Jay Smith. Close of symposium. Statistics in Medicine, 15(18):1985– 1986, September 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mq

[1673] Anonymous. Masthead. Statistics in Medicine, 15(19):fmi, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mittlbock:1996:EVL

[1674] Martina Mittlböck and Michael Schemper. Explained variation for logistic regression. Statistics in Medicine, 15(19):1987–1997, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:1996:EVS

[1675] Michael Schemper and Janez Stare. Explained variation in survival analysis. Statistics in Medicine, 15(19):1999–2012, October 15, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:1996:EPV

[1676] Nancy R. Cook. Estimating predictive values for blood pressure measurements from multivariate regression models with covariates. Statistics in Medicine, 15(19):2013–2028, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weinberg:1996:SES

[1677] C. R. Weinberg, R. Skjærven, and A. J. Wilcox. Statistical evidence for shared transient causes of anatomically distinct birth defects. Statistics in Medicine, 15(19):2029–2036, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:1996:CGS

[1678] Ming Tan and Xiaoping Xiong. Continuous and group sequential conditional probability ratio tests for phase ii clinical trials. Statistics in Medicine, 15(19):2037–2051, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:1996:AUT

[1679] Hung-Mo Lin and Michael D. Hughes. Analysis of uncontrolled treatment changes in HIV clinical trials. Statistics in Medicine, 15(19):2053–2067, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Law:1996:SAR

[1680] Matthew G. Law and John M. Kaldor. Survival analyses of randomized clinical trials adjusted for patients who switch treatments. Statistics in Medicine, 15(19):2069–2076, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2002] and reply [2003].

Greenhouse:1996:AMS

[1681] Joel B. Greenhouse and Nancy Paul Silliman. Applications of a mixture survival model with covariates to the analysis of a depression prevention trial. Statistics in Medicine, 15(19):2077–2094, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Podgor:1996:ERT

[1682] Marvin J. Podgor, Joseph L. Gastwirth, and Cyrus R. Mehta. Efficiency robust tests of independence in contingency tables with ordered classifications. Statistics in Medicine, 15(19):2095–2105, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirst:1996:GMM

[1683] Kathryn Hirst, Gary O. Zerbe, and William W. Hay, Jr. A generalized Michaelis–Menten response surface. Statistics in Medicine, 15(19):2107– 2119, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wellek:1996:CBI

[1684] S. Wellek. Clinical biostatistics. an introduction to evidence based medicine. Statistics in Medicine, 15(19):2121–2123, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Best:1996:BDA

[1685] Nicky Best. Bayesian data analysis. Statistics in Medicine, 15(19):2123– 2124, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1996:BT

[1686] Michael D. Hughes. Bayesian theory. Statistics in Medicine, 15(19):2124– 2125, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molenberghs:1996:BRS

[1687] Geert Molenberghs. Book review: The statistical analysis of categorical data (3rd edn). Statistics in Medicine, 15(19):2125–2126, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:A

[1688] Anonymous. Announcement. Statistics in Medicine, 15(19):i, October 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mr

[1689] Anonymous. Masthead. Statistics in Medicine, 15(20):fmi, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesser:1996:GBB

[1690] Martin L. Lesser. Guidelines for budgeting biostatistics involvement in research projects. Statistics in Medicine, 15(20):2127–2133, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hallstrom:1996:LPW

[1691] Al Hallstrom, Maria Mori Brooks, and Monika Peckova. Logrank, play the winner, power and ethics. Statistics in Medicine, 15(20):2135–2142, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:1996:SBT

[1692] Robert G. Newcombe. Sequentially balanced three-squares cross-over designs. Statistics in Medicine, 15(20):2143–2147, October 30, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seillier-Moiseiwitsch:1996:PDL

[1693] Françoise Seillier-Moiseiwitsch. Predictive diagnostics for logistic models. Statistics in Medicine, 15(20):2149–2160, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mendoza-Blanco:1996:BIP

[1694] José R. Mendoza-Blanco, Xin M. Tu, and Satish Iyengar. Bayesian inference on prevalence using a missing-data approach with simulation-based techniques: applications to HIV screening. Statistics in Medicine, 15 (20):2161–2176, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carstensen:1996:RMI

[1695] Bendix Carstensen. Regression models for interval censored survival data: application to HIV infection in Danish homosexual men. Statistics in Medicine, 15(20):2177–2189, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1996:MCD

[1696] Janet Andersen, Els Goetghebeur, and Louise Ryan. Missing cause of death information in the analysis of survival data. Statistics in Medicine, 15(20):2191–2201, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faraggi:1996:SSC

[1697] David Faraggi and Richard Simon. A simulation study of cross-validation for selecting an optimal cutpoint in univariate survival analysis. Statistics in Medicine, 15(20):2203–2213, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edmunds:1996:TDC

[1698] W. J. Edmunds, G. F. Medley, and D. J. Nokes. The transmission dynamics and control of hepatitis B virus in The Gambia. Statistics in Medicine, 15(20):2215–2233, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Auranen:1996:SMT

[1699] Kari Auranen, Jukka Ranta, Aino K. Takala, and Elja Arjas. A statistical model of transmission of Hib bacteria in a family. Statistics in Medicine, 15(20):2235–2252, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saei:1996:TMM

[1700] A. Saei, J. Ward, and C. A. Mcgilchrist. Threshold models in a methadone programme evaluation. Statistics in Medicine, 15(20):2253– 2260, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauerbrei:1996:BRE

[1701] Willi Sauerbrei. Book review: Evaluating clinical evidence: a handbook for managers. Statistics in Medicine, 15(20):2261–2262, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:1996:BRC

[1702] Emery N. Brown. Biologic rhythms in clinical and laboratory medicine. Statistics in Medicine, 15(20):2262–2263, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wilson:1996:ICB

[1703] Susan R. Wilson. Introduction to computational biology: maps, sequences and genomes. Statistics in Medicine, 15(20):2264, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Winkel:1996:BRM

[1704] Per Winkel. Book review: Measurement in laboratory medicine: a primer on control and interpretation. Statistics in Medicine, 15(20):2265–2266, October 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schafer:1996:SBR

[1705] Helmut Schäfer. Statistical basis of reference values in laboratory medicine. Statistics in Medicine, 15(20):2266, October 30, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Ms

[1706] Anonymous. Masthead. Statistics in Medicine, 15(21):fmi, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foulkes:1996:P

[1707] Mary A. Foulkes, Dennis O. Dixon, Victor Degruttola, and Steven Self. Preface. Statistics in Medicine, 15(21):2269–2270, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Killen:1996:I

[1708] John Y. Killen, Jr. Introduction. Statistics in Medicine, 15(21):2271– 2272, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:1996:UOD

[1709] Mitchell H. Gail. Use of observational data, including surveillance studies, for evaluating AIDS therapies. Statistics in Medicine, 15(21):2273– 2288, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lavalley:1996:MEB

[1710] Michael P. Lavalley and Victor Degruttola. Models for empirical Bayes estimators of longitudinal CD4 counts. Statistics in Medicine, 15(21): 2289–2305, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hoover:1996:UMD

[1711] Donald R. Hoover, Yun Peng, Alfred J. Saah, Roger R. Detels, Roger S. Day, and John P. Phair. Using multiple decrement models to estimate risk and morbidity from specific AIDS illnesses. Statistics in Medicine, 15(21):2307–2321, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gange:1996:PMA

[1712] Stephen J. Gange, Lisa P. Jacobson, and Alvaro Muñoz. Parametric models to assess tracking of categorical high-risk sexual behaviour for HIV infection. Statistics in Medicine, 15(21):2323–2336, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shiboski:1996:D

[1713] Stephen Shiboski. Discussion. Statistics in Medicine, 15(21):2337–2340, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williams:1996:SMC

[1714] Paige L. Williams. Sequential monitoring of clinical trials with multiple survival endpoints. Statistics in Medicine, 15(21):2341–2357, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schoenfeld:1996:LTF

[1715] David A. Schoenfeld. Long-term follow-up in AIDS clinical trials. Statistics in Medicine, 15(21):2359–2366, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Follmann:1996:D

[1716] Dean Follmann. Discussion. Statistics in Medicine, 15(21):2367–2370, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1996:CCT

[1717] Jeffrey M. Albert. A control-conditional test for assessing HIV vaccine efficacy in small sample preclinical studies. Statistics in Medicine, 15 (21):2371–2378, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ewell:1996:CMC

[1718] Marian Ewell. Comparing methods for calculating confidence intervals for vaccine efficacy. Statistics in Medicine, 15(21):2379–2392, November 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mt

[1719] Anonymous. Masthead. Statistics in Medicine, 15(22):fmi, November 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rida:1996:AEH

[1720] Wasima N. Rida. Assessing the effect of HIV vaccination on infectiousness. Statistics in Medicine, 15(22):2393–2404, November 30, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Halloran:1996:DVE

[1721] M. Elizabeth Halloran. Discussion: vaccine effect on susceptibility. Statistics in Medicine, 15(22):2405–2412, November 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yao:1996:PWP

[1722] Q. Yao and L. J. Wei. Play the winner for Phase II/III clinical trials. Statistics in Medicine, 15(22):2413–2423, November 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elashoff:1996:HTS

[1723] Michael Elashoff and Stephen Lagakos. Hiv treatment strategies utilizing virologic and immunologic markers as criteria for changing treatments. Statistics in Medicine, 15(22):2425–2443, November 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Larntz:1996:DAI

[1724] Kinley Larntz, James D. Neaton, Deborah N. Wentworth, and Teresa Yurik. Data analysis issues for protocols with overlapping enrollment.

Statistics in Medicine, 15(22):2445–2453, November 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cornell:1996:D

[1725] Richard G. Cornell. Discussion. Statistics in Medicine, 15(22):2455– 2458, November 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Munoz:1996:MIA

[1726] Alvaro Muñoz and Jane Xu. Models for the incubation of AIDS and variations according to age and period. Statistics in Medicine, 15(22): 2459–2473, November 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kim:1996:NPI

[1727] Hyungjin Myra Kim and Stephen W. Lagakos. Non-parametric inference of a failure time distribution when the failure times are estimated. Statistics in Medicine, 15(22):2475–2490, November 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mu

[1728] Anonymous. Masthead. Statistics in Medicine, 15(23):fmi, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gleser:1996:MEN

[1729] Leon J. Gleser and Ingram Olkin. Models for estimating the number of unpublished studies. Statistics in Medicine, 15(23):2493–2507, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Davis:1996:NPM

[1730] Charles S. Davis. Non-parametric methods for comparing multiple treatment groups to a control group, based on incomplete non-decreasing repeated measurements. Statistics in Medicine, 15(23):2509–2521, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eastwood:1996:OSA

[1731] Brian J. Eastwood. Optimal sample allocation in clinical trials designed to investigate relative risks. Statistics in Medicine, 15(23):2523–2538, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kafadar:1996:SGD

[1732] Karen Kafadar. Smoothing geographical data, particularly rates of disease. Statistics in Medicine, 15(23):2539–2560, December 15, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ranta:1996:DOS

[1733] Jukka Ranta, Janne Pitkäniemi, Marjatta Karvonen, Esa Virtala, Jarmo Rusanen, Alfred Colpaert, Arvo Naukkarinen, and Jaakko Tuomilehto. Detection of overall space-time clustering in a non-uniformly distributed population. Statistics in Medicine, 15(23):2561–2572, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [2005, 2004].

TenHave:1996:PAC

[1734] Thomas R. Ten Have, J. Richard Landis, and Jonathan Hartzel. Population-averaged and cluster-specific models for clustered ordinal response data. Statistics in Medicine, 15(23):2573–2588, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heinzl:1996:AIB

[1735] Harald Heinzl, Alexandra Kaider, and Gerhard Zlabinger. Assessing interactions of binary time-dependent covariates with time in Cox proportional hazards regression models using cubic spline functions. Statistics in Medicine, 15(23):2589–2601, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barry:1996:DRB

[1736] Daniel Barry. Differential recall bias and spurious associations in case/ control studies. Statistics in Medicine, 15(23):2603–2616, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hassani:1996:USM

[1737] Habib Hassani and Alan Ebbutt. Use of a stochastic model for repeated binary assessment. Statistics in Medicine, 15(23):2617–2627, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1996:OS

[1738] Sander Greenland. Observational studies. Statistics in Medicine, 15(23): 2629–2632, December 15, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1996:Mv

[1739] Anonymous. Masthead. Statistics in Medicine, 15(24):fmi, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1996:P

[1740] Ted Colton, Tony Johnson, and David Machin. Preface. Statistics in Medicine, 15(24):2633, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vuorinen:1996:TSP

[1741] Jouni Vuorinen and Jari Turunen. A three-step procedure for assessing bioequivalence in the general mixed model framework. Statistics in Medicine, 15(24):2635–2655, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bonellie:1996:CDA

[1742] S. R. Bonellie and G. M. Raab. A comparison of different approaches for fitting centile curves to birthweight data. Statistics in Medicine, 15(24): 2657–2667, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mariotti:1996:SUE

[1743] Sergio Mariotti and Raffaella Cascioli. Sources of uncertainty in estimating HIV infection rates by back-calculation: an application to Italian data. Statistics in Medicine, 15(24):2669–2687, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1996:SSI

[1744] Stephen Senn. Some statistical issues in project prioritization in the pharmaceutical industry. Statistics in Medicine, 15(24):2689–2702, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1996:SDE

[1745] John Whitehead. Sequential designs for equivalence studies. Statistics in Medicine, 15(24):2703–2715, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beacon:1996:MLM

[1746] Heather J. Beacon and Simon G. Thompson. Multi-level models for repeated measurement data: application to quality of life data in clinical trials. Statistics in Medicine, 15(24):2717–2732, December 30, 1996. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Higgins:1996:BSE

[1747] Julian P. T. Higgins and Anne Whitehead. Borrowing strength from external trials in a meta-analysis. Statistics in Medicine, 15(24):2733– 2749, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berridge:1996:AMP

[1748] Damon M. Berridge. An application of a marked point process in preclinical medicine. Statistics in Medicine, 15(24):2751–2762, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Valsecchi:1996:ELT

[1749] M. G. Valsecchi, D. Silvestri, and P. Sasieni. Evaluation of long-term survival: use of diagnostics and robust estimators with Cox’s proportional hazards model. Statistics in Medicine, 15(24):2763–2780, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanDerTweel:1996:COS

[1750] Ingeborg Van Der Tweel, Rudolf Kaaks, and Paul A. H. Van Noord. Comparison of one-sample two-sided sequential t-tests for application in epidemiological studies. Statistics in Medicine, 15(24):2781–2795, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyse:1996:RBR

[1751] Marc Buyse and Pascal Piedbois. On the relationship between response to treatment and survival time. Statistics in Medicine, 15(24):2797–2812, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goetghebeur:1996:ANC

[1752] Els J. T. Goetghebeur and Stanley H. Shapiro. Analysing non-compliance in clinical trials: ethical imperative or mission impossible? Statistics in Medicine, 15(24):2813–2826, December 30, 1996. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schumacher:1997:I

[1753] Martin Schumacher, Werner Vach, and Maria Blettner. Introduction. Statistics in Medicine, 16(1):1–2, January 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reilly:1997:RBH

[1754] Marie Reilly and Margaret Pepe. The relationship between hot-deck multiple imputation and weighted likelihood. Statistics in Medicine, 16 (1):5–19, January 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:1997:NRMa

[1755] James M. Robins. Non-response models for the analysis of non-monotone non-ignorable missing data. Statistics in Medicine, 16(1):21–37, January 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:1997:NRMb

[1756] James M. Robins and Richard D. Gill. Non-response models for the analysis of non-monotone ignorable missing data. Statistics in Medicine, 16(1):39–56, January 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vach:1997:SIE

[1757] Werner Vach. Some issues in estimating the effect of prognostic factors from incomplete covariate data. Statistics in Medicine, 16(1):57–72, January 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:1997:PIR

[1758] Michael Schemper and Georg Heinze. Probability imputation revisited for prognostic factor studies. Statistics in Medicine, 16(1):73–80, January 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rotnitzky:1997:ASP

[1759] Andrea Rotnitzky and James Robins. Analysis of semi-parametric regression models with non-ignorable non-response. Statistics in Medicine, 16(1):81–102, January 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Breslow:1997:WLP

[1760] Norman E. Breslow and Richard Holubkov. Weighted likelihood, pseudolikelihood and maximum likelihood methods for logistic regression analysis of two-stage data. Statistics in Medicine, 16(1):103–116, January 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schill:1997:LAS

[1761] Walter Schill and Karsten Drescher. Logistic analysis of studies with twostage sampling: a comparison of four approaches. Statistics in Medicine, 16(2):117–132, January 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benichou:1997:AEE

[1762] Jacques Benichou, Celia Byrne, and Mitchell Gail. An approach to estimating exposure-specific rates of breast cancer from a two-stage casecontrol study within a cohort. Statistics in Medicine, 16(2):133–151, January 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whittemore:1997:MSS

[1763] Alice S. Whittemore and Jerry Halpern. Multi-stage sampling in genetic epidemiology. Statistics in Medicine, 16(2):153–167, January 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

KUChenhoff:1997:SRE

[1764] H. Küchenhoff and R. J. Carroll. Segmented regression with errors in predictors: semi-parametric and parametric methods. Statistics in Medicine, 16(2):169–188, January 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuha:1997:EDA

[1765] Jouni Kuha. Estimation by data augmentation in regression models with continuous and discrete covariates measured with error. Statistics in Medicine, 16(2):189–201, January 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:1997:SCM

[1766] Sylvia Richardson and Laurent Leblond. Some comments on misspecification of priors in Bayesian modelling of measurement error problems. Statistics in Medicine, 16(2):203–213, January 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1997:RFM

[1767] Niels Keiding, Per Kragh Andersen, and John P. Klein. The role of frailty models and accelerated failure time models in describing heterogeneity due to omitted covariates. Statistics in Medicine, 16(2):215–224, January 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmoor:1997:ECO

[1768] Claudia Schmoor and Martin Schumacher. Effects of covariate omission and categorization when analysing randomized trials with the Cox model. Statistics in Medicine, 16(3):225–237, February 15, 1997. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hogan:1997:MMJ

[1769] Joseph W. Hogan and Nan M. Laird. Mixture models for the joint distribution of repeated measures and event times. Statistics in Medicine, 16(3):239–257, February 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hogan:1997:MBA

[1770] Joseph W. Hogan and Nan M. Laird. Model-based approaches to analysing incomplete longitudinal and failure time data. Statistics in Medicine, 16(3):259–272, February 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Martinussen:1997:MWM

[1771] Torben Martinussen and Niels Keiding. The Manton–Woodbury model for longitudinal data with dropouts. Statistics in Medicine, 16(3):273– 283, February 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:1997:TCD

[1772] James M. Robins and Ya’Acov Ritov. Toward a curse of dimensionality appropriate (CODA) asymptotic theory for semi-parametric models. Statistics in Medicine, 16(3):285–319, February 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Broemeling:1997:BAR

[1773] Lyle D. Broemeling and Peyton Cook. A Bayesian analysis of regression models with continuous errors with application to longitudinal studies. Statistics in Medicine, 16(4):321–332, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2066].

Thompson:1997:CMC

[1774] Mary Lou Thompson and L. Paul Fatti. Construction of multivariate centile charts for longitudinal measurements. Statistics in Medicine, 16 (4):333–345, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heitjan:1997:MRS

[1775] Daniel F. Heitjan and Divakar Sharma. Modelling repeated-series longitudinal data. Statistics in Medicine, 16(4):347–355, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:1997:MMC

[1776] Bernard Rosner and Robert J. Glynn. Multivariate methods for clustered ordinal data with applications to survival analysis. Statistics in Medicine, 16(4):357–372, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rose:1997:ALF

[1777] M. Sarah Rose and Gordon H. Fick. Assessment of lack of fit in simple linear regression: an application to serologic response to treatment for syphilis. Statistics in Medicine, 16(4):373–384, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tibshirani:1997:LMV

[1778] Robert Tibshirani. The lasso method for variable selection in the Cox model. Statistics in Medicine, 16(4):385–395, February 28, 1997. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Warszawski:1997:EMP

[1779] Josiane Warszawski, Antoine Messiah, Joseph Lellouch, Laurence Meyer, and Jean-Claude Deville. Estimating means and percentages in a complex sampling survey: application to a French national survey on sexual behaviour (acsf). Statistics in Medicine, 16(4):397–423, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Capocaccia:1997:ECP

[1780] R. Capocaccia and R. De Angelis. Estimating the completeness of prevalence based on cancer registry data. Statistics in Medicine, 16(4):425– 440, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauck:1997:PAP

[1781] Walter W. Hauck, Frederic Y. Bois, Terry Hyslop, Lauren Gee, and Sharon Anderson. A parametric approach to population bioequivalence. Statistics in Medicine, 16(4):441–454, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Russek-Cohen:1997:ETW

[1782] Estelle Russek-Cohen and Richard M. Simon. Evaluating treatments when a gender by treatment interaction may exist. Statistics in Medicine, 16(4):455–464, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Betensky:1997:CPC

[1783] Rebecca A. Betensky. Conditional power calculations for early acceptance of H_{0} embedded in sequential tests. Statistics in Medicine, 16(4): 465–477, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sahai:1997:CFT

[1784] H. Sahai and A. Khurshid. Corrections: Formulae and tables for the determination of sample sizes and power in clinical trials for testing differences in proportions for the two-sample design. Statistics in Medicine, 16(4):479–480, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karrison:1997:CCI

[1785] T. Karrison. Correction: Confidence intervals for median survival times under a piecewise exponential model with proportional hazards covariate effects. Statistics in Medicine, 16(4):480, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Francis:1997:BRE

[1786] Brian Francis. Book review: The Elements of Graphing Data. William S. Cleveland, Hobart Press, Summit, New Jersey, 1994. No. of pages: 297. Price: £35 hardback. ISBN: 0-9634884-1-4. Statistics in Medicine, 16 (4):481, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parsons:1997:BRP

[1787] Van L. Parsons. Book review: Practical Methods for Design and Analysis of Complex Surveys. Risto Lehtonen and Erkki J. Pahkinen, Wiley, Chichester, 1995. No. of pages: ix + 337. Price: £34.95. ISBN: 0-471-93934-X. Statistics in Medicine, 16(4):482–483, February 28, 1997. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanDerHeijden:1997:BRS

[1788] Peter G. M. Van Der Heijden, Jos Dessens, and Wim Jansen. Book review: Statistical Models for Ordinal Variables. Advanced Quantitative Techniques in the Social Sciences 4. Clifford C. Clogg and Edward S. Shihadeh, Sage Publications, Thousand Oaks, 1994. No. of pages: xiii + 192. Price: £26.50. ISBN: 0-8039-3676-1. Statistics in Medicine, 16 (4):483–484, February 28, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1997:MIC

[1789] Philip J. Smith, Theodore J. Thompson, and John A. Jereb. A model for interval-censored tuberculosis outbreak data. Statistics in Medicine, 16(5):485–496, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sun:1997:RAI

[1790] Jianguo Sun. Regression analysis of interval-censored failure time data. Statistics in Medicine, 16(5):497–504, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vacek:1997:AEI

[1791] Pamela M. Vacek. Assessing the effect of intensity when exposure varies over time. Statistics in Medicine, 16(5):505–513, March 15, 1997. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1997:SSL

[1792] Sander Greenland. Second-stage least squares versus penalized quasilikelihood for fitting hierarchical models in epidemiologic analyses. Statistics in Medicine, 16(5):515–526, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pitard:1997:SMA

[1793] A. Pitard and J. F. Viel. Some methods to address collinearity among pollutants in epidemiological time series. Statistics in Medicine, 16(5): 527–544, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bull:1997:JBR

[1794] Shelley B. Bull, Celia M. T. Greenwood, and Walter W. Hauck. Jackknife bias reduction for polychotomous logistic regression. Statistics in Medicine, 16(5):545–560, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coste:1997:ODA

[1795] Joël Coste, Eric Walter, Daniel Wasserman, and Alain Venot. Optimal discriminant analysis for ordinal responses. Statistics in Medicine, 16(5): 561–569, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zackin:1997:ARV

[1796] Robert Zackin and L. J. Wei. Analysis of repeated virological measurements based on cell dilution assays. Statistics in Medicine, 16(5):571–582, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Akazawa:1997:PLT

[1797] Kouhei Akazawa, Tsuyoshi Nakamura, and Yuko Palesch. Power of logrank test and Cox regression model in clinical trials with heterogeneous samples. Statistics in Medicine, 16(5):583–597, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Axelsen:1997:BRS

[1798] Nils Axelsen. Book review: Scientific Deception: an Overview and Guide to the Literature of Misconduct and Fraud in Scientific Research. Lesley Grayson, The British Library, London, 1995. No. of pages: xii + 107. Price: £24.50. ISBN: 0-7123-0831-8. Statistics in Medicine, 16(5):599, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fitzmaurice:1997:BRP

[1799] Garrett Fitzmaurice. Book review: Practical Biostatistical Methods. Steve Selvin, Duxbury Press, Belmont, California, 1995. No. of pages: xv + 503. Price: £22.95. ISBN: 0-534-23802-5. Statistics in Medicine, 16 (5):600, March 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haber:1997:EPE

[1800] Michael Haber. Estimation of the population effectiveness of vaccination. Statistics in Medicine, 16(6):601–610, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

NgAndu:1997:ECS

[1801] Nicholas H. Ng’Andu. An empirical comparison of statistical tests for assessing the proportional hazards assumption of Cox’s model. Statistics in Medicine, 16(6):611–626, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Etzioni:1997:EAD

[1802] Ruth Etzioni and Yu Shen. Estimating asymptomatic duration in cancer: the AIDS connection. Statistics in Medicine, 16(6):627–644, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bacchetti:1997:IHR

[1803] Peter Bacchetti. Incidence of HIV-related deaths in the United States: seasonality and trend. Statistics in Medicine, 16(6):645–652, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachin:1997:GSM

[1804] John M. Lachin. Group sequential monitoring of distribution-free analyses of repeated measures. Statistics in Medicine, 16(6):653–668, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walsh:1997:LRB

[1805] Stephen J. Walsh. Limitations to the robustness of binormal ROC curves: effects of model misspecification and location of decision thresholds on bias, precision, size and power. Statistics in Medicine, 16(6):669–679, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brittain:1997:BSR

[1806] Erica Brittain, Jolie Palensky, Jason Blood, and Janet Wittes. Blinded subjective rankings as a method of assessing treatment effect: a large sample example from the systolic hypertension in the elderly program (SHEP). Statistics in Medicine, 16(6):681–693, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sposto:1997:PGL

[1807] Richard Sposto, Donald Stablein, and Shelly Carter-Campbell. A partially grouped logrank test. Statistics in Medicine, 16(6):695–704, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:1997:FPD

[1808] J. N. S. Matthews. A formula for the probability of discordant classification in method comparison studies. Statistics in Medicine, 16(6): 705–710, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:1997:LEC

[1809] Qin Liu. Letter to the Editor: Confidence limits based on the first occurrence of an event, by V. T. George and R. C. Elston, Statistics in Medicine, 12, 685–690 (1993). Statistics in Medicine, 16(6):711, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [727, 2126, 2127].

Elston:1997:AR

[1810] Robert Elston and Varghese George. Authors’ reply. Statistics in Medicine, 16(6):712, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1997:BRH

[1811] Anne Whitehead. Book review: The handbook of research synthesis. H. Cooper and L. V. Hedges (eds), Russell Sage Foundation, New York, 1994. No of Pages: xvi + 573. Price: $49.95. ISBN: 0-87154-226-9. Statistics in Medicine, 16(6):713–714, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1997:BRS

[1812] Per Kragh Andersen. Book review: Survival Analysis: a Practical Approach. M. K. B. Parmar and D. Machin, Wiley, Chichester, 1995. No. of pages: 255. Price: £29·95. ISBN: 0-471-93460-5. Statistics in Medicine, 16(6):714–715, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Men:1997:BRS

[1813] Tamara Men. Book review: Statistical Methods in Cancer Research. Vol. IV. Descriptive Epidemiology. J. Estève, E. Benhamou and L. Raymond, IARC Scientific Publications, Lyon, 1994. No. of pages: 302. Price: £25. ISBN: 92-832-21281. Statistics in Medicine, 16(6):715–716, March 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crichton:1997:MDC

[1814] Nicola J. Crichton, John P. Hinde, and Jonathan Marchini. Models for diagnosing chest pain: is CART helpful? Statistics in Medicine, 16(7): 717–727, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1997:CED

[1815] Wen-Chung Lee. Characterizing exposure-disease association in human populations using the Lorenz curve and Gini index. Statistics in Medicine, 16(7):729–739, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernadinelli:1997:DME

[1816] L. Bernadinelli, C. Pascutto, N. G. Best, and W. R. Gilks. Disease mapping with errors in covariates. Statistics in Medicine, 16(7):741–752, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biggerstaff:1997:IVE

[1817] B. J. Biggerstaff and R. L. Tweedie. Incorporating variability in estimates of heterogeneity in the random effects model in meta-analysis. Statistics in Medicine, 16(7):753–768, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joseph:1997:BMB

[1818] Lawrence Joseph, Roxane Du Berger, and Patrick Bélisle. Bayesian and mixed Bayesian/likelihood criteria for sample size determination. Statistics in Medicine, 16(7):769–781, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1997:CIL

[1819] Xiao-Hua Zhou and Sujuan Gao. Confidence intervals for the log-normal mean. Statistics in Medicine, 16(7):783–790, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:1997:CID

[1820] Michael P. Fay and Eric J. Feuer. Confidence intervals for directly standardized rates: a method based on the gamma distribution. Statistics in Medicine, 16(7):791–801, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:1997:TIA

[1821] Eileen B. Brown, Hari K. Iyer, and C. Ming Wang. Tolerance intervals for assessing individual bioequivalence. Statistics in Medicine, 16(7): 803–820, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vargha:1997:LEC

[1822] P. Vargha. Letter to the Editor: A critical discussion of intraclass correlation coefficients, by R. Müller and P. Büttner, Statistics in Medicine, 13, 2465–2476 (1995). Statistics in Medicine, 16(7):821–822, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1176].

Milne:1997:BRB

[1823] Robin K. Milne. Book review: Biostatistics: Perspectives in Health Care Research and Practice. B. L. Verma, G. D. Shukla and R. N. Srivastava (Eds), CBS Publishers and Distributors, Delhi, 1993. No. of pages: xii + 204. Price: 80 Rupees (paperback). ISBN: 81-239-0161-5. Statistics in Medicine, 16(7):825–826, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaye:1997:BRI

[1824] D. H. Kaye. Book review: Interpreting Evidence: Evaluating Forensic Science in the Courtroom. Bernard Robertson and G. A. Vignaux, Wiley, 1995. No. of pages: xxi + 240. Price: £24.95. ISBN: 0-471-96026-8. Statistics in Medicine, 16(7):826–828, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1997:BRH

[1825] E. A. Thompson. Book review: Human Identification: The Use of DNA Markers. Bruce S. Weir (ed.), Kluwer Academic Publishers, 1995. No. of pages: 218. Price: $76. ISBN: 0-7923-3520-1. Statistics in Medicine, 16 (7):828–829, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

BOHning:1997:BRE

[1826] Dankmar Böhning. Book review: Epidemiologische Methoden. Lothar Kreienbrock and Siegfried Schach, Gustav Fischer Verlag, Stuttgart / Jena, 1995. No of pages: 256. Price: DM 64. ISBN: 3-437-20511-0. Statistics in Medicine, 16(7):829–830, April 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anbar:1997:P

[1827] Dan Anbar and Stephen Lagakos. Preface. Statistics in Medicine, 16 (8):831, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wei:1997:OSM

[1828] L. J. Wei and David V. Glidden. An overview of statistical methods for multiple failure time data in clinical trials. Statistics in Medicine, 16(8): 833–839, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:1997:DPW

[1829] R. J. Cook, J. F. Lawless, David Oakes, Dennis O. Dixon, R. J. Spiegel, Katherine H. Lipschutz, Steven M. Snapinn, Cyrus R. Mehta, and J. P. Siegel. Discussion of paper by Wei and Glidden. Statistics in Medicine, 16 (8):841–851, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Turnbull:1997:RMR

[1830] Bruce W. Turnbull, Wenxin Jiang, and Larry C. Clark. Regression models for recurrent event data: parametric random effects models with measurement error. Statistics in Medicine, 16(8):853–864, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hughes:1997:PCC

[1831] Michael D. Hughes. Power considerations for clinical trials using multivariate time-to-event data. Statistics in Medicine, 16(8):865–882, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Catalano:1997:BMC

[1832] Paul J. Catalano. Bivariate modelling of clustered continuous and ordered categorical outcomes. Statistics in Medicine, 16(8):883–900, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:1997:NPI

[1833] D. Y. Lin. Non-parametric inference for cumulative incidence functions in competing risks studies. Statistics in Medicine, 16(8):901–910, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:1997:MAR

[1834] Richard J. Cook and Jerald F. Lawless. Marginal analysis of recurrent events and a terminating event. Statistics in Medicine, 16(8):911–924, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:1997:UWL

[1835] Qian H. Li and Stephen W. Lagakos. Use of the Wei–Lin–Weissfeld method for the analysis of a recurring and a terminating event. Statistics in Medicine, 16(8):925–940, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barai:1997:MSM

[1836] Usha Barai and Nick Teoh. Multiple statistics for multiple events, with application to repeated infections in the growth factor studies. Statistics in Medicine, 16(8):941–949, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finkelstein:1997:AMF

[1837] Dianne M. Finkelstein, David A. Schoenfeld, and Eva Stamenovic. Analysis of multiple failure time data from an AIDS clinical trial. Statistics in Medicine, 16(8):951–961, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2164].

Rennolls:1997:BRC

[1838] Keith Rennolls. Book review: Case Studies in Biometry. Nicholas Lange, Louise Ryan, Lynne Billard, David Brillinger, Loveday Conquest and Joel Greenhouse (ed), Wiley Chichester 1994. No. of pages: 496 with a data disc. Price: £24.95. ISBN 0-471-58885-7(cloth), 0-471-58925-X (paper). Statistics in Medicine, 16(8):963–964, April 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hosmer:1997:CGF

[1839] D. W. Hosmer, T. Hosmer, S. Le Cessie, and S. Lemeshow. A comparison of goodness-of-fit tests for the logistic regression model. Statistics in Medicine, 16(9):965–980, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brenner:1997:VSS

[1840] Hermann Brenner and Olaf Gefeller. Variation of sensitivity, specificity, likelihood ratios and predictive values with disease prevalence. Statistics in Medicine, 16(9):981–991, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2128].

Neuhauser:1997:IOR

[1841] Markus Neuhäuser and Heiko Becher. Improved odds ratio estimation by post hoc stratification of case-control data. Statistics in Medicine, 16(9): 993–1004, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Austin:1997:ERR

[1842] Harland Austin, Laura L. Perkins, and David O. Martin. Estimating a relative risk across sparse case-control and follow-up studies: a method for meta-analysis. Statistics in Medicine, 16(9):1005–1015, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cuzick:1997:ANC

[1843] Jack Cuzick, Robert Edwards, and Nereo Segnan. Adjusting for noncompliance and contamination in randomized clinical trials. Statistics in Medicine, 16(9):1017–1029, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chow:1997:SMT

[1844] Shein-Chung Chow and Jun Shao. Statistical methods for two-sequence three-period cross-over designs with incomplete data. Statistics in Medicine, 16(9):1031–1039, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bull:1997:TBS

[1845] Kate Bull and David J. Spiegelhalter. Tutorial in biostatistics survival analysis in observational studies. Statistics in Medicine, 16(9):1041–1074, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Valsecchi:1997:BRI

[1846] Maria Grazia Valsecchi. Book review: Introduction to Biostatistics. A Guide to Design, Analysis, and Discovery. R. N. Forthofer and E. S. Lee, Academic Press, San Diego, 1995. No. of pages: 567. Price: $59.95. ISBN: 0-12-262270-7. Statistics in Medicine, 16(9):1075–1076, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lundbye-Christensen:1997:BRL

[1847] SØRen Lundbye-Christensen. Book review: Linear Statistical Models. James H. Stapleton, Wiley, New York, 1995. No. of pages: 449. Price: £50. ISBN 0-471-57150-4. Statistics in Medicine, 16(9):1076–1077, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klein:1997:BRR

[1848] John P. Klein. Book review: Recent Advances in Clinical Trial Design and Analysis. Peter F. Thall (ed), Kluwer Academic Publishers, Massachusetts, 1995. No. of pages: 258. Price: $175. ISBN 0-7923-3235-0. Statistics in Medicine, 16(9):1077–1078, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mcclean:1997:BRS

[1849] Sally Mcclean. Book review: Statistical Analysis of Behavioural Data — an Approach Based on Time-Structured Models. Patsy Haccou and Evert Meelis, Oxford University Press, 1994. No. of pages: xvii + 396. Price: £21.50. ISBN: 0-19-854850-8. Statistics in Medicine, 16(9):1079, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Richardson:1997:BRE

[1850] Sylvia Richardson. Book review: Elsevier’s Dictionary of Biometry. D. Rasch, M. L. Tiku and D. Sumpf (eds), Elsevier Science, Amsterdam, 1994. No. of pages: 887. Price: $225.75. ISBN: 0-444-81495-7. Statistics in Medicine, 16(9):1080, May 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

May:1997:VPT

[1851] Warren L. May and William D. Johnson. The validity and power of tests for equality of two correlated proportions. Statistics in Medicine, 16(10): 1081–1096, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2130].

Donner:1997:HAI

[1852] Allan Donner and Michael Eliasziw. A hierarchical approach to inferences concerning interobserver agreement for multinomial data. Statistics in Medicine, 16(10):1097–1106, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pichlmeier:1997:CAA

[1853] Uwe Pichlmeier and Olaf Gefeller. Conceptual aspects of attributable risk with recurrent disease events. Statistics in Medicine, 16(10):1107–1120, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Erkanli:1997:BIT

[1854] Alaattin Erkanli, Refik Soyer, and Dalene Stangl. Bayesian inference in two-phase prevalence studies. Statistics in Medicine, 16(10):1121–1133, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wakefield:1997:PAI

[1855] Jon Wakefield and Stephen Walker. A population approach to initial dose selection. Statistics in Medicine, 16(10):1135–1149, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaiser:1997:CNP

[1856] Mark S. Kaiser and David Siev. Comparison of non parallel immunoassay curves resulting from mixtures of competing antigens. Statistics in Medicine, 16(10):1151–1166, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shao:1997:PBR

[1857] Jun Shao and Ling Chen. Prediction bounds for random shelf-lives. Statistics in Medicine, 16(10):1167–1173, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bjorling:1997:RBR

[1858] Lisa E. Bjorling and James S. Hodges. Rule-based ranking schemes for antiretroviral trials. Statistics in Medicine, 16(10):1175–1191, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fidler:1997:BRS

[1859] Václav Fidler. Book review: Statistics in Clinical Practice. David Coggon, BMJ Publishing Group, 1995. No of pages: vii + 116. Price: £12.95 (paperback). ISBN: 0-7279-0907-X. Statistics in Medicine, 16(10):1193, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Regan:1997:BRE

[1860] Myra O’ Regan. Book review: Essential Statistics for Medical Practice — a Case-Study Approach. D. G. Rees, Chapman and Hall, 1994. No of pages: 223. Price: £12.99. ISBN: 0-412-59930-9. Statistics in Medicine, 16(10):1194, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Salome:1997:BRS

[1861] D. Salomé. Book review: Statistical Methods for Non-Precise Data. Reinhard Viertl, CRC Press, 1996. No. of pages: 191. Price: $49.95 (£39). ISBN: 0-8493-8242-4. Statistics in Medicine, 16(10):1194–1195, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holstein-Rathlou:1997:BRC

[1862] Niels-Henrik Holstein-Rathlou. Book review: Computer Analysis of Cardiovascular Signals. M. Di Rienzo, G. Mancia, G. Parati, A. Pedotti, and A. Zanchetti (eds.), IOS Press, 1995. No. of pages: 350. Price: £74. ISBN: 90-5199-158-4. Statistics in Medicine, 16(10):1195–1196, May 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fairclough:1997:SMS

[1863] Diane L. Fairclough. Summary measures and statistics for comparison of quality of life in a clinical trial of cancer therapy. Statistics in Medicine, 16(11):1197–1209, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diehr:1997:ESP

[1864] Paula Diehr, Bruce M. Psaty, and Donald L. Patrick. Effect size and power for clinical trials that measure years of healthy life. Statistics in Medicine, 16(11):1211–1223, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:1997:TTD

[1865] Weichung Joseph Shih and Hui Quan. Testing for treatment differences with dropouts present in clinical trials — a composite approach. Statistics in Medicine, 16(11):1225–1239, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2143].

Roe:1997:CPP

[1866] Denise J. Roe, Edward F. Vonesh, Russell D. Wolfinger, Florence Mesnil, and Alain Mallet. Comparison of population pharmacokinetic modeling methods using simulated data: results from the population modeling workgroup. Statistics in Medicine, 16(11):1241–1262, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leisenring:1997:MRM

[1867] Wendy Leisenring, Margaret Sullivan Pepe, and Gary Longton. A marginal regression modelling framework for evaluating medical diagnostic tests. Statistics in Medicine, 16(11):1263–1281, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacqmin-Gadda:1997:TGC

[1868] Hélène Jacqmin-Gadda, Daniel Commenges, Chakib Nejjari, and Jean-François Dartigues. Tests of geographical correlation with adjustment for explanatory variables: an application to dyspnoea in the elderly. Statistics in Medicine, 16(11):1283–1297, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1997:CGL

[1869] Philip J. Smith, Theodore J. Thompson, Michael M. Engelgau, and William H. Herman. Correction: A Generalized Linear Model for Analysing Receiver Operating Characteristic Curves, by Philip J. Smith, Theodore J. Thompson, Michael M. Engelgau and William H. Herman, Statistics in Medicine, 15, 323–333 (1996). Statistics in Medicine, 16 (11):1299, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1495].

Hsieh:1997:CSS

[1870] F. Y. Hsieh. Correction: Sample Size Formulae for Intervention Studies with the Cluster as Unit of Randomization, by F. Y. Hsieh, Statistics in Medicine, 7, 1195–1201 (1988). Statistics in Medicine, 16(11):1300, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1].

Carlin:1997:CRB

[1871] Bradley P. Carlin and Daniel J. Sargent. Correction: Robust Bayesian Approaches for Clinical Trial Monitoring, by Bradley P. Carlin and Daniel J. Sargent, Statistics in Medicine, 15, 1093–1106 (1996). Statistics in Medicine, 16(11):1300, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1571].

Zhou:1997:LEC

[1872] Hong Zhou and Patrick S. Romano. Letter to the Editor: Confidence Interval Estimates of an Index of Quality Performance Based on Logistic Regression Models, by David W. Hosmer and Stanley Lemeshow, Statistics in Medicine, 14, 2161–2172 (1995). Statistics in Medicine, 16(11): 1301–1303, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1413].

Hosmer:1997:AR

[1873] David W. Hosmer and Stanley Lemeshow. Authors’ reply. Statistics in Medicine, 16(11):1303, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1997:LET

[1874] Stephen Senn. Letter to the Editor: Testing for Individual and Population Equivalence Based on the Proportion of Similar Responses, by D. M. Rom and E. Hwang, Statistics in Medicine, 15, 1489–1505 (1996). Statistics in Medicine, 16(11):1303–1305, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rom:1997:AR

[1875] Dror Rom and Eunhee Hwang. Authors’ reply. Statistics in Medicine, 16 (11):1306, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:1997:BRB

[1876] M. A. A. Cox. Book review: Basic Biostatistics: Concepts for the Health Sciences: The Almost No Math Stats Book. Frank D. Bell, W. C. Brown Publishers, Oxford, England, 1995. No. of pages: xiv + 176. Price: £23.95. ISBN: 0-697-14979-X. Statistics in Medicine, 16(11):1307–1308, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rushton:1997:BRI

[1877] Lesley Rushton. Book review: Introductory Biological Statistics. Raymond E. Hampton, W. C. Brown Publishers, 1994. No. of pages: 233. Price: £17.99. ISBN: 0-697-20209-7. Statistics in Medicine, 16(11):1308– 1309, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gottschau:1997:BRM

[1878] Adam Gottschau. Book review: Multilevel Statistical Models (2nd edn). Harvey Goldstein, Edward Arnold, London, 1995. No. of pages: 178. Price: £29.99. ISBN: 0-340-59529-9. Statistics in Medicine, 16(11):1309– 1310, June 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hashemi:1997:BAS

[1879] Lobat Hashemi, Balgobin Nandram, and Robert Goldberg. Bayesian analysis for a single 2×2 table. Statistics in Medicine, 16(12):1311–1328, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2248].

Wallenstein:1997:NIA

[1880] Sylvan Wallenstein. A non-iterative accurate asymptotic confidence interval for the difference between two proportions. Statistics in Medicine, 16(12):1329–1336, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mcshane:1997:CSP

[1881] Lisa Meier Mcshane, Kristen L. Meier, and Eric M. Wassermann. A comparison of spatial prediction techniques for an exploratory analysis of human cortical motor representations. Statistics in Medicine, 16(12): 1337–1355, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wang:1997:USS

[1882] Yuedong Wang, Grace Wahba, Chong Gu, Ronald Klein, and Barbara Klein. Using smoothing spline anova to examine the relation of risk factors to the incidence and progression of diabetic retinopathy. Statistics in Medicine, 16(12):1357–1376, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1997:RUD

[1883] A. J. Lee, Lovina Mcmurchy, and A. J. Scott. Re-using data from casecontrol studies. Statistics in Medicine, 16(12):1377–1389, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hadgu:1997:BED

[1884] Alula Hadgu. Bias in the evaluation of DNA-amplification tests for detecting Chlamydia trachomatis. Statistics in Medicine, 16(12):1391–1399, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2190, 2145].

Chen:1997:ESC

[1885] Rina Chen, Roger R. Connelly, and Nathan Mantel. The efficiency of the sets and the cuscore techniques under biased baseline rates. Statistics in Medicine, 16(12):1401–1411, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fayers:1997:TBB

[1886] Peter M. Fayers, Deborah Ashby, and Mahesh K. B. Parmar. Tutorial in biostatistics: Bayesian data monitoring in clinical trials. Statistics in Medicine, 16(12):1413–1430, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeLeeuw:1997:BRR

[1887] Jan De Leeuw. Book review: Rasch Models. Foundations, Recent Developments, and Applications. Gerhard H. Fischer and Ivo W. Molenaar (eds). Springer-Verlag, Berlin, 1995. No. of pages: 436. Price: DM 88. ISBN: 0-387-94499-0. Statistics in Medicine, 16(12):1431–1433, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holm:1997:BRH

[1888] Sture Holm. Book review: Handbook of Statistical Modeling for the Social and Behavioral Sciences. G. Arminger, C. C. Clogg and M. E. Sobel (eds), Plenum Press, New York, 1995. No. of pages: 592. Price: $59.50. ISBN: 0-306-44805-X. Statistics in Medicine, 16(12):1433–1434, June 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zahl:1997:LTS

[1889] Per-Henrik Zahl and Steinar Tretli. Long-term survival of breast cancer in Norway by age and clinical stage. Statistics in Medicine, 16(13):1435– 1449, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yarnold:1997:PHM

[1890] Paul R. Yarnold, Robert C. Soltysik, and Charles L. Bennett. Predicting in-hospital mortality of patients with AIDS-related Pneumocystis carinii pneumonia: an example of hierarchically optimal classification tree analysis. Statistics in Medicine, 16(13):1451–1463, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mehrotra:1997:NIR

[1891] Devan V. Mehrotra. Non-iterative robust estimators of variance components in within-subject designs. Statistics in Medicine, 16(13):1465–1479, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunstan:1997:DRF

[1892] F. D. J. Dunstan, J. C. Gray, A. B. J. Nix, and T. Reynolds. Detection rates and false positive rates for Down’s syndrome screening: how precisely can they be estimated and what factors influence their value? Statistics in Medicine, 16(13):1481–1495, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mian:1997:SAI

[1893] I. U. H. Mian and M. M. Shoukri. Statistical analysis of intraclass correlations from multiple samples with applications to arterial blood pressure data. Statistics in Medicine, 16(13):1497–1514, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:1997:EPT

[1894] D. Y. Lin, T. R. Fleming, and V. De Gruttola. Estimating the proportion of treatment effect explained by a surrogate marker. Statistics in Medicine, 16(13):1515–1527, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2331].

Obuchowski:1997:SSD

[1895] Nancy A. Obuchowski and Donna Katzman Mcclish. Sample size determination for diagnostic accuracy studies involving binormal ROC curve indices. Statistics in Medicine, 16(13):1529–1542, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kligiene:1997:BRL

[1896] Nerute Kligiene. Book review: Learning Statistics Through Playing Cards. Thomas R. Knapp, Sage Publications, Thousand Oaks, 1996. No. of pages: xii + 99. Price: £13.95 (paperback). ISBN: 0-7619-0109-4. Statistics in Medicine, 16(13):1543–1544, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olsen:1997:BRM

[1897] J. H. Olsen. Book review: Mortality from Smoking in Developed Countries 1950–2000. Richard Peto, Alan D. Lopez, Jillian Boreham, Michael Thun and Clark Heath Jr, Oxford University Press, 1994. No. of pages:

553. Price: £35. ISBN: 0-19-262619-1. Statistics in Medicine, 16(13): 1544–1545, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guttorp:1997:BRM

[1898] Peter Guttorp. Book review: Models for Infectious Human Diseases: Their Structure and Relation to Data. V. Isham and G. Medley (eds) Publications of the Newton Institute, vol. 6, Cambridge University Press, 1996. No. of pages: xxiii + 490. Price: $59.95. ISBN: 0-521-45339-9. Statistics in Medicine, 16(13):1545–1546, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:1997:BRT

[1899] Jørgen Hilden. Book review: Tree Models of Similarity and Association. Sage University Paper series on Quantitative Applications in the Social Sciences, No. 07-112. James E. Corter, Sage Publications, Thousand Oaks, CA, 1996. No. of pages: vi + 65. Price: £7.95. ISBN: 0-8039-5707-6. Statistics in Medicine, 16(13):1546–1547, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rehm:1997:BRC

[1900] Jürgen Rehm. Book review: Causal Analysis with Panel Data. Steven E. Finkel, Sage Publications, Thousand Oaks, California, 1995. No. of pages: 98. Price: £7.95. ISBN: 0-8039-3896-9. Statistics in Medicine, 16(13):1547–1548, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abrams:1997:BRM

[1901] Keith R. Abrams. Book review: Modern Applied Statistics with S-PLUS. W. N. Venables and B. D. Ripley, Springer-Verlag, New York, 1994. No. of pages: xiii + 462 + 3.5 inch diskette. Price: £24.50. ISBN: 0-387-94350-1. Statistics in Medicine, 16(13):1548–1550, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Commenges:1997:BRA

[1902] Daniel Commenges and Alioum Ahmadou. Book review: Analyse Statistique des Données de Survie. 2nd edn. Catherine Hill, Catherine Com-Nougué, Andrew Kramar, Thierry Moreau, John O’Quigley, Rachid Senoussi and Claude Chastang, INSERM Médecine–Sciences Flammarion, Paris, France, 1996. No. of pages: 190. Price: 222 FF. ISBN: 2-257-12310-7. Statistics in Medicine, 16(13):1550, July 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Couper:1997:MPC

[1903] David Couper and Margaret Sullivan Pepe. Modelling prevalence of a condition: chronic graft-versus-host disease after bone marrow transplantation. Statistics in Medicine, 16(14):1551–1571, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zahl:1997:FME

[1904] Per-Henrik Zahl. Frailty modelling for the excess hazard. Statistics in Medicine, 16(14):1573–1585, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsuyama:1997:MMB

[1905] Yutaka Matsuyama and Yasuo Ohashi. Mixed models for bivariate response repeated measures data using Gibbs sampling. Statistics in Medicine, 16(14):1587–1601, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Segal:1997:MSD

[1906] Mark R. Segal, Suzanne Wight, John P. Hanrahan, and Ira B. Tager. Maternal smoking during pregnancy and birth outcomes with weight gain adjustments via varying-coefficient models. Statistics in Medicine, 16(14):1603–1616, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Langer:1997:DDI

[1907] Stefan F. J. Langer. Data-dependent interval partition of naturally ordered individuals by complete cluster analysis in epidemiological and cardiac data processing. Statistics in Medicine, 16(14):1617–1628, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gamel:1997:CPN

[1908] John W. Gamel and Robert L. Vogel. Comparison of parametric and nonparametric survival methods using simulated clinical data. Statistics in Medicine, 16(14):1629–1643, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:1997:TQI

[1909] Guohua Pan and Douglas A. Wolfe. Test for qualitative interaction of clinical significance. Statistics in Medicine, 16(14):1645–1652, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeKlerk:1997:FAD

[1910] Erik De Klerk, Sjef Van Der Linden, Desiree Van Der Heijde, and John Urquhart. Facilitated analysis of data on drug regimen compliance. Statistics in Medicine, 16(14):1653–1664, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gilks:1997:BRB

[1911] W. R. Gilks. Book review: Bayesian Biostatistics. Donald A. Berry and Dalene K. Stangl, Marcel Dekker, New York, 1996. No. of pages:

696. Price: $165.00 ISBN: 0-8247-9334-X. Statistics in Medicine, 16(14): 1665–1666, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tomba:1997:BRE

[1912] Gianpaolo Scalia Tomba. Book review: Epidemic Models: Their Structure and Relation to Data. Denis Mollison (ed), Cambridge University Press, U.K., 1995. No. of pages: xviii + 424. Price: $49.95. ISBN: 0-521-47536-8. Statistics in Medicine, 16(14):1666, July 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carey:1997:UHI

[1913] Vincent J. Carey. Using hypertext and the Internet for structure and management of observational studies. Statistics in Medicine, 16(15): 1667–1682, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauck:1997:TPS

[1914] Walter W. Hauck, Linda J. McKee, and Barbara J. Turner. Two-part survival models applied to administrative data for determining rate of and predictors for maternal-child transmission of HIV. Statistics in Medicine, 16(15):1683–1694, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elswick:1997:IOR

[1915] R. K. Elswick, Jr., Pamela F. Schwartz, and Josephine A. Welsh. Interpretation of the odds ratio from logistic regression after a transformation of the covariate vector. Statistics in Medicine, 16(15):1695–1703, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Voit:1997:DLL

[1916] Eberhard O. Voit and Rebecca G. Knapp. Derivation of the linear-logistic model and Cox’s proportional hazard model from a canonical system description. Statistics in Medicine, 16(15):1705–1729, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Umbach:1997:DAC

[1917] David M. Umbach and Clarice R. Weinberg. Designing and analysing case-control studies to exploit independence of genotype and exposure. Statistics in Medicine, 16(15):1731–1743, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Clark:1997:RAM

[1918] I. Clark, R. Biscay, and J. C. Jiménez. Reparametrization of autoregressive models with coefficients depending on covariables: application to EEG spectrum maturation. Statistics in Medicine, 16(15):1745–1752, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Klar:1997:MMC

[1919] Neil Klar and Allan Donner. The merits of matching in community intervention trials: a cautionary tale. Statistics in Medicine, 16(15):1753– 1764, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2250].

Zeng:1997:THI

[1920] Qi Zeng and Marie Davidian. Testing homogeneity of intra-run variance parameters in immunoassay. Statistics in Medicine, 16(15):1765–1776, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1997:EET

[1921] Kung-Jong Lui. Exact equivalence test for risk ratio and its sample size determination under inverse sampling. Statistics in Medicine, 16 (15):1777–1786, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biggeri:1997:BRR

[1922] Annibale Biggeri. Book review: Recent Advances in Descriptive Multivariate Analysis. Wojtek J. Krzanowski, Oxford University Press, 1995. No. of pages: 384. Price: £35. ISBN: 0-19-852285-1. Statistics in Medicine, 16(15):1787–1788, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jensen:1997:BRR

[1923] Peter Jensen. Book review: Regression Models: Censored, Sampleselected, or Truncated Data. Richard Breen, Sage University Paper series on Quantitative Applications in the Social Sciences, series no. 07-111, Thousand Oaks, 1996. No. of pages: viii + 80. Price: £7.95. ISBN: 0-8039-5710-6. Statistics in Medicine, 16(15):1789, August 15, 1997. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1997:BRS

[1924] J. K. Lindsey. Book review: Statistical Applications Using Fuzzy Sets. K. G. Manton, M. A. Woodbury and H. D. Tolley, Wiley, New York, 1994. No. of pages: xiv + 312. Price: £53. ISBN: 0-471-54561-9. Statistics in Medicine, 16(15):1789–1790, August 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heitjan:1997:BIA

[1925] Daniel F. Heitjan. Bayesian interim analysis of phase II cancer clinical trials. Statistics in Medicine, 16(16):1791–1802, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Atkinson:1997:BSE

[1926] E. Neely Atkinson. A Bayesian strategy for evaluating treatments applicable only to a subset of patients. Statistics in Medicine, 16(16):1803– 1815, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Larose:1997:GRE

[1927] Daniel T. Larose and Dipak K. Dey. Grouped random effects models for Bayesian meta-analysis. Statistics in Medicine, 16(16):1817–1829, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mackenzie:1997:NPH

[1928] Gilbert Mackenzie. On a non-proportional hazards regression model for repeated medical random counts. Statistics in Medicine, 16(16):1831– 1843, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xie:1997:TPR

[1929] Tailiang Xie and Mikel Aickin. A truncated Poisson regression model with applications to occurrence of adenomatous polyps. Statistics in Medicine, 16(16):1845–1857, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:1997:TMC

[1930] Taesung Park and Seung-Yeoun Lee. A test of missing completely at random for longitudinal data with missing observations. Statistics in Medicine, 16(16):1859–1871, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1997:CLC

[1931] Rina Chen, Jose Iscovich, and Uri Goldbourt. Clustering of leukaemia cases in a city in Israel. Statistics in Medicine, 16(16):1873–1887, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lu:1997:CCG

[1932] Ying Lu, Keying Ye, Ashwini K. Mathur, Siu Hui, Thomas P. Fuerst, and Harry K. Genant. Comparative calibration without a gold standard. Statistics in Medicine, 16(16):1889–1905, August 30, 1997. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [2165, 2638].

MacEachern:1997:BRK

[1933] Steve MacEachern. Book review: Kendall’s Advanced Theory Of Statistics. Vol 2B: Bayesian Inference. Anthony O’Hagan, Edward Arnold, London, 1994. No. of pages: 330. Price: £35. ISBN: 0-340-52922-9. Statistics in Medicine, 16(16):1907–1908, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1997:BRQ

[1934] Peter Armitage. Book review: Quantification and the Quest For Medical Certainty. J. Rosser Matthews, Princeton University Press, Princeton, New Jersey, 1995. No. of pages: x + 195. Price: $39.50. ISBN: 0-691-03794-9. Statistics in Medicine, 16(16):1908–1909, August 30, 1997. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanStrik:1997:BRS

[1935] Roel van Strik. Book review: Statistics in the Pharmaceutical Industry. (2nd edn). C. R. Buncher and J. Y. Tsay (eds), Marcel Dekker, New York, 1994. No. of pages: xviii + 563. Price: $150. ISBN: 0-8247-9073-1. Statistics in Medicine, 16(16):1909–1910, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gehan:1997:BRC

[1936] Edmund A. Gehan. Book reviews: The Cambridge Dictionary of Statistics in the Medical Sciences. B. S. Everitt, Cambridge University Press, 1995. No. of pages: 277. Price: $54.95 (hardback), $19.95 (paperback). ISBN 0-521-47382-9 (hardback). ISBN: 0-521-47928-2 (paperback). Statistics in Medicine, 16(16):1910–1911, August 30, 1997. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sylvester:1997:BRC

[1937] Richard Sylvester. Book review: Clinical Trials Dictionary: Terminology and Usage Recommendations. Curtis L. Meinert, The John Hopkins Center for Clinical Trials, Maryland, 1996. No. of pages: cviii + 363. Price: $95. ISBN: 0-9646424-0-9. Statistics in Medicine, 16(16):1911, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Arminger:1997:BRL

[1938] Gerhard Arminger. Book review: LISREL — Issues, Debates, And Strategies. Leslie A. Hayduk, The Johns Hopkins University Press, 1996. No. of pages: xxii + 256. ISBN: 0-8018-5336-2. Statistics in Medicine, 16 (16):1911–1912, August 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:1997:DSS

[1939] Weichung Joseph Shih and Peng-Liang Zhao. Design for sample size reestimation with interim data for double-blind clinical trials with binary outcomes. Statistics in Medicine, 16(17):1913–1923, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lazaridis:1997:CMS

[1940] Emmanuel N. Lazaridis and René Gonin. Continuously monitored stopping boundary methodologies: the issues of sensitivity, association and trial suspension. Statistics in Medicine, 16(17):1925–1941, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matts:1997:GAP

[1941] John P. Matts, Cynthia A. Launer, Eileen T. Nelson, Carol Miller, and Bradley Dain. A graphical assessment of the potential impact of losses to follow-up on the validity of study results. Statistics in Medicine, 16(17): 1943–1954, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1997:TPI

[1942] J. K. Lindsey and B. Jones. Treatment-patient interactions for diagnostics of cross-over trials. Statistics in Medicine, 16(17):1955–1964, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Daniels:1997:MAE

[1943] Michael J. Daniels and Michael D. Hughes. Meta-analysis for the evaluation of potential surrogate markers. Statistics in Medicine, 16(17): 1965–1982, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xue:1997:RAD

[1944] Xiaonan Xue and Ron Brookmeyer. Regression analysis of discrete time survival data under heterogeneity. Statistics in Medicine, 16(17):1983– 1993, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

TenHave:1997:MPH

[1945] Thomas R. Ten Have and Edward O. Bixler. Modelling population heterogeneity in sensitivity and specificity of a multi-stage screen for obstructive sleep apnoea. Statistics in Medicine, 16(17):1995–2008, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cai:1997:EMH

[1946] Jianwen Cai, Haibo Zhou, and C. E. Davis. Estimating the mean hazard ratio parameters for clustered survival data with random clusters. Statistics in Medicine, 16(17):2009–2020, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1997:LEC

[1947] Stephen Senn. Letter to the Editor: The case for cross-over trials in phase III, by B. J. Jones and J. Lewis, Statistics in Medicine, 14, 1025–1038 (1995). Statistics in Medicine, 16(17):2021–2022, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1299].

Marshall:1997:LEH

[1948] Roger Marshall. Letter to the Editor: How independent are multiple ‘independent’ diagnostic classifications?, by H. Brenner, Statistics in Medicine, 15, 1377–1386 (1996). Statistics in Medicine, 16(17):2022– 2023, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1603].

Brenner:1997:AR

[1949] H. Brenner. Author’s reply. Statistics in Medicine, 16(17):2024, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moolgavkar:1997:BRW

[1950] Suresh H. Moolgavkar. Book review: Wings of Death: Nuclear Pollution and Human Health. Chris Busby, Green Audit Books, Aberystwyth, Wales, 1995. No. of pages: ix + 340. Price: £10.99. ISBN 1-897761-03-1. Statistics in Medicine, 16(17):2025–2026, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liski:1997:BRP

[1951] Erkki Liski. Book review: Practical Longitudinal Data Analysis. David Hand and Martin Crowder, Chapman and Hall, London, 1996. No of pages: x + 232. Price: £35. ISBN: 0-412-59940-6. Statistics in Medicine, 16(17):2026–2027, September 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Therneau:1997:RER

[1952] Terry M. Therneau and Scott A. Hamilton. rhDNase as an example of recurrent event analysis. Statistics in Medicine, 16(18):2029–2047, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:1997:ECT

[1953] Sujuan Gao and Xiao-Hua Zhou. An empirical comparison of two semiparametric approaches for the estimation of covariate effects from multivariate failure time data. Statistics in Medicine, 16(18):2049–2062, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1997:DAP

[1954] Simon G. Thompson, Stephen D. M. Pyke, and Rebecca J. Hardy. The design and analysis of paired cluster randomized trials: an application of meta-analysis techniques. Statistics in Medicine, 16(18):2063–2079, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rogerson:1997:SSM

[1955] Peter A. Rogerson. Surveillance systems for monitoring the development of spatial patterns. Statistics in Medicine, 16(18):2081–2093, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:1997:CFP

[1956] Qing Liu, Yulan Li, and James M. Boyett. Controlling false positive rates in prognostic factor analyses with small samples. Statistics in Medicine, 16(18):2095–2101, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rose:1997:FYS

[1957] M. Sarah Rose, Gordon H. Fick, Barbara Romanowski, and Edgar J. Love. First year serologic response to treatment for syphilis: a model for prediction of seroreversion. Statistics in Medicine, 16(18):2103–2115, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Milner:1997:TYF

[1958] P. C. Milner. Ten-year follow-up of ARIMA forecasts of attendances at accident and emergency departments in the trent region. Statistics in Medicine, 16(18):2117–2125, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

May:1997:CID

[1959] Warren L. May and William D. Johnson. Confidence intervals for differences in correlated binary proportions. Statistics in Medicine, 16(18): 2127–2136, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [2332, 2237, 2635].

VanHouwelingen:1997:BRA

[1960] Hans C. Van Houwelingen. Book review: Advances in Biometry. P. Armitage and H. A. David (eds), Wiley, New York, 1996. No. of pages: xii + 473. Price: $64.95. ISBN 0-471-16018-0. Statistics in Medicine, 16(18): 2137–2138, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:1997:BRA

[1961] Odd O. Aalen. Book review: Analysing Survival Data from Clinical Trials and Observational Studies. E. Marubini and M. G. Valsecchi, Wiley, Chichester, 1995. No. of Pages: xvi + 414. Price: £45. ISBN 0-471-939870. Statistics in Medicine, 16(18):2138–2139, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edler:1997:BRB

[1962] Lutz Edler. Book review: Biometrische Methodik Pharmakologischer Tierexperimente. Eberhard Gäbler, Gustav Fischer Verlag, Stuttgart, 1996. No. of pages: iv + 435. Price: DM 118. ISBN 3-437-20536-6. Statistics in Medicine, 16(18):2139–2142, September 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:1997:SNP

[1963] Kelly H. Zou, W. J. Hall, and David E. Shapiro. Smooth non-parametric receiver operating characteristic (ROC) curves for continuous diagnostic tests. Statistics in Medicine, 16(19):2143–2156, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Torrance-Rynard:1997:EDE

[1964] Vicki L. Torrance-Rynard and Stephen D. Walter. Effects of dependent errors in the assessment of diagnostic test performance. Statistics in Medicine, 16(19):2157–2175, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dear:1997:ECE

[1965] Keith B. G. Dear, Martin L. Puterman, and Annette J. Dobson. Estimating correlations from epidemiological data in the presence of measurement error. Statistics in Medicine, 16(19):2177–2189, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aalen:1997:MMH

[1966] Odd O. Aalen, Vernon T. Farewell, Daniela de Angelis, Nicholas E. Day, and O. Nöel Gill. A Markov model for HIV disease progression including the effect of HIV diagnosis and treatment: Application to AIDS prediction in England and Wales. Statistics in Medicine, 16(19):2191–2210, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laska:1997:CEC

[1967] Eugene M. Laska, Morris Meisner, and Dei-In Tang. Classification of the effectiveness of combination treatments. Statistics in Medicine, 16 (19):2211–2228, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vuorinen:1997:PAA

[1968] Jouni Vuorinen. A practical approach for the assessment of bioequivalence under selected higher-order cross-over designs. Statistics in Medicine, 16(19):2229–2243, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberger:1997:SDP

[1969] William F. Rosenberger and Stephen E. Grill. A sequential design for psychophysical experiments: an application to estimating timing of sensory events. Statistics in Medicine, 16(19):2245–2260, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holst:1997:BRI

[1970] Claus Holst. Book review: Item Response Theory. Parameter Estimation Techniques. Frank B. Baker, Marcel Dekker, New York, 1992. No. of pages: 440. Price: $150. ISBN 0-8247-8636-X. Statistics in Medicine, 16 (19):2261–2262, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:1997:BRI

[1971] R. J. Prescott. Book review: an Introduction To Medical Statistics. Martin Bland, Oxford University Press, Oxford, 1995. No of pages: xiv + 396. Price: £14.95. ISBN 0-19-262428-8. Statistics in Medicine, 16(19):2262– 2263, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachenbruch:1997:BRG

[1972] Peter A. Lachenbruch. Book review: a Guide to Chi-Squared Testing. Priscilla E. Greenwood and Mikhail S. Nikulin, Wiley, New York, 1996. No. of pages: xii + 280. Price: £29.95. ISBN 0-471-55779-X. Statistics in Medicine, 16(19):2263–2265, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osterlind:1997:BRM

[1973] Kell Østerlind. Book review: MEDSNAP. A Statistical Toolkit for Health Professionals. Version 2. Paul Strike, Butterworth–Heinemann, Oxford, 1996. 2 Disks Price: £70.50. Statistics in Medicine, 16(19):2265–2266, October 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moye:1997:SCT

[1974] Lemuel A. Moyé. Sizing clinical trials with variable endpoint event rates. Statistics in Medicine, 16(20):2267–2282, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lau:1997:LCM

[1975] Tai-Shing Lau. The latent class model for multiple binary screening tests. Statistics in Medicine, 16(20):2283–2295, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dyba:1997:SNL

[1976] Tadeusz Dyba, Timo Hakulinen, and Lassi Päivärinta. A simple nonlinear model in incidence prediction. Statistics in Medicine, 16(20):2297– 2309, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Daniels:1997:HPR

[1977] Michael J. Daniels and Constantine Gatsonis. Hierarchical polytomous regression models with applications to health services research. Statistics in Medicine, 16(20):2311–2325, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stromberg:1997:MDE

[1978] Ulf Strömberg. A method for deciding early stopping of inconclusive case-control studies in settings where data are stratified. Statistics in Medicine, 16(20):2327–2337, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2354].

Becker:1997:EDD

[1979] Niels G. Becker and Ji-Sheng Cui. Estimating a delay distribution from incomplete data, with application to reporting lags for AIDS cases. Statistics in Medicine, 16(20):2339–2347, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cnaan:1997:UGL

[1980] Avital Cnaan, Nan M. Laird, and Peter Slasor. Using the general linear mixed model to analyse unbalanced repeated measures and longitudinal data. Statistics in Medicine, 16(20):2349–2380, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schulgen:1997:BRE

[1981] Gabi Schulgen and Sören Kristiansen. Book review: Epidemiological Research Methods. Don McNeil, Chichester, 1996. No of pages: x + 305. Price: £19.99. ISBN 0-471-96195-7. Statistics in Medicine, 16(20):2381– 2382, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Plummer:1997:BRM

[1982] Martyn Plummer. Book review: Multivariable Analysis: an Introduction. Alvan R. Feinstein, Yale University Press, 1996. No. of pages: xvi + 630. Price: £45. ISBN 0-300-06299-0. Statistics in Medicine, 16(20):2382– 2383, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Christmann:1997:BRI

[1983] Andreas Christmann. Book review: an Introduction to Categorical Data Analysis. Alan Agresti, Wiley, New York, 1996. No. of pages: vii + 290. Price: £24.95. ISBN 0-471-11338-7. Statistics in Medicine, 16(20):2383– 2385, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bartlett:1997:BRN

[1984] Peter Bartlett. Book review: Neural Networks for Pattern Recognition. Christopher M. Bishop, Oxford University Press, New York, 1995. No. of pages: xvii + 512. Price: £55. ISBN 0-19-853864-2. Statistics in Medicine, 16(20):2385–2386, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alho:1997:BRA

[1985] Juha M. Alho. Book review: The Analysis of Time Series — An Introduction. Chris Chatfield, Chapman and Hall, London, 1996. No. of pages: xiii + 283. Price: £18.99. Statistics in Medicine, 16(20):2386–2387, October 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:1997:SPE

[1986] Michael P. Fay and Eric J. Feuer. A semi-parametric estimate of extra-Poisson variation for vital rates. Statistics in Medicine, 16(21):2389– 2401, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2649].

Hadgu:1997:AEP

[1987] A. Hadgu, G. Koch, and L. Westrom. Analysis of ectopic pregnancy data using marginal and conditional models. Statistics in Medicine, 16(21): 2403–2417, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gao:1997:LRM

[1988] Sujuan Gao and Siu L. Hui. Logistic regression models with missing covariate values for complex survey data. Statistics in Medicine, 16(21): 2419–2428, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Betensky:1997:LES

[1989] Rebecca A. Betensky. Local estimation of smooth curves for longitudinal data. Statistics in Medicine, 16(21):2429–2445, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailey:1997:RAH

[1990] Norman T. J. Bailey. A revised assessment of the HIV/AIDS incubation period, assuming a very short early period of high infectivity and using only San Francisco public health data on prevalence and incidence. Statistics in Medicine, 16(21):2447–2458, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seaman:1997:PAM

[1991] Shaun R. Seaman, Ray P. Brettle, and Sheila M. Gore. Pre-AIDS mortality in the Edinburgh city hospital HIV cohort. Statistics in Medicine, 16(21):2459–2474, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morrell:1997:CHP

[1992] Christopher H. Morrell, Jay D. Pearson, Larry J. Brant, and Sandra Gordon-Salant. Construction of hearing percentiles in women with nonconstant variance from the linear mixed-effects model. Statistics in Medicine, 16(21):2475–2488, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunnett:1997:MTE

[1993] Charles W. Dunnett and Ajit C. Tamhane. Multiple testing to establish superiority/equivalence of a new treatment compared with k standard treatments. Statistics in Medicine, 16(21):2489–2506, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hirji:1997:BRD

[1994] Karim F. Hirji. Book review: Deadly Medicine: Why Tens Of Thousands Of Patients Died In America’s Worst Drug Disaster. Thomas J. Moore, Simon and Shuster, New York, 1995. No. of pages: 349. Price: U.S.

$23.00. ISBN 0-684-80417-4. Statistics in Medicine, 16(21):2507–2510, November 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:1997:MCM

[1995] Barry W. Brown and Kathy Russell. Methods correcting for multiple testing: operating characteristics. Statistics in Medicine, 16(22):2511– 2528, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sankoh:1997:SCF

[1996] A. J. Sankoh, M. F. Huque, and S. D. Dubey. Some comments on frequently used multiple endpoint adjustment methods in clinical trials. Statistics in Medicine, 16(22):2529–2542, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lele:1997:DDR

[1997] Chitra Lele and Alice S. Whittemore. Different disease rates in two populations: how much is due to differences in risk factors? Statistics in Medicine, 16(22):2543–2554, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McNeil:1997:BEI

[1998] Alexander J. McNeil. Bayes estimates for immunological progression rates in HIV disease. Statistics in Medicine, 16(22):2555–2572, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Faraggi:1997:LSB

[1999] David Faraggi and Richard Simon. Large sample Bayesian inference on the parameters of the proportional hazard models. Statistics in Medicine, 16(22):2573–2585, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Betensky:1997:EMS

[2000] Rebecca A. Betensky and Camlin Tierney. An examination of methods for sample size recalculation during an experiment. Statistics in Medicine, 16(22):2587–2598, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chuang-Stein:1997:TBR

[2001] Christy Chuang-Stein and Alan Agresti. Tutorial in biostatistics a review of tests for detecting a monotone dose-response relationship with ordinal response data. Statistics in Medicine, 16(22):2599–2618, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:1997:LES

[2002] I. R. White. Letter to the Editor: Survival analysis of randomized clinical trials adjusted for patients who switch treatments, by M. G. Law and J. M. Kaldor, Statistics in Medicine, 15, 2069–2076 (1996). Statistics in Medicine, 16(22):2619–2620, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1680] and reply [2003].

Law:1997:ARS

[2003] Matthew G. Law and John M. Kaldor. Author’s reply: Survival analysis of randomized clinical trials adjusted for patients who switch treatments by M. G. Law and J. M. Kaldor, Statistics in Medicine, 15, 2069–2076 (1996). Statistics in Medicine, 16(22):2620–2621, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1680, 2002].

Tango:1997:LDO

[2004] T. Tango. Letter: Detection of overall space-time clustering in a nonuniformly distributed population, by Jukka Ranta, J. Pitkäniemi, M. Karvonen, E. Virtala, J. Rusanen, A. Colpaert, A. Naukkarinen and J. Tuomilehto, Statistics in Medicine, 15, 2561–2572 (1996). Statistics in Medicine, 16(22):2621–2623, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1733].

Ranta:1997:ARD

[2005] Jukka Ranta. Author’s reply: Detection of overall space-time clustering in a non-uniformly distributed population by Jukka Ranta, J. Pitkäniemi, M. Karvonen, E. Virtala, J. Rusanen, A. Colpaert, A. Naukkarinen and J. Tuomilehto, Statistics in Medicine, 15, 2561–2572 (1996). Statistics in Medicine, 16(22):2623–2625, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1733].

Hilden:1997:BRU

[2006] Jørgen Hilden. Book review: Using Designed Experiments to Shrink Health Care Costs. M. Daniel Sloan, Quality Press, Milwaukee WI, 1996. No. of pages: xvi + 215. Price: US$40.00. ISBN 0-87389-367-0. Statistics in Medicine, 16(22):2627–2628, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuha:1997:BRM

[2007] Jouni Kuha. Book review: Measurement Error in Nonlinear Models. R. J. Carroll, D. Ruppert and L. A. Stefanski, Chapman and Hall, London, 1995. No. of pages: xxiv + 305. Price: £29.95. ISBN 0-412-4721-7. Statistics in Medicine, 16(22):2628–2629, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wechsler:1997:BRS

[2008] Sergio Wechsler. Book review: Statistics at Square One. Ninth Edition, revised by M. J. Campbell, T. D. V. Swinscow, BMJ Publ. Group, London, 1996. No. of pages: 140. Price: £11. ISBN 0-7279-0916-9. Statistics in Medicine, 16(22):2629–2630, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dobrzeniecki:1997:BRM

[2009] A. B. Dobrzeniecki. Book review: Medical Imaging — Analysis of Multimodality 2D/3D Image. L. Beolchi and M. H. Kuhn, ISO Press, Amsterdam, 1994. No. of pages: 222. Price: £71. ISBN 90-5199-210-6. Statistics in Medicine, 16(22):2630–2631, November 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brenner:1997:ERL

[2010] Hermann Brenner, Irene Schmidtmann, and Christa Stegmaier. Effects of record linkage errors on registry-based follow-up studies. Statistics in Medicine, 16(23):2633–2643, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeLong:1997:CRA

[2011] Elizabeth R. DeLong, Eric D. Peterson, David M. DeLong, Lawrence H. Muhlbaier, Suzanne Hackett, and Daniel B. Mark. Comparing riskadjustment methods for provider profiling. Statistics in Medicine, 16 (23):2645–2664, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:1997:MLM

[2012] Huiqi Pan and Harvey Goldstein. Multi-level models for longitudinal growth norms. Statistics in Medicine, 16(23):2665–2678, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Langbehn:1997:DAU

[2013] Douglas R. Langbehn and Robert F. Woolson. Discriminant analysis using the unweighted sum of binary variables: a comparison of model selection methods. Statistics in Medicine, 16(23):2679–2700, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1997:OTS

[2014] T. Timothy Chen. Optimal three-stage designs for phase II cancer clinical trials. Statistics in Medicine, 16(23):2701–2711, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:1997:IMN

[2015] Nancy R. Cook. An imputation method for non-ignorable missing data in studies of blood pressure. Statistics in Medicine, 16(23):2713–2728, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:1997:PNP

[2016] James F. Troendle, R. Clifford Blair, Deborah Rumsey, and Pamela Moke. Parametric and non-parametric tests for the overall comparison of several treatments to a control when treatment is expected to increase variability. Statistics in Medicine, 16(23):2729–2739, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1997:IUR

[2017] Simon G. Thompson, Teresa C. Smith, and Stephen J. Sharp. Investigating underlying risk as a source of heterogeneity in meta-analysis. Statistics in Medicine, 16(23):2741–2758, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2333].

Ersboll:1997:BRI

[2018] Bjarne Kjær Ersbøll. Book review: Image Analysis for the Biological Sciences. Chris A. Glasbey and Graham W. Horgan, Wiley, Chichester, 1995. No. of pages: 218. Price: £29.95. ISBN 0-471-93726-6. Statistics in Medicine, 16(23):2759–2760, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shanmugam:1997:BRG

[2019] Ram Shanmugam. Book review: Graphing Statistics and Data. Anders Wallgren, Britt Wallgren, Rolf Persson, Ulf Jorner and Jan-Aage Haaland, Sage Publications, London, United Kingdom, 1996. No. of pages:

94. Price: £12.95. ISBN 0-7619-0599-5. Statistics in Medicine, 16(23): 2760–2761, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanderLaan:1997:BRP

[2020] Mark J. Van der Laan. Book review: Probabilistic Causality in Longitudinal Studies. Mervi Eerola, Springer-Verlag, New York, 1994. No. of pages: 131. Price: DM49. ISBN 3-540-94367-6. Statistics in Medicine, 16 (23):2761–2762, December 15, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1997:P

[2021] Ted Colton, Tony Johnson, and David Machin. Preface. Statistics in Medicine, 16(24):2765, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyse:1997:BTI

[2022] Marc Buyse. A biostatistical tribute to Ignaz Philip Semmelweis. Statistics in Medicine, 16(24):2767–2772, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanHouwelingen:1997:FBE

[2023] Hans C. Van Houwelingen. The future of biostatistics: expecting the unexpected. Statistics in Medicine, 16(24):2773–2784, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wright:1997:SEA

[2024] E. M. Wright and P. Royston. Simplified estimation of age-specific reference intervals for skewed data. Statistics in Medicine, 16(24):2785–2803, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Singer:1997:ESS

[2025] Júlia Singer. Estimating sample size for continuous outcomes, comparing more than two parallel groups with unequal sizes. Statistics in Medicine, 16(24):2805–2811, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schumacher:1997:RCV

[2026] Martin Schumacher, Norbert Holländer, and Willi Sauerbrei. Resampling and cross-validation techniques: a tool to reduce bias caused by model building? Statistics in Medicine, 16(24):2813–2827, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zwinderman:1997:SMC

[2027] Aeilko H. Zwinderman, J. Wouter Jukema, Ad J. Van Boven, and Johan H. C. Reiber. On the statistical modelling of coronary arteriographic data: dynamics of coronary atherosclerosis related to systemic and focal parameters. Statistics in Medicine, 16(24):2829–2841, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morel:1997:CBL

[2028] Jorge G. Morel and Nagaraj K. Neerchal. Clustered binary logistic regression in teratology data using a finite mixture distribution. Statistics in Medicine, 16(24):2843–2853, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frison:1997:LDT

[2029] Lars J. Frison and Stuart J. Pocock. Linearly divergent treatment effects in clinical trials with repeated measures: efficient analysis using summary statistics. Statistics in Medicine, 16(24):2855–2872, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1997:SMR

[2030] J. K. Lindsey, B. Jones, and A. F. Ebbutt. Simple models for repeated ordinal responses with an application to a seasonal rhinitis clinical trial. Statistics in Medicine, 16(24):2873–2882, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1997:VBR

[2031] S. D. Walter. Variation in baseline risk as an explanation of heterogeneity in meta-analysis. Statistics in Medicine, 16(24):2883–2900, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2343].

Whitehead:1997:PPC

[2032] Anne Whitehead. A prospectively planned cumulative meta-analysis applied to a series of concurrent clinical trials. Statistics in Medicine, 16 (24):2901–2913, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Todd:1997:IST

[2033] Susan Todd. Incorporation of sequential trials into a fixed effects metaanalysis. Statistics in Medicine, 16(24):2915–2925, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1997:C

[2034] A. Donner. Correction. Statistics in Medicine, 16(24):2927, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1997:C

[2035] Richard H. Jones. Correction. Statistics in Medicine, 16(24):2927–2928, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jorgensen:1997:C

[2036] Bent Jørgensen. Correction. Statistics in Medicine, 16(24):2928, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bull:1997:C

[2037] S. B. Bull. Correction. Statistics in Medicine, 16(24):2928, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grandjean:1997:BRY

[2038] Philippe Grandjean. Book review: Up to Your Armpits in Alligators? How to Sort Out What Risks are Worth Worrying About! J. Paling, Risk Communication and Environmental Institute, Gainesville, 1997. No. of pages: 153. ISBN 0-9642236-6-X. Statistics in Medicine, 16(24):2929, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DerSimonian:1997:BRS

[2039] Rebecca DerSimonian. Book review: Systematic Review. I. Chalmers and D. G. Altman (eds), BMJ Publishing Group, London, 1995. No. of pages: 117. Price: £17. ISBN 0-7279-0904-5. Statistics in Medicine, 16(24):2930, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:1997:BRS

[2040] Nick Longford. Book review: Seeing Through Statistics. Jessica M. Utts, Duxbury Press, Belmont, CA, 1996. No. of pages: xvi + 464. Price: £21.95. ISBN 0-534-25776-3. Statistics in Medicine, 16(24):2931–2932, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fortuin:1997:BRS

[2041] F. David Fortuin and Joan F. Hilton. Book review: Statistics from Scratch: an Introduction for Health Care Professionals. David Bowers, Wiley, Chichester, 1996. No. of pages: xi + 185. Price: £12.99. ISBN 0-471-96325-9. Statistics in Medicine, 16(24):2932–2933, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goetghebeur:1997:BRS

[2042] Els Goetghebeur. Book review: Survival Analysis: a Self-learning Text. David G. Kleinbaum, Springer series in Statistics, Statistics in the Health Sciences, 1996. No. of pages: xii + 324. Price: DM 58. ISBN 0-387-94543-1. Statistics in Medicine, 16(24):2933–2935, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Quistorff:1997:BRN

[2043] Bjørn Quistorff. Book review: NMR Data Processing. Jeffrey C. Hoch and Alan S. Stern. Wiley, New York, 1996. No. of pages: 195. Price: £50. ISBN 0471-03900-4. Statistics in Medicine, 16(24):2935–2936, December 30, 1997. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Foulkes:1998:AHA

[2044] Mary A. Foulkes. Advances in HIV/AIDS statistical methodology over the past decade. Statistics in Medicine, 17(1):1–25, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsodikov:1998:CMT

[2045] A. Tsodikov, M. Loeffler, and A. Yakovlev. A cure model with time-changing risk factor: an application to the analysis of secondary leukaemia. a report from the international database on Hodgkin’s disease. Statistics in Medicine, 17(1):27–40, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Langford:1998:MLM

[2046] Ian H. Langford, Graham Bentham, and Anné-Lise McDonald. Multilevel modelling of geographically aggregated health data: a case study on malignant melanoma mortality and UV exposure in the European community. Statistics in Medicine, 17(1):41–57, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1998:CAG

[2047] J. K. Lindsey and B. Jones. Choosing among generalized linear models applied to medical data. Statistics in Medicine, 17(1):59–68, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Regal:1998:MVC

[2048] Ronald R. Regal and Ernest B. Hook. Marginal versus conditional versus ‘structural source’ models: a rationale for an alternative to log-linear methods for capture-recapture estimates. Statistics in Medicine, 17(1): 69–74, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scharfstein:1998:USB

[2049] Daniel O. Scharfstein and Anastasios A. Tsiatis. The use of simulation and bootstrap in information-based group sequential studies. Statistics in Medicine, 17(1):75–87, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:1998:MDO

[2050] A. Lawrence Gould and W. Joseph Shih. Modifying the design of ongoing trials without unblinding. Statistics in Medicine, 17(1):89–100, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1998:SSO

[2051] S. D. Walter, M. Eliasziw, and A. Donner. Sample size and optimal designs for reliability studies. Statistics in Medicine, 17(1):101–110, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Duchateau:1998:PES

[2052] L. Duchateau, B. McDermott, and G. J. Rowlands. Power evaluation of small drug and vaccine experiments with binary outcomes. Statistics in Medicine, 17(1):111–120, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Santner:1998:BRI

[2053] Thomas Santner. Book review: Introduction to graphical modelling. David Edwards, Springer, New York, 1995. No. of pages: 274. Price: DM68. ISBN 0-387-94483-4. Statistics in Medicine, 17(1):121–122, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ulmer:1998:BRC

[2054] Hanno Ulmer. Book review: a casebook for a first course in statistics and data analysis. Samprit Chatterjee, Mark S. Handcock and Jeffrey S. Simonoff, Wiley, New York, 1995. No. of pages: 314. Price: £19.95. ISBN 3-471-11030-2. Statistics in Medicine, 17(1):122, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brunner:1998:BRU

[2055] Edgar Brunner and Olaf Gefeller. Book review: Using and understanding medical statistics. D. E. Matthews and V. T. Farewell, Karger, Switzerland, 1996. No. of pages: 246. Price: $35. ISBN 3-8055-6276-4. Statistics in Medicine, 17(1):122–123, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ulm:1998:BRG

[2056] Kurt Ulm. Book review: Grundbegriffe der Biometrie. Rolf J. Lorenz, Gustav Fischer, Stuttgart, 1996. No. of pages: 228. Price: DM 48. ISBN 3-437-25100-7. Statistics in Medicine, 17(1):123–124, January 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karon:1998:P

[2057] John M. Karon. Preface. Statistics in Medicine, 17(2):125, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Karon:1998:CSM

[2058] John M. Karon, Meena Khare, and Philip S. Rosenberg. The current status of methods for estimating the prevalence of human immunodeficiency virus in the United States of America. Statistics in Medicine, 17(2):127–142, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Green:1998:USD

[2059] Timothy A. Green. Using surveillance data to monitor trends in the AIDS epidemic. Statistics in Medicine, 17(2):143–154, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberg:1998:ECI

[2060] Philip S. Rosenberg and James J. Goedert. Estimating the cumulative incidence of HIV infection among persons with haemophilia in the United States of America. Statistics in Medicine, 17(2):155–168, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Byers:1998:PAH

[2061] Robert H. Byers, Jr., M. Blake Caldwell, Susan Davis, Marta Gwinn, and Mary Lou Lindegren. Projection of AIDS and HIV incidence among children born infected with HIV. Statistics in Medicine, 17(2):169–181, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Datta:1998:AHV

[2062] Susmita Datta, M. Elizabeth Halloran, and Ira M. Longini, Jr. Augmented HIV vaccine trial design for estimating reduction in infectiousness and protective efficacy. Statistics in Medicine, 17(2):185–200, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Abrams:1998:ABI

[2063] Keith Abrams and Bruno Sansó. Approximate Bayesian inference for random effects meta-analysis. Statistics in Medicine, 17(2):201–218, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1998:MIC

[2064] Jane C. Lindsey and Louise M. Ryan. Methods for interval-censored data. Statistics in Medicine, 17(2):219–238, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flandre:1998:LEA

[2065] Philippe Flandre. Letter to the Editor: Analysing survival in the presence of an auxiliary variable, by D. M. Finkelstein and D. A. Schoenfeld, Statistics in Medicine, 13, 1747–1754 (1994). Statistics in Medicine, 17 (2):239–241, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1103].

Nelder:1998:LEB

[2066] J. A. Nelder. Letter to the Editor: A Bayesian analysis of regression models with continuous errors with application to longitudinal studies, by L. D. Broemeling and P. Cook, Statistics in Medicine, 16, 321–332 (1997). Statistics in Medicine, 17(2):241, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1773].

Becher:1998:BRM

[2067] Heiko Becher. Book review: Medical Biostatistics and Epidemiology. Diane Essex-Saulie, Prentice Hall, 1995. No. of pages: 408. Price: $37.95. ISBN 0-8385-6214-0. Statistics in Medicine, 17(2):243, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kneip:1998:BRG

[2068] Alois Kneip. Book review: Growth Curves. A. M. Kshirsagar and W. B. Smith, New York, 1995. No. of pages: 359. Price: $135. ISBN 0-8247-9341-2. Statistics in Medicine, 17(2):244, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tjur:1998:BRP

[2069] Tue Tjur. Book review: Practical Data Analysis for Designed Experiments. Brian S. Yandell, Chapman and Hall, London, 1997. No. of pages: xii + 437. Price: US$59.95, UK£32.50. ISBN 0-412-06341-7. Statistics in Medicine, 17(2):244–245, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gather:1998:BRC

[2070] Ursula Gather. Book review: a Course in Probability and Statistics. Charles J. Stone, Duxbury Press, Belmont, 1996. No. of pages: 838. Price: £24.95. ISBN 0-534-23328-7. Statistics in Medicine, 17(2):246, January 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goetghebeur:1998:P

[2071] Els Goetghebeur. Preface. Statistics in Medicine, 17(3):247, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Efron:1998:FLC

[2072] Bradley Efron. Foreword: Limburg Compliance Symposium. Statistics in Medicine, 17(3):249–250, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Urquhart:1998:CPI

[2073] John Urquhart and Erik De Klerk. Contending paradigms for the interpretation of data on patient compliance with therapeutic drug regimens. Statistics in Medicine, 17(3):251–267, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robins:1998:CNC

[2074] James M. Robins. Correction for non-compliance in equivalence trials. Statistics in Medicine, 17(3):269–302, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pocock:1998:HHU

[2075] Stuart J. Pocock and Mona Abdalla. The hope and the hazards of using compliance data in randomized controlled trials. Statistics in Medicine, 17(3):303–317, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:1998:CTC

[2076] Ian R. White and Els J. T. Goetghebeur. Clinical trials comparing two treatment policies: which aspects of the treatment policies make a difference? Statistics in Medicine, 17(3):319–339, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goetghebeur:1998:ECE

[2077] Els Goetghebeur, Geert Molenberghs, and Joanne Katz. Estimating the causal effect of compliance on binary outcome in randomized controlled trials. Statistics in Medicine, 17(3):341–355, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1998:CAH

[2078] David M. Smith and Peter J. Diggle. Compliance in an anti-hypertension trial: a latent process model for binary longitudinal data. Statistics in Medicine, 17(3):357–370, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubin:1998:MPR

[2079] Donald B. Rubin. More powerful randomization-based p-values in double-blind trials with non-compliance. Statistics in Medicine, 17(3): 371–385, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cox:1998:D

[2080] David Cox. Discussion. Statistics in Medicine, 17(3):387–389, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heesterbeek:1998:BRI

[2081] Hans Heesterbeek. Book review: an Introduction to the Mathematics of Biology: With Computer Algebra Models. E. K. Yeargers, R. W. Shonkwiler and J. V. Herod, Birkhäuser, Boston, 1996. No. of pages: 433. Price: DM118. ISBN 0-8176-3809-1. Statistics in Medicine, 17(3):391– 392, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:1998:BRP

[2082] Michael A. Lewis. Book review: Pharmacoepidemiology. Brian L. Strom, Wiley, Chichester, 1994. No. of pages: 741. Price: £80. ISBN 047-194-0585. Statistics in Medicine, 17(3):392–393, February 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chetwynd:1998:MMS

[2083] Amanda G. Chetwynd, Peter J. Diggle, Robert F. Drewett, and Bridget Young. A mixture model for sucking patterns of breast-fed infants. Statistics in Medicine, 17(4):395–405, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:1998:BGR

[2084] T. J. Cole, J. V. Freeman, and M. A. Preece. British 1990 growth reference centiles for weight, height, body mass index and head circumference fitted by maximum penalized likelihood. Statistics in Medicine, 17(4): 407–429, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:1998:ABB

[2085] Chris Robertson and Peter Boyle. Age at any birth and breast cancer risk. Statistics in Medicine, 17(4):431–445, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1998:AMM

[2086] J. K. Lindsey and P. Lambert. On the appropriateness of marginal models for repeated measurements in clinical trials. Statistics in Medicine, 17(4):447–469, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:1998:MKT

[2087] Chris Roberts and Roseanne McNamee. A matrix of kappa-type coefficients to assess the reliability of nominal scales. Statistics in Medicine, 17(4):471–488, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rogel:1998:GPA

[2088] A. Rogel, P. Y. Boëlle, and J. Y. Mary. Global and partial agreement among several observers. Statistics in Medicine, 17(4):489–501, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Blomqvist:1998:LEE

[2089] Nils Blomqvist. Letter to the Editor: Estimating confidence intervals for cost-effectiveness ratios: an example from a randomized trial, by M. A. Chaudhary and S. C. Stearns, Statistics in Medicine, 15, 1447–1458 (1996). Statistics in Medicine, 17(4):503, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1609].

Bayer:1998:BRP

[2090] Thomas Bayer. Book review: Principles of Statistical Data Handling. Fred Davidson, SAGE Publications, U.S.A., 1996. No. of pages: xii + 319. Price: $125. ISBN 0-7619-0103-5. Statistics in Medicine, 17(4):505– 506, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kent:1998:BRS

[2091] John T. Kent. Book review: The Statistical Theory of Shape. Christopher G. Small, Springer-Verlag, New York, 1996. No. of pages: x + 227. Price: DM 78. ISBN 0-387-94729-9. Statistics in Medicine, 17(4):506– 507, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Commenges:1998:BRS

[2092] Daniel Commenges. Book review: Survival Analysis: Techniques for Censored and Truncated Data. John P. Klein and Melvin L.

Moeschberger, Springer Verlag, 1997. No. of pages: 502. Price: DM 94. ISBN 0-387-94829-5. Statistics in Medicine, 17(4):507, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pigeot:1998:BRM

[2093] Iris Pigeot. Book review: Modelling Frequency and Count Data. J. K. Lindsey, Oxford University Press, New York, 1995. No. of pages: ix + 291. Price: £30. ISBN 0-19-852331-9. Statistics in Medicine, 17(4):507– 509, February 28, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernhard:1998:P

[2094] Jürg Bernhard, Richard D. Gelber, and Christoph Hürny. Preface. Statistics in Medicine, 17(5-7):511–512, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernhard:1998:MQL

[2095] Jürg Bernhard, David F. Cella, Alan S. Coates, Lesley Fallowfield, Patricia A. Ganz, Carol M. Moinpour, Paola Mosconi, David Osoba, John Simes, and Christoph Hürny. Missing quality of life data in cancer clinical trials: serious problems and challenges. Statistics in Medicine, 17(5-7): 517–532, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coates:1998:QLS

[2096] Alan Coates and Val Gebski. Quality of life studies of the Australian new Zealand breast cancer trials group: approaches to missing data. Statistics in Medicine, 17(5-7):533–540, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fallowfield:1998:CIQ

[2097] Lesley Fallowfield. Compliance issues in quality of life assessment: experiences of two cancer research campaign sponsored groups. Statistics in Medicine, 17(5-7):541–546, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hahn:1998:MDQ

[2098] Elizabeth A. Hahn, Kimberly A. Webster, David Cella, and Diane L. Fairclough. Missing data in quality of life research in Eastern cooperative oncology group (ECOG) clinical trials: problems and solutions. Statistics in Medicine, 17(5-7):547–559, March–April 15–15, 1998. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kiebert:1998:QLE

[2099] Gwendoline M. Kiebert, Desmond Curran, and Neil K. Aaronson. Quality of life as an endpoint in EORTC clinical trials. Statistics in Medicine, 17(5-7):561–569, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olschewski:1998:CQA

[2100] Manfred Olschewski. Compliance with QOL assessment in multi-centre German breast cancer trials. Statistics in Medicine, 17(5-7):571–575, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mosconi:1998:MCA

[2101] Paola Mosconi, Valter Torri, Stefania Cifani, Resy Ruggiata, Beth E. Meyerowitz, Giovanni Apolone, and Alessandro Liberati. The multicentre assessment of quality of life: the interdisciplinary group for cancer care evaluation (GIVIO) experience in Italy. Statistics in Medicine, 17 (5-7):577–585, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernhard:1998:QLA

[2102] Jürg Bernhard, Harriet F. Peterson, Alan S. Coates, Heidi Gusset, Mary Isley, Rita Hinkle, Richard D. Gelber, Monica Castiglione-Gertsch, and Christoph Hürny. Quality of life assessment in international breast cancer study group (IBCSG) trials: practical issues and factors associated with missing data. Statistics in Medicine, 17(5-7):587–601, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osoba:1998:CRH

[2103] David Osoba and Benny Zee. Completion rates in health-related qualityof-life assessment: approach of the national cancer institute of Canada clinical trials group. Statistics in Medicine, 17(5-7):603–612, March– April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ganz:1998:CQL

[2104] Patricia A. Ganz, Richard Day, and Joseph Costantino. Compliance with quality of life data collection in the national surgical adjuvant breast and bowel project (NSABP) breast cancer prevention trial. Statistics in Medicine, 17(5-7):613–622, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kaasa:1998:CQL

[2105] Stein Kaasa, Marianne Jensen Hjermstad, Marit S. Jordhøy, Finn Wisløff, and Jon Håvard Loge. Compliance in quality of life data: a Norwegian experience. Statistics in Medicine, 17(5-7):623–632, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernhard:1998:PIQ

[2106] Jürg Bernhard, Heidi Gusset, and Christoph Hürny. Practical issues in quality of life assessment in multicentre trials conducted by the Swiss group for clinical cancer research. Statistics in Medicine, 17(5-7):633– 639, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Moinpour:1998:EQQ

[2107] Carol M. Moinpour and Laura C. Lovato. Ensuring the quality of quality of life data: the Southwest Oncology Group experience. Statistics in Medicine, 17(5-7):641–651, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troxel:1998:SAQ

[2108] Andrea B. Troxel, Diane L. Fairclough, Desmond Curran, and Elizabeth A. Hahn. Statistical analysis of quality of life with missing data in cancer clinical trials. Statistics in Medicine, 17(5-7):653–666, March– April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fairclough:1998:WMQ

[2109] Diane L. Fairclough, Harriet F. Peterson, and Victor Chang. Why are missing quality of life data a problem in clinical trials of cancer therapy? Statistics in Medicine, 17(5-7):667–677, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fayers:1998:IQL

[2110] Peter M. Fayers, Desmond Curran, and David Machin. Incomplete quality of life data in randomized trials: missing items. Statistics in Medicine, 17(5-7):679–696, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curran:1998:IQL

[2111] D. Curran, G. Molenberghs, P. M. Fayers, and D. Machin. Incomplete quality of life data in randomized trials: missing forms. Statistics in Medicine, 17(5-7):697–709, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Machin:1998:SPQ

[2112] David Machin and Simon Weeden. Suggestions for the presentation of quality of life data from clinical trials. Statistics in Medicine, 17(5-7): 711–724, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Simes:1998:PAM

[2113] R. John Simes, V. Greatorex, and V. J. Gebski. Practical approaches to minimize problems with missing quality of life data. Statistics in Medicine, 17(5-7):725–737, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Curran:1998:ITM

[2114] D. Curran, M. Bacchi, S. F. Hsu Schmitz, G. Molenberghs, and R. J. Sylvester. Identifying the types of missingness in quality of life data from clinical trials. Statistics in Medicine, 17(5-7):739–756, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zee:1998:GCM

[2115] Benny C. Zee. Growth curve model analysis for quality of life data. Statistics in Medicine, 17(5-7):757–766, March–April 15–15, 1998. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troxel:1998:CAQ

[2116] Andrea B. Troxel. A comparative analysis of quality of life data from a Southwest Oncology Group randomized trial of advanced colorectal cancer. Statistics in Medicine, 17(5-7):767–779, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fairclough:1998:CSM

[2117] Diane L. Fairclough, Harriet F. Peterson, David Cella, and Phil Bonomi. Comparison of several model-based methods for analysing incomplete quality of life data in cancer clinical trials. Statistics in Medicine, 17(5-7): 781–796, March–April 15–15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guggenmoos-Holzmann:1998:KLI

[2118] Irene Guggenmoos-Holzmann and Richard Vonk. Kappa-like indices of observer agreement viewed from a latent class perspective. Statistics in Medicine, 17(8):797–812, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Peng:1998:GMM

[2119] Yingwei Peng, Keith B. G. Dear, and J. W. Denham. A generalized F mixture model for cure rate estimation. Statistics in Medicine, 17(8): 813–830, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gieser:1998:MCR

[2120] Peter W. Gieser, Myron N. Chang, P. V. Rao, Jonathan J. Shuster, and Jeanette Pullen. Modelling cure rates using the Gompertz model with covariate information. Statistics in Medicine, 17(8):831–839, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hardy:1998:DDH

[2121] Rebecca J. Hardy and Simon G. Thompson. Detecting and describing heterogeneity in meta-analysis. Statistics in Medicine, 17(8):841–856, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:1998:TSC

[2122] Robert G. Newcombe. Two-sided confidence intervals for the single proportion: comparison of seven methods. Statistics in Medicine, 17(8): 857–872, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2648].

Newcombe:1998:IED

[2123] Robert G. Newcombe. Interval estimation for the difference between independent proportions: comparison of eleven methods. Statistics in Medicine, 17(8):873–890, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:1998:ETC

[2124] Toshiro Tango. Equivalence test and confidence interval for the difference in proportions for the paired-sample design. Statistics in Medicine, 17 (8):891–908, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Harrell:1998:DCP

[2125] Frank E. Harrell, Jr., Peter A. Margolis, Sandy Gove, Karen E. Mason, E. Kim Mulholland, Deborah Lehmann, Lulu Muhe, Salvacion Gatchalian, and Heinz F. Eichenwald. Development of a clinical prediction model for an ordinal outcome: the world health organization multicentre study of clinical signs and etiological agents of pneumonia, sepsis and meningitis in Young infants. Statistics in Medicine, 17(8): 909–944, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yu:1998:LEC

[2126] Kai Fun Yu. Letter to the Editor: Confidence limits based on the first occurrence of an event, by V. T. George and R. C. Elston, Statistics in Medicine, 12, 685–690 (1993). Statistics in Medicine, 17(8):945, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [727].

Elston:1998:ARC

[2127] Robert Elston and Varghese George. Authors’ reply: Confidence limits based on the first occurrence of an event. V. T. George and R. C. Elton. Statistics in Medicine, 12, 685–690 (1993). Statistics in Medicine, 17 (8):945, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [727].

Bender:1998:LEV

[2128] Ralf Bender, Stefan Lange, Gudrun Freitag, and Hans Joachim Trampisch. Letter to the Editor: Variation of sensitivity, specificity, likelihood ratios and predictive values with disease prevalence, by H. Brenner and O. Gefeller, Statistics in Medicine, 16, 981–991 (1997). Statistics in Medicine, 17(8):946–947, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1840] and reply [2129].

Brenner:1998:ARV

[2129] Hermann Brenner and Olaf Gefeller. Authors’ reply: Variation of sensitivity, specificity, likelihood ratios and predictive values with disease prevalence by H. Brenner and O. Gefeller, Statistics in Medicine, 16, 981–991 (1997). Statistics in Medicine, 17(8):947–948, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1840, 2128].

Channon:1998:LET

[2130] E. J. Channon. Letter to the Editor: Tests for equality of two correlated proportions, by W. L. May and W. D. Johnson, Statistics in Medicine, 16, 1081–1096 (1997). Statistics in Medicine, 17(8):949–950, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1851].

Dalgaard:1998:BRSa

[2131] Peter Dalgaard. Book review: Statistical Computing Environments for Social Research. Robert Stine and John Fox, Sage Publications, Thousand Oaks, California, 1997. No. of pages: 250. Price: £39.95. ISBN 0-7619-0269-4. Statistics in Medicine, 17(8):951–952, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laake:1998:BRA

[2132] Petter Laake. Book review: Analysis of Change. Advanced Techniques in Panel Data Analysis. Uwe Engel and Jost Reinecke (eds), Walter de Gruyter, Berlin 1996. No. of pages: x + 296. Price: DM 148. ISBN 3-11-014936-2. Statistics in Medicine, 17(8):952–953, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nilsson:1998:BRB

[2133] Per E. W. Nilsson and Hans Wedel. Book review: Basic Epidemiological Methods and Biostatistics. A Practical Handbook. Randy M. Page, Galen O. Cole and Thomas C. Timmereck, Jones and Bartlett Publishers, Boston–London, 1995. No. of pages: 433. Price: £19.95. ISBN 0-86720-869-4. Statistics in Medicine, 17(8):954–955, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Granfeldt:1998:BRA

[2134] Jorgen Granfeldt. Book review: Analysis of Variance, Design and Regression. Ronald Christensen, Chapman and Hall, 1996. No. of pages: xvi + 587. Price: £45. ISBN 0-412-06291-7. Statistics in Medicine, 17 (8):955–956, April 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:1998:MOR

[2135] Frank B. Hu, Jack Goldberg, Donald Hedeker, and William G. Henderson. Modelling ordinal responses from co-twin control studies. Statistics in Medicine, 17(9):957–970, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Strauss:1998:EKM

[2136] David Strauss and Robert Shavelle. An extended Kaplan–Meier estimator and its applications. Statistics in Medicine, 17(9):971–982, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1998:ACE

[2137] Eunyoung Lee and Lisa A. Weissfeld. Assessment of covariate effects in Aalen’s additive hazard model. Statistics in Medicine, 17(9):983–998, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:1998:PLM

[2138] Yu Shen and Peter F. Thall. Parametric likelihoods for multiple nonfatal competing risks and death. Statistics in Medicine, 17(9):999–1015, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marschner:1998:FAT

[2139] Ian C. Marschner and Ronald J. Bosch. Flexible assessment of trends in age-specific HIV incidence using two-dimensional penalized likelihood. Statistics in Medicine, 17(9):1017–1031, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Metz:1998:MLE

[2140] Charles E. Metz, Benjamin A. Herman, and Jong-Her Shen. Maximum likelihood estimation of receiver operating characteristic (ROC) curves from continuously-distributed data. Statistics in Medicine, 17(9):1033– 1053, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:1998:LEA

[2141] T. Tango. Letter to the Editor: Adjusting Moran’s I for population density, by N. Oden, Statistics in Medicine, 14, 17–26 (1995). Statistics in Medicine, 17(9):1055–1058, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1625].

Oden:1998:AR

[2142] Neal Oden, Geoffrey Jacquez, and Roger Grimson. Authors’ reply. Statistics in Medicine, 17(9):1058–1062, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:1998:LET

[2143] James F. Troendle. Letter to the Editor: Testing for treatment differences with dropouts present in clinical trials — a composite approach, by Weichung J. Shih and Hui Quan, Statistics in Medicine, 16, 1225–1239 (1997). Statistics in Medicine, 17(9):1062–1063, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1865].

Shih:1998:AR

[2144] Weichung Joe Shih and Hui Quan. Authors’ reply. Statistics in Medicine, 17(9):1063–1064, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chernesky:1998:LEB

[2145] Max Chernesky, John Sellors, and James Mahony. Letter to the Editor: Bias in the evaluation of DNA-amplification tests for detecting Chlamydia trachomatis, by A. Hadgu, Statistics in Medicine, 16, 1391–1399 (1997). Statistics in Medicine, 17(9):1064–1065, May 15, 1998. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1884, 2190].

Hadgu:1998:ARa

[2146] Alula Hadgu. Author’s reply. Statistics in Medicine, 17(9):1065–1066, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bech:1998:BRH

[2147] Per Bech. Book reviews: Health Measurement Scales: a Practical Guide to their Development and Use. David L. Streiner and Geoffrey R. Norman, Oxford University Press, Oxford, 1995, second edition. No. of pages:

231. Price: £40 (paperback) £18.50 (hardback). ISBN 0-19-262637X (Hbk) ISBN 0-19-262670-1 (pbk). Statistics in Medicine, 17(9):1067– 1068, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:1998:BRS

[2148] Vern Farewell. Book review: Survival Analysis with Long-term Survivors. Ross Maller and Xian Zhou, Wiley, 1996. No. of pages: xvi + 278. Price: £39.95. ISBN 0-471-96201-5. Statistics in Medicine, 17(9):1068–1069, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sundberg:1998:BRP

[2149] Rolf Sundberg. Book review: Parametric Statistical Inference. J. K. Lindsey, Oxford University Press, 1996. No. of pages: xviii + 490. Price £35. ISBN 0-19-852359-9. Statistics in Medicine, 17(9):1069–1071, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pickles:1998:BRT

[2150] Andrew Pickles. Book review: Textbook in Psychiatric Epidemiology. Ming T. Tsuang, Mauricio Tohen and Gwendolyn E. P. Zahner (eds), Wiley, New York, 1995. No. of pages: xii + 483. Price: £45. ISBN 0-471-59375-3. Statistics in Medicine, 17(9):1071, May 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joffe:1998:EEZ

[2151] Marshall M. Joffe, Donald R. Hoover, Lisa P. Jacobson, Lawrence Kingsley, Joan S. Chmiel, Barbara R. Visscher, and James M. Robins. Estimating the effect of zidovudine on Kaposi’s sarcoma from observational data using a rank preserving structural failure-time model. Statistics in Medicine, 17(10):1073–1102, May 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Babb:1998:CPC

[2152] James Babb, André Rogatko, and Shelemyahu Zacks. Cancer phase I clinical trials: efficient dose escalation with overdose control. Statistics in Medicine, 17(10):1103–1120, May 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longini:1998:OVT

[2153] Ira M. Longini, Jr., Karen Sagatelian, Wasima N. Rida, and M. Elizabeth Halloran. Optimal vaccine trial design when estimating vaccine efficacy for susceptibility and infectiousness from multiple populations. Statistics in Medicine, 17(10):1121–1136, May 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ecochard:1998:MLM

[2154] Rene Ecochard and David G. Clayton. Multi-level modelling of conception in artificial insemination by donor. Statistics in Medicine, 17(10): 1137–1156, May 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Donner:1998:SSR

[2155] Allan Donner. Sample size requirements for the comparison of two or more coefficients of inter-observer agreement. Statistics in Medicine, 17 (10):1157–1168, May 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Biganzoli:1998:FFN

[2156] Elia Biganzoli, Patrizia Boracchi, Luigi Mariani, and Ettore Marubini. Feed forward neural networks for the analysis of censored survival data: a partial logistic regression approach. Statistics in Medicine, 17(10): 1169–1186, May 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heitjan:1998:BRE

[2157] Daniel F. Heitjan. Book review: The EM Algorithm and Extensions. Geoffrey J. McLachlan and Thriyambakam Krishnan, Wiley, New York, 1997. No. of pages: xvii + 274. Price: $59.95. ISBN 0-471-12358-7. Statistics in Medicine, 17(10):1187, May 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lie:1998:ESC

[2158] Stein Atle Lie, Rolv Terje Lie, and Cecilie Svanes. Expected survival compared with survival of peptic ulcer patients. Statistics in Medicine, 17(11):1189–1199, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yau:1998:MRE

[2159] K. K. W. Yau and C. A. McGilchrist. ML and REML estimation in survival analysis with time dependent correlated frailty. Statistics in Medicine, 17(11):1201–1213, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glasziou:1998:QAS

[2160] Paul P. Glasziou, Bernard F. Cole, Richard D. Gelber, Jorgen Hilden, and R. John Simes. Quality adjusted survival analysis with repeated quality of life measures. Statistics in Medicine, 17(11):1215–1229, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:1998:EAS

[2161] Toshiro Tango. Estimation of age-specific reference ranges via smoother AVAS. Statistics in Medicine, 17(11):1231–1243, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:1998:COS

[2162] James F. Troendle and Julie M. Legler. A comparison of one-sided methods to identify significant individual outcomes in a multiple outcome setting: stepwise tests or global tests with closed testing. Statistics in Medicine, 17(11):1245–1260, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Burton:1998:ESL

[2163] Paul Burton, Lyle Gurrin, and Peter Sly. Extending the simple linear regression model to account for correlated responses: an introduction to generalized estimating equations and multi-level mixed modelling. Statistics in Medicine, 17(11):1261–1291, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahe:1998:LEA

[2164] Cédric Mahé. Letter to the Editor: Analysis of multiple failure time data from an AIDS clinical trial, by Dianne M. Finkelstein, David A. Schoenfeld and Eva Stamenovic, Statistics in Medicine, 16, 951–961 (1997). Statistics in Medicine, 17(11):1293–1294, June 15, 1998. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1837].

Dunn:1998:LEC

[2165] Graham Dunn. Letter to the Editor: Comparative calibration without a gold standard, by Ying Lu, Keying Ye, Ashwini K. Mathur, Siu Hui, Thomas P. Fuerst and Harry K. Genant, Statistics in Medicine, 16, 1889– 1905 (1997). Statistics in Medicine, 17(11):1294–1296, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1932].

Lu:1998:AR

[2166] Ying Lu, Keying Ye, and Ashwini K. Mathur. Authors’ reply. Statistics in Medicine, 17(11):1297–1298, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vonRosen:1998:BRM

[2167] Dietrich von Rosen. Book review: Modern Regression Methods. Thomas P. Ryan, Wiley, New York, 1997. No. of pages: xix + 515. Price: $71.50. ISBN 0-471-52912-5. Statistics in Medicine, 17(11):1299–1300, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schneider:1998:BRN

[2168] Berthold Schneider. Book review: Networks and Chaos — Statistical and Probabilistic Aspects. O. E. Barndorff-Nielsen, J. L. Jensen and W. S. Kendall (eds), Chapman and Hall, London, 1993. No. of pages: vii + 307. Price: £30. ISBN 0-412-46530-2. Statistics in Medicine, 17(11): 1300–1301, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:1998:BRM

[2169] Valen E. Johnson. Book review: Markov Chain Monte Carlo in Practice. W. R. Gilks, S. Richardson and D. J. Spiegelhalter (eds), Chapman and Hall, London, 1995. No. of pages: xv + 481. Price: £38. ISBN 0-412-05551-1. Statistics in Medicine, 17(11):1301–1302, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scheike:1998:BRS

[2170] Thomas H. Scheike. Book review: Smoothing Methods in Statistics. J. S. Simonoff, Springer, 1996. No. of pages: xii + 338. Price: DM84. ISBN 0-38794716-7. Statistics in Medicine, 17(11):1303, June 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:1998:APCa

[2171] Chris Robertson and Peter Boyle. Age-period-cohort analysis of chronic disease rates. i: modelling approach. Statistics in Medicine, 17(12):1305– 1323, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Robertson:1998:APCb

[2172] Chris Robertson and Peter Boyle. Age–period-cohort models of chronic disease rates. II: graphical approaches. Statistics in Medicine, 17(12): 1325–1340, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2473].

Little:1998:VTA

[2173] M. P. Little, F. De Vathaire, M. W. Charles, M. M. Hawkins, and C. R. Muirhead. Variations with time and age in the risks of solid cancer incidence after radiation exposure in childhood. Statistics in Medicine, 17(12):1341–1355, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roy:1998:URR

[2174] P. Roy and J. Esteve. Using relative risk models for estimating synergy between two risk factors. Statistics in Medicine, 17(12):1357–1373, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Devine:1998:ESS

[2175] Owen J. Devine and James M. Smith. Estimating sample size for epidemiologic studies: the impact of ignoring exposure measurement uncertainty. Statistics in Medicine, 17(12):1375–1389, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verbeke:1998:DRS

[2176] Geert Verbeke, Emmanuel Lesaffre, and Larry J. Brant. The detection of residual serial correlation in linear mixed models. Statistics in Medicine, 17(12):1391–1402, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chan:1998:ETE

[2177] Ivan S. F. Chan. Exact tests of equivalence and efficacy with a nonzero lower bound for comparative studies. Statistics in Medicine, 17(12): 1403–1413, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2474].

Wichmann:1998:BRF

[2178] H.-Erich Wichmann. Book review: Field Epidemiology. Michael B. Gregg (ed.), Oxford University Press, New York, 1996. No. of pages: xii + 275. Price: £29.50. ISBN 0-1950-7207-3. Statistics in Medicine, 17(12): 1415–1416, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hilden:1998:BRH

[2179] Jørgen Hilden. Book review: Handbook of Medical Informatics. J. H. van Bemmel and M. A. Musen (eds), Springer-Verlag, Heidelberg, 1997. No. of pages: 621 + XL. Price: DM 98. ISBN 3-540-63351-0. Statistics in Medicine, 17(12):1416–1417, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Verdooren:1998:BRM

[2180] L. R. Verdooren. Book review: Multiple Tests und Auswahlverfahren. Manfred Horn and Rudinger Vollandt, Gustav Fischer Verlag, Stuttgart, Jena, New York, 1995. No. of pages: 289. Price: DM 58. ISBN 3-437-20510-2. Statistics in Medicine, 17(12):1417–1418, June 30, 1998. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schemper:1998:BRF

[2181] M. Schemper. Book review: Fundamentals of Biostatistical Inference. Chap T. Le, Marcel Dekker Inc., New York, 1992. No. of pages: v + 254. Price: $55. ISBN 0-8247-8674-2. Statistics in Medicine, 17(12): 1418–1419, June 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Johnson:1998:EA

[2182] Tony Johnson and David Machin. Editorial announcement. Statistics in Medicine, 17(13):1421–1424, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newton:1998:SAA

[2183] Michael A. Newton, Michael N. Gould, Catherine A. Reznikoff, and Jill D. Haag. On the statistical analysis of allelic-loss data. Statistics in Medicine, 17(13):1425–1445, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zenios:1998:PTH

[2184] Stefanos A. Zenios and Lawrence M. Wein. Pooled testing for HIV prevalence estimation: exploiting the dilution effect. Statistics in Medicine, 17(13):1447–1467, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Liu:1998:DAS

[2185] Ping-Yu Liu, Wei Yann Tsai, and Michael Wolf. Design and analysis for survival data under order restrictions with a modified logrank test. Statistics in Medicine, 17(13):1469–1479, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1998:TSM

[2186] Paul S. Albert and Myron A. Waclawiw. A two-state Markov chain for heterogeneous transitional data: a quasi-likelihood approach. Statistics in Medicine, 17(13):1481–1493, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Obuchowski:1998:CCP

[2187] Nancy A. Obuchowski. On the comparison of correlated proportions for clustered data. Statistics in Medicine, 17(13):1495–1507, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1998:NIE

[2188] Kung-Jong Lui. A note on interval estimation of the relative difference in data with matched pairs. Statistics in Medicine, 17(13):1509–1515, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Makuch:1998:PIM

[2189] Robert W. Makuch, Zhongxin Zhang, Peter A. Charpentier, and Sharon K. Inouye. Prospective individual matching: covariate balance and power in a comparative study. Statistics in Medicine, 17(13):1517– 1526, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schachter:1998:LEB

[2190] Julius Schachter. Letter to the Editor: Bias in the evaluation of DNAamplification tests for detecting Chlamydia trachomatis, by A. Hadgu, Statistics in Medicine, 16, 1391–1399 (1997). Statistics in Medicine, 17 (13):1527–1528, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1884, 2145].

Hadgu:1998:ARb

[2191] Alula Hadgu. Author’s reply. Statistics in Medicine, 17(13):1528–1530, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Smith:1998:BRV

[2192] Daniel Smith. Book review: Visual Explanations: Images and Quantities, Evidence and Narrative. Edward R. Tufte, Graphics Press, Cheshire CT, 1997. No. of pages: 156. Price: £45. ISBN 0-961-3921-2-6. Statistics in Medicine, 17(13):1531–1532, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Waller:1998:BRM

[2193] Lance A. Waller. Book review: Methods for investigating localized clustering of disease. F. E. Alexander and P. Boyle (eds), International Agency for Research on Cancer Scientific Publication No. 135, Lyon, France. 1996. No. of pages: xvii + 247. Price: £45. ISBN 92-832-2135-4. Statistics in Medicine, 17(13):1532–1534, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuha:1998:BRR

[2194] Jouni Kuha. Book review: Regression Models for Categorical and Limited Dependent Variables. J. Scott Long, Sage Publications, U.S.A., 1997. No. of pages: xxx + 297. Price: £37.00. ISBN 0-8039-7374-8. Statistics in Medicine, 17(13):1534–1535, July 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahn:1998:EPC

[2195] Chul Ahn. An evaluation of phase I cancer clinical trial designs. Statistics in Medicine, 17(14):1537–1549, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fisher:1998:SDC

[2196] Lloyd D. Fisher. Self-designing clinical trials. Statistics in Medicine, 17(14):1551–1562, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thall:1998:SEA

[2197] Peter F. Thall and Hsi-Guang Sung. Some extensions and applications of a Bayesian strategy for monitoring multiple outcomes in clinical trials. Statistics in Medicine, 17(14):1563–1580, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Murray:1998:ADG

[2198] David M. Murray, Peter J. Hannan, Russell D. Wolfinger, William L. Baker, and James H. Dwyer. Analysis of data from group-randomized trials with repeat observations on the same groups. Statistics in Medicine, 17(14):1581–1600, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1998:EEE

[2199] Haibo Zhou and Clarice R. Weinberg. Evaluating effects of exposures on embryo viability and uterine receptivity in in vitro fertilization. Statistics in Medicine, 17(14):1601–1612, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmer:1998:BOD

[2200] J. Lynn Palmer and Peter Müller. Bayesian optimal design in population models for haematologic data. Statistics in Medicine, 17(14):1613–1622, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsieh:1998:SMS

[2201] F. Y. Hsieh, Daniel A. Bloch, and Michael D. Larsen. A simple method of sample size calculation for linear and logistic regression. Statistics in Medicine, 17(14):1623–1634, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Julious:1998:SSC

[2202] S. A. Julious and M. J. Campbell. Sample size calculations for paired or matched ordinal data. Statistics in Medicine, 17(14):1635–1642, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rochon:1998:AGP

[2203] James Rochon. Application of GEE procedures for sample size calculations in repeated measures experiments. Statistics in Medicine, 17(14): 1643–1658, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwards:1998:BRL

[2204] A. W. F. Edwards. Book review: Leading Personalities in Statistical Sciences from the Seventeenth Century to the Present. N. L. Johnson and S. Kotz, (eds), Wiley, New York, 1997. No. of pages: xxiii + 399. Price: £39.95. ISBN 0-4711-6381-3. Statistics in Medicine, 17(14):1659, July 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1998:P

[2205] Byron Jones. Preface. Statistics in Medicine, 17(15-16):1661, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:1998:SSL

[2206] John A. Lewis and Karen M. Facey. Statistical shortcomings in licensing applications. Statistics in Medicine, 17(15-16):1663–1673, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koch:1998:MAP

[2207] Gary G. Koch, Sonia M. Davis, and Randy L. Anderson. Methodological advances and plans for improving regulatory success for confirmatory studies. Statistics in Medicine, 17(15-16):1675–1690, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ebbutt:1998:PIE

[2208] A. F. Ebbutt and L. Frith. Practical issues in equivalence trials. Statistics in Medicine, 17(15-16):1691–1701, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rohmel:1998:TEI

[2209] Joachim Röhmel. Therapeutic equivalence investigations: statistical considerations. Statistics in Medicine, 17(15-16):1703–1714, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Grieve:1998:ISP

[2210] A. P. Grieve. Issues for statisticians in pharmaco-economic evaluations. Statistics in Medicine, 17(15-16):1715–1723, August 15–30, 1998. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hunt:1998:FAA

[2211] C. Anthony Hunt, Serge Guzy, and Daniel L. Weiner. A forecasting approach to accelerate drug development. Statistics in Medicine, 17(15-16):1725–1740, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Seldrup:1998:D

[2212] Jorgen Seldrup. Discussion. Statistics in Medicine, 17(15-16):1741–1743, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1998:CTE

[2213] J. K. Lindsey. Counts and times to events. Statistics in Medicine, 17(15-16):1745–1751, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1998:SCP

[2214] Stephen Senn. Some controversies in planning and analysing multi-centre trials. Statistics in Medicine, 17(15-16):1753–1765, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1998:CVE

[2215] B. Jones, D. Teather, J. Wang, and J. A. Lewis. A comparison of various estimators of a treatment difference for a multi-centre clinical trial. Statistics in Medicine, 17(15-16):1767–1777, August 15–30, 1998. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:1998:MCT

[2216] A. Lawrence Gould. Multi-centre trial analysis revisited. Statistics in Medicine, 17(15-16):1779–1797, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freeman:1998:D

[2217] Peter R. Freeman. Discussion. Statistics in Medicine, 17(15-16):1799– 1800, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Facey:1998:MIA

[2218] Karen M. Facey and John A. Lewis. The management of interim analyses in drug development. Statistics in Medicine, 17(15-16):1801–1809, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2506].

Whitehead:1998:D

[2219] John Whitehead. Discussion. Statistics in Medicine, 17(15-16):1811– 1812, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ruberg:1998:PEE

[2220] Stephen Ruberg and Victoria Cairns. Providing evidence of efficacy for a new drug. Statistics in Medicine, 17(15-16):1813–1823, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gould:1998:D

[2221] A. Lawrence Gould. Discussion. Statistics in Medicine, 17(15-16):1825– 1827, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Huster:1998:FEC

[2222] William J. Huster and Gregory G. Enas. A framework establishing clear decision criteria for the assessment of drug efficacy. Statistics in Medicine, 17(15-16):1829–1838, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ashby:1998:SIP

[2223] Deborah Ashby, Rosalind L. Smyth, and Philip J. Brown. Statistical issues in pharmacoepidemiological case-control studies. Statistics in Medicine, 17(15-16):1839–1850, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

ONeill:1998:BCP

[2224] Robert T. O’Neill. Biostatistical considerations in pharmacovigilance and pharmacoepidemiology: linking quantitative risk assessment in premarket licensure application safety data, post-market alert reports and formal epidemiological studies. Statistics in Medicine, 17(15-16):1851– 1858, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Evans:1998:D

[2225] Stephen J. W. Evans. Discussion. Statistics in Medicine, 17(15-16):1859– 1862, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koch:1998:ICA

[2226] Gary G. Koch, Catherine M. Tangen, Jin-Whan Jung, and Ingrid A. Amara. Issues for covariance analysis of dichotomous and ordered categorical data from randomized clinical trials and non-parametric strategies for addressing them. Statistics in Medicine, 17(15-16):1863–1892, August 15–30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsuyama:1998:BHS

[2227] Yutaka Matsuyama, Junichi Sakamoto, and Yasuo Ohashi. A Bayesian hierarchical survival model for the institutional effects in a multi-centre cancer clinical trial. Statistics in Medicine, 17(17):1893–1908, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tsodikov:1998:RBO

[2228] A. Tsodikov, D. Hasenclever, and M. Loeffler. Regression with bounded outcome score: evaluation of power by bootstrap and simulation in a chronic myelogenous leukaemia clinical trial. Statistics in Medicine, 17 (17):1909–1922, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schmid:1998:ESE

[2229] Christopher H. Schmid, Joseph Lau, Martin W. McIntosh, and Joseph C. Cappelleri. An empirical study of the effect of the control rate as a predictor of treatment efficacy in meta-analysis of clinical trials. Statistics in Medicine, 17(17):1923–1942, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Crown:1998:ASS

[2230] William H. Crown, Robert L. Obenchain, Luella Englehart, Tamra Lair, Don P. Buesching, and Thomas Croghan. The application of sample selection models to outcomes research: the case of evaluating the effects of antidepressant therapy on resource utilization. Statistics in Medicine, 17(17):1943–1958, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tosteson:1998:CME

[2231] Tor D. Tosteson, John P. Buonaccorsi, and Eugene Demidenko. Covariate measurement error and the estimation of random effect parameters in a mixed model for longitudinal data. Statistics in Medicine, 17(17):1959– 1971, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Commenges:1998:MAS

[2232] Daniel Commenges, Luc Letenneur, Pierre Joly, Ahmadou Alioum, and Jean-François Dartigues. Modelling age-specific risk: application to dementia. Statistics in Medicine, 17(17):1973–1988, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wade:1998:ICB

[2233] A. M. Wade and A. E. Ades. Incorporating correlations between measurements into the estimation of age-related reference ranges. Statistics in Medicine, 17(17):1989–2002, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fung:1998:SSC

[2234] W. K. Fung and T. S. Tsang. A simulation study comparing tests for the equality of coefficients of variation. Statistics in Medicine, 17(17): 2003–2014, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

May:1998:C

[2235] Warren L. May and William D. Johnson. Correction. Statistics in Medicine, 17(17):2015, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:1998:C

[2236] Guohua Pan. Correction. Statistics in Medicine, 17(17):2015–2016, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1998:LEC

[2237] Kung-Jong Lui. Letter to the Editor: Confidence intervals for differences in correlated binary proportions, by Warren L. May and William D. Johnson, Statistics in Medicine, 16, 2127–2136 (1997). Statistics in Medicine, 17(17):2017–2020, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1959, 2332, 2635].

May:1998:AR

[2238] Warren L. May and William D. Johnson. Authors’ reply. Statistics in Medicine, 17(17):2020–2021, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Svensson:1998:BRU

[2239] Elisabeth Svensson. Book review: Understanding Statistics: an Introduction For The Social Sciences. Daniel B. Wright, Sage Publications, London, 1997. No. of pages: 228. Price: £12.95. ISBN 0-8039-7918-5. Statistics in Medicine, 17(17):2023–2024, September 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xia:1998:STM

[2240] Hong Xia and Bradley P. Carlin. Spatio-temporal models with errors in covariates: mapping Ohio lung cancer mortality. Statistics in Medicine, 17(18):2025–2043, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Knorr-Held:1998:MRD

[2241] Leonhard Knorr-Held and Julian Besag. Modelling risk from a disease in time and space. Statistics in Medicine, 17(18):2045–2060, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bycott:1998:CST

[2242] Paul Bycott and Jeremy Taylor. A comparison of smoothing techniques for CD4 data measured with error in a time-dependent Cox proportional hazards model. Statistics in Medicine, 17(18):2061–2077, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghani:1998:SBM

[2243] A. C. Ghani, C. A. Donnelly, and G. P. Garnett. Sampling biases and missing data in explorations of sexual partner networks for the spread of sexually transmitted diseases. Statistics in Medicine, 17(18):2079– 2097, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Figueiras:1998:RMC

[2244] Adolfo Figueiras, Jose Maria Domenech-Massons, and Carmen Cadarso. Regression models: calculating the confidence interval of effects in the presence of interactions. Statistics in Medicine, 17(18):2099–2105, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Westfall:1998:UPI

[2245] Peter H. Westfall, Alok Krishen, and S. Stanley Young. Using prior information to allocate significance levels for multiple endpoints. Statistics in Medicine, 17(18):2107–2119, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brookmeyer:1998:PTA

[2246] Ron Brookmeyer and Ying-Qing Chen. Person-time analysis of paired community intervention trials when the number of communities is small. Statistics in Medicine, 17(18):2121–2132, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:1998:TSM

[2247] P. Bauer, J. Röhmel, W. Maurer, and L. Hothorn. Testing strategies in multi-dose experiments including active control. Statistics in Medicine, 17(18):2133–2146, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goodman:1998:LEB

[2248] Steven Goodman and Haydar Sengul. Letter to the Editor: Bayesian analysis for a single 2 × 2 table, by L. Hashemi, B. Nandram and R. Goldberg, Statistics in Medicine, 16, 1311–1328 (1997). Statistics in Medicine, 17(18):2147–2148, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1879].

Hashemi:1998:AR

[2249] Lobat Hashemi. Author’s reply. Statistics in Medicine, 17(18):2148, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1998:LEM

[2250] Simon G. Thompson. Letter to the Editor: The merits of matching in community intervention trials: a cautionary tale, by N. Klar and A. Donner, Statistics in Medicine, 16, 1753–1764 (1997). Statistics in Medicine, 17(18):2149–2151, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1919].

Klar:1998:AR

[2251] Neil Klar and Allan Donner. Authors’ reply. Statistics in Medicine, 17 (18):2151–2152, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1998:BRP

[2252] Peter Armitage. Book review: The Progress of Experiment: Science and Therapeutic Reform in the United States, 1900–1990. Harry M. Marks, Cambridge University Press, Cambridge, 1997. No. of pages: xi + 258. Price: £45. ISBN 0-521-58142-7. Statistics in Medicine, 17(18):2153– 2154, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersson:1998:BRM

[2253] Steen Arne Andersson. Book review: Multivariate Models and Dependence Concepts. Harry Joe, Chapman and Hall, 1997. No. of pages: xiii + 399. Price: £39. ISBN 0-412-07331-5. Statistics in Medicine, 17(18): 2154–2155, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

English:1998:BRW

[2254] Dallas English and Sascha Dublin. Book review: Workbook of Epidemiology. Staffan E. Norell, Oxford University Press, New York, 1995. No. of pages: 317. Price: $49.95 (cloth), $26.50 (paper). ISBN 0-19-507490-4. Statistics in Medicine, 17(18):2155–2156, September 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reeves:1998:SAM

[2255] G. K. Reeves, D. R. Cox, S. C. Darby, and E. Whitley. Some aspects of measurement error in explanatory variables for continuous and binary regression models. Statistics in Medicine, 17(19):2157–2177, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bull:1998:RMM

[2256] Shelley B. Bull. Regression models for multiple outcomes in large epidemiologic studies. Statistics in Medicine, 17(19):2179–2197, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1998:DTA

[2257] Andy H. Lee, Jianguo Xiao, Siva Ram Vemuri, and Yuejen Zhao. A discordancy test approach to identify outliers of length of hospital stay. Statistics in Medicine, 17(19):2199–2206, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lachenbruch:1998:AST

[2258] Peter A. Lachenbruch and Cornelius J. Lynch. Assessing screening tests: extensions of McNemar’s test. Statistics in Medicine, 17(19):2207–2217, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Baker:1998:PTD

[2259] Stuart G. Baker, Robert J. Connor, and Larry G. Kessler. The partial testing design: a less costly way to test equivalence for sensitivity and specificity. Statistics in Medicine, 17(19):2219–2232, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Joffe:1998:RMR

[2260] Marshall M. Joffe and Graham A. Colditz. Restriction as a method for reducing bias in the estimation of direct effects. Statistics in Medicine, 17 (19):2233–2249, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1998:ELN

[2261] Xiao-Hua Zhou. Estimation of the log-normal mean. Statistics in Medicine, 17(19):2251–2264, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:1998:PSM

[2262] Ralph B. D’Agostino, Jr. Propensity score methods for bias reduction in the comparison of a treatment to a non-randomized control group. Statistics in Medicine, 17(19):2265–2281, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Guilbaud:1998:BRS

[2263] Olivier Guilbaud. Book review: Statistical Issues in Drug Development. Stephen Senn, Wiley, Chichester, 1997. No. of pages: xvii + 423. Price: £45. ISBN 0-4719-7488-9. Statistics in Medicine, 17(19):2283–2285, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reilly:1998:BRP

[2264] Marie Reilly. Book review: Primer of Biostatistics. Fourth Edition, Stanton A. Glantz, McGraw Hill, 1997. No. of pages: 455. Price: £51.95. ISBN 0-0702-4268-2. Statistics in Medicine, 17(19):2285–2287, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:1998:BRN

[2265] Ian R. White. Book review: Nonparametric Statistics for Health Care Research. Statistics for Small Samples and Unusual Distributions. Marjorie A. Pett, Page Publications. Thousand Oaks, 1997. No. of pages: 305. Price: £22. ISBN 0-8039-7039-0. Statistics in Medicine, 17(19):2287– 2288, October 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parker:1998:CAS

[2266] Robert A. Parker and Nancy G. Berman. Criteria for authorship for statisticians in medical papers. Statistics in Medicine, 17(20):2289–2299, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1998:OFD

[2267] T. Timothy Chen and Tie-Hua Ng. Optimal flexible designs in phase II clinical trials. Statistics in Medicine, 17(20):2301–2312, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Girard:1998:MME

[2268] Pascal Girard, Terrence F. Blaschke, Helen Kastrissios, and Lewis B. Sheiner. A Markov mixed effect regression model for drug compliance. Statistics in Medicine, 17(20):2313–2333, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Golm:1998:SPM

[2269] Gregory T. Golm, M. Elizabeth Halloran, and Ira M. Longini, Jr. Semiparametric models for mismeasured exposure information in vaccine trials. Statistics in Medicine, 17(20):2335–2352, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berry:1998:UTH

[2270] Scott M. Berry. Understanding and testing for heterogeneity across 2×2 tables: application to meta-analysis. Statistics in Medicine, 17(20):2353– 2369, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hannan:1998:AFT

[2271] Peter J. Hannan, Xiao O. Shu, Daniel Weisdorf, and Anne Goldman. Analysis of failure times for multiple infections following bone marrow transplantation: an application of the multiple failure time proportional hazards model. Statistics in Medicine, 17(20):2371–2380, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Taylor:1998:DCS

[2272] Jeremy M. G. Taylor and Ngayee Law. Does the covariance structure matter in longitudinal modelling for the prediction of future CD4 counts? Statistics in Medicine, 17(20):2381–2394, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zilber:1998:SMQ

[2273] N. Zilber, O. Manor, R. Inzelberg, E. Kahana, and A. D. Korczyn. A statistical method for quantitative evaluation of the progression of chronic diseases: the mean score graph (MSG). Statistics in Medicine, 17(20): 2395–2403, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yarnold:1998:PHM

[2274] Paul R. Yarnold, Robert C. Soltysik, Frank Lefevre, and Gary J. Martin. Predicting in-hospital mortality of patients receiving cardiopulmonary resuscitation: unit-weighted MultiODA for binary data. Statistics in Medicine, 17(20):2405–2414, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Eerola:1998:BRA

[2275] Mervi Eerola. Book review: The Art of Causal Conjecture. Glenn Shafer, MIT Press, Massachusetts, 1996. No. of pages: 511. ISBN 0-262-19368-X. Statistics in Medicine, 17(20):2415–2416, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bickeboller:1998:BRG

[2276] Heike Bickeböller. Book review: Genetic Mapping and DNA-Sequencing. Terry Speed and Michael S. Waterman (Eds), IMA Vol. 81, Springer, New York, 1996. No. of pages: 215. Price: $65.95. ISBN 0-387-94803-1. Statistics in Medicine, 17(20):2416–2418, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bjorner:1998:BRM

[2277] Jakob Bue Bjorner. Book reviews: Multivariate Analysis Techniques in Social Science Research. Jacques Tacq, Sage Publications, London, 1997. No. of pages: 411. Price: $29.95 (paperback). ISBN 0-7619-5273-X. Price: $79.95 (hardcover), ISBN 0-7619-5272-1. Statistics in Medicine, 17 (20):2418–2419, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kala:1998:BRQ

[2278] Radosław Kala. Book review: QUECUS: Statistics for Bioscientists. A Student Guidebook. Moya McCloskey, Steve Blythe and Chris Robertson, Arnold, U.K., 1997. No. of pages: 160. No. of discs: 3. QUERCUS consists of software that can be downloaded free to all educational estab-

lishments form http://www.stams.strath.ac.uk/external/quercus

as well as a ‘A student guidebook’, list price £12.99. ISBN 0-3406-768-6. Statistics in Medicine, 17(20):2419–2420, October 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). URL http:

//www.stams.strath.ac.uk/external/quercus.

Marschner:1998:DHV

[2279] Ian C. Marschner. Design of HIV viral dynamics studies. Statistics in Medicine, 17(21):2421–2434, November 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1998:SIH

[2280] Jeffrey M. Albert, John P. A. Ioannidis, Patricia Reichelderfer, Brian Conway, Robert W. Coombs, Lawrence Crane, Ralph Demasi, Dennis O. Dixon, Phillipe Flandre, Michael D. Hughes, Leslie A. Kalish, Kinley Larntz, Danyu Lin, Ian C. Marschner, Alvaro Muñoz, Jeff Murray, Jim Neaton, Carla Pettinelli, Wasima Rida, Jeremy M. G. Taylor, and Seth L. Welles. Statistical issues for HIV surrogate endpoints: point/counterpoint. Statistics in Medicine, 17(21):2435–2462, November 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wu:1998:EHD

[2281] Hulin Wu, A. Adam Ding, and Victor De Gruttola. Estimation of HIV dynamic parameters. Statistics in Medicine, 17(21):2463–2485, November 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gibbons:1998:ECP

[2282] Robert D. Gibbons and John V. Lavigne. Emergence of childhood psychiatric disorders: a multivariate probit analysis. Statistics in Medicine, 17(21):2487–2499, November 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ennis:1998:CSL

[2283] Marguerite Ennis, Geoffrey Hinton, David Naylor, Mike Revow, and Robert Tibshirani. A comparison of statistical learning methods on the GUSTO database. Statistics in Medicine, 17(21):2501–2508, November 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cronin:1998:AUM

[2284] Kathleen A. Cronin, Julie M. Legler, and Ruth D. Etzioni. Assessing uncertainty in microsimulation modelling with application to cancer screening interventions. Statistics in Medicine, 17(21):2509–2523, November 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ahnn:1998:SSD

[2285] Sang Ahnn and Stewart J. Anderson. Sample size determination in complex clinical trials comparing more than two groups for survival endpoints. Statistics in Medicine, 17(21):2525–2534, November 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hatzinger:1998:BRH

[2286] Reinhold Hatzinger. Book review: Handbook of Modern Item Response Theory. W. J. Van der Linden and R. K. Hambleton, Springer, New York, 1997. No. of pages: xv + 510. Price: DM 108. ISBN 0-3879-4661-6. Statistics in Medicine, 17(21):2535–2536, November 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Berkey:1998:MAM

[2287] C. S. Berkey, D. C. Hoaglin, A. Antczak-Bouckoms, F. Mosteller, and G. A. Colditz. Meta-analysis of multiple outcomes by regression with random effects. Statistics in Medicine, 17(22):2537–2550, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Beacon:1998:ACP

[2288] Heather J. Beacon, Simon G. Thompson, and Peter D. England. The analysis of complex patterns of longitudinal binary response: an example of transient dysphagia following radiotherapy. Statistics in Medicine, 17 (22):2551–2561, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Emir:1998:ARM

[2289] Birol Emir, Sam Wieand, John Q. Su, and Steve Cha. Analysis of repeated markers used to predict progression of cancer. Statistics in Medicine, 17(22):2563–2578, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kleinman:1998:SPB

[2290] Ken P. Kleinman and Joseph G. Ibrahim. A semi-parametric Bayesian approach to generalized linear mixed models. Statistics in Medicine, 17 (22):2579–2596, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yanez:1998:EME

[2291] N. David Yanez III, Richard A. Kronmal, and Lynn R. Shemanski. The effects of measurement error in response variables and tests of association of explanatory variables in change models. Statistics in Medicine, 17(22): 2597–2606, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yang:1998:OMM

[2292] Mark C. K. Yang, Yeanok Y. Namgung, Taewoon Kang, Ingvar Magnusson, and Arthur Hefti. Option-3 median measurement scheme in outlier reduction. Statistics in Medicine, 17(22):2607–2615, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haber:1998:EIP

[2293] Michael Haber and Andre Ndikuyeze. Estimation of individual and population vaccination effectiveness from time-to-event data. Statistics in Medicine, 17(22):2617–2623, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coory:1998:NMR

[2294] Michael Coory and Robert Gibberd. New measures for reporting the magnitude of small-area variation in rates. Statistics in Medicine, 17 (22):2625–2634, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:1998:ICI

[2295] Robert G. Newcombe. Improved confidence intervals for the difference between binomial proportions based on paired data. Statistics in Medicine, 17(22):2635–2650, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2633].

Haines:1998:BRE

[2296] Linda M. Haines. Book review: Experiment Design Made Easy. Annotated Instructor Notes. Instructors from Stat-Ease, STAT-EASE, Minneapolis, 1996. No. of Pages: 480. Price: $50. Statistics in Medicine, 17 (22):2651–2652, November 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:1998:I

[2297] John Matthews and Harvey Goldstein. Introduction. Statistics in Medicine, 17(23):2653–2654, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tanner:1998:F

[2298] James Mourilyan Tanner. Foreword. Statistics in Medicine, 17(23):2655– 2659, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1998:SRM

[2299] Douglas G. Altman. Statistical reviewing for medical journals. Statistics in Medicine, 17(23):2661–2674, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1998:ACT

[2300] Peter Armitage. Attitudes in clinical trials. Statistics in Medicine, 17 (23):2675–2683, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Campbell:1998:CRE

[2301] N. A. Campbell, H. P. Lopuhaä, and P. J. Rousseeuw. On the calculation of a robust S-estimator of a covariance matrix. Statistics in Medicine, 17 (23):2685–2695, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cole:1998:PIG

[2302] T. J. Cole. Presenting information on growth distance and conditional velocity in one chart: practical issues of chart design. Statistics in Medicine, 17(23):2697–2707, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gower:1998:RCD

[2303] J. C. Gower. The role of constraints in determining optimal scores. Statistics in Medicine, 17(23):2709–2721, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kenward:1998:SMR

[2304] Michael G. Kenward. Selection models for repeated measurements with non-random dropout: an illustration of sensitivity. Statistics in Medicine, 17(23):2723–2732, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matthews:1998:ACO

[2305] J. N. S. Matthews. Alternative criteria for optimal designs of limiting dilution assays. Statistics in Medicine, 17(23):2733–2746, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nelder:1998:LCM

[2306] John A. Nelder. A large class of models derived from generalized linear models. Statistics in Medicine, 17(23):2747–2753, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:1998:MLR

[2307] Huiqi Pan and Harvey Goldstein. Multi-level repeated measures growth modelling using extended spline functions. Statistics in Medicine, 17 (23):2755–2770, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:1998:ETC

[2308] S. B. Tan and A. F. M. Smith. Exploratory thoughts on clinical trials with utilities. Statistics in Medicine, 17(23):2771–2791, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1998:BPM

[2309] Anonymous. Bibliography: Publications of Michael Healy up to May 1998. Statistics in Medicine, 17(23):2793–2802, December 15, 1998. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1998:BRS

[2310] Peter Armitage. Book review: Statistics and Truth: Putting Chance to Work. C. Radhakrishna Rao, World Scientific, Singapore, 2nd edn, 1997. No. of pages: xvii + 192. Price: £25. ISBN 9-8102-3111-3. Statistics in Medicine, 17(23):2803–2804, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osterlind:1998:BRI

[2311] Kell Østerlind. Book review: Intuitive Biostatistics. Harvey Motulsky, Oxford University Press, New York, 1995. No. of pages: xxi + 386. Price: $26. ISBN 0-1950-8607-4. Statistics in Medicine, 17(23):2804–2805, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brunner:1998:BRB

[2312] Edgar Brunner. Book review: Biostatistics. A Methodology for the Health Sciences. L. D. Fisher and G. van Belle, Wiley, New York, 1993. No. of pages: xxii + 991. Price: £37.50 (Paperback). ISBN 0-4711-6609-X. Statistics in Medicine, 17(23):2805–2806, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leon:1998:BRM

[2313] Andrew C. Leon. Book review: Mathematics: From the Birth of Numbers, Jan Gullberg, W. W. Norton and Company, New York, 1997. No. of pages: 1093. Price: $50.00. ISBN 0-393-04002-X. Statistics in Medicine, 17(23):2806–2807, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dalgaard:1998:BRSb

[2314] Peter Dalgaard. Book review: Statistical Tools for Nonlinear Regression. A Practical Guide with S-Plus examples. Sylvie Huet, Annie Bouvier, Marie-Anne Gruet and Emmanuel Jolivet, Springer-Verlag, New York, 1996. No. of pages: ix + 154. Price: £28. ISBN 0-387-94727-2. Statistics in Medicine, 17(23):2807, December 15, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Vail:1998:EBU

[2315] Andy Vail. Experiences of a biostatistician on a U.K. Research Ethics Committee. Statistics in Medicine, 17(24):2811–2814, December 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parmar:1998:ESS

[2316] Mahesh K. B. Parmar, Valter Torri, and Lesley Stewart. Extracting summary statistics to perform meta-analyses of the published literature for survival endpoints. Statistics in Medicine, 17(24):2815–2834, December 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bolland:1998:SSR

[2317] K. Bolland, M. R. Sooriyarachchi, and J. Whitehead. Sample size review in a head injury trial with ordered categorical responses. Statistics in Medicine, 17(24):2835–2847, December 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Senn:1998:RRA

[2318] Stephen Senn and Dimitrios Lambrou. Robust and realistic approaches to carry-over. Statistics in Medicine, 17(24):2849–2864, December 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Houwing-Duistermaat:1998:IFH

[2319] Jeanine J. Houwing-Duistermaat and Hans C. Van Houwelingen. Incorporation of family history in logistic regression models. Statistics in Medicine, 17(24):2865–2882, December 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Boutitie:1998:ATT

[2320] F. Boutitie, F. Gueyffier, S. J. Pocock, and J. P. Boissel. Assessing treatment-time interaction in clinical trials with time to event data: a meta-analysis of hypertension trials. Statistics in Medicine, 17(24):2883– 2903, December 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1998:GSL

[2321] Isabelle Albert and Jean-Philippe Jais. Gibbs sampler for the logistic model in the analysis of longitudinal binary data. Statistics in Medicine, 17(24):2905–2921, December 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Svensson:1998:OIM

[2322] Elisabeth Svensson. Ordinal invariant measures for individual and group changes in ordered categorical data. Statistics in Medicine, 17(24):2923– 2936, December 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Houwing-Duistermaat:1998:MCD

[2323] Jeanine J. Houwing-Duistermaat, Hans C. Van Houwelingen, and Annemarie Terhell. Modelling the cause of dependency with application to filaria infection. Statistics in Medicine, 17(24):2939–2954, December 30, 1998. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenland:1999:UAA

[2324] Sander Greenland. A unified approach to the analysis of case-distribution (case-only) studies. Statistics in Medicine, 18(1):1–15, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:1999:DTA

[2325] M. P. Little, C. R. Muirhead, and M. W. Charles. Describing time and age variations in the risk of radiation-induced solid tumour incidence in the Japanese atomic bomb survivors using generalized relative and absolute risk models. Statistics in Medicine, 18(1):17–33, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:1999:CIB

[2326] Robert H. Lyles and Jane Xu. Classifying individuals based on predictors of random effects. Statistics in Medicine, 18(1):35–52, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longini:1999:MMM

[2327] Ira M. Longini, Jr., Michael G. Hudgens, M. Elizabeth Halloran, and Karen Sagatelian. A Markov model for measuring vaccine efficacy for both susceptibility to infection and reduction in infectiousness for prophylactic HIV vaccines. Statistics in Medicine, 18(1):53–68, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Steiner:1999:MPB

[2328] Stefan H. Steiner, Richard J. Cook, and Vern T. Farewell. Monitoring paired binary surgical outcomes using cumulative sum charts. Statistics in Medicine, 18(1):69–86, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schouten:1999:SSF

[2329] Hubert J. A. Schouten. Sample size formula with a continuous outcome for unequal group sizes and unequal variances. Statistics in Medicine, 18(1):87–91, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hauschke:1999:SSD

[2330] Dieter Hauschke, Meinhard Kieser, Edgar Diletti, and Martin Burke. Sample size determination for proving equivalence based on the ratio of two means for normally distributed data. Statistics in Medicine, 18(1): 93–105, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Flandre:1999:LEE

[2331] Philippe Flandre and Yacine Saidi. Letter to the Editor: Estimating the proportion of treatment effect explained, by a surrogate marker by D. Y. Lin, T. R. Fleming and V. De Gruttola, Statistics in Medicine, 16, 1515– 1527 (1997). Statistics in Medicine, 18(1):107–109, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1894].

Broemeling:1999:LEC

[2332] Lyle D. Broemeling. Letter to the Editor: Confidence intervals for differences in correlated binary proportions, by W. L. May and W. D. Johnson, Statistics in Medicine, 16, 2127–2136 (1997). Statistics in Medicine, 18 (1):109, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1959, 2237, 2635].

vanHouwelingen:1999:LEI

[2333] Hans van Houwelingen and Stephen Senn. Letter to the Editor: Investigating underlying risk as a source of heterogeneity in meta-analysis, by S. G. Thompson, T. C. Smith and S. J. Sharp, Statistics in Medicine, 16, 2741–2758 (1997). Statistics in Medicine, 18(1):110–115, January 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2017].

Landrum:1999:ABI

[2334] Mary Beth Landrum and Sharon-Lise T. Normand. Applying Bayesian ideas to the development of medical guidelines. Statistics in Medicine, 18(2):117–137, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rose:1999:EBU

[2335] M. S. Rose, A. M. Gillis, and R. S. Sheldon. Evaluation of the bias in using the time to the first event when the inter-event intervals have a Weibull distribution. Statistics in Medicine, 18(2):139–154, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xu:1999:EPL

[2336] Jian-Lun Xu, Richard M. Fagerstrom, and Philip C. Prorok. Estimation of post-lead-time survival under dependence between lead-time and post-lead-time survival. Statistics in Medicine, 18(2):155–162, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pepe:1999:ECU

[2337] Margaret Sullivan Pepe, Robert C. Whitaker, and Kristy Seidel. Estimating and comparing univariate associations with application to the prediction of adult obesity. Statistics in Medicine, 18(2):163–173, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cook:1999:TSM

[2338] Richard J. Cook, Edmund T. M. Ng, Jayanti Mukherjee, and David Vaughan. Two-state mixed renewal processes for chronic disease. Statistics in Medicine, 18(2):175–188, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Morton-Jones:1999:IEE

[2339] Tony Morton-Jones, Peter Diggle, and Paul Elliott. Investigation of excess environmental risk around putative sources: Stone’s test with covariate adjustment. Statistics in Medicine, 18(2):189–197, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Barnhart:1999:RMM

[2340] Huiman X. Barnhart, Andrzej S. Kosinski, and Allan R. Sampson. A regression model for multivariate random length data. Statistics in Medicine, 18(2):199–211, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Horton:1999:GFG

[2341] Nicholas J. Horton, Judith D. Bebchuk, Cheryl L. Jones, Stuart R. Lipsitz, Paul J. Catalano, Gwendolyn E. P. Zahner, and Garrett M. Fitzmaurice. Goodness-of-fit for GEE: an example with mental health service utilization. Statistics in Medicine, 18(2):213–222, January 30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mayo:1999:BLM

[2342] Matthew S. Mayo, Eustache Paramithiotis, and Max. D. Cooper. B lymphocyte migration to the bone marrow of humans is not random. Statistics in Medicine, 18(2):223–231, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bernsen:1999:LEV

[2343] R. M. D. Bernsen, M. J. A. Tasche, and N. J. D. Nagelkerke. Letter to the Editor: Variation in baseline risk as an explanation of heterogeneity in meta-analysis, by S. D. Walter, Statistics in Medicine, 16, 2883–2900 (1997). Statistics in Medicine, 18(2):233–237, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2031].

Walter:1999:AR

[2344] S. D. Walter. Author’s reply. Statistics in Medicine, 18(2):237–238, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacobsen:1999:BRC

[2345] Jacob M. Jacobsen. Book review: Computational Biology of the Heart. A. V. Panfilov and A. V. Holden (eds), Wiley, Chichester, 1997. No. of pages: xii + 416. Price: $120. ISBN 0-4719-6020-9. Statistics in Medicine, 18(2):239–240, January 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Meinert:1999:RTU

[2346] Curtis L. Meinert. Redesign of trials under different enrolment mixes. Statistics in Medicine, 18(3):241–251, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Geller:1999:CRT

[2347] Nancy L. Geller. Commentary on ‘Redesign of trials under different enrolment mixes’ by Curtis L. Meinert. Statistics in Medicine, 18(3): 253–254, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Schouten:1999:PGS

[2348] Hubert J. A. Schouten. Planning group sizes in clinical trials with a continuous outcome and repeated measures. Statistics in Medicine, 18 (3):255–264, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prescott:1999:CUB

[2349] Philip Prescott. Construction of uniform-balanced cross-over designs for any odd number of treatments. Statistics in Medicine, 18(3):265–272, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:1999:CSS

[2350] Michael P. Fay. Comparing several score tests for interval censored data. Statistics in Medicine, 18(3):273–285, February 15, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See correction [2559].

Albert:1999:CHM

[2351] James H. Albert. Criticism of a hierarchical model using Bayes factors. Statistics in Medicine, 18(3):287–305, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:1999:SPE

[2352] Nico Nagelkerke, Siem Heisterkamp, Martien Borgdorff, Jaap Broekmans, and Hans Van Houwelingen. Semi-parametric estimation of agetime specific infection incidence from serial prevalence data. Statistics in Medicine, 18(3):307–320, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Normand:1999:MAF

[2353] Sharon-Lise T. Normand. Meta-analysis: formulating, evaluating, combining, and reporting. Statistics in Medicine, 18(3):321–359, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2637].

vanderTweel:1999:LEM

[2354] Ingeborg van der Tweel and Paul A. H. van Noord. Letter to the Editor: A method for deciding early stopping of inconclusive case-control studies in settings where data are stratified, by U. Strömberg, Statistics in Medicine, 16, 2327–2337 (1997). Statistics in Medicine, 18(3):361– 362, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1978].

Stromberg:1999:AR

[2355] Ulf Strömberg. Author’s reply. Statistics in Medicine, 18(3):362–363, February 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lin:1999:ISA

[2356] Zhengning Lin. An issue of statistical analysis in controlled multi-centre studies: how shall we weight the centres? Statistics in Medicine, 18 (4):365–373, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ash:1999:UMM

[2357] Arlene Ash and Michael Shwartz. R^{2}: a useful measure of model performance when predicting a dichotomous outcome. Statistics in Medicine, 18(4):375–384, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2636].

Palta:1999:LVM

[2358] Mari Palta and Chin-Yu Lin. Latent variables, measurement error and methods for analysing longitudinal binary and ordinal data. Statistics in Medicine, 18(4):385–396, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fraser:1999:ELR

[2359] Gary E. Fraser and David J. Shavlik. The estimation of lifetime risk and average age at onset of a disease using a multivariate exponential hazard rate model. Statistics in Medicine, 18(4):397–410, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1999:PBE

[2360] Haibo Zhou and Clarice R. Weinberg. Potential for bias in estimating human fecundability parameters: a comparison of statistical models. Statistics in Medicine, 18(4):411–422, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Parrish:1999:MAA

[2361] Rudolph S. Parrish, Linda J. Hazlett, Kerry D. Bridges, and P. Jean Henslee-Downey. A multivariate approach for assessing severity of acute graft-versus-host disease in bone marrow transplantation. Statistics in Medicine, 18(4):423–440, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeAngelis:1999:MMC

[2362] R. De Angelis, R. Capocaccia, T. Hakulinen, B. Soderman, and A. Verdecchia. Mixture models for cancer survival analysis: application to population-based data with covariates. Statistics in Medicine, 18 (4):441–454, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1999:PAG

[2363] Wen-Chung Lee. Probabilistic analysis of global performances of diagnostic tests: interpreting the Lorenz curve-based summary measures. Statistics in Medicine, 18(4):455–471, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Leong:1999:UMD

[2364] Traci Leong, Stuart R. Lipsitz, and Joseph G. Ibrahim. Using missing data methods in genetic studies with missing mutation status. Statistics in Medicine, 18(4):473–485, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Altman:1999:SML

[2365] Douglas G. Altman. Statistics in the medical literature: 3. Statistics in Medicine, 18(4):487–490, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saporta:1999:BRA

[2366] Gilbert Saporta. Book review: The Analysis of Proximity Data. B. S. Everitt and S. Rabe-Hesketh, Arnold, London, 1997. No. of pages: ix + 178. Price: £35. ISBN 0-340-67776-7. Statistics in Medicine, 18(4):491– 492, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Borgan:1999:BRA

[2367] Ørnulf Borgan. Book review: Applied Survival Analysis. Chap T. Le, Wiley, New York, 1997. No. of pages: xiii + 257. Price: £34.95. ISBN 0-471-17085-2. Statistics in Medicine, 18(4):492–494, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Benner:1999:BRS

[2368] Axel Benner. Book review: Sample Size Tables For Clinical Studies. (2nd edn). David Machin, Michael J. Campbell, Peter M. Fayers and Alain P. Y. Pinol, Blackwell Science Ltd., Oxford, 1997. No. of pages: x + 315. Price: £45. ISBN 0-86542-870-0. Statistics in Medicine, 18(4):494– 495, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goldstein:1999:BRS

[2369] Richard Goldstein. Book review: Statistics Software Qualification: Reference Data Sets. B. P. Butler, M. G. Cox, S. L. R. Ellison and W. A. Hardcastle (eds), Royal Society of Chemistry, Cambridge, 1996. No. of pages: xvii + 317. Price: £99.50. ISBN 0-85404-422-1. Statistics in Medicine, 18(4):495–496, February 28, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armstrong:1999:GMH

[2370] Marc P. Armstrong, Gerard Rushton, and Dale L. Zimmerman. Geographically masking health data to preserve confidentiality. Statistics in Medicine, 18(5):497–525, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cui:1999:EAI

[2371] Jisheng Cui. Estimating AIDS incidence and jack-knife variance from a continuous delay distribution and incomplete data. Statistics in Medicine, 18(5):527–537, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feng:1999:ECL

[2372] Ziding Feng, Paula Diehr, Yutaka Yasui, Brent Evans, Shirley Beresford, and Thomas D. Koepsell. Explaining community-level variance in group randomized trials. Statistics in Medicine, 18(5):539–556, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1999:MCR

[2373] Xiao-Hua Zhou, E. J. Brizendine, and M. B. Pritz. Methods for combining rates from several studies. Statistics in Medicine, 18(5):557–566, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rubenstein:1999:EAC

[2374] Linda M. Rubenstein and Charles S. Davis. Estimation of the average correlation coefficient for stratified bivariate data. Statistics in Medicine, 18(5):567–580, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yafune:1999:BACa

[2375] Akifumi Yafune and Makio Ishiguro. Bootstrap approach for constructing confidence intervals for population pharmacokinetic parameters. i: a use of bootstrap standard error. Statistics in Medicine, 18(5):581–599, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yafune:1999:BACb

[2376] Akifumi Yafune and Makio Ishiguro. Bootstrap approach for constructing confidence intervals for population pharmacokinetic parameters. II: a bootstrap modification of standard two-stage (STS) method for phase I trial. Statistics in Medicine, 18(5):601–612, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kosorok:1999:ESC

[2377] Michael R. Kosorok and Roger Qu. Exact simultaneous confidence bands for a collection of univariate polynomials in regression analysis. Statistics in Medicine, 18(5):613–620, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Inskip:1999:BRP

[2378] Hazel Inskip. Book review: People, Environment, Disease and Death. G. Melvyn Howe, University of Wales Press, Cardiff, 1997. No. of pages: xx + 328. Price: £50. ISBN 0-7083-1373-6. Statistics in Medicine, 18(5): 621–622, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gotzsche:1999:BRP

[2379] Peter C. Gøtzsche. Book review: Probability Without Equations — Concepts For Clinicians. Bart K. Holland, Johns Hopkins University Press, 1998. No. of pages: 113. Price: $20.50 (paperback) $54.00 (hardcover). ISBN 0-8018-5759-7. Statistics in Medicine, 18(5):622–623, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McCullagh:1999:BRP

[2380] Peter McCullagh. Book review: Principles of Statistical Inference from a Neo-Fisherian Perspective. Luigi Pace and Alessandra Salvan, World Scientific Publishing Co., Singapore, 1997. No of pages: xvii + 556. Price: £47. ISBN 9-8102-3066-4. Statistics in Medicine, 18(5):623–624, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tiit:1999:BRC

[2381] Ene-Margit Tiit. Book review: Computer-Aided Multivariate Analysis. Third edition. A. A. Afifi and V. Clark, Chapman and Hall, U.K. 1996 (reprinted 1998). No. of pages: xxi + 455. Price: £45. ISBN 0-4127-3060-X. Statistics in Medicine, 18(5):624–625, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wild:1999:BRB

[2382] Pascal Wild. Book review: Bayes and Empirical Bayes Methods for Data Analysis. B. C. Carlin and T. A. Louis, Chapman and Hall, London, 1996, No. of pages: xvi + 399. Price: £30. ISBN 0-412-05611-9. Statistics in Medicine, 18(5):625–626, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sauerbrei:1999:BRM

[2383] Willi Sauerbrei. Book review: Management of Data in Clinical Trials. Eleanor McFadden, Wiley, 1998. No. of pages: xi + 210. Price: £50. ISBN 0-471-30316-X. Statistics in Medicine, 18(5):626–627, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scheike:1999:BRE

[2384] Thomas H. Scheike. Book review: Essential Wavelets for Statistical Applications and Data Analysis: R. Todd Ogden, Birkhäuser, 1997. No. of pages: xii + 198. Price: DM 78. ISBN 3-7643-3864-4. Statistics in Medicine, 18(5):627–628, March 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kallen:1999:DRS

[2385] Anders Källén and Per Larsson. Dose response studies: how do we make them conclusive? Statistics in Medicine, 18(6):629–641, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Platt:1999:GLM

[2386] Robert W. Platt, Brian G. Leroux, and Norman Breslow. Generalized linear mixed models for meta-analysis. Statistics in Medicine, 18(6): 643–654, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jacobs:1999:IAP

[2387] David R. Jacobs, Jr., Peter J. Hannan, Dennis Wallace, Kiang Liu, O. Dale Williams, and Cora E. Lewis. Interpreting age, period and cohort effects in plasma lipids and serum insulin using repeated measures regression analysis: the CARDIA study. Statistics in Medicine, 18(6): 655–679, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

vanBuuren:1999:MIM

[2388] S. van Buuren, H. C. Boshuizen, and D. L. Knook. Multiple imputation of missing blood pressure covariates in survival analysis. Statistics in Medicine, 18(6):681–694, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gooley:1999:EFP

[2389] Ted A. Gooley, Wendy Leisenring, John Crowley, and Barry E. Storer. Estimation of failure probabilities in the presence of competing risks: new representations of old estimators. Statistics in Medicine, 18(6):695–706, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osnes:1999:IRA

[2390] Kåre Osnes. Iterative random aggregation of small units using regional measures of spatial autocorrelation for cluster localization. Statistics in Medicine, 18(6):707–725, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shannon:1999:CCT

[2391] William D. Shannon and David Banks. Combining classification trees using MLE. Statistics in Medicine, 18(6):727–740, March 30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Elston:1999:SCG

[2392] Robert C. Elston, Ramana M. Idury, Lon R. Cardon, and Jay B. Lichter. The study of candidate genes in drug trials: sample size considerations. Statistics in Medicine, 18(6):741–751, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DerSimonian:1999:LEC

[2393] Rebecca DerSimonian. Letter to the Editor: Conference on meta-analysis in the design and monitoring of clinical trials, by N. Geller, L. Freedman, Y. J. Lee and R. DerSimonian (eds), Statistics in Medicine, 15, 1233– 1324 (1996). Statistics in Medicine, 18(6):753–754, March 30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bailey:1999:BRE

[2394] Norman Bailey. Book review: The Eradication of Infectious Diseases. W. R. Dowdle and D. R. Hopkins (eds.), Wiley, U.K., 1998. No. of pages: x + 222. Price: £60. ISBN 0-4719-8089-7. Statistics in Medicine, 18(6): 755–756, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Skovgaard:1999:BRB

[2395] Lene Theil Skovgaard. Book review: Basic Principles of Structural Equation Modeling. An Introduction to LISREL and EQS. R. O. Mueller, Springer for Science, Netherlands, 1996. No. of pages: xxviii + 229. Price: DM58. ISBN 0-387-94516-4. Statistics in Medicine, 18(6):757–758, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Glad:1999:BRI

[2396] Ingrid K. Glad. Book review: Introductory Statistics: a Modelling Approach. Jim K. Lindsey, Oxford University Press; Oxford, 1995. No. of pages: xi + 214. Price: £19.95. ISBN 0-19-852345-9. Statistics in Medicine, 18(6):759, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Axelsen:1999:BRS

[2397] Nils Axelsen. Book review: Scientific Deception: an Update. Lesley Grayson, The British Library, London, 1997. No. of pages: ix + 71. Price: £10. ISBN 0-7123-0841-5. Statistics in Medicine, 18(6):760, March 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coad:1999:CRP

[2398] D. S. Coad and William F. Rosenberger. A comparison of the randomized play-the-winner rule and the triangular test for clinical trials with binary responses. Statistics in Medicine, 18(7):761–769, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edwards:1999:MPC

[2399] David Edwards. On model prespecification in confirmatory randomized studies. Statistics in Medicine, 18(7):771–785, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Proschan:1999:MCP

[2400] Michael A. Proschan. A multiple comparison procedure for three- and four-armed controlled clinical trials. Statistics in Medicine, 18(7):787– 798, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letters [2641, 2640].

Lee:1999:ADD

[2401] Jack C. Lee, Dung-Tsa Chen, Hui-Nien Hung, and James J. Chen. Analysis of drug dissolution data. Statistics in Medicine, 18(7):799–814, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chouquet:1999:THI

[2402] Cécile Chouquet, Sylvia Richardson, Marianne Burgard, Stéphane Blanche, Marie-Jeanne Mayaux, Christine Rouzioux, and Dominique Costagliola. Timing of human immunodeficiency virus type 1 (HIV-1) transmission from mother to child: Bayesian estimation using a mixture. Statistics in Medicine, 18(7):815–833, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lesaffre:1999:AGF

[2403] Emmanuel Lesaffre, Makonnen Asefa, and Geert Verbeke. Assessing the goodness-of-fit of the Laird and Ware model — an example: the Jimma Infant Survival Differential Longitudinal Study. Statistics in Medicine, 18(7):835–854, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sullivan:1999:IHL

[2404] Lisa M. Sullivan, Kimberly A. Dukes, and Elena Losina. An introduction to hierarchical linear modelling. Statistics in Medicine, 18(7):855–888, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1999:C

[2405] Xiao-Hua Zhou. Correction. Statistics in Medicine, 18(7):889, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1999:C

[2406] Jane C. Lindsey. Correction. Statistics in Medicine, 18(7):890, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longini:1999:C

[2407] Ira M. Longini, Jr. Correction. Statistics in Medicine, 18(7):890, April 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cohen:1999:BRN

[2408] Steven B. Cohen. Book review: Non-random Reflections on Health Services Research: On the 25th Anniversary of Archie Cochrane’s Effectiveness and Efficiency. Alan Maynard and Iain Chalmers (eds), BMJ Publishing Group, London, 1997. No. of pages: xv + 303. Price: $25. ISBN 0-7279-1151-1. Statistics in Medicine, 18(7):891–892, April 15, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reis:1999:EAT

[2409] Isildinha M. Reis, Karim F. Hirji, and Abdelmonem A. Afifi. Exact and asymptotic tests for homogeneity in several 2 × 2 tables. Statistics in Medicine, 18(8):893–906, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Xue:1999:AHC

[2410] Xiaonan Xue and Ye Ding. Assessing heterogeneity and correlation of paired failure times with the bivariate frailty model. Statistics in Medicine, 18(8):907–918, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mezzetti:1999:HBA

[2411] Maura Mezzetti and Chris Robertson. A hierarchical Bayesian approach to age-specific back-calculation of cancer incidence rates. Statistics in Medicine, 18(8):919–933, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zaslavsky:1999:SAO

[2412] Alan M. Zaslavsky, Thomas V. Williams, and Paul D. Cleary. Sample allocation for overlapping domains in a physician survey with a limited population. Statistics in Medicine, 18(8):935–946, April 30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

TenHave:1999:CME

[2413] Thomas R. Ten Have, Allen R. Kunselman, and Luan Tran. A comparison of mixed effects logistic regression models for binary response data with two nested levels of clustering. Statistics in Medicine, 18(8): 947–960, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shih:1999:PAR

[2414] Weichung J. Shih and Hui Quan. Planning and analysis of repeated measures at key time-points in clinical trials sponsored by pharmaceutical companies. Statistics in Medicine, 18(8):961–973, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Miller:1999:VCD

[2415] William E. Miller. A visualization of cross-over data using linear functions. Statistics in Medicine, 18(8):975–987, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jung:1999:MNP

[2416] Jin-Whan Jung and Gary G. Koch. Multivariate non-parametric methods for Mann-Whitney statistics to analyse cross-over studies with two treatment sequences. Statistics in Medicine, 18(8):989–1017, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:1999:BRN

[2417] Nick Longford. Book review: The new statistical analysis of data. T. W. Anderson and J. D. Finn, Springer-Verlag, New York, 1996. No. of pages: xxi + 712. Price: £38. ISBN 0-387-94619-5. Statistics in Medicine, 18(8): 1019–1021, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:1999:BRS

[2418] Nick Longford. Book review: The SPSS guide to the new statistical analysis of data. S. B. Gerber and K. E. Voelkl, Springer-Verlag, New York, 1997. No. of pages: vii + 203. Price: US$32.95. ISBN 0-387-94821-X. Statistics in Medicine, 18(8):1021–1022, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuchenhoff:1999:BRL

[2419] H. Küchenhoff. Book review: Lineare Regression und Verwandtes. Hans Riedwyl, Birkhäuser Verlag, Switzerland, 1997. No. of pages: 133. Price: SFR 32. ISBN 3-7643-5495-X. Statistics in Medicine, 18(8):1022, April 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wohlfahrt:1999:MCR

[2420] Jan Wohlfahrt, Per Kragh Andersen, and Mads Melbye. Multivariate competing risks. Statistics in Medicine, 18(9):1023–1030, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Deuffic:1999:AIT

[2421] Sylvie Deuffic and Dominique Costagliola. Is the AIDS incubation time changing? A back-calculation approach. Statistics in Medicine, 18(9): 1031–1047, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gupta:1999:GTP

[2422] Diwakar Gupta and Regina Malina. Group testing in presence of classification errors. Statistics in Medicine, 18(9):1049–1068, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lyles:1999:AME

[2423] Robert H. Lyles, Alvaro Munõz, Jane Xu, Jeremy M. G. Taylor, and Joan S. Chmiel. Adjusting for measurement error to assess health effects of variability in biomarkers. Statistics in Medicine, 18(9):1069–1086, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Begg:1999:IEC

[2424] Melissa D. Begg and Katherine S. Panageas. Interval estimation of the common odds ratio from k (2×2) tables under cluster sampling. Statistics in Medicine, 18(9):1087–1100, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hin:1999:DCM

[2425] L. Y. Hin, T. K. Lau, M. S. Rogers, and A. M. Z. Chang. Dichotomization of continuous measurements using generalized additive modelling — application in predicting intrapartum caesarean delivery. Statistics in Medicine, 18(9):1101–1110, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oldehinkel:1999:TRM

[2426] Albertine J. Oldehinkel. A transition rate model for first admissions to psychiatric institutions. Statistics in Medicine, 18(9):1111–1118, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosen:1999:MPH

[2427] Ori Rosen and Martin Tanner. Mixtures of proportional hazards regression models. Statistics in Medicine, 18(9):1119–1131, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jiang:1999:CTV

[2428] Guoyong Jiang and Sanat K. Sarkar. Comparing treatment variances in repeated measures bioavailability trials. Statistics in Medicine, 18(9): 1133–1149, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Singer:1999:LEM

[2429] Júlia Singer. Letter to the Editor: A method for determining the size of internal pilot studies, by L. Sandvik, J. Erikssen, P. Mowinckel and E. A. Rødland, Statistics in Medicine, 15, 1587–1590 (1996). Statistics in Medicine, 18(9):1151–1153, May 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1624].

Thall:1999:ASP

[2430] Peter F. Thall, J. Jack Lee, Chi-Hong Tseng, and Elihu H. Estey. Accrual strategies for phase I trials with delayed patient outcome. Statistics in Medicine, 18(10):1155–1169, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zhou:1999:CTP

[2431] Xiao-Hua Zhou, Pete Castelluccio, Siu L. Hui, and Cynthia A. Rodenberg. Comparing two prevalence rates in a two-phase design study. Statistics in Medicine, 18(10):1171–1182, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sposto:1999:SVR

[2432] Richard Sposto and Daniel O. Stram. A strategic view of randomized trial design in low-incidence paediatric cancer. Statistics in Medicine, 18(10):1183–1197, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Coffey:1999:ETS

[2433] Christopher S. Coffey and Keith E. Muller. Exact test size and power of a Gaussian error linear model for an internal pilot study. Statistics in Medicine, 18(10):1199–1214, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Touloumi:1999:ECR

[2434] G. Touloumi, S. J. Pocock, A. G. Babiker, and J. H. Darbyshire. Estimation and comparison of rates of change in longitudinal studies with informative drop-outs. Statistics in Medicine, 18(10):1215–1233, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farrington:1999:ICS

[2435] C. P. Farrington and N. J. Gay. Interval-censored survival data with informative examination times: parametric models and approximate inference. Statistics in Medicine, 18(10):1235–1248, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tomasko:1999:DAE

[2436] Lisa Tomasko, Ronald W. Helms, and Steven M. Snapinn. A discriminant analysis extension to mixed models. Statistics in Medicine, 18(10):1249– 1260, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albrektsen:1999:JEC

[2437] Grethe Albrektsen, Ivar Heuch, and Gunnar Kvåle. Joint effects on cancer risk of age at childbirth, time since birth and attained age: circumventing the problem of collinearity. Statistics in Medicine, 18(10): 1261–1277, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finch:1999:PCP

[2438] S. J. Finch, J. B. Chen, C. H. Chen, M. M. Hall, V. Matkovich, B. Wenz, H. Brandwein, and J. Whitbread. Process control procedures to augment quality control of leukocyte-reduced red cell blood products. Statistics in Medicine, 18(10):1279–1289, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olsen:1999:BRS

[2439] Sjúrdur F. Olsen. Book review: Statistics further from scratch: for health care professionals. David Bowers, Wiley, 1997. No. of Pages: vi + 173. Price: £15.99. ISBN 0-471-97136-7. Statistics in Medicine, 18(10):1291– 1292, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Newcombe:1999:C

[2440] Robert G. Newcombe. Correction. Statistics in Medicine, 18(10):1293, May 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1999:E

[2441] Ted Colton and Laurence Freedman. Editorial. Statistics in Medicine, 18(11):1295–1296, June 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Choi:1999:IAD

[2442] Sung C. Choi and Young Jack Lee. Interim analyses with delayed observations in clinical trials. Statistics in Medicine, 18(11):1297–1306, June 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heyd:1999:ADI

[2443] Julie M. Heyd and Bradley P. Carlin. Adaptive design improvements in the continual reassessment method for phase I studies. Statistics in Medicine, 18(11):1307–1321, June 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hanfelt:1999:OMS

[2444] John J. Hanfelt. Optimal multi-stage designs for a phase II trial that permits one dose escalation. Statistics in Medicine, 18(11):1323–1339, June 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Finkelstein:1999:CML

[2445] Dianne M. Finkelstein and David A. Schoenfeld. Combining mortality and longitudinal measures in clinical trials. Statistics in Medicine, 18 (11):1341–1354, June 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Craig:1999:BAM

[2446] Bruce A. Craig, Dennis G. Fryback, Ronald Klein, and Barbara E. K. Klein. A Bayesian approach to modelling the natural history of a chronic condition from observations with intervention. Statistics in Medicine, 18 (11):1355–1371, June 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rao:1999:SMA

[2447] J. N. K. Rao and A. J. Scott. A simple method for analysing overdispersion in clustered Poisson data. Statistics in Medicine, 18(11):1373–1385, June 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosner:1999:UMW

[2448] B. Rosner and D. Grove. Use of the Mann–Whitney U-test for clustered data. Statistics in Medicine, 18(11):1387–1400, June 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Narula:1999:MSA

[2449] Subhash C. Narula, Paulo H. N. Saldiva, Carmen D. S. Andre, Silvia N. Elian, Aurea Favero Ferreira, and Vera Capelozzi. The minimum sum of absolute errors regression: a robust alternative to the least squares regression. Statistics in Medicine, 18(11):1401–1417, June 15, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stiger:1999:TPP

[2450] Thomas R. Stiger, Huiman X. Barnhart, and John M. Williamson. Testing proportionality in the proportional odds model fitted with GEE. Statistics in Medicine, 18(11):1419–1433, June 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Farewell:1999:CIP

[2451] V. T. Farewell and D. A. Sprott. Conditional inference for predictive agreement. Statistics in Medicine, 18(12):1435–1449, June 30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Perisic:1999:CMI

[2452] Igor Perisic and Bernard Rosner. Comparisons of measures of interclass correlations: the general case of unequal group size. Statistics in Medicine, 18(12):1451–1466, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:1999:SBT

[2453] Nicholas T. Longford. Selection bias and treatment heterogeneity in clinical trials. Statistics in Medicine, 18(12):1467–1474, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:1999:SSD

[2454] Zhiying Pan and Lawrence L. Kupper. Sample size determination for multiple comparison studies treating confidence interval width as random. Statistics in Medicine, 18(12):1475–1488, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1999:TCE

[2455] Per Kragh Andersen, John P. Klein, and Mei-Jie Zhang. Testing for centre effects in multi-centre survival studies: a Monte Carlo comparison of fixed and random effects tests. Statistics in Medicine, 18(12):1489– 1500, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wray:1999:CSS

[2456] Naomi R. Wray, Freda E. Alexander, Colin R. Muirhead, Eero Pukkala, Irene Schmidtmann, and Charles Stiller. A comparison of some simple methods to identify geographical areas with excess incidence of a rare disease such as childhood leukaemia. Statistics in Medicine, 18(12):1501– 1516, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Siegmund:1999:FAM

[2457] Kimberly D. Siegmund, Alexandre A. Todorov, and Michael A. Province. A frailty approach for modelling diseases with variable age of onset in families: the NHLBI family heart study. Statistics in Medicine, 18(12): 1517–1528, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Andersen:1999:MCA

[2458] Per Kragh Andersen, Mary M. Horowitz, John P. Klein, Gerard Socie, Judith Veum Stone, and Mei-Jie Zhang. Modelling covariate adjusted mortality relative to a standard population. Statistics in Medicine, 18 (12):1529–1540, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Shen:1999:EMQ

[2459] L. Z. Shen, E. Pulkstenis, and M. Hoseyni. Estimation of mean quality adjusted survival time. Statistics in Medicine, 18(12):1541–1554, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Weiss:1999:BCM

[2460] Robert E. Weiss, Meehyung Cho, and Michael Yanuzzi. On Bayesian calculations for mixture likelihoods and priors. Statistics in Medicine, 18(12):1555–1570, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Larsen:1999:BRB

[2461] Klaus Larsen. Book review: Basics of Structural Equation Modeling. Geoffrey M. Maruyama, Sage Publication, 1997. No. of pages: xvi + 298. Price: £47.99. ISBN 0-8039-7409-4. Statistics in Medicine, 18(12): 1571–1572, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Armitage:1999:BRH

[2462] Peter Armitage. Book review: A History of Mathematical Statistics from 1750 to 1930. Anders Hald, Wiley, New York, 1998. No. of pages: xvii + 795. Price: $125.00/£80.95. ISBN 0-471-17912-4. Statistics in Medicine, 18(12):1572–1573, June 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Denne:1999:ESS

[2463] Jonathan S. Denne and Christopher Jennison. Estimating the sample size for a t-test using an internal pilot. Statistics in Medicine, 18(13): 1575–1585, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Omar:1999:ARM

[2464] Rumana Z. Omar, Eileen M. Wright, Rebecca M. Turner, and Simon G. Thompson. Analysing repeated measurements data: a practical comparison of methods. Statistics in Medicine, 18(13):1587–1603, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Diggle:1999:CCI

[2465] Peter Diggle, Sara Morris, and Tony Morton-Jones. Case-control isotonic regression for investigation of elevation in risk around a point source. Statistics in Medicine, 18(13):1605–1613, July 15, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1999:TCM

[2466] Jeffrey M. Albert. A threshold causal model for clinical trials with departures from intended treatment. Statistics in Medicine, 18(13):1615–1626, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hwang:1999:MCE

[2467] Jing-Shiang Hwang and Jung-Der Wang. Monte Carlo estimation of extrapolation of quality-adjusted survival for follow-up studies. Statistics in Medicine, 18(13):1627–1640, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhalgh:1999:TDR

[2468] David Greenhalgh and Murray T. Doyle. A test to detect replication in HIV serological data labelled by birth date based on the number of matching pairs in a sample. Statistics in Medicine, 18(13):1641–1656, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Frost:1999:UHM

[2469] Chris Frost, Robert Clarke, and Heather Beacon. Use of hierarchical models for meta-analysis: experience in the metabolic ward studies of diet and blood cholesterol. Statistics in Medicine, 18(13):1657–1676, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Young:1999:ABM

[2470] Philip J. Young, Simon Weeden, and John R. Kirwan. The analysis of a bivariate multi-state Markov transition model for rheumatoid arthritis with an incomplete disease history. Statistics in Medicine, 18(13):1677– 1690, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Esbjerg:1999:RDC

[2471] Sille Esbjerg, Niels Keiding, and Nils Koch-Henriksen. Reporting delay and corrected incidence of multiple sclerosis. Statistics in Medicine, 18 (13):1691–1706, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Albert:1999:LDA

[2472] Paul S. Albert. Longitudinal data analysis (repeated measures) in clinical trials. Statistics in Medicine, 18(13):1707–1732, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1999:LEA

[2473] Niels Keiding. Letter to the Editor: Age–period-cohort models of chronic disease rates. II: graphical approaches, by C. Robertson and P. Boyle, Statistics in Medicine, 17, 1325–1340 (1998). Statistics in Medicine, 18 (13):1733, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2172].

Rohmel:1999:LEE

[2474] Joachim Röhmel and Ulrich Mansmann. Letter to the Editor: Exact tests of equivalence and efficacy with a non-zero lower bound for comparative studies, by I. S. F. Chan, Statistics in Medicine, 17, 1403–1413 (1998). Statistics in Medicine, 18(13):1734–1735, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2177].

Chan:1999:AR

[2475] Ivan S. F. Chan. Author’s reply. Statistics in Medicine, 18(13):1735– 1737, July 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Scherrer:1999:P

[2476] Bruno Scherrer and Stephen Senn. Preface. Statistics in Medicine, 18 (14):1739, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Atkinson:1999:OBC

[2477] Anthony C. Atkinson. Optimum biased-coin designs for sequential treatment allocation with covariate information. Statistics in Medicine, 18 (14):1741–1752, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2643].

Anonymous:1999:Da

[2478] Anonymous. Discussion. Statistics in Medicine, 18(14):1753–1755, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rosenberger:1999:BMA

[2479] William F. Rosenberger and Feifang Hu. Bootstrap methods for adaptive designs. Statistics in Medicine, 18(14):1757–1767, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Muller:1999:OTT

[2480] Hans-Helge Müller and Helmut Schäfer. Optimization of testing times and critical values in sequential equivalence testing. Statistics in Medicine, 18(14):1769–1788, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1999:Db

[2481] Anonymous. Discussion. Statistics in Medicine, 18(14):1789, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Broet:1999:UCT

[2482] Philippe Broët, Thierry Moreau, Joseph Lellouch, and Bernard Asselain. Unobserved covariates in the two-sample comparison of survival times: a maximin efficiency robust test. Statistics in Medicine, 18(14):1791–1800, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1999:Dc

[2483] Anonymous. Discussion. Statistics in Medicine, 18(14):1801, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

LeCorfec:1999:VDE

[2484] Emmanuelle Le Corfec, Sylvie Chevret, and Dominique Costagliola. Visit-driven endpoints in randomized HIV/AIDS clinical trials: impact of missing data on treatment difference measured on summary statistics. Statistics in Medicine, 18(14):1803–1817, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1999:Dd

[2485] Anonymous. Discussion. Statistics in Medicine, 18(14):1819, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mahe:1999:ETE

[2486] Cédric Mahé and Sylvie Chevret. Estimation of the treatment effect in a clinical trial when recurrent events define the endpoint. Statistics in Medicine, 18(14):1821–1829, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1999:De

[2487] Anonymous. Discussion. Statistics in Medicine, 18(14):1831, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bauer:1999:CDP

[2488] Peter Bauer and Meinhard Kieser. Combining different phases in the development of medical treatments within a single trial. Statistics in Medicine, 18(14):1833–1848, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kramar:1999:CRM

[2489] Andrew Kramar, Anne Lebecq, and Emmanuel Candalh. Continual reassessment methods in phase I trials of the combination of two drugs in oncology. Statistics in Medicine, 18(14):1849–1864, July 30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunne:1999:NAM

[2490] Adrian Dunne, Thomas O’Hara, and John Devane. A new approach to modelling the relationship between in vitro and in vivo drug dissolution/absorption. Statistics in Medicine, 18(14):1865–1876, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1999:Df

[2491] Anonymous. Discussion. Statistics in Medicine, 18(14):1877, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Penny:1999:MOD

[2492] Kay I. Penny and Ian T. Jolliffe. Multivariate outlier detection applied to multiply imputed laboratory data. Statistics in Medicine, 18(14): 1879–1895, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anonymous:1999:Dg

[2493] Anonymous. Discussion. Statistics in Medicine, 18(14):1897, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oja:1999:BRR

[2494] Hannu Oja. Book review: Robust nonparametric statistical methods. Thomas P. Hettmansperger and Joseph McKean, Arnold/Wiley, London/New York, 1998. No. of pages: xi + 467. Price £45. ISBN 0-340-54937-8 (Arnold) and 0-471-19479-4 (Wiley). Statistics in Medicine, 18 (14):1899–1900, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hsiung:1999:BRA

[2495] Chao Agnes Hsiung. Book review: Analysis of censored data. H. L. Koul and J. V. Deshpande (eds), Institute of Mathematical Statistics, 1995. No. of pages: vi + 291. Price: $18. ISBN 0-940600-39-0. Statistics in Medicine, 18(14):1900–1901, July 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lewis:1999:SPC

[2496] John A. Lewis. Statistical principles for clinical trials (ICH E9): an introductory note on an international guideline. Statistics in Medicine, 18(15):1903–1942, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rottman:1999:UMD

[2497] Jeffrey N. Rottman, Bruce Levin, Myunghee Cho Paik, Wei Yann Tsai, and J. Thomas Bigger, Jr. Using missing data techniques to explore the lack of survival effect: illustration with the CABG patch trial. Statistics in Medicine, 18(15):1943–1959, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Yamaguchi:1999:ICE

[2498] Takuhiro Yamaguchi and Yasuo Ohashi. Investigating centre effects in a multi-centre clinical trial of superficial bladder cancer. Statistics in Medicine, 18(15):1961–1971, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1999:SMC

[2499] Peter N. Lee. Simple methods for checking for possible errors in reported odds ratios, relative risks and confidence intervals. Statistics in Medicine, 18(15):1973–1981, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tan:1999:BHM

[2500] Ming Tan, Yinsheng Qu, Ed Mascha, and Armin Schubert. A Bayesian hierarchical model for multi-level repeated ordinal data: analysis of oral practice examinations in a large anaesthesiology training programme. Statistics in Medicine, 18(15):1983–1992, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

He:1999:MML

[2501] Xuming He and Wing K. Fung. Method of medians for lifetime data with Weibull models. Statistics in Medicine, 18(15):1993–2009, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ananth:1999:BLR

[2502] Cande V. Ananth and John S. Preisser. Bivariate logistic regression: modelling the association of small for gestational age births in twin gestations. Statistics in Medicine, 18(15):2011–2023, August 15, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Cardinal:1999:AIV

[2503] Mitsi Cardinal, Roch Roy, and Jean Lambert. On the application of integer-valued time series models for the analysis of disease incidence. Statistics in Medicine, 18(15):2025–2039, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lui:1999:NIE

[2504] Kung-Jong Lui and Colleen Kelly. A note on interval estimation of kappa in a series of 2 × 2 tables. Statistics in Medicine, 18(15):2041–2049, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zou:1999:HMA

[2505] Kelly H. Zou and Michael P. McDermott. Higher-moment approaches to approximate interval estimation for a certain intraclass correlation coefficient. Statistics in Medicine, 18(15):2051–2061, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anderson:1999:LEM

[2506] James R. Anderson and David Harrington. Letter to the Editor: The management of interim analyses in drug development, by Karen M. Facey and John A. Lewis, Statistics in Medicine, 17, 1801–1809 (1998). Statistics in Medicine, 18(15):2063–2065, August 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2218].

Tritchler:1999:RAD

[2507] David Tritchler. Reasoning about data with directed graphs. Statistics in Medicine, 18(16):2067–2076, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Saez:1999:GMA

[2508] Marc Saez, Aurelio Tobias, Pilar Muñoz, and M. J. Campbell. A GEE moving average analysis of the relationship between air pollution and mortality for asthma in Barcelona, Spain. Statistics in Medicine, 18 (16):2077–2086, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Osnes:1999:SSC

[2509] Kåre Osnes and Odd O. Aalen. Spatial smoothing of cancer survival: a Bayesian approach. Statistics in Medicine, 18(16):2087–2099, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Haber:1999:EDI

[2510] Michael Haber. Estimation of the direct and indirect effects of vaccination. Statistics in Medicine, 18(16):2101–2109, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rossi:1999:ACP

[2511] G. Rossi, L. Lampugnani, and M. Marchi. An approximate CUSUM procedure for surveillance of health events. Statistics in Medicine, 18 (16):2111–2122, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aydemir:1999:ATD

[2512] Ülker Aydemir, Sibel Aydemir, and Peter Dirschedl. Analysis of timedependent covariates in failure time data. Statistics in Medicine, 18 (16):2123–2134, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tritchler:1999:MSQ

[2513] David Tritchler. Modelling study quality in meta-analysis. Statistics in Medicine, 18(16):2135–2145, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Assuncao:1999:NPA

[2514] Renato M. Assunção and Edna A. Reis. A new proposal to adjust Moran’s I for population density. Statistics in Medicine, 18(16):2147– 2162, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brumback:1999:AEC

[2515] Babette A. Brumback, Lewis B. Holmes, and Louise M. Ryan. Adverse effects of chorionic villus sampling: a meta-analysis. Statistics in Medicine, 18(16):2163–2175, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Oppenheimer:1999:IME

[2516] Leonard Oppenheimer and Uma Kher. The impact of measurement error on the comparison of two treatments using a responder analysis. Statistics in Medicine, 18(16):2177–2188, August 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jones:1999:P

[2517] Byron Jones. Preface. Statistics in Medicine, 18(17-18):2189, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Agresti:1999:MOC

[2518] Alan Agresti. Modelling ordered categorical data: recent advances and future challenges. Statistics in Medicine, 18(17-18):2191–2207, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Collett:1999:SPI

[2519] D. Collett and K. Stepniewska. Some practical issues in binary data analysis. Statistics in Medicine, 18(17-18):2209–2221, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lindsey:1999:RSE

[2520] J. K. Lindsey. A review of some extensions to generalized linear models. Statistics in Medicine, 18(17-18):2223–2236, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molenberghs:1999:MMM

[2521] Geert Molenberghs and Emmanuel Lesaffre. Marginal modelling of multivariate categorical data. Statistics in Medicine, 18(17-18):2237–2255, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jorgensen:1999:DML

[2522] Bent Jørgensen and Min Tsao. Dispersion models and longitudinal data analysis. Statistics in Medicine, 18(17-18):2257–2270, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1999:UTS

[2523] John Whitehead. A unified theory for sequential clinical trials. Statistics in Medicine, 18(17-18):2271–2286, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hu:1999:GSA

[2524] X. Joan Hu and Stephen W. Lagakos. Group sequential analyses for the mean function of a repeated measure process. Statistics in Medicine, 18 (17-18):2287–2299, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Anello:1999:ERI

[2525] Charles Anello. Emerging and recurrent issues in drug development. Statistics in Medicine, 18(17-18):2301–2309, September 15–30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Longford:1999:SVS

[2526] Nicholas T. Longford and John A. Nelder. Statistics versus statistical science in the regulatory process. Statistics in Medicine, 18(17-18): 2311–2320, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rohmel:1999:CAS

[2527] Joachim Röhmel. Controversies about sponsor initiated re-analyses of clinical trial data in the licensing process. Statistics in Medicine, 18 (17-18):2321–2330, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olkin:1999:DSP

[2528] Ingram Olkin. Diagnostic statistical procedures in medical metaanalyses. Statistics in Medicine, 18(17-18):2331–2341, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Aitkin:1999:MAR

[2529] Murray Aitkin. Meta-analysis by random effect modelling in generalized linear models. Statistics in Medicine, 18(17-18):2343–2351, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Keiding:1999:EHA

[2530] Niels Keiding. Event history analysis and inference from observational epidemiology. Statistics in Medicine, 18(17-18):2353–2363, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lawless:1999:SDM

[2531] J. F. Lawless and D. Y. T. Fong. State duration models in clinical and observational studies. Statistics in Medicine, 18(17-18):2365–2376, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wakefield:1999:ISA

[2532] Jon Wakefield and Paul Elliott. Issues in the statistical analysis of small area health data. Statistics in Medicine, 18(17-18):2377–2399, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lange:1999:STF

[2533] Nicholas Lange. Statistical thinking in functional and structural magnetic resonance neuroimaging. Statistics in Medicine, 18(17-18):2401– 2407, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Matsui:1999:ARE

[2534] Shigeyuki Matsui and Yasuo Ohashi. Analysis of recurrent events: application to a clinical trial of colony stimulating factor with the endpoint of febrile neutropenia. Statistics in Medicine, 18(17-18):2409–2420, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Heitjan:1999:IBC

[2535] Daniel F. Heitjan. Ignorability and bias in clinical trials. Statistics in Medicine, 18(17-18):2421–2434, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lipsitz:1999:NIM

[2536] Stuart R. Lipsitz, Joseph G. Ibrahim, Ming-Hui Chen, and Harriet Peterson. Non-ignorable missing covariates in generalized linear models. Statistics in Medicine, 18(17-18):2435–2448, September 15–30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Molenberghs:1999:MDP

[2537] Geert Molenberghs, Els J. T. Goetghebeur, Stuart R. Lipsitz, Michael G. Kenward, Emmanuel Lesaffre, and Bart Michiels. Missing data perspectives of the fluvoxamine data set: a review. Statistics in Medicine, 18 (17-18):2449–2464, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Little:1999:SMA

[2538] Roderick J. Little and Trivellore Raghunathan. On summary measures analysis of the linear mixed effects model for repeated measures when data are not missing completely at random. Statistics in Medicine, 18 (17-18):2465–2478, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sammel:1999:MLM

[2539] Mary Sammel, Xihong Lin, and Louise Ryan. Multivariate linear mixed models for multiple outcomes. Statistics in Medicine, 18(17-18):2479– 2492, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Louis:1999:IBE

[2540] Thomas A. Louis and Wei Shen. Innovations in Bayes and empirical Bayes methods: estimating parameters, populations and ranks. Statistics in Medicine, 18(17-18):2493–2505, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tierney:1999:SAM

[2541] Luke Tierney and Antonietta Mira. Some adaptive Monte Carlo methods for Bayesian inference. Statistics in Medicine, 18(17-18):2507–2515, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Prentice:1999:NPM

[2542] Ross L. Prentice. On non-parametric maximum likelihood estimation of the bivariate survivor function. Statistics in Medicine, 18(17-18): 2517–2527, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Graf:1999:ACP

[2543] Erika Graf, Claudia Schmoor, Willi Sauerbrei, and Martin Schumacher. Assessment and comparison of prognostic classification schemes for survival data. Statistics in Medicine, 18(17-18):2529–2545, September 15– 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1999:BRI

[2544] Jae Chang Lee. Book review: Introduction to the Statistical Analysis of Categorical Data. E. B. Andersen, Springer for Science, 1997. No. of pages: 265. Price: $44.95. ISBN 3-5406-2399-X. Statistics in Medicine, 18(17-18):2547, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ghitany:1999:BRP

[2545] Mohamed E. Ghitany. Book review: Primer of Biostatistics Fourth edition, Stanton A. Glantz, McGraw-Hill Inc., New York, 1997. No. of pages: xvi + 473 + computer program. Price: £51.95. ISBN 0-07-024268-2. Statistics in Medicine, 18(17-18):2548–2549, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1999:BRS

[2546] Moo-Song Lee. Book review: Statistics, the Conceptual Approach. Gudmund R. Iversen and Mary Gergen, Springer, New York, 1997. No. of pages: xviii + 735. Price: $54.95. ISBN 0-387-94610-1. Statistics in Medicine, 18(17-18):2549–2550, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Sabroe:1999:BRA

[2547] Svend Sabroe. Book review: Adventures in Social Research. Data Analysis Using SPSS for Windows. E. Babbie and F. Halley, Pine Forge Press, 1995. pages: 346 + diskette. Price: £25. ISBN 0-8039-9055-3. Statistics in Medicine, 18(17-18):2550, September 15–30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Snr:1999:EQC

[2548] Ralph B. D’Agostino Snr. Editorial: quantifying the comparison of two groups. Statistics in Medicine, 18(19):2551–2555, October 15, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feinstein:1999:ICQ

[2549] Alvan R. Feinstein. Indexes of contrast and quantitative significance for comparisons of two groups. Statistics in Medicine, 18(19):2557–2581, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Freedman:1999:WQS

[2550] Laurence S. Freedman. What is quantitative significance and is it useful? Statistics in Medicine, 18(19):2583–2584, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Koch:1999:DIC

[2551] Gary G. Koch. Discussion for ‘Indexes of contrast and quantitative significance for comparisons of two groups’. Statistics in Medicine, 18 (19):2585–2589, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Feinstein:1999:RPF

[2552] Alvan R. Feinstein. Response to professors Freedman and Koch. Statistics in Medicine, 18(19):2591–2592, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brant:1999:EMI

[2553] Rollin Brant, Lloyd Sutherland, and Robert Hilsden. Examining the minimum important difference. Statistics in Medicine, 18(19):2593–2603, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Roberts:1999:IVO

[2554] Chris Roberts. The implications of variation in outcome between health professionals for the design and analysis of randomized controlled trials. Statistics in Medicine, 18(19):2605–2615, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

White:1999:RBM

[2555] Ian R. White, Abdel G. Babiker, Sarah Walker, and Janet H. Darbyshire. Randomization-based methods for correcting for treatment changes: examples from the Concorde trial. Statistics in Medicine, 18(19):2617– 2634, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

LeBlanc:1999:UBE

[2556] Michael LeBlanc and John Crowley. Using the bootstrap for estimation in group sequential designs: an application to a clinical trial for nasopharyngeal cancer. Statistics in Medicine, 18(19):2635–2644, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Lee:1999:OWS

[2557] Wen-Chung Lee and Yung-Po Liaw. Optimal weighting systems for direct age-adjustment of vital rates. Statistics in Medicine, 18(19):2645–2654, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Carlin:1999:ABO

[2558] John B. Carlin, Rory Wolfe, Carolyn Coffey, and George C. Patton. Analysis of binary outcomes in longitudinal studies using weighted estimating equations and discrete-time survival methods: prevalence and incidence of smoking in an adolescent cohort. Statistics in Medicine, 18 (19):2655–2679, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Fay:1999:CSM

[2559] Michael P. Fay. Correction to Statistics in Medicine, 18, 273–285 (1999). Statistics in Medicine, 18(19):2681, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2350].

Platt:1999:CSM

[2560] Robert Platt. Correction to Statistics in Medicine, 18, 643–654 (1999). Statistics in Medicine, 18(19):2681–2682, October 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

OQuigley:1999:ALT

[2561] John O’Quigley. Another look at two phase I clinical trial designs. Statistics in Medicine, 18(20):2683–2690, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korn:1999:C

[2562] Edward L. Korn, Douglas Midthune, T. Timothy Chen, Lawrence V. Rubinstein, Michaele C. Christian, and Richard M. Simon. Commentary.

Statistics in Medicine, 18(20):2691–2692, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Thompson:1999:EHM

[2563] Simon G. Thompson and Stephen J. Sharp. Explaining heterogeneity in meta-analysis: a comparison of methods. Statistics in Medicine, 18 (20):2693–2708, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mander:1999:AMD

[2564] A. P. Mander, M. D. Hughes, S. J. Sharp, and C. J. Lamm. Autoregressive models for describing non-linear changes in biological parameters fitted using BUGS. Statistics in Medicine, 18(20):2709–2722, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Marshall:1999:COC

[2565] Roger J. Marshall. Classification to ordinal categories using a search partition methodology with an application in diabetes screening. Statistics in Medicine, 18(20):2723–2735, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Goggins:1999:ACP

[2566] William B. Goggins, Dianne M. Finkelstein, and Alan M. Zaslavsky. Applying the Cox proportional hazards model for analysis of latency data with interval censoring. Statistics in Medicine, 18(20):2737–2747, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:1999:WTC

[2567] Wanzhu Tu and Xiao-Hua Zhou. A Wald test comparing medical costs based on log-normal distributions with zero valued costs. Statistics in Medicine, 18(20):2749–2761, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Troendle:1999:APW

[2568] James F. Troendle. Approximating the power of Wilcoxon’s rank-sum test against shift alternatives. Statistics in Medicine, 18(20):2763–2773, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

McIntosh:1999:IVW

[2569] Martin W. McIntosh. Instrumental variables when evaluating screening trials: estimating the benefit of detecting cancer by screening. Statistics in Medicine, 18(20):2775–2794, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kosinski:1999:EED

[2570] Andrzej S. Kosinski and W. Dana Flanders. Evaluating the exposure and disease relationship with adjustment for different types of exposure misclassification: a regression approach. Statistics in Medicine, 18 (20):2795–2808, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Palmer:1999:BRH

[2571] Chris Palmer. Book reviews: How to Report Statistics in Medicine: Annotated Guidelines for Authors, Editors and Reviewers. Thomas A. Lang and Michelle Secic, American College of Physicians, Philadelphia PA, 1997. No. of pages: 367. Price: $39.95. ISBN 0-943126-44-4. Publishing Your Medical Research Paper: What They Don’t Teach in Medical School. Daniel W. Byrne, Williams and Wilkins, Philadelphia, PA, 1998. No. of pages: 298. Price: $23.95. ISBN 0-683-30074-1. Statistics in Medicine, 18 (20):2809–2811, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Laud:1999:BRB

[2572] Purushottam W. (Prakash) Laud. Book review: Bayesian Statistics: an Introduction, second edition. Peter M. Lee, Arnold, 1997. No. of pages: xii + 344. Price: £19.99. ISBN 0-340-67785-6. Statistics in Medicine, 18 (20):2812–2813, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Ibrahim:1999:BRP

[2573] Joseph G. Ibrahim. Book review: The Practice of Bayesian Analysis. S. French and J. Q. Smith (eds), Wiley, New York, 1997. No. of pages: ix + 284. Price: £29.99. ISBN 0-340-66240-9. Statistics in Medicine, 18 (20):2813–2814, October 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:1999:MEEa

[2574] M. Y. Wong, N. E. Day, S. A. Bashir, and S. W. Duffy. Measurement error in epidemiology: the design of validation studies i: univariate situation. Statistics in Medicine, 18(21):2815–2829, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wong:1999:MEEb

[2575] M. Y. Wong, N. E. Day, and N. J. Wareham. Measurement error in epidemiology: the design of validation studies II: bivariate situation. Statistics in Medicine, 18(21):2831–2845, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Holcomb:1999:RCO

[2576] John P. Holcomb, Jr. Regression with covariates and outcome calculated from a common set of variables measured with error: estimation using the SIMEX method. Statistics in Medicine, 18(21):2847–2862, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Dunn:1999:PME

[2577] Graham Dunn. The problem of measurement error in modelling the effect of compliance in a randomized trial. Statistics in Medicine, 18(21):2863– 2877, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Korhonen:1999:CNC

[2578] Pasi A. Korhonen, Nan M. Laird, and Juni Palmgren. Correcting for non-compliance in randomized trials: an application to the ATBC study. Statistics in Medicine, 18(21):2879–2897, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bandeen-Roche:1999:MDP

[2579] Karen Bandeen-Roche, Charles B. Hall, Walter F. Stewart, and Scott L. Zeger. Modelling disease progression in terms of exposure history. Statistics in Medicine, 18(21):2899–2916, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Douglas:1999:IRM

[2580] Jeffrey A. Douglas. Item response models for longitudinal quality of life data in clinical trials. Statistics in Medicine, 18(21):2917–2931, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Park:1999:SPM

[2581] Taesung Park and Seung-Yeoun Lee. Simple pattern-mixture models for longitudinal data with missing observations: analysis of urinary incontinence data. Statistics in Medicine, 18(21):2933–2941, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stanek:1999:WRU

[2582] Edward J. Stanek III, Arnold Well, and Ira Ockene. Why not routinely use best linear unbiased predictors (BLUPs) as estimates of cholesterol, per cent fat from kcal and physical activity? Statistics in Medicine, 18 (21):2943–2959, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Olson:1999:GMC

[2583] Jane M. Olson, John S. Witte, and Robert C. Elston. Genetic mapping of complex traits. Statistics in Medicine, 18(21):2961–2981, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Almond:1999:BRG

[2584] Russell Almond. Book review: Graphical Models. Steffen L. Lauritzen, Oxford University Press, 1996. No. of pages: 298. ISBN 0-19-852219-3. Statistics in Medicine, 18(21):2983–2984, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Petersen:1999:BRS

[2585] Jørgen Holm Petersen. Book review: SPSS Analysis Without Anguish. S. J. Coakes and L. G. Steed, Wiley, Singapore, 1997. No. of pages: x + 276 + 2 diskettes. Price: £16.99. ISBN 0-471-33869-9. Statistics in Medicine, 18(21):2984–2985, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Jorgensen:1999:BRG

[2586] Merete Jørgensen. Book review: Good Clinical Practice. Standard Operating Procedures for Clinical Researchers. Josef Kolman, Paul Meng and Graeme Scott (eds), Wiley, Chichester, 1998, No of Pages: xi + 177. Price: £29.25. ISBN 0-471-96936-2. Statistics in Medicine, 18(21):2985– 2986, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nagelkerke:1999:BRP

[2587] Nico Nagelkerke. Book review: Principles of Experimental Design for the Life Sciences. Murray R. Selwyn, CRC Press, Boca Raton, 1996. No. of pages: 160. Price: $ 49.95. ISBN 0-8493-9461-9. Statistics in Medicine, 18(21):2986, November 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Alonzo:1999:UCR

[2588] Todd A. Alonzo and Margaret S. Pepe. Using a combination of reference tests to assess the accuracy of a new diagnostic test. Statistics in Medicine, 18(22):2987–3003, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See letter [2642].

Matsui:1999:DCT

[2589] Shigeyuki Matsui and Hideaki Miyagishi. Design of clinical trials for recurrent events with periodic monitoring. Statistics in Medicine, 18 (22):3005–3020, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Bilker:1999:EDV

[2590] Warren B. Bilker, Jesse A. Berlin, Mitchell H. Gail, and Brian L. Strom. An efficient design for verifying disease outcome status in large cohorts with rare exposures and low disease rates. Statistics in Medicine, 18(22): 3021–3036, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Putt:1999:MEM

[2591] Mary Putt and Vernon M. Chinchilli. A mixed effects model for the analysis of repeated measures cross-over studies. Statistics in Medicine, 18(22):3037–3058, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tu:1999:BAP

[2592] Xin M. Tu, Jeanne Kowalski, and Gang Jia. Bayesian analysis of prevalence with covariates using simulation-based techniques: applications to HIV screening. Statistics in Medicine, 18(22):3059–3073, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Hess:1999:HFE

[2593] Kenneth R. Hess, Dan M. Serachitopol, and Barry W. Brown. Hazard function estimators: a simulation study. Statistics in Medicine, 18(22): 3075–3088, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Betensky:1999:NPM

[2594] Rebecca A. Betensky and Dianne M. Finkelstein. A non-parametric maximum likelihood estimator for bivariate interval censored data. Statistics in Medicine, 18(22):3089–3100, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Betensky:1999:EKC

[2595] Rebecca A. Betensky and Dianne M. Finkelstein. An extension of Kendall’s coefficient of concordance to bivariate interval censored data. Statistics in Medicine, 18(22):3101–3109, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pan:1999:LRB

[2596] Wei Pan and Charles Kooperberg. Linear regression for bivariate censored data via multiple imputation. Statistics in Medicine, 18(22):3111– 3121, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Belin:1999:PGL

[2597] Thomas R. Belin, Ming-Yi Hu, Alexander S. Young, and Oscar Grusky. Performance of a general location model with an ignorable missing-data assumption in a multivariate mental health services study. Statistics in Medicine, 18(22):3123–3135, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Brown:1999:APC

[2598] Barry W. Brown, James Lovato, and Kathy Russell. Asymptotic power calculations: description, examples, computer code. Statistics in Medicine, 18(22):3137–3151, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kuchler:1999:BRD

[2599] Ingeborg Küchler. Book review: Data Driven Statistical Methods. Peter Sprent, Chapman and Hall, London 1998. No. of pages: 406. Price. £35. ISBN 0-412-79540-X. Statistics in Medicine, 18(22):3153–3154, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Rychtarikova:1999:BRD

[2600] Jitka Rychtaríková. Book review: Demographic Methods. Andrew Hinde, Arnold, London, 1998. No. of pages: xiii + 305. Price: £21.99. ISBN 0-340-71892-7. Statistics in Medicine, 18(22):3154–3155, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kendall:1999:BRL

[2601] Wilfrid S. Kendall. Book review: Local Stereology. Eva B. Vedel Jensen, World Scientific, Singapore, 1998. No. of pages: xv + 247. Price: £37. ISBN 981-02-2454-0. Statistics in Medicine, 18(22):3155–3156, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DeGunst:1999:BRS

[2602] Mathisca De Gunst. Book review: Stochastic Models of Tumor Latency and their Biostatistical Applications. A. Y. Yakovlev and A. D. Tsodikov, World Scientific, Singapore, 1996. No. of pages. xvi + 269. Price: £44. ISBN 981-02-1831-1. Statistics in Medicine, 18(22):3156–3157, November 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Falter:1999:P

[2603] K. H. Falter. Preface. Statistics in Medicine, 18(23):3159–3160, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Broome:1999:IW

[2604] Claire V. Broome. Introduction and welcome. Statistics in Medicine, 18(23):3165, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Kafadar:1999:SSA

[2605] Karen Kafadar. Simultaneous smoothing and adjusting mortality rates in U.S. counties: melanoma in white females and white males. Statistics in Medicine, 18(23):3167–3188, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Malec:1999:MBS

[2606] Donald Malec, William W. Davis, and Xin Cao. Model-based small area estimates of overweight prevalence using sample selection adjustment. Statistics in Medicine, 18(23):3189–3200, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Mungiole:1999:AWH

[2607] Michael Mungiole, Linda W. Pickle, and Katherine Hansen Simonson. Application of a weighted head-banging algorithm to mortality data maps. Statistics in Medicine, 18(23):3201–3209, December 15, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Pickle:1999:ESP

[2608] Linda W. Pickle, Michael Mungiole, Gretchen K. Jones, and Andrew A. White. Exploring spatial patterns of mortality: the new Atlas of United States Mortality. Statistics in Medicine, 18(23):3211–3220, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gelman:1999:AMP

[2609] Andrew Gelman and Phillip N. Price. All maps of parameter estimates are misleading. Statistics in Medicine, 18(23):3221–3234, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Russell:1999:MCE

[2610] Louise B. Russell. Modelling for cost-effectiveness analysis. Statistics in Medicine, 18(23):3235–3244, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Briggs:1999:CCI

[2611] Andrew H. Briggs, Christopher Z. Mooney, and David E. Wonderling. Constructing confidence intervals for cost-effectiveness ratios: an evaluation of parametric and non-parametric techniques using Monte Carlo simulation. Statistics in Medicine, 18(23):3245–3262, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Edmunds:1999:ECE

[2612] W. J. Edmunds, G. F. Medley, and D. J. Nokes. Evaluating the cost-effectiveness of vaccination programmes: a dynamic perspective. Statistics in Medicine, 18(23):3263–3282, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Williamson:1999:MSD

[2613] G. David Williamson and Ginner Weatherby Hudson. A monitoring system for detecting aberrations in public health surveillance reports. Statistics in Medicine, 18(23):3283–3298, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Chen:1999:CIC

[2614] Rina Chen. The cumulative q interval curve as a starting point in disease cluster investigation. Statistics in Medicine, 18(23):3299–3307, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

VanBrackle:1999:SAR

[2615] Lewis VanBrackle and G. David Williamson. A study of the average run length characteristics of the national notifiable diseases surveillance system. Statistics in Medicine, 18(23):3309–3319, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gong:1999:EGI

[2616] Gail Gong and Alice S. Whittemore. Estimating genetic influence on disease from population-based case-control data: application to cancers of the breast and ovary. Statistics in Medicine, 18(23):3321–3336, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Li:1999:ALM

[2617] Jianli Li. An application of lifetime models in estimation of expected length of stay of patients in hospital with complexity and age adjustment. Statistics in Medicine, 18(23):3337–3344, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Nobre:1999:MCD

[2618] Flávio Fonseca Nobre, Lilian Terezinha Ferreira Trotta, and Luiz Flávio Autran Monteiro Gomes. Multi-criteria decision making — an approach to setting priorities in health care. Statistics in Medicine, 18 (23):3345–3354, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wassell:1999:RIE

[2619] James T. Wassell, William C. Wojciechowski, and Deborah D. Landen. Recurrent injury event-time analysis. Statistics in Medicine, 18(23): 3355–3363, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Stroup:1999:CR

[2620] Donna F. Stroup. Closing remarks. Statistics in Medicine, 18(23):3373– 3375, December 15, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

DAgostino:1999:P

[2621] Ralph D’Agostino, David Machin, and Chris Palmer. Preface. Statistics in Medicine, 18(24):3377, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Colton:1999:HNM

[2622] Theodore Colton and Janet Wittes. In honour of Nathan Mantel’s 80th birthday. Statistics in Medicine, 18(24):3379, December 30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:1999:MU

[2623] Janet Wittes. Mantel unhyphenated. Statistics in Medicine, 18(24): 3381–3388, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Gail:1999:SNM

[2624] Mitchell H. Gail. Some of Nathan Mantel’s contributions to epidemiology. Statistics in Medicine, 18(24):3389–3400, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Greenhouse:1999:SMC

[2625] Samuel W. Greenhouse. A selection of Mantel’s contributions to laboratory research. Statistics in Medicine, 18(24):3401–3408, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Walter:1999:MAS

[2626] S. D. Walter and A. R. Jadad. Meta-analysis of screening data: a survey of the literature. Statistics in Medicine, 18(24):3409–3424, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Whitehead:1999:BSD

[2627] John Whitehead. On being the statistician on a data and safety monitoring board. Statistics in Medicine, 18(24):3425–3434, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Buyse:1999:RBP

[2628] Marc Buyse, Stephen L. George, Stephen Evans, Nancy L. Geller, Jonas Ranstam, Bruno Scherrer, Emmanuel Lesaffre, Gordon Murray, Lutz Edler, Jane Hutton, Theodore Colton, Peter Lachenbruch, and Babu L. Verma. The role of biostatistics in the prevention, detection and treatment of fraud in clinical trials. Statistics in Medicine, 18(24):3435–3451, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Reitmeir:1999:RBM

[2629] Peter Reitmeir and Gernot Wassmer. Resampling-based methods for the analysis of multiple endpoints in clinical trials. Statistics in Medicine, 18 (24):3453–3462, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

LeStrat:1999:MES

[2630] Yann Le Strat and Fabrice Carrat. Monitoring epidemiologic surveillance data using hidden Markov models. Statistics in Medicine, 18(24):3463– 3478, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Wittes:1999:IPS

[2631] Janet Wittes, Oliver Schabenberger, David Zucker, Erica Brittain, and Michael Proschan. Internal pilot studies I: type I error rate of the naive t-test. Statistics in Medicine, 18(24):3481–3491, December 30, 1999. CO-DEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Zucker:1999:IPS

[2632] David M. Zucker, Janet T. Wittes, Oliver Schabenberger, and Erica Brittain. Internal pilot studies II: comparison of various procedures. Statistics in Medicine, 18(24):3493–3509, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:1999:LEI

[2633] T. Tango. Letter to the Editor: Improved confidence intervals for the difference between binomial proportions based on paired data, by Robert G. Newcombe, Statistics in Medicine, 17, 2635–2650 (1998). Statistics in Medicine, 18(24):3511–3513, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2295].

Newcombe:1999:AR

[2634] Robert G. Newcombe. Author’s reply. Statistics in Medicine, 18(24): 3513, December 30, 1999. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic).

Tango:2000:LEC

[2635] T. Tango. Letter to the Editor: Confidence intervals for differnces in correlated binary proportions by W. L. May and W. D. Johnson, Statistics in Medicine, 16, 2127–2136 (1997). Statistics in Medicine, 19(1): 133–136, January 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1959, 2332, 2237].

Buyse:2000:LEU

[2636] Marc Buyse. Letter to the Editor: R^{2}: a useful measure of model performance when predicting a dichotomous outcome, by A. Ash and M. Schwartz, Statistics in Medicine, 18, 375–384 (1999). Statistics in Medicine, 19(2):271–274, January 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2357].

Stijnen:2000:LET

[2637] Theo Stijnen. Letter to the Editor: Tutorial in Biostatistics. Metaanalysis: formulating, evaluating, combining, and reporting, by S-L. Normand, Statistics in Medicine, 18, 321–359 (1999). Statistics in Medicine, 19(5):759–761, March 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2353].

Bolfarine:2000:LEC

[2638] Heleno Bolfarine and Filidor Vilca-Labra. Letter to the Editor: Comparative calibration without a gold standard Y. Lu, K. Ye, A. Mathur, S. Hul, T. Fuerst and H. Genant, Statistics in Medicine 1997; 16:1889– 1905. Statistics in Medicine, 19(9):1253–1254, May 15, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1932].

Julious:2000:LER

[2639] Steven A. Julious. Letter to the Editor: Repeated measures in clinical trials: analysis using mean summary statistics and its implications for design by L. Frison and S. J. Pocock, Statistics in Medicine 1992; 12: 1685–1704. Statistics in Medicine, 19(22):3133–3135, November 30, 2000. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [603].

Tamhane:2001:LEM

[2640] A. C. Tamhane and L. A. Hothorn. Letter to the Editor: A multiple comparison procedure for three- and four-armed controlled clinical trials by M. A. Proschan, Statistics in Medicine 1999; 18:787–798. Statistics in Medicine, 20(2):317–318, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2400].

Bauer:2001:LEM

[2641] Peter Bauer. Letter to the Editor: A multiple comparison procedure for three- and four-armed controlled clinical trials by M. A. Proschan, Statistics in Medicine 1999; 18:787–798. Statistics in Medicine, 20(2): 319–320, January 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2400].

Hadgu:2001:LEU

[2642] Alula Hadgu and William Miller. Letter to the Editor: Using a combination of reference tests to assess the accuracy of a diagnostic test by A. Alonzo and M. Pepe, Statistics in Medicine 1999; 18: 2987–3003. Statistics in Medicine, 20(4):656–658, February 28, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2588].

Taves:2001:LEO

[2643] Donald R. Taves. Letter to the Editor: Optimum biased-coin designs for sequential treatment allocation with covariate information by A. C. Atkinson, Statistics in Medicine 1999; 18:1741–1752. Statistics in Medicine, 20(5):813–816, March 15, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2477].

Chen:2001:LEM

[2644] Tuhao Chen. Letter to the Editor: Methods of Adjustment for Estimating the Attributable Risk in Case-Control Studies: a Review, by Jacques Benichou, Statistics in Medicine 1991; 10:1753–1773. Statistics in Medicine, 20(6):979–981, March 30, 2001. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [398].

Schoder:2002:LET

[2645] Volker Schoder. Letter to the Editor: Test statistics and sample size formulae for comparative binomial trials with null hypothesis of non-zero risk difference or non-unity relative risk by C. P. Farrington and G. Manning, Statistics in Medicine 1990; 9:1447–1454. Statistics in Medicine, 21(13):1958–1960, July 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [174].

Senn:2002:LEC

[2646] Stephen Senn. Letter to the Editor: A comment on replication, p-values and evidence, S. N. Goodman, Statistics in Medicine 1992; 11:875–879. Statistics in Medicine, 21(16):2437–2444, August 30, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [516].

Lin:2002:LER

[2647] Chii-Dean Lin and Kung-Jong Lui. Letter to the Editor: Randomized consent designs for clinical trials: an update by M. Zelen, Statistics in Medicine 1990; 9: 645–656. Statistics in Medicine, 21(17):2601–2603, September 15, 2002. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [66].

Julious:2005:LET

[2648] Steven A. Julious. Letter to the Editor: Two-sided confidence intervals for the single proportion: comparison of seven methods by Robert G. Newcombe, Statistics in Medicine 1998; 17:857–872. Statistics in Medicine, 24(21):3383–3384, November 15, 2005. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [2122].

Anonymous:2007:LES

[2649] Anonymous. Letter to the Editor: A semi-parametric estimate of extra-Poisson variation for vital rates, by M. Fay and E. Feuer, Statistics in Medicine 1997; 16:2389–2401. Statistics in Medicine, 26(9):2131–2133, April 30, 2007. CODEN SMEDDA. ISSN 0277-6715 (print), 1097-0258 (electronic). See [1986].